A Bibliography of Papers in Lecture Notes in Computer Science (2001) (Part 4 of 4)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

14 October 2017 Version 1.15

Title word cross-reference

4 [1370]. (3, 1, 2) [1853]. (α, β) [2788]. (g, f) [661]. (k, [k]P) [2593]. (p^{6m}) [2604]. + [2726]. {0, 1} [2647]. 1 [1399, 2621, 269, 2751]. 10 [1853]. 2 [1067, 912, 658, 1040, 1399, 2621, 1572, 1043, 794, 1407, 1996, 755, 878, 1672, 1005, 370, 1009, 145]. 24 [1539]. 2^{nd} [334]. 2K^{Q+} [1288]. 2 × 2 [1396]. 3 [1854, 934, 110, 884, 925, 149, 1067, 901, 912, 1992, 1002, 1064, 1993, 932, 971, 1682, 986, 876, 134, 443, 1046, 1379, 111, 1043, 794, 826, 861, 972, 943, 1017, 985, 1681, 1979, 1027, 386, 941, 94, 937, 527, 824, 2427, 878, 910, 954, 147, 105, 1997, 1029, 953, 1999, 2000, 848, 1672, 1005, 243, 416, 990, 1009, 907, 871, 1856, 2426, 109]. 32 [1539]. 40 [1539]. _{2} [1682, 125]. _{k} [669]. _{λ} [2572]. a [2251]. α [2786, 2368]. c [2364]. d [2464]. E_{0} [2584]. ℓ [207]. η [725]. f8 [2146]. F_{q} [1540]. F_{q^{m}} [1540]. GF (2) [1539]. ≫ [2672]. i [1410]. J [1781]. K [2815, 161, 2480, 1107, 1797, 1371, 661, 667, 1911, 1026]. k, + [656]. kDNF_{n}^{s} [1112]. L [2621, 1098]. L_{1} [2364]. L_{2}(R) [2246]. ≪ [2672]. log n [669]. M [2748, 1721, 1631, 194, 1344, 917, 2285, 1557]. F_{2^{607}} [2137]. F_{2^{r}} [1544]. F_{q} [1556]. Z_{4} [1539, 1546]. Z_{m} [558]. M_{\×∫} (F_{⨿}) [2612]. T [2043]. IR^{2} [966]. IR^{3} [966]. μ [982]. N [564, 230, 2585, 1187, 1352, 2751]. n ≡ 1 mod 4 [2136]. O(n^{2k+1}) [669]. O(n log^{2} n) [1404]. ω [1485]. P [1785, 616, 1790, 1796, 1798, 1782, 1800]. Ψ [195]. Q [2491, 628]. r [2767]. S^{3} [50]. t [2587]. τ [2627]. ∨ [794]. X [1817]. x + c [1556].

* [2096].

-Approximative [669]. -Band [2285]. -Binomial [628]. -CNF [1911]. -Codes [1546]. -Commerce [194, 1721, 1631]. -Comparability [1370]. -Component [161]. -Conjugate [1853]. -Convex [1344]. -Coverings [2786, 2368]. -D [1002, 861, 269, 1681, 1996, 990, 1009]. -Decomposition [669]. -Designs [2627]. -Dimensional [443, 755, 2464]. -Discrepancy [1396]. -Distance-Hereditary [656]. -Distributive [794]. -Dominating [667]. -Exclusion [207]. -Factorizations [661]. -Fold [2815]. -Hardness [1782]. -Languages [1485]. -Learning [2491]. -Level [658]. -Linear [1540]. -Mall [1721]. -Manifolds [1979]. -Means [2364, 2480]. -Median [1407]. -Metric [2647]. -mode [1410]. -Monotonicity [1781]. -Out- [2751]. -Party [564]. -Pervasive [195]. -Queen [230]. -reps [917]. -Resilient [2587]. -SAT [1854, 1856]. -Satisfiability [1107]. -Sequences [1557]. -Space [1026, 2364]. -Subset [1797]. -systems [1098]. -tegration [1817]. -Terminal [1371]. -Time [1130].

1-st [2748]. 1V [552].

2 [225, 2055]. 2000 [2757]. 2001 [2131]. 2005 [215].

3.0 [1630]. 3GPP [2146, 554].

4 [1239, 2301, 365, 330, 225, 229, 1233, 231]. 4.4BSD [574].

7 [228].

= [184].

A. [2005, 2004]. A.D. [2180]. A1 [1810]. A2 [1811]. ABACUS [1954]. Abbreviations [1811]. ABC [425]. Abduction [2222, 792]. Abductive [640, 1839]. Ablation [951]. AbsCon [1904]. Absolute [544]. Abstract [1816, 203, 2015, 2534, 2006, 1792, 2212, 2048, 2703, 1359, 1320, 699, 1800, 1465, 1398, 1400, 2231]. Abstracting [1276]. Abstraction [175, 2202, 1904]. Abstractions [2692, 1140]. ABT [1864]. AC [1884, 1875, 2084]. AC-Matching [1875]. Academic [2035]. Academy [254]. Acceleration [761]. Accelerator [552, 555]. Access [1748, 1284, 538, 1653, 1659, 473, 453, 1236, 1937, 1204, 768, 1771, 522, 440, 592, 352, 560, 1749, 1182, 1647, 1654]. Accessing [1429, 235]. Accomplishments [373]. Account [1970]. Accountability [1641]. Accounting [1971]. Accuracy [1043, 1045, 910, 960]. Accurate [2013]. Accurately [1159]. Acetabular [1071]. Achievable [1602]. Achieve [2140]. Achieving [1396, 434]. ACI [2198]. ACI-Unification [2198]. ACISP2001 [2126]. ACL2 [2521]. ACMD [515]. Acoustic [2801, 289, 907]. Acquired [111]. Acquirement [2392]. Acquiring [2479, 697, 124]. Acquisition [918, 2547, 1070, 2449, 2485, 930, 954]. Acronyms [1811]. across [604]. Act [2087]. Actant [86]. Actin [105]. Action [191, 2919, 2673, 2397, 193]. Actions [633, 2495, 617, 2350, 2504, 1361]. Activation [2550]. Active [810, 808, 1687, 804, 856, 127, 916, 1048, 807, 805, 813, 1798, 63, 809, 1025, 811, 587, 286, 1050, 2291]. Active-Network-Powered [804]. Activities [594, 405]. Activity [1686, 1279, 983, 734, 271, 2392]. Actor [1422]. Actors [393, 2096]. Actuator [68]. Actuators [1068]. Acyclic [784, 2195, 1200]. Ad [1248, 1176]. Adaptability [163]. Adaptable [481, 381, 2379]. Adaptation [100, 1936, 159, 2546, 2738, 2450, 2479, 771].

Adapted [2853, 1114]. Adapting [96]. Adaptive [1581, 1011, 723, 1172, 2435, 1277, 203, 396, 161, 112, 922, 2287, 2478, 2480, 2620, 337, 196, 1964, 720, 465, 2282, 2150, 2317, 2242, 486, 2777, 2791, 2492, 1350, 2091, 2758, 160, 941, 1135, 2014, 828, 1142, 354, 64, 811, 2553, 852, 346, 1863, 223, 426, 301]. Adapts [2295]. Adder [2625]. Adding [2530]. Addition [2075, 2874]. additive [2370]. Addressable [1740]. Adjoining [2872]. adjoint [1839, 2569]. Admission [1235, 339]. Adoption [2447]. ADST [1771]. ADT [2027]. Adult [921]. Advanced [1656, 741, 527]. Advances [1758, 1025]. Advantage [1745]. Adversarial [1575]. Adversary [199]. Advertisement [570]. Advertising [631]. Advice [374]. AEGIS [804]. Aerial [102, 2372]. AES [2135, 2713]. Affect [2436]. Affected [204]. Affects [2466]. Affine [880, 779, 2640, 116]. Affinities [1521]. Affordance [1329]. After [1786, 396]. against [804, 2151, 2760, 1643, 2717, 559, 2496, 2631, 1613, 566, 2142]. Age [460]. Agendas [590]. Agent [80, 81, 200, 2541, 84, 78, 1255, 2313, 2314, 918, 2355, 2561, 1947, 2099, 1256, 389, 1940, 1933, 637, 83, 1946, 1941, 1934, 208, 2295, 2356, 2343, 463, 460, 2568, 464, 2299, 465, 2830, 401, 2348, 2177, 2171, 2173, 2168, 2091, 2493, 642, 461, 2564, 1931, 457, 2344, 1932, 1447, 162, 1945, 467, 2507, 2510, 873, 247, 90, 898, 646, 969, 1770, 2337, 2559, 640, 687, 2307, 2308, 2354, 2562, 2095, 2563, 638, 1104, 448]. Agent-Based [2314, 2356, 2343, 464, 2348, 2091, 2344]. Agent-Enhanced [1255]. Agent-Mediated [2313]. Agent-Oriented [642, 1447]. Agents [2302, 79, 2090, 2558, 1936, 2315, 462, 1731, 2560, 396, 2086, 2347, 1937, 198, 1944, 1235, 2345, 1603, 413, 2307, 196, 74, 2489, 82, 2306, 2170, 94, 647, 87, 73, 1942, 2346, 2041, 77, 2350, 1939, 2565, 88, 1655, 2509, 2092, 1948, 2352]. Agentspace [645]. Aggregate [1431]. Aggregated [1751]. Aggregates [1835]. Aggregation [1158, 722, 435, 1528]. Aggregations [1441]. Agile [1900]. AGORA [408]. Agreement [1755, 1255, 2633, 1361]. AGV [727]. AI [400, 2503, 775]. AI-Based [400]. Aided [968, 497, 1057, 1053, 994, 721, 419, 1056, 889, 510, 511, 2132, 951, 797]. AIDS [520]. AIMD [1746]. Aiming [217]. Air [215]. Aircraft [2775, 1914, 2323]. Alarm [1077]. Album [220]. Alert [1156]. Alerts [1158]. Algebra [2893, 2154]. Algebraic [169, 1566, 2692, 2907, 2695, 2589, 1573, 2038, 1551, 2706, 1537]. Algebraic-Geometry [1551]. Algebras [1213]. Algorithm [2610, 1563, 1620, 2853, 1619, 723, 750, 737, 1732, 1347, 2727, 52, 867, 202, 759, 744, 1246, 1940, 106, 2582, 1399, 2347, 2158, 1466, 2928, 1523, 65, 2074, 347, 1776, 1101, 748, 2598, 2304, 2261, 2325, 240, 2682, 1051, 2404, 1215, 1579, 2826, 905, 1858, 2818, 524, 676, 308, 338, 2122, 924, 1924, 2101, 735, 773, 2248, 2367, 307, 1357, 1100, 796, 212, 1199, 1374, 1570, 2712, 327, 766, 441, 246, 1365, 2324, 222, 1247]. Algorithmic [627, 2007, 1211]. Algorithms [1373, 2544, 2863, 1401, 2686, 655, 1371, 2067, 2016, 1954, 1846, 663, 1494, 1372, 2688, 1564, 1385, 668, 2072, 207, 1586, 1403, 613, 2486, 1336, 333, 1568, 1576, 1366, 1978, 1407, 510, 2260, 900, 1383, 1392, 1916, 2575, 2678, 554, 1950, 2594, 2685, 519, 1370, 1387, 1109, 1876, 726, 1356, 2128, 1092, 2141, 2552, 1578, 1537, 1146, 1520, 774, 2011, 846, 1200, 1361, 1386, 2061]. Aligning [2870, 817]. Alignment [886, 328]. All-Optical [1351]. Alloarthroplasty [1071]. Allocation [350, 1127, 1131, 2526, 273, 1919, 1901, 2083, 1590, 1966, 286, 1335]. ALM [1173]. Almost [2622, 1555]. Almost- [1555]. Almost-Certainly [2622]. Along [696]. Alpha1 [2634]. Alphabet [91]. Alphabets [2183]. Alternate [2763]. Alternating [2807, 1836]. Alternation [1499]. Alternatives [1833]. Altruistic [1625].

Alzheimer [1053, 1016, 1022, 894, 820]. AMBAR [1653]. Ambients [1203, 608]. Ambiguity [697]. Ammunition [2277]. Among [250, 1363, 975, 1355, 2366, 2499, 785]. Amplification [539]. Amsterdam [2669]. AN-Codes [2621]. Anaerobic [714]. Analog [1092]. Analogue [1100]. Analogy [1512]. Analyses [2260]. Analysing [755]. Analysis [733, 495, 1377, 925, 1222, 2442, 1187, 499, 2263, 819, 491, 2739, 1189, 962, 1002, 1626, 1679, 1830, 2042, 493, 2231, 918, 2375, 971, 800, 68, 1256, 502, 922, 2584, 2384, 2534, 1236, 2820, 2688, 2226, 2227, 986, 541, 2156, 823, 2745, 995, 1466, 268, 2022, 459, 331, 344, 1046, 2403, 1226, 1012, 2709, 507, 130, 2289, 827, 54, 908, 1349, 2482, 137, 1777, 1645, 1269, 920, 2753, 310, 1240, 2278, 74, 2808, 1614, 1611, 1280, 516, 2116, 2519, 2321, 282, 815, 2701, 795, 104, 2372, 1454, 520, 522, 523, 461, 2257, 2274, 525, 2795, 841, 263, 2198, 1630, 2854, 1517]. Analysis [716, 1100, 919, 2228, 2114, 981, 2712, 573, 236, 125, 535, 158, 915, 5, 2351, 907, 364, 2283, 1322, 1532]. Analytic [2894, 2268]. Analytical [1180]. Analyzer [2220, 300, 2884]. Analyzing [57, 2043, 784, 1497, 1403, 2343, 1144, 435]. Anaphora [2859, 2858]. Anaplastology [885]. Anatomical [962, 939, 505, 1681, 997]. Anchor [327]. Andorra [2221]. Anesthesia [730]. Angiograms [1027]. Angiography [836]. Angioplasty [854]. Angles [1680]. Animal [1036]. Animated [397, 78]. Animating [94]. Animation [76, 1188, 365, 407, 399, 2503, 1662]. Animations [2426]. Anisotropic [977, 105, 1689]. Ankle [960]. Annealing [867, 2685, 771, 2367]. Annex [1304, 1305, 1306, 1308, 1309, 1310, 1315, 1313, 1311, 1312, 1314, 1307, 1303, 1316]. Annotation [326]. Annotations [2225, 2228]. Anomalies [2280]. Anomaly [176, 1251]. Anonymous [2100, 2753]. Answer [10, 2879]. Ant [1100, 90, 2305].

ANT-Based [2305]. Antenatal [921]. Antennas [1108]. Anthologies [1521]. Anticritical [677]. Ants [1896]. Any [2714, 1221]. Any-Lock [1221]. Anytime [2679]. AOCS [1693, 1706, 1699]. Aortoiliac [898]. Aperiodic [1557]. API [1182]. APIs [2616]. App [2525]. Apparatus [842]. Appearance [874, 971, 129, 97]. Appearance-Based [971, 129]. Appendix [1811, 1810, 2177, 1715]. Apples [2530]. Appliances [1283, 2431, 605]. Applicable [1449]. Application [745, 2626, 1154, 2691, 686, 2424, 808, 449, 2258, 2833, 1068, 1256, 1021, 389, 2218, 31, 1958, 1892, 2245, 2486, 1046, 1915, 511, 1238, 2278, 2265, 2887, 1134, 1028, 1902, 2106, 1241, 1344, 2167, 1683, 647, 406, 1740, 1016, 2367, 1025, 2503, 716, 215, 774, 1972, 458, 1182, 2284, 1061, 2280]. Application-Integrated [1154]. Application-Level [1740]. Application-Specific [1241]. Applications [758, 810, 563, 2302, 1237, 1787, 2739, 1720, 165, 570, 136, 427, 1270, 1282, 2792, 1961, 667, 2434, 1586, 2401, 194, 1403, 1603, 2767, 2765, 1015, 2600, 2050, 1730, 1065, 2279, 2077, 1870, 2926, 1116, 1272, 924, 1938, 2127, 2819, 1800, 731, 1130, 2253, 2000, 592, 1980, 162, 1155, 1818, 1636, 369, 1647]. Applied [1871, 1069, 918, 508, 510, 1392, 1684, 519]. Applying [2423, 255, 492, 1630]. Approach [2462, 57, 1445, 1734, 417, 934, 26, 627, 646, 633, 1412, 2326, 2870, 2802, 548, 547, 1126, 378, 2355, 724, 1947, 1976, 1840, 659, 840, 2309, 2015, 2029, 2794, 2347, 1861, 2828, 2788, 1844, 2420, 1615, 967, 807, 383, 1414, 255, 2533, 2914, 469, 2774, 1141, 509, 2821, 1610, 2817, 2304, 1327, 2381, 444, 2700, 2282, 1681, 2835, 1839, 3, 2413, 1901, 2824, 2781, 1684, 2170, 757, 2091, 2298, 941, 521, 2826, 2858, 461, 1432, 2455, 1443, 171, 849, 1419, 2041, 529, 2197, 2665, 1231, 865, 2393, 1624, 2379, 139, 1878, 2841, 1433, 2565, 638, 446, 877, 1009]. Approach [1288, 781, 766, 2040, 1339, 34, 109]. Approaches [1969, 2915, 2793, 2318, 2825, 913, 790, 251, 1196]. Approximability [1765]. Approximate [558, 652, 2689, 1049, 793, 65, 1501, 1406, 2830, 2378]. Approximately [1775]. Approximating [286, 1400]. Approximation [2058, 1992, 1358, 1814, 1345, 664, 668, 1346, 1384, 2836, 1209, 878, 1997, 1357, 948, 2061]. Approximations [966, 2837, 1359]. Approximative [1772, 669]. Apriori [2404, 2548]. APRIORI-C [2548]. Aqueous [2928]. Arbitrary [2716, 2742, 151, 1897]. Arbitrary-Length [2742]. Arc [1884, 2080]. Architectural [2698]. Architecture [1969, 1099, 808, 89, 1171, 2559, 1174, 161, 1106, 1083, 344, 1944, 2386, 1281, 2036, 1964, 950, 1902, 2493, 2456, 2145, 1167, 2510, 1251, 780]. Architectures [2415, 1148, 1139, 1823]. Archival [2253]. Area [242]. Argumentation [1913, 2562, 2388]. Arithmetic [553, 2621, 2102, 2592, 2242]. Arithmetic-Modular [2621]. ARM7 [555]. Arms [644]. Arrangements [2071]. Array [2468, 361, 1093]. Arrays [2075]. Art [407, 1267, 2411, 1727]. Artefacts [585]. Arterial [849, 898]. Artery [833, 1027, 1072]. Arthroplasty [863]. Articulated [874]. Artifacts [2237, 848, 1008]. Artificial [2462, 636, 75, 2086, 2347, 1266, 1265, 1652, 2550, 1096, 1896, 1800]. Arts [408]. ARX [2803]. Asiacrypt [2757]. ASIC [2713]. ASL [2701]. Aspatial [1417]. Aspect [169, 156, 374]. Aspect-Oriented [169, 156, 374]. AspectJ [175]. Aspects [2907, 577, 1385, 1828, 1783, 180]. Assembling [576]. Assembly [119, 1227, 1379, 139, 2932]. Assertion [2066]. Assess [42, 846]. Assessing [876, 279]. Assessment [982, 2679, 753, 12, 2483, 1684, 1683, 7, 11, 1052, 979]. Assignment [1376, 666, 1914, 1406, 1605].

Assist [2450]. Assistance [1054]. Assistant [1069, 1018, 452, 438, 487, 588]. Assisted [1067, 860, 857, 508, 855, 1019, 1072, 949]. Assisting [1506]. Association [42, 61, 2402, 752, 1501, 67, 2381, 2383, 278, 2513]. Associative [618, 1585]. Assumptions [2149]. Astrophysics [1503]. Asymmetric [295, 2117]. Asymmetry [893]. Asynchronicity [617]. Asynchronous [1081, 623, 1095, 2074, 1922, 1928, 1596]. Asynchrony [1124]. AT-Free [670]. Atherosclerotic [908]. Atlas [1039, 903, 961, 864]. ATM [768]. Atom [2200]. Atomic [2223]. Atrium [1690]. Attachment [811]. Attack [1639, 2597, 1643, 2116, 2717, 2758, 1665, 2596, 2759]. Attacking [2640]. Attacks [2764, 2616, 804, 2638, 1159, 2151, 2143, 2760, 2620, 2757, 559, 2144, 2631, 1613, 1161, 1622, 1646, 2615]. Attempt [588]. Attention [2310]. Attitude [1693]. Attracting [14]. Attractiveness [1328]. Attributes [2366, 1312]. Auction [2748, 1731, 2113]. Auctions [1732, 565]. Audience [408]. Audio [325, 2733, 288, 283, 239, 1274, 589, 1283, 277, 429, 312]. Audiophonic [2232]. Auditable [570]. Augmentation [1372, 1065, 1354]. Augmentative [290]. Augmented [895, 897, 2421, 154, 1060, 2433, 843, 899, 974]. Augmenting [17, 2338]. Aura [2409]. Authenticated [563, 2148, 2147]. Authentication [1748, 2154, 2745, 1666, 2112, 2110, 343]. Author [32, 1201, 1230, 1260, 1268, 1289, 1330, 1342, 1408, 1458, 1486, 1533, 1574, 1612, 1669, 1691, 1738, 1753, 1784, 1802, 1813, 1845, 1930, 1949, 1957, 1975, 2003, 71, 2056, 2085, 2097, 2130, 2165, 2230, 2286, 2331, 2407, 2437, 2461, 2518, 2540, 2581, 2609, 2648, 2676, 2690, 2707, 2729, 2740, 2768, 2842, 2891, 2909, 2933, 95, 155, 181, 197, 213, 371, 384, 409, 490, 537, 575, 607, 635, 648, 678, 712, 814, 1073, 1087, 1117, 1152, 1165, 1184, 400, 489]. Authoring [599, 2317, 392, 299]. Authorities [551]. Authorization [1653, 1971]. Autocorrelation [2120, 1554, 2157]. Autocovariance [2233]. AutoMarkup [2888]. Automata [1186, 1768, 1188, 1788, 2924, 1218, 2534, 2797, 1585, 1194, 1366, 1792, 615, 1219, 1216, 1781, 2182, 1198, 2834, 1199, 1200]. Automated [2033, 1123, 548, 1887, 867, 825, 2020, 998, 2850, 1503, 1063, 1269, 2201, 1016, 984, 948, 573, 2884]. Automatic [216, 923, 497, 901, 89, 962, 319, 1595, 1907, 2871, 986, 823, 2401, 1058, 2481, 1020, 2261, 249, 515, 241, 2269, 444, 2857, 2685, 439, 964, 2880, 824, 988, 1029, 949, 835, 445, 2876, 277, 907, 2323, 322, 1051]. Automatically [2888, 2503]. Automaton [1192, 2182]. Autonomic [1161]. Autonomous [397, 2558, 1731, 396, 389, 1262, 1102, 1103, 2774, 782, 617, 87, 1762, 77, 138, 2459, 1147, 1020]. Autonomy [2557, 421]. Auxiliary [2860, 2659]. Availability [1081]. Avatamsaka [2349]. Avatar [83]. Avatar-Agent [83]. Avatars [87, 77]. Average [2590, 420]. Average-Clicks [420]. Aviation [60]. Avoiding [1600]. Aware [165, 164, 688, 598, 484, 589]. Awareness [2395, 1183]. Axiomatic [1976]. Axiomatizing [423]. Axis [113, 867, 915].

B [1271, 1048, 2657, 1304, 948, 2117]. B-spline [948]. B-splines [1048]. B2B [183, 1412]. Baby [1462, 1491]. Back [725, 1124, 77]. Backbone [1248]. Backbones [1856]. Backfilling [1339]. Background [89, 2354, 2656, 129]. Backscatter [2801]. Backtracking [1918]. Backwards [2226]. Bacterial [737]. Bag [1035]. Bag-of-Particles [1035]. Bagging [50]. Baire [1362]. Bakery [1776]. Balance [624]. Balanced [2910, 626, 1348]. Balancing [1602, 863, 1110, 421]. Balloon [854, 997]. Ballot [2750]. Band [2238, 2285].

Bandwidth [1745, 1242, 1173, 1606, 286]. Bangla [2799]. Banks [2258, 2271, 2272, 2273, 2262]. Basal [903]. Base [2, 13, 1024, 532]. Based [2438, 1445, 1112, 745, 110, 495, 1451, 26, 430, 1278, 2720, 1686, 1879, 2422, 1222, 2090, 2691, 267, 778, 1885, 2785, 499, 1057, 2263, 750, 294, 2065, 2616, 1720, 2469, 2536, 1729, 921, 362, 1653, 61, 400, 2802, 2801, 320, 1687, 2303, 228, 462, 365, 356, 418, 2435, 311, 2314, 982, 740, 2641, 2042, 2154, 2140, 1983, 2355, 571, 1865, 1659, 262, 2099, 971, 2447, 1256, 1246, 1259, 577, 1435, 321, 2029, 1940, 1039, 106, 2715, 1933, 1262, 481, 1890, 1899, 2721, 2534, 1585, 367, 1891, 1048, 772, 1083, 2296, 2102, 2256, 1833, 2235, 2525, 1013, 1213, 478]. Based [96, 1254, 721, 268, 2556, 1522, 2385, 2779, 1457, 2849, 344, 292, 260, 1047, 2898, 2356, 1046, 221, 2786, 1414, 1012, 289, 1131, 255, 1757, 333, 483, 101, 413, 238, 2034, 507, 187, 1030, 2308, 998, 2591, 2481, 328, 454, 2480, 1625, 1502, 137, 2364, 67, 2252, 682, 1610, 1645, 863, 2343, 2304, 1760, 1718, 463, 434, 460, 1133, 1240, 720, 314, 38, 351, 464, 1660, 861, 683, 2126, 1258, 1327, 958, 515, 950, 2269, 2240, 269, 1652, 1318, 1663, 1681, 881, 2242, 2316, 574, 1001, 486, 2348, 975, 2491, 348, 1319, 2791, 315, 1045, 2793, 7, 488, 1650, 2091, 2318, 2758, 936, 489]. Based [1215, 2123, 2576, 2277, 287, 1921, 359, 340, 285, 844, 714, 1109, 2032, 46, 961, 308, 338, 334, 2455, 2754, 1273, 129, 1272, 1799, 754, 849, 2014, 526, 828, 817, 568, 2672, 762, 878, 529, 138, 883, 2234, 1231, 735, 8, 363, 318, 2127, 2305, 457, 2378, 2552, 141, 1144, 556, 435, 2854, 1516, 139, 2344, 1527, 811, 1447, 122, 1181, 274, 2804, 919, 2789, 2823, 2399, 1088, 643, 2565, 2810, 2811, 965, 560, 785, 1820, 2243, 1163, 2045, 279, 1945, 467, 458, 277, 243, 2031, 370, 1288, 868, 2092, 278, 1361, 730, 1622, 907, 2324, 2806, 369, 364, 2511]. Based [23, 1244, 2319, 1647, 2283, 1247, 2885, 1654, 2061, 1128, 34, 1322, 2606, 1420, 482, 1455, 1419]. Baseline [2852]. Bases [2222, 2246, 2207, 1353, 2773, 764]. Basic [1190, 1891, 1390]. Basis [2517, 690, 1501, 1633, 2599, 2615, 2441]. Basket [2403, 2321]. Batch [1375, 1345, 543, 1638]. Batched [2077]. Bayes [2358, 2783]. Bayesian [494, 423, 65, 483, 463, 2678, 2378, 2841, 2508, 370, 536, 34]. Bayesian-Network [536]. Baysian [724]. BCH [297]. BDDs [1854]. BDI [2495]. Be [1218]. Beam [1045, 2468]. Bean [1271]. Beans [1271]. Beats [1831]. Bed [688]. before [2552, 535]. Behaved [426]. Behavior [30, 1873, 2384, 2652, 2926, 2675, 2394, 1972, 573, 2031]. Behaviour [2694, 1739, 2001, 2419, 66]. Behaviours [1325]. Belief [2090, 1626, 2573, 2517, 2899, 2562, 791]. Believable [2301]. Benchmark [1080]. Benchmarks [1888]. Benders [1846]. Benefit [137]. Bent [1559]. Beowulf [1823]. Best [1523, 1500, 254]. Better [2438, 1084, 2532, 2631]. Between [2071, 2410, 1872, 1081, 1404, 1010, 657, 482, 585, 1281, 929, 717, 470, 1267, 2780, 2793, 404, 2128, 935, 1474]. Beyond [186, 2070, 2078, 1761]. Bi [1667, 1749, 1214]. Bi-directional [1667, 1749]. Bias [911, 984]. Biaxial [960]. Bibliography [2660]. Bid [565]. Bidders [565]. Bidding [1731]. Bidomain [1689]. Bijections [2928]. Bijective [2585]. Bilevel [2781]. Bilingual [2873]. Bin [1401]. Binary [758, 2064, 1566, 1886, 1192, 2008, 2928, 1363, 2190, 624, 142, 2321, 1554, 2917, 530, 2800, 846, 2206, 2646, 1857, 1867, 1924]. Binding [1213, 2022, 176, 1212, 2228]. Binding-Time [2022, 2228]. Binomial [628]. Bio [1754, 1785]. Bio-language [1785]. Bio-sciences [1754]. Biogas [714]. Biological [1428, 1266]. Biologically [716]. Biomechanical [957, 962]. Biomedical [1005]. Biometric [2792]. Biometric-Applications [2792]. Biometrics [2792, 367]. Biopsies [843].

Biopsy [855, 864, 842, 996]. Bioreactor [2806]. Biorthogonal [2246, 2285]. BioWatch [1097]. Bipartite [652, 655, 1374]. Biped [1101]. Biplane [1672]. Bisimulation [2920, 1832, 1220]. Bit [350, 273, 1561, 2127, 1232]. Bite [2088]. Biting [1539]. Bits [2336]. BizBuilder [188]. Blacklisting [2613]. Blending [976]. Blind [2164, 313, 1617, 347, 269, 2747, 1613]. Blindable [2161]. Bloat [760]. Block [2610, 1619, 2098, 619, 2143, 347, 2646, 270, 1596, 2142, 272]. Block-DCT [1619]. Block-Deterministic [619]. Block-Matching [272]. Blocking [2237, 2238]. Blood [916, 101, 2481, 849, 873, 888, 898]. Bluetooth [1176]. Board [552]. Body [767, 961, 1029, 990]. Bone [955, 973]. Bookmarks [419]. Boolean [1112, 2219, 2121, 1893, 1829, 2199, 1836, 2120, 2082, 2204, 1468, 2157, 2011, 1636, 1465, 2585]. Boosted [250]. Booster [1609]. Boosters [2496]. Boosting [1466, 1896]. Bootstrapped [2821]. Both [2237, 2346]. Bottom [2209, 2346]. Bottom-Up [2209]. Bound [2049, 2314, 1869, 2590, 2627, 632]. Boundaries [682, 988, 877]. Boundary [1673, 1001, 2234, 1672, 222]. Bounded [1358, 2068, 660, 2181, 2639]. Boundedly [1990]. Boundedness [2013]. Bounding [1494, 1776]. Bounds [2057, 1369, 538, 2688, 668, 2073, 2929, 613, 615, 1867, 1579, 1868, 2110, 557, 2710]. Box [770, 2129, 2145]. Boxed [1203, 2356]. Boxes [1379, 557]. Brachytherapy [856]. Braid [2139]. Brain [923, 921, 902, 944, 1039, 818, 995, 845, 821, 729, 983, 943, 941, 911, 875, 953, 919]. Branch [1593, 2314, 1954, 1955, 1950, 1847]. Branch-and-Bound [2314]. Branch-and-Check [1847]. Branch-and-Cut [1954, 1950]. Branching [2476, 1466, 2914, 2680]. Breaking [1852, 1917]. Breast [1003, 842, 1055, 963].

Bricks [2336]. BRIDGE [1281]. Bridging [2288, 2830]. Brief [2380, 2377]. Bringing [2086]. Broadband [1082]. Broadcast [283]. Broadcasting [1169, 1349, 1170, 2118]. Broker [571, 1242]. Broker-Based [571]. Brokering [1083]. Brokers [1722]. Browser [325]. Browsing [1444, 606]. Bruijn [675]. BSC [1570]. BSP [1767]. BTTC [2818]. Buffer [349, 1231, 286]. Buffers [1607]. Build [1266, 469, 2519, 72]. Building [2302, 2, 13, 397, 481, 2448, 1827, 62, 2774, 38, 2317, 1744, 1137, 1923, 1938, 141, 1426, 1818, 2092, 2426, 1722]. Bureau [140]. Bursty [1236]. Bus [1139]. BUSHWHACK [1357]. Business [1735, 31, 1456, 184, 2298, 2274, 190, 1447, 1409, 1719, 23]. Business-Oriented [2298]. Buyer [2112]. Bypass [858, 1072]. Byte [2527].

C [2548, 729, 2701, 2658, 953, 1305, 2810, 981]. C-Means [729, 2810, 981]. Cache [417, 421, 301]. Caching [1271, 1270, 285, 471]. Cactus [658]. CAD [826, 1054, 517, 2046]. CAD/CAM [2046]. Cadaver [952]. Calculation [753, 2886]. Calculations [1595]. Calculi [2359]. Calculus [2566, 1223, 610, 1204, 2023, 632, 1866, 707, 1228, 1224]. Calibration [926, 895, 992, 2822, 1043, 2683, 1090, 949]. Call [478, 612]. Call-by-Value [612]. Caller [2735]. Calls [1660]. CAM [2046]. Camellia [1639, 2142]. Camera [136, 152, 107, 815, 364]. Cameras [361]. Can [396, 767, 1824, 1764]. Canal [982]. Cancellation [2823]. Cancellative [2084]. Cancer [993, 509, 1054, 2809, 864, 947]. Candidate [2508]. Candidates [556]. Cannot [2008]. Canonical [1377, 2694]. Cantonese [283]. Capabilities [2450, 140, 88, 1973]. Capability [12]. Capable [2316]. Capacitors [756].

Capacity [1602, 1242]. Capitalization [2731]. Captions [328]. Capture [586, 592]. Capturing [1855]. Carcinomas [994, 522]. Cardiac [1670, 1686, 1011, 1687, 1031, 872, 920, 858, 958, 1681, 881, 1685, 1677, 1066, 841]. Cardinality [1401, 1825, 742, 2908]. Cardinality-Path [1825]. Cardio [988]. Cardiographic [986]. Cardiovascular [854]. Cards [2645, 2767, 1360, 1728]. Care [478]. Career [4]. Career-Long [4]. Caring [2318]. Carotid [833]. Cartan [1571]. Cartesian [651]. Cartographic [699]. Carving [355]. CAS [945]. Cascade [2405]. Case [2148, 1649, 1934, 1498, 2483, 7, 439, 1109, 11, 2464, 1587, 529, 8, 530, 1517, 1528, 2585, 606, 1425]. Case-Based [1109, 529]. Cases [1695]. Cash [569, 1663]. CASL [2699]. Casual [2410]. Categoric [774]. Categorical [52, 521]. Categories [2209, 636]. Categorization [769, 319, 444]. Categorizing [473]. Category [1362]. Catheter [869, 997]. Catheters [1023]. Causal [1598, 210]. Causality [1285]. CBR [2479]. CBS [2438, 2438]. CDIS [1160]. CDMA [773]. CDR [1076]. CE [898]. CE-MRA [898]. CeBASE [2, 13]. Cell [130, 2481, 1977, 1991]. Cells [1891, 523, 1096]. Cellular [1602, 2797, 1585, 1235, 218]. Census [2218, 332]. Cent [1405]. Cent-dian [1405]. Center [1618, 630, 478, 979]. Centered [2290, 219, 2413]. Centerline [948]. Centre [960]. Centres [715]. Centric [1126, 1647]. Century [403]. Cerebral [1013, 980]. Certain [2893, 2763, 2762]. Certainly [2622]. Certificate [550]. Certificates [2719, 2161]. Certification [1210, 551]. Certifications [2049]. Certified [2720, 2111]. Certifying [2213]. CGI [2531]. Chain [1548, 1878, 1736, 1343, 5]. Chaining [2102, 801]. Challenge [70, 2400]. Challenges [2439, 597, 1675, 154, 73, 4, 373, 1143, 1145].

Challenging [2218]. Chance [2395, 2388, 2398, 2387, 2389, 2394, 2399]. Change [9, 118, 318, 2874]. Changes [1604, 172, 821]. Changing [4]. Channel [2625, 2764, 353, 2338, 2760, 539, 559, 1220, 302, 369]. Channelling [2205]. Channels [1169, 2611]. Chaotic [2628, 2127]. Chapter [1805, 1808, 1809, 1804, 1806, 1803, 1807]. Character [397, 2261, 2260, 399]. Characterise [535]. Characteristic [1571, 1363, 2116]. Characteristics [1172, 1340, 1314]. Characterization [33, 1540, 2797, 1193, 1058, 830, 2247]. Characterizations [1752, 1636]. Characters [397, 1571, 400, 92, 398]. Cheating [2105, 1628]. Check [2636, 1847]. Checkerboard [1396]. Checking [1766, 2184, 2065, 1217, 2070, 2665, 1927, 2009]. Checkpoint [287]. Checkpointing [158]. Checks [2705]. Chemical [1786, 798, 970]. Chinese [2885, 356, 2261, 2260, 56, 2754, 291, 233, 1654]. Chip [1091, 2275]. Choco [1892]. Choice [2557, 2351, 1850]. Choosing [2282, 2552]. Chordal [1355]. Chosen [2725, 2742, 2151, 2620, 2757, 2758]. Chosen-Ciphertext [2725, 2742, 2151, 2620, 2758]. Chosen-Plaintext [2757]. CHR [1863]. Church [114]. Ciao [2541]. Ciliates [2931]. Cine [1683]. Cipher [2618, 1639, 2634, 2127, 2142]. Ciphers [2716, 2098, 2618, 2143]. Ciphertext [2725, 2742, 2151, 2620, 2744, 2758]. Circuit [1100]. Circuits [2062, 1105, 1198, 1100]. Circulant [1546]. Circulation [1670, 211]. Cited [432]. City [2288, 2484]. Civil [2278]. Clarification [738]. Class [1185, 2696, 2668, 179, 1552, 1800, 2787, 128, 133]. Classes [2417, 2414, 1645, 1478]. Classic [1403]. Classical [2892, 2203, 541, 1521]. Classification [2276, 745, 1810, 2470, 362, 2801, 1438, 746, 2548, 130, 2481, 2488, 310, 306, 893, 518, 2453, 523, 526, 29, 2809, 762, 528, 529, 492, 2378, 2501, 894, 796, 2789, 445, 1055, 2324, 2283, 2885]. Classification-Driven [893]. Classifier [724, 2809, 50]. Classifiers [2545, 2802, 754, 128, 44, 53, 224]. Classroom [298]. Clause [1893, 2673]. Clauses [1907, 1829]. Clicks [420]. Client [1748, 2435, 2749, 1865, 427, 1946, 2318]. Client-Server [1748, 1865, 427]. Client/Server [2749, 1946]. Clinical [1067, 500, 860, 1013, 885, 998, 899, 2524, 835]. Clipping [1497]. Clique [653, 657, 660, 1355]. Clique-Width [657, 660]. Cliques [1387]. Clock [1880, 2124]. Clock-Controlled [2124]. Close [1820]. Closed [328, 67, 1454, 1990]. Closest [905, 1380]. Closing [2035]. Closure [2194, 1792, 2084]. CLP [1908, 1878]. Cluster [57, 750, 366, 359, 1076, 766, 447]. Cluster-Based [750]. Cluster-Defining [57]. Clustered [344, 511]. Clustering [792, 2812, 473, 106, 2059, 2385, 459, 2732, 2480, 729, 2364, 2252, 514, 1525, 38, 484, 2813, 2549, 1514, 2551, 2810, 2811, 774, 981, 2324, 1398]. Clustering-Based [2810, 2811]. Clusters [57, 1081, 1589, 509, 1580, 1333]. Clutter [115]. CMM [10]. CNF [1911]. CNFs [1363]. Co [1335, 2352, 2561, 2722, 2562, 2091, 1966]. Co-Allocation [1966, 1335]. Co-evolution [2352]. Co-operation [2091]. Co-operative [2562]. Co-operatively [2722]. Co-ordination [2561]. Cocyclic [1563]. Code [2625, 1159, 1106, 1256, 2073, 2745, 2752, 2527, 1658]. Codec [293]. Coded [1573, 297]. Coder [267, 341]. Codes [2610, 620, 1566, 1538, 1541, 1544, 294, 353, 1595, 1540, 1560, 2621, 1550, 1546, 1545, 1547, 2614, 1548, 2587, 2930, 1549, 1542, 2611, 2646, 2622, 2635, 1534, 2639, 2586, 1535, 2636, 2110, 1551, 2109, 1543, 1536, 2612, 1537, 1622, 2623, 1539]. Codes-Theme [1535]. codewords [1568]. Coding [353, 268, 252, 1179, 2241, 2628, 273, 226, 270, 229, 2242, 340, 338, 334, 2234, 530, 231, 234, 2613, 227, 1247]. Coefficients [317, 2249, 628, 2157]. Coercive [2196]. Cogenerators [2832]. Cognition [704]. Cognitive [687, 694, 1526, 5, 2089]. Cognitively [688]. Cognitively-Aware [688]. Cograph [2183]. Coherence [2196]. Coherency [417]. Coherent [2222, 1470]. Coil [948]. Colinearity [118]. Collaborated [935]. Collaboration [2422, 482, 2293, 2499, 584]. Collaborative [969, 1409, 2924, 1849, 83, 506, 1254, 459, 2321, 2433, 586, 2335, 93]. Collaborators [2410]. Collecting [41, 428, 2512]. Collection [1598, 1154, 1828, 2537]. Collections [2738]. Collectors [628]. Collide [2057]. Collinearity [2002]. Colonies [90]. Colonography [889]. Colonoscopy [891, 835, 892]. Color [1396, 763, 320, 360, 330, 221, 361, 720, 1642, 151, 976, 533]. Color- [720]. Coloring [652, 662, 671]. Colorings [1386]. Coloured [2030, 1322]. Colourer [2679]. Combination [149, 2812, 790]. Combinatorial [1377, 1951, 1982, 1873, 1954]. Combinators [2528]. Combine [588]. Combined [2443, 2576, 2759]. Combines [383]. Combining [1884, 2895, 2693, 2296, 2102, 872, 130, 748, 2855, 1464, 1490, 251, 224]. Command [2457]. Comments [266, 2126]. Commerce [1735, 2313, 194, 2829, 1623, 1737, 436, 2302, 1721, 541, 1631, 1603, 1730, 2294, 480, 1736]. Commercial [2439, 811]. Commercial-Off-The-Shelf [2439]. Commit [1271]. Commitment [2564]. Commitments [2153]. Commodity [1283, 1959]. Common [2442, 250, 2706]. Commons [644]. Commonsense [2498]. Communal [2343, 605]. Communicable [1493]. Communicating [75]. Communication [2626, 1253, 1748, 1766, 2927, 1335, 623, 203, 1604, 2071, 1742, 2677, 2484, 1250, 2490, 589, 1661, 1704, 2563, 290, 1654]. Communicational [2293]. Communications [1808, 336, 234, 305]. Communicators [2394]. Communities [43, 509, 470, 432, 2390, 1511, 1724, 428]. Community [2338, 425, 2334, 1900]. Commutative [2840]. COMOR [2322]. Compact [1577, 1068, 2023, 600, 1990, 2145]. Compaction [1198]. Compactness [2839]. Companies [23]. Company [21]. Comparability [1370]. Comparative [2878, 1054, 2793, 2305]. Compared [697]. Comparing [802, 1910, 2485, 1378, 1927]. Comparison [2064, 2868, 565, 1064, 2543, 2374, 943, 1683, 937, 291, 53, 324, 1478, 1722]. Comparison-Shopping [1722]. Compatibility [2696, 185]. Compensated [2242]. Compensation [995, 244, 369]. Competency [1454]. Competitive [1349]. Compilation [2869, 1826, 1299, 583, 1288]. Compile [1828]. Compiler [1123, 2069, 2526, 2523, 2052, 1135, 1196]. Compiler-Cooperative [2052]. Compilers [2037, 166]. Compiling [2069, 812, 2527]. Complementary [266, 1509]. Complete [2183, 629, 863, 2575, 2612]. Completeness [1840, 1209, 676]. Completion [1375, 1345]. Complex [633, 1415, 1324, 40, 1103, 347, 1418]. Complexes [1977, 1991, 1983]. Complexities [1473, 1474]. Complexity [2057, 550, 1539, 2066, 1190, 2155, 1376, 1368, 1388, 663, 2071, 31, 629, 1403, 2345, 624, 1390, 2597, 671, 1779, 1210, 1834, 2192, 1378, 1211, 632, 674, 2128, 707, 1578, 2515, 1398, 341]. Compliant [552]. Component [1277, 161, 2447, 1961, 459, 130, 1071, 1777, 310, 1133, 1240, 809, 1144, 1527, 1288, 1128]. Component-Based [2447, 1133, 1527, 1288, 1128]. Components [1120, 2443, 1862, 1958, 2448, 2449, 2452, 2454, 1273, 29, 2458, 2379, 2206]. Composing [186, 1274]. Composite [2108, 189, 209, 2116]. Composition [2222, 199, 2053, 89, 1892, 2401, 2836, 192]. Compositional [2700, 2531]. Comprehensive [1417]. Compressed [358, 276, 1586, 1195, 334, 271, 327, 364]. Compression [377, 2236, 2239, 2242, 757, 2818, 1115, 2243]. Computability [2066, 2050, 2048, 612]. Computation [2208, 1188, 1869, 1204, 1998, 713, 2137, 2932]. Computational [1205, 1843, 496, 1492, 833, 412, 1656, 1463, 1488, 684, 1493, 2071, 2185, 1913, 1967, 2918, 1950, 1685, 754, 1515, 707, 2874, 1526, 713, 1966, 781, 775]. Computations [2281, 2539, 1762]. Compute [1589]. Computed [968, 901, 985]. Computed-Aided [968]. Computer [497, 1057, 547, 2520, 860, 1053, 994, 857, 904, 721, 407, 508, 1056, 889, 1978, 510, 511, 516, 1979, 2214, 855, 154, 817, 143, 576, 590, 949, 215, 951, 1764, 1160, 797]. Computer-Aided [497, 1057, 1056, 510, 511, 951]. Computer-Assisted [857, 855]. Computer-Based [817]. Computers [1785, 1767, 1583, 765, 1938, 2784]. Computing [1563, 966, 2785, 1601, 1404, 2435, 1960, 597, 1192, 382, 2792, 2409, 793, 629, 1970, 214, 2829, 2598, 1759, 2381, 1318, 2430, 154, 2360, 1579, 2798, 582, 734, 1552, 1968, 590, 2784, 588, 584, 2357, 1570, 2147, 606]. Concealment [272]. Concentration [125]. Concept [778, 1721, 474, 2361, 2482, 2354, 2229, 440]. Concepts [686, 1470, 2898, 1471, 1319, 2318, 2005]. Conceptual [1451, 1429, 1409, 2473, 1438, 2849, 454, 2038, 1432, 1416, 1514, 1423, 441, 1420]. Concerning [1221]. Concerns [167, 171, 162]. Conclusion [2655].

Conclusions [2773]. Concurrency

[1202, 1820]. Concurrent [1770, 2155, 2926, 1866, 2025, 2041, 1206, 1929]. Conditional [2927, 423, 61, 2368, 2899, 736, 2383, 632, 2187, 984]. Conditions [715, 2226, 1568, 2224, 426]. Cone [1045]. Cone-Beam [1045]. Conference [438, 588]. Confidence [1593, 2517, 2136]. Confidentiality [554, 1243]. Configurable [1281, 487, 583, 1157]. Configuration [430, 1890, 1891, 2486]. Configurational [251]. Configurations [2624, 1648]. Conflation [2882]. Conflict [2375, 2562, 2371]. Conflicting [2039]. Conflicts [1430]. Confluence [610]. Conform [1956]. Conformal [998]. Conformation [1508]. Congested [238]. Congestion [1746, 2778, 2292, 1180]. Congruence [1221]. Conjecture [620, 2727, 2923]. Conjugate [1853]. Conjunctions [1894]. Connected [2432, 2261, 2206]. Connection [339]. Connections [585, 1167]. Connectivity [1372, 1354, 673]. Connector [2700]. Conquer [1512, 2631]. Cons [150]. Consciousness [1261]. Consensus [1391, 488, 2371, 477]. Consequences [2193]. Conservative [1831]. Considerations [2389, 300, 256]. Considering [2778, 2182]. Consistency [1884, 2574, 2421, 2902, 2705, 1864]. Consistent [820]. Constant [2062, 1771]. Constrained [1401, 652, 655, 625, 2522, 1876, 1868, 2427, 1050]. Constraint [1848, 1883, 1879, 1832, 1906, 2466, 1860, 1825, 1871, 1850, 1885, 2862, 1859, 1770, 2315, 2664, 2681, 1814, 1887, 2574, 1869, 1865, 1849, 1846, 1889, 1890, 1899, 1827, 1891, 1909, 1861, 1892, 1427, 1912, 1826, 1913, 2674, 1914, 2211, 1902, 1867, 1919, 1901, 2926, 2576, 1866, 1921, 1858, 1868, 1922, 828, 1923, 1926, 2046, 1847, 1820, 2045, 1929, 1905, 1900, 1863]. Constraint-Based [1879, 1865, 1890, 1899, 1891, 2576, 1820].

Constraints [1872, 1382, 1790, 2908, 2670, 1334, 2667, 1894, 2671, 1164, 2900, 2199, 1870, 2079, 1881, 1413, 1925, 1025, 2205]. Constructing [620, 944, 822, 2121, 2714, 2072, 2849, 2898, 1453, 1501, 1379, 512, 2887, 2586, 1144, 1200]. Construction [2555, 385, 2625, 449, 2258, 1277, 2016, 2347, 2156, 2268, 2523, 2700, 903, 824, 2426]. Constructions [538, 1573, 1547, 2073, 1364, 2639, 2110, 2613]. Constructs [1435, 2296]. Consumer [1281]. Consumption [772]. Contained [2928]. Container [439]. Containment [1668]. Content [1239, 228, 418, 2435, 450, 451, 330, 321, 331, 289, 604, 328, 1452, 351, 1173, 334, 1740, 279, 458, 429, 284, 1247]. Content-Addressable [1740]. Content-Based [321, 289, 328, 351, 334, 279, 458, 1247]. Context [2572, 165, 2432, 2434, 102, 2898, 2242, 589, 2397, 1692, 2703, 1321, 1273, 937]. Context-Aware [165, 589]. Context-Based [2242]. Context-Free [2572]. Context-Sensitive [2432, 1321]. Contexts [8, 2889]. Contextual [442, 46]. Contingency [3]. Continuing [1265]. Continuity [685]. Continuos [783]. Continuous [2574, 83, 2245, 771, 2497, 919, 1980, 1005, 233]. Contour [1011, 916, 1048, 139]. Contour-Based [139]. Contract [564]. Contraction [957, 366, 1687]. Contracts [187, 1424, 1972]. Contrast [918, 522, 873]. Contrast-Enhanced [522]. Contribute [1764]. Contributions [1110]. Control [1748, 1284, 1222, 2922, 778, 1746, 1124, 547, 1659, 68, 506, 1204, 2775, 2308, 2428, 1020, 1760, 2646, 768, 1744, 2780, 232, 1693, 2457, 2683, 339, 2795, 440, 354, 1079, 352, 560, 1234, 1233, 1749, 1180, 2538, 730, 2806, 1647, 1654]. Control-Flow [1222, 547]. Controllable [885, 614]. Controlled [267, 2895, 946, 2124, 1006, 843, 2426].

Controller

[779, 207, 2774, 2777, 600, 1713, 1088]. Controllers [20, 1262, 2825, 727]. Controlling [2778, 617]. Convergence [1085, 1341, 1994, 2798, 735]. Conversation [247]. Conversational [463, 2507, 591]. Conversations [184, 2734]. Conversion [1605, 1420]. Convex [1872, 1395, 2833, 142, 1344, 1998, 143]. Convexity [1988, 1991]. Conveying [2316]. Conway [2914]. Cooperating [198, 1235, 1922]. Cooperation [2306]. Cooperations [2047]. Cooperative [1069, 2352, 1065, 185, 2052, 405, 138, 477, 1175]. Cooperatively [2406]. Coordinates [126]. Coordinating [617]. Coordination [2558, 1725, 2682, 1273]. Coordinator [465]. Coping [172, 1427, 176]. Copyright [1560, 312]. Coq [1216]. CORA [796]. CORBA [1960, 1279, 1280, 1959]. Core [1336, 1244]. Core-Stateless [1244]. Cores [1879]. Corner [2821]. Coronary [1027, 959, 1072]. Corpora [2547, 39]. Corporate [2447]. Corpus [2869]. Correct [2016]. Correcting [944, 2636, 1537, 1622, 1658]. Correction [923, 850, 294, 41, 1179, 837, 2631, 984, 1029, 848, 1026]. Correctness [1222, 1840]. Correlated [69, 573, 1465]. Correlation [1253, 313, 245, 1158, 2120, 2717, 46, 295, 984, 2157, 1557, 1156, 1636, 329]. Correlation-Based [46]. Correlation-Immune [1636]. Correspondence [2913, 2817]. Cortex [1047]. Cortical [921, 505]. COS [2618]. Coscheduling [1334, 1333, 1337]. Cost [1598, 2186, 2471, 1040, 137, 580, 1923, 1359]. Cost-Distance [1359]. Costs [1600, 1771]. COTS [2439, 2438, 2440, 2441, 2442, 2443, 2446, 2445, 2447, 2449, 2450, 2453, 2454, 2457, 2455, 2451, 2460]. COTS-Based [2438, 2455]. COTS/GOTS [2457]. Count [504]. Counter [1768, 2068]. Counterexamples [2395]. Countermodel [1214]. Counting [2158, 1911, 2190, 2077, 1355, 2905]. Counts [1499]. Couple [2127]. Coupling [1685, 919]. Coupon [628]. coupons [631]. Courses [481]. Cover [2607, 1399]. Covering [2587, 1384]. Coverings [2786, 2368]. Covers [655, 1388]. Cranio [871]. Cranio-maxillofacial [871]. Crash [204]. Crash-Affected [204]. Crawled [472]. Crawlets [1939]. Create [2089, 602]. Creating [868]. Creation [24, 531, 1181, 866]. Creative [431, 2393, 2394, 2392]. Credential [2161]. Crime [792]. Crisp [1418]. Criteria [2098, 493, 96, 2508, 1478, 1735, 2782]. Criterion [118]. Critical [166, 1139, 677]. CRM [1454]. CRML [418]. Cross [199, 500, 449, 2815, 1817, 2765, 187, 269, 528, 2886, 357]. Cross-Enterprise [1817]. Cross-Lingual [449, 2886]. Cross-Organizational [187]. Cross-Over [199]. Cross-Security [500]. Cross-Trees [2765]. Cross-Validation [528]. Crosscuts [168]. Crosscutting [172, 167]. Crossing [707]. Crossover [2689]. Crowding [2827]. Cryptanalysis [2115, 1617, 2131, 2583, 2647, 556, 2142, 2644, 2117, 1658]. CRYPTIM [1562]. Cryptographic [2616, 2098, 2129, 1931, 2101]. Cryptography [2107, 2642, 2108, 2159, 2595, 2591, 1657, 2628, 2608, 2127]. Cryptologic [1615]. Cryptology [2605]. Cryptosystem [2637, 2603, 1650, 2117]. Cryptosystems [2151, 2709, 2743, 2150, 2741, 2631, 555]. crystalline [755]. C’s [1271]. CSP [1915, 1920, 1924]. CSPAMM [987]. CSPs [1857, 1917, 1904]. CT [982, 932, 827, 1056, 1054, 863, 900, 1045, 844, 522, 878]. Cube [1992, 151]. Cubic [1048, 2104, 2550]. Cues [595]. Cultural [705, 2436, 404, 2330, 2489]. Culture [26, 408]. Cumulative [2344]. Current [2442, 837, 2454, 182]. Curvature [921, 1989]. Curve [2601, 553, 2108, 1572, 2760, 559, 555, 1640]. Curves [148, 966, 2107, 2764, 1992, 1993, 2593, 2159, 2158, 2312, 2592, 2106, 1995, 2763, 2762, 1552, 1621]. Customer [262, 483]. Customization [806, 1287]. Customized [2459]. Cut [1954, 1851, 375, 1955, 1950, 674, 2554]. Cuts [1956, 940, 1371, 832]. Cutsets [672]. Cutting [1001, 865, 797]. Cyber [2086, 254, 1665]. Cyber-Attack [1665]. Cyber-Societies [2086]. Cycle [2065, 1353, 1880, 2274]. Cycles [1565]. Cyclic [1620, 1566, 1541, 2840, 1540, 2621].

D [1067, 1043, 972, 878, 953, 1005, 934, 110, 884, 925, 149, 1067, 901, 912, 940, 1992, 1002, 1064, 1993, 932, 1040, 971, 1682, 367, 986, 876, 134, 1046, 1379, 111, 1043, 826, 861, 269, 943, 1017, 1681, 1027, 2659, 386, 1996, 941, 94, 937, 527, 824, 2427, 878, 910, 954, 147, 105, 1997, 1029, 1999, 1306, 2000, 848, 1672, 1005, 243, 416, 990, 370, 1009, 145, 907, 871, 2426, 109]. D-to- [1005]. D/ [1067, 1043, 878]. D/Statistics [1067]. D0 [1974]. Daemon [1179]. Damage [574]. Data [57, 1734, 1075, 1592, 2018, 1154, 495, 1482, 928, 1428, 1172, 940, 41, 2066, 1429, 264, 1595, 1064, 70, 792, 927, 777, 1807, 472, 840, 36, 502, 49, 2542, 2218, 2384, 1083, 2039, 847, 175, 1934, 209, 2385, 1524, 1431, 728, 2403, 939, 1453, 1162, 507, 719, 2396, 54, 1978, 513, 514, 2038, 38, 972, 2376, 1974, 2239, 1681, 2813, 1421, 1611, 650, 518, 2321, 1684, 58, 488, 2048, 1367, 930, 521, 2360, 1642, 1705, 1771, 523, 1965, 1419, 2880, 527, 69, 263, 1464, 1490, 2551, 141, 2367, 117, 953, 2841, 1518, 1433, 1690, 774, 868]. Data [437, 2538, 275, 2511, 2291, 2512, 1647, 2406, 898, 2615, 435]. Data-Centric [1647]. Data-Mining [728]. Database [2894, 1425, 1659, 259, 2906, 1625, 1793, 351, 1080, 2880, 560, 1077, 2512]. Databases [1445, 1843, 2896, 1074, 498, 250, 1084, 2908, 1844, 2211, 35, 2902, 1430, 2041, 307, 1821, 2904, 2905, 429, 441, 437, 2514]. Dataset [2401]. Datasets [33]. Davis [2204]. DCT [1619, 313, 317]. DDoS [804]. De-blocking [2238]. Dead [1815]. Deadlines [1850]. Deal [2040]. Dealing [1626, 2376]. Dealmaker [1719]. Debugger [1944]. Debugging [2536]. Decentralised [2562]. Decentralized [1725, 2679, 1963, 2308, 1286]. Decidability [2476, 707, 2060]. Decimation [230]. Decimators [2272]. Decision [2544, 1566, 2471, 2782, 2543, 715, 2788, 2008, 1466, 690, 1363, 718, 2087, 572, 2454, 2298, 503, 524, 531, 749, 60]. Decision-Making [2454, 2298]. Decisional [2727]. Declarations [2042, 2657, 1292]. Declarative [2536, 1840, 2420, 2519, 1153, 2524]. Declarative/Procedural [2420]. Declustering [450]. Decodable [2621]. Decoder [2614]. Decoding [2610, 1566, 1567, 2073, 1568, 1569, 2647, 2624, 1551, 2141]. Decomposable [1767, 1367]. Decomposed [2803]. Decomposition [2311, 1846, 669, 914, 2903, 2763, 2280]. Deconvolution [2262]. Decreasing [626]. Decremental [627]. Decryption [2620]. Deduction [1507]. Deductive [1843]. Deep [955, 1498]. Default [2463, 1623, 2772]. Defeasible [2463]. Defeating [1646]. Defect [98]. Defence [1099]. Defense [804, 2460]. Defibrillation [931]. Defined [1620, 686, 1313, 1312, 1314]. Defining [57, 2871]. Definition [2691, 168, 2054, 1015, 2453]. Definitions [2200, 2088]. Definitory [2889]. Deflection [929]. Defocused [2257]. Deformable [110, 884, 895, 938, 1039, 823, 1046, 881, 917, 913]. Deformation [1033, 944, 1688, 971, 879, 1006, 1677, 832, 828, 824, 820]. Deformations [923, 1003, 883, 1004]. Defuzzification [741]. Degradation [1619]. Degraded [2820]. Degree [654, 2771, 2076, 2116, 329]. Degrees [1358]. Delay [769, 2042, 1348, 1391]. Delegation [1641, 2474]. Deliberation [2186]. Delineation [961]. Delivering [1239, 2435]. Delivery [261, 1940, 568, 1736]. Delsarte [2627]. Delta [2912]. Demand [349, 1171, 345]. Dementia [1053, 820]. Demo [2739, 787]. Democracy [2340]. Demonstration [1099]. Demosaic [361]. Dempster [2774, 1684]. Dempster-Shafer [2774]. Denial [1649, 1161, 1646]. Denial-of-Service [1646]. Dense [581]. Density [2438, 2636]. Dental [1018, 860, 1000]. Department [2460]. Departments [11]. Dependence [2183]. Dependencies [2219, 2043, 2897, 2366]. Dependency [2882]. Dependency-Pairs [2882]. Depictions [697]. Deploying [808]. Deployment [583]. Depolarization [1670]. Depth [2062, 211, 2257]. Depth-First [211]. Deputy [1430]. Derandomization [2057]. Derivation [1516]. Derivative [2833, 1623]. Derived [112, 953]. Deriving [1889]. Descent [2471, 2116]. Describing [87, 25]. Description [2194, 228, 1888, 1414, 794, 1437, 1452, 1318, 429]. Description-Based [1414]. Descriptions [618, 253, 221, 2195, 697]. Descriptive [27, 1530]. Descriptor [307, 271]. Descriptors [324]. Design [2438, 1441, 1118, 1981, 1377, 1222, 1140, 1732, 2098, 349, 2067, 1816, 2629, 1425, 1324, 691, 2618, 2341, 1958, 2434, 207, 1179, 255, 2386, 2396, 2322, 2737, 1141, 598, 572, 2753, 1133, 1442, 431, 2269, 2436, 574, 2413, 1319, 2332, 945, 359, 1697, 1700, 1109, 1416, 2411, 2275, 1741, 605, 1418, 1090, 8, 1536, 1147, 1763, 2502, 1100, 1433, 592, 2890, 638, 1959, 380, 300, 2270, 2096, 2319, 1082, 1128]. Designed [159]. Designing [1120, 2802, 2420, 1157]. Designs [2627, 2419]. Desk [2850]. destructive [751]. Detect [989]. Detecting [493, 1159, 2417, 2709, 152, 509, 820, 1160, 357].

Detection [891, 497, 1906, 1011, 2065, 41, 228, 317, 242, 1158, 1673, 2311, 743, 114, 204, 825, 102, 101, 1162, 2289, 111, 2821, 511, 2261, 239, 2310, 1660, 890, 517, 118, 123, 1709, 964, 98, 2631, 295, 773, 318, 141, 947, 265, 988, 146, 1163, 1157, 533, 327, 877, 1155, 2233, 284, 222, 322]. Detection/Correction [2631]. Detector [313]. Determination [1149, 2254, 915]. Determining [926]. Deterministic [2067, 619, 210, 1952, 1199, 1200, 2534, 2048, 2366]. Developed [1563]. Developing [2445, 1440, 1258, 589, 1436]. Development [20, 1804, 2424, 2033, 2425, 394, 453, 28, 18, 2416, 255, 1020, 2484, 516, 3, 2413, 2005, 2706, 2509, 996, 1337, 888]. Device [854, 602]. Devices [1126, 595, 1281, 2634, 1318, 2429, 584, 583, 2147]. DFT [1540]. DGML [443]. Diacritics [2877]. Diagnosis [2276, 496, 497, 1057, 491, 2790, 1053, 721, 1056, 718, 889, 510, 511, 512, 1350, 2405]. Diagnostic [993, 522, 1022]. Diagonal [309, 1667]. Diagram [1787]. Diagrams [2668, 2008, 1363, 1207, 1436]. Dialectical [2388]. Dialogs [2337, 2500]. Dialogue [2507]. dian [1405]. DICOM [522]. Dictionaries [563, 2868]. Dictionary [2871]. Didactic [348]. Difference [1026]. Differences [820]. Different [1437, 2793, 1995, 911]. Differentiability [1362]. Differential [2833, 2245, 1882, 2596, 1536, 2142, 2615]. Differentiated [1245]. Differentiation [2685]. Differently [695]. Difficulty [747]. Diffie [2148, 2727, 2129]. Diffserv [1242]. Diffusion [838, 837, 836, 105]. Digestion [714]. Digit [2511, 787]. Digital [220, 1785, 1635, 561, 276, 2140, 262, 1976, 1988, 2745, 596, 1719, 252, 1638, 2288, 2290, 2484, 511, 361, 2628, 2490, 336, 854, 2112, 1623, 1902, 489, 1661, 309, 2122, 849, 265, 588, 1088, 1055, 2002, 1991, 1664, 312, 275, 1622, 1640, 1985]. Digitization [1996]. Digitized [497]. Digraphs [1193]. Dijkstra [202]. Dilation [2251]. Dilemmas [2087]. Dimension [1441, 249]. Dimensional [1676, 443, 827, 1056, 54, 2304, 985, 836, 959, 2570, 755, 2464, 842, 223, 2568, 1014]. Dirac [2247]. Direct [2911, 840]. Directed [627, 1577, 201, 784, 378, 2076, 2195, 2787]. Directed-Divergence [2787]. Directing [2300]. Direction [692, 775]. directional [1667, 235, 1749]. Directions [2390, 698, 1143, 1130]. Disagreeing [488]. Disambiguation [2853, 2855, 2852]. Disappearing [590]. Disc [533]. Discarding [2238]. Disconnection [2189]. Discourse [2339]. Discourses [2340]. Discovered [472]. Discovering [61, 1463, 1488, 432, 1511, 1521]. Discovery [1460, 1487, 1492, 1962, 1493, 2401, 2395, 1498, 1963, 1504, 1505, 67, 1506, 470, 2388, 2404, 2398, 475, 2397, 2387, 2405, 1513, 2389, 69, 1464, 1490, 2394, 1517, 1528, 1519, 2399, 785, 1764, 1532, 2513, 2514, 1530]. Discrepancy [1396, 1553]. Discrete [1986, 186, 1993, 1989, 1984, 2598, 2728, 2255, 2275, 1613, 1597, 2137, 1980]. Discrimination [717]. Discriminative [128]. DisCSPs [1928]. Disease [968, 519, 1016, 1022, 894]. Disjunctive [2218, 1870]. Dispatching [2779, 1655]. Displacement [987]. Display [2428, 861, 899]. Displays [604, 603, 602]. DISSECT [546]. Dissection [1036]. Distance [148, 1451, 1404, 1539, 311, 656, 625, 249, 2316, 420, 975, 601, 1359, 1999, 919, 2000, 2585]. Distances [2816]. DISTIL [1416]. Distillation [770]. Distinct [1860]. Distinctions [711]. Distinguishing [2643]. Distortion [837]. Distortions [566]. Distraction [2409]. Distraction-Free [2409]. Distributed [430, 2222, 1284, 1918, 538, 2065, 1769, 2337, 1755, 418, 1171, 540, 1084, 1270, 2015, 1105, 2149, 1287, 1204, 1913, 719, 1132, 1392, 2562, 2095, 1919, 1901, 1241, 488, 2080, 1771, 1272, 1922, 1762, 1142, 192, 440, 1147, 2119, 1161, 1307, 212, 638, 1183, 1973]. Distribution [1396, 538, 630, 2689, 2364, 1173, 2678, 2464, 546]. Distributions [1331, 2361, 2274]. Distributive [794]. Distrust [2088]. Dithering [231]. Divergence [2787]. Diverse [593, 1513]. Divide [1512, 2631]. dividing [2128]. Division [1576]. DJ [160]. DLS [755]. DLS-Spectra [755]. DNA [1379, 2930, 675, 2931, 1096, 2283]. Do [1956, 2711, 2643]. DocToon\begin{matrix}c \\ ◯ \end{matrix} [403]. Document [500, 2873, 1774, 1627, 2482, 1502, 2391, 1510, 309, 2551, 2886, 2259, 291, 2883]. Documents [2888, 561, 2733, 1440, 1449, 2820, 474, 1499, 1509, 475, 1667, 1166, 2253, 587, 2040, 2885, 2884]. Does [2089, 2430]. DOF [887]. Doing [2642]. Doll [1877]. Domain [2555, 1884, 1619, 313, 276, 2235, 1498, 469, 2482, 314, 2237, 2112, 711, 334, 549, 271, 2259, 1426, 2221, 364]. Domain-Specific [2482, 1426]. Domains [2716, 1242, 1915, 2879, 536]. Dominating [667]. Done [2505]. Doppler [1677, 2277]. Double [1546, 240, 707]. Double-Crossing [707]. Doubt [2087]. Downs [1343]. DPA [556]. DPA-Based [556]. DPDA [2060]. DPO [2919]. Dramatised [86]. Drawn [2327]. Driven [1451, 1168, 1428, 365, 2924, 2738, 1031, 2850, 893, 1096, 2011, 2529]. Drivers [499]. Driving [782, 2552]. Drowsy [499]. DRT [2847]. Drug [1530]. DS [773]. DS/CDMA [773]. DSMs [1600]. DT [1061]. DTD [1420]. Dual [1895, 877, 1538, 1539, 1535]. Duality [1543]. Dump [2620]. During [983, 1028, 30, 1430]. Dynamic [1445, 2010, 1906, 1918, 2558, 940, 1936, 2148, 1732, 856, 1255, 2472, 938, 1246, 1325, 2432, 1833, 1390, 257, 1242, 424, 1101, 1452, 1964, 858, 2568, 163, 1028, 1901, 2578, 160, 1921, 1965, 1799, 849, 303, 2305, 1528, 811, 374, 877, 747, 227]. Dynamical [2026]. Dynamically [473, 381]. Dynamics [833, 2089, 1257, 38, 958, 1028]. dynP [1580]. e-AGORA [408]. E-business [1409, 1719, 23]. E-Commerce [2829, 2302, 541, 1603, 1730, 2294, 480, 1736]. E-coupons [631]. E-Mail [2720, 2111]. E-Partner [247]. E-Service [187, 192, 1723]. e-Services [185, 186, 188, 182, 1720, 1718]. E-unifiability [634]. E2E [1455]. E3 [24]. Each [2008]. EAGLES [2867]. EAGLES/ISLE [2867]. eaLib [726]. Easier [581]. Easily [585]. Easy [609, 200, 1977]. EasyVision [1000]. Eavesdropping [1205]. EBPDT [2815]. ECC [1620]. ECDLP [2116]. ECG [1686, 1690, 534]. Echo [986]. Echo-Cardiographic [986]. Echocardiography [907]. Echoes [990]. Echographic [1673]. echography [853]. ECIES [2617]. ECLiPSe [2535]. ECOC [2553]. Ecological [38]. Ecology [395]. Economic [1725, 2356, 2348, 1783, 2346]. Economics [2354, 1344]. Economy [2355, 2356, 1411]. ECT [980]. Eddy [837]. Eddy-Current [837]. Edge [651, 652, 1372, 838, 872, 2252, 720, 2257, 145, 2247]. Edge-Based [720]. Edge-Connectivity [1372]. Edge-Isoperimetric [651]. Edges [1565, 943, 677]. Editor [1459, 2054]. Education [482, 73, 4, 298]. Educational [404, 4, 392]. Effect [850, 325, 1102, 1008]. Effective [621, 548, 463, 272]. Effects [355, 2254, 980, 526, 1338, 2350]. Efficiency [1747, 2185]. Efficient [1482, 2208, 1884, 2470, 2139, 1617, 1371, 565, 2125, 1277, 2154, 540, 822, 1270, 1940, 564, 259, 207, 2745, 1586, 1431, 2075, 333, 569, 257, 111, 2634, 1407, 1645, 2113, 237, 2881, 2077, 269, 2023, 1402, 2212, 1633, 2123, 193, 1579, 1750, 2122, 1514, 376, 2141, 1516, 583, 2751, 2712, 2147, 555, 2538, 1333, 2513, 1337, 1386]. Efficiently [2711]. Effort [30]. EHBT [1750]. EHW [1098]. Eigenspace [2793]. Eigenspace-Based [2793]. Elaboration [2327]. Elastic [879, 1042]. Elasticity [2248]. Elderly [785]. Election [2074]. Elections [2749, 2750]. Electrical [1680, 1690]. Electro [1686, 841]. Electro-Mechanical [1686, 841]. Electroencephalogram [716]. Electromechanical [1685]. Electronic [1735, 448, 1595, 483, 1424, 479, 1663, 1623, 1737, 1723, 2564, 568, 1097, 436, 591, 408, 193]. Electronics [1113]. Electrophysiological [2795]. Electroshock [526]. Element [1031, 830, 1001, 832, 865, 756, 1672]. Elementary [1483, 1516]. Elements [1039, 1291, 278]. Elicitation [27]. Eligibility [2780]. Eliminating [2545, 1499]. Elimination [1893, 2508]. Elliptic [2601, 553, 2107, 2764, 2108, 2593, 2760, 2592, 2106, 559, 2763, 2762, 2602, 555, 1640]. Elog [2217]. Elucidating [59]. Embedded [1118, 1119, 2466, 1120, 1121, 1148, 1122, 1140, 1124, 1125, 1126, 1127, 1898, 1129, 1131, 2241, 1141, 1134, 1838, 1136, 1137, 2527, 1139, 1142, 1144, 1145, 1146, 1079, 1150, 1151]. Embodying [2493]. Embryonics [1095, 1096, 1094]. Emergence [258, 2343]. Emergency [494]. Emergent [2474]. Emerging [422, 470]. Emission [985]. Emotion [2785, 2469, 75, 284]. Emotion/Intention [2785]. Empirical [2037, 1525, 2887, 2234, 53]. Employ [1769]. Employing [2794, 2796]. Empower [2291]. Emulator [1938]. Enabled [507]. Enabling [2560, 21, 1730, 2339, 480]. Encapsulated [1969]. Enciphering [2100]. Encoded [1192]. Encoding [1194, 1857, 1985]. Encodings [2930, 2571, 2197, 1927]. Encrypted [2638, 568]. Encryption [2748, 2726, 2720, 2163, 264, 2641, 2742, 2584, 2146, 2744, 2761, 1629, 1562, 2725]. End [261, 1242, 1137, 2554, 1183, 5]. End-Cut [2554]. End-To-End [1137, 261, 1242]. End-User [1183]. Endings [2874]. Endodontics [982].

Endoscope [1043, 815]. Endoscopes [1023]. Endoscopy [1063, 815]. Energy [1282, 772]. Enforced [2416]. Enforcement [199, 2902]. Enforcing [1632]. Engine [20, 1178, 453, 40, 351, 1824, 1971, 1905]. Engineering [2, 13, 639, 2315, 637, 1385, 482, 2522, 19, 1610, 414, 1267, 642, 2455, 2411, 2035, 2040, 469]. Engines [2539, 1939]. English [2858]. Enhance [431]. Enhanced [325, 1255, 2901, 522]. Enhancement [30, 2808, 906]. Enhancing [1238, 387, 105]. Enrich [284]. Ensemble [412, 46, 1966, 53]. Entailment [1219]. Entanglement [2639]. Enterprise [183, 1817, 572, 1086, 415, 2563, 1447, 1647, 1737, 1271]. Entities [11]. Entity [1271]. Entropy [1192, 922, 984, 2854, 846]. Entropy-Based [2854]. Entry [2867]. Enumerating [1352, 1374]. Enumerator [2612, 2623]. Envelope [2072, 1028]. Environment [496, 810, 1186, 2873, 1448, 806, 389, 1961, 1334, 1941, 2395, 1102, 187, 783, 2214, 185, 2334, 406, 1416, 2563, 586, 1276, 440, 1303, 2505, 243, 587, 1183, 2419, 298, 606]. Environmental [2348, 282]. Environments [1936, 578, 1255, 75, 202, 704, 83, 1970, 1431, 1963, 82, 582, 1272, 999, 584]. EPIC [1135]. Epicardial [1680]. Epidemiology [509]. Epileptic [983]. Epilog [2044]. Epipolar [815]. Episode [1523]. Episodes [52, 59, 525]. Epistemic [2487]. Equation [2536, 1246, 112, 1836, 1211, 1519]. Equation-Based [1246]. Equational [2695]. Equations [2755, 2245, 1882]. Equilibrium [2571, 2795]. Equipment [2782]. Equipping [78]. Equivalence [2365, 1208, 1366, 2083, 1782, 2060]. Equivalences [2920, 1220]. Equivalent [2018, 2017, 2190, 1636]. Erasures [2073, 1569]. Erlang [2526]. Erroneous [317]. Error [355, 294, 1179, 2611, 297, 338, 2636, 2631, 1467, 1537, 272, 1622, 227, 1658]. Error-Correcting [1537, 1658]. Error-Prone [297]. Errors [41, 2417, 2897, 1569, 2646]. Errors-and-Erasures [1569]. Ershov [2005, 2004]. Escape [1104]. Establish [585]. ESTIA [193]. Estimates [121, 960]. Estimating [659, 1030, 1644]. Estimation [1593, 1686, 30, 245, 1993, 136, 987, 1673, 108, 2255, 2678, 837, 1623, 1677, 751, 1995, 911, 151, 138, 913, 2266, 370, 230]. Estimator [2249]. Euclidean [150, 1357, 1558]. Eurocrypt [2131]. Europe [385]. European [2761, 408, 2442]. EUROVOC [2886]. EuroWordNet [2868]. Evaluate [2341]. Evaluating [2090, 2415, 689, 1933, 12]. Evaluation [1078, 2422, 1069, 2895, 1619, 2558, 2337, 2539, 2152, 2446, 1331, 378, 1040, 1947, 1898, 1604, 31, 453, 885, 845, 24, 1603, 908, 2883, 2299, 985, 2429, 836, 887, 945, 177, 899, 143, 2234, 44, 1971, 1328, 976, 274, 1034, 835, 370, 727, 5, 2796, 1082, 1322, 2011]. Evaluators [379]. Even [1554]. EVENT [2316, 1253, 228, 434, 239, 1273, 1741, 1597, 2923, 2024, 2011, 2529]. Event-Driven [2011, 2529]. Events [217, 235]. Eventually [1218]. Ever [396]. Everyday [581, 594, 2430, 2784]. Everywhere [602]. Evidence [702]. Evolution [2439, 6, 1074, 1262, 1265, 1102, 1103, 1114, 1759, 2550, 1089, 1111, 323, 1088, 93, 2352]. Evolutionary [2462, 2863, 2686, 750, 1404, 2326, 2313, 867, 28, 1107, 1108, 444, 2682, 2825, 1109, 2683, 726, 1090, 713, 749, 2805, 774, 90]. Evolutions [827]. Evolvability [2444]. Evolvable [1091, 1116, 1115]. Evolve [2828]. Evolved [1263]. Evolving [2559, 2207, 1585, 2728, 1089, 989, 1433, 1761]. eWorkshop [2, 13]. Exact [1380, 2617]. Example [279, 90]. Examples [144, 2658]. Exceptions [1300, 180]. Exchange [2148, 1729, 1644, 1728, 2147, 1664]. Exchanging [428]. Excision [955]. Excitation [1675]. Exclusion [207]. Executable [2420]. Execution [183, 189, 1431, 2173, 2578]. Execution-Monitoring [2578]. Exhaustive [307]. Exhibition [215]. existence [1853, 1555]. Existing [1805]. Expansion [292, 117]. Expectations [705]. Expected [791]. Experience [385, 2, 13, 17, 19, 218, 14, 8]. Experiences [14, 480, 428, 1163]. Experiencing [485]. Experiential [2092]. Experiment [2349, 2260, 1974, 163]. Experimental [552, 2679, 1685, 2519, 727]. Experimentation [2795]. Experimenting [1076]. Experiments [159, 2873, 1114, 1274, 2092]. Expert [495, 424, 512, 513, 532]. Experts [36]. Explanations [1889, 1500]. Explanatory [2871]. Explicit [1227, 2495]. Exploit [1824]. Exploiting [1920, 984, 2538]. Exploration [1524]. Exploratory [486]. Exploring [1011, 594, 2418, 2051]. EXPO [215]. Exponentially [626]. Exponentiation [2712, 2594]. Exposed [1969]. Expressing [2668]. Expression [745, 134, 2289, 54, 518, 2809, 1196]. Expressions [2469, 1837, 2312, 2571, 2843, 1293, 1518, 1521, 1365]. Expressive [1886, 2009]. Expressively [2183]. Extendable [1032]. Extended [2200, 2907, 203, 2848, 1279, 2308, 1504, 2604, 2404, 2563, 1359, 706, 2367, 1365, 1465, 1398, 1400]. Extendible [1196]. Extending [26, 1483, 2101, 643]. Extension [2018, 1620, 2010, 2107, 2033, 2670, 2158, 157]. Extensions [2598]. External [1003, 2535]. Extra [2635]. Extract [57, 1367, 2889]. Extracting [110, 884, 737, 2733, 2402, 2735, 1510, 2377]. Extraction [2217, 2309, 2878, 2828, 2256, 92, 454, 2825, 732, 1109, 2879, 2880, 63, 2551, 271]. Extractor [1531]. Extrapolation [1757]. Extremal [2904]. Extreme [23]. Extremum [1743]. Eye [267, 74]. Eye-Movement [267]. Eyeball [870].

Eyeglasses [368].

F [1308]. FA [1185]. Fabrics [1607]. Face [220, 79, 365, 2311, 244, 120, 249, 248, 2255, 2793, 816, 363, 246]. Face-Unlock [363]. Faces [2289, 104, 2605, 94]. Facial [2469, 2312, 134, 306, 357]. Facilitate [792]. Facility [2664]. Factor [1577, 1682, 2251]. Factoring [2715, 2123]. Factorization [1798]. Factorizations [661]. Factors [9]. Fading [302]. Fages [1837]. Fail [2709]. Failure [641, 2396, 2210, 1280, 2229, 1709, 1710]. FaintPop [2333]. Fair [2164, 2320, 1663, 1728, 1664, 286, 1244]. Fairness [199, 1348, 1175]. Fall [2326]. False [1600, 573]. False-Positive [573]. Falsification [2081]. Families [2607, 372, 659, 1792, 1554]. Family [20, 1860, 1825, 2108, 2664, 589, 1621]. Family-Oriented [20]. Fanky [94]. Farms [1590]. Fast [2601, 1879, 2108, 2593, 806, 2311, 743, 473, 1937, 2760, 2600, 2104, 1777, 2488, 2238, 248, 229, 211, 2599, 1551, 735, 307, 50, 1042, 1374, 1044, 327, 272, 447, 230, 1621, 2206, 1897]. Faster [2763]. Fatigue [499]. FAucS [1731]. Fault [1581, 1755, 2307, 1250, 1089, 1350, 29, 66, 1251]. Fault-Prone [29]. Fault-Tolerant [1581, 1755, 2307]. Faults [2276, 2502]. FBCC [308]. FCC [1731]. FCM [729]. Feasibility [1019, 2393]. Feasible [1362]. Feature [2555, 33, 880, 2802, 2303, 311, 1756, 2235, 238, 1777, 2252, 2260, 2255, 2903, 2791, 519, 46, 2809, 153, 2854, 370, 278, 322]. Feature-Based [238]. Features [902, 250, 102, 1994, 711, 98, 1311, 1055, 277, 246, 357]. Featuring [863]. Federations [1659]. Feedback [2125, 452, 2299, 2430, 1252, 303, 1743, 278, 34]. FeedForward [2115]. Feistel [2144]. Female [514]. Femur [817]. FGS [350]. Fiber [1680]. Fibre [1674]. Fibrosis [508]. Fickle [621]. Fiducials [900]. Field [987, 2086, 994, 992, 2075, 1093, 1677, 1633, 911, 984]. Field-of-View [992]. Fields [891, 1620, 2107, 1059, 2727, 2108, 971, 709, 62, 1547, 972, 2116]. FIFO [1575]. Fifteen [1786]. Fighting [406]. Figure [917]. Figures [2661, 1899, 1194, 99]. Filament [105, 906]. File [1937, 1229, 2459]. Files [237]. Filesystems [1596]. Film [394, 282]. Films [386]. Filter [609, 149, 2258, 741, 361, 2647, 2271, 2272, 2273, 2262]. Filtered [983]. Filtering [1884, 319, 112, 506, 2525, 459, 890, 2321, 2791, 466, 1858, 1876, 105, 64, 875, 146, 467, 906, 2267]. Filters [2268, 2270]. Filtration [2382]. Financial [2555]. Find [1523]. Finding [474, 2928, 45, 58]. Fine [1592, 1204, 1065, 1247]. Fine-Grain [1592]. Fine-Grained [1204]. Fingerprinting [262, 567, 2100]. Finite [1185, 620, 2219, 1906, 2466, 2716, 1768, 1188, 1190, 2258, 2727, 2108, 1192, 1039, 1773, 2534, 1031, 1548, 1194, 830, 1195, 1366, 1792, 1607, 2116, 615, 832, 2627, 1633, 1197, 1111, 1782, 865, 2923, 1199, 1200, 2687, 1478, 1196]. Finite-Automaton-Encoded [1192]. Finite-State [1195, 1782, 1196]. Finite-Tree [2219]. Fire [406]. Firewall [1648]. Firewalls [2034, 2708]. First [2200, 2063, 2, 13, 2566, 2180, 1498, 2202, 211, 1212, 2362, 1263, 898]. First-Order [2200, 2063, 2180, 1498]. First-Pass [898]. Fit [58]. Fitness [1102]. Fitting [148, 150]. Five [2447]. Fixed [2211, 1836, 2756, 1834, 1209, 1980]. Fixed-Parameter [1834]. Fixed-Pattern [2756]. Fixed-Point [2211, 1209]. Fixtures [1066, 952]. FLASH [556]. Flashover [2790]. FLC [2778]. Flexibility [7, 1446]. Flexible [2686, 194, 156, 1596, 352, 560]. Flip [1880]. Floating [1881]. Floating-Point [1881]. Floorplans [1352]. Flow [891, 1222, 1859, 547, 1048, 1651, 1989, 1275, 1645, 1344, 849, 2025, 1035]. Flows [1245, 1606]. Fluctuation [2292]. Fluent [59]. Fluid [833, 1016]. Fluorescence [909, 105]. Fluorescent [523]. Fluoroscopic [949, 973]. Fluoroscopy [997, 1672]. Flux [121]. Fly [2123]. Flying [1938]. fMRI [922]. Fold [2815]. Folding [921]. Forbidden [671]. Force [929]. Forceps [887]. Forearm [888]. Forecasting [2486, 2804, 60]. Foreground [2309]. Forgery [1618, 2756, 265]. Form [2601, 2470, 1006, 828, 404, 883]. Formal [1205, 2746, 1729, 1873, 427, 168, 709, 2864, 1934, 2482, 1791, 1483, 2656, 167, 475, 2019, 1763, 1516]. Formalisms [2029]. Formalizing [1148]. Formation [1887, 2229, 477]. Formed [2862, 2722]. Forms [2839, 1792, 2531]. Formula [2887]. Formulae [2185, 2204]. Formulas [1112, 1911]. Formulation [1503]. Formulations [1952]. Forward [1179]. Forwarding [1227]. Foundation [2017, 2356, 414]. Foundations [169, 2894, 1057, 2490, 72]. Founded [2208, 1835]. Fourier [504, 918, 324]. FPGA [552]. FPGAs [1105]. Fractal [803, 249, 25, 2243, 2267]. Fractals [2362]. Fractional [2266]. Fractured [817]. Fractures [946]. Fragile [317, 315, 1667, 343]. Fragment [1842]. Fragmentary [1403]. Fragments [2180, 817]. Frame [2256, 268, 1424, 2653, 1231]. Frame-Rate [268]. Framelet [1706, 1713, 1709, 1710, 1704, 1708, 1702, 1703, 1707, 1705, 1701, 1711, 1712]. Framelets [1695]. Frameproof [2109]. Frames [242, 692]. Framework [76, 1253, 1748, 2896, 1451, 42, 2739, 2425, 418, 36, 1725, 1851, 31, 194, 1456, 1279, 2356, 1012, 1032, 54, 1504, 188, 1569, 1250, 958, 1902, 348, 2318, 1697, 1714, 1696, 1699, 582, 2564, 726, 2706, 883, 1971, 1527, 1847, 643, 839, 1288, 868, 2096, 1400]. Frameworks [2561, 1910, 1694, 1136]. Free [2893, 2213, 2572, 2068, 2409, 1348, 670, 2260, 1006, 828, 883, 235]. Free-Form [828, 883]. Freehand [848, 907]. Frequency [2887, 2274, 549, 2243, 1044]. Frequent [1494, 521]. Friendly [1233, 1180]. Friends [585, 1181]. Frog [1996]. Front [915]. Frontends [1442]. Frontiers [1368]. Fronto [820]. FRP [2529]. Full [617, 1029, 990]. Fully [1171, 2149, 1963, 1134, 2857]. Function [501, 938, 2714, 2102, 2786, 2708, 2129, 976, 1690, 1557]. Functional [2438, 1843, 1067, 1002, 927, 2897, 1495, 48, 2533, 2532, 505, 2210, 1681, 2652, 790, 1130, 919, 380, 2626, 159]. Functionality [31, 1170]. Functions [966, 2219, 737, 1362, 2098, 1559, 1626, 2833, 2588, 1040, 2121, 665, 1829, 2898, 2587, 1480, 1473, 2129, 2202, 37, 505, 2120, 2550, 2586, 2082, 1468, 2101, 2370, 2157, 2011, 1636, 1465, 2585]. Fundus [533]. Furniture [2682]. Further [1854, 2613, 2513]. Fusion [968, 127, 367, 375, 464, 1681, 1684, 903]. Future [1121, 1255, 1410, 1169, 2390, 2176, 256, 974]. Fuzzy [2838, 2831, 725, 778, 763, 2840, 2839, 501, 737, 2469, 789, 802, 360, 449, 2315, 2837, 740, 2833, 777, 2790, 2782, 724, 2812, 2769, 800, 799, 742, 715, 2770, 739, 2820, 721, 2776, 459, 753, 752, 2898, 2771, 2775, 2822, 719, 508, 786, 2774, 729, 2364, 2866, 2817, 35, 460, 720, 2808, 2835, 2836, 1795, 2363, 2824, 2781, 2777, 2791, 2780, 751, 2549, 2825, 714, 2818, 776, 765, 2772, 754, 1418, 722, 801, 2832, 2834, 790, 731, 2841, 764, 2789, 2823, 2399, 2810, 2814, 2805, 467, 981, 727, 730, 797, 723, 2802, 2794, 2803, 2804, 2828]. Fuzzy- [725]. Fuzzy-Based [778]. Fuzzy-Neural [2790]. Fuzzy-Neuro [460]. Fuzzy-Similarity-Based [2823].

G [1309]. GA [1112, 772, 2828, 1110, 735]. GA-Based [1112, 772, 735]. Gabor [2264]. GAC [1894]. Gain [1602, 2777]. Gallager [2623]. Galois [1567, 1557]. GAMA [1724]. GAMA-Mall [1724]. Game [2349, 1389, 1729, 1739, 2351].

Game-Theoretic [1389]. Games [2184, 665, 391, 1778, 2353, 392]. Gaming [2355]. Gamma [1786, 867]. Gang [1339]. Gang-Scheduling [1339]. Ganglia [903]. Gap [1756, 2035]. Garbage [1598, 1828, 2537]. Gas [2284, 793]. Gate [1091, 2075, 480]. Gate-Level [1091]. Gates [1089, 1088]. Gathering [462, 464, 439, 447]. Gaussian [306]. Gaussians [762]. Gaussians-Based [762]. Gaze [91]. Gazetteers [701]. GCAP [1174]. GDT [2406]. Geffe [2124]. GeLog [748]. GEM [2725]. Gemmation [616]. Gene [745, 54, 518, 2809]. General [2610, 538, 1869, 2721, 65, 2268, 1950, 1507, 2836, 1700, 1699, 2599, 1290, 2270]. Generalisation [2821]. Generalised [2050, 1561, 2814]. Generalization [2815, 2786, 2368, 1640]. Generalizations [1441, 1795, 2363]. Generalized [1078, 501, 2588, 1548, 2365, 2124, 1407, 2846, 269, 1925, 342, 364, 1251]. Generate [2643, 2367]. Generated [665, 166, 1925, 2503, 441]. Generating [1890, 2637, 35, 2328]. Generation [1239, 1188, 2469, 2873, 1887, 2593, 1907, 321, 1214, 2185, 474, 1589, 2235, 2872, 2766, 2533, 2104, 2199, 2189, 229, 414, 123, 386, 1215, 521, 2858, 334, 2122, 376, 2182, 393, 976, 2876]. Generational [2537]. Generations [1263]. Generative [372]. Generator [2115, 2124, 2127]. Generators [2062, 2647]. Generic [1871, 500, 102, 1977, 2528, 941, 2144, 676, 883, 2535, 2617, 1301, 1361, 2725]. Genetic [759, 1106, 1105, 2347, 2486, 507, 1101, 748, 2304, 2325, 2824, 760, 2404, 2826, 1092, 761, 773, 2501, 1104]. Genetic-Fuzzy [2824]. Genomics [1512]. Genotype [1098]. Genus [1572, 1552]. Geographer [679]. Geographic [703, 711]. Geographical [700]. Geographically [408, 1081]. Geography [1783]. Geomatics [688]. Geometric [2058, 902, 938, 1194, 1994, 1384, 682, 2699, 566, 868]. Geometrical [1671, 1674, 878, 246]. Geometry [815, 1551]. GeoNotes [577]. German [795, 153]. Germany [935]. GESA [867]. Gesture [2428, 152, 310, 91]. Gesture/Speech [2428]. Gestures [236]. Getting [2505]. GF [2103, 1615, 2592, 2604, 2105]. GH [2603]. GHS [2116]. Gigabit [552]. Giotto [1129]. Girard [1225]. Girard-Reynolds [1225]. Given [1396]. Giving [698]. Glenohumeral [979]. Global [2574, 2908, 382, 1851, 268, 221, 694, 1052, 1942]. Globule [1166]. Glomerular [508]. Glossary [1315]. GLOWER [2543]. GMM [369]. GMM-Based [369]. Go [1387]. Goal [378, 2781]. Goal-Directed [378]. Goals [1601]. Gödel [2193]. Goemans [1347]. Gong [2637]. Gong-Harn [2637]. Good [1440]. Goods [1635, 1736]. GOTS [2457]. GPRS [1611]. Gradient [2471, 1048, 975]. Gradient-Based [975]. Gradients [145]. Grading [44]. Grafting [1072]. Grain [1592]. Grained [1204]. Grammar [2922, 2572]. Grammars [2920, 2910, 2925, 2872, 1792]. Grammatical [2826, 2500]. Grams [2885]. Grand [1377]. Granular [681, 2381, 2382, 2357, 1247]. Granules [2377, 2359]. Graph [2920, 763, 940, 784, 659, 660, 608, 663, 2919, 1208, 2918, 1355, 2624, 2639, 2195, 143, 441]. Graph-Matching [143]. Graphic [443, 243]. Graphical [2047, 1870, 602]. Graphics [154]. Graphs [651, 2862, 654, 655, 1371, 656, 2473, 660, 662, 1388, 2679, 664, 667, 1390, 670, 671, 672, 2078, 2373, 1354, 1355, 1370, 673, 1579, 1387, 675, 676, 55, 1562, 677, 1386]. Gravity [2280]. Gray [770, 1040]. Greedy [1399, 676, 1543]. Grey [1047, 2821]. Grey-Level [1047, 2821]. Grid [1969, 1962, 705, 1960, 1961, 1958, 1970, 1963, 1964, 1974, 1965, 1968, 680, 1959, 1972, 1966, 1973, 1962, 1754, 1967, 1759, 805].

Grid-Wide [1967]. Grids [1971]. Ground [2084]. Groundness [1830]. Group [1748, 2442, 2148, 1618, 2837, 203, 1751, 1742, 1325, 482, 569, 424, 2753, 2722, 1663, 1917, 1750, 1536, 2617, 1361, 1654]. GroupCast [603]. Grouping [102, 142, 2079]. Groups [2139, 482, 543, 1360, 1743]. Groupware [26, 2422, 1809]. Growing [972, 1051]. Growth [126]. Guarantee [1590]. Guaranteed [1440, 1]. Guaranteeing [2172]. Guarded [2184, 2185]. Guards [1383]. Guidance [856, 858, 2317, 2452, 461, 843, 899]. Guide [1442]. Guidebooks [591]. Guided [867, 904, 1893, 942, 882, 1032, 1006, 855, 842, 996, 899]. Guidelines [2423]. Guruswami [1551]. Gyrokinetic [1583].

H [1682, 890, 1310]. H.263 [338]. Haar [2244]. Hacking [573]. HAL [1827]. Half [918]. Halftone [1115]. Hamiltonian [1372]. Hamming [2614, 1548, 2611]. Hand [850, 2736, 2327, 1088]. Hand-Drawn [2327]. Handheld [2431]. Handlers [1599]. Handling [2782, 1826, 1346, 1430, 1863]. Handprinted [2799]. Handset [369]. Handwriting [100, 2253]. Handwritten [436, 787]. Happily [396]. Haptic [935, 888]. Haptics [2290]. Hard [2062, 2058, 1467, 660]. Hardness [1782]. Hardware [2642, 2711, 1099, 1091, 214, 554, 1115, 2145, 1763, 555]. Harmonics [147]. Harn [2637]. HARP [500]. Harvesting [851]. Hash [2099, 2102, 2708, 2101]. Hash-Based [2099]. Hausdorff [311, 249, 1990]. Having [514, 58]. Head [245, 2300, 899]. Head-Mounted [899]. Heads [2301]. Healing [1097]. Health [1164, 2278]. Heart [1676, 1675, 1017, 1671, 1680, 841, 1672, 535]. Hearts [1683]. Heat [121]. Heavy [1873, 1331]. Heavy-Tailed [1873]. Heel [850]. Helical [1054]. Hellman [2148, 2727, 2129]. Help [1733, 2850, 2430]. HEp [523]. HEp-2 [523]. Hepatic [951]. Herbrand [2193]. Hereditary [656]. Heterogeneous [1118, 498, 1936, 1601, 1575, 1243, 129, 1276, 1926, 809, 584]. Heuristic [2438, 2422, 1328, 1903]. Heuristics [1185, 1885, 1898, 1915, 761]. Hexapod [1024]. Hidden [593, 2133, 134, 1378, 1857, 1600, 63, 1561, 2259, 2814, 1520]. Hiding [1642, 263, 2291]. Hierarchical [1119, 874, 1947, 1586, 2773, 1141, 514, 240, 1360, 2579, 323, 1520, 2045, 1662, 224, 1654]. Hierarchies [1441]. Hierarchy [259, 1164, 2182]. High [1081, 1962, 1595, 1960, 916, 2921, 1604, 1912, 2526, 2477, 54, 1437, 2100, 2238, 1684, 2136, 2586, 1965, 1594, 2187, 1588, 2535, 1606, 953, 1939, 223, 2270]. High-Bandwidth [1606]. High-Dimensional [54, 223]. High-Level [2921, 1912, 1437, 2187, 2535]. High-Pass [2270]. High-Performance [1595, 2526, 1965, 1588]. High-Resolution [916]. High-Speed [1604]. Higher [2625, 2216, 1830, 2695, 2308, 2523, 2197]. Higher-Order [2216, 2197]. Higher-Precision [1830]. Highlight [451]. Highly [2121, 69, 1157]. Highly-Configurable [1157]. Hilbert [2266]. HIMS [2779]. Hip [1071]. Hippocampal [819, 1016]. HiQoS [1780]. Histogram [320]. Histological [903]. Historical [114, 1417]. History [593, 404, 456]. HLA [260]. HLA-Based [260]. HMM [310, 38]. Hobby [402]. Hoc [1248, 1176]. Hole [140]. Hollywood [396]. Home [597, 1281, 1283]. homed [1168]. Homework [479]. Homogeneity [2905]. Homogeneous [1559, 2565]. Homomorphic [2748, 2724]. Homonymy [2868]. hop [305]. Hopfield [2828]. Horse [402]. Horses [1496]. Hospital [403, 525]. Host [1177]. Hostile [2505]. Hot [792]. Hotlink [666, 1406]. Hough [149, 1987, 364]. House [408, 706]. HP [1879]. HP-Sequences [1879]. HPC [1580]. HPC-Clusters [1580]. HPSG [2860, 2847]. HST [2457]. Hull [2574, 2590]. Hulls [649]. Human [934, 1678, 1676, 2134, 1675, 2293, 92, 1047, 2289, 2290, 2481, 1101, 219, 249, 2413, 1674, 104, 903, 941, 2346, 72, 1690, 533, 236]. Human-Centered [2290, 219, 2413]. Human-Like [1101]. Human-Robot [2293]. Humanized [713]. Humans [2086, 2288]. Hurdles [2447]. Hurts [761]. Hwang [1658]. Hwang-Rao [1658]. Hybrid [1119, 2566, 2870, 759, 1846, 1218, 83, 252, 939, 2050, 510, 2579, 171, 762, 954, 1689, 1847, 2232]. Hydrophobic [1879]. Hyper [2041]. Hyper-Set-Theoretic [2041]. Hyperacute [1007]. Hyperalgebraic [2838]. Hyperbolic [1550]. Hyperelliptic [2595, 1621]. Hypergraphs [627]. Hyperlink [1522]. Hypermedia [426, 299]. Hyperpolyhedron [1924]. Hyperspectral [757]. Hypotheses [2841]. Hypothesis [2844, 2511]. Hyppocrates [542]. Hysteresis [1231].

I/O [2077, 1596]. I/O-Efficient [2077]. Icon [290]. ICP [904]. ICrafter [582]. ICSPACE [2330]. IDA’01 [70]. Ideal [2549, 1537]. Ideas [1368]. Idempotent [1853]. Identification [733, 2231, 262, 1565, 2134, 2753, 1643, 2791, 526, 2889, 948, 369]. Identify [767, 2502]. Identifying [2444, 1560, 2361, 1542, 692, 1965, 2379, 1606]. Identity [2641, 363]. Identity-Specific [363]. IDS [1153, 2499]. IF [764]. II [1544, 530]. ILF [2819]. Ill [2862]. Ill-Formed [2862]. Illumination [118, 115]. Illumination-Invariant [118]. ILP [2376]. Im [204]. Image [266, 758, 926, 1538, 267, 1619, 313, 1057, 969, 294, 1010, 710, 320, 316, 2258, 1382, 250, 856, 325, 1064, 493, 258, 2816, 2311, 121, 1070, 103, 2235, 829, 823, 96, 252, 1046, 221, 942, 882, 2241, 257, 508, 130, 1032, 998, 142, 1043, 2481, 2236, 1063, 782, 2252, 720, 314, 270, 269, 2239, 717, 1006, 123, 2372, 936, 523, 1014, 2818, 308, 334, 954, 899, 2234, 566, 841, 2819, 722, 146, 949, 274, 2823, 965, 342, 352, 560, 2243, 1044, 279, 1005, 124, 346, 243, 231, 2329, 302, 125, 278, 447, 251, 996, 2328, 341, 357, 280, 1008]. Image-Based [243]. Image-Feature [2235]. Image-Gathering [447]. Image-Guided [882, 1032, 1006, 996]. Imageries [2804]. Imagery [912, 453, 757, 913]. IMages [97, 745, 884, 1686, 925, 968, 1011, 149, 362, 250, 1679, 1062, 938, 932, 1040, 916, 987, 1673, 1192, 1049, 986, 721, 102, 101, 872, 111, 827, 1056, 2481, 1054, 306, 890, 943, 881, 903, 941, 1642, 309, 522, 911, 2257, 335, 849, 964, 527, 1115, 878, 910, 105, 970, 988, 875, 894, 1029, 948, 2000, 2243, 533, 973]. Imaging [1678, 838, 909, 993, 1015, 1681, 837, 1677, 522, 911, 937, 1690, 981, 868]. Imai [2117]. Imbalanced [36]. Imitation [1261, 2565]. Imitation-Based [2565]. Immersive [390, 218]. Immune [1099, 2120, 1160, 1636, 1628]. Immunity [2157]. Immunology [1652]. Impact [1747, 1631, 1162, 1437, 1243, 2025]. Imperfect [2732, 2376]. Implant [1000]. Implantation [1071]. Implantology [1018, 860, 904]. Implementation [2544, 1138, 2053, 1099, 2139, 1381, 2125, 2108, 724, 1954, 468, 2776, 1179, 1091, 2054, 1242, 2322, 2592, 254, 1134, 554, 574, 2824, 2177, 1633, 2758, 359, 1695, 522, 2818, 178, 1094, 726, 1092, 393, 2875, 1314, 811, 2890, 1959, 1823, 2713, 555, 1647, 2419]. Implementation-Defined [1314]. Implementations [552, 1256, 813, 1280]. Implemented [1198]. Implementing [2893, 1833, 2460]. Implications [1889, 2484, 1838]. Implicit [148, 452, 2204, 1050]. Implicitly [1841]. Implied [1907]. Implies [2752]. Importance [2620, 1014, 2644]. Impossibility [2714]. Impossible [2142].

Improve [875, 2882, 471]. Improved [1617, 655, 2309, 779, 818, 375, 2709, 2477, 2592, 836, 2762, 231, 227, 871, 898]. Improvement [266, 1347, 22, 1522, 255, 19, 10, 517, 2627, 339, 766, 2613]. Improvements [934]. Improving [110, 884, 876, 1599, 2590, 15, 1916, 2903, 1587, 2631, 2182, 796, 446]. Improvisational [86]. iMRI [899]. iMRI-guided [899]. In-Vitro [849]. Incentives [1726]. Including [2363]. Incomplete [633]. Incompressibility [828]. Inconsistent [2896, 1844, 1427]. Incontinence [514]. Increase [1159]. Incremental [611, 2102, 1103, 1924, 1433, 1088]. Independence [1668]. Independencies [423, 61]. Independent [916, 946, 130, 1777, 2153, 2701]. Index [32, 1201, 1230, 1260, 1268, 1289, 1330, 1342, 1408, 1458, 1486, 1533, 1574, 1612, 1669, 1691, 1738, 1753, 1784, 1802, 1813, 1845, 1930, 1949, 1957, 1975, 2003, 71, 2056, 2085, 2097, 2130, 2165, 2230, 2286, 2331, 2407, 2437, 2461, 2518, 2540, 2581, 2609, 2648, 2676, 2690, 2707, 2729, 2740, 2768, 2842, 2891, 2909, 2933, 95, 155, 181, 197, 213, 371, 384, 409, 490, 537, 575, 607, 635, 648, 678, 712, 814, 1073, 1087, 1117, 1152, 1165, 1184, 1317, 1717, 2662, 2178, 2408, 258, 259, 1367, 223]. Indexed [2209]. Indexes [2816]. Indexical [282]. Indexing [220, 216, 326, 451, 321, 259, 2252, 291, 224]. Indicators [499]. Indices [450]. Individual [504]. Individualized [486]. Indoor [137, 580]. Induced [740, 2368, 671]. Induction [2402, 1507, 2197, 50]. Inductive [1848, 2401, 748, 2373, 1478]. Industrial [2413, 154, 406, 1763, 2035, 1322]. Inequalities [1874, 1474]. Infarction [1683]. Inference [789, 360, 611, 721, 383, 65, 1757, 741, 751, 2826, 776, 2224, 2500, 2789, 1478, 1485]. Inferring [912, 2226]. Infinite [1223, 1908, 2913, 2917]. Infinite-State [1908]. Infiniteness [2363]. Inflation [854]. Influence [1335]. Info [2830]. Info-Miner [2830]. InfoPipes [194]. Information [2439, 2730, 2896, 2090, 968, 880, 462, 450, 2907, 577, 2878, 1494, 1774, 679, 1946, 2434, 172, 596, 1651, 2735, 443, 1414, 1475, 1162, 872, 700, 2396, 217, 1549, 736, 998, 454, 1502, 708, 1275, 2483, 1645, 464, 239, 861, 2881, 2830, 336, 2915, 1086, 2316, 438, 579, 2362, 2880, 2431, 2795, 605, 2025, 2366, 63, 1450, 701, 303, 1464, 1490, 457, 2377, 2359, 435, 2502, 2392, 1308, 1527, 2516, 2615]. Information-Based [998]. Information-Retrieval [1946]. Information-Theoretic [2615]. Informative [2276, 45, 1501, 2883, 1315, 1313, 1312, 1314, 1316]. InfoScope [596]. Infrared [729, 981]. Infrastructure [1656, 2278, 195, 1741, 1596, 584]. Infrastructures [639]. Inhabitants [2572]. Inherent [2450, 751]. Inherently [2426]. Inherently- [2426]. Inheritance [2031]. Initial [1114, 1093]. Initiative [2867, 2734, 2706]. Injective [2832]. Injurious [2384]. Innovation [2335]. Input [289, 2429, 1607, 1531]. Input/Soft [2614]. Inputs [2862]. INs [1343]. InSAR [141]. Insecure [2710]. Insecurity [2132]. Inseparable [2646]. Insertion [51, 1023]. Inspiration [1266]. Inspired [716]. Instability [1986]. Instance [1472]. Instances [2637]. Instantaneous [617]. Instantiation [1714, 34]. Institution [2701]. Institutions [2694, 2561, 2703, 2564]. Instruction [1897]. Instrument [930]. Instrumentation [982]. Insulation [2790]. int [890]. Integer [1564, 1950, 1798, 2123, 2763, 1847]. Integers [2600]. Integrability [1996]. Integral [861, 461]. Integrals [2380]. Integrated [1154, 1679, 453, 1961, 2872, 1504, 1273, 1052, 1288, 1339]. Integrating [2896, 379, 2440, 2039, 2396, 2413, 1323, 586, 1847, 246, 322]. Integration [645, 1428, 1412, 1429, 2444, 1084, 1844, 885, 2244, 2402, 1414, 486, 2456, 1430, 2457, 1419, 1736]. Integrations [2360]. Integrity [174, 1934, 554, 2761, 2114]. Intellectual [571, 1422]. Intelligence [412, 2341, 2338, 2332, 439, 754, 713, 781, 410]. Intelligent [758, 80, 81, 430, 79, 68, 2287, 478, 483, 782, 1610, 460, 74, 1250, 464, 1614, 768, 487, 73, 2359, 60, 2509, 305, 298, 2806, 392]. Intensionality [2845]. Intensity [902, 977, 828, 878]. Intensity-Based [828, 878]. Intensive [478]. Intention [2785]. Inter [183, 1751, 978, 845, 1737]. Inter-Enterprise [183, 1737]. Inter-Group [1751]. Inter-slice [978]. Inter-subject [845]. Interacting [601, 2173, 243]. Interaction [1883, 84, 1324, 2336, 2410, 219, 2318, 2353]. Interactions [1038, 427, 1226, 1267]. Interactive [1765, 2691, 2301, 1059, 870, 2303, 394, 1326, 1890, 391, 2434, 885, 1524, 442, 288, 292, 2416, 86, 1238, 869, 972, 958, 1870, 390, 466, 2412, 1320, 1613, 713, 1761, 2714]. Interactivity [2290]. Interchangeability [1906, 2466, 1920]. Intercom [589]. Intercomponent [1704]. Interconnection [1589, 2827]. Interdependent [1107]. Interdisciplinary [2413]. Interest [216, 497, 43, 432]. Interesting [2884]. Interests [428]. Interface [173, 2287, 2047, 599, 2299, 1318, 2436, 1319, 887, 218, 945, 522, 2411, 2458, 2535, 1304, 785, 1128, 1322]. Interfaces [2302, 430, 593, 2425, 2420, 1132, 184, 602, 1321, 1320, 279]. Interfacing [1155]. Intermediate [1471]. Internal [1752, 931]. Internet [500, 448, 1239, 80, 1609, 1720, 1064, 427, 191, 1259, 288, 719, 188, 1644, 196, 2340, 813, 1232, 2392, 2292, 2330, 1423, 1167, 785, 436, 1646, 227, 1634]. Internet-Based [1720, 1259]. Internets [1610]. Interoperability [1281, 2632].

Interoperable [500]. Interplay [404]. Interpolation [1757, 977, 2773, 361, 1551]. Interpretability [2770]. Interpretation [1205, 1816, 2534, 152, 2212, 732]. Interpretations [764]. Interpreters [379]. Interpreting [1899]. Interrupt [1599]. Intersection [1223, 1356]. Interstitial [508]. Interval [2607, 2374, 45, 2201, 2370]. Intervals [802]. Intervention [1683, 1024]. Interventional [956, 869, 854]. Intimate [2410]. Intra [923, 861, 952]. Intra-operative [923, 861, 952]. Intraoperative [944, 943, 997]. Intravascular [959, 527]. Introducing [1825, 1673, 2221]. Introduction [1459, 1982, 1803, 2086, 2649, 2166, 1692, 2380, 2663, 2377, 2060, 2328]. Intruder [1614]. Intrusion [1158, 1162, 1164, 1660, 574, 1163, 1157, 1155]. Intrusion-Detection [1158]. Intrusions [1160]. Intuitionistic [1842, 2195, 731]. Invariant [744, 2255, 118, 98, 115]. Invariants [1572, 2654]. Invasive [955, 858, 1001, 1683, 1025]. Inventing [1464, 1490]. Invention [1526]. Inverse [2281, 830]. Inversion [2133]. Investigating [911, 66]. Investigation [2478, 2635, 2023, 795]. Investigations [702, 528]. Investment [2543]. Invited [649, 650]. InViWo [87]. Involutions [2930]. Involvement [1373]. Involving [1235, 393]. iOpt [1903]. IP [1178, 293, 2638, 1243, 1249, 811, 1234, 1749, 256]. IPP [2109]. IPSec [552]. IPSec-Compliant [552]. IPv6 [1168]. IR [2738]. Iris [2231, 743, 2477]. Irreducible [2104]. iSAGAs [191]. Ischaemic [1688]. ISLE [2867]. Isomorphism [2063, 1225]. Isomorphisms [1355]. Isoperimetric [651]. ISP [1168]. Issue [1897]. Issues [1120, 2445, 468, 752, 1575, 1950, 182, 590, 1299, 1302, 592]. Items [277]. Iterated [2001]. Iteration [1789]. Iterations [1553]. Iterative [905, 1999]. ITMS [486]. ITSEP [2439]. ITTALKS [647]. IV [2143]. IVUS [969]. IWShopper [460].

J [1311]. JAC [156]. JACK [2493]. Jacobian [1572, 1621]. JADE [460]. Japan [1460, 1487, 935]. Japanese [785, 1521]. Jaundice [503]. Java [2672, 1774, 621, 1138, 2033, 30, 1271, 164, 1064, 1270, 1933, 2020, 1944, 1269, 2052, 160, 177, 156, 179, 726, 1588, 157, 583, 1945, 1863]. Java-Based [1933, 1945]. JavaCloak [179]. JavaPod [159]. Jaw [934]. JIT [583]. Jitter [2263]. JKDD01 [2400]. Job [1375, 1341, 613, 614, 1580, 1338, 2505]. JobQueue [1967]. Join [259, 374]. Joins [1078]. Joint [353, 1002, 2274, 960, 979]. Jordan [1997]. JSAI [2400]. Judgment [2849]. JULIUS [1032]. Juridical [551]. Just [1815, 14]. Justification [1831].

Kalman [149, 2791, 146, 2267]. Kansei [2492, 2329]. Kansei-Oriented [2329]. KASUMI [2146]. KDD [468, 2400]. Kedlaya [2158]. Kerberized [540]. Kernel [131, 1777, 128]. Key [2557, 2163, 550, 2607, 538, 2148, 2616, 264, 630, 2582, 2603, 2766, 2600, 2124, 2743, 1644, 2752, 2744, 2135, 2728, 1650, 2633, 1750, 2632, 2741, 1426, 2147, 1361]. Key-Evolving [2728]. Key-Privacy [2163]. Key-Stream [2124]. Keyblock [280]. Keynote [154]. Keyphrases [2733]. Keys [1360, 2128]. Keywords [1510]. KeyWorld [1510]. Kidney [844]. Killer [2525]. Kind [2306]. Kinetic [1678]. KISS [594]. Kit [1959]. KM [8]. Knapsack [1765, 1850]. Knee [886, 863]. Knife [867]. Knowing [1618]. Knowledge [2726, 2555, 495, 2222, 633, 1492, 703, 119, 27, 1962, 2907, 2155, 2154, 2560, 472, 1493, 2207, 2188, 2401, 411, 18, 2402, 2485, 507, 2479, 15, 1760, 434, 1509, 2857, 486, 414, 16, 2405, 503, 2579, 2498, 2353, 761, 435, 433, 2399, 785, 532, 534, 1764, 1531, 2040, 1962].

Knowledge-Based [507, 1760, 435]. Knowledge-Bases [2222]. Knowledge-Poor [2857]. Known [2895, 2143, 2717]. Known-IV [2143]. Koblitz [2593]. Kolmogorov [2057, 1474]. Korea [1453]. Kosko [2794]. Kosko-Neuro-fuzzy [2794]. Kripke [1209]. Krylov [1584]. Krylov-Type [1584]. Kumar [2613]. Kumar-Rajagopalan-Sahai [2613].

L [1313]. Labeling [625, 1397, 1394]. Labelling [1915]. Laboratories [1754]. LAD [1030, 1367]. Lagrangian [1953]. Lambda [1223, 610, 2023, 1224]. Lambda-mu [1224]. Lamport [1776]. LAN [1175]. Landing [2775]. Landmarks [110, 884, 698, 2552, 1042]. Landscape [1377, 2918, 335]. Landscapes [1756]. Lanes [122]. Language [1492, 2691, 2217, 2053, 2066, 618, 418, 2738, 1213, 2872, 2020, 381, 2848, 1129, 24, 2054, 2850, 455, 1318, 2519, 348, 2021, 1479, 642, 2498, 2847, 2784, 1313, 1312, 1303, 695, 1819, 93, 1196, 1785, 233]. Language-Based [348]. Language-Defined [1313, 1312]. Languages [1185, 2910, 2536, 1190, 2664, 1484, 1772, 619, 40, 1477, 567, 1792, 2535, 1304, 1485]. LANs [1604]. Laparoscopic [897, 1068, 1020, 887, 816, 1019, 954]. Laplace [112]. Large [41, 1539, 548, 1340, 1742, 1564, 2637, 2449, 2428, 964, 1286, 1741, 605, 1624, 2460, 307, 233, 1743, 66, 1662, 1333, 2585]. Large-Scale [548, 1742, 2449, 1286, 1741, 2460]. Laser [111, 844]. Laser-Based [844]. Latencies [1897]. Latency [1739, 1591]. Latency-Tolerant [1591]. Latent [795]. Latin [1853]. Lattice [2638, 1984, 1645]. Lattices [2605]. Laws [1055]. Layer [1448, 1272, 2281]. Layered [624, 1099, 399]. Layers [99]. Layout [1752, 691]. Lazy [1223, 2508]. LDAP [1084]. Leader [2074, 2004]. League [2354]. Leak [2162]. Leap [1996]. Leap-Frog [1996]. Learn [397]. Learned [2440, 1496]. Learner [2483]. Learning [1112, 2544, 1451, 1482, 2093, 397, 723, 2467, 1188, 2152, 59, 1484, 2543, 1472, 1849, 36, 1772, 1470, 144, 474, 825, 1582, 1497, 482, 2423, 2485, 2395, 1476, 1477, 1480, 1475, 413, 2548, 736, 783, 782, 2821, 1471, 1110, 515, 485, 2317, 2903, 1507, 768, 1508, 2316, 1481, 2877, 486, 2491, 2094, 1479, 466, 1512, 1468, 1111, 2346, 761, 2367, 445, 93, 534, 536, 1104, 1465, 2324, 2806, 2687, 1478, 780]. Least [148, 2554]. LECTURE [2316]. Left [1678, 1011, 986, 872, 1017, 1674, 907]. Legendre [1473]. Legislation [1162]. Lemmas [2659]. Lemmatizer [1774]. Length [1539, 2303, 1993, 2742, 613, 1625, 1607, 1995, 1554, 1244]. Length-Sensitive [1625]. Lengths [794]. Lenient [543]. Lenient/Strict [543]. LERS [2384]. Lesions [908, 718, 826, 989, 913, 963]. Lesk [2853]. Lessons [2440, 1496, 1526, 1900]. Letters [2877]. Level [2466, 808, 1255, 658, 2921, 506, 1912, 1047, 1091, 1437, 1065, 2821, 1133, 1684, 1740, 1594, 2019, 2187, 2535, 1596, 1973, 2296, 883, 180]. Levels [1140, 1437]. Leveraging [2471, 484]. Lexical [1189, 2867, 2880, 1291, 2851]. Lexicographic [2049]. Lexicon [91]. Lexicons [2876]. LFSR [2126]. Liability [551]. Libraries [561, 275]. Library [325, 489]. Life [581, 1149, 1266, 1261, 1800, 1818]. Lifelike [78]. Lifetime [909]. Lifting [1951, 1567]. Light [2778, 909, 244, 1022]. Lighting [2254]. Lighting-Effects [2254]. Lightweight [810, 17, 1742, 1226, 64, 23]. Like [2027, 1101, 2041, 2510, 909]. Likelihood [1568, 911, 1009, 2323]. Likely [1568]. Lime [1935]. Limit [1761]. Limitations [2454]. Limiting [2622, 2001]. Limits [786, 910]. Line [1401, 2625, 1987, 2802, 1375, 311, 114, 2072, 1351, 2260, 672, 139, 781, 2261]. Line-Feature-Based [311]. Linear [2610, 1874, 1842, 1886, 2258, 2125, 1584, 2154, 1540, 1106, 2071, 2526, 2590, 2752, 2628, 683, 2192, 2835, 2635, 1215, 2586, 1799, 2110, 1543, 2182, 1042, 923, 927, 2797, 1004, 1474]. Lines [1988, 1397, 108, 107]. Lingual [449, 2886]. Linguistic [548, 253, 719]. Linguistics [2455]. Link [596, 454]. Link-Based [454]. Linkages [1343]. Links [2432]. Linux [214]. LIPE [23]. Lisp [2069]. List [2661, 2073, 2887, 1551, 2141]. Listen [2520]. Lists [1349]. Literary [1403]. Literature [795]. Live [1733, 396, 407, 235]. Liveness [2068, 2911]. Liver [1033, 1006]. Living [1264]. LLEC [341]. Load [1602, 421]. Local [733, 1883, 1832, 2203, 320, 245, 2681, 927, 2574, 2183, 2546, 21, 96, 909, 1893, 221, 2919, 1568, 37, 1354, 1514, 2187, 1896, 1199, 1044]. Local-Vertex-Connectivity [1354]. Locality [1595, 1587]. Localization [140, 706, 109]. Locating [826]. Location [926, 578, 1649, 2689, 715, 577, 579, 1286, 960]. Location-Based [577]. Lock [1221]. Locking [1625]. Locomotes [1266]. Log [1613]. Logarithm [2728]. Logarithms [2598, 2137]. Logging [1598]. Logic [1445, 1843, 2209, 2463, 2892, 2194, 2010, 2193, 1842, 1832, 2566, 2203, 2043, 1202, 2472, 2042, 1835, 2183, 800, 1846, 2218, 2621, 2534, 1214, 2226, 2296, 1844, 1833, 2525, 1213, 2898, 719, 748, 1815, 184, 2568, 1666, 2523, 345, 1834, 2210, 2373, 1839, 2569, 2777, 2791, 2383, 1209, 632, 1838, 2825, 2571, 1212, 2083, 2005, 2182, 1925, 2198, 1630, 1821, 1847, 1088, 643, 2538, 730, 797, 2515, 1530]. Logic-Based [2296]. Logical [2186, 2896, 545, 567, 2490]. Logics [2184, 2693, 738, 2476, 2180, 2308, 2211, 2201, 2009, 2905]. Logs [473, 1528, 471]. Long [2242, 4, 2393, 1606, 973]. Long-Term [2242, 2393, 1606]. Look [1820]. Looks [679]. Loop [1454]. Loops [2224]. Loose [2702]. Loss [1473, 1348]. Loss-Free [1348]. Lossless [1115]. Lossy [1220]. Lotion [767]. Lottery [1634]. Low [1598, 1396, 1539, 283, 2268, 1554, 580, 2636, 2147, 341]. Low-Complexity [341]. Low-Cost [1598]. Low-Memory [341]. Low-Passed [2268]. Lower [2057, 2837, 1869, 2688, 668, 2361, 1437, 1867, 1579, 1868]. LSA [448, 2859]. LTL [2065, 2911, 1217]. LTS [608]. Lübeck [2695]. Lumen [916]. Lump [565]. Lung [901, 1054, 1052]. LUPS [2567]. Lyapunov [2005]. Lying [2895].

M [1314, 2712]. MAC [2606]. Machine [2276, 2543, 1106, 1497, 2006, 2485, 1261, 782, 2079, 1512, 1094, 445, 368, 2324]. Machines [1469, 1264, 2068, 2015, 132, 1468, 1111, 1588, 133]. Macroset [1791]. MACs [2616]. Made [2495]. MAF [914]. MAgentA [89]. Magic [2932]. Magnetic [1679, 1062, 916, 941, 910, 991, 948, 862]. Magnetization [2623]. Magneto [954]. Magneto-optic [954]. Magnets [14]. Magnified [896]. Mail [2720, 2111]. Mailbox [1940]. Mailbox-Based [1940]. Main [1368, 539, 881, 1671]. Maintenance [2036, 11, 1864]. Majority [2621, 1778]. Majority-Logic-Decodable [2621]. Make [1977]. Makers [1725, 2087]. Making [581, 2782, 2543, 718, 2454, 2298]. Mall [1721, 1724]. MAMBO [61]. Mammogram [826]. Mammograms [497, 511, 1055]. Mammographic [516, 517, 963]. Managed [1599]. Management [810, 6, 500, 1653, 1822, 1656, 2474, 1807, 1970, 1254, 1719, 255, 2396, 15, 1964, 572, 1250, 1258, 1086, 650, 2052, 438, 1241, 16, 1713, 1709, 1710, 1708, 1703, 1707, 1701, 1711, 1712, 1455, 1750, 1231, 433, 1288, 286, 1422, 301]. Manager [1167]. Managing [2039, 18, 1248]. Managment [579]. Mandarin [291]. Manifesto [2669]. Manifolds [1979]. Manipulation [1069, 1065, 456, 2427].

Manipulations [2931]. Manipulator [1068, 887]. Manoeuvre [1708]. Manufacturing [1905]. Many [1252, 643]. Many-to-One [1252]. Map [690, 694, 1453, 2774, 484]. Map-Building [2774]. Mapped [1452]. Mapping [1789, 1902, 1680, 1690]. Mappings [1987, 829]. Maps [1192, 1756, 130, 2628, 2683, 699, 1044, 125]. Mark [1442]. Mark-Up [1442]. Marked [2913]. Markers [1245]. Market [2347, 2486, 2321, 1136]. Marketing [1454]. Marketplaces [1723]. Markets [1725, 2345]. Marking [2888, 376]. Markov [940, 134, 972, 1378, 63, 2259, 2814, 1520]. Marks [2890]. Markup [418, 567, 1319]. Marriage [1861]. Mart [2346]. Mass [829, 517]. Massive [565, 1076]. Massively [1106, 1583]. Match [966]. Matching [622, 2470, 311, 1040, 330, 1039, 1899, 474, 1403, 1875, 2260, 2135, 879, 674, 143, 153, 307, 1042, 116, 272, 281, 963]. Matching-Cut [674]. Matchings [1369, 1374]. Material [2782, 486]. Materialization [1527]. Materials [755, 265]. Mathematical [1688, 2214, 1344, 1952]. Mathematics [1801]. MATLAB^{T M} [1100]. Matrices [1563, 2121]. Matrix [1396, 2374, 1564, 1549, 335]. Matsui [2597]. Matsumoto [2117]. Maturity [11]. Maude [2697]. Maui [1336]. Max [1353]. Maxillofacial [932, 999, 871]. Maximal [620, 1369, 1556]. Maximum [1369, 653, 1602, 1568, 2590, 911, 376, 2854, 1009, 2323, 1856]. Maxwellian [1529]. MB [225]. McEliece [2140, 2757]. McEliece-Based [2140]. MD [1078]. MD-Joins [1078]. Mean [921]. Meaning [2848, 2667]. Meaningful [52]. Means [769, 2121, 1105, 2480, 729, 2364, 1197, 2810, 5, 1104, 1023, 981]. Means-End [5]. Measure [245, 2732, 420, 1998, 846]. Measurement [135, 1676, 1680, 930, 849, 817, 910, 2026, 990, 1256]. Measurements [126, 1173]. Measures [752, 893, 2380, 115, 2370, 2787]. Measuring [573]. Mechanical [1686, 1003, 930, 842, 841]. Mechanically [2069]. Mechanics [1670, 2422]. Mechanism [1168, 1081, 1172, 804, 442, 1599, 887, 2632, 811, 357]. Mechanisms [2558, 1099, 1463, 1488, 2099, 2172]. Mechatronic [1071]. Media [385, 909, 301, 2291, 1082]. Medial [819, 113, 867]. Median [649, 1407]. Mediated [2313, 2293]. Mediation [1457, 1723]. Mediator [403]. Medical [494, 884, 496, 491, 1064, 493, 938, 823, 1046, 1015, 510, 512, 154, 2405, 1060, 522, 1014, 524, 964, 937, 529, 531, 731, 2000, 352, 560, 868, 536, 2884]. Medicine [492, 2524, 774]. Meeting [2439, 586]. MEG [983]. MELISSA [448]. Melody [281]. Membership [737, 203, 1742, 2898, 2786, 2549, 1650]. Membranes [616, 1798]. Memetic [2462]. Memorizing [119]. Memory [1138, 1822, 1340, 1585, 1589, 1334, 2620, 2611, 2242, 2052, 1587, 1338, 341]. Menger [673]. Menger-Connectivity [673]. Meningitis [2402, 2403, 2406]. Meningoencephalitis [2401, 2405]. Meniscal [721]. Mental [703, 717]. Mercury [1828]. Merge [210]. Merging [2456]. Merit [46]. Merit-Based [46]. Mesh [366, 1012, 1576]. Mesh-Based [1012]. Meshes [1058]. Message [1598, 1940, 1269, 1285, 2078]. Message-Oriented [1285]. Messages [2742, 1941]. Messy [1089]. Meta [780, 2664, 161, 382, 2296, 178, 2215, 373, 435, 180, 1532]. Meta-analysis [1532]. Meta-data [435]. Meta-learning [780]. Meta-level [2296, 180]. Meta-model [2664, 161, 178]. Meta-programming [382, 373]. Meta-programs [2215]. Metabolic [506]. Metabolism [909].

Metacomputations [381].

Metacomputing [1759]. Metadata [1426]. MetaFace [2301]. MetaKlaim [382]. METALA [780]. Metalinguistic [2880]. Metamodel [1439]. MetaObject [166]. Metaphor [689]. Metaprogramming [173]. Metering [570]. Method [926, 498, 2470, 320, 326, 1584, 982, 2681, 2473, 822, 860, 825, 541, 807, 2423, 2385, 347, 217, 2036, 120, 2278, 2901, 2239, 881, 2857, 815, 1511, 1683, 2576, 1642, 2762, 340, 844, 1513, 2257, 1443, 817, 1115, 1551, 549, 1198, 1022, 2266, 2800, 2644, 50, 2890, 892, 2031, 2725]. Methodological [637, 2389]. Methodology [851, 847, 1970, 512, 2852, 1648, 1137, 754, 380, 866]. Methods [1003, 1601, 493, 2782, 2543, 2864, 847, 1586, 1001, 2413, 2321, 2106, 2855, 488, 1060, 999, 1763, 445, 1903, 53, 2328]. Metric [925, 1189, 2647, 2627]. Metrics [1341, 31, 314, 1587]. Microaggregation [2901]. Microbiological [513]. Microcalcifications [511, 890]. Microcontroller [714, 730]. Microcontroller-Based [714]. Microcontrollers [2776]. Micropayments [2718]. Microprocessor [2697]. Microscope [992]. Microscopy [508]. Microsurgery [831, 859]. Microsurgical [1069]. Middleware [645, 1278, 1806, 174, 165, 1277, 1947, 2449, 1281, 1275, 1657, 1285, 1273, 1142, 1288]. Middlewares [1960]. Midpoint [1107]. Migration [1278, 1941, 1944, 196, 1339]. MIKAS [2479]. Min [1353]. Min-Max [1353]. Mind [78]. Miner [2830]. Miniaturized [887]. Minimal [2908, 2185, 2767, 632, 1357, 1200]. Minimally [955, 858, 1001, 1025]. Minimally-Invasive [1001]. Minimisation [1185]. Minimization [2475, 984]. Minimize [1375, 1442]. Minimizing [1345, 1467]. Minimum [1860, 1539, 655, 1358, 772, 2076, 1354, 557].

Minimum-Degree [2076]. Mining [472, 49, 2542, 2384, 2402, 728, 2396, 1502, 1608, 476, 2360, 523, 1464, 1490, 1518, 1077, 471, 2369, 2512, 2406]. Minority [474]. MinRank [2154]. Minsky [2464]. MIR [351]. Misclassification [2471]. Mismatch [225, 2456, 369]. Misregistrations [964]. Miss [1599]. Missing [1439, 283]. Misuse [1153]. Mitkov [2857]. Mitosis [1028]. Mixed [385, 362, 2420, 1950, 393, 1847, 301]. Mixed-Mode [362]. Mixed-Reality [393]. Mixture [36, 2364, 306, 970, 2259, 1529]. Mixture-of-Experts [36]. Mixture-State [2259]. MLD [1570]. MLP [2815]. Mobile [616, 1936, 622, 639, 1947, 2099, 1256, 1940, 1933, 772, 198, 2432, 1946, 1941, 1934, 595, 1169, 1944, 1603, 2774, 2634, 196, 460, 1383, 1318, 1274, 82, 1267, 154, 1661, 461, 617, 2433, 1931, 1942, 1248, 1762, 1938, 588, 1932, 162, 1945, 1655, 1727, 448]. Mobile-Agent [1947, 1946]. Mobility [1943, 608, 1208, 1177]. Modal [633, 2898, 2308, 2302, 1067, 967, 1141, 2430, 290]. Modality [2293, 1209, 275]. Modally [1006]. Mode [362, 2146, 1280, 1703, 1410]. Model [2726, 494, 733, 1843, 1766, 1377, 1686, 1786, 2208, 2184, 957, 1033, 1237, 2065, 1123, 1729, 1732, 870, 245, 1687, 2303, 356, 1448, 1277, 1688, 658, 916, 1659, 1217, 106, 2070, 2738, 2696, 2797, 1048, 2185, 83, 62, 2803, 1417, 1031, 96, 2779, 1651, 545, 2898, 2356, 2775, 24, 1913, 929, 86, 2364, 2199, 1575, 1437, 2484, 1610, 1645, 510, 2189, 1327, 505, 1614, 1017, 476, 2373, 1506, 1681, 881, 1319, 2171, 2827, 2549, 2318, 2405, 520, 1921, 2798, 1430, 849, 69, 1587, 1968, 147, 1320, 841, 1927, 2009, 440, 2874, 960, 1689, 791, 2617, 531, 1035, 913, 1181]. Model [1034, 965, 352, 560, 1529, 60, 437, 866, 2221, 1662, 2511, 1647, 280, 34, 1530, 2664, 161, 178, 813]. Model-Based [1686, 96, 1610, 1327, 1681, 1319, 1921, 849, 1181]. Model-Checking [1217]. Model-Class [2696]. Model-Discovery [1506].

Modeling [1445, 1118, 1678, 1119, 1981, 2892, 1670, 1003, 1492, 27, 874, 501, 2536, 1412, 962, 1409, 2517, 504, 1891, 1457, 2312, 260, 1414, 882, 2522, 483, 1424, 463, 2038, 1442, 869, 237, 345, 248, 1421, 1671, 1685, 1027, 1444, 1413, 1432, 1416, 1143, 970, 1447, 1423, 2565, 234, 766, 2796]. Modelings [911]. Modelisation [1216]. Modelling [1733, 1451, 2027, 2010, 646, 770, 2561, 1325, 2434, 1912, 1675, 24, 54, 1437, 2699, 1674, 1321, 2507, 2031, 1422, 2096, 1322]. Models [609, 110, 884, 1735, 2799, 1389, 940, 789, 944, 1873, 1438, 822, 938, 1326, 502, 2770, 144, 2249, 2776, 1363, 134, 1046, 1131, 424, 1437, 1836, 9, 37, 38, 306, 1327, 1378, 1086, 2659, 1323, 2195, 1952, 1594, 2497, 824, 1515, 63, 1925, 2499, 865, 2378, 1895, 2259, 1426, 1146, 2814, 1520, 2045, 1672, 868, 1050, 2426, 2553]. Modes [2143, 1020]. Modification [2026]. Modified [1399, 2767, 2757, 2567, 249, 773]. Modifiers [740]. Modula [2055]. Modula-2 [2055]. Modular [2133, 743, 2621, 28, 2102, 1070, 2821, 2652, 2653, 2654, 2710]. Modulation [266, 1573, 2250]. Modules [1670, 2893, 1558]. Moduli [2721, 2710]. Mojave [2650, 2651]. Molecular [649]. MolèNE [163]. Moments [147]. Monitor [1660, 981]. Monitoring [1154, 119, 857, 807, 728, 1164, 2278, 798, 2578, 1702, 1249, 1252, 1972, 256]. Monitors [1632]. Monodic [2180]. Monoids [2911]. Monomials [2585]. Monomodal [1008]. Monotonic [2216]. Monotonicity [1382, 665, 1781]. Montgomery [2711, 2710, 2601]. MOP [164]. Morphing [2304]. Morphogenesis [1046]. Morphogenesis-Based [1046]. Morphological [2481, 2001, 533]. Morphologically [40]. Morphology [818, 1047, 2875]. Morphometry [820]. Mortality [502]. Mortem [928]. MOS [756]. Mosaic [123]. MosaicPlasty [886]. Mosaics [240]. Most [1568]. Motion [1678, 1033, 228, 106, 268, 333, 1030, 2480, 920, 2242, 1680, 2824, 815, 1029, 271, 122, 852, 862, 1026, 990, 332, 230, 1008]. Motion-Based [228]. Motion-Compensated [2242]. Motion-Robust [862]. Motor [717, 2683]. Mounted [899]. Movement [267, 92, 692]. Movie [239, 299]. Moving [604, 241, 2240, 123, 1343]. MPEG [1239, 2301, 228, 365, 262, 330, 225, 1238, 229, 297, 1233, 231, 364]. MPEG-2 [225]. MPEG-4 [1239, 2301, 365, 330, 225, 229, 1233, 231]. MPEG-7 [228]. MPI [1960, 1582]. MR [1678, 928, 968, 1011, 1062, 721, 872, 972, 943, 1681, 881, 837, 903, 911, 988, 875, 913, 953]. MR-Anatomical [1681]. MR-Tagging [1678]. MRA [873, 898]. MRF [264]. MRI [923, 1059, 902, 978, 847, 1031, 1013, 1047, 956, 908, 900, 1051, 1016, 984, 1029, 919, 1026, 990, 1061]. MRI-Ultrasound [923]. MRI/SPECT [847]. MRIs [1007]. MRM [2644]. MS [989]. MSC [2798]. MT [2858]. MTEACH [348]. mu [1224]. Much [761]. Mucosa [718]. Muller [1541, 2587, 1534]. Multi [745, 758, 2626, 2302, 1168, 646, 1735, 1067, 1850, 778, 750, 969, 2281, 1099, 1770, 2337, 2315, 2352, 1983, 2782, 2559, 640, 2542, 687, 21, 2347, 793, 967, 1466, 746, 101, 2307, 1346, 2308, 1141, 2883, 2744, 2354, 2568, 1660, 2489, 2430, 2562, 2095, 185, 1839, 2569, 399, 2594, 2780, 887, 104, 1014, 2563, 128, 376, 883, 133, 235, 638, 2031, 305, 290, 1948, 1636, 1104, 275, 227]. Multi-adjoint [1839, 2569]. Multi-agent [646, 969, 1770, 2337, 2559, 640, 687, 2307, 2308, 2354, 2562, 2095, 2563, 638, 1104]. Multi-agents [2352]. Multi-choice [1850]. Multi-class [128, 133]. Multi-complexes [1983]. Multi-criteria [1735, 2782]. Multi-cultural [2489]. Multi-dimensional [2568, 1014]. Multi-document [2883]. Multi-DOF [887]. Multi-exponentiation [2594]. Multi-functional [2626].

Multi-homed [1168]. Multi-hop [305]. Multi-layer [2281]. Multi-layered [1099, 399]. Multi-level [883]. Multi-marking [376]. Multi-modal [2302, 1067, 967, 1141, 2430, 290]. Multi-modality [275]. Multi-national [21]. Multi-objective [778, 750, 2347, 793, 2780]. Multi-output [1636]. Multi-platform [185]. Multi-processes [1660]. Multi-purpose [746]. Multi-recipient [2744]. Multi-reference [227]. Multi-relational [2542]. Multi-resolution [104]. Multi-task [1948]. Multi-threshold [101]. Multi-tier [2031]. Multi-user [2315]. Multi-users [235]. Multi-valued [745, 758]. Multi-vehicle [1346, 1141]. Multi-way [1466]. Multiagent [2474, 1725]. Multiattribute [55]. Multiband [2808]. Multibrowsing [604]. Multicast [1752, 1745, 1172, 1746, 540, 1174, 1751, 304, 1740, 1252, 1624, 811, 1749, 1743]. Multicasting [1351]. Multicomputer [1579]. Multicounter [2068]. Multicriterially [1500]. Multidimensional [1734, 35, 47]. Multidoubling [2602]. Multifactorial [2820]. Multifunctions [1980]. Multigrid [1994]. Multiheaded [1829]. Multilayered [2848]. Multilevel [1651, 1456, 1625, 828]. Multilingual [2867, 449, 455, 2551, 2886]. Multimedia [1239, 1172, 349, 325, 319, 258, 1174, 259, 1254, 1169, 292, 1235, 255, 1175, 1452, 351, 237, 348, 287, 359, 303, 354, 352, 560, 234, 1180, 284, 2040, 2516]. Multimedia-Based [255]. Multimedia-on-Demand [349]. Multimodal [882, 956, 1044]. Multimodality [964]. Multiplayer [1739]. Multiple [2896, 1140, 1438, 2417, 1869, 127, 1472, 738, 1449, 666, 331, 1638, 604, 1242, 2598, 432, 1444, 1430, 153, 2644, 907, 963]. Multiple-Instance [1472]. Multiple-Person [331]. Multiple-Record [1449]. Multiple-Valued [738]. Multiples [2103]. Multiplication [2601, 1814, 2760, 2600, 1576, 559, 1633, 2763, 2762, 2599, 2602, 2710, 1621]. Multiplicative [2195]. Multiply [65]. Multiprocess [1393]. Multiprocessing [396]. Multiprocessor [1581, 1350, 1333]. Multirate [1746]. Multireceiver [2745]. Multiresolution [1058, 2257]. Multiresolutional [340]. Multiscale [2267]. Multisensorial [722]. Multisequence [1558]. Multiset [1787, 1788, 1790, 1792, 1797, 1197, 1799]. Multisets [1793, 1794, 1795, 2363, 1543, 1800, 1801]. Multispectral [941]. Multistage [776]. Multisubject [902]. Multitask [524]. Multithreading [1591]. Multiuser [773]. Multivariate [2755, 1553, 2244, 54, 58, 1529]. Multiwavelet [2236]. Multiwavelets [2251]. Multiway [1400]. Multiword [2870, 2843]. Muscle [957, 1687]. Music [89, 289, 296, 329]. Mutagenes [1532]. Mutagenicity [1498]. Mutation [2678]. Mutual [1041, 880, 998, 2094, 2564, 2351]. Mutually [2153, 2147]. mVITAL [81]. Myocardial [1682, 1680, 1684, 1683, 988]. Myocardium [1688, 987, 1689].

N [940, 1315, 2885]. N-Grams [2885]. Naive [2467, 2318]. Names [1212, 1293]. Napster [2326]. Narrative [394, 396, 395, 404, 393, 2040]. Narrowing [634]. NAS [344]. NASA [2457]. National [2004, 21]. Natural [2872, 2428, 2850, 1819, 1196, 1442]. Naturalistic [2293]. Navigation [690, 956, 1032, 1509, 707, 693]. Navigational [577]. Navigator [990]. NCP [1630]. Near [729, 981, 1107]. Nearest [2546]. Nearness [684, 786]. Necessary [1360]. Neck [2843]. Need [708]. Needle [929, 950, 843]. Needle-Based [950]. Needs [1502, 2436]. Negaperiodic [1554]. Negation [2212]. Negative [2065, 1494, 1841, 1923]. Negotiating [1756]. Negotiation [1428, 2313, 571, 2295, 1913, 2632, 2564, 643]. Negotiations [2314]. Neighbor [2546, 1479]. Neighborhood [1885]. Neighborhood-Based [1885]. Neonatal [503]. NESSIE [2761, 556]. Nested [1837, 2571, 2221]. Net [2922, 2921, 2029, 739, 515, 2031]. Net-Based [2031]. NETLA [2800]. NetPeas [485]. Nets [743, 2308, 2030, 788, 803, 2080, 2195, 755, 2923, 1322]. Network [1752, 810, 2799, 1859, 723, 808, 2801, 804, 1649, 365, 1255, 1085, 1126, 807, 1651, 2486, 1281, 572, 463, 434, 1245, 813, 2318, 287, 2798, 1092, 2499, 1624, 354, 801, 2800, 2292, 716, 811, 1146, 2805, 1160, 536, 2796, 1662, 301, 1251, 2319, 34, 787]. Network-Adaptive [301]. Network-Centric [1126]. Networked [578, 431, 1283, 1938]. Networking [805, 1228, 809]. Networks [2462, 494, 1752, 745, 1405, 725, 1389, 769, 491, 423, 1577, 201, 581, 770, 1602, 724, 2311, 2828, 1589, 1726, 1351, 2848, 65, 1235, 1575, 1250, 2808, 717, 768, 2550, 650, 798, 732, 415, 1740, 1243, 1249, 1248, 303, 1605, 2378, 791, 66, 1244, 47]. Neural [2462, 745, 733, 2799, 769, 491, 2801, 2152, 365, 770, 2790, 724, 2311, 743, 1262, 2486, 788, 803, 2808, 515, 717, 2550, 757, 798, 732, 2798, 755, 1092, 2800, 2805, 2319, 47, 787]. Neuro [723, 2802, 2794, 2803, 460, 2804]. Neuro-fuzzy [723, 2802, 2803, 2804]. Neuroendocrinology [1506]. Neuroimage [893]. Neuron [1467, 756]. Neuron-MOS [756]. Neuronal [2795]. Neurons [745, 758]. Neuropharmacology [1530]. Neuropsychological [702]. Neurosurgery [1067, 882, 899]. News [1497, 328, 277, 322]. Newsletters [776]. Next [1239, 1720, 1589, 414]. Next-Generation [1589]. NIDS [1652]. NIRS [729]. NLC [669]. NLP [2848, 642, 2843]. NMR [1683]. No [708, 125]. Node [808]. Nodes [1589]. Nodule [827]. Nodules [1056]. Noise [514, 1996, 2496, 2266, 2823, 2811]. Noisy [539, 2808, 105, 2504]. Nomadic [1272]. Nominal [2385, 58, 1212]. Nominative [1661]. Non [923, 1872, 1871, 159, 1010, 902, 2258, 927, 1853, 2546, 2714, 2534, 2797, 821, 1175, 863, 920, 2646, 2271, 2273, 1857, 562, 1006, 1867, 1027, 751, 1219, 2048, 1555, 2880, 828, 2795, 824, 2366, 1924, 883, 1004, 2370, 1044, 1474]. Non-additive [2370]. Non-binary [2646, 1857, 1867, 1924]. Non-cooperative [1175]. Non-CT [863]. Non-destructive [751]. Non-deterministic [2534, 2048, 2366]. Non-equilibrium [2795]. Non-existence [1853, 1555]. Non-functional [159]. Non-interactive [2714]. Non-linear [923, 927, 2797, 1004, 1474]. Non-overlapping [1872, 1871]. Non-parametric [2546]. Non-repudiation [562]. Non-rigid [1010, 902, 821, 920, 1006, 828, 824, 883, 1044]. Non-separable [2258]. Non-standard [2880]. Non-structural [1219]. Non-uniform [2271, 2273, 883]. Nondeterminism [2067]. Nondeterministic [2916]. Nonlinear [2349, 2115, 131, 2121, 2256, 2776, 2647, 875, 2796]. Nonlinearity [2586]. Nonmetric [47]. Nonmonotonic [2207]. Nonrigid [924, 846]. Nonsymmetric [1584]. Nonuniform [2727]. Norm [2343]. Normal [2572, 1840, 2773, 1792, 1633, 1838, 2599, 2464, 894, 1347]. Normalization [2303, 132, 1224]. normative [1304, 1305, 1306, 1308, 1309, 1310, 1311, 1307, 1303]. Norms [794, 2351]. North [706]. Notation [1320]. Notations [2656]. Note [2383, 2382]. Notes [2807, 2684, 531]. Notification [595, 1741]. Notion [2915]. Novel [627, 316, 2249, 2732, 217, 509, 900, 1642, 924, 849, 859, 278, 1247]. Nowhere [1362]. NP [2058, 660, 676]. NP-Completeness [676]. NP-Hard [2058, 660]. NSS [2131]. NT [1076]. NTRU [2131]. Nuclear [1503, 1028]. Nucleus [968]. Number [1860, 2133, 1547, 2190, 626, 2835, 2023, 1355, 2082, 1561, 1537]. Number-Theoretic [1537]. Numbers [565, 1085, 1553, 2643, 1881, 1360]. Numerals [2799]. Numeric [58, 1928]. Numerical [2385, 2224, 2248]. Numerics [1309]. NURBS [2312]. Nursing [2318].

O [1682, 1596]. O-Efficient [2077]. Oberon [2055]. Oberon-2 [2055]. Obfuscation [547]. Object [1278, 173, 362, 358, 1448, 127, 136, 1270, 2924, 321, 2029, 876, 2828, 1287, 1456, 1226, 126, 2310, 2240, 140, 1702, 1430, 129, 1286, 115, 139, 234, 416, 109]. Object-Based [362, 2029]. Object-Oriented [1448, 2924]. Objective [1052, 2092, 778, 750, 2347, 793, 2780]. Objective-Trust [2092]. Objects [1445, 895, 2686, 2521, 622, 834, 1084, 330, 709, 1417, 1456, 2522, 241, 711, 1623, 123, 2570, 917, 151, 706, 1999, 243]. Obligation [1632]. Oblivious [2152, 2588, 566, 2751, 1616]. Obscurity [1668]. Observability [2709]. Observable [1111]. Observation [2467]. Observational [1208, 2024]. Observations [1900]. Observed [2394]. Obsolescent [1311]. Obstacle [146]. Occlusion [115]. Occlusions [129]. Occupation [286]. OCL [2010, 2669, 2670, 2668, 2667, 2673, 2666, 2672, 2665, 2675]. OCL-Syntax [2672]. OCR [2820]. Octopus [813]. Octrees [1586]. Odd [2585]. Off [2439, 1010, 2459]. Off-The-Shelf [2459]. Office [1282]. Offline [1664]. OLAP [1432]. Omni [235]. Omni-directional [235]. Omniscience [2186]. On-Line [1401, 2802, 1375, 1351, 2260, 781, 2261].

One [1768, 113, 2714, 1880, 2765, 2757, 1366, 2113, 2079, 2088, 615, 1551, 1252, 643]. One-Point [1551]. One-Sided [113]. One-Time [2113]. One-to-Many [643]. One-to-One [643]. One-Way [1768, 2714, 2765, 1366, 615]. One-Wayness [2757]. Online [1850, 135, 100, 631, 2749, 474, 425, 436]. Only [1815, 2231]. Onset [535]. onto [1902]. Ontological [2489, 711, 414]. Ontologically [1435]. Ontologies [2473, 700, 469, 477]. Ontology [462, 1449, 518, 2494, 475, 457]. Ontology-Applicable [1449]. Ontology-Based [462]. OOCSMP [90]. OOM [2322]. Open [408, 1601, 1102, 1177, 2095, 182, 178, 166, 157, 215, 638, 1182, 1727]. Open-Environment [1102]. Opening [813]. OpenORB [174]. Operating [396, 992, 1283, 1597, 974]. Operation [2912, 2143, 2091]. Operational [2211, 1703, 2510]. Operations [1190, 1259, 626, 364]. operative [923, 861, 2562, 952]. operatively [2722]. Operator [2010, 1825]. Operators [1892, 2001]. Ophthalmic [1057]. Ophthalmologic [870]. Opinion [2338]. Opportunistic [606]. Opportunities [73, 1145]. Optic [533, 954]. Optical [891, 1609, 928, 1351, 1116, 1605]. Optimal [2064, 2107, 2742, 822, 136, 2577, 1546, 2072, 108, 1882, 226, 107, 2678, 2809, 2580, 1072, 864, 354, 2119, 2800, 212, 2011, 1897]. Optimisation [1883, 2468, 2486, 771]. Optimistic [2374, 564, 562, 1597]. Optimization [1951, 2686, 778, 2807, 750, 1595, 1869, 1954, 1909, 383, 1383, 1392, 1108, 1902, 399, 1867, 2145, 143, 2305, 947, 749]. Optimizations [1391, 1135]. Optimize [714]. Optimizer [1005]. Optimizing [2314, 2556, 1826, 754]. Optimum [2828, 1917]. Options [1271]. Or-Parallel [1823]. Or-Parallelism [1824]. Oracle [2726]. Oral [904, 718, 522]. Oranges [2530].

Orbit [1693]. Order

[2200, 2063, 2107, 2216, 2695, 1853, 2180, 1498, 1841, 2308, 2620, 2202, 2523, 2126, 1771, 1212, 2197, 2537, 1224]. Order-Sorted [1841]. Ordered [1909, 1363, 676]. Ordering [1885, 2049, 2216, 2369]. Orderings [683]. Orders [1215]. Ordinary [2214]. ordination [2561]. Organ [930, 1008]. Organisation [503]. Organisms [823]. Organization [484, 1674, 2579, 2796]. Organizational [2423, 187]. Organizations [1754, 639, 2413]. Organized [1107]. Organizing [130]. Organs [961, 1035]. Orientation [107, 1370, 2570, 697]. Oriented [20, 6, 169, 1448, 186, 2924, 1285, 2210, 2298, 642, 156, 1447, 421, 374, 2329, 90]. Origin [2844, 93]. Oropharyngeal [522]. Orthogonal [148, 2258, 1853, 661, 1775, 1026, 2251]. Osteotomy [871]. Other [2543, 2008, 711, 1220, 1304]. Otoneurological [532]. Ought [397]. OutCast [603]. Outcomes [504]. Outdoor [727]. Outerplanar [664]. Outliers [514]. Output [2072, 2614, 1198, 1636]. Output-Sensitive [2072]. Outs [1343]. Over-Constrained [1876, 1868]. Overconstrained [2465]. Overhead [152]. Overlapping [99, 1872, 1871]. Overlay [1744]. Overoptimism [528]. Overview [82, 2297, 2168, 2169, 2342, 2332, 256].

P [2005, 2004, 2105, 2927, 1370, 1316]. Paced [1690]. Packages [1296]. Packet [806, 252, 1348, 2250, 270, 1607, 2283]. Packets [2254, 2233]. Packing [1401, 2686, 1590]. PACS [522]. Padding [2756]. Page [444, 466, 445]. Pages [432]. Pain [2843]. Pair [2102]. Pairing [2160, 2161]. Pairs [2593, 659, 2882]. PALS [1823]. Panoramas [973]. Papous [85]. Paraconsistent [2892]. Paradigm [1956, 2006, 289, 2819, 722]. Paradigms [623, 468]. Paradox [2464]. Parallel [1592, 2544, 1584, 450, 1106, 1341, 1105, 2072, 1586, 2760, 1608, 1955, 1583, 2325, 1902, 2101, 2580, 1338, 1597, 1823, 1655, 1339, 2319, 1386]. Parallelism [1607, 1824]. Parallelized [2827]. Parameter [925, 136, 1834, 2282, 771]. Parameterizations [2285]. Parameterized [1368, 739]. Parameters [1034, 2051, 797]. Parametric [884, 870, 1882, 1578, 1973, 2546]. Parametrized [2704]. Parchments [2693]. Pareto [2462]. Parikh [1789]. Parity [2636]. Parkinson [968, 519]. Parser [2532]. Parsing [2862, 1787, 1819]. Part [2863, 1439, 2864, 2494, 1413, 55]. Part-of [2494]. Part-of-Speech [2863, 2864, 55]. Part-Whole [1413]. Partial [379, 2470, 611, 1388, 629, 177, 1014, 2204, 1925, 1516, 839]. Partially [1560, 1909, 1238, 2747, 676, 1111]. Partially-Ordered [676]. Participants [588]. Particle [1504]. Particles [1035]. Partition [1884, 2771, 1400, 1884]. Partition-k-AC [1884]. Partitioning [142]. Partitions [681, 1985]. Partner [247]. Parts [1439, 2640, 1499]. Party [564, 562, 1664]. Pass [2270, 898]. Passage [1502]. Passage-Based [1502]. Passed [2268]. Passing [1206]. Past [217, 696]. Pastry [1286]. Patching [285]. Path [1405, 1825, 2216, 1385, 1346, 696, 1051, 1518, 835]. Pathological [1058, 893]. Paths [1769, 659, 1372, 772, 1923, 1357]. Pathways [1503]. Pathwidth [664]. Patient [928, 944, 833, 958, 931, 521, 868]. Patient-Specific [833, 958, 931]. Patients [894]. Pattern [1373, 2797, 1585, 1403, 510, 2756, 74, 1258, 2360, 524, 133, 2644, 1528, 116, 2514]. Pattern-Based [1258]. Patterns [745, 2231, 705, 1324, 28, 2423, 1523, 1456, 2492, 2398, 521, 1109, 2917, 734, 2889, 2377, 1077, 66]. Pawlak [2907]. Payment [568].

Payments [1727]. PC [447]. PDA [594]. Pebbling [1779]. Pedagogical [413]. Peeling [1381]. Peer [644, 189, 1742, 1726, 1941, 1286]. Peer-to-Peer [644, 189, 1742, 1726, 1941, 1286]. Pelvic [946]. Pelvis [994]. Penetration [929]. People [391]. Per-Packet [806]. Perception [401]. Perceptions [775]. Perceptrons [767]. Perceptual [142, 2492, 231]. Percutaneous [844, 1004]. Perfect [2750, 2271, 2273, 1555, 1374, 677]. Performance [2276, 2466, 1069, 174, 1335, 2471, 1962, 2815, 1595, 1331, 1960, 1947, 1604, 1946, 1631, 344, 407, 1603, 2526, 2477, 1250, 1611, 2624, 1965, 1587, 1588, 263, 1338, 949, 1939, 2882, 1972, 370]. Performances [393]. Perfusion [1682, 988, 981]. Periaxial [817]. Periodized [2279]. Periods [1556]. Peripheral [2481, 603]. Permissive [2573]. Permutation [2643, 1895, 2205]. Permutations [1373]. Persistent [563]. Person [331]. Personal [220, 2302, 2231, 743, 92, 438, 605, 428]. Personalisation [2297]. Personalization [2296, 2298, 776, 2294]. Personalized [2299, 433, 952]. Perspective [1463, 1488, 1602, 2769, 2542, 637, 336, 1444, 1698]. Perspectives [2730, 1125]. Perturbation [909]. Perturbations [1003]. Pervasive [2424, 195]. PET [1682, 1681, 881, 980]. PET-Functional [1681]. Petri [2922, 2029, 739, 2308, 2030, 2080, 2923, 1322]. Phantom [959, 953]. Phantoms [843]. Phase [1874, 2258, 1031, 1875, 1107, 936, 991, 1026, 1856]. Phase-Based [936]. Phase-Driven [1031]. Phased [2468]. Phe [2491]. Phe-Q [2491]. Phenomenology [141]. Phenomenology-Based [141]. Pheromone [2491]. Philosophy [1463, 1488]. Phone [218]. Photo [247]. Photo-Realistic [247]. Photoelectrons [1529]. Photograph [326]. Photographs [932, 976]. Photon [985]. Phrasing [356].

Physical [926, 789, 1783]. Physically [965]. Physically-Based [965]. Physics [2467, 1504, 958, 2636]. Physics-Based [958]. Physiological [499, 852]. Physiologically [1687]. Physiologically-Based [1687]. Physiology [980]. Pictorial [2816]. Pictures [2000]. Piecewise [2628, 2279]. PING [206]. Pioneer [2004]. Pipelined [1767, 2697]. Pitch [281]. Pitfall [2424]. Pixel [230]. PKC [2757]. Place [700]. Placement [1072, 1605, 1878]. Placing [1105, 151]. Plaintext [2757, 2717]. PLAN [812]. Planar [1371, 667, 1977]. Planarity [658]. Plane [926, 2248, 2243, 2002]. Planes [108]. Planetary [576]. Planned [952]. Planner [2580]. Planning [867, 2577, 1032, 998, 137, 858, 1000, 2824, 2229, 1952, 2579, 961, 999, 1024, 2504, 1905, 88, 868, 871]. Plants [798]. Plasma [1583]. Plastic [2425]. Plasticity [1320]. Plate [106]. Platform [1075, 500, 159, 1944, 214, 1166, 88, 185]. Plausible [2517]. Play [2320]. Player [2478, 2119]. Playing [217, 2269, 1634]. Plays [86]. Plot [1854]. Pneumatic [1068]. Pockets [1446]. Poems [1521]. Point [110, 884, 1987, 1395, 106, 2158, 904, 2914, 1984, 2592, 2211, 551, 1881, 559, 1209, 2570, 2763, 905, 1551, 1980, 1900, 116]. Point-Counting [2158]. Point-Surface [904]. Point-to-Line [1987]. Pointer [2016]. Pointing [152, 2429]. Points [2831, 1415, 108, 1836, 107, 768, 1394, 374]. Poker [2478]. Policies [2895, 2234]. Policy [1254, 1971]. Policy-Based [1254]. Politeness [2874]. Pollard [2598]. Poly [755]. Poly-crystalline [755]. Polya [2349]. Polygonal [245, 1383, 1587, 1343]. Polygons [1992, 1343]. Polyhedra [1998]. Polyhedrization [1984]. Polyhedron [1358]. Polyline [966]. Polymorphic [2022, 1113]. Polymorphically [2198]. Polynomial [2152, 2727, 1345, 660, 1218, 1372, 1346, 2591, 1366, 2627, 1798, 2082, 69].

Polynomial-Time [1366]. Polynomially [2008]. Polynomials [2103, 1615, 1553, 2104, 58]. Polyominoes [1775]. Polyp [891, 825]. Polypectomy [1037]. Polyproline [530]. Polysomnographic [2265]. Polytopes [1872]. Pool [873, 898]. Poor [2857]. Population [1075, 759]. Port [1072]. Portable [1945]. Portal [878]. Pos [2043, 2865]. Pose [355, 135, 245, 151]. Positioning [580]. Positive [1829, 573]. Possessing [2635]. Possibilities [204, 796]. Possibility [738, 2361, 736]. Post [928, 1688, 2913]. Post-Systolic [1688]. Postdiction [2472]. Postmortem [17]. Pot [2932]. Potential [1754, 395, 2451]. Potentially [2710]. Potpourri [412]. Power [1886, 558, 799, 411, 2602, 2759, 2009, 2789, 2712, 2147, 2615]. Powered [804, 600]. Powers [651]. Practical [810, 1237, 1746, 548, 540, 1602, 724, 2156, 2583, 1523, 515, 1828]. Practice [2711, 2032, 1434, 1530]. Practiced [2413]. Practices [1805, 2442]. Pragmas [1313]. Pragmatic [1432]. Pragmatism [691]. Pre [2419, 2010]. Pre-implementation [2419]. Precedence [2079, 1579]. Precise [1149, 2667, 1599, 2710, 866]. Precision [1830, 476, 2373]. Precolored [625]. Preconditioning [2236]. Predefined [2657, 1303]. Predicate [204]. Predicates [1941, 1418]. Predicting [1003, 506, 37, 1972]. Prediction [2555, 1593, 792, 2242, 762, 999, 530, 55, 2284, 227]. Predictive [339, 1474]. Predictors [1592]. Prefer [2560]. Preference [480, 2554]. Preference-Specific [480]. Preferences [2295]. Preferred [1836]. Preferring [640]. Prefetching [285, 471]. Prefixes [1485]. Preliminary [968, 54, 38, 817]. Premature [432]. Prenex [2193]. Preoperative [943]. Prepositions [696]. Preprocessing [2240]. Presence [2897, 83, 103]. Present [1121]. Presentation [649, 1979, 650, 2672].

Presentations [325, 2703]. Preservation [2899]. Preserving [779, 838, 829, 1285, 1771, 2537, 906]. Pressure [848]. Prevention [2396, 574, 2105]. Price [2748, 1257, 1955]. Primal [921]. Prime [2107, 2136]. Primitive [2103, 1615, 2741]. Primitives [2156]. Principal [459, 310, 1240]. Principle [2899]. Principles [598, 2360, 1700, 2346]. Printed [265]. Priorities [2026]. Priority [2374]. Priors [2678]. Privacy [544, 1635, 2163, 545, 1162, 539, 1730, 598, 2901]. Privacy-Aware [598]. Privacy-Enhanced [2901]. Privilege [1159]. Probabilistic [1768, 423, 578, 724, 2835, 128, 138, 1821, 1156, 2514, 109]. Probability [738, 799, 753, 2368, 2590, 2611, 2370, 1570]. Probable [2136]. Probe [848]. Probing [101]. Problem [1874, 1405, 2466, 763, 1885, 2133, 2281, 2727, 2154, 715, 1888, 2715, 1399, 1861, 668, 2913, 2914, 1346, 1875, 1384, 2076, 2354, 1344, 1650, 1798, 2576, 1512, 674, 143, 1359, 2248, 1561, 1878, 2506, 747, 1104, 2613]. Problems [1765, 2058, 627, 1883, 1401, 2465, 1906, 2442, 651, 1859, 1601, 1755, 2472, 1125, 1898, 1376, 660, 1889, 1909, 1564, 1380, 1912, 1390, 614, 1407, 2077, 1867, 2781, 1876, 2082, 1952, 1868, 128, 1356, 1111, 376, 133, 1926, 1895, 2205, 747, 1727, 1398, 1400]. Procedural [2420, 2569]. Procedure [2044, 2204]. Procedures [882, 858, 950, 854, 844, 2366]. Process [2439, 20, 6, 169, 26, 2349, 2090, 119, 27, 778, 183, 921, 2446, 2474, 1682, 28, 2449, 2450, 690, 1204, 728, 24, 1390, 2416, 255, 1107, 15, 1718, 10, 9, 2452, 7, 798, 1714, 911, 1905, 23]. Process-Algebraic [169]. Process-Monitoring [728]. Processes [1770, 789, 1409, 2921, 2925, 19, 2488, 1028, 2413, 16, 1782, 190, 25, 1660]. Processing [758, 703, 1057, 1375, 806, 704, 2849, 614, 1058, 508, 2481, 782, 1902, 414, 2169, 1215, 1921, 1705, 1197, 2795, 2819, 722, 1079, 346, 1076, 1196]. Processor [1149, 2386, 1594]. Processors [1592, 1591, 1135, 1579, 2599, 1146, 1897]. Procurement [2454]. Product [6, 1810, 430, 2446, 1807, 2805]. Production [1340, 739, 2682, 348, 1952, 1973]. Production-Level [1973]. Products [2439, 1563, 2445, 49, 487]. Professionals [4]. Profile [309]. Profiles [2371, 1667]. Profiling [1594, 948]. Program [2017, 1222, 30, 1816, 2020, 383, 2005, 1594, 2009, 1299, 158, 2515, 322]. Programmable [807, 2075, 1093, 813, 1134]. Programming [169, 1883, 1842, 1832, 1850, 1121, 1859, 2541, 1412, 1732, 1886, 372, 1106, 1846, 389, 1105, 2218, 1861, 1844, 1833, 2525, 1213, 2020, 1129, 1914, 2532, 748, 1977, 2568, 2523, 2210, 2373, 2835, 1950, 2569, 1919, 2781, 760, 2021, 2926, 2404, 160, 1866, 156, 1952, 2004, 1923, 2535, 2501, 2500, 1821, 1305, 1130, 1847, 2524, 1206, 1929, 374, 877, 1150, 1104, 23, 2061, 382, 373]. Programs [2209, 1227, 30, 2043, 547, 2042, 1835, 1840, 1837, 1908, 2534, 2226, 1645, 1258, 1834, 2048, 1838, 177, 2571, 2704, 2527, 2025, 2187, 376, 2680, 761, 1925, 2198, 583, 1079, 380, 2538, 2215]. Progress [1512, 2706, 2451, 2344]. Project [1460, 1487, 388, 1085, 22, 2346, 1181]. Projected [1049]. Projection [1951, 2516]. Projections [1775]. Projective [1541, 2592, 1543]. Projector [602]. Projects [1804]. Prolog [2541, 1822, 2044, 2224, 2537, 1823, 1818]. Prone [297, 29]. Pronominal [2859, 2858]. Pronoun [2857]. Pronunciation [2876]. Proof [2726, 1214, 2189, 2229, 2195, 2201]. Proof-Search [1214]. Proofs [2746, 2190, 1210, 2214, 2659, 2704]. Propagation [725, 2208, 1892, 574, 1689, 915]. Properties [51, 2616, 159, 618, 2911, 1440, 2668, 1934, 1792, 2628, 2279, 2120, 2635, 2013, 2653, 2080]. Property [571, 930, 1422].

Proportional [2549]. Proposal [2263, 2453, 2397, 2856]. Proposed [2757, 2126, 2170]. Proposition [2191]. Propositional [2476, 1214, 2009, 2206]. Proprietary [2719]. Pros [150]. Prosodic [356]. Prospect [2292]. Prospects [2685]. Prostate [856, 993, 942, 864, 947, 996]. Prosthetic [1088]. Protecting [2615]. Protection [1635, 2099, 1560, 545, 1162, 1175, 1730, 1931, 312]. Protein [1377, 1878]. Protocol [1178, 1653, 349, 564, 1236, 1625, 1177, 2320, 1614, 562, 1744, 2132, 287, 2633, 1750, 2632, 568]. Protocols [1766, 183, 1865, 2714, 541, 2134, 2677, 2753, 2728, 1931, 166, 2119, 2114, 1361, 933]. Prototype [1660, 1904, 2177]. Prototypes [2831]. Prototyping [2419]. Provable [2146, 2202]. Provably [2164, 2148, 1643, 2575, 2747]. Proved [375]. Provider [2458]. Providers [1269]. Providing [1287, 1170, 1252]. Proving [2193, 1840]. Provision [1723]. Provisioning [1182]. Proxies [179]. Proximity [2077, 1356]. Proxy [417, 1274, 1661, 285, 421]. PRS [2510]. PRS-Like [2510]. Pruning [1860, 1871, 1882, 1916, 53, 2513]. PSD [2250]. Pseudo [2062, 2156, 134, 2628, 2127, 1357]. Pseudo- [134]. Pseudo-Euclidean [1357]. Pseudo-Random [2628, 2062, 2127]. Pseudo-Randomness [2156]. Pseudorandom [1553]. PSK [1573]. Psychological [2318]. Public [646, 2163, 550, 2338, 2603, 595, 2600, 2743, 2752, 2744, 2113, 2135, 2339, 1650, 11, 2741]. Public-Key [2163, 550, 2603, 2600, 2743, 2752, 2744, 2741]. Publish [190]. Publish/Subscribe [190]. Publishing [561, 418]. Pulmonary [827, 1056]. Pulsating [959]. Punch [140]. Punishment [2350]. Pupillary [1022]. Pure [2023]. Purified [1511]. Purpose [24, 2599, 746]. Purposes [389].

Pushbutton [600]. Pushdown [1217, 1364]. Putnam [2204]. Putting [1592]. PVA [953]. Pyramids [1982].

Q [317, 914, 2491]. QA [2883]. QoS

[1168, 1172, 1174, 806, 1083, 1287, 287, 339, 1249, 1590, 1167, 1288, 1183]. QoS-Awareness [1183]. QoS-Based [1083, 287]. QoS-Driven [1168]. QOSMIC [1167]. QTk [2420]. Quadrangulations [1395]. Quadratic [2641, 2755, 1467, 557]. Qualification [2506]. Qualitative [2467, 685, 683, 2570, 701, 707, 2787, 766]. Quality [1428, 350, 752, 1238, 1241, 1243, 1, 1231, 953, 2789, 5, 2512]. Quality-Driven [1428]. Quantification [916, 508, 2846, 2845, 1028, 836, 535]. Quantified [2476, 2204]. Quantifying [1649, 909, 821]. Quantitative [2868, 1679, 908, 2915, 943, 985, 1515, 2787]. Quantitative-Qualitative [2787]. Quantization [274]. Quantum [1768, 1773, 1366, 615, 765, 1356]. Quartet [1404]. Quasi [504, 1215]. Quasi-Fourier [504]. Quaternary [1555]. QuDAS [1083]. Queen [230]. Queries [1461, 1489, 1482, 319, 1472, 1431, 1465]. Query [2895, 250, 319, 474, 292, 289, 2900, 1215, 456, 1079, 279, 329]. Query-Session-Based [292]. Querying [689, 1844, 2041, 1450]. Queue [1393, 286, 1244]. Queued [1607]. Queueing [1575, 1244]. Queues [1392]. Queuing [1967, 1348]. Quorum [1391]. Quotation [489]. Quotients [1572].

Rabin [2619]. Rabin-Type [2619]. Race [644]. Radar [2277]. Radiation [961]. Radio [1602, 283, 1234]. Radiofrequency [1673]. Radiographs [850]. Radiology [869]. Radiosurgery [867]. Radiotherapy [1040, 942, 998]. Radius [2587]. Radix [2767]. Radix- [2767]. Railway [2576]. Rajagopalan [2613]. Random [2726, 1854, 1469, 1236, 62, 2059, 2643, 2598, 2628, 972, 1387, 2684, 2062, 2127]. Randomats [1629]. Randomized [2677, 1392, 843, 2680, 1433, 2712]. Randomness [2156, 2766, 2743]. Range [427, 1431, 111, 2077, 2083, 117]. Range-Aggregate [1431]. Rank [2556]. Ranking [446]. Rankings [2545]. Rao [1658]. Rapid [358, 1938]. Rapidly [4]. Rapidly-Changing [4]. Rat [1683]. Rate [355, 1676, 268, 1131, 2100, 283, 2678, 232, 1231, 735, 1232, 1233, 535]. Rate-Based [1131]. Rates [2625, 922, 1549]. Rating [1522]. Ratio [2556, 2323]. Rational [1729]. ray [849, 973]. RBAC [1627, 440]. RC4 [2582]. RC5 [2717]. Re [2475, 469, 2563, 2128, 2011]. Re-dividing [2128]. Re-engineering [469]. Re-evaluation [2011]. Re-scheduling [2563]. Re-visited [2475]. Reachability [2227, 1364, 1252]. Reaction [1786]. Reactions [79, 1503]. Reactive [1187, 770, 1273]. Reactors [714]. Reading [2785, 1054]. Real [1581, 895, 408, 1138, 1362, 2739, 1038, 89, 2231, 2911, 1127, 1472, 1149, 1604, 2670, 596, 1281, 2480, 1132, 120, 2325, 861, 1080, 398, 1001, 399, 815, 154, 816, 1142, 151, 1228, 2394, 530, 1143, 896, 935, 1306, 1130, 122, 859, 88, 573, 233, 747, 1818, 1150]. Real-Life [1149, 1818]. Real-Time [1581, 895, 1138, 2739, 1038, 2231, 1127, 1604, 2670, 1281, 2480, 1132, 120, 861, 1080, 1001, 399, 154, 816, 1142, 1143, 896, 935, 1306, 122, 859, 88, 573, 233, 1150]. Real-Valued [1472, 747]. Realistic [847, 536, 247]. Reality [895, 385, 897, 2421, 407, 154, 1060, 2433, 843, 404, 899, 393, 974, 1061]. Realization [2279, 2271]. Realizing [2451]. Really [1815]. Reasoning [76, 2892, 641, 633, 2465, 1787, 2473, 1641, 1439, 1326, 2517, 2769, 2207, 2188, 2223, 1891, 793, 708, 683, 2490, 1839, 2382, 2570, 1921, 1109, 2772, 2014, 790, 1900, 2516]. Receipt [2111]. Recipient [2720, 2744]. Recognition [2799, 100, 2303, 127, 136, 744, 2738, 2797, 1585, 244, 134, 2289, 2477, 729, 510, 2261, 2260, 310, 249, 2793, 1370, 2360, 524, 129, 133, 153, 2367, 115, 699, 139, 2814, 124, 416, 233, 1365, 2323, 2511, 787, 109]. Recognition-Based [2303]. Recognizing [1449]. Recollection [217]. Recombination [784]. Recommendation [465]. Recommender [1734, 424]. Reconciliation [2489]. Reconciling [488]. Reconfigurable [1114, 1288]. Reconfiguration [1095, 1707, 1343]. Reconstructed [928]. Reconstructing [1672]. Reconstruction [934, 355, 1002, 1439, 1049, 1775, 978, 103, 92, 2591, 983, 2271, 2273, 907, 2511]. Reconstructional [2270]. Record [1449, 1531]. Recording [283]. Recovery [2616, 2731, 1448, 191, 2036, 1710, 2257, 2632, 342]. Rectangles [1871]. Rectification [1063]. Rectilinear [649]. Recurrence [535]. Recurrent [994, 2889]. Recursive [1480, 1568, 1645, 2553]. Reduced [1639, 2643, 2596, 1923, 2142]. Reduced-Round [2596]. Reducing [2545, 1282, 760]. Reduction [1078, 2441, 1653, 2788, 946, 783, 2237, 7, 1023]. Reductions [2588]. Redundancy [1668]. Redundant [2545]. Reed [1541, 2587, 1569, 1534]. Reengineering [1443, 2460]. Reexecution [2042]. Reexecution-Based [2042]. Reference [2425, 692, 693, 227]. References [1812, 2179, 1716]. Refinement [1438, 2188]. Refining [966, 532]. Reflecting [179]. Reflection [177, 896, 158]. Reflections [103, 402, 170]. Reflective [1278, 174, 165, 1142, 157]. Reflex [1022, 157]. Reformulation [739, 1912]. Refusal [2895]. Refutable [1479]. Refutably [1480]. Refuting [1481]. Region [872, 2480, 1383, 972]. Region-Based [2480]. Regional [1688, 1052, 1016].

Regions [497, 1415, 493, 308, 1418, 699]. Register [2526]. Registering [925]. Registers [2125, 114, 2124]. Registration [1007, 923, 928, 901, 880, 1010, 902, 927, 1062, 932, 106, 904, 847, 829, 845, 821, 946, 998, 1043, 920, 2113, 900, 881, 1006, 1045, 844, 924, 828, 824, 878, 1016, 883, 965, 1044, 1005, 846, 1009, 1008]. Regression [48, 69, 2554]. Regular [651, 1484, 1217, 1772, 2070, 619, 2826, 2923, 1365, 1196]. Regularisation [1682]. Regularization [1010, 838, 2282]. Regulation [1737, 303]. Reification [1945]. Reinforcement [783, 768, 93]. Rejection [1570]. Related [1380, 2129, 2082]. Relating [2463, 629]. Relation [802, 740, 2928, 1457, 786]. Relational [173, 466, 2905, 2542]. Relations [2861, 2517, 684, 2071, 2368, 2365, 929, 2381, 557]. Relationship [657, 1438]. Relationships [1415, 1439, 2333, 1413, 1932, 1422]. Relative [706, 2552]. Relaxation [1953]. Relaxations [142]. Relaxed [624]. Release [1346, 2608]. Relevance [452, 446, 34]. Relevant [1805, 2689]. Reliable [1149, 121, 1238, 190]. Remainder [2754, 1654]. Remote [1064, 807, 995, 2429, 757, 461]. Removing [1851, 1364, 177]. Renal [985]. Rendering [834, 977, 1780]. Renee [1211]. Reordering [1833, 270]. Repair [2218]. Repairing [1844, 1097]. Repairs [2894]. Repeatability [960]. Repetitive [1521]. Replaceable [2458]. Replacement [417]. Replanning [2578]. Replica [1964]. Replicating [1166, 1094]. Replication [1081, 1430, 1965]. Report [2, 13, 1493, 19, 1512, 1455, 2761]. Reports [17]. Represent [2568]. Representation [702, 2572, 2907, 2589, 2715, 83, 2848, 728, 2767, 1402, 701, 1302, 697, 346, 2426]. Representations [2467, 694, 2023, 2082, 557, 2232, 124]. Representative [2869, 67]. Represented [2473]. Representing [1983, 736, 2570]. reps [917]. repudiation [562]. Reputation [2090]. Request [418]. Required [2306]. Requirement [341]. Requirements [6, 1122, 2450, 19, 1416]. Research [1804, 982, 2086, 1368, 214, 1652, 373, 590, 1411]. Resection [994]. Reservation [1242]. Reset [1643]. Residential [1082]. Residue [2644]. Residues [2641]. Resilience [338, 227]. Resilient [2121, 2587, 2586]. Resistance [1023]. Resistant [2760]. Resisting [2712]. Resolution [916, 978, 2008, 2190, 1841, 2859, 2562, 2857, 104]. Resolving [1430, 2464]. Resonance [1679, 916, 941, 910, 991, 948, 862]. Resource [1204, 1963, 1131, 455, 1919, 1901, 1241, 489, 1590, 1206, 1234, 1945, 1966]. Resource-Passing [1206]. Resources [1067, 2435]. Respect [753]. Respiratory [1033]. Response [1256, 1161, 919]. Responsibility [2474]. Responsive [2155]. Rest [717]. Restarting [1781]. Restoration [1538, 1862, 269, 2877, 937]. Restricted [659, 2929, 1355]. Restrictions [2547]. Restrictive [2747]. Results [627, 2840, 2066, 54, 1503, 2260, 1796, 2175, 1221, 817]. Resuming [2539]. Retargetable [2037, 1594]. RETE [801]. Rethinking [2452]. Retina [533]. Retinal [101]. Retireval [333]. Retransmission [1747]. Retrieval [220, 2730, 320, 449, 250, 253, 450, 258, 2738, 1774, 1946, 225, 443, 221, 289, 700, 257, 217, 328, 2482, 2737, 1502, 2483, 2252, 893, 2881, 2830, 2398, 334, 2736, 147, 701, 323, 457, 274, 279, 2882, 467, 458, 291, 2329, 446, 290, 278, 251, 324, 281, 280, 34]. Retrieval-Filtering [467]. Retrieving [335]. Retrospective [850, 845]. Reusable [481, 1136]. Reuse [2322, 15, 179, 8]. Reusing [1456]. Reveal [593]. Revenue [1257]. Reversal [1593, 2068]. Reversal-Bounded [2068]. Reverse [1769]. Reversible [1779, 1999]. Reviews [17].

Revised [2358, 2221]. Revision [2090, 2573, 2517, 2899, 2562, 1515, 1519]. Revisited [1461, 1489, 2693, 1935, 1757, 1349, 2598, 1481, 2718, 2824, 1998, 2932]. Revisiting [1825]. Revocation [2164, 2607]. Revolution [1759]. Rewrite [2084]. Rewriting [2927, 1787, 2227, 2900, 1208, 2225, 1800]. Reynolds [318, 1225]. RFSA [1484]. Rho [2598]. Rich [385, 325, 2381, 1426]. Rights [571, 1719]. Rigid [895, 904, 1029, 990, 1010, 902, 821, 920, 1006, 828, 824, 883, 1044]. Rigorous [2824, 1228]. Rijndael [552, 2589, 2590, 2145]. Ring [1567, 946, 1557]. Rings [2893, 1548, 1549]. RIPEMD [2606]. Rise [2326]. Risk [2441, 1626, 7]. Risks [1623]. River [706]. RMI [1270]. RNA [51]. RNA-Sequence-Structure [51]. Robot [1011, 1067, 1069, 2785, 1018, 70, 1462, 1491, 1021, 772, 2293, 2774, 1101, 2824, 2187, 1072, 1104, 888]. Robot-Assisted [1072]. Robotic [1070, 853, 885, 858, 855, 1267, 1066, 844, 1019, 393, 1024, 852]. Robotics [2290]. Robotize [851]. Robots [135, 1262, 1102, 1103, 461, 1263, 617, 1762, 138, 139]. Robust [245, 311, 938, 261, 473, 114, 244, 900, 248, 229, 837, 2118, 1051, 642, 1667, 549, 566, 146, 2001, 862, 124, 302]. Robustness [110, 884, 733, 2172, 1931, 949, 274]. Role [183, 1659, 2395, 176, 693, 480, 697, 1411, 560, 591, 1647]. Role- [480]. Role-Based [1659, 560, 1647]. Role-Binding [176]. Roles [2194, 2628, 1273]. Room [578, 1281]. ROOM-BRIDGE [1281]. Rooms [1352]. Root [982, 126]. Roots [558]. Rotation [744, 817, 979]. Rotational [836]. Rough [57, 2375, 2788, 2361, 2385, 2786, 2368, 2365, 419, 2386, 476, 2373, 2358, 2380, 2783, 461, 2379, 2357, 785]. Round [2155, 2643, 599, 2596]. Rounding [1396]. Route [2099]. Router [1606]. Routers [1609]. Routes [1655]. Routing [1577, 201, 1105, 2779, 1758, 1286, 1624, 2305, 1176].

RS [2376, 2406]. RSA

[2642, 264, 2638, 2149, 2721, 2766, 2756, 2132, 2758, 2741, 2644, 2710, 1637]. RSA-Based [2721]. RSA-Encrypted [2638]. RSA-Primitive [2741]. RSBR [2406]. Rule [1278, 2691, 737, 2469, 2802, 792, 2402, 1757, 2548, 2773, 434, 2404, 521, 2825, 732, 1517, 764, 467, 2513, 1420]. Rule-Based [1278, 2691, 2802, 1757, 467, 1420]. Rule-Bases [2773]. Rules [57, 42, 61, 2769, 1826, 45, 752, 1501, 2087, 67, 1525, 2381, 1508, 2383, 526, 801, 1297, 1447, 2508, 2369, 322]. Run [1288]. Run-Time [1288]. Runlength [2622]. Runlength-Limiting [2622]. Runtime [1318, 583]. Russian [1877].

s [178, 367, 2145, 557]. S-Box [2145]. S-Boxes [557]. S.M.A.R.T [852]. SAC [2643]. SADES [171]. Safe [2018, 851, 522]. Safety [2911, 800, 799, 2416, 166, 1139, 1310]. Safety-Critical [166, 1139]. SAGA [1265]. SAGAs [191]. Sahai [2613]. Sales [2302, 487]. SAM [1974]. Same [928, 1680]. Sample [2908, 1110]. Samples [888]. Sampling [355, 1469, 2688, 2488, 1990, 69, 2552]. Sampling-Based [2552]. Samplings [1995]. SAT [1854, 1907, 1916, 1927, 1856]. Satellite [2804, 125]. Satellites [1900]. Satisfaction [1906, 2466, 1885, 2315, 1889, 1912, 1913, 1901, 1922, 2045]. Satisfiability [1855, 1893, 1107, 2575, 2684]. Satisfiable [1911]. Saturation [2014]. Saturation-Based [2014]. Saver [363]. SBoxes [2713]. SC2000 [2596]. Scalability [1946, 1644]. Scalable [1081, 1171, 1285, 1737, 1771, 1942, 1286, 1252, 1588, 302, 1749, 1333, 1247]. Scalar [2601, 2602, 1621]. Scale [1041, 41, 548, 689, 112, 1742, 2449, 1286, 1741, 605, 1624, 2460]. Scaling [2821, 566, 47]. SCALPP [851].

Scamp [1742]. Scan [2526, 878]. Scans [928, 901]. Scattered [2239]. Scenario [2167, 7]. Scenario-Based [7]. Scenarios [1123, 2414, 1323]. Scene [335, 2427, 151, 322]. Scenes [137]. Scents [75]. Scheduler [1336, 303, 1580]. Scheduling [1581, 2315, 1375, 1345, 1898, 1341, 2582, 1332, 2779, 613, 614, 1346, 1348, 304, 2079, 2576, 1579, 2563, 1968, 1580, 1578, 1338, 1900, 1966, 1948, 1339, 1897]. Schedulings [1347]. Schema [2872, 1426, 1420, 1452]. Schema-Tree [2872]. Schemas [1429, 1421]. Scheme [1581, 1678, 2536, 316, 570, 317, 2641, 2140, 1869, 1345, 2812, 261, 2131, 1833, 2645, 2307, 1346, 337, 2100, 1643, 2126, 304, 1663, 2747, 2827, 1661, 285, 339, 2754, 2632, 29, 354, 1232, 342, 272, 1637, 286, 2117]. Schemes [2058, 538, 1567, 293, 2724, 2708, 2728, 2150, 2723, 2123, 2144, 2005, 2761, 2751, 1622, 1640, 301, 1658]. School [254]. Schützenberger [620]. Science [1460, 1487, 1463, 1488, 1764]. sciences [1754]. Scientific [33, 1496, 1513, 1515]. Scintigraphy [985]. SCOOP [1531]. Scope [2387]. Scoped [1138]. Screen [363]. Scribe [1741]. Scripting [2531]. Scripts [399]. SDK [300]. SDL [2012]. Sea [121, 1186]. Seafloor [2801]. Sealed [565]. Sealed-Bid [565]. Search [2064, 1883, 2466, 448, 1769, 400, 1873, 1862, 319, 2681, 2574, 1898, 452, 1214, 772, 40, 1893, 442, 292, 624, 454, 1875, 283, 1916, 2575, 1877, 2576, 2683, 1514, 1924, 1928, 1896, 307, 2644, 1939, 1903, 1005, 223, 2221, 275, 97]. Searching [1513, 766]. Second [2235, 1224]. Secondary [530]. Secrecy [2750]. Secret [1618, 2714, 1360, 2105, 2162, 342, 1658, 1628, 2615]. Section [1292, 1300, 1290, 1301, 1291, 1293, 1296, 1299, 1302, 1294, 1295, 1298, 1297]. Sectioned [65]. Sections [903]. Secure [2164, 561, 500, 538, 2108, 631, 2725, 2742, 571, 806, 2151, 2745, 1651, 2134, 1625, 2743, 572, 1643, 2490, 2728, 2747, 2118, 2633, 1942, 2025, 2741, 1624, 2119, 1655, 1616, 1654]. Securing [2726, 316, 559]. SECurity [546, 1154, 2746, 500, 164, 1656, 2445, 1933, 206, 2745, 1651, 2146, 1645, 1657, 572, 2753, 2135, 1614, 2150, 2456, 1432, 2754, 1632, 1613, 192, 2499, 1561, 2617, 2142, 1310, 352, 1662]. Security-Aware [164]. SECUSIM [1665]. Seeded [321]. Seeding [867]. Seeds [432]. seeing [2918]. Segment [872, 2537]. Segmentation [968, 149, 969, 940, 2258, 1382, 1993, 938, 987, 971, 840, 1682, 1048, 876, 96, 967, 728, 1047, 1046, 939, 2480, 720, 972, 2808, 241, 2240, 2881, 881, 99, 1051, 936, 941, 56, 917, 1014, 961, 1016, 970, 141, 875, 913, 2253, 2259, 122, 839, 533, 277, 332, 873, 109]. Segmentation/Enhancement [2808]. Segmented [834]. Segmenting [874, 52, 2734]. Segments [2072, 2261]. Seizures [983]. Selecting [2452, 1900]. Selection [1168, 2441, 2443, 2802, 293, 2782, 2474, 136, 2547, 2871, 347, 519, 46, 809, 2854, 797, 2884]. Selective [2111, 2756]. Selector [369]. Selenocysteine [51]. Self [79, 1538, 1539, 2337, 203, 202, 161, 2384, 1668, 207, 96, 208, 130, 1379, 2750, 1107, 2483, 484, 401, 600, 56, 211, 1166, 1535, 1094, 693, 1097, 2243, 2161, 2796, 2932]. Self-Adapting [96]. Self-Adaptive [161]. Self-Assembly [1379, 2932]. Self-Assessment [2483]. Self-Blindable [2161]. Self-Containment [1668]. Self-dual [1538, 1539, 1535]. Self-Evaluation [2337]. Self-Healing [1097]. Self-Injurious [2384]. Self-Organization [484, 2796]. Self-Organized [1107]. Self-Organizing [130]. Self-Perception [401]. Self-Powered [600]. Self-Reference [693]. Self-Repairing [1097]. Self-Replicating [1166, 1094]. Self-Similarity [2243]. Self-Stabilizing [203, 202, 207, 208, 211]. Self-Supervised [56]. Self-tallying [2750].

Semantic [2216, 1412, 228, 1448, 1816, 689, 321, 1756, 2848, 411, 690, 752, 1421, 601, 795, 647, 1419, 2856, 224]. Semantically [2381, 2741, 1426]. Semantically-Rich [1426]. Semantics [2209, 686, 1835, 1435, 608, 2525, 2673, 2211, 2567, 1666, 1834, 2846, 2845, 2569, 2021, 2651, 1838, 2702, 1434, 2666, 2019, 1228, 2024, 374, 2510]. Semantics-Based [2525]. Semi [1482, 1347, 1499, 1020, 315, 1051, 343, 1518, 441]. Semi-automatic [1051]. Semi-autonomous [1020]. Semi-fragile [343]. Semi-normal [1347]. Semi-structured [1482, 1499, 1518, 441]. Semidefinite [2061]. Semidirect [1563]. Semiotic [2484]. Senary [1545]. Sender [1749]. Sensation [935]. Sense [2853, 2855, 2852]. Sensed [2434]. Sensing [757]. Sensitive [319, 2432, 2072, 1625, 1321]. Sensor [581, 68, 111, 1114, 461, 954]. Sensors [1157]. Sensuous [2849]. Separability [612]. separable [2258]. Separate [1476]. Separated [1081]. Separation [96, 171, 162, 1529, 2264]. Sepe [2865]. Septic [521]. Sequence [1377, 51, 649, 2241, 827, 2260, 2078, 2682, 1554, 1436, 125]. Sequences [2626, 1879, 121, 986, 45, 238, 2304, 920, 2240, 2372, 2639, 1555, 117, 146, 1557, 124, 2283, 357]. Sequential [2686, 547, 2688, 1955, 2488, 1705, 2580, 1198, 1077]. Serendipity [606]. Serial [1054, 1016]. Series [2555, 52, 918, 502, 922, 2244, 1524, 728, 54, 2398, 529, 919, 1386]. Series-Parallel [1386]. Server [1748, 448, 2749, 1865, 427, 1946, 1739, 2132, 2459, 1590, 440, 2531]. Server-Aided [2132]. Server-Side [2531]. Servers [1332, 421]. Service [645, 430, 1969, 561, 2785, 578, 1649, 1255, 203, 715, 1742, 1279, 187, 1269, 1238, 1964, 304, 1280, 1241, 582, 1243, 1942, 2458, 192, 1161, 1181, 1182, 587, 1183, 1646, 1251, 1723]. Services [494, 1284, 1962, 186, 189, 1604, 1958, 1083, 1254, 188, 434, 1245, 1611, 1274, 185, 195, 193, 182, 1276, 2114, 1082, 1720, 1718]. Serving [1745]. Servoing [1020]. Session [186, 292, 1272]. Session-Oriented [186]. Sessions [2531]. Set [2375, 106, 1399, 2534, 2361, 2385, 419, 2386, 1984, 2866, 1384, 67, 2272, 2273, 2210, 476, 2373, 1454, 2358, 2783, 617, 2772, 2041, 2367, 953, 2357, 785]. Set-Aided [419]. Set-Based [2534]. Set-Oriented [2210]. Sets [57, 2062, 1571, 940, 1395, 2837, 742, 36, 1494, 667, 2368, 2365, 719, 508, 1107, 1542, 1525, 1990, 2379, 731]. Setting [540, 2447, 1175, 2150]. Seven [597]. Several [1814, 543, 1235]. Sevofluorane [730]. SFLASH [2640, 556]. Shadow [2087]. Shafer [2774, 1684]. Shape [874, 819, 250, 311, 131, 822, 779, 144, 818, 1012, 1058, 101, 872, 914, 869, 1017, 323, 906, 145, 324]. Shape-Based [101]. Shape-Preserving [779]. Shaping [1246]. Shared [1340, 1589, 1496, 2293, 1392]. Sharing [2625, 1751, 1726, 1162, 2343, 1600, 1360, 2105, 2431, 342, 1628]. Sharpe [2556]. Sheets [296]. Shelf [2439, 2459]. Shift [2125, 995, 2124, 970]. Shock [521]. Shooting [389]. Shop [613, 614]. Shopping [2295, 483, 460, 1724, 2394, 480, 1722]. Short [2160, 375, 54, 324]. Short-Time [324]. Shortened [2744, 2611]. Shortest [1769, 1992, 659, 1385]. Shot [318, 327, 222]. Should [1442]. Shrinking [1098]. sic [724]. Side [2764, 482, 2760, 336, 559, 1878, 2531]. Side-Chain [1878]. Side-Channel [2764, 559]. Sided [113]. Sieve [2106]. Sight [2918]. Sight-seeing [2918]. Sigmoid [1467]. Sign [313, 2844, 2847, 233, 290]. Signal [1058, 1902, 461, 2262, 2264]. Signaling [1177]. Signals [1686, 499, 2238, 2265, 2266, 2232, 2267]. Signature [1618, 2140, 2794, 2131, 2724, 2645, 2708, 2126, 2150, 1663, 2747, 2723, 2123, 1661, 2761, 1637, 781, 436, 1622, 1640].

Signatures [2164, 2160, 2630, 1617, 2721, 2745, 1638, 569, 2619, 2756, 2722, 1153, 2122, 1613, 1664]. Signcryption [2746]. Signed [695]. Significant [493]. Signing [564]. Signs [282]. SIMBA [97]. SIMD [553]. SimHuman [88]. Similar [2398, 329]. Similarity [1010, 250, 2732, 2771, 1525, 2887, 2780, 308, 1998, 2856, 115, 2886, 2823, 2243, 223]. SimPlanner [2578]. Simple [895, 1992, 164, 1688, 2589, 610, 1833, 2697, 2614, 1638, 2238]. Simple-Deformable [895]. Simpler [1466]. Simplex [1545]. Simplicial [1983]. Simplification [366, 1587]. Simplified [992, 1520]. Simplifying [360, 2509, 2206]. Simulate [1785, 2008]. Simulated [1036, 867, 847, 2685, 771, 2367]. Simulating [1688, 2310, 950, 1035]. Simulation [2686, 1178, 831, 870, 833, 932, 2355, 687, 260, 1030, 2290, 2343, 869, 958, 1001, 2348, 931, 2175, 2174, 832, 2091, 1665, 2055, 1004, 2499, 1689, 935, 1597, 2344, 868, 1662, 90]. Simulation-Based [1030]. Simulations [2485, 2356, 1583]. Simulator [1731]. Simultaneous [2601, 1732, 1027, 109]. Single [2521, 1769, 985, 2576, 917, 1897]. Single-Figure [917]. Single-Issue [1897]. Single-Threaded [2521]. Single-Track [2576]. Singular [742]. Singularity [2233]. Sinograms [840]. Site [1745, 451, 1522]. SiteLang [1423]. Sites [1168, 1423]. Situated [2306]. Situation [2566, 2846, 2845]. Size [916, 818, 383, 508, 615, 1367]. Skeletonization [2000]. Skeletons [1999]. Sketch [921]. Sketches [2327]. Skewness [2264]. Skimming [1240]. Skin [851]. Skull [1024, 866]. Skull-Base [1024]. SLAAC [552]. SLAAC-1V [552]. Sleeping [1815]. SLG [2539]. slice [978]. Slitlamp [1057]. Slitlamp-Based [1057]. Slope [281].

Sloppy [1899]. Small

[30, 1395, 667, 821, 509, 39, 2391, 1510].

Smallest [629]. Smart

[886, 585, 2645, 2767, 1728, 298, 2291]. Smart-Its [585]. Smartcard [568]. Smartcard-Based [568]. SMIL [325]. Smooth [2683]. Smoother [350]. Smoothness [1382]. SMS [1656, 599]. Smushing [777]. Snakes [1050]. SNAP [812]. Snarfing [601]. Soccer [138]. Social [577, 2341, 2338, 2332, 2333, 415, 2353, 2350]. Societal [2568]. Societies [80, 81, 636, 2561, 2086, 637, 2565]. Society [72, 2092]. Socio [2089]. Socio-cognitive [2089]. Soft [1581, 1566, 2785, 932, 2816, 2679, 2792, 1910, 857, 2614, 830, 2829, 863, 2360, 999, 1004, 1034]. Soft-Biometrics [2792]. Soft-Computing [2792]. Soft-Input [2614]. Soft-Input/Soft-Output [2614]. Soft-Output [2614]. Soft-Tissue [857, 863, 1034]. Software [20, 26, 2441, 2, 13, 2415, 27, 1148, 2053, 1188, 1140, 2125, 2446, 1125, 1126, 372, 2445, 161, 1256, 1958, 2448, 28, 172, 1586, 381, 482, 18, 1599, 1032, 1141, 2036, 15, 2452, 1134, 1600, 3, 2413, 1283, 2454, 1633, 1694, 1783, 1136, 1, 2706, 2346, 8, 4, 2502, 1145, 584, 2035, 1146, 1903, 868, 1151, 5]. Software-Managed [1599]. Soil [2281]. Solid [930]. Solo [1207]. Solomon [1569]. Solution [156, 2248]. Solutions [2439, 165, 1380, 629, 2913, 2456]. Solve [2727, 660, 1904]. Solver [2047, 1924]. Solvers [1827, 1922, 2046]. Solving [1848, 2465, 2681, 1814, 2755, 2574, 1849, 1893, 1857, 1870, 1881, 2781, 1798, 1111, 376, 133, 2506, 1104]. Some [1563, 1120, 2807, 2840, 2839, 618, 1368, 1817, 1390, 2260, 2915, 2635, 1796, 2082, 1931, 2684, 2832, 796, 1900, 2710, 731]. Song [281]. SOR [2796]. Sorted [1841]. Sorting [419]. Sorts [1841]. Sound [387, 458, 429]. Sounds [282]. Source [1428, 1769, 353, 1159, 1414, 2264]. Source/Channel [353]. Sources [2090]. SPA [2758]. SPA-Based [2758]. Space [355, 925, 2049, 880, 2185, 1573, 596, 1989, 2364, 1779, 1402, 2492, 1513, 1990, 1536, 695, 212, 2243, 1026, 2330]. Space-Efficient [1402]. Space-Frequency [2243]. Space-Optimal [212]. Space-Time [1573, 1536]. Spaces [649, 1814, 131, 689, 1777, 2836, 2627, 1357, 1980]. Spacetime [399]. SpaMod [1981]. Spanish [2868, 2861, 2869, 2865, 2860, 2858, 55, 2876, 2882]. Spanning [2076]. Sparse [2679, 1195]. Spatial [1981, 702, 686, 2467, 1202, 2066, 691, 704, 687, 679, 1417, 1414, 708, 683, 2237, 273, 2112, 692, 387, 1413, 701, 693, 875, 697]. Spatially [983, 820]. Spatio [1031, 685, 104, 1416]. Spatio-Temporal [685, 104, 1031, 1416]. Speaker [2810, 2811, 369]. Special [502, 2121]. Specialization [1908]. Specializing [1877]. Specific [944, 833, 2482, 958, 1134, 140, 931, 1241, 1876, 363, 480, 1426, 868]. Specification [2691, 2033, 2069, 427, 1934, 2034, 1132, 2652, 2653, 2654, 2492, 1696, 2032, 1229, 1434, 2028, 1446, 2706, 1971, 1163, 2890]. Specification-Based [2034, 1163]. Specifications [2695, 1427, 2698, 2012, 2181, 2704, 2705]. Specify [2675]. Specifying [1941]. SPECT [847, 894]. Spectra [755]. Spectral [1673]. Spectrally [2639]. Spectroscopy [729]. Spectrum [294, 1731]. Specular [103]. Speculative [1831]. Speech [2730, 2863, 2739, 2303, 326, 365, 356, 2738, 2864, 2428, 2737, 2808, 2269, 154, 2736, 2367, 2814, 55, 2262, 284]. Speech-Driven [2738]. Speeches [93]. Speed [938, 1604, 568]. Speeding [2602, 2138]. Speedup [1607]. SPEM [232]. Sphere [879]. Spherical [1496, 147, 990]. SPI [21, 8]. SPIHT [270, 340]. Spiking [1262]. Spine [962, 844, 965, 877]. Spiral [932]. Splicing [2929]. Spline [2239, 2550, 948]. Splines [779, 106, 1048]. Splitting [99, 2191]. SPM [989]. Spoken [2733, 2500, 695, 291]. Spontaneously [99]. Spot [792]. Spread [294, 432]. Sprite [229]. SQL [1078]. Square [1853, 1639]. Squares [148, 2554]. SRPT [1332]. SSC2 [2583]. SSL [1630]. st [2748]. Stability [1986, 113, 2246, 1575, 528]. Stabilization [200, 201, 198, 205]. Stabilized [286]. Stabilizing [203, 202, 207, 209, 208, 210, 211]. Stable [1835, 1861, 1836, 672, 1925]. Stack [1364]. Staged [381]. Staggered [1170]. STAIR [1746]. Stairs [1571]. Standard [2867, 2149, 2354, 2192, 1027, 2459, 2880]. Standards [1451]. Standards-Driven [1451]. Stapedotomy [1069]. Star [662, 673]. Start [22]. Start-Up [22]. Starting [19]. Startup [23]. State [1190, 1862, 136, 2015, 1908, 2475, 1773, 2006, 1195, 729, 1197, 1111, 1782, 138, 2259, 1727, 898, 1196]. Stateless [2720, 286, 1244]. Stateless-Recipient [2720]. Statements [1294]. States [2487, 717, 2639, 2460]. Static [377, 2794, 2701]. Station [288]. Statistical [2610, 874, 1033, 822, 2738, 2671, 510, 314, 2628, 893, 1017, 118, 129, 824, 2636, 864, 913, 839, 1465, 251]. Statistically [2714]. Statistically-Secret [2714]. Statistics [1067, 131, 280]. Status [2761]. Steady [898]. Steady-State [898]. Steered [936]. Steganography [316, 548]. Steiner [649, 668]. STeLLa [2580]. Stenting [833]. Step [1007, 2012, 771, 2362, 1088]. Steps [2566, 1264, 2621, 603]. Stereo [2817, 235, 109]. Stereoscopic [146, 2328]. Stereotypes [1434]. Stick [1899]. Sticks [1629]. Still [10]. Stimulating [1505]. Stimulation [1062, 919]. Stimulation-Response [919]. Stochastic [1850, 1326, 2679, 2575, 905, 2266, 2687]. stochastICP [905]. Stock [2486]. Storage [1127, 344, 1195]. Stored [350]. Stores [1627]. Stories [391, 398]. Storing [14].

Story [401, 387, 171, 404, 2459]. Storyboard [2450]. Storyteller [85]. Storytelling [84, 400, 395, 391, 406, 405]. Straight [107]. Strain [1676]. Strategic [1422, 2096]. Strategies [2878, 2906, 1965, 2631, 590, 749]. Strategy [2629, 196, 2225]. Stream [2618, 2124, 2634, 1590, 2127, 811]. Stream-Cipher [2127]. Stream-Packing [1590]. Streaming [350, 261, 1231, 1232, 1247, 1082]. Streamlining [2449]. Streams [1175, 1180]. Strength [495]. Stretch [1577, 1858]. Strict [543]. STRING [1380]. Strings [2912]. Stroke [1007, 2260, 913]. Stroll [2433]. Strong [1943, 2620, 673, 1224]. Structural [618, 259, 1891, 1549, 335, 1198, 820, 2502, 534, 1219]. Structure [51, 1855, 912, 1747, 1595, 59, 684, 103, 209, 1522, 2271, 2272, 2273, 1920, 1699, 1771, 1052, 755, 530, 1299, 695, 346, 223]. Structured [133, 1482, 1499, 1518, 441]. Structures [2018, 1785, 149, 874, 538, 2066, 962, 818, 967, 1498, 2522, 1978, 881, 2048, 1367, 997, 2923, 2024, 5, 2538]. Structuring [1444]. Student [482]. Studied [2405]. Studies [500, 1682, 1093, 2348, 959, 1683, 980, 11, 29]. Study [968, 1850, 2037, 921, 1649, 2878, 2347, 1934, 1498, 1030, 2236, 2737, 2483, 1245, 38, 505, 2174, 2793, 7, 439, 2795, 1587, 530, 2305, 1526, 1528, 2554, 952, 1180, 1973, 1398]. Stumps [2471]. Style [2471, 2734, 2379]. Sub [2468, 1568]. Sub-beam [2468]. Sub-codewords [1568]. Subacute [1007]. Subband [2238]. Subcolorings [663]. Subcortical [983]. Subdefinite [2045]. Subdivision [104]. Subfamilies [670]. Subgraph [1899]. Subgraphs [661, 671]. Subgroup [1650]. Subgroups [2840]. Subject [2178, 2408, 319, 2079, 845]. Subjective [216, 753]. Subjects [894]. Submodular [665, 1344]. Subprograms [1295]. Subsampling [50]. Subscribe [190]. Subset [1797]. Subsethood [2805]. Subspace [120]. Subspaces [363]. Subthalamic [968]. Subtle [595]. Subtype [1219]. Subtyping [622, 2196]. Subway [1397]. Subword [291]. Success [1582, 9, 2459]. Successful [2447]. Successfully [2413]. Successor [2055]. Sudan [1551]. Sufficient [1568, 2743, 1360, 426]. Suffix [1402, 55]. Sugeno [779, 766]. Suggestion [442]. Suitability [541]. Suitable [2108]. Suite [1323]. Suited [1092]. Summaries [35]. Summarization [216, 276, 719, 2883, 1240]. Summary [1810, 2176, 2366, 1316, 2357]. Super [978]. Super-Resolution [978]. Superelliptic [2158]. Superimposition [895]. Superior [1412]. Supermedia [1259]. Supersingular [2159]. Supervaluation [686, 682]. Supervised [2903, 2855, 56, 492, 2324]. Supervised-Unsupervised [2855]. Supervising [1021]. Supervision [1736]. Supply [1736]. Support [1118, 496, 1469, 897, 1425, 688, 132, 2401, 482, 217, 572, 805, 1525, 594, 486, 1241, 503, 524, 1416, 2201, 1468, 1320, 133, 2393, 2392, 1597, 1528, 60, 587, 368, 1422]. Supporting [28, 2410, 1279, 1250, 167, 832, 1272, 190, 588, 477, 584]. Supports [2188]. Surface [2861, 1039, 121, 904, 117, 1999]. Surfaces [148, 901, 819, 360, 840, 1039, 104, 1997]. Surgeon [816]. Surgeries [1004]. Surgery [968, 886, 897, 870, 932, 1021, 994, 857, 1020, 1065, 1006, 855, 1001, 887, 832, 816, 1066, 1019, 954, 1024, 1025, 949, 868, 871]. Surgical [882, 1032, 858, 950, 935, 852]. Surveillance [513]. Survey [2058, 2677, 2915]. Survival [520]. Susceptibility [2371]. Suspending [2539]. SVD [783]. SVM [2483]. Sweep [1871]. Sweeping [1383]. Switches [1607]. Switching [1475, 756]. Symbiosis [16]. Symbol [2270]. Symbolic [57, 1186, 1832, 728, 2813, 2516]. Symbols [2191]. Symmetric [665, 2900, 1562, 1224, 1333]. Symmetries [1851]. Symmetry [1852, 1917, 1329]. Symposium [1493]. Synchronisation [2025]. Synchronization [1120, 608, 1298]. Synchronized [1208]. Synchronizer [1229]. Synchronous [1121, 623, 317, 584]. Synchrony [2213, 1124]. Syntactic [2893, 2861, 1816, 2882]. Syntax [2666, 2672, 1316]. Synthesis [2017, 2626, 1595, 2020, 2304, 1648, 1248, 2019, 2800, 1558, 312]. System [81, 2720, 1178, 737, 1099, 831, 2203, 897, 2469, 422, 2337, 349, 360, 1375, 1656, 372, 2474, 2695, 800, 799, 1898, 1021, 2584, 2384, 2338, 739, 2577, 1070, 853, 721, 482, 746, 885, 2556, 2779, 1457, 2849, 2293, 331, 260, 443, 221, 1257, 2416, 255, 419, 483, 238, 2307, 956, 1967, 2477, 1071, 908, 998, 2322, 826, 1164, 748, 1054, 1569, 863, 2883, 513, 572, 460, 1133, 1245, 1660, 861, 465, 249, 431, 2269, 345, 248, 516, 2316, 574, 1828, 855, 517, 2860, 438, 486, 1267, 815, 1479, 2650, 589, 1283, 887, 1323, 2578, 2277, 359, 1701, 714, 2457, 1780, 503, 524]. System [2455, 580, 1052, 1249, 1019, 762, 2019, 899, 2675, 693, 1024, 2367, 435, 2392, 1597, 1932, 235, 949, 2804, 2789, 279, 1234, 532, 233, 158, 1160, 447, 730, 2806, 2096, 996, 299, 1337, 1082, 486]. System-Level [1133]. System-Theoretic [1569]. Systematic [27, 2029]. Systematically [2552]. Systems [76, 1733, 2438, 1734, 2626, 1581, 1253, 1118, 1119, 496, 2838, 1748, 1785, 2440, 646, 2922, 1120, 1187, 778, 1121, 616, 1148, 2927, 1770, 1122, 1429, 1822, 1755, 356, 1584, 2907, 2755, 640, 1326, 1127, 396, 1790, 2421, 1217, 577, 1725, 1940, 1908, 1933, 687, 1774, 204, 772, 2227, 2432, 1833, 468, 1287, 2848, 1169, 2074, 2919, 1131, 1757, 2308, 424, 2482, 2236, 1625, 2737, 1132, 2050, 1836, 512, 1760, 1793, 1483, 598, 2135, 1080, 2317, 2562, 1134, 2095, 2013, 1796, 1283, 488, 1797, 1350, 1116, 1798, 2702, 287, 945, 1921, 1579, 1693, 2412, 1197, 2362, 1137, 1799, 1286]. Systems [166, 1139, 1252, 1968, 1450, 1142, 1320, 1220, 773, 2127, 1143, 1147, 2359, 1144, 1516, 1800, 1305, 1306, 1308, 1307, 2046, 583, 2399, 560, 638, 1945, 1155, 1150, 426, 2796, 224, 1761, 1098]. Systems-Challenges [1143]. Systolic [1688].

t [1027, 2105, 2838]. T-Fuzzy [2838]. T1 [721, 1047]. T2 [721]. Table [2788, 2366, 312]. Tableaux [2894, 2223, 2189]. Tabled [1832, 1833]. Tables [173, 683]. Tabling [1824]. Tabu [1005]. Tabular [1819]. Tachycardia [535]. Tagged [1679, 1031]. Tagger [2865]. Tagging [1678, 2863, 2864, 2866, 39, 55]. Tail [1539]. Tail-Biting [1539]. Tailed [1873, 1331]. Tailored [2044, 2902]. Tailoring [2044]. Takagi [779]. Takagi-Sugeno [779]. Take [888]. Taking [852]. Talk [2520]. Talking [2301, 2300]. tallying [2750]. Tapping [289]. Target [2738]. Targeted [188, 2221]. Targets [1941]. Tartar [2875]. Task [1946, 1103, 1065, 1327, 140, 1323, 1321, 1948]. Task-Level [1065]. Tasks [1581, 2474, 613, 1065, 2430, 1298, 2426]. Taxonomy [1735, 681, 1200]. TBNA [933]. TBNA-protocols [933]. TCP [1178, 2638, 1245, 1231, 2026, 1233, 1180]. TCP-Friendly [1233, 1180]. TCP/IP [1178]. TD [2626]. TD-Sequences [2626]. TEA [2643, 2884]. Teacher [482]. Teaching [486]. Team [2924, 586]. Team-Automata-Driven [2924]. Teamwork [587]. Tears [721]. Technical [2879]. Technique [1469, 1871, 1884, 2790, 992, 585, 2044, 2252, 283, 2238, 1591, 2818, 859, 1333, 2615]. Techniques [1538, 1121, 777, 1497, 2396, 2477, 1977, 720, 1916, 516, 2079, 836, 1683, 2454, 1921, 1667, 937, 2294, 143, 190, 492, 2804, 533, 291].

Technological [1496]. Technologies [895, 688, 2792, 2288]. Technology [2439, 2445, 460, 2830, 2168, 182, 4, 713, 1076, 852]. tegration [1817]. Tele [853, 298]. Tele-echography [853]. Tele-education [298]. Telecom [1453]. Telecommand [1711]. Telecommunication [1077]. Telecommunications [1074, 186, 1080]. Telemanipulation [1019]. Telemetry [1712]. Teleoperation [2293]. Telephony [66]. Telerobotic [1259]. Tell [683]. Telling [401, 387]. Template [1956, 173, 2051, 246]. Templates [952]. Temporal [1842, 2465, 1425, 2183, 1324, 2577, 685, 2180, 1132, 1437, 38, 683, 273, 345, 104, 2926, 1866, 2497, 2182, 820, 1929, 66, 1031, 1416]. Temporal-Spatial [273]. Temporomandibular [1002]. Tensor [1059, 838, 837]. TER [853]. Term [2893, 2227, 442, 2242, 2393, 1606, 34]. Terminal [1371]. Termination [2049, 2226, 1210, 2225, 2224, 2215]. Terms [609, 2870, 319]. Test [2033, 688, 2136, 1455, 2552, 1570, 2323]. Test-Driving [2552]. Testability [1199]. Testable [2448]. Testing [552, 2034, 1043, 2887, 2824, 751, 2032, 166, 1938, 1782, 2841, 2511]. Text [2888, 2731, 449, 242, 356, 2738, 2256, 567, 1502, 2488, 2881, 309, 64, 2884]. Text-to-Speech [356]. Textual [1442]. Textural [98]. Texture [253, 2533, 98, 2501, 976, 1055, 231]. Textured [245]. TH [1656]. TH-SMS [1656]. Theatre [405, 974]. Their [2626, 563, 2910, 1954, 1341, 108, 2765, 2628, 1795, 1116, 1025, 557, 2000, 2243, 2560]. Them [14]. Theme [1535]. Theorem [2063, 2193, 1837, 539, 2358, 2783, 1654, 2754, 318]. Theorems [2904]. Theoretic [1389, 1569, 2082, 2041, 2197, 1537, 2615]. Theoretical [2017, 556]. Theories [2192, 2084, 1128, 2206]. Theory [1186, 1729, 1887, 2375, 738, 2188, 709, 367, 679, 205, 2385, 736, 1791, 682, 2902, 1507, 1917, 1219, 2318, 1513, 1434, 1212, 2005, 1799, 2795, 1536, 1519, 1980, 2357, 446, 775]. Therapeutic [993]. Therapy [961]. Thermometry [991, 862]. Thesaurus [449, 2886]. Thickening [1688]. Thickens [1854]. Thickness [929, 910]. Thin [106]. Thin-Plate [106]. Things [711]. Think [2087]. Thinking [1496, 2393]. Thinning [1999]. Third [2126, 562, 1664]. Thompson [1193]. Thoracic [827, 1056]. Thorax [881]. Thoughts [1817]. Thread [1275]. Threaded [2521]. Threading [1879]. Three [808, 1448, 684, 1933, 2249, 827, 1056, 2304, 603, 836, 959, 2570, 1263, 842, 2328]. Three-Dimensional [827, 1056, 836, 959, 2570, 842]. Three-Layer [1448]. Three-Level [808]. Three-Valued [684]. Threshold [2630, 2149, 2151, 2249, 628, 2150, 2754, 1199, 756, 1570, 1637, 101]. Thresholding [2480]. Thresholds [2064, 1251]. tier [2031]. Tight [2164, 1369, 2702]. Timbuk [2227]. Time [1581, 2555, 895, 1138, 2625, 769, 1404, 1393, 2739, 1038, 89, 1375, 2727, 52, 2231, 918, 1345, 1127, 2188, 660, 502, 1604, 922, 2670, 2476, 1372, 2577, 986, 1573, 2022, 1129, 1524, 728, 669, 1281, 1346, 54, 328, 2480, 1132, 1366, 1779, 120, 2113, 37, 2325, 861, 1080, 1828, 2088, 1001, 399, 154, 816, 2398, 1798, 1215, 1579, 2274, 529, 1142, 1536, 1143, 2026, 896, 935, 1306, 1130, 2292, 122, 919, 859, 2228, 88, 573, 233, 1288, 1150, 324]. Time-Frequency [2274]. Time-Optimal [2577]. Time-Series [922, 728, 2398]. Time-Sharing [2625]. Time-Triggered [1129]. Time-Varying [37]. Timed [2074, 2608, 2013, 2080, 1216, 2024]. Timed-Arc [2080]. Timed-Release [2608]. Timers [1187]. Times [614, 1346, 10]. Timetable [2315]. Timing [2759]. Tiny [1126]. Tissue [932, 857, 909, 863, 1677, 999, 1004, 875, 1035, 1034, 981]. Tissue-like [909]. Tissues [830]. TLB [1599]. Together [2086]. Token [211]. Tokenization [2864]. Tolerance [1794, 1089]. Tolerant [1581, 1755, 2307, 1591]. Tomographic [840, 896]. Tomography [1986, 901, 803, 985]. Too [761]. Tool [2888, 1981, 886, 1064, 770, 1425, 876, 207, 2401, 18, 993, 24, 2129, 1502, 1452, 484, 515, 2317, 2519, 1323, 94, 1665, 278, 392, 2426, 546]. Toolbox [1377, 1100]. Toolkit [87, 1903]. Tools [1038, 644, 1015, 1134, 1028, 2666, 1464, 1490, 2335, 1562]. Toolset [1144]. Toothbrushes [1496]. Topic [422, 2734, 2551]. Topics [2595, 470]. Topological [1415, 1381, 1413, 1248, 1418]. Topologies [740, 536]. Topology [874, 1976, 1604, 1978, 2832, 2834]. Total [1375, 1345, 1675, 1071, 863]. Touch [2333, 91]. Tour [2918]. Tourists [193]. Tournament [1888]. Toy [2303]. Traced [189]. Traces [2911]. Tracing [2752, 489, 2141]. Track [2576, 2606, 1747]. Tracker [835]. Tracking [422, 358, 1031, 838, 331, 238, 152, 120, 241, 123, 815, 816, 1060, 997, 1942, 370, 281, 109]. Tract [2303]. Trade [1010, 2354]. Tradeoff [273]. Tradeoffs [2749, 1890]. Trading [2347, 2556, 2353]. Traditional [2868]. Traffic [1237, 2778, 1602, 1236, 807, 238, 1175, 1610, 768, 439, 122, 256]. Traffics [2292]. Tragedy [644]. Trail [2629]. Trails [2606]. Train [1376]. Trained [2815, 732]. Training [1469, 857, 255, 39, 406, 73, 1467, 1520]. Traitor [2752, 2141]. Trajectories [2916]. Tram [594]. Trans [993]. Trans-urethral [993]. Transaction [1631, 1625, 2171, 2169]. Transactional [1278, 2170, 190]. Transactions [1635, 1279]. Transcoding [225]. Transcranial [1062]. Transcribed [2736]. Transcripts [2732]. Transducers [1773, 1195, 1197]. Transduction [859]. Transfer [2588, 2751, 952, 1233, 1616]. Transferable [569, 791]. Transfiction [390]. Transform [149, 2263, 113, 2220, 2245, 252, 511, 2279, 2269, 2255, 2265, 717, 315, 2266, 919, 2789, 332, 2284, 2247, 364, 2267, 2280]. Transformation [2018, 2017, 2695, 2012, 2919, 257, 1473, 2918, 2702, 1136, 2275, 2028, 878, 2000]. Transformational [2016, 383]. Transformations [1222, 1816, 2696, 2019, 2551]. Transformed [2237]. Transformers [2790]. Transforming [1797]. Transforms [1987, 1999]. Transistor [1093]. Transition [1874, 1785, 377, 1875, 1107]. Transitions [1856]. Transitive [2194, 2723]. Transitivity [2196]. Translating [1943]. Translation [621, 2911, 1363, 1437, 2255, 2847]. Translation-Invariant [2255]. Translations [1910]. Transmission [297, 354, 1232, 2292, 234, 302]. Transparency [172, 1275]. Transparent [1179, 1944, 562, 231]. Transport [1238, 318]. Transportation [646]. Transports [1179]. Transversals [653]. Trapdoor [2714]. Trapezoidal [737]. Traveling [1888]. Traversal [208, 2528, 212]. Treatment [1191, 867, 998]. Tree [1752, 2544, 2219, 1405, 1747, 654, 356, 1862, 1751, 2534, 668, 2872, 1466, 2076, 2271, 2272, 2273, 1744, 1402, 1027, 521, 456, 1359, 133, 2259, 533, 346]. Tree-Approximations [1359]. Tree-Based [356]. Tree-Degree [654]. Tree-Structure [346]. Tree-Structured [133]. Trees [2064, 1404, 2471, 2912, 49, 625, 1495, 48, 624, 626, 2765, 2674, 1778, 1407, 1173, 528, 2554, 1749]. Treewidth [657]. TrekTrack [599]. Trellis [353]. Trellises [1539]. Tremor [859]. Trial [843]. Triangular [794]. Triangulation [1358]. Triggered [1129]. Triple [2143, 552]. Tripod [1417]. Trous [257]. True [1218, 304, 386, 680]. Truncated [2142]. Trust [1284, 2090, 2093, 2086, 2089, 2087, 2490, 2095, 2088, 2094, 2091, 1932, 2092, 2096].

Trusted [562, 1723, 1155, 1664]. Trustworthy [2705]. TSP [1956]. Tubular [149, 834, 967]. Tumors [955, 1055]. Tumour [951]. Tuning [2851]. Turbid [909]. Turbo [353]. Turbulence [1583]. Turbulent [10]. Turing [1094, 1761]. Turn [692]. Turning [2411]. Twice [2206]. Twin [2122]. TWLinuX [1597]. Two [1007, 1577, 1676, 759, 2595, 1638, 696, 1407, 1245, 2429, 717, 2116, 959, 2605, 1642, 1781, 11, 937, 2366, 1088, 1622, 1398, 2606]. Two- [959]. Two-Color [1642]. Two-Dimensional [1676]. Two-Population [759]. Two-Step [1007, 1088]. Two-Track- [2606]. Two-Way [1781]. Type [1544, 1584, 1438, 1053, 2828, 1475, 2619, 2657, 2654, 2650, 1434, 2197, 2665, 530]. Type-Theoretic [2197]. Typed [2043, 1204, 2528, 2198, 2531]. Types [1223, 2572, 611, 610, 2844, 2022, 1133, 2549, 1292]. Typical [1619]. Typing [1227, 179].

U [501, 2346]. U-Mart [2346]. Ubiquitous [597, 2409, 1281, 598, 603, 602, 582, 590, 588, 584, 583, 587, 606]. Ubiquity [1534]. Uda [1108]. UDP [1245, 1228]. Ultimate [1835]. Ultrafast [2618]. Ultrasonic [1676, 751]. Ultrasound [926, 923, 1686, 1070, 993, 920, 943, 855, 959, 527, 954, 842, 848, 965]. Ultrasound-Guided [855, 842]. UML [1441, 2027, 2664, 1435, 2668, 2412, 170, 1434, 178, 2028, 2675, 1436, 25]. UML-Like [2027]. UMTS [1182]. Unbelievable [2135]. Unbounded [1345]. Uncertainties [1626]. Uncertainty [2517]. Unconditionally [538, 2745, 2633, 1616]. Unconditionally-Secure [1616]. Unconstrained [2427, 2511]. Undecidable [1220]. Undeniable [2721, 2645, 1637]. Underdefined [2755]. Underground [704]. Understanding [30, 1822, 818]. Undetachable [2630]. Undetected [2611]. Unearthing [593]. Unexpected [692].

Unfair [199]. Unfolding [2204]. Unfoldings [2030]. Unicast [811, 603]. unifiability [634]. Unification [2194, 2893, 2198, 1516]. Unified [659, 2688, 1648, 1288, 1400]. Uniform [1478, 2271, 2273, 883]. Uniformed [1327]. Unifying [498, 2425, 378, 1569, 1537]. Union [1223, 1477]. Uniqueness [2906, 764]. Unison [1229]. Unit [478, 1893, 613, 1706, 2683]. United [2460]. Units [1301]. Universal [758, 1085, 1457, 1349, 1086, 1094]. Universality [1796]. Universe [2786, 2368, 2650]. UNIX [574]. Unknown [1191, 137]. Unlock [363]. Unmanned [2372]. Unreliable [1179]. Unsatisfiability [628]. Unsolvability [2007]. Unsupervised [2855, 941, 236]. Unsupportive [202]. Untidy [1089]. Untyped [610]. Unwrapping [991]. Update [2208, 2560, 2906]. Updates [2223, 2515]. Updating [702, 640, 2487, 444]. Upper [2837, 2361, 2072, 2929, 2590, 615]. Ups [1343]. UPT [1085]. Uptake [918]. Urban [646, 705]. urethral [993]. Urinary [514]. Urn [2349]. Usability [2415, 689, 2423, 12, 2416, 1328, 279]. Usage [1123, 1282, 2414, 476, 2373, 785]. Use [2555, 1284, 2107, 1067, 294, 998, 2737, 1977, 2568, 237, 82, 1273, 404]. Used [800, 698, 692]. Useful [1660]. Useless [1499]. User [1733, 79, 2442, 2425, 1617, 2417, 191, 83, 2420, 1739, 2295, 1235, 217, 1502, 463, 237, 1318, 2436, 1319, 936, 1698, 1321, 1320, 1596, 279, 1183, 2426, 1322, 2315]. User-Controlled [2426]. User-Level [1596]. User-Steered [936]. Users [2442, 767, 585, 1274, 77, 235]. Uses [2418]. Using [2748, 891, 494, 266, 884, 733, 553, 1538, 2853, 968, 1832, 2686, 2208, 149, 593, 2799, 819, 2281, 940, 1038, 2469, 1140, 2470, 902, 353, 2870, 833, 2801, 245, 264, 326, 2067, 1626, 552, 1085, 581, 938, 1484, 2311, 743, 452, 779, 1039, 106, 2738, 31, 1266, 259, 1958, 2820, 2226, 506, 904, 2434, 978, 172, 847, 1497, 1941, 96, 2059, 2697, 807, 2401, 2776, 102, 1893, 821, 459, 1565, 292, 2312, 2732, 1499, 1944, 347, 2645, 257, 827, 1056, 1914, 977, 1349, 1101, 2480, 2592, 2859, 1164, 729, 509, 1020, 1054, 2817, 2883, 920, 1964, 2488, 310, 1245, 1525, 2135, 720]. Using [38, 239, 306, 249, 431, 893, 879, 1916, 2255, 2678, 881, 1857, 1421, 2242, 516, 717, 2388, 1917, 2519, 399, 118, 975, 1677, 1319, 2212, 601, 751, 2827, 798, 104, 488, 2404, 218, 309, 1667, 177, 844, 1109, 523, 2818, 776, 335, 1740, 1594, 179, 989, 2599, 526, 828, 997, 1923, 824, 1468, 1115, 954, 528, 2041, 98, 151, 1231, 2675, 363, 153, 1022, 1536, 960, 2141, 457, 2377, 2198, 2856, 2500, 2886, 820, 2502, 2800, 146, 235, 2259, 2046, 583, 2804, 55, 965, 445, 1055, 1044, 2882, 1728, 533, 231, 2262, 158, 781, 915, 312, 447, 230, 1465, 2323]. Using [797, 2096, 324, 1339, 2516, 2267, 357, 2406, 280, 898, 109]. Utility [694, 791]. Utilization [1257, 2451].

Vague [682]. Vaguely [686]. Validate [847]. Validated [1122]. Validation [550, 2815, 860, 876, 2620, 513, 269, 2813, 1045, 849, 528, 1016, 999, 875, 435, 952, 1182]. Valmet [876]. Valuations [1217]. Value [1592, 1040, 612, 530]. Valued [2833, 1472, 738, 684, 747, 745, 758]. Values [1860]. Variability [535]. Variable [1885, 268, 476, 2373, 1857, 1607, 1232]. Variables [2689, 1107, 58, 2221]. Variant [904]. Variants [1796]. Variations [1535]. Varieties [2021]. Various [2048]. Varying [37]. Vascular [925, 912, 908, 997, 533, 868]. Vasculature [898]. VBR [1237]. VCR [1170]. Vector [1469, 1059, 1048, 132, 1402, 1468, 133, 530, 274, 368]. Vehicle [102, 2779, 238, 2372, 122, 1346, 1141]. Ventricle [1678, 1011, 986, 872, 1017, 1674, 1690, 907]. Ventricles [1013]. Ventricular [821, 1680, 535]. Verb [2860]. Verbalizing [2784]. Verifiability [2113]. Verifiable [2720, 2766]. Verification [1125, 2068, 1865, 743, 2015, 1908, 2794, 2450, 1934, 543, 2012, 2074, 2671, 2652, 2653, 2654, 167, 2081, 1648, 363, 1763, 2810, 2811, 781, 436, 368, 246, 357]. Verified [2069]. Verifier [2220]. Verify [1199]. Verifying [2697, 2698, 1638]. Version [2857, 2665, 2142]. Versions [2643]. versus [1969, 1768, 818, 1946, 2877, 894]. Vertex [366, 655, 1983, 111, 1354]. Vertex-Based [1983]. Verticality [704]. Vertically [1281]. Vertices [1565, 1542]. Very [1081, 2136, 735, 1743]. Vessel [916, 101, 836, 959, 915, 873]. Vessels [1671]. VEST [1144]. VHML [2300]. via [940, 365, 1887, 131, 608, 610, 1913, 2343, 2271, 2273, 2229, 522, 2847, 2041, 477, 1529, 634, 2233, 359]. Viability [1684]. Vibration [1023]. Video [1445, 216, 1237, 228, 350, 358, 242, 325, 276, 1171, 261, 262, 330, 321, 268, 344, 288, 333, 238, 217, 328, 337, 1063, 1170, 1238, 1240, 273, 345, 2240, 240, 229, 2242, 855, 297, 1444, 338, 2418, 1231, 1232, 1233, 277, 300, 327, 1151, 305, 227, 299, 1247, 224]. Video-Assisted [855]. Video-Endoscopy [1063]. Video-on-Demand [1171, 345]. Videography [861]. VideoSpace [2418]. View [1492, 992, 187, 137, 152, 2210, 551, 2358, 2783, 4, 1900]. View-Based [187]. Viewer [1064]. Viewpoint [2038]. Viewpoints [2559, 2568, 235]. Views [127, 2906, 963]. Violations [1868]. Virtual [891, 1754, 80, 81, 408, 385, 79, 1011, 593, 1038, 84, 400, 962, 2355, 75, 688, 389, 395, 83, 1589, 946, 407, 889, 74, 254, 398, 401, 82, 815, 387, 2334, 218, 1066, 406, 87, 73, 404, 1588, 405, 85, 1527, 72, 835, 892, 88, 1037, 2419, 1061]. Virus [54]. Visemes [153]. Visibility [1297]. Vision [1262, 411, 219, 516, 154, 143, 138]. Vision-Based [1262, 138]. VISIONS [388]. visited [2475]. Visitors [473]. Visual [2691, 119, 1787, 2467, 468, 746, 708, 1020, 239, 2310, 698, 1283, 816, 279, 231, 322]. Visualisation [1059, 1100]. Visualization [897, 962, 1064, 1062, 219, 484, 2325, 958, 1028, 836, 975, 527, 1464, 1490, 435, 892, 416, 873, 1061, 898]. Visualizing [472, 1509, 428]. Viterbi [2647]. Vitro [849]. Vivo [1676, 1683, 930]. VizWear [219]. VLIW [1591]. VLSI [1093, 1092]. VML [1492]. Vocabulary [2871, 233]. Vocal [2303]. VoD [304]. Voice [808, 293, 1243]. Voicemail [2735]. VoIP [293]. Volume [355, 834, 821, 1015, 977, 975, 913, 839, 1008]. Volumes [1682, 1061]. Volumetric [1039, 1049]. Volumetry [1013]. Voluming [1014]. Voting [544, 1778]. Voxel [984]. Voxels [975]. vs [2303, 753, 617, 2427, 2035, 2623]. VSS [2806]. Vulnerability [206].

Walking [1101]. Walks [2059, 2598]. Wall [1678, 1680, 1677]. WAP [1272]. WAP-Based [1272]. Warehouse [1075, 1431, 1433]. Warehousing [1734]. WASABI [2739]. WASH [2531]. WASH/CGI [2531]. Watch [214, 1097]. Watermark [295]. Watermarking [266, 1619, 313, 294, 317, 2235, 347, 337, 314, 336, 2112, 296, 315, 309, 1667, 549, 566, 343, 265, 312]. Watermarks [297]. WAV [312]. WAV-Table [312]. Wave [1670, 1720, 751, 915]. Waveform [2266]. Waveforms [849]. Wavelength [1758, 1605]. Wavelet [267, 2263, 2281, 2246, 2231, 2235, 2249, 2245, 2244, 252, 2241, 257, 2254, 120, 2252, 511, 314, 2278, 2268, 2250, 2279, 2269, 2240, 270, 2255, 2265, 2242, 717, 315, 2277, 2257, 2275, 2234, 2248, 2266, 2253, 2259, 919, 2789, 2262, 2233, 2284, 2247, 2270, 2283, 2267, 2280]. Wavelet-Based [2252, 2240, 2242, 2277, 2234]. Wavelet-Domain [314, 2259]. Wavelet-Packet [2250]. Wavelets [2276, 2282, 2239, 2243, 2264, 2285]. Way [2642, 1768, 2714, 2765, 1366, 485, 615, 1781, 2784, 1466]. Wayfinding [705]. Wayness [2757]. WCET [1149]. WDM [1605]. Weak [1943, 205, 2002]. Weaker [2839]. Weakest [1221]. Weaknesses [2582]. Wear [79]. Wearable [216, 214, 217, 219, 215]. Weather [2804, 60]. Web [417, 1451, 430, 1745, 2217, 412, 570, 2435, 450, 451, 1256, 1449, 452, 481, 1958, 474, 2296, 478, 1332, 482, 411, 1522, 442, 292, 604, 413, 507, 1608, 434, 460, 484, 455, 444, 2317, 476, 2373, 2390, 420, 486, 1511, 2297, 488, 489, 2298, 475, 466, 439, 522, 647, 1166, 415, 2294, 2041, 1450, 1590, 440, 457, 435, 588, 433, 1528, 1939, 2531, 445, 1157, 467, 458, 416, 446, 447, 471, 410, 1722]. Web-Aware [484]. Web-Based [1451, 430, 481, 478, 413, 434, 486, 488, 457, 435, 482]. Web-Enabled [507]. Web-Like [2041]. Web-Page [445]. WEBARM [1256]. Website [1328]. Websites [2301]. Weierstraß [2764]. Weight [2612, 2623]. Weighted [1375, 350, 721, 143, 1359, 1386]. Weights [2374, 1399, 1548, 1549, 1394, 1543, 2119]. Weil [2160, 2116, 2161]. Weld [1591]. Well [2208, 1835, 1215, 426]. Well-Behaved [426]. Well-Founded [2208, 1835]. Well-Quasi-Orders [1215]. Where [140, 592]. Which [1956, 2426, 383]. Who [79, 592]. Whole [1439, 1413, 961]. WI [410]. Wide [2629, 427, 1589, 1498, 1967, 455, 420, 447]. Widely [1284]. Widenings [2013]. Widget [2414]. Width [657, 660, 771]. Will [2520, 1218]. Winding [2790]. Window [1620]. Windows [907]. Winners [1387]. Wireless [578, 293, 2634, 1175, 1177, 768, 600, 461, 1248, 302]. within [644, 1641, 2457, 586, 2460, 1690, 66, 606]. Without [1042, 1618, 987, 1725, 671, 1683, 2026, 2228, 1531]. Women [785]. Word [2853, 2887, 2855, 2852, 56, 2874, 291]. WordNet [2853]. Words [1191, 793, 2883, 2877, 2917]. Work [1592, 2924, 1424, 2176, 405, 2451, 440, 1143]. Worker [2307]. Workflow [26, 1448, 1807, 187, 188, 1446]. Working [2442, 1828]. Workload [1340]. Works [1403, 935]. Workshop [1496, 431, 2357]. Workstation [962]. World [596, 2354, 2489, 2391, 1510, 218, 1100, 455, 420, 447]. World-Wide [447]. Worlds [2057, 593, 2578]. Worst [1517]. Worst-Case [1517]. Worth [1235]. Wrapper [2445]. Wrapping [2496]. Wristwatch [599]. Writer [100]. Writing [1442, 2214]. WSD [2856, 2854]. WTLS [1631]. WWW [1064, 419, 479, 470, 690]. WYSIWYG [2424].

X [849, 973]. X-ray [849, 973]. XML [561, 418, 777, 1440, 1627, 2532, 1452, 1318, 1421, 2316, 1455, 1419, 437]. XML-Based [418, 1318, 2316, 1455, 1419]. XML-schema [1452]. XTG [2028]. XTR [2104, 2604, 1561, 2138]. XXIst [403].

Yagi [1108]. Yagi-Uda [1108]. Yasukawa [766]. Years [1786]. Yima [1082].

zag [2588]. Zero [2155, 2154]. Zero-Knowledge [2155, 2154]. Zerotree [2241]. Zig [2588]. Zig-zag [2588]. Zones [2284]. Zygomatic [952].

References

Rombach:2001:GSQ

[1] Dieter Rombach. Guaranteed software quality. Lecture Notes in Computer Science, 2188:1–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2188/21880001.
htm;   http://link.springer-ny.

com/link/service/series/0558/papers/
2188/21880001.pdf.

Basili:2001:BEBa

[2] Victor Basili. Building an experience base for software engineering: a report on the first CeBASE eWorkshop. Lecture Notes in Computer Science, 2188: 3–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2188/21880003.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2188/21880003.
pdf.

Mellis:2001:CAS

[3] Werner Mellis. A contingency approach to software development. Lecture Notes in Computer Science, 2188: 4–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2188/21880004.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2188/21880004.
pdf.

Shaw:2001:CLE

[4] Mary Shaw. Career-long education for software professionals: a US view of the educational challenges in a rapidly-changing technology. Lecture Notes in Computer Science, 2188:5– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2188/21880005.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2188/21880005.
pdf.

Wong:2001:CSS

[5] Bernard Wong and Ross Jeffery. Cognitive structures of software evaluation: a means-end chain analysis of quality. Lecture Notes in Computer Science, 2188:6–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2188/21880006.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2188/21880006.
pdf.

Anderson:2001:REP

[6] Stuart Anderson and Massimo Felici. Requirements evolution from process to product oriented management. Lecture Notes in Computer Science, 2188: 27–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2188/21880027.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2188/21880027.
pdf.

Nedstam:2001:CSS

[7] Josef Nedstam, Martin Höst, Björn Regnell, and Jennie Nilsson. A case study on scenario-based process flexibility assessment for risk reduction. Lecture Notes in Computer Science, 2188:42–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2188/21880042.htm;   http:

                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2188/21880042.
                                        pdf.

Seppanen:2001:CKB

[8] V. Seppänen, S. Komi-Sirviö, A. Mäntyniemi, T. Rahikkala, and M. Pikkarainen. Contexts of KM based SPI: The case of software design experience reuse. Lecture Notes in Computer Science, 2188: 57–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2188/21880057.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2188/21880057.
                                        pdf.

Lepasaar:2001:MSF

[9] Marion Lepasaar, Timo Varkoi, and Hannu Jaakkola. Models and success factors of process change. Lecture Notes in Computer Science, 2188:68– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2188/21880068.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2188/21880068.
                                        pdf.

Lebsanft:2001:PIT

[10] Karl Lebsanft. Process improvement in turbulent times — is CMM still an answer ? Lecture Notes in Computer Science, 2188:78–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2188/21880078.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2188/21880078.
pdf.

Polo:2001:AMM

[11] Macario Polo, Mario Piattini, Francisco Ruiz, and Mar Jiménez. Assessment of maintenance maturity in IT departments of public entities: Two case studies. Lecture Notes in Computer Science, 2188:86– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2188/21880086.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2188/21880086.
pdf.

Iivari:2001:EUC

[12] Netta Iivari and Timo Jokela. Evaluating a usability capability assessment. Lecture Notes in Computer Science, 2188:98–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2188/21880098.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2188/21880098.
pdf.

Basili:2001:BEBb

[13] Victor Basili, Roseanne Tesoriero, Patricia Costa, Mikael Lindvall, Ioana Rus, Forrest Shull, and Marvin Zelkowitz. Building an experience base for software engineering: a report on the first CeBASE eWorkshop. Lecture Notes in Computer Science, 2188:110– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2188/21880110.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2188/21880110.
                                        pdf.

Schneider:2001:EMA

[14] Kurt Schneider. Experience magnets: Attracting experiences, not just storing them. Lecture Notes in Computer Science, 2188:126–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2188/21880126.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2188/21880126.
                                        pdf.

Kucza:2001:IKM

[15] Timo Kucza, Minna Nättinen, and Päivi Parviainen. Improving knowledge management in software reuse process. Lecture Notes in Computer Science, 2188:141–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2188/21880141.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2188/21880141.
                                        pdf.

Nagel:2001:PKM

[16] Christof Nagel. Processes and knowledge management: a symbiosis. Lecture Notes in Computer Science, 2188: 153–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2188/21880153.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2188/21880153.
pdf.

Dingsoyr:2001:AER

[17] Torgeir Dingsøyr, Nils Brede Moe, and Øystein Nytrø. Augmenting experience reports with lightweight postmortem reviews. Lecture Notes in Computer Science, 2188:167–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2188/21880167.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2188/21880167.
pdf.

Henninger:2001:TMS

[18] Scott Henninger and Jason Schlabach. A tool for managing software development knowledge. Lecture Notes in Computer Science, 2188:182–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2188/21880182.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2188/21880182.
pdf.

Kauppinen:2001:SIR

[19] Marjo Kauppinen and Sari Kujala. Starting improvement of requirements engineering processes: An experience report. Lecture Notes in Computer Science, 2188:196–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2188/21880196.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2188/21880196.
                                        pdf.

Allenby:2001:FOS

[20] Karen Allenby, Simon Burton, Darren Buttle, John McDermid, John Murdoch, Alan Stephenson, Mike Bardill, and Stuart Hutchesson. A familyoriented software development process for engine controllers. Lecture Notes in Computer Science, 2188:210– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2188/21880210.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2188/21880210.
                                        pdf.

Frohlich:2001:ELS

[21] Peter Fröhlich, Horst Lichter, and Manfred Zeller. Enabling local SPI in a multi-national company. Lecture Notes in Computer Science, 2188:227– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2188/21880227.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2188/21880227.
                                        pdf.

Dekkers:2001:PIS

[22] Ton Dekkers. Project improvement as start-up. Lecture Notes in Computer Science, 2188:240–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2188/21880240.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2188/21880240.
pdf.

Zettel:2001:LLP

[23] Jörg Zettel, Frank Maurer, Jürgen Münch, and Les Wong. LIPE: a lightweight process for E-business startup companies based on extreme programming. Lecture Notes in Computer Science, 2188:255–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2188/21880255.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2188/21880255.
pdf.

Jaccheri:2001:EEP

[24] M. Letizia Jaccheri and Tor Stålhane. Evaluation of the E3 process modelling language and tool for the purpose of model creation. Lecture Notes in Computer Science, 2188:271–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2188/21880271.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2188/21880271.
pdf.

Storrle:2001:DFP

[25] Harald Störrle. Describing fractal processes with UML. Lecture Notes in Computer Science, 2188:282–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2188/21880282.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2188/21880282.
                                        pdf.

Araujo:2001:ESP

[26] Renata Araujo and Marcos Borges. Extending the software process culture — an approach based on groupware and workflow. Lecture Notes in Computer Science, 2188:297–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2188/21880297.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2188/21880297.
                                        pdf.

Becker-Kornstaedt:2001:TSK

[27] Ulrike Becker-Kornstaedt. Towards systematic knowledge elicitation for descriptive software process modeling. Lecture Notes in Computer Science, 2188:312–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2188/21880312.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2188/21880312.
                                        pdf.

Gnatz:2001:MPP

[28] Michael Gnatz, Frank Marschall, Gerhard Popp, Andreas Rausch, and Wolfgang Schwerin. Modular process patterns supporting an evolutionary software development process. Lecture Notes in Computer Science, 2188:326– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2188/21880326.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2188/21880326.
pdf.

Runeson:2001:CSS

[29] Per Runeson, Magnus C. Ohlsson, and Claes Wohlin. A classification scheme for studies on faultprone components. Lecture Notes in Computer Science, 2188:341–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2188/21880341.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2188/21880341.
pdf.

Bratthall:2001:PUB

[30] Lars Bratthall, Erik Arisholm, and Magne Jørgensen. Program understanding behavior during estimation of enhancement effort on small Java programs. Lecture Notes in Computer Science, 2188:356–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2188/21880356.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2188/21880356.
pdf.

Fujiwara:2001:EBA

[31] Hikaru Fujiwara, Shinji Kusumoto, Katsuro Inoue, Toshifusa Ootsubo, and Katsuhiko Yuura. Evaluation of a business application framework using complexity and functionality metrics. Lecture Notes in Computer Science, 2188: 371–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2188/21880371.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2188/21880371.
                                        pdf.

Anonymous:2001:AIa

[32] Anonymous. Author index. Lecture Notes in Computer Science, 2188:381– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2188/2188auth.pdf.

Bradley:2001:FCS

[33] Elizabeth Bradley, Nancy Collins, and W. Philip Kegelmeyer. Feature characterization in scientific datasets. Lecture Notes in Computer Science, 2189: 1–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2189/21890001.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2189/21890001.
                                        pdf.

deCampos:2001:RFB

[34] Luis M. de Campos, Juan M. Fernández-Luna, and Juan F. Huete.

Relevance feedback in the Bayesian network retrieval model: An approach based on term instantiation. Lecture Notes in Computer Science, 2189:13– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2189/21890013.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2189/21890013.
pdf.

Laurent:2001:GFS

[35] Anne Laurent. Generating fuzzy summaries from fuzzy multidimensional databases. Lecture Notes in Computer Science, 2189:24–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2189/21890024.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2189/21890024.
pdf.

Estabrooks:2001:MEF

[36] Andrew Estabrooks and Nathalie Japkowicz. A mixture-of-experts framework for learning from imbalanced data sets. Lecture Notes in Computer Science, 2189:34–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2189/21890034.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2189/21890034.
pdf.

Lewandowski:2001:PTV

[37] Achim Lewandowski and Peter Protzel. Predicting time-varying functions with local models. Lecture Notes in Computer Science, 2189:44–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2189/21890044.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2189/21890044.
                                        pdf.

Li:2001:BME

[38] Cen Li, Gautam Biswas, Mike Dale, and Pat Dale. Building models of ecological dynamics using HMM based temporal data clustering — A preliminary study. Lecture Notes in Computer Science, 2189:53–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2189/21890053.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2189/21890053.
                                        pdf.

Marques:2001:TST

[39] Nuno C. Marques and Gabriel Pereira Lopes. Tagging with small training corpora. Lecture Notes in Computer Science, 2189:63–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2189/21890063.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2189/21890063.pdf.

Hahn:2001:SEM

[40] Udo Hahn, Martin Honeck, and Stefan Schulz. A search engine for morphologically complex languages. Lecture Notes in Computer Science, 2189:73– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2189/21890073.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2189/21890073.
pdf.

Bruni:2001:EDC

[41] Renato Bruni and Antonio Sassano. Errors detection and correction in large scale data collecting. Lecture Notes in Computer Science, 2189:84– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2189/21890084.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2189/21890084.
pdf.

Berzal:2001:NFA

[42] Fernando Berzal, Ignacio Blanco, Daniel Sánchez, and María-Amparo Vila. A new framework to assess association rules. Lecture Notes in Computer Science, 2189:95–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2189/21890095.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2189/21890095.
pdf.

Cortes:2001:CI

[43] Corinna Cortes, Daryl Pregibon, and Chris Volinsky. Communities of interest. Lecture Notes in Computer Science, 2189:105–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2189/21890105.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2189/21890105.
                                        pdf.

Seewald:2001:EGC

[44] Alexander K. Seewald and Johannes Fürnkranz. An evaluation of grading classifiers. Lecture Notes in Computer Science, 2189:115–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2189/21890115.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2189/21890115.
                                        pdf.

Hoppner:2001:FIR

[45] Frank Höppner and Frank Klawonn. Finding informative rules in interval sequences. Lecture Notes in Computer Science, 2189:125–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2189/21890125.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2189/21890125.
                                        pdf.

Puuronen:2001:CBC

[46] Seppo Puuronen, Alexey Tsymbal, and Iryna Skrypnyk. Correlationbased and contextual merit-based ensemble feature selection. Lecture Notes in Computer Science, 2189:135– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2189/21890135.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2189/21890135.
pdf.

vanWezel:2001:NMS

[47] Michiel C. van Wezel, Walter A. Kosters, Peter van der Putten, and Joost N. Kok. Nonmetric multidimensional scaling with neural networks. Lecture Notes in Computer Science, 2189:145–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2189/21890145.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2189/21890145.
pdf.

Gama:2001:FTR

[48] João Gama. Functional trees for regression. Lecture Notes in Computer Science, 2189:156–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2189/21890156.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2189/21890156.pdf.

Ferreira:2001:DMP

[49] José Tomé A. S. Ferreira, David G. T. Denison, and David J. Hand. Data mining with products of trees. Lecture Notes in Computer Science, 2189: 167–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2189/21890167.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2189/21890167.
pdf.

Terabe:2001:BFC

[50] Masahiro Terabe, Takashi Washio, and Hiroshi Motoda. S^{3} bagging: Fast classifier induction method with subsampling and bagging. Lecture Notes in Computer Science, 2189:177–186, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2189/21890177.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2189/21890177.
pdf.

Backofen:2001:RSS

[51] Rolf Backofen. RNA-sequence-structure properties and selenocysteine insertion. Lecture Notes in Computer Science, 2189:187–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2189/21890187.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2189/21890187.
pdf.

Cohen:2001:ASC

[52] Paul Cohen and Niall Adams. An algorithm for segmenting categorical time series into meaningful episodes. Lecture Notes in Computer Science, 2189: 198–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2189/21890198.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2189/21890198.
pdf.

Windeatt:2001:ECP

[53] Terry Windeatt and Gholamreza Ardeshir. An empirical comparison of pruning methods for ensemble classifiers. Lecture Notes in Computer Science, 2189:208–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2189/21890208.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2189/21890208.
pdf.

Kellam:2001:FMS

[54] Paul Kellam, Xiaohui Liu, Nigel Martin, Christine Orengo, Stephen Swift, and Allan Tucker. A framework for modelling short, high-dimensional multivariate time series: Preliminary results in virus gene expression data analysis. Lecture Notes in Computer Science, 2189:218–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2189/21890218.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2189/21890218.
                                        pdf.

Trivino-Rodriguez:2001:UMP

[55] José L. Triviño-Rodriguez and Rafael Morales-Bueno. Using multiattribute prediction suffix graphs for Spanish part-of-speech tagging. Lecture Notes in Computer Science, 2189:228– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2189/21890228.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2189/21890228.
                                        pdf.

Peng:2001:SSC

[56] Fuchun Peng and Dale Schuurmans. Self-supervised Chinese word segmentation. Lecture Notes in Computer Science, 2189:238–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2189/21890238.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2189/21890238.
                                        pdf.

Abidi:2001:ADC

[57] Syed Sibte Raza Abidi, Kok Meng Hoe, and Alwyn Goh. Analyzing data clusters: a rough sets approach to extract cluster-defining symbolic rules. Lecture Notes in Computer Science, 2189:248– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2189/21890248.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2189/21890248.
pdf.

Nakano:2001:FPF

[58] Ryohei Nakano and Kazumi Saito. Finding polynomials to fit multivariate data having numeric and nominal variables. Lecture Notes in Computer Science, 2189:258–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2189/21890258.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2189/21890258.
pdf.

Cohen:2001:FLE

[59] Paul R. Cohen. Fluent learning: Elucidating the structure of episodes. Lecture Notes in Computer Science, 2189: 268–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2189/21890268.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2189/21890268.
pdf.

Viademonte:2001:IDS

[60] Sérgio Viademonte and Frada Burstein. An intelligent decision support model for aviation weather forecasting. Lecture Notes in Computer Science, 2189: 278–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2189/21890278.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2189/21890278.
                                        pdf.

Castelo:2001:MDA

[61] Robert Castelo, Ad Feelders, and Arno Siebes. MAMBO: Discovering association rules based on conditional independencies. Lecture Notes in Computer Science, 2189:289–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2189/21890289.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2189/21890289.
                                        pdf.

Glendinning:2001:MBR

[62] R. H. Glendinning. Model building for random fields. Lecture Notes in Computer Science, 2189:299–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2189/21890299.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2189/21890299.
                                        pdf.

Scheffer:2001:AHM

[63] Tobias Scheffer, Christian Decomain, and Stefan Wrobel. Active hidden Markov models for information extraction. Lecture Notes in Computer Science, 2189:309–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2189/21890309.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2189/21890309.
pdf.

Somlo:2001:ALT

[64] Gabriel L. Somlo and Adele E. Howe. Adaptive lightweight text filtering. Lecture Notes in Computer Science, 2189: 319–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2189/21890319.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2189/21890319.
pdf.

Hongwei:2001:GAA

[65] Zhang Hongwei, Tian Fengzhan, and Lu Yuchang. A general algorithm for approximate inference in multiply sectioned Bayesian networks. Lecture Notes in Computer Science, 2189:330– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2189/21890330.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2189/21890330.
pdf.

Yearling:2001:ITP

[66] Dave Yearling and David J. Hand. Investigating temporal patterns of fault behaviour within large telephony networks. Lecture Notes in Computer Science, 2189:340–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2189/21890340.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2189/21890340.
                                        pdf.

Kryszkiewicz:2001:CSB

[67] Marzena Kryszkiewicz. Closed set based discovery of representative association rules. Lecture Notes in Computer Science, 2189:350–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2189/21890350.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2189/21890350.
                                        pdf.

Easley:2001:ISA

[68] Matthew Easley and Elizabeth Bradley. Intelligent sensor analysis and actuator control. Lecture Notes in Computer Science, 2189:360–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2189/21890360.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2189/21890360.
                                        pdf.

Rumantir:2001:SHC

[69] Grace W. Rumantir and Chris S. Wallace. Sampling of highly correlated data for polynomial regression and model discovery. Lecture Notes in Computer Science, 2189:370– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2189/21890370.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2189/21890370.
pdf.

Cohen:2001:IRD

[70] Paul Cohen, Niall Adams, and David J. Hand. The IDA’01 robot data challenge. Lecture Notes in Computer Science, 2189:378–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2189/21890378.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2189/21890378.
pdf.

Anonymous:2001:AIb

[71] Anonymous. Author index. Lecture Notes in Computer Science, 2189:383– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2189/2189indx.pdf.

Thalmann:2001:FBV

[72] Daniel Thalmann. The foundations to build a virtual human society. Lecture Notes in Computer Science, 2190: 1–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2190/21900001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2190/21900001.
pdf.

Rickel:2001:IVA

[73] Jeff Rickel. Intelligent virtual agents for education and training: Opportunities and challenges. Lecture Notes in Computer Science, 2190:15– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2190/21900015.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2190/21900015.
pdf.

Li:2001:EPA

[74] Michael Li and Ted Selker. Eye pattern analysis in intelligent virtual agents. Lecture Notes in Computer Science, 2190:23–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2190/21900023.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2190/21900023.
pdf.

Delgado-Mata:2001:CEV

[75] Carlos Delgado-Mata and Ruth Aylett. Communicating emotion in virtual environments through artificial scents. Lecture Notes in Computer Science, 2190:36–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2190/21900036.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2190/21900036.
pdf.

Aaron:2001:FRA

[76] Eric Aaron, Dimitris Metaxas, Franjo Ivančić, and Oleg Sokolsky. A framework for reasoning about animation systems. Lecture Notes in Computer Science, 2190:47–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2190/21900047.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2190/21900047.
pdf.

Scheutz:2001:AAU

[77] Matthias Scheutz and Brigitte Römmer. Autonomous avatars? from users to agents and back. Lecture Notes in Computer Science, 2190:61–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2190/21900061.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2190/21900061.
pdf.

Chen:2001:ELA

[78] L. Chen, K. Bechkoum, and G. Clapworthy. Equipping a lifelike animated agent with a mind. Lecture Notes in Computer Science, 2190:72– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2190/21900072.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2190/21900072.
pdf.

Bailenson:2001:IAW

[79] Jeremy N. Bailenson, Andrew C. Beall, Jim Blascovich, Mike Raimundo, and Max Weisbuch. Intelligent agents who wear your face: Users’ reactions to the virtual self. Lecture Notes in Computer Science, 2190:86– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2190/21900086.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2190/21900086.
pdf.

Amo:2001:IVA

[80] F. Alonso Amo, F. Fernández Velasco, G. López Gómez, J. P. Rojas Jiménez, and F. J. Soriano Camino. Intelligent virtual agent societies on the Internet. Lecture Notes in Computer Science, 2190:100– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2190/21900100.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2190/21900100.
pdf.

Anastassakis:2001:VAS

[81] George Anastassakis, Themis Panayiotopoulos, and Tim Ritchings. Virtual agent societies with the mVITAL intelligent agent system. Lecture Notes in Computer Science, 2190:112–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2190/21900112.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2190/21900112.
pdf.

Mendez:2001:OUM

[82] Gonzalo Méndez, Pedro Pérez, and Angélica de Antonio. An overview of the use of mobile agents in virtual environments. Lecture Notes in Computer Science, 2190:126–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2190/21900126.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2190/21900126.
pdf.

Gerhard:2001:CPC

[83] M. Gerhard, D. J. Moore, and D. J. Hobbs. Continuous presence in collaborative virtual environments: Towards a hybrid avatar-agent model for user representation. Lecture Notes in Computer Science, 2190:137–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2190/21900137.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2190/21900137.
pdf.

Cavazza:2001:AIV

[84] Marc Cavazza, Fred Charles, and Steven J. Mead. Agents’ interaction in virtual storytelling. Lecture Notes in Computer Science, 2190:156– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2190/21900156.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2190/21900156.
                                        pdf.

Silva:2001:PVS

[85] André Silva, Marco Vala, and Ana Paiva. Papous: The virtual storyteller. Lecture Notes in Computer Science, 2190:171–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2190/21900171.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2190/21900171.
                                        pdf.

Klesen:2001:DAM

[86] Martin Klesen, Janek Szatkowski, and Niels Lehmann. A dramatised actant model for interactive improvisational plays. Lecture Notes in Computer Science, 2190:181–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2190/21900181.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2190/21900181.
                                        pdf.

Richard:2001:ITD

[87] Nadine Richard, Philippe Codognet, and Alain Grumbach. The InViWo toolkit: Describing autonomous virtual agents and avatars. Lecture Notes in Computer Science, 2190:195– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2190/21900195.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2190/21900195.
pdf.

Vosinakis:2001:SPR

[88] Spyros Vosinakis and Themis Panayiotopoulos. SimHuman: a platform for real-time virtual agents with planning capabilities. Lecture Notes in Computer Science, 2190:210–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2190/21900210.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2190/21900210.
pdf.

Casella:2001:MAR

[89] Pietro Casella and Ana Paiva. MAgentA: An architecture for real time automatic composition of background music. Lecture Notes in Computer Science, 2190:224–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2190/21900224.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2190/21900224.pdf.

deLara:2001:AOS

[90] Juan de Lara and Manuel Alfonseca. Agent oriented simulation with OOCSMP. an example in evolutionary ant colonies. Lecture Notes in Computer Science, 2190:233–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2190/21900233.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2190/21900233.
                                        pdf.

Poggi:2001:LAG

[91] Isabella Poggi. The lexicon and the alphabet of gesture, gaze, and touch. Lecture Notes in Computer Science, 2190: 235–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2190/21900235.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2190/21900235.
                                        pdf.

Hoshino:2001:ERP

[92] Junichi Hoshino. Extraction and reconstruction of personal characters from human movement. Lecture Notes in Computer Science, 2190:237–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2190/21900237.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2190/21900237.
                                        pdf.

Walshe:2001:OSL

[93] Ray Walshe. The origin of the speeches: Language evolution through collaborative reinforcement learning. Lecture Notes in Computer Science, 2190:239– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2190/21900239.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2190/21900239.
pdf.

Paradiso:2001:FTA

[94] Aldo Paradiso, Fabio Zambetta, and Fabio Abbattista. Fanky: a tool for animating faces of 3D agents. Lecture Notes in Computer Science, 2190:242– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2190/21900242.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2190/21900242.
pdf.

Anonymous:2001:AIc

[95] Anonymous. Author index. Lecture Notes in Computer Science, 2190:245– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2190/2190auth.pdf.

Hanek:2001:MBI

[96] Robert Hanek. Model-based image segmentation using local selfadapting separation criteria. Lecture Notes in Computer Science, 2191:1– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2191/21910001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2191/21910001.
pdf.

Siggelkow:2001:SSI

[97] Sven Siggelkow, Marc Schael, and Hans Burkhardt. SIMBA — Search IMages By Appearance. Lecture Notes in Computer Science, 2191:9– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2191/21910009.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2191/21910009.
pdf.

Schael:2001:TDD

[98] Marc Schael. Texture defect detection using invariant textural features. Lecture Notes in Computer Science, 2191: 17–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2191/21910017.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2191/21910017.
pdf.

Massad:2001:SSS

[99] Amin Massad and Bärbel Mertsching. Segmentation of spontaneously splitting figures into overlapping layers. Lecture Notes in Computer Science, 2191:25–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2191/21910025.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2191/21910025.
pdf.

Brakensiek:2001:WAO

[100] Anja Brakensiek, Andreas Kosmala, and Gerhard Rigoll. Writer adaptation for online handwriting recognition. Lecture Notes in Computer Science, 2191:

32–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2191/21910032.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2191/21910032.
pdf.

Jiang:2001:BVD

[101] Xiaoyi Jiang and Daniel Mojon. Blood vessel detection in retinal images by shape-based multi-threshold probing. Lecture Notes in Computer Science, 2191:38–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2191/21910038.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2191/21910038.
pdf.

Hinz:2001:VDA

[102] Stefan Hinz and Albert Baumgartner. Vehicle detection in aerial images using generic features, grouping, and context. Lecture Notes in Computer Science, 2191:45–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2191/21910045.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2191/21910045.
pdf.

Groger:2001:RIS

[103] Martin Gröger, Wolfgang Sepp, Tobias Ortmaier, and Gerd Hirzinger. Reconstruction of image structure in presence of specular reflections. Lecture Notes in Computer Science, 2191:53– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2191/21910053.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2191/21910053.
                                        pdf.

Neumann:2001:STA

[104] Jan Neumann and Yiannis Aloimonos. Spatio-temporal analysis of human faces using multi-resolution subdivision surfaces. Lecture Notes in Computer Science, 2191:61–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2191/21910061.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2191/21910061.
                                        pdf.

Scharr:2001:ADF

[105] Hanno Scharr and Dietmar Uttenweiler. 3D anisotropic diffusion filtering for enhancing noisy actin filament fluorescence images. Lecture Notes in Computer Science, 2191:69– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2191/21910069.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2191/21910069.
                                        pdf.

Fieres:2001:PSR

[106] J. Fieres, J. Mattes, and R. Eils. A point set registration algorithm using a motion model based on thin-plate splines and point clustering. Lecture Notes in Computer Science, 2191:76– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2191/21910076.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2191/21910076.
pdf.

Luxen:2001:OCO

[107] Marc Luxen and Wolfgang Förstner. Optimal camera orientation from points and straight lines. Lecture Notes in Computer Science, 2191:84– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2191/21910084.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2191/21910084.
pdf.

Heuel:2001:PLP

[108] Stephan Heuel. Points, lines, and planes and their optimal estimation. Lecture Notes in Computer Science, 2191:92–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2191/21910092.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2191/21910092.
pdf.

vonWichert:2001:PAS

[109] Georg von Wichert. A probabilistic approach to simultaneous segmentation, object recognition, 3D localization, and tracking using stereo. Lecture Notes in Computer Science, 2191: 100–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2191/21910100.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2191/21910100.
pdf.

Alker:2001:IREa

[110] Manfred Alker, Sönke Frantz, Karl Rohr, and H. Siegfried Stiehl. Improving the robustness in extracting 3D point landmarks based on deformable models. Lecture Notes in Computer Science, 2191:108–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2191/21910108.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2191/21910108.
pdf.

Katsoulas:2001:EVD

[111] Dimitrios Katsoulas and Lothar Bergen. Efficient 3D vertex detection in range images acquired with a laser sensor. Lecture Notes in Computer Science, 2191:116–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2191/21910116.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2191/21910116.
pdf.

Felsberg:2001:SAF

[112] Michael Felsberg and Gerald Sommer. Scale adaptive filtering derived from the Laplace equation. Lecture Notes in Computer Science, 2191:124– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2191/21910124.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2191/21910124.
pdf.

Choi:2001:OSS

[113] Sung Woo Choi and Hans-Peter Seidel. One-sided stability of medial axis transform. Lecture Notes in Computer Science, 2191:132–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2191/21910132.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2191/21910132.
pdf.

Feldbach:2001:RLD

[114] M. Feldbach and K. D. Tönnies. Robust line detection in historical church registers. Lecture Notes in Computer Science, 2191:140–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2191/21910140.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2191/21910140.
pdf.

Steger:2001:SMO

[115] Carsten Steger. Similarity measures for occlusion, clutter, and illumination invariant object recognition. Lecture Notes in Computer Science, 2191:148– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2191/21910148.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2191/21910148.
                                        pdf.

Voss:2001:APP

[116] K. Voss and H. Suesse. Affine point pattern matching. Lecture Notes in Computer Science, 2191:155–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2191/21910155.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2191/21910155.
                                        pdf.

Spies:2001:SER

[117] Hagen Spies, Bernd Jähne, and John L. Barron. Surface expansion from range data sequences. Lecture Notes in Computer Science, 2191:163–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2191/21910163.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2191/21910163.
                                        pdf.

Mester:2001:IIC

[118] Rudolf Mester, Til Aach, and Lutz Dümbgen. Illumination-invariant change detection using a statistical colinearity criterion. Lecture Notes in Computer Science, 2191:170–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2191/21910170.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2191/21910170.
pdf.

Bauckhage:2001:MVK

[119] Christian Bauckhage, Jannik Fritsch, and Gerhard Sagerer. Memorizing visual knowledge for assembly process monitoring. Lecture Notes in Computer Science, 2191:178–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2191/21910178.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2191/21910178.
pdf.

Kruger:2001:WSM

[120] Volker Krüger and Rogerio S. Feris. Wavelet subspace method for realtime face tracking. Lecture Notes in Computer Science, 2191:186–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2191/21910186.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2191/21910186.
pdf.

Garbe:2001:RES

[121] Christoph S. Garbe and Bernd Jähne. Reliable estimates of the sea surface heat flux from image sequences. Lecture Notes in Computer Science, 2191: 194–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2191/21910194.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2191/21910194.
                                        pdf.

Techmer:2001:RTM

[122] Axel Techmer. Real-time motion based vehicle segmentation in traffic lanes. Lecture Notes in Computer Science, 2191:202–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2191/21910202.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2191/21910202.
                                        pdf.

Moller:2001:DTM

[123] Birgit Möller and Stefan Posch. Detection and tracking of moving objects for mosaic image generation. Lecture Notes in Computer Science, 2191:208– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2191/21910208.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2191/21910208.
                                        pdf.

Wallraven:2001:ARR

[124] Christian Wallraven and Heinrich Bülthoff. Acquiring robust representations for recognition from image sequences. Lecture Notes in Computer Science, 2191:216–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2191/21910216.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2191/21910216.
pdf.

Wenig:2001:ISA

[125] M. Wenig, C. Leue, S. Kraus, T. Wagner, U. Platt, and B. Jähne. Image sequence analysis of satellite NO_{2} concentration maps. Lecture Notes in Computer Science, 2191:223–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2191/21910223.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2191/21910223.
pdf.

Kirchgessner:2001:RGM

[126] Norbert Kirchgeßner, Hagen Spies, Hanno Scharr, and Uli Schurr. Root growth measurements in object coordinates. Lecture Notes in Computer Science, 2191:231–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2191/21910231.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2191/21910231.
                                        pdf.

Deinzer:2001:FMV

[127] Frank Deinzer, Joachim Denzler, and Heinrich Niemann. On fusion of multiple views for active object recognition. Lecture Notes in Computer Science, 2191:239–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2191/21910239.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2191/21910239.
                                        pdf.

Roth:2001:PDK

[128] Volker Roth. Probabilistic discriminative kernel classifiers for multiclass problems. Lecture Notes in Computer Science, 2191:246–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2191/21910246.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2191/21910246.
                                        pdf.

Reinhold:2001:ABS

[129] Michael Reinhold, Dietrich Paulus, and Heinrich Niemann. Appearancebased statistical object recognition by heterogeneous background and occlusions. Lecture Notes in Computer Science, 2191:254–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2191/21910254.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2191/21910254.
pdf.

Kampfe:2001:CIC

[130] Tanja Kämpfe, Tim W. Nattkemper, and Helge Ritter. Combining independent component analysis and self-organizing maps for cell image classification. Lecture Notes in Computer Science, 2191:262–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2191/21910262.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2191/21910262.
pdf.

Cremers:2001:NSS

[131] Daniel Cremers, Timo Kohlberger, and Christoph Schnörr. Nonlinear shape statistics via kernel spaces. Lecture Notes in Computer Science, 2191:269– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2191/21910269.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2191/21910269.
pdf.

Graf:2001:NSV

[132] Arnulf B. A. Graf and Silvio Borer. Normalization in support vector machines. Lecture Notes in Computer Science, 2191:277–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/

                                        series/0558/bibs/2191/21910277.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2191/21910277.pdf.

Schwenker:2001:SMC

[133] Friedhelm Schwenker. Solving multiclass pattern recognition problems with tree-structured support vector machines. Lecture Notes in Computer Science, 2191:283–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2191/21910283.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2191/21910283.pdf.

Hulsken:2001:FER

[134] Frank Hülsken, Frank Wallhoff, and Gerhard Rigoll. Facial expression recognition with pseudo-3D hidden Markov models. Lecture Notes in Computer Science, 2191:291–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2191/21910291.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2191/21910291.
                                        pdf.

Bohm:2001:TOP

[135] Jan Böhm, Jürgen Hefele, and Dieter Fritsch. Towards online pose measurement for robots. Lecture Notes in Computer Science, 2191:298– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2191/21910298.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2191/21910298.
pdf.

Denzler:2001:OCP

[136] J. Denzler, C. M. Brown, and H. Niemann. Optimal camera parameter selection for state estimation with applications in object recognition. Lecture Notes in Computer Science, 2191:305– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2191/21910305.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2191/21910305.
pdf.

Klein:2001:VPU

[137] Konrad Klein and Vítor Sequeira. View planning for unknown indoor scenes based on a cost benefit analysis. Lecture Notes in Computer Science, 2191: 313–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2191/21910313.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2191/21910313.
pdf.

Schmitt:2001:CPS

[138] Thorsten Schmitt, Robert Hanek, Sebastian Buck, and Michael Beetz. Cooperative probabilistic state estimation for vision-based autonomous soccer robots. Lecture Notes in Computer Science, 2191:321–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2191/21910321.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2191/21910321.pdf.

Suing:2001:NCB

[139] Markus Suing, Lothar Hermes, and Joachim M. Buhmann. A new contourbased approach to object recognition for assembly line robots. Lecture Notes in Computer Science, 2191:329– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2191/21910329.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2191/21910329.
                                        pdf.

Michaelsen:2001:WHP

[140] E. Michaelsen, U. Ahlrichs, U. Stilla, D. Paulus, and H. Niemann. Where is the hole punch? object localization capabilities on a specific bureau task. Lecture Notes in Computer Science, 2191: 337–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2191/21910337.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2191/21910337.
                                        pdf.

Soergel:2001:PBS

[141] Uwe Soergel, Karsten Schulz, and Ulrich Thoennessen. Phenomenologybased segmentation of InSAR data for building detection. Lecture Notes in Computer Science, 2191:345–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2191/21910345.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2191/21910345.
pdf.

Keuchel:2001:CRB

[142] Jens Keuchel, Christian Schellewald, Daniel Cremers, and Christoph Schnörr. Convex relaxations for binary image partitioning and perceptual grouping. Lecture Notes in Computer Science, 2191:353–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2191/21910353.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2191/21910353.
pdf.

Schellewald:2001:ECO

[143] Christian Schellewald, Stefan Roth, and Christoph Schnörr. Evaluation of convex optimization techniques for the weighted graph-matching problem in computer vision. Lecture Notes in Computer Science, 2191:361– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2191/21910361.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2191/21910361.
pdf.

Gavrila:2001:LSM

[144] D. M. Gavrila, J. Giebel, and H. Neumann. Learning shape models from examples. Lecture Notes in Computer Science, 2191:369–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2191/21910369.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2191/21910369.
                                        pdf.

Winkelbach:2001:SEG

[145] S. Winkelbach and F. M. Wahl. Shape from 2D edge gradients. Lecture Notes in Computer Science, 2191:377– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2191/21910377.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2191/21910377.
                                        pdf.

Suppes:2001:ROD

[146] Alexander Suppes, Frank Suhling, and Michael Hötter. Robust obstacle detection from stereoscopic image sequences using Kalman filtering. Lecture Notes in Computer Science, 2191: 385–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2191/21910385.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2191/21910385.
                                        pdf.

Saupe:2001:MRS

[147] Dietmar Saupe and Dejan V. Vranić. 3D model retrieval with spherical harmonics and moments. Lecture Notes in Computer Science, 2191:392–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2191/21910392.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2191/21910392.
pdf.

Ahn:2001:LSO

[148] Sung Joon Ahn, Wolfgang Rauh, and Matthias Recknagel. Least squares orthogonal distance fitting of implicit curves and surfaces. Lecture Notes in Computer Science, 2191:398–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2191/21910398.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2191/21910398.
pdf.

Behrens:2001:STS

[149] Thorsten Behrens, Karl Rohr, and H. Siegfried Stiehl. Segmentation of tubular structures in 3D images using a combination of the Hough transform and a Kalman filter. Lecture Notes in Computer Science, 2191:406– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2191/21910406.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2191/21910406.
pdf.

Faber:2001:PCE

[150] P. Faber and R. B. Fisher. Pros and cons of Euclidean fitting. Lecture Notes in Computer Science, 2191:414– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2191/21910414.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2191/21910414.
                                        pdf.

Schmidt:2001:PAO

[151] Jochen Schmidt, Ingo Scholz, and Heinrich Niemann. Placing arbitrary objects in a real scene using a color cube for pose estimation. Lecture Notes in Computer Science, 2191:421– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2191/21910421.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2191/21910421.
                                        pdf.

Kolesnik:2001:DTI

[152] Marina Kolesnik and Thomas Kuleßa. Detecting, tracking, and interpretation of a pointing gesture by an overhead view camera. Lecture Notes in Computer Science, 2191:429–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2191/21910429.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2191/21910429.
                                        pdf.

Shdaifat:2001:RGV

[153] Islam Shdaifat, Rolf-Rainer Grigat, and Stefan Lütgert. Recognition of the German visemes using multiple feature matching. Lecture Notes in Computer Science, 2191:437–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2191/21910437.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2191/21910437.
pdf.

Navab:2001:KSM

[154] Nassir Navab. Keynote speech: Medical & industrial augmented reality: Challenges for real-time vision, computer graphics and mobile computing. Lecture Notes in Computer Science, 2191:443–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2191/21910443.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2191/21910443.
pdf.

Anonymous:2001:AId

[155] Anonymous. Author index. Lecture Notes in Computer Science, 2191:453– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2191/2191auth.pdf.

Pawlak:2001:JFS

[156] Renaud Pawlak, Lionel Seinturier, Laurence Duchien, and Gérard Florin. JAC: a flexible solution for aspectoriented programming in Java. Lecture Notes in Computer Science, 2192: 1–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2192/21920001.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2192/21920001.
                                        pdf.

Tanter:2001:RTO

[157] Éric Tanter, Noury M. N. Bouraqadi-Saâdani, and Jacques Noyé. Reflex — towards an open reflective extension of Java. Lecture Notes in Computer Science, 2192:25–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2192/21920025.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2192/21920025.
                                        pdf.

Whaley:2001:SCU

[158] John Whaley. System checkpointing using reflection and program analysis. Lecture Notes in Computer Science, 2192:44–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2192/21920044.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2192/21920044.
                                        pdf.

Bruneton:2001:EJP

[159] Eric Bruneton and Michel Riveill. Experiments with JavaPod, a platform designed for the adaptation of non-functional properties. Lecture Notes in Computer Science, 2192:52– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2192/21920052.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2192/21920052.
pdf.

Orleans:2001:DDA

[160] Doug Orleans and Karl Lieberherr. DJ: Dynamic adaptive programming in Java. Lecture Notes in Computer Science, 2192:73–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2192/21920073.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2192/21920073.
pdf.

Dowling:2001:KCA

[161] Jim Dowling and Vinny Cahill. The K-component architecture meta-model for self-adaptive software. Lecture Notes in Computer Science, 2192:81– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2192/21920081.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2192/21920081.
pdf.

Ubayashi:2001:SCM

[162] Naoyasu Ubayashi and Tetsuo Tamai. Separation of concerns in mobile agent applications. Lecture Notes in Computer Science, 2192:89–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2192/21920089.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2192/21920089.
                                        pdf.

Malenfant:2001:DAM

[163] Jacques Malenfant, Maria-Teresa Segarra, and Françoise André. Dynamic adaptability: The MolèNE experiment. Lecture Notes in Computer Science, 2192:110–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2192/21920110.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2192/21920110.pdf.

Caromel:2001:SSA

[164] Denis Caromel, Fabrice Huet, and Julien Vayssière. A simple securityaware MOP for Java. Lecture Notes in Computer Science, 2192:118–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2192/21920118.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2192/21920118.
                                        pdf.

Capra:2001:RMS

[165] Licia Capra, Wolfgang Emmerich, and Cecilia Mascolo. Reflective middleware solutions for contextaware applications. Lecture Notes in Computer Science, 2192:126–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2192/21920126.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2192/21920126.
pdf.

Ruiz:2001:TMP

[166] Juan Carlos Ruiz, Jean-Charles Fabre, and Pascale Thévenod-Fosse. Testing MetaObject protocols generated by open compilers for safetycritical systems. Lecture Notes in Computer Science, 2192:134–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2192/21920134.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2192/21920134.
pdf.

Nelson:2001:SFV

[167] Torsten Nelson, Donald Cowan, and Paulo Alencar. Supporting formal verification of crosscutting concerns. Lecture Notes in Computer Science, 2192: 153–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2192/21920153.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2192/21920153.
pdf.

Douence:2001:FDC

[168] Rémi Douence, Olivier Motelet, and Mario Südholt. A formal definition of crosscuts. Lecture Notes in Computer Science, 2192:170–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2192/21920170.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2192/21920170.
                                        pdf.

Andrews:2001:PAF

[169] James H. Andrews. Processalgebraic foundations of aspectoriented programming. Lecture Notes in Computer Science, 2192:187–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2192/21920187.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2192/21920187.
                                        pdf.

Pennaneach:2001:UR

[170] François Pennaneac’h, Jean-Marc Jézéquel, Jacques Malenfant, and Gerson Sunyé. UML reflections. Lecture Notes in Computer Science, 2192:210– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2192/21920210.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2192/21920210.
                                        pdf.

Rashid:2001:HAS

[171] Awais Rashid. A hybrid approach to separation of concerns: The story of SADES. Lecture Notes in Computer Science, 2192:231–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2192/21920231.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2192/21920231.
pdf.

Griswold:2001:CCS

[172] William G. Griswold. Coping with crosscutting software changes using information transparency. Lecture Notes in Computer Science, 2192:250– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2192/21920250.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2192/21920250.
pdf.

Attardi:2001:TMO

[173] Giuseppe Attardi and Antonio Cisternino. Template metaprogramming an object interface to relational tables. Lecture Notes in Computer Science, 2192:266–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2192/21920266.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2192/21920266.
pdf.

Blair:2001:PIO

[174] Gordon S. Blair, Geoff Coulson, Michael Clarke, and Nikos Parlavantzas. Performance and integrity in the OpenORB reflective middleware. Lecture Notes in Computer Science, 2192:268–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2192/21920268.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2192/21920268.
                                        pdf.

Gudmundson:2001:DAA

[175] Stephan Gudmundson and Gregor Kiczales. Data abstraction in AspectJ. Lecture Notes in Computer Science, 2192:270–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2192/21920270.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2192/21920270.
                                        pdf.

Lee:2001:TCR

[176] Joon-Sang Lee, Doo-Hwan Bae, and Ikju Han. Towards coping with role-binding anomaly. Lecture Notes in Computer Science, 2192:272–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2192/21920272.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2192/21920272.
                                        pdf.

Park:2001:RRJ

[177] Jung Gyu Park and Arthur H. Lee. Removing reflection from Java programs using partial evaluation. Lecture Notes in Computer Science, 2192: 274–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2192/21920274.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2192/21920274.
pdf.

Rapicault:2001:OIU

[178] Pascal Rapicault and Jean-Paul Rigault. Open implementation of UML meta-model(s). Lecture Notes in Computer Science, 2192:276–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2192/21920276.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2192/21920276.
pdf.

Renaud:2001:JRJ

[179] Karen Renaud and Huw Evans. Java-Cloak: Reflecting on Java typing for class reuse using proxies. Lecture Notes in Computer Science, 2192:278– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2192/21920278.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2192/21920278.
pdf.

Welch:2001:AEM

[180] Ian Welch, Robert J. Stroud, and Alexander Romanovsky. Aspects of exceptions at the meta-level. Lecture Notes in Computer Science, 2192:280– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2192/21920280.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2192/21920280.
                                        pdf.

Anonymous:2001:AIe

[181] Anonymous. Author index. Lecture Notes in Computer Science, 2192:283– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2192/2192auth.pdf.

Pilioura:2001:SCT

[182] Thomi Pilioura and Aphrodite Tsalgatidou. E-services: Current technology and open issues. Lecture Notes in Computer Science, 2193:1– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2193/21930001.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2193/21930001.
                                        pdf.

Bussler:2001:RBP

[183] Christoph Bussler. The role of B2B protocols in inter-enterprise process execution. Lecture Notes in Computer Science, 2193:16–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2193/21930016.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2193/21930016.
pdf.

Kuno:2001:CIB

[184] Harumi Kuno, Mike Lemon, Alan Karp, and Dorothea Beringer. Conversations + interfaces = business logic. Lecture Notes in Computer Science, 2193:30–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2193/21930030.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2193/21930030.
pdf.

Mecella:2001:CSC

[185] Massimo Mecella, Barbara Pernici, and Paolo Craca. Compatibility of e-services in a cooperative multiplatform environment. Lecture Notes in Computer Science, 2193:44–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2193/21930044.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2193/21930044.
pdf.

Christophides:2001:BDS

[186] Vassilis Christophides, Richard Hull, Gregory Karvounarakis, Akhil Kumar, Geliang Tong, and Ming Xiong. Beyond discrete E-services: Composing session-oriented services in telecommunications. Lecture Notes in Computer Science, 2193:58–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2193/21930058.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2193/21930058.
                                        pdf.

Kafeza:2001:VBC

[187] Eleanna Kafeza, Dickson K. W. Chiu, and Irene Kafeza. Viewbased contracts in an E-service crossorganizational workflow environment. Lecture Notes in Computer Science, 2193:74–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2193/21930074.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2193/21930074.
                                        pdf.

Krithivasan:2001:BSF

[188] Raja Krithivasan and Abdelsalam (Sumi) Helal. BizBuilder — an E-services framework targeted for Internet workflow. Lecture Notes in Computer Science, 2193:89–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2193/21930089.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2193/21930089.pdf.

Fauvet:2001:PPT

[189] Marie-Christine Fauvet, Marlon Dumas, Boualem Benatallah, and Hye-Young Paik. Peer-to-peer traced execution of composite services. Lecture Notes in Computer Science, 2193:103– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2193/21930103.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2193/21930103.
pdf.

Schuler:2001:SRT

[190] Christoph Schuler, Heiko Schuldt, and Hans-Jörg Schek. Supporting reliable transactional business processes by publish/subscribe techniques. Lecture Notes in Computer Science, 2193: 118–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2193/21930118.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2193/21930118.
pdf.

Dutta:2001:UAR

[191] Kaushik Dutta, Debra VanderMeer, Anindya Datta, and Krithi Ramamritham. User action recovery in Internet SAGAs (iSAGAs). Lecture Notes in Computer Science, 2193:132– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2193/21930132.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2193/21930132.
pdf.

Seltzsam:2001:SDS

[192] Stefan Seltzsam, Stephan Börzsöpnyi, and Alfons Kemper. Security for distributed E-service composition. Lecture Notes in Computer Science, 2193: 147–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2193/21930147.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2193/21930147.
                                        pdf.

Ortiz:2001:EEE

[193] Guillermo Fernandez Ortiz, Ana Sofia Caires Branco, Paz Ruiz Sancho, and José Luis Castillo. ES-TIA — efficient Electronic Services for Tourists in Action. Lecture Notes in Computer Science, 2193:163–174, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2193/21930163.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2193/21930163.
                                        pdf.

Herzog:2001:IFF

[194] Marcus Herzog and Georg Gottlob. InfoPipes: a flexible framework for M-commerce applications. Lecture Notes in Computer Science, 2193:175– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2193/21930175.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2193/21930175.
                                        pdf.

Milojicic:2001:PSI

[195] Dejan Milojicic, Alan Messer, Philippe Bernadat, Ira Greenberg, Olaf Spinczyk, Danilo Beuche, and Wolfgang Schröder-Preikschat. Ψ-pervasive services infrastructure. Lecture Notes in Computer Science, 2193:187–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2193/21930187.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2193/21930187.pdf.

Lee:2001:AMS

[196] DongChun Lee and JeomGoo Kim. Adaptive migration strategy for mobile agents on Internet. Lecture Notes in Computer Science, 2193:201–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2193/21930201.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2193/21930201.
pdf.

Anonymous:2001:AIf

[197] Anonymous. Author index. Lecture Notes in Computer Science, 2193:213– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2193/21930001.pdf.

Ghosh:2001:CMA

[198] Sukumar Ghosh. Cooperating mobile agents and stabilization. Lecture Notes in Computer Science, 2194:1– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2194/21940001.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2194/21940001.
                                        pdf.

Beauquier:2001:CCE

[199] Joffroy Beauquier, Maria Gradinariu, and Colette Johnen. Cross-over composition — enforcement of fairness under unfair adversary. Lecture Notes in Computer Science, 2194:19– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2194/21940019.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2194/21940019.
                                        pdf.

Beauquier:2001:ESA

[200] Joffroy Beauquier, Thomas Hérault, and Elad Schiller. Easy stabilization with an agent. Lecture Notes in Computer Science, 2194:35–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2194/21940035.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2194/21940035.
                                        pdf.

Cobb:2001:SRD

[201] Jorge A. Cobb and Mohamed G. Gouda. Stabilization of routing in directed networks. Lecture Notes in Computer Science, 2194:51–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2194/21940051.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2194/21940051.
pdf.

Dolev:2001:DSS

[202] Shlomi Dolev and Ted Herman. Dijkstra’s self-stabilizing algorithm in unsupportive environments. Lecture Notes in Computer Science, 2194:67– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2194/21940067.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2194/21940067.
pdf.

Dolev:2001:CAS

[203] Shlomi Dolev and Elad Schiller. Communication adaptive self-stabilizing group membership service (extended abstract). Lecture Notes in Computer Science, 2194:82–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2194/21940082.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2194/21940082.pdf.

Gartner:2001:IPP

[204] Felix C. Gärtner and Stefan Pleisch. (im)possibilities of predicate detection in crash-affected systems. Lecture Notes in Computer Science, 2194:98– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2194/21940098.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2194/21940098.
                                        pdf.

Gouda:2001:TWS

[205] Mohamed G. Gouda. The theory of weak stabilization. Lecture Notes in Computer Science, 2194:114–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2194/21940114.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2194/21940114.
                                        pdf.

Gouda:2001:SVP

[206] Mohamed G. Gouda, Chin-Tser Huang, and Anish Arora. On the security and vulnerability of PING. Lecture Notes in Computer Science, 2194: 124–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2194/21940124.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2194/21940124.
                                        pdf.

Hadid:2001:NET

[207] Rachid Hadid and Vincent Villain. A new efficient tool for the design of self-stabilizing ℓ-exclusion algorithms: The controller. Lecture Notes in Computer Science, 2194:136– 151, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2194/21940136.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2194/21940136.
pdf.

Herman:2001:SSA

[208] Ted Herman and Toshimitsu Masuzawa. Self-stabilizing agent traversal. Lecture Notes in Computer Science, 2194:152–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2194/21940152.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2194/21940152.
pdf.

Herman:2001:CSD

[209] Ted Herman and Imran Pirwani. A composite stabilizing data structure. Lecture Notes in Computer Science, 2194:167–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2194/21940167.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2194/21940167.
pdf.

Kulkarni:2001:SCD

[210] Sandeep S. Kulkarni and Ravikant. Stabilizing causal deterministic merge. Lecture Notes in Computer Science, 2194:183–199, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2194/21940183.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2194/21940183.pdf.

Petit:2001:FSS

[211] Franck Petit. Fast self-stabilizing depth-first token circulation. Lecture Notes in Computer Science, 2194:200– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2194/21940200.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2194/21940200.
                                        pdf.

Tixeuil:2001:SOD

[212] Sébastien Tixeuil. On a space-optimal distributed traversal algorithm. Lecture Notes in Computer Science, 2194: 216–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2194/21940216.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2194/21940216.
                                        pdf.

Anonymous:2001:AIg

[213] Anonymous. Author index. Lecture Notes in Computer Science, 2194:229– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2194/21940229.pdf.

Kamijoh:2001:LWH

[214] Noboru Kamijoh, Tadanobu Inoue, Kohichiroh Kishimoto, and Ken Tamagawa. Linux watch: Hardware platform for wearable computing research. Lecture Notes in Computer Science, 2195: 1–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2195/21950001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2195/21950001.
pdf.

Ueoka:2001:WCA

[215] Ryoko Ueoka, Michitaka Hirose, Kengo Kuma, Michie Sone, Kenji Kohiyama, Tomohiro Kawamura, and Kenichiro Hiroto. Wearable computer application for open air exhibition in EXPO 2005. Lecture Notes in Computer Science, 2195:8– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2195/21950008.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2195/21950008.
pdf.

Aizawa:2001:ASW

[216] Kiyoharu Aizawa, Kenichiro Ishijima, and Makoto Shiina. Automatic summarization of wearable video — indexing subjective interest. Lecture Notes in Computer Science, 2195:16– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2195/21950016.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2195/21950016.
pdf.

Kawamura:2001:NVR

[217] Tatsuyuki Kawamura, Yasuyuki Kono, and Masatsugu Kidode. A novel video retrieval method to support a user’s recollection of past events aiming for wearable information playing. Lecture Notes in Computer Science, 2195:24– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2195/21950024.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2195/21950024.
pdf.

Ogi:2001:EIV

[218] Tetsuro Ogi, Koji Yamamoto, Toshio Yamada, and Michitaka Hirose. Experience of immersive virtual world using cellular phone interface. Lecture Notes in Computer Science, 2195:32– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2195/21950032.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2195/21950032.
pdf.

Kurata:2001:VTH

[219] Takeshi Kurata, Takashi Okuma, Masakatsu Kourogi, Takekazu Kato, and Katsuhiko Sakaue. VizWear: Toward human-centered interaction through wearable vision and visualization. Lecture Notes in Computer Science, 2195:40–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2195/21950040.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2195/21950040.pdf.

Ai:2001:FIR

[220] Haizhou Ai, Luhong Liang, Xipan Xiao, and Guangyou Xu. Face indexing and retrieval in personal digital album. Lecture Notes in Computer Science, 2195:48–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2195/21950048.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2195/21950048.
pdf.

Idrissi:2001:IRS

[221] K. Idrissi, J. Ricard, A. Anwander, and A. Baskurt. An image retrieval system based on local and global color descriptions. Lecture Notes in Computer Science, 2195:55– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2195/21950055.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2195/21950055.
pdf.

Zhang:2001:NSB

[222] Dong Zhang, Wei Qi, and Hong Jiang Zhang. A new shot boundary detection algorithm. Lecture Notes in Computer Science, 2195:63–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2195/21950063.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2195/21950063.
pdf.

Wu:2001:AIS

[223] P. Wu, B. S. Manjunath, and S. Chandrasekaran. An adaptive index structure for high-dimensional similarity search. Lecture Notes in Computer Science, 2195:71–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2195/21950071.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2195/21950071.
pdf.

Zhou:2001:CHC

[224] Wensheng Zhou and Son K. Dao. Combining hierarchical classifiers with video semantic indexing systems. Lecture Notes in Computer Science, 2195: 78–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2195/21950078.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2195/21950078.
pdf.

Guo:2001:MMR

[225] Wang Xing Guo, Liu Ji Lin, Wei Guo Zheng, and Ishfaq. Mismatch MB retrieval for MPEG-2 to MPEG-4 transcoding. Lecture Notes in Computer Science, 2195:86–93, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2195/21950086.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2195/21950086.
pdf.

Long:2001:OC

[226] Dongyang Long and Weijia Jia. On the optimal coding. Lecture Notes in Computer Science, 2195:94–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2195/21950094.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2195/21950094.
pdf.

Yu:2001:DMR

[227] Yang Yu and Xuelong Zhu. Dynamic multi-reference prediction in video coding for improved error resilience over Internet. Lecture Notes in Computer Science, 2195:102–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2195/21950102.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2195/21950102.
pdf.

Chen:2001:MBS

[228] Duan-Yu Chen and Suh-Yin Lee. Motion-based semantic event detection for video content description in MPEG-7. Lecture Notes in Computer Science, 2195:110–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2195/21950110.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2195/21950110.
pdf.

Lu:2001:FRS

[229] Yan Lu, Wen Gao, and Feng Wu. Fast and robust sprite generation for MPEG-4 video coding. Lecture Notes in Computer Science, 2195:118– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2195/21950118.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2195/21950118.
                                        pdf.

Yang:2001:FME

[230] Shin-Wei Yang, Chung-Neng Wang, Chi-Min Liu, and Tihao Chiang. Fast motion estimation using Nqueen pixel decimation. Lecture Notes in Computer Science, 2195:126–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2195/21950126.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2195/21950126.
                                        pdf.

Wang:2001:IMV

[231] Chung-Neng Wang, Chi-Min Liu, and Tihao Chiang. Improved MPEG-4 visual texture coding using perceptual dithering for transparent image coding. Lecture Notes in Computer Science, 2195:134–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2195/21950134.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2195/21950134.
pdf.

Nguyen:2001:SRC

[232] Anthony G. Nguyen and Jenq-Neng Hwang. SPEM rate control. Lecture Notes in Computer Science, 2195:142– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2195/21950142.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2195/21950142.
pdf.

Wang:2001:RTL

[233] Chunli Wang, Wen Gao, and Zhaoguo Xuan. A real-time large vocabulary continuous recognition system for Chinese Sign Language. Lecture Notes in Computer Science, 2195:150– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2195/21950150.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2195/21950150.
pdf.

Wang:2001:OMC

[234] Dongmei Wang and Russell M. Mersereau. Object modeling, coding, and transmission for multimedia communications. Lecture Notes in Computer Science, 2195:158–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2195/21950158.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2195/21950158.
                                        pdf.

Tanahashi:2001:LEA

[235] Hideki Tanahashi, Caihua Wang, Yoshinori Niwa, and Kazuhiko Yamamoto. Live events accessing for multi-users with free viewpoints using stereo omni-directional system. Lecture Notes in Computer Science, 2195: 166–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2195/21950166.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2195/21950166.
                                        pdf.

Wang:2001:UAH

[236] Tian-Shu Wang, Heung-Yeung Shum, Ying-Qing Xu, and Nan-Ning Zheng. Unsupervised analysis of human gestures. Lecture Notes in Computer Science, 2195:174–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2195/21950174.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2195/21950174.pdf.

Lin:2001:UME

[237] Fan Lin, Liu Wenyin, Zheng Chen, Hongjiang Zhang, and Tang Long. User modeling for efficient use of multimedia files. Lecture Notes in Computer Science, 2195:182–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2195/21950182.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2195/21950182.
pdf.

Jung:2001:FBV

[238] Young-Kee Jung and Yo-Sung Ho. A feature-based vehicle tracking system in congested traffic video sequences. Lecture Notes in Computer Science, 2195:190–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2195/21950190.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2195/21950190.
pdf.

Li:2001:MED

[239] Ying Li and C.-C. Jay Kuo. Movie event detection by using audio visual information. Lecture Notes in Computer Science, 2195:198–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2195/21950198.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2195/21950198.
pdf.

Lu:2001:DHA

[240] Peizhong Lu and Lide Wu. Double hierarchical algorithm for video mosaics. Lecture Notes in Computer Science, 2195:206–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2195/21950206.htm;   http:
//link.springer-ny.com/link/service/

                                        series/0558/papers/2195/21950206.
                                        pdf.

Liu:2001:AST

[241] Ming-Gang Liu and Chao-Huan Hou. Automatic segmentation and tracking of moving objects. Lecture Notes in Computer Science, 2195:214–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2195/21950214.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2195/21950214.
                                        pdf.

Chen:2001:TAD

[242] Xiangrong Chen and Hongjiang Zhang. Text area detection from video frames. Lecture Notes in Computer Science, 2195:222–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2195/21950222.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2195/21950222.
                                        pdf.

Wang:2001:IGO

[243] Yong Wang, Jiang Li, Kun Zhou, and Heung-Yeung Shum. Interacting with 3D graphic objects in an image-based environment. Lecture Notes in Computer Science, 2195:229– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2195/21950229.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2195/21950229.
pdf.

Huang:2001:RFR

[244] Yea-Shuan Huang, Yao-Hong Tsai, and Jun-Wei Shieh. Robust face recognition with light compensation. Lecture Notes in Computer Science, 2195:237– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2195/21950237.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2195/21950237.
pdf.

Chang:2001:RHP

[245] Yao-Jen Chang and Yung-Chang Chen. Robust head pose estimation using textured polygonal model with local correlation measure. Lecture Notes in Computer Science, 2195:245–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2195/21950245.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2195/21950245.
pdf.

Xie:2001:FVA

[246] Feng Xie, Guangyou Xu, and Eckart Hundt. A face verification algorithm integrating geometrical and template features. Lecture Notes in Computer Science, 2195:253–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2195/21950253.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2195/21950253.
                                        pdf.

Zhang:2001:PPR

[247] Bo Zhang, Changbo Hu, Qingsheng Cai, Baining Guo, and Harry Shum. E-partner: a photo-realistic conversation agent. Lecture Notes in Computer Science, 2195:261–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2195/21950261.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2195/21950261.
                                        pdf.

Liu:2001:RFF

[248] Zicheng Liu, Zhengyou Zhang, Dennis Adler, Erik Hanson, and Michael Cohen. A robust and fast face modeling system. Lecture Notes in Computer Science, 2195:269–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2195/21950269.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2195/21950269.
                                        pdf.

Lin:2001:AHF

[249] Kwan-Ho Lin, Baofeng Guo, Kin-Man Lam, and Wan-Chi Siu. Automatic human face recognition system using fractal dimension and modified Hausdorff distance. Lecture Notes in Computer Science, 2195:277– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2195/21950277.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2195/21950277.
pdf.

Chen:2001:SRI

[250] Jiann-Jone Chen, Cheng-Yi Liu, Yea-Shuan Huang, and Jun-Wei Hsieh. Similarity retrieval in image databases by boosted common shape features among query images. Lecture Notes in Computer Science, 2195:285–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2195/21950285.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2195/21950285.
pdf.

Yu:2001:CCS

[251] Huizhen Yu and W. Eric L. Grimson. Combining configurational and statistical approaches in image retrieval. Lecture Notes in Computer Science, 2195: 293–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2195/21950293.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2195/21950293.
pdf.

Huang:2001:DIC

[252] Pei-Yuan Huang and Long-Wen Chang. Digital image coding with hybrid wavelet packet transform. Lecture Notes in Computer Science, 2195:301– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2195/21950301.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2195/21950301.
                                        pdf.

Chiu:2001:TRL

[253] Chih-Yi Chiu, Hsin-Chih Lin, and Shi-Nine Yang. Texture retrieval with linguistic descriptions. Lecture Notes in Computer Science, 2195:308– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2195/21950308.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2195/21950308.
                                        pdf.

Lin:2001:IVS

[254] Chi-Syan Lin. Implementation of the virtual school: Best cyber academy. Lecture Notes in Computer Science, 2195:316–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2195/21950316.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2195/21950316.
                                        pdf.

Jehng:2001:API

[255] Jihn-Chang J. Jehng and Jia-Sheng Heh. Applying process improvement approach to the design and development of a multimedia-based training management system. Lecture Notes in Computer Science, 2195:324– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2195/21950324.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2195/21950324.
pdf.

Wei:2001:ITM

[256] Dong Wei and Nirwan Ansari. IP traffic monitoring: An overview and future considerations. Lecture Notes in Computer Science, 2195:335–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2195/21950335.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2195/21950335.
pdf.

Joubert:2001:EDI

[257] G. R. Joubert and O. Kao. Efficient dynamic image retrieval using the Á trous wavelet transformation. Lecture Notes in Computer Science, 2195:343– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2195/21950343.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2195/21950343.
pdf.

Deb:2001:IRM

[258] Sagarmay Deb and Yanchun Zhang. Image retrieval in multimedia with emergence index. Lecture Notes in Computer Science, 2195:351–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2195/21950351.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2195/21950351.
                                        pdf.

Fung:2001:EMD

[259] Chi wai Fung and Qing Li. Efficient multimedia database indexing using structural join index hierarchy. Lecture Notes in Computer Science, 2195: 359–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2195/21950359.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2195/21950359.
                                        pdf.

Huang:2001:MHB

[260] Jiung-Yao Huang and Lawrence Y. Deng. Modeling of the HLA-based simulation system. Lecture Notes in Computer Science, 2195:367–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2195/21950367.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2195/21950367.
                                        pdf.

Ding:2001:EED

[261] Jen-Wen Ding, Yueh-Min Huang, and Cheng-Chung Chu. An endto-end delivery scheme for robust video streaming. Lecture Notes in Computer Science, 2195:375–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2195/21950375.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2195/21950375.
pdf.

Dittmann:2001:CIM

[262] Jana Dittmann, Enrico Hauer, Claus Vielhauer, Jörg Schwenk, and Eva Saar. Customer identification for MPEG video based on digital fingerprinting. Lecture Notes in Computer Science, 2195:383–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2195/21950383.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2195/21950383.pdf.

Shimizu:2001:PAD

[263] Shuichi Shimizu. Performance analysis of data hiding. Lecture Notes in Computer Science, 2195:391–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2195/21950391.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2195/21950391.
pdf.

Chen:2001:DEU

[264] Chaur-Chin Chen. Data encryption using MRF with an RSA key. Lecture Notes in Computer Science, 2195:399– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2195/21950399.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2195/21950399.
                                        pdf.

Song:2001:DWF

[265] Y. J. Song, R. Z. Liu, and T. N. Tan. Digital watermarking for forgery detection in printed materials. Lecture Notes in Computer Science, 2195:403– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2195/21950403.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2195/21950403.
                                        pdf.

Agung:2001:ICI

[266] I. Wiseto Agung and Peter Sweeney. Improvement and comments on image watermarking using complementary modulation. Lecture Notes in Computer Science, 2195:411–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2195/21950411.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2195/21950411.
                                        pdf.

Bergstrom:2001:EMC

[267] Peter Bergström. An eye-movement controlled wavelet based image coder. Lecture Notes in Computer Science, 2195:418–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2195/21950418.htm;   http:

//link.springer-ny.com/link/service/
series/0558/papers/2195/21950418.
pdf.

He:2001:VFR

[268] Yuwen He, Xuejun Zhao, Shiqiang Yang, and Yuzhuo Zhong. Variable frame-rate video coding based on global motion analysis. Lecture Notes in Computer Science, 2195:426– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2195/21950426.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2195/21950426.
pdf.

Lun:2001:EBI

[269] Daniel P. K. Lun, Tommy C. L. Chan, T. C. Hsung, and David D. Feng. Efficient blind image restoration based on 1-D generalized cross validation. Lecture Notes in Computer Science, 2195: 434–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2195/21950434.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2195/21950434.
pdf.

Lu:2001:BRW

[270] Ta-Te Lu, Kuo-Wei Wen, and Pao-Chi Chang. Block reordering wavelet packet SPIHT image coding. Lecture Notes in Computer Science, 2195:442– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2195/21950442.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2195/21950442.
                                        pdf.

Sun:2001:MAD

[271] Xinding Sun, Ajay Divakaran, and B. S. Manjunath. A motion activity descriptor and its extraction in compressed domain. Lecture Notes in Computer Science, 2195:450–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2195/21950450.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2195/21950450.
                                        pdf.

Wang:2001:FEB

[272] Jian Wang and Chang Wen Chen. A fast and effective block-matching error concealment scheme. Lecture Notes in Computer Science, 2195:458– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2195/21950458.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2195/21950458.
                                        pdf.

Liu:2001:BAV

[273] Shan Liu, JongWon Kim, and C.-C. Jay Kuo. Bit allocation for video coding with temporalspatial tradeoff. Lecture Notes in Computer Science, 2195:466–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2195/21950466.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2195/21950466.
pdf.

Teng:2001:ERI

[274] Shyhwei Teng and Guojun Lu. An evaluation of the robustness of image retrieval based on vector quantization. Lecture Notes in Computer Science, 2195:474–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2195/21950474.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2195/21950474.
pdf.

Yang:2001:SMM

[275] Jun Yang, Yueting Zhuang, and Qing Li. Search for multi-modality data in digital libraries. Lecture Notes in Computer Science, 2195:482–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2195/21950482.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2195/21950482.
pdf.

Chew:2001:CDS

[276] C. M. Chew and M. S. Kankanhalli. Compressed domain summarization of digital video. Lecture Notes in Computer Science, 2195:490– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2195/21950490.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2195/21950490.
                                        pdf.

Wang:2001:ASN

[277] Weiqiang Wang and Wen Gao. Automatic segmentation of news items based on video and audio features. Lecture Notes in Computer Science, 2195: 498–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2195/21950498.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2195/21950498.
                                        pdf.

Xu:2001:AFN

[278] Yin Xu and Yujin Zhang. Association feedback: a novel tool for feature elements based image retrieval. Lecture Notes in Computer Science, 2195:506– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2195/21950506.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2195/21950506.
                                        pdf.

Venters:2001:QVE

[279] Colin C. Venters, Richard J. Hartley, Matthew D. Cooper, and William T. Hewitt. Query by visual example: Assessing the usability of content-based image retrieval system user interfaces. Lecture Notes in Computer Science, 2195:514–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2195/21950514.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2195/21950514.
pdf.

Zhu:2001:UKS

[280] Lei Zhu, Chung Tang, and Aidong Zhang. Using keyblock statistics to model image retrieval. Lecture Notes in Computer Science, 2195:522–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2195/21950522.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2195/21950522.
pdf.

Zhu:2001:PTM

[281] Yongwei Zhu, Mohan S. Kankanhalli, and Changsheng Xu. Pitch tracking and melody slope matching for song retrieval. Lecture Notes in Computer Science, 2195:530–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2195/21950530.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2195/21950530.
pdf.

Moncrieff:2001:AES

[282] Simon Moncrieff, Chitra Dorai, and Svetha Venkatesh. Analysis of environmental sounds as indexical signs in film. Lecture Notes in Computer Science, 2195:538–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2195/21950538.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2195/21950538.
                                        pdf.

Lam:2001:LMR

[283] H. S. Lam, Tan Lee, and P. C. Ching. A low missing rate audio search technique for cantonese radio broadcast recording. Lecture Notes in Computer Science, 2195:546–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2195/21950546.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2195/21950546.
                                        pdf.

Yu:2001:EDS

[284] Feng Yu, Eric Chang, Yingqing Xu, and Heung-Yeung Shum. Emotion detection from speech to enrich multimedia content. Lecture Notes in Computer Science, 2195:550–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2195/21950550.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2195/21950550.
                                        pdf.

Park:2001:PCB

[285] Yong Woon Park, Keon Hyo Baek, and Ki Dong Chung. Proxy caching based on patching scheme and prefetching. Lecture Notes in Computer Science, 2195:558–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2195/21950558.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2195/21950558.
pdf.

Yan:2001:SAQ

[286] Hu Yan and Zhang Guangzhao. A stateless active queue management scheme for approximating fair bandwidth allocation and stabilized buffer occupation. Lecture Notes in Computer Science, 2195:566–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2195/21950566.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2195/21950566.
pdf.

Osada:2001:QBC

[287] Shinji Osada and Hiroaki Higaki. QoSbased checkpoint protocol for multimedia network systems. Lecture Notes in Computer Science, 2195:574– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2195/21950574.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2195/21950574.
pdf.

Huang:2001:IAV

[288] Chung-Ming Huang, Pei-Chuan Liu, and Pi-Fung Shih. An interactive audio and video station over Internet. Lecture Notes in Computer Science, 2195: 582–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2195/21950582.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2195/21950582.
pdf.

Jang:2001:QTN

[289] Jyh-Shing Roger Jang, Hong-Ru Lee, and Chia-Hui Yeh. Query by tapping: a new paradigm for content-based music retrieval from acoustic input. Lecture Notes in Computer Science, 2195: 590–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2195/21950590.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2195/21950590.
pdf.

Wu:2001:MMS

[290] Chung-Hsien Wu, Yu-Hsien Chiu, and Kung-Wei Cheng. Multi-modal sign icon retrieval for augmentative communication. Lecture Notes in Computer Science, 2195:598–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2195/21950598.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2195/21950598.
pdf.

Wang:2001:CWS

[291] Hsin min Wang and Berlin Chen. Comparison of word and subword indexing techniques for Mandarin Chinese spoken document retrieval. Lecture Notes in Computer Science, 2195:606– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2195/21950606.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2195/21950606.
pdf.

Huang:2001:IWM

[292] Chien-Kang Huang, Lee-Feng Chien, and Yen-Jen Oyang. Interactive Web multimedia search using query-session-based query expansion. Lecture Notes in Computer Science, 2195:614–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2195/21950614.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2195/21950614.
pdf.

Chao:2001:CSS

[293] Han-Chieh Chao, Y. M. Chu, and T. G. Tsuei. Codec schemes selection for wireless voice over IP (VoIP). Lecture Notes in Computer Science, 2195:622– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2195/21950622.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2195/21950622.
pdf.

Brandao:2001:UEC

[294] T. Brandão, M. P. Queluz, and A. Rodrigues. On the use of error correction codes in spread spectrum based image watermarking. Lecture Notes in Computer Science, 2195:630–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2195/21950630.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2195/21950630.
                                        pdf.

Seo:2001:CDA

[295] Jin S. Seo and Chang D. Yoo. Correlation detection of asymmetric watermark. Lecture Notes in Computer Science, 2195:638–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2195/21950638.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2195/21950638.
                                        pdf.

Monsignori:2001:WMS

[296] M. Monsignori, P. Nesi, and M. B. Spinu. Watermarking music sheets. Lecture Notes in Computer Science, 2195:646–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2195/21950646.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2195/21950646.
                                        pdf.

Miaou:2001:BCW

[297] Shaou-Gang Miaou, Tzung-Shian Lee, and Chih-Ming Chen. BCH coded watermarks for error-prone transmission of MPEG video. Lecture Notes in Computer Science, 2195:654–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2195/21950654.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2195/21950654.
pdf.

Xie:2001:SCI

[298] Weikai Xie, Yuanchun Shi, Guanyou Xu, and Dong Xie. Smart classroom — an intelligent environment for tele-education. Lecture Notes in Computer Science, 2195:662–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2195/21950662.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2195/21950662.
pdf.

Zhang:2001:MVH

[299] Wenli Zhang, Yoshitomo Yaginuma, and Masao Sakauchi. Movie video hypermedia authoring system. Lecture Notes in Computer Science, 2195:669– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2195/21950669.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2195/21950669.
pdf.

Wang:2001:DCV

[300] Frank Wang, Grace Lian, and Jeff Zheng. Design considerations for the video analyzer SDK. Lecture Notes in Computer Science, 2195:677– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2195/21950677.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2195/21950677.
                                        pdf.

Yu:2001:NAC

[301] Fang Yu, Qian Zhang, Wenwu Zhu, and Ya-Qin Zhang. Networkadaptive cache management schemes for mixed media. Lecture Notes in Computer Science, 2195:685–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2195/21950685.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2195/21950685.
                                        pdf.

Wang:2001:RSI

[302] GuiJin Wang and Xinggang Lin. Robust scalable image transmission over wireless fading channel. Lecture Notes in Computer Science, 2195:693–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2195/21950693.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2195/21950693.
                                        pdf.

Shih:2001:DRS

[303] Hsiang-Ren Shih, Chun-Liang Hou, I-Chieh Lin, and Shie-Jue Lee. A dynamic regulation with scheduler feedback information for multimedia networks. Lecture Notes in Computer Science, 2195:700–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2195/21950700.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2195/21950700.
pdf.

Ma:2001:NSS

[304] Huadong Ma and Kang G. Shin. A new scheduling scheme for multicast true VoD service. Lecture Notes in Computer Science, 2195:708–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2195/21950708.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2195/21950708.
pdf.

Wu:2001:IMH

[305] Yunnan Wu, Anthony Vetro, Huifang Sun, and S.-Y. Kung. Intelligent multihop video communications. Lecture Notes in Computer Science, 2195:716– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2195/21950716.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2195/21950716.
pdf.

Liao:2001:CFI

[306] Pin Liao, Wen Gao, Li Shen, Xilin Chen, Shiguang Shan, and Wenbing Zeng. Classification of facial images using Gaussian mixture models. Lecture Notes in Computer Science, 2195: 724–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2195/21950724.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2195/21950724.
                                        pdf.

Song:2001:FDM

[307] B. C. Song, M. J. Kim, and J. B. Ra. A fast descriptor matching algorithm for exhaustive search in large databases. Lecture Notes in Computer Science, 2195:732–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2195/21950732.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2195/21950732.
                                        pdf.

Qian:2001:FIS

[308] Fang Qian, Lei Zhang, Fuzong Lin, and Bo Zhang. FBCC: An image similarity algorithm based on regions. Lecture Notes in Computer Science, 2195: 740–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2195/21950740.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2195/21950740.
                                        pdf.

Park:2001:NDW

[309] Ji Hwan Park, Sook Ee Jeong, and Young Huh. A new digital watermarking for text document images using diagonal profile. Lecture Notes in Computer Science, 2195:748–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2195/21950748.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2195/21950748.
pdf.

Lee:2001:GCR

[310] Hyun-Ju Lee, Yong-Jae Lee, and Chil-Woo Lee. Gesture classification and recognition using principal component analysis and HMM. Lecture Notes in Computer Science, 2195:756– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2195/21950756.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2195/21950756.
pdf.

Choi:2001:RLF

[311] Wai-Pak Choi, Kin-Man Lam, and Wan-Chi Siu. A robust linefeature-based Hausdorff distance for shape matching. Lecture Notes in Computer Science, 2195:764–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2195/21950764.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2195/21950764.
                                        pdf.

Xu:2001:CPW

[312] Changsheng Xu, David Dagan Feng, and Yongwei Zhu. Copyright protection for WAV-table synthesis audio using digital watermarking. Lecture Notes in Computer Science, 2195:772– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2195/21950772.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2195/21950772.
                                        pdf.

Bo:2001:SCD

[313] Xiaochen Bo, Lincheng Shen, and Wensen Chang. Sign correlation detector for blind image watermarking in the DCT domain. Lecture Notes in Computer Science, 2195:780– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2195/21950780.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2195/21950780.
                                        pdf.

Leung:2001:WDI

[314] K. H. Leung and Bing Zeng. Waveletdomain image watermarking based on statistical metrics. Lecture Notes in Computer Science, 2195:788–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2195/21950788.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2195/21950788.
pdf.

Nakai:2001:SFW

[315] Yuichi Nakai. Semi fragile watermarking based on wavelet transform. Lecture Notes in Computer Science, 2195: 796–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2195/21950796.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2195/21950796.
pdf.

Chang:2001:NSS

[316] Chin-Chen Chang, Jyh-Chiang Yeh, and Ju-Yuan Hsiao. A novel scheme for securing image steganography. Lecture Notes in Computer Science, 2195: 804–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2195/21950804.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2195/21950804.
pdf.

Chen:2001:SFW

[317] Minghua Chen and Yun He. A synchronous fragile watermarking scheme for erroneous Q–DCT coefficients detection. Lecture Notes in Computer Science, 2195:812–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2195/21950812.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2195/21950812.
                                        pdf.

Shih:2001:SCD

[318] C. C. Shih, H. R. Tyan, and H. Y. Mark Liao. Shot change detection based on the Reynolds Transport Theorem. Lecture Notes in Computer Science, 2195: 819–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2195/21950819.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2195/21950819.
                                        pdf.

Chuang:2001:ASC

[319] Shui-Lung Chuang, Lee-Feng Chien, and Hsiao-Tieh Pu. Automatic subject categorization of query terms for filtering sensitive queries in multimedia search. Lecture Notes in Computer Science, 2195:825–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2195/21950825.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2195/21950825.
                                        pdf.

Chang:2001:CIR

[320] Chin-Chen Chang, Chi-Shiang Chan, and Ju-Yuan Hsiao. A color image retrieval method based on local histogram. Lecture Notes in Computer Science, 2195:831–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2195/21950831.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2195/21950831.
pdf.

Fan:2001:SSO

[321] Jianping Fan, Xingquan Zhu, and Lide Wu. Seeded semantic object generation toward content-based video indexing. Lecture Notes in Computer Science, 2195:837–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2195/21950837.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2195/21950837.
pdf.

Zhu:2001:ASD

[322] Xingquan Zhu, Lide Wu, Xiangyang Xue, Xiaoye Lu, and Jianping Fan. Automatic scene detection in news program by integrating visual feature and rules. Lecture Notes in Computer Science, 2195:843–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2195/21950843.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2195/21950843.
pdf.

Shui-Hua:2001:SRH

[323] Hans Shui-Hua and Lu Zheng-Ding. Shape retrieval by hierarchical evolution. Lecture Notes in Computer Science, 2195:849–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2195/21950849.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2195/21950849.
                                        pdf.

Zhang:2001:CSR

[324] Dengsheng Zhang and Guojun Lu. A comparison of shape retrieval using Fourier descriptors and short-time Fourier descriptors. Lecture Notes in Computer Science, 2195:855–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2195/21950855.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2195/21950855.
                                        pdf.

Cheung:2001:SBE

[325] Kathy Cheung, Cindy Hung, C. W. Kok, and Mansun Chan. A SMIL browser with an enhanced image, audio and video effect library for multimedia rich presentations. Lecture Notes in Computer Science, 2195:861– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2195/21950861.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2195/21950861.
                                        pdf.

Chen:2001:MPI

[326] Jiayi Chen, Tele Tan, and Philippe Mulhem. A method for photograph indexing using speech annotation. Lecture Notes in Computer Science, 2195:

867–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2195/21950867.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2195/21950867.
pdf.

Wang:2001:FAS

[327] WeiQiang Wang and Wen Gao. A fast anchor shot detection algorithm on compressed video. Lecture Notes in Computer Science, 2195:873–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2195/21950873.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2195/21950873.
pdf.

Kim:2001:CBN

[328] Young tae Kim, Jae-Gon Kim, Hyun Sung Chang, Kyeongok Kang, and Jinwoong Kim. Content-based news video retrieval with closed captions and time alignment. Lecture Notes in Computer Science, 2195:879– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2195/21950879.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2195/21950879.
pdf.

Yahzong:2001:QSM

[329] Feng Yahzong, Zhuang Yueting, and Pan Yunhe. Query similar music by correlation degree. Lecture Notes in Computer Science, 2195:885–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2195/21950885.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2195/21950885.
                                        pdf.

Erol:2001:CCM

[330] Berna Erol and Faouzi Kossentini. Color content matching of MPEG-4 video objects. Lecture Notes in Computer Science, 2195:891–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2195/21950891.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2195/21950891.
                                        pdf.

Hsieh:2001:MPT

[331] Jun-Wei Hsieh, L. W. Huang, and Yea-Shuan Huang. Multiple-person tracking system for content analysis. Lecture Notes in Computer Science, 2195: 897–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2195/21950897.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2195/21950897.
                                        pdf.

Yamada:2001:MSC

[332] Kunio Yamada, Kenji Mochizuki, Kiyoharu Aizawa, and Takahiro Saito. Motion segmentation with census transform. Lecture Notes in Computer Science, 2195:903–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2195/21950903.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2195/21950903.
pdf.

Jeong:2001:EAM

[333] Jong Myeon Jeong and Young Shik Moon. Efficient algorithms for motion based video retireval. Lecture Notes in Computer Science, 2195:909– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2195/21950909.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2195/21950909.
pdf.

Qiu:2001:CBR

[334] G. Qiu. Content-based retrieval on the compressed domain of 2^{nd} generation image coding. Lecture Notes in Computer Science, 2195:915–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2195/21950915.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2195/21950915.
pdf.

Qiu:2001:RLI

[335] G. Qiu and S. Sudirman. Retrieving landscape images using scene structural matrix. Lecture Notes in Computer Science, 2195:921–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2195/21950921.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2195/21950921.
                                        pdf.

Lu:2001:DWC

[336] Chun-Shien Lu and Hong-Yuan Mark Liao. Digital watermarking: a communications with side information perspective. Lecture Notes in Computer Science, 2195:927–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2195/21950927.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2195/21950927.
                                        pdf.

Kong:2001:AVW

[337] Xiangwei Kong, Yu Liu, and Huajian Liu. Adaptive video watermarking scheme. Lecture Notes in Computer Science, 2195:933–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2195/21950933.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2195/21950933.
                                        pdf.

Qiang:2001:ERA

[338] Li Qiang, Cui Huijuan, and Tang Kun. An error resilience algorithm for H.263++ based video coding. Lecture Notes in Computer Science, 2195:939– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2195/21950939.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2195/21950939.
pdf.

Qiu:2001:PCA

[339] Bin Qiu and Xiaoxiang Guan. A predictive connection admission control scheme for QoS improvement. Lecture Notes in Computer Science, 2195:945– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2195/21950945.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2195/21950945.
pdf.

Park:2001:MCM

[340] Keun hyeong Park, Chul Soo Lee, and HyunWook Park. A multiresolutional coding method based on SPIHT. Lecture Notes in Computer Science, 2195: 951–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2195/21950951.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2195/21950951.
pdf.

Zhao:2001:LIC

[341] Debin Zhao, Wen Gao, Shiguang Shan, and Y. K. Chan. LLEC: An image coder with low-complexity and low-memory requirement. Lecture Notes in Computer Science, 2195:957– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2195/21950957.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2195/21950957.
                                        pdf.

Tsai:2001:GSI

[342] Chwei-Shyong Tsai and Chin-Chen Chang. A generalized secret image sharing and recovery scheme. Lecture Notes in Computer Science, 2195:963– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2195/21950963.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2195/21950963.
                                        pdf.

Shi:2001:NSF

[343] Jiaoying Shi and Kaixiang Yi. New semi-fragile authentication watermarking. Lecture Notes in Computer Science, 2195:969–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2195/21950969.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2195/21950969.pdf.

Hu:2001:PAV

[344] Yong Hu and Chang sheng Xie. Performance analysis of video storage based on clustered NAS architecture. Lecture Notes in Computer Science, 2195: 975–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2195/21950975.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2195/21950975.
pdf.

Liu:2001:MVD

[345] Weimin Liu and Huadong Ma. Modeling video-on-demand system in temporal logic. Lecture Notes in Computer Science, 2195:983–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2195/21950983.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2195/21950983.
pdf.

Wang:2001:APT

[346] Zhiyong Wang, Zheru Chi, Dagan Feng, and S. Y. Cho. Adaptive processing of tree-structure image representation. Lecture Notes in Computer Science, 2195:989–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2195/21950989.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2195/21950989.
pdf.

Jang:2001:BWA

[347] Yongwon Jang, Intaek Kim, Hwan Il Kang, Kab Il Kim, and Seung-Soo Han. Blind watermarking algorithm using complex block selection method. Lecture Notes in Computer Science, 2195: 996–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2195/21950996.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2195/21950996.
                                        pdf.

Montessoro:2001:MLB

[348] P. L. Montessoro, D. Pierattoni, and D. Cortolezzis. MTEACH: a languagebased framework for didactic multimedia production. Lecture Notes in Computer Science, 2195:1002–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2195/21951002.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2195/21951002.
                                        pdf.

Chan:2001:PBD

[349] Siu-Ping Chan and Chi-Wah Kok. Protocol and buffer design for multimediaon-demand system. Lecture Notes in Computer Science, 2195:1010–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2195/21951010.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2195/21951010.
                                        pdf.

Chen:2001:QWB

[350] Zhibo Chen and Yun He. Quality weighted bit allocation for smoother streaming of stored FGS video. Lecture Notes in Computer Science, 2195:1016– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2195/21951016.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2195/21951016.
pdf.

Li:2001:CBM

[351] Guohui Li, Jun Zhang, and Defeng Wu. A content-based multimedia database engine: MIR. Lecture Notes in Computer Science, 2195:1024–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2195/21951024.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2195/21951024.
pdf.

Tzelepi:2001:FAC

[352] Sofia Tzelepi and George Pangalos. A flexible access control model for multimedia medical image security. Lecture Notes in Computer Science, 2195:1030– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2195/21951030.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2195/21951030.
pdf.

Cai:2001:JSC

[353] Zhaohui Cai, T. H. Cheng, and Chao Lu. Joint source/channel coding using turbo trellis codes. Lecture Notes in Computer Science, 2195:1036–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2195/21951036.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2195/21951036.
                                        pdf.

Shyu:2001:AOM

[354] Mei-Ling Shyu, Shu-Ching Chen, and Hongli Luo. An adaptive optimal multimedia network transmission control scheme. Lecture Notes in Computer Science, 2195:1042–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2195/21951042.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2195/21951042.
                                        pdf.

Avanaki:2001:ESR

[355] Alireza Nasiri Avanaki, Babak Hamidzadeh, and Faouzi Kossentini. Effects of sampling rate and pose error on volume reconstruction by space carving. Lecture Notes in Computer Science, 2195:1048– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2195/21951048.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2195/21951048.
                                        pdf.

Chen:2001:TBM

[356] Weijun Chen, Fuzong Lin, Jianmin Li, and Bo Zhang. A tree-based model of prosodic phrasing for Chinese text-to-speech systems. Lecture Notes in Computer Science, 2195:1054– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2195/21951054.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2195/21951054.
pdf.

Zhenyun:2001:DFF

[357] Peng Zhenyun, Hong Wei, Liang Luhong, Xu Guangyou, and Zhang Hongjian. Detecting facial features on image sequences using cross– verification mechanism. Lecture Notes in Computer Science, 2195:1060–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2195/21951060.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2195/21951060.
pdf.

Chen:2001:ROT

[358] Hanfeng Chen, Yiqiang Zhan, and Feihu Qi. Rapid object tracking on compressed video. Lecture Notes in Computer Science, 2195:1066–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2195/21951066.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2195/21951066.
pdf.

Park:2001:DIM

[359] Si-Yong Park, Seung-Ho Park, and Ki-Dong Chung. Design and implementation of the multimedia cluster system based on VIA. Lecture Notes in Computer Science, 2195:1072– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2195/21951072.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2195/21951072.
                                        pdf.

Chang:2001:FIS

[360] Chin-Chen Chang, Shu-Kai Yang, JunWei Hsieh, Ding-Zhou Duan, and Ming-Fen Lin. A fuzzy inference system for simplifying surfaces with color. Lecture Notes in Computer Science, 2195:1078–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2195/21951078.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2195/21951078.
                                        pdf.

Lam:2001:DCF

[361] Suk-Han Lam and Chi-Wah Kok. Demosaic: Color filter array interpolation for digital cameras. Lecture Notes in Computer Science, 2195:1084– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2195/21951084.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2195/21951084.
                                        pdf.

Cai:2001:OBC

[362] Hua Cai and Bing Zeng. Object-based classification of mixed-mode images. Lecture Notes in Computer Science, 2195:1090–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2195/21951090.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2195/21951090.
pdf.

Shan:2001:FUS

[363] Shiguang Shan, Wen Gao, Xilin Chen, Bo Cao, and Wenbin Zeng. A face-unlock screen saver by using face verification based on identityspecific subspaces. Lecture Notes in Computer Science, 2195:1096–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2195/21951096.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2195/21951096.
pdf.

Yoo:2001:ACO

[364] Won-Young Yoo and Joonwhoan Lee. Analysis of camera operations in MPEG compressed domain based on generalized Hough transform. Lecture Notes in Computer Science, 2195:1102– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2195/21951102.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2195/21951102.
pdf.

Chen:2001:SDM

[365] Yiqiang Chen, Wen Gao, Zhaoqi Wang, and Li Zuo. Speech driven MPEG-4 based face animation via neural network. Lecture Notes in Computer Science, 2195:1108–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2195/21951108.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2195/21951108.
                                        pdf.

Chan:2001:MSV

[366] Kam-Fai Chan and Chi-Wah Kok. Mesh simplification by vertex cluster contraction. Lecture Notes in Computer Science, 2195:1114–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2195/21951114.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2195/21951114.
                                        pdf.

Gao:2001:FBB

[367] Xiujuan Gao, Hongxun Yao, Wen Gao, and Wei Zeng. Fusion of biometrics based on D–S theory. Lecture Notes in Computer Science, 2195:1120– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2195/21951120.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2195/21951120.
                                        pdf.

Wu:2001:EVS

[368] Chenyu Wu, Ce Liu, and Jie Zhou. Eyeglasses verification by support vector machine. Lecture Notes in Computer Science, 2195:1126–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2195/21951126.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2195/21951126.
pdf.

Yiu:2001:GBH

[369] K. K. Yiu, M. W. Mak, and S. Y. Kung. A GMM-based handset selector for channel mismatch compensation with applications to speaker identification. Lecture Notes in Computer Science, 2195:1132–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2195/21951132.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2195/21951132.
pdf.

Wenyin:2001:PEF

[370] Yan Li Liu Wenyin and Heung-Yeung Shum. Performance evaluation of 2D feature tracking based on Bayesian estimation. Lecture Notes in Computer Science, 2195:1138–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2195/21951138.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2195/21951138.
                                        pdf.

Anonymous:2001:AIh

[371] Anonymous. Author index. Lecture Notes in Computer Science, 2195:1145– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2195/2195ind.pdf.

Czarnecki:2001:GPS

[372] Krzysztof Czarnecki. Generative programming and software system families. Lecture Notes in Computer Science, 2196:1–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2196/21960001.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2196/21960001.pdf.

Sheard:2001:ARC

[373] Tim Sheard. Accomplishments and research challenges in metaprogramming. Lecture Notes in Computer Science, 2196:2–??, 2001. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer-ny.com/link/
                                        service/series/0558/bibs/2196/21960002.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2196/21960002.pdf.

Wand:2001:SAD

[374] Mitchell Wand. A semantics for advice and dynamic join points in aspect-oriented programming. Lecture Notes in Computer Science, 2196:45– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2196/21960045.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2196/21960045.
pdf.

Johann:2001:SCF

[375] Patricia Johann. Short cut fusion: Proved and improved. Lecture Notes in Computer Science, 2196:47– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2196/21960047.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2196/21960047.
pdf.

Sasano:2001:GEP

[376] Isao Sasano, Zhenjiang Hu, and Masato Takeichi. Generation of efficient programs for solving maximum multimarking problems. Lecture Notes in Computer Science, 2196:72–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2196/21960072.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2196/21960072.
pdf.

Damian:2001:STC

[377] Daniel Damian and Olivier Danvy. Static transition compression. Lecture Notes in Computer Science, 2196:92– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2196/21960092.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2196/21960092.
                                        pdf.

Danvy:2001:UAG

[378] Olivier Danvy, Bernd Grobauer, and Morten Rhiger. A unifying approach to goal-directed evaluation. Lecture Notes in Computer Science, 2196:108– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2196/21960108.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2196/21960108.
                                        pdf.

Asai:2001:IPE

[379] Kenichi Asai. Integrating partial evaluators into interpreters. Lecture Notes in Computer Science, 2196:126– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2196/21960126.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2196/21960126.
                                        pdf.

Wakeling:2001:DMF

[380] David Wakeling. A design methodology for functional programs. Lecture Notes in Computer Science, 2196:146– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2196/21960146.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2196/21960146.
pdf.

Harrison:2001:DAS

[381] Bill Harrison and Tim Sheard. Dynamically adaptable software with metacomputations in a staged language. Lecture Notes in Computer Science, 2196:163–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2196/21960163.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2196/21960163.
pdf.

Ferrari:2001:MMP

[382] Gianluigi Ferrari, Eugenio Moggi, and Rosario Pugliese. MetaKlaim: Metaprogramming for global computing. Lecture Notes in Computer Science, 2196:183–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2196/21960183.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2196/21960183.
pdf.

Herrmann:2001:TAW

[383] Christoph A. Herrmann and Christian Lengauer. A transformational approach which combines size inference and program optimization. Lecture Notes in Computer Science, 2196:199– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2196/21960199.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2196/21960199.
                                        pdf.

Anonymous:2001:AIi

[384] Anonymous. Author index. Lecture Notes in Computer Science, 2196:219– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2196/2196auth.pdf.

Badique:2001:UCE

[385] Eric Badiqué. Under construction in Europe: Virtual and mixed reality for a rich media experience. Lecture Notes in Computer Science, 2197:3– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2197/21970003.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2197/21970003.
                                        pdf.

Nebel:2001:GTF

[386] Jean-Christophe Nebel. Generation of true 3D films. Lecture Notes in Computer Science, 2197:10–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2197/21970010.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2197/21970010.
                                        pdf.

Murphy:2001:SSE

[387] David Murphy and Ian Pitt. Spatial sound enhancing virtual story telling. Lecture Notes in Computer Science, 2197:20–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2197/21970020.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2197/21970020.
pdf.

Balet:2001:VP

[388] Olivier Balet, Paul Kafno, Fred Jordan, and Tony Polichroniadis. The VISIONS project. Lecture Notes in Computer Science, 2197:30–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2197/21970030.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2197/21970030.
pdf.

Favier:2001:PAP

[389] Pierre-Alexandre Favier, Pierre De Loor, and Jacques Tisseau. Programming agent with purposes: Application to autonomous shooting in virtual environment. Lecture Notes in Computer Science, 2197:40–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2197/21970040.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2197/21970040.
pdf.

Nandi:2001:IIT

[390] Alok Nandi and Xavier Marichal. Interactive immersive transfiction. Lecture Notes in Computer Science, 2197: 44–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2197/21970044.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2197/21970044.
                                        pdf.

Glassner:2001:ISP

[391] Andrew Glassner. Interactive storytelling: People, stories, and games. Lecture Notes in Computer Science, 2197:51–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2197/21970051.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2197/21970051.
                                        pdf.

Zancanaro:2001:ATI

[392] Massimo Zancanaro, Alessandro Cappelletti, Claudio Signorini, and Carlo Strapparava. An authoring tool for intelligent educational games. Lecture Notes in Computer Science, 2197:61– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2197/21970061.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2197/21970061.
                                        pdf.

Sgouros:2001:GIM

[393] Nikitas M. Sgouros and Sophia Kousidou. Generation and implementation of mixed-reality, narrative performances involving robotic actors. Lecture Notes in Computer Science, 2197: 69–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2197/21970069.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2197/21970069.
pdf.

Clarke:2001:FDI

[394] Andy Clarke and Grethe Mitchell. Film and the development of interactive narrative. Lecture Notes in Computer Science, 2197:81–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2197/21970081.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2197/21970081.
pdf.

Fencott:2001:VSN

[395] Clive Fencott. Virtual storytelling as narrative potential: Towards an ecology of narrative. Lecture Notes in Computer Science, 2197:90–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2197/21970090.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2197/21970090.
pdf.

Douglas:2001:ANH

[396] Jay Douglas and Jonathan Gratch. Adaptive narrative: How autonomous agents, hollywood, and multiprocessing operating systems can live happily ever after. Lecture Notes in Computer Science, 2197:100–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2197/21970100.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2197/21970100.
                                        pdf.

Blumberg:2001:LCB

[397] Bruce M. Blumberg. Learning in character: Building autonomous animated characters that learn what they ought to learn. Lecture Notes in Computer Science, 2197:113–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2197/21970113.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2197/21970113.
                                        pdf.

Machado:2001:RCV

[398] Isabel Machado, Ana Paiva, and Paul Brna. Real characters in virtual stories. Lecture Notes in Computer Science, 2197:127–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2197/21970127.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2197/21970127.
pdf.

Menou:2001:RTC

[399] Eric Menou. Real-time character animation using multi-layered scripts and spacetime optimization. Lecture Notes in Computer Science, 2197:135– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2197/21970135.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2197/21970135.
pdf.

Cavazza:2001:CSA

[400] Marc Cavazza, Fred Charles, and Steven J. Mead. Characters in search of an author: AI-based virtual storytelling. Lecture Notes in Computer Science, 2197:145–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2197/21970145.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2197/21970145.
pdf.

Maffre:2001:VAS

[401] Eric Maffre, Jacques Tisseau, and Marc Parenthoën. Virtual agents’ selfperception in story telling. Lecture Notes in Computer Science, 2197:155– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2197/21970155.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2197/21970155.
                                        pdf.

Kafno:2001:RHH

[402] Paul Kafno. Reflections from a hobby horse. Lecture Notes in Computer Science, 2197:161–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2197/21970161.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2197/21970161.
                                        pdf.

Labaye:2001:DMH

[403] Benoît Labaye, Nadia Guérin, and Sébastien Dohogne. DocToon\begin{matrix}c \\ ◯ \end{matrix} — A mediator in the hospital of the XXIst century. Lecture Notes in Computer Science, 2197:171–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2197/21970171.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2197/21970171.
                                        pdf.

Roussou:2001:IBF

[404] Maria Roussou. The interplay between form, story, and history: The use of narrative in cultural and educational virtual reality. Lecture Notes in Computer Science, 2197:181–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2197/21970181.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2197/21970181.
pdf.

Schafer:2001:VSC

[405] Leonie Schäfer, Wolfgang Prinz, Amanda Oldroyd, and Lesley Gavin. Virtual storytelling of cooperative activities in a theatre of work. Lecture Notes in Computer Science, 2197:191– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2197/21970191.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2197/21970191.
pdf.

Querrec:2001:VST

[406] Ronan Querrec and Pierre Chevaillier. Virtual storytelling for training: An application to fire fighting in industrial environment. Lecture Notes in Computer Science, 2197:201– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2197/21970201.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2197/21970201.
pdf.

Jessel:2001:CAV

[407] Jean-Pierre Jessel, Céline Jaspart, and Jean-Jacques Flores. Computer animation and virtual reality for live art performance. Lecture Notes in Computer Science, 2197:205– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2197/21970205.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2197/21970205.
                                        pdf.

Audience:2001:VHE

[408] Jaroslav Adamec, Jiri Cizek, Michal Masa, Pascal Silondi, Pavel Smetana, and Jiri Zara. Virtual house of European culture: e-AGORA (Electronic Arts for Geographically Open Real Audience). Lecture Notes in Computer Science, 2197:208–211, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2197/21970208.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2197/21970208.
                                        pdf.

Anonymous:2001:AIj

[409] Anonymous. Author index. Lecture Notes in Computer Science, 2197:213– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2197/2197auth.pdf.

Yao:2001:WIW

[410] Y. Y. Yao, Ning Zhong, Jiming Liu, and Setsuo Ohsuga. Web intelligence (WI). Lecture Notes in Computer Science, 2198:1–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2198/21980001.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2198/21980001.
pdf.

Hendler:2001:KPS

[411] James Hendler and Edward A. Feigenbaum. Knowledge is power: The semantic Web vision. Lecture Notes in Computer Science, 2198:18–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2198/21980018.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2198/21980018.
pdf.

Cercone:2001:CIW

[412] Nick Cercone. From computational intelligence to Web intelligence: An ensemble from potpourri. Lecture Notes in Computer Science, 2198:30– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2198/21980030.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2198/21980030.
pdf.

Johnson:2001:PAW

[413] W. Lewis Johnson. Pedagogical agents for Web-based learning. Lecture Notes in Computer Science, 2198:43– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2198/21980043.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2198/21980043.
pdf.

Mizoguchi:2001:OEF

[414] Riichiro Mizoguchi. Ontological engineering: Foundation of the next generation knowledge processing. Lecture Notes in Computer Science, 2198:44– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2198/21980044.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2198/21980044.
                                        pdf.

Raghavan:2001:SNW

[415] Prabhakar Raghavan. Social networks on the Web and in the enterprise. Lecture Notes in Computer Science, 2198: 58–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2198/21980058.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2198/21980058.
                                        pdf.

Wang:2001:ORV

[416] Patrick S. P. Wang. 3D object recognition and visualization on the Web. Lecture Notes in Computer Science, 2198: 61–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2198/21980061.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2198/21980061.
                                        pdf.

Aguilar:2001:WPC

[417] Jose Aguilar and Ernst Leiss. A Web proxy cache coherency and replacement approach. Lecture Notes in Computer Science, 2198:75–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2198/21980075.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2198/21980075.
pdf.

Chiu:2001:CRM

[418] Chi-Huang Chiu, Kai-Chih Liang, and Shyan-Ming Yuan. Content request markup language (CRML): a distributed framework for XMLbased content publishing. Lecture Notes in Computer Science, 2198:85– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2198/21980085.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2198/21980085.
pdf.

Jensen:2001:RSA

[419] Richard Jensen and Qiang Shen. A rough set-aided system for sorting WWW bookmarks. Lecture Notes in Computer Science, 2198:95–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2198/21980095.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2198/21980095.
pdf.

Matsuo:2001:ACN

[420] Yutaka Matsuo, Yukio Ohsawa, and Mitsuru Ishizuka. Average-clicks: a new measure of distance on the World Wide Web. Lecture Notes in Computer Science, 2198:106–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2198/21980106.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2198/21980106.
                                        pdf.

Tsui:2001:AOL

[421] Kwok Ching Tsui, Jiming Liu, and Hiu Lo Liu. Autonomy oriented load balancing in proxy cache servers. Lecture Notes in Computer Science, 2198: 115–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2198/21980115.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2198/21980115.
                                        pdf.

Bun:2001:ETT

[422] Khoo Khyou Bun and Mitsuru Ishizuka. Emerging topic tracking system. Lecture Notes in Computer Science, 2198:125–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2198/21980125.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2198/21980125.pdf.

Butz:2001:APC

[423] C. J. Butz. On axiomatizing probabilistic conditional independencies in Bayesian networks. Lecture Notes in Computer Science, 2198:131–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2198/21980131.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2198/21980131.
pdf.

Kim:2001:DEG

[424] DaeEun Kim and Sea Woo Kim. Dynamic expert group models for recommender systems. Lecture Notes in Computer Science, 2198:136–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2198/21980136.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2198/21980136.
pdf.

McArthur:2001:AOC

[425] Robert McArthur and Peter Bruza. The ABC’s of online community. Lecture Notes in Computer Science, 2198: 141–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2198/21980141.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2198/21980141.
pdf.

Wu:2001:SCW

[426] Hongjing Wu and Paul De Bra. Sufficient conditions for well-behaved adaptive hypermedia systems. Lecture Notes in Computer Science, 2198:148– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2198/21980148.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2198/21980148.
                                        pdf.

Doubrovski:2001:TFS

[427] Vadim Doubrovski. Towards formal specification of client-server interactions for a wide range of Internet applications. Lecture Notes in Computer Science, 2198:153–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2198/21980153.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2198/21980153.
                                        pdf.

Sumi:2001:CVE

[428] Yasuyuki Sumi and Kenji Mase. Collecting, visualizing, and exchanging personal interests and experiences in communities. Lecture Notes in Computer Science, 2198:163–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2198/21980163.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2198/21980163.
                                        pdf.

Wieczorkowska:2001:ACD

[429] Alicja A. Wieczorkowska and Zbigniew W. Ras. Audio content description in sound databases. Lecture Notes in Computer Science, 2198:175– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2198/21980175.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2198/21980175.
pdf.

Ardissono:2001:IID

[430] L. Ardissono, A. Felfernig, G. Friedrich, D. Jannach, R. Schäfer, and M. Zanker. Intelligent interfaces for distributed Web-based product and service configuration. Lecture Notes in Computer Science, 2198:184–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2198/21980184.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2198/21980184.pdf.

Lin:2001:UNW

[431] Sunny S. J. Lin, Eric Z.-F. Liu, M. C. Cheng, and S. M. Yuan. Using networked workshop system to enhance creative design. Lecture Notes in Computer Science, 2198:189–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2198/21980189.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2198/21980189.
pdf.

Matsumura:2001:DSN

[432] Naohiro Matsumura, Yukio Ohsawa, and Mitsuru Ishizuka. Discovering seeds of new interest spread from premature pages cited by multiple communities. Lecture Notes in Computer Science, 2198:195–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2198/21980195.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2198/21980195.
                                        pdf.

Takeda:2001:PWK

[433] Koichi Takeda and Hiroshi Nomiyama. Personalized Web knowledge management. Lecture Notes in Computer Science, 2198:200–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2198/21980200.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2198/21980200.pdf.

Lee:2001:ERS

[434] Minsoo Lee, Stanley Y. W. Su, and Herman Lam. Event and rule services for achieving a Web-based knowledge network. Lecture Notes in Computer Science, 2198:205–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2198/21980205.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2198/21980205.
                                        pdf.

Stuckenschmidt:2001:KBV

[435] Heiner Stuckenschmidt and Frank van Harmelen. Knowledge-based validation, aggregation, and visualization of meta-data: Analyzing a Web-based information system. Lecture Notes in Computer Science, 2198:217–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2198/21980217.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2198/21980217.
pdf.

Wijesoma:2001:OHS

[436] W. Sardha Wijesoma, K. W. Yue, K. L. Chien, and T. K. Chow. Online handwritten signature verification for electronic commerce over the Internet. Lecture Notes in Computer Science, 2198: 227–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2198/21980227.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2198/21980227.
pdf.

Wuwongse:2001:DMX

[437] Vilas Wuwongse, Kiyoshi Akama, Chutiporn Anutariya, and Ekawit Nantajeewarawat. A data model for XML databases. Lecture Notes in Computer Science, 2198:237–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2198/21980237.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2198/21980237.
pdf.

Mine:2001:CIM

[438] Tsunenori Mine, Makoto Amamiya, and Teruko Mitamura. Conference information management system: Towards a personal assistant system. Lecture Notes in Computer Science, 2198: 247–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2198/21980247.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2198/21980247.
                                        pdf.

Perdigao:2001:AIG

[439] José Perdigao, Anjula Garg, Thomas Barbas, Stefan Scheer, Giuseppe Mastrangelo, and Giovanna Rubino. Automatic intelligence gathering from the Web: a case study in container traffic. Lecture Notes in Computer Science, 2198:254–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2198/21980254.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2198/21980254.
                                        pdf.

Shim:2001:WCR

[440] Won Bo Shim and Seog Park. The work concept RBAC model for the access control of the distributed Web server environment. Lecture Notes in Computer Science, 2198:262–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2198/21980262.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2198/21980262.
pdf.

Wong:2001:NCG

[441] Kam Fai Wong, Yat Fan Su, Dongqing Yang, and Shiwei Tang. A new conceptual graph generated algorithm for semi-structured databases. Lecture Notes in Computer Science, 2198:267– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2198/21980267.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2198/21980267.
pdf.

Huang:2001:CTS

[442] Chien-Kang Huang, Yen-Jen Oyang, and Lee-Feng Chien. A contextual term suggestion mechanism for interactive Web search. Lecture Notes in Computer Science, 2198:272–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2198/21980272.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2198/21980272.
pdf.

Hwang:2001:DGI

[443] Jong Ha Hwang, Keung Hae Lee, and Soochan Hwang. 3DGML: a 3-dimensional graphic information retrieval system. Lecture Notes in Computer Science, 2198:282–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2198/21980282.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2198/21980282.
                                        pdf.

Loia:2001:EAA

[444] Vincenzo Loia and Paolo Luongo. An evolutionary approach to automatic Web page categorization and updating. Lecture Notes in Computer Science, 2198:292–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2198/21980292.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2198/21980292.
                                        pdf.

Tsukada:2001:AWP

[445] Makoto Tsukada, Takashi Washio, and Hiroshi Motoda. Automatic Webpage classification by using machine learning methods. Lecture Notes in Computer Science, 2198:303–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2198/21980303.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2198/21980303.
                                        pdf.

Wang:2001:TAI

[446] Z. W. Wang and R. B. Maguire. A theory and approach to improving relevance ranking in Web retrieval. Lecture Notes in Computer Science, 2198: 314–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2198/21980314.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2198/21980314.
pdf.

Yanai:2001:FIG

[447] Keiji Yanai, Masaya Shindo, and Kohei Noshita. A fast image-gathering system on the World-Wide Web using a PC cluster. Lecture Notes in Computer Science, 2198:324–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2198/21980324.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2198/21980324.
pdf.

Brian:2001:MME

[448] Hal D. Brian and Max H. Garzon. MELISSA: Mobile Electronic LSA Internet Server Search Agent. Lecture Notes in Computer Science, 2198:335– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2198/21980335.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2198/21980335.
pdf.

Chau:2001:CFM

[449] Rowena Chau and Chung-Hsing Yeh. Construction of a fuzzy multilingual thesaurus and its application to crosslingual text retrieval. Lecture Notes in Computer Science, 2198:340–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2198/21980340.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2198/21980340.
                                        pdf.

Chung:2001:DWC

[450] Yoojin Chung, Hyuk-Chul Kwon, Sang-Hwa Chung, and Kwang Ryel Ryu. Declustering Web content indices for parallel information retrieval. Lecture Notes in Computer Science, 2198: 346–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2198/21980346.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2198/21980346.
                                        pdf.

Desmontils:2001:IWS

[451] E. Desmontils and C. Jacquin. Indexing a Web site to highlight its content. Lecture Notes in Computer Science, 2198:351–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2198/21980351.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2198/21980351.
                                        pdf.

Fasli:2001:UIR

[452] Maria Fasli and Udo Kruschwitz. Using implicit relevance feedback in a Web search assistant. Lecture Notes in Computer Science, 2198:356–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2198/21980356.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2198/21980356.
pdf.

Fukumoto:2001:DEI

[453] Toru Fukumoto and Kanji Akahori. The development and evaluation of an integrated imagery access engine. Lecture Notes in Computer Science, 2198: 361–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2198/21980361.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2198/21980361.
pdf.

Kim:2001:CIE

[454] Kyung-Joong Kim and Sung-Bae Cho. Conceptual information extraction with link-based search. Lecture Notes in Computer Science, 2198:367–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2198/21980367.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2198/21980367.
pdf.

Li:2001:WWW

[455] Fang Li, Huanye Sheng, and Wilhelm Weisweber. World Wide Web — A multilingual language resource. Lecture Notes in Computer Science, 2198: 373–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2198/21980373.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2198/21980373.
                                        pdf.

Sakairi:2001:QHT

[456] Takashi Sakairi and Hiroshi Nomiyama. Query by history tree manipulation. Lecture Notes in Computer Science, 2198:379–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2198/21980379.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2198/21980379.
                                        pdf.

Sim:2001:WBI

[457] Kwang Mong Sim and Pui Tak Wong. Web-based information retrieval using agent and ontology. Lecture Notes in Computer Science, 2198:384– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2198/21980384.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2198/21980384.
                                        pdf.

Wan:2001:CBS

[458] Chunru Wan, Mingchun Liu, and Lipo Wang. Content-based sound retrieval for Web application. Lecture Notes in Computer Science, 2198:389– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2198/21980389.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2198/21980389.
pdf.

Honda:2001:CFU

[459] Katsuhiro Honda, Nobukazu Sugiura, Hidetomo Ichihashi, and Shoichi Araki. Collaborative filtering using principal component analysis and fuzzy clustering. Lecture Notes in Computer Science, 2198:394–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2198/21980394.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2198/21980394.
pdf.

Lee:2001:JIN

[460] Raymond S. T. Lee. i JADE IW-Shopper: a new age of intelligent mobile Web shopping system based on fuzzy-neuro agent technology. Lecture Notes in Computer Science, 2198:403– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2198/21980403.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2198/21980403.
pdf.

Peters:2001:WAG

[461] J. F. Peters, S. Ramanna, A. Skowron, and M. Borkowski. Wireless agent guidance of remote mobile robots: Rough integral approach to sensor signal analysis. Lecture Notes in Computer Science, 2198:413–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2198/21980413.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2198/21980413.
                                        pdf.

Chen:2001:OBI

[462] Yi-Jia Chen and Von-Wun Soo. Ontology-based information gathering agents. Lecture Notes in Computer Science, 2198:423–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2198/21980423.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2198/21980423.pdf.

Lee:2001:ECA

[463] Seung-Ik Lee, Chul Sung, and Sung-Bae Cho. An effective conversational agent with user modeling based on Bayesian network. Lecture Notes in Computer Science, 2198:428–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2198/21980428.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2198/21980428.
                                        pdf.

Li:2001:IFI

[464] Yuefeng Li. Information fusion for intelligent agent-based information gathering. Lecture Notes in Computer Science, 2198:433–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2198/21980433.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2198/21980433.
pdf.

Lim:2001:ARS

[465] Myungeun Lim and Juntae Kim. An adaptive recommendation system with a coordinator agent. Lecture Notes in Computer Science, 2198:438–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2198/21980438.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2198/21980438.
pdf.

Okabe:2001:IWP

[466] Masayuki Okabe and Seiji Yamada. Interactive Web page filtering with relational learning. Lecture Notes in Computer Science, 2198:443–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2198/21980443.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2198/21980443.
pdf.

Vrettos:2001:FRB

[467] S. Vrettos and A. Stafylopatis. A fuzzy rule-based agent for Web retrieval-filtering. Lecture Notes in Computer Science, 2198:448–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2198/21980448.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2198/21980448.
                                        pdf.

Han:2001:IIP

[468] Jianchao Han and Nick Cercone. Implementation issues and paradigms of visual KDD systems. Lecture Notes in Computer Science, 2198:454–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2198/21980454.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2198/21980454.
                                        pdf.

Kayed:2001:REA

[469] Ahmad Kayed and Robert M. Colomb. Re-engineering approach to build domain ontologies. Lecture Notes in Computer Science, 2198:464–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2198/21980464.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2198/21980464.
                                        pdf.

Matsumura:2001:DET

[470] Naohiro Matsumura, Yukio Ohsawa, and Mitsuru Ishizuka. Discovery of emerging topics between communities on WWW. Lecture Notes in Computer Science, 2198:473–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2198/21980473.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2198/21980473.
pdf.

Yang:2001:MWL

[471] Qiang Yang, Henry Haining Zhang, Ian T. Y. Li, and Ye Lu. Mining Web logs to improve Web caching and prefetching. Lecture Notes in Computer Science, 2198:483–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2198/21980483.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2198/21980483.
pdf.

Dubois:2001:MCD

[472] V. Dubois, M. Quafafou, and B. Habegger. Mining crawled data and visualizing discovered knowledge. Lecture Notes in Computer Science, 2198:493– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2198/21980493.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2198/21980493.
pdf.

Estivill-Castro:2001:CVD

[473] Vladimir Estivill-Castro and Jianhua Yang. Categorizing visitors dynamically by fast and robust clustering of access logs. Lecture Notes in Computer Science, 2198:498–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2198/21980498.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2198/21980498.
                                        pdf.

Ghani:2001:OLW

[474] Rayid Ghani, Rosie Jones, and Dunja Mladenic. Online learning for Web query generation: Finding documents matching a minority concept on the Web. Lecture Notes in Computer Science, 2198:508–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2198/21980508.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2198/21980508.
                                        pdf.

Ogata:2001:FOD

[475] Norihiro Ogata. A formal ontology discovery from Web documents. Lecture Notes in Computer Science, 2198:514– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2198/21980514.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2198/21980514.
                                        pdf.

Maheswari:2001:VPRa

[476] V. Uma Maheswari, Arul Siromoney, and K. M. Mehata. The variable precision rough set model for Web usage mining. Lecture Notes in Computer Science, 2198:520–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2198/21980520.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2198/21980520.
pdf.

Sumi:2001:SCC

[477] Kaoru Sumi and Riichiro Mizoguchi. Supporting cooperative consensus formation via ontologies. Lecture Notes in Computer Science, 2198:525–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2198/21980525.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2198/21980525.
pdf.

Han:2001:WBI

[478] Kyungsook Han and Dongkyu Lee. Web-based intelligent call center for an intensive care unit. Lecture Notes in Computer Science, 2198:530–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2198/21980530.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2198/21980530.
pdf.

Liu:2001:EHW

[479] Chunnian Liu, Lei Zheng, Junzhong Ji, Chengzhong Yang, Jingyue Li, and Wensheng Yang. Electronic homework on the WWW. Lecture Notes in Computer Science, 2198:540–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2198/21980540.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2198/21980540.
                                        pdf.

Stolze:2001:SGE

[480] Markus Stolze and Michael Ströbel. The shopping gate — enabling roleand preference-specific E-commerce shopping experiences. Lecture Notes in Computer Science, 2198:549–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2198/21980549.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2198/21980549.
                                        pdf.

Forcheri:2001:BRA

[481] Paola Forcheri, Maria Teresa Molfino, Stefano Moretti, and Alfonso Quarati. Building reusable and adaptable Web-based courses. Lecture Notes in Computer Science, 2198:562–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2198/21980562.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2198/21980562.
                                        pdf.

Hazeyama:2001:GLS

[482] Atsuo Hazeyama, Akiko Nakako, Sachiko Nakajima, and Keiji Osada. Group learning support system for software engineering education — Web-based collaboration support between the teacher side and the student groups. Lecture Notes in Computer Science, 2198:568–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2198/21980568.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2198/21980568.
pdf.

Ji:2001:IES

[483] Junzhong Ji, Lei Zheng, and Chunnian Liu. The intelligent electronic shopping system based on Bayesian customer modeling. Lecture Notes in Computer Science, 2198:574–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2198/21980574.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2198/21980574.
pdf.

Li:2001:LWA

[484] Sheng-Tun Li. Leveraging a Webaware self-organization map tool for clustering and visualization. Lecture Notes in Computer Science, 2198:579– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2198/21980579.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2198/21980579.
pdf.

Liu:2001:ENA

[485] Eric Zhi-Feng Liu, Sunny S. J. Lin, and Shyan-Ming Yuan. Experiencing Net-Peas: Another way of learning. Lecture Notes in Computer Science, 2198: 584–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2198/21980584.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2198/21980584.
pdf.

Mitsuhara:2001:IIT

[486] Hiroyuki Mitsuhara, Youji Ochi, and Yoneo Yano. ITMS: Individualized Teaching Material System — Web-based exploratory learning support system by adaptive knowledge integration. Lecture Notes in Computer Science, 2198:589–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2198/21980589.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2198/21980589.
pdf.

Molina:2001:ISA

[487] Martin Molina. An intelligent sales assistant for configurable products. Lecture Notes in Computer Science, 2198: 596–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2198/21980596.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2198/21980596.
pdf.

Nguyen:2001:RDD

[488] Ngoc Thanh Nguyen. Reconciling of disagreeing data in Webbased distributed systems using consensus methods. Lecture Notes in Computer Science, 2198:601–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2198/21980601.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2198/21980601.
pdf.

Ochi:2001:WBD

[489] Youji Ochi, Yoneo Yano, and Riko Wakita. Web based digital resource library tracing author’s quotation. Lecture Notes in Computer Science, 2198: 606–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2198/21980606.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2198/21980606.
pdf.

Anonymous:2001:AIk

[490] Anonymous. Author index. Lecture Notes in Computer Science, 2198:613– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2198/2198auth.pdf.

Brause:2001:MAD

[491] R. W. Brause. Medical analysis and diagnosis by neural networks. Lecture Notes in Computer Science, 2199:1– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2199/21990001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2199/21990001.
pdf.

Sierra:2001:ASC

[492] B. Sierra, I. Inza, and P. Larrañaga. On applying supervised classification techniques in medicine. Lecture Notes in Computer Science, 2199:14– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2199/21990014.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2199/21990014.
pdf.

Crespo:2001:MCD

[493] J. Crespo, H. Billhardt, J. Rodríguez-Pedrosa, and J. A. Sanandrés. Methods and criteria for detecting significant regions in medical image analysis. Lecture Notes in Computer Science, 2199: 20–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2199/21990020.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2199/21990020.
pdf.

Acid:2001:UBN

[494] S. Acid, L. M. de Campos, S. Rodríguez, J. M. Rodríguez, and J. L. Salcedo. Using Bayesian networks to model emergency medical services. Lecture Notes in Computer Science, 2199:28– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2199/21990028.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2199/21990028.
pdf.

Alonso:2001:ASD

[495] F. Alonso, Á. López-Illescas, L. Martínez, C. Montes, and J. P. Valente. Analysis of strength data based on expert knowledge. Lecture Notes in Computer Science, 2199:35–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2199/21990035.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2199/21990035.
pdf.

Alves:2001:CEM

[496] V. Alves, J. Neves, M. Maia, and L. Nelas. A computational environment for medical diagnosis support systems. Lecture Notes in Computer Science, 2199:42–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2199/21990042.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2199/21990042.
pdf.

Barreiro:2001:ADR

[497] J. M. Barreiro, A. Carrascal, D. Manrique, J. Ríos, and A. Vilarrasa. Automatic detection of regions of interest in digitized mammograms for computer-aided diagnosis. Lecture Notes in Computer Science, 2199:48– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2199/21990048.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2199/21990048.
pdf.

Billhardt:2001:NMU

[498] H. Billhardt, J. Crespo, V. Maojo, F. Martin, and J. L. Maté. A new method for unifying heterogeneous databases. Lecture Notes in Computer Science, 2199:54–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2199/21990054.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2199/21990054.
pdf.

Bittner:2001:FID

[499] R. Bittner, P. Smrcka, M. Pavelka, P. Vysoký, and L. Pousek. Fatigue indicators of drowsy drivers based on analysis of physiological signals. Lecture Notes in Computer Science, 2199: 62–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2199/21990062.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2199/21990062.
pdf.

Blobel:2001:SID

[500] B. Blobel, G. Stassinopoulos, and P. Hoepner. Secure and interoperable document management over the internet — the generic HARP cross-security platform for clinical studies. Lecture Notes in Computer Science, 2199:69– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2199/21990069.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2199/21990069.
pdf.

Bolotin:2001:GUF

[501] A. Bolotin. A generalized U. function and fuzzy modeling. Lecture Notes in Computer Science, 2199:75– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2199/21990075.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2199/21990075.
pdf.

Fazekas:2001:STS

[502] M. Fazekas. Special time series models for analysis of mortality data. Lecture Notes in Computer Science, 2199: 81–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2199/21990081.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2199/21990081.
pdf.

Pozo:2001:KON

[503] J. A. F. del Pozo, C. Bielza, and M. Gómez. Knowledge organisation in a neonatal jaundice decision support system. Lecture Notes in Computer Science, 2199:88–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2199/21990088.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2199/21990088.
pdf.

Friger:2001:QFM

[504] M. Friger, Y. Novack, and U. Ranft. Quasi-fourier modeling for individual and count outcomes. Lecture Notes in Computer Science, 2199:95– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2199/21990095.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2199/21990095.
pdf.

Linares:2001:AFM

[505] A. J. García de Linares and L. de la Peña Fernández. An anatomical and functional model for the study of cortical functions. Lecture Notes in Computer Science, 2199:101–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2199/21990101.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2199/21990101.
pdf.

Grabert:2001:PLM

[506] M. Grabert, R. W. Holl, U. Krause, I. Melzer, and F. Schweiggert. Predicting the level of metabolic control using collaborative filtering. Lecture Notes in Computer Science, 2199:108– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2199/21990108.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2199/21990108.
pdf.

Juvan:2001:WEK

[507] P. Juvan, B. Zupan, J. Demsar, I. Bratko, John A. Halter, A. Kuspa, and G. Shaulsky. Webenabled knowledge-based analysis of genetic data. Lecture Notes in Computer Science, 2199:113–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2199/21990113.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2199/21990113.
pdf.

Kaczmarek:2001:FSA

[508] E. Kaczmarek, A. Wozniak, and W. Salwa-Zurawska. Fuzzy sets applied to image processing and quantification of interstitial fibrosis and glomerular size in computer assisted microscopy. Lecture Notes in Computer Science, 2199:120–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2199/21990120.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2199/21990120.
                                        pdf.

Kordysh:2001:CES

[509] E. Kordysh, A. Bolotin, M. Barchana, and R. Chen. Cancer epidemiology of small communities: Using a novel approach to detecting clusters. Lecture Notes in Computer Science, 2199:126– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2199/21990126.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2199/21990126.
                                        pdf.

Kurzynski:2001:HPR

[510] M. Kurzynski, E. Puchala, and J. Sas. Hybrid pattern recognition algorithms with the statistical model applied to the computer-aided medical diagnosis. Lecture Notes in Computer Science, 2199:133–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2199/21990133.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2199/21990133.
                                        pdf.

Lado:2001:CAD

[511] M. J. Lado, A. J. Méndez, P. G. Tahoces, M. Souto, and J. J. Vidal. Computer-aided diagnosis: Application of wavelet transform to the detection of clustered microcalcifications in digital mammograms. Lecture Notes in Computer Science, 2199:140– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2199/21990140.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2199/21990140.
pdf.

Laita:2001:MCE

[512] L. M. Laita, G. González-Páez, E. Roanes-Lozano, V. Maojo, L. de Ledesma, and L. Laita. A methodology for constructing expert systems for medical diagnosis. Lecture Notes in Computer Science, 2199:146–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2199/21990146.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2199/21990146.
pdf.

Lamma:2001:ESM

[513] E. Lamma, P. Mello, A. Nanetti, G. Poli, F. Riguzzi, and S. Storari. An expert system for microbiological data validation and surveillance. Lecture Notes in Computer Science, 2199: 153–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2199/21990153.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2199/21990153.
pdf.

Laurikkala:2001:HCF

[514] J. Laurikkala and M. Juhola. Hierarchical clustering of female urinary incontinence data having noise and outliers. Lecture Notes in Computer Science, 2199:161–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2199/21990161.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2199/21990161.
pdf.

Linder:2001:APT

[515] R. Linder and S. J. Pöppl. ACMD: a practical tool for automatic neural net based learning. Lecture Notes in Computer Science, 2199:168–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2199/21990168.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2199/21990168.
pdf.

Marti:2001:DMA

[516] J. Martí, P. Planiol, J. Freixenet, J. Español, and E. Golobardes. Development of a mammographic analysis system using computer vision techniques. Lecture Notes in Computer Science, 2199:174–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2199/21990174.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2199/21990174.
pdf.

Mendez:2001:IMC

[517] A. J. Méndez, P. G. Tahoces, C. Varela, M. J. Lado, M. Souto, and J. J. Vidal. Improvement of a mammographic CAD system for mass detection. Lecture Notes in Computer Science, 2199: 181–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2199/21990181.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2199/21990181.
pdf.

Midelfart:2001:CGE

[518] H. Midelfart, A. Lægreid, and J. Komorowski. Classification of gene expression data in an ontology. Lecture Notes in Computer Science, 2199:186– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2199/21990186.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2199/21990186.
pdf.

Navio:2001:FSA

[519] M. Navío, J. J. Aguilera, M. J. del Jesus, R. González, F. Herrera, and C. Iríbar. Feature selection algorithms applied to Parkinson’s disease. Lecture Notes in Computer Science, 2199:195– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2199/21990195.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2199/21990195.
pdf.

Orbe:2001:NMA

[520] J. Orbe, E. Ferreira, and V. Núñez-Antón. A new model for AIDS survival analysis. Lecture Notes in Computer Science, 2199:201–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2199/21990201.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2199/21990201.
                                        pdf.

Paetz:2001:FPT

[521] J. Paetz and R. Brause. A frequent patterns tree approach for rule generation with categorical septic shock patient data. Lecture Notes in Computer Science, 2199:207–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2199/21990207.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2199/21990207.
                                        pdf.

Pereira:2001:AMD

[522] J. Pereira, A. Lamelo, J. M. Vázquez-Naya, M. Fernández, J. Dorado, A. Santos, J. Teijeiro, and A. Pazos. Analysis of medical diagnostic images via the implementation and access to a safe DICOM PACS with a Web interface: Analysis of contrastenhanced CT imaging of oral and oropharyngeal carcinomas. Lecture Notes in Computer Science, 2199:213– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2199/21990213.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2199/21990213.
pdf.

Perner:2001:CHC

[523] P. Perner. Classification of HEp-2 cells using fluorescent image analysis and data mining. Lecture Notes in Computer Science, 2199:219–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2199/21990219.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2199/21990219.
pdf.

Puchala:2001:MPR

[524] E. Puchala and M. Kurzynski. Multitask pattern recognition algorithm for the medical decision support system. Lecture Notes in Computer Science, 2199:225–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2199/21990225.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2199/21990225.
pdf.

Riano:2001:AHE

[525] D. Riaño and S. Prado. The analysis of hospital episodes. Lecture Notes in Computer Science, 2199:231– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2199/21990231.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2199/21990231.
                                        pdf.

Rodas:2001:EEI

[526] J. Rodas, K. Gibert, and J. E. Rojo. Electroshock effects identification using classification based on rules. Lecture Notes in Computer Science, 2199: 238–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2199/21990238.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2199/21990238.
                                        pdf.

Rotger:2001:AVD

[527] D. Rotger, C. Cañero, P. Radeva, J. Mauri, E. Fernandez, A. Tovar, and V. Valle. Advanced visualization of 3D data of intravascular ultrasound images. Lecture Notes in Computer Science, 2199:245–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2199/21990245.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2199/21990245.
                                        pdf.

Sauerbrei:2001:ISO

[528] W. Sauerbrei. Investigations on stability and overoptimism of classification trees by using cross-validation. Lecture Notes in Computer Science, 2199: 251–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2199/21990251.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2199/21990251.
pdf.

Schlaefer:2001:CBA

[529] A. Schlaefer, K. Schröter, and L. Fritsche. A case-based approach for the classification of medical time series. Lecture Notes in Computer Science, 2199:258–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2199/21990258.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2199/21990258.
pdf.

Siermala:2001:BVR

[530] M. Siermala, M. Juhola, and M. Vihinen. Binary vector or real value coding for secondary structure prediction? A case study of polyproline type II prediction. Lecture Notes in Computer Science, 2199:264–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2199/21990264.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2199/21990264.
pdf.

Sprogar:2001:NMD

[531] M. Sprogar, P. Kokol, M. Zorman, V. Podgorelec, L. Lhotska, and J. Klema. Notes on medical decision model creation. Lecture Notes in Computer Science, 2199:270–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2199/21990270.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2199/21990270.
                                        pdf.

Viikki:2001:RKB

[532] K. Viikki and M. Juhola. Refining the knowledge base of an otoneurological expert system. Lecture Notes in Computer Science, 2199:276–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2199/21990276.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2199/21990276.
                                        pdf.

Walter:2001:SCF

[533] T. Walter and J.-C. Klein. Segmentation of color fundus images of the human retina: Detection of the optic disc and the vascular tree using morphological techniques. Lecture Notes in Computer Science, 2199:282– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2199/21990282.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2199/21990282.
                                        pdf.

Wang:2001:LSK

[534] F. Wang, R. Quiniou, G. Carrault, and M.-O. Cordier. Learning structural knowledge from the ECG. Lecture Notes in Computer Science, 2199: 288–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2199/21990288.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2199/21990288.
pdf.

Wessel:2001:RQA

[535] N. Wessel, N. Marwan, U. Meyerfeldt, A. Schirdewan, and J. Kurths. Recurrence quantification analysis to characterise the heart rate variability before the onset of ventricular tachycardia. Lecture Notes in Computer Science, 2199:295–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2199/21990295.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2199/21990295.
pdf.

Wu:2001:LBN

[536] X. Wu, P. Lucas, S. Kerr, and R. Dijkhuizen. Learning Bayesian-network topologies in realistic medical domains. Lecture Notes in Computer Science, 2199:302–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2199/21990302.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2199/21990302.
pdf.

Anonymous:2001:AIl

[537] Anonymous. Author index. Lecture Notes in Computer Science, 2199:309– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2199/2199ind.pdf.

Blundo:2001:BCU

[538] Carlo Blundo, Paolo D’Arco, Vanesa Daza, and Carles Padró. Bounds and constructions for unconditionally secure distributed key distribution schemes for general access structures. Lecture Notes in Computer Science, 2200:1–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2200/22000001.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2200/22000001.
                                        pdf.

Korjik:2001:PAT

[539] Valeri Korjik, Guillermo Morales-Luna, and Vladimir B. Balakirsky. Privacy amplification theorem for noisy main channel. Lecture Notes in Computer Science, 2200:18–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2200/22000018.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2200/22000018.
                                        pdf.

Crescenzo:2001:EKM

[540] Giovanni Di Crescenzo and Olga Kornievskaia. Efficient kerberized multicast in a practical distributed setting. Lecture Notes in Computer Science, 2200:27–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2200/22000027.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2200/22000027.
pdf.

Gurgens:2001:SCA

[541] Sigrid Gürgens and Javier Lopez. Suitability of a classical analysis method for E-commerce protocols. Lecture Notes in Computer Science, 2200:46– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2200/22000046.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2200/22000046.
pdf.

Blundo:2001:H

[542] Carlo Blundo, Paolo D’Arco, Alfredo De Santis, and Clemente Galdi. Hyppocrates. Lecture Notes in Computer Science, 2200:63–80, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2200/22000063.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2200/22000063.
pdf.

Hoshino:2001:LSB

[543] Fumitaka Hoshino, Masayuki Abe, and Tetsutaro Kobayashi. Lenient/strict batch verification in several groups. Lecture Notes in Computer Science, 2200:81–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2200/22000081.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2200/22000081.
                                        pdf.

Asonov:2001:APV

[544] Dmitri Asonov, Markus Schaal, and Johann-Christoph Freytag. Absolute privacy in voting. Lecture Notes in Computer Science, 2200:95– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2200/22000095.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2200/22000095.
                                        pdf.

Hsu:2001:LMP

[545] Tsan sheng Hsu, Churn-Jung Liau, and Da-Wei Wang. A logical model for privacy protection. Lecture Notes in Computer Science, 2200:110–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2200/22000110.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2200/22000110.
                                        pdf.

Valdez:2001:DDS

[546] Enriquillo Valdez and Moti Yung. DIS-SECT: DIStribution for SECurity Tool. Lecture Notes in Computer Science, 2200:125–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2200/22000125.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2200/22000125.
pdf.

Chow:2001:AOC

[547] Stanley Chow, Yuan Gu, Harold Johnson, and Vladimir A. Zakharov. An approach to the obfuscation of control-flow of sequential computer programs. Lecture Notes in Computer Science, 2200:144–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2200/22000144.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2200/22000144.
pdf.

Chapman:2001:PEA

[548] Mark Chapman, George I. Davida, and Marc Rennhard. A practical and effective approach to large-scale automated linguistic steganography. Lecture Notes in Computer Science, 2200: 156–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2200/22000156.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2200/22000156.
pdf.

Sanchez:2001:RNM

[549] David Sánchez, Agustín Orfila, Julio César Hernández, and José María Sierra. Robust new method in frequency domain watermarking. Lecture Notes in Computer Science, 2200:166–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2200/22000166.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2200/22000166.
                                        pdf.

Berbecaru:2001:CPK

[550] Diana Berbecaru, Antonio Lioy, and Marius Marian. On the complexity of public-key certificate validation. Lecture Notes in Computer Science, 2200: 183–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2200/22000183.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2200/22000183.
                                        pdf.

Martinez-Nadal:2001:LCA

[551] Apol·lònia Martínez-Nadal and Josep L. Ferrer-Gomila. Liability of certification authorities: a juridical point of view. Lecture Notes in Computer Science, 2200:204–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2200/22000204.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2200/22000204.
                                        pdf.

Chodowiec:2001:ETG

[552] Pawel Chodowiec, Kris Gaj, Peter Bellows, and Brian Schott. Experimental testing of the Gigabit IPSeccompliant implementations of Rijndael and Triple DES using SLAAC-1V FPGA accelerator board. Lecture Notes in Computer Science, 2200:220– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2200/22000220.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2200/22000220.
pdf.

Aoki:2001:ECA

[553] Kazumaro Aoki, Fumitaka Hoshino, Tetsutaro Kobayashi, and Hiroaki Oguro. Elliptic curve arithmetic using SIMD. Lecture Notes in Computer Science, 2200:235–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2200/22000235.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2200/22000235.pdf.

Marinis:2001:HIC

[554] Kostas Marinis, Nikos K. Moshopoulos, Fotis Karoubalis, and Kiamal Z. Pekmestzi. On the hardware implementation of the 3GPP confidentiality and integrity algorithms. Lecture Notes in Computer Science, 2200:248– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2200/22000248.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2200/22000248.
                                        pdf.

Xu:2001:EIE

[555] Sheng-Bo Xu and Lejla Batina. Efficient implementation of elliptic curve cryptosystems on an ARM7 with hardware accelerator. Lecture Notes in Computer Science, 2200:266–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2200/22000266.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2200/22000266.
                                        pdf.

Steinwandt:2001:TDB

[556] Rainer Steinwandt, Willi Geiselmann, and Thomas Beth. A theoretical DPAbased cryptanalysis of the NESSIE candidates FLASH and SFLASH. Lecture Notes in Computer Science, 2200:280– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2200/22000280.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2200/22000280.
                                        pdf.

Terada:2001:QRB

[557] Routo Terada and Paulo G. Pinheiro. Quadratic relations for Sboxes: Their minimum representations and bounds. Lecture Notes in Computer Science, 2200:294–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2200/22000294.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2200/22000294.
pdf.

Calvo:2001:APR

[558] Ismael Jiménez Calvo and German Sáez Moreno. Approximate power roots in Z_{m}. Lecture Notes in Computer Science, 2200:310–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2200/22000310.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2200/22000310.
pdf.

Moller:2001:SEC

[559] Bodo Möller. Securing elliptic curve point multiplication against side-channel attacks. Lecture Notes in Computer Science, 2200:324–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2200/22000324.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2200/22000324.
pdf.

Tzelepi:2001:FRB

[560] Sofia Tzelepi and George Pangalos. A flexible role-based access control model for multimedia medical image database systems. Lecture Notes in Computer Science, 2200:335–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2200/22000335.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2200/22000335.
                                        pdf.

Bertino:2001:SPS

[561] Elisa Bertino, Barbara Carminati, and Elena Ferrari. A secure publishing service for digital libraries of XML documents. Lecture Notes in Computer Science, 2200:347–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2200/22000347.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2200/22000347.
                                        pdf.

Markowitch:2001:ONR

[562] Olivier Markowitch and Steve Kremer. An optimistic non-repudiation protocol with transparent trusted third party. Lecture Notes in Computer Science, 2200:363–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2200/22000363.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2200/22000363.
                                        pdf.

Anagnostopoulos:2001:PAD

[563] Aris Anagnostopoulos, Michael T. Goodrich, and Roberto Tamassia. Persistent authenticated dictionaries and their applications. Lecture Notes in Computer Science, 2200:379–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2200/22000379.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2200/22000379.
pdf.

Ferrer-Gomila:2001:EOP

[564] Josep L. Ferrer-Gomila, Magadalena Payeras-Capellà, and Llorenç Huguet-Rotger. Efficient optimistic N-party contract signing protocol. Lecture Notes in Computer Science, 2200:394– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2200/22000394.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2200/22000394.
pdf.

Chida:2001:ESB

[565] Koji Chida, Kunio Kobayashi, and Hikaru Morita. Efficient sealed-bid auctions for massive numbers of bidders with lump comparison. Lecture Notes in Computer Science, 2200:408– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2200/22000408.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2200/22000408.
pdf.

Sebe:2001:OIW

[566] Francesc Sebé and Josep Domingo-Ferrer. Oblivious image watermarking robust against scaling and geometric distortions. Lecture Notes in Computer Science, 2200:420–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2200/22000420.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2200/22000420.
                                        pdf.

Jensen:2001:FTL

[567] Christian D. Jensen. Fingerprinting text in logical markup languages. Lecture Notes in Computer Science, 2200: 433–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2200/22000433.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2200/22000433.
                                        pdf.

Ruiz:2001:SPS

[568] Antonio Ruiz, Gregorio Martínez, Oscar Cánovas, and Antonio F. Gómez. SPEED protocol: Smartcardbased payment with encrypted electronic delivery. Lecture Notes in Computer Science, 2200:446–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2200/22000446.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2200/22000446.
                                        pdf.

Jeong:2001:ETC

[569] Ik Rae Jeong, Dong Hoon Lee, and Jong In Lim. Efficient transferable cash with group signatures. Lecture Notes in Computer Science, 2200:462– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2200/22000462.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2200/22000462.
pdf.

Chen:2001:AMS

[570] Liqun Chen and Wenbo Mao. An auditable metering scheme for Web advertisement applications. Lecture Notes in Computer Science, 2200:475– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2200/22000475.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2200/22000475.
pdf.

Delgado:2001:BBS

[571] Jaime Delgado, Isabel Gallego, and Xavier Perramon. Broker-based secure negotiation of intellectual property rights. Lecture Notes in Computer Science, 2200:486–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2200/22000486.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2200/22000486.
pdf.

Lee:2001:DDS

[572] Jae Seung Lee and Sang Choon Kim. Design of the decision support system for network security management to secure enterprise network. Lecture Notes in Computer Science, 2200:497– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2200/22000497.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2200/22000497.
                                        pdf.

Wang:2001:MFP

[573] Jia Wang and Insup Lee. Measuring false-positive by automated real-time correlated hacking behavior analysis. Lecture Notes in Computer Science, 2200:512–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2200/22000512.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2200/22000512.
                                        pdf.

Masui:2001:DUS

[574] Kenji Masui, Masahiko Tomoishi, and Naoki Yonezaki. Design of UNIX system for the prevention of damage propagation by intrusion and its implementation based on 4.4BSD. Lecture Notes in Computer Science, 2200:536– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2200/22000536.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2200/22000536.
                                        pdf.

Anonymous:2001:AIm

[575] Anonymous. Author index. Lecture Notes in Computer Science, 2200:553– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2200/2200auth.pdf.

Smarr:2001:APC

[576] Larry Smarr. Assembling the planetary computer. Lecture Notes in Computer Science, 2201:1–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2201/22010001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2201/22010001.
pdf.

Espinoza:2001:GSN

[577] Fredrik Espinoza, Per Persson, Anna Sandin, Hanna Nyström, Elenor Cacciatore, and Markus Bylund. GeoNotes: Social and navigational aspects of location-based information systems. Lecture Notes in Computer Science, 2201:2–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2201/22010002.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2201/22010002.pdf.

Castro:2001:PRL

[578] Paul Castro, Patrick Chiu, Ted Kremenek, and Richard Muntz. A probabilistic room location service for wireless networked environments. Lecture Notes in Computer Science, 2201: 18–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2201/22010018.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2201/22010018.
                                        pdf.

Naguib:2001:LIM

[579] Hani Naguib and George Coulouris. Location information managment. Lecture Notes in Computer Science, 2201: 35–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2201/22010035.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2201/22010035.
                                        pdf.

Randell:2001:LCI

[580] Cliff Randell and Henk Muller. Low cost indoor positioning system. Lecture Notes in Computer Science, 2201: 42–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2201/22010042.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2201/22010042.
                                        pdf.

Conner:2001:MEL

[581] W. Steven Conner, Lakshman Krishnamurthy, and Roy Want. Making everyday life easier using dense sensor networks. Lecture Notes in Computer Science, 2201:49–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2201/22010049.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2201/22010049.
pdf.

Ponnekanti:2001:ISF

[582] Shankar R. Ponnekanti, Brian Lee, Armando Fox, Pat Hanrahan, and Terry Winograd. ICrafter: a service framework for ubiquitous computing environments. Lecture Notes in Computer Science, 2201:56–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2201/22010056.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2201/22010056.
pdf.

Teodorescu:2001:UJC

[583] Radu Teodorescu and Raju Pandey. Using JIT compilation and configurable runtime systems for efficient deployment of Java programs on ubiquitous devices. Lecture Notes in Computer Science, 2201:76–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2201/22010076.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2201/22010076.
pdf.

Tandler:2001:SIU

[584] Peter Tandler. Software infrastructure for ubiquitous computing environments: Supporting synchronous collaboration with heterogeneous devices. Lecture Notes in Computer Science, 2201:96–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2201/22010096.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2201/22010096.
pdf.

Holmquist:2001:SFT

[585] Lars Erik Holmquist, Friedemann Mattern, Bernt Schiele, Petteri Alahuhta, Michael Beigl, and Hans-W. Gellersen. Smart-its friends: a technique for users to easily establish connections between smart artefacts. Lecture Notes in Computer Science, 2201:116– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2201/22010116.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2201/22010116.
pdf.

Richter:2001:IMC

[586] Heather Richter, Gregory D. Abowd, Werner Geyer, Ludwin Fuchs, Shahrokh Daijavad, and Steven Poltrock. Integrating meeting capture within a collaborative team environment. Lecture Notes in Computer Science, 2201:123– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2201/22010123.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2201/22010123.
pdf.

Werle:2001:USE

[587] Patrik Werle, Fredrik Kilander, Martin Jonsson, Peter Lönnqvist, and Carl Gustaf Jansson. A ubiquitous service environment with active documents for teamwork support. Lecture Notes in Computer Science, 2201:139– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2201/22010139.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2201/22010139.
pdf.

Sumi:2001:DAS

[588] Yasuyuki Sumi and Kenji Mase. Digital assistant for supporting conference participants: An attempt to combine mobile, ubiquitous and Web computing. Lecture Notes in Computer Science, 2201:156–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2201/22010156.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2201/22010156.
pdf.

Nagel:2001:FID

[589] Kris Nagel, Cory D. Kidd, Thomas O’Connell, Anind Dey, and Gregory D. Abowd. The family intercom: Developing a context-aware audio communication system. Lecture Notes in Computer Science, 2201:176–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2201/22010176.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2201/22010176.
                                        pdf.

Streitz:2001:UCD

[590] Norbert Streitz. Ubiquitous computing and the disappearing computer — research agendas, issues, and strategies. Lecture Notes in Computer Science, 2201:184–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2201/22010184.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2201/22010184.
                                        pdf.

Woodruff:2001:CRE

[591] Allison Woodruff, Margaret H. Szymanski, Paul M. Aoki, and Amy Hurst. The conversational role of electronic guidebooks. Lecture Notes in Computer Science, 2201:187–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2201/22010187.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2201/22010187.
                                        pdf.

Truong:2001:WWW

[592] Khai N. Truong, Gregory D. Abowd, and Jason A. Brotherton. Who, what, when, where, how: Design issues of capture & access applications. Lecture Notes in Computer Science, 2201:209– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2201/22010209.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2201/22010209.
pdf.

Benford:2001:UVH

[593] Steve Benford, John Bowers, Paul Chandler, Luigina Ciolfi, Martin Flintham, Mike Fraser, Chris Greenhalgh, Tony Hall, Sten Olof Hellström, Shahram Izadi, Tom Rodden, Holger Schnädelbach, and Ian Taylor. Unearthing virtual history: Using diverse interfaces to reveal hidden virtual worlds. Lecture Notes in Computer Science, 2201:225–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2201/22010225.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2201/22010225.
pdf.

Lunde:2001:KTE

[594] Thorstein Lunde and Arve Larsen. KISS the tram: Exploring the PDA as support for everyday activities. Lecture Notes in Computer Science, 2201: 232–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2201/22010232.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2201/22010232.
pdf.

Hansson:2001:SPN

[595] Rebecca Hansson, Peter Ljungstrand, and Johan Redström. Subtle and public notification cues for mobile devices. Lecture Notes in Computer Science, 2201:240–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2201/22010240.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2201/22010240.
                                        pdf.

Haritaoglu:2001:ILR

[596] Ismail Haritaoglu. InfoScope: Link from real world to digital information space. Lecture Notes in Computer Science, 2201:247–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2201/22010247.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2201/22010247.
                                        pdf.

Edwards:2001:HUC

[597] W. Keith Edwards and Rebecca E. Grinter. At home with ubiquitous computing: Seven challenges. Lecture Notes in Computer Science, 2201:256– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2201/22010256.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2201/22010256.
                                        pdf.

Langheinrich:2001:PDP

[598] Marc Langheinrich. Privacy by design — principles of privacy-aware ubiquitous systems. Lecture Notes in Computer Science, 2201:273–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2201/22010273.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2201/22010273.
pdf.

Kirkeby:2001:TRW

[599] Anders Kirkeby, Rasmus Zacho, Jock Mackinlay, and Polle Zellweger. Trek-Track: a round wristwatch interface for SMS authoring. Lecture Notes in Computer Science, 2201:292–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2201/22010292.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2201/22010292.
pdf.

Paradiso:2001:CWS

[600] Joseph A. Paradiso and Mark Feldmeier. A compact, wireless, selfpowered pushbutton controller. Lecture Notes in Computer Science, 2201: 299–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2201/22010299.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2201/22010299.
pdf.

Myers:2001:IDU

[601] Brad A. Myers, Choon Hong Peck, Jeffrey Nichols, Dave Kong, and Robert Miller. Interacting at a distance using semantic snarfing. Lecture Notes in Computer Science, 2201:305–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2201/22010305.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2201/22010305.
pdf.

Pinhanez:2001:EDP

[602] Claudio Pinhanez. The everywhere displays projector: a device to create ubiquitous graphical interfaces. Lecture Notes in Computer Science, 2201: 315–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2201/22010315.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2201/22010315.
pdf.

McCarthy:2001:UOG

[603] Joseph F. McCarthy, Tony J. Costa, and Edy S. Liongosari. UniCast, Out-Cast & GroupCast: Three steps toward ubiquitous, peripheral displays. Lecture Notes in Computer Science, 2201: 332–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2201/22010332.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2201/22010332.
pdf.

Johanson:2001:MMW

[604] Brad Johanson, Shankar Ponnekanti, Caesar Sengupta, and Armando Fox. Multibrowsing: Moving Web content across multiple displays. Lecture Notes in Computer Science, 2201:346– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2201/22010346.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2201/22010346.
pdf.

Russell:2001:DPC

[605] Daniel M. Russell and Rich Gossweiler. On the design of personal & communal large information scale appliances. Lecture Notes in Computer Science, 2201: 354–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2201/22010354.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2201/22010354.
pdf.

deBruijn:2001:SWU

[606] Oscar de Bruijn and Robert Spence. Serendipity within a ubiquitous computing environment: a case for opportunistic browsing. Lecture Notes in Computer Science, 2201:362–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2201/22010362.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2201/22010362.
pdf.

Anonymous:2001:AIn

[607] Anonymous. Author index. Lecture Notes in Computer Science, 2201:371– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2201/2201ind.pdf.

Ferrari:2001:LSA

[608] GianLuigi Ferrari, Ugo Montanari, and Emilio Tuosto. A LTS semantics of ambients via graph synchronization with mobility. Lecture Notes in Computer Science, 2202:1– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2202/22020001.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2202/22020001.
                                        pdf.

Alessi:2001:FME

[609] Fabio Alessi, Mariangiola Dezani-Ciancaglini, and Furio Honsell. Filter models and easy terms. Lecture Notes in Computer Science, 2202:17– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2202/22020017.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2202/22020017.
                                        pdf.

Ghilezan:2001:CUL

[610] Silvia Ghilezan and Viktor Kuncak. Confluence of untyped lambda calculus via simple types. Lecture Notes in Computer Science, 2202:38– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2202/22020038.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2202/22020038.
pdf.

Coppo:2001:IIP

[611] Mario Coppo and Daniel Hirschkoff. Incremental inference of partial types. Lecture Notes in Computer Science, 2202:50–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2202/22020050.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2202/22020050.
pdf.

Paolini:2001:CVS

[612] Luca Paolini. Call-by-value separability and computability. Lecture Notes in Computer Science, 2202:74– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2202/22020074.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2202/22020074.
pdf.

Hromkovic:2001:JSS

[613] Juraj Hromkovic, Kathleen Steinhöfel, and Peter Widmayer. Job shop scheduling with unit length tasks: Bounds and algorithms. Lecture Notes in Computer Science, 2202:90– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2202/22020090.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2202/22020090.
                                        pdf.

Jansen:2001:JSS

[614] Klaus Jansen, Monaldo Mastrolilli, and Roberto Solis-Oba. Job shop scheduling problems with controllable processing times. Lecture Notes in Computer Science, 2202:107–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2202/22020107.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2202/22020107.
                                        pdf.

Mereghetti:2001:UBS

[615] Carlo Mereghetti and Beatrice Palano. Upper bounds on the size of oneway quantum finite automata. Lecture Notes in Computer Science, 2202:123– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2202/22020123.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2202/22020123.
                                        pdf.

Besozzi:2001:PSG

[616] Daniela Besozzi, Claudio Zandron, Giancarlo Mauri, and Nicoletta Sabadini. P systems with gemmation of mobile membranes. Lecture Notes in Computer Science, 2202:136–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2202/22020136.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2202/22020136.
pdf.

Prencipe:2001:IAV

[617] Giuseppe Prencipe. Instantaneous actions vs. full asynchronicity: Controlling and coordinating a set of autonomous mobile robots. Lecture Notes in Computer Science, 2202:154– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2202/22020154.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2202/22020154.
pdf.

Cherubini:2001:SSP

[618] Alessandra Cherubini, Stefano Crespi Reghizzi, and Pierluigi San Pietro. Some structural properties of associative language descriptions. Lecture Notes in Computer Science, 2202:172– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2202/22020172.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2202/22020172.
pdf.

Giammarresi:2001:BDR

[619] Dora Giammarresi, Rosa Montalbano, and Derick Wood. Block-deterministic regular languages. Lecture Notes in Computer Science, 2202:184–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2202/22020184.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2202/22020184.
                                        pdf.

Anselmo:2001:CFM

[620] Marcella Anselmo. Constructing finite maximal codes from schützenberger conjecture. Lecture Notes in Computer Science, 2202:197–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2202/22020197.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2202/22020197.pdf.

Ancona:2001:ETF

[621] Davide Ancona, Christopher Anderson, Ferruccio Damiani, Sophia Drossopoulou, Paola Giannini, and Elena Zucca. An effective translation of Fickle into Java. Lecture Notes in Computer Science, 2202:215– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2202/22020215.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2202/22020215.
                                        pdf.

Bugliesi:2001:SMM

[622] Michele Bugliesi, Giuseppe Castagna, and Silvia Crafa. Subtyping and matching for mobile objects. Lecture Notes in Computer Science, 2202:235– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2202/22020235.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2202/22020235.
pdf.

Cacciagrano:2001:SAC

[623] Diletta Cacciagrano and Flavio Corradini. On synchronous and asynchronous communication paradigms. Lecture Notes in Computer Science, 2202:256–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2202/22020256.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2202/22020256.
pdf.

Jacobsen:2001:CLB

[624] Lars Jacobsen and Kim S. Larsen. Complexity of layered binary search trees with relaxed balance. Lecture Notes in Computer Science, 2202:269– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2202/22020269.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2202/22020269.
pdf.

Fiala:2001:DCL

[625] Jirí Fiala, Jan Kratochvíl, and Andrzej Proskurowski. Distance constrained labeling of precolored trees. Lecture Notes in Computer Science, 2202:285– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2202/22020285.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2202/22020285.
                                        pdf.

Jacobsen:2001:EDN

[626] Lars Jacobsen and Kim S. Larsen. Exponentially decreasing number of operations in balanced trees. Lecture Notes in Computer Science, 2202:293– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2202/22020293.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2202/22020293.
                                        pdf.

Ausiello:2001:DHP

[627] Giorgio Ausiello, Paolo G. Franciosa, and Daniele Frigioni. Directed hypergraphs: Problems, algorithmic results, and a novel decremental approach. Lecture Notes in Computer Science, 2202: 312–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2202/22020312.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2202/22020312.
                                        pdf.

Kaporis:2001:CCB

[628] Alexis C. Kaporis, Lefteris M. Kirousis, Yannis C. Stamatiou, Malvina Vamvakari, and Michele Zito. Coupon collectors, q-binomial coefficients and the unsatisfiability threshold. Lecture Notes in Computer Science, 2202:328– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2202/22020328.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2202/22020328.
pdf.

Grosse:2001:RPC

[629] André Große, Jörg Rothe, and Gerd Wechsung. Relating partial and complete solutions and the complexity of computing smallest solutions. Lecture Notes in Computer Science, 2202:339– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2202/22020339.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2202/22020339.
pdf.

DArco:2001:DKD

[630] Paolo D’Arco. On the distribution of a key distribution center. Lecture Notes in Computer Science, 2202:357– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2202/22020357.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2202/22020357.
pdf.

Cimato:2001:OAS

[631] Stelvio Cimato and Annalisa De Bonis. Online advertising: Secure E-coupons. Lecture Notes in Computer Science, 2202:370–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2202/22020370.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2202/22020370.
                                        pdf.

Olivetti:2001:CCB

[632] Nicola Olivetti and Camilla B. Schwind. A calculus and complexity bound for minimal conditional logic. Lecture Notes in Computer Science, 2202:384–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2202/22020384.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2202/22020384.
                                        pdf.

Baldoni:2001:RAC

[633] Matteo Baldoni, Laura Giordano, Alberto Martelli, and Viviana Patti. Reasoning about complex actions with incomplete knowledge: a modal approach. Lecture Notes in Computer Science, 2202:405–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2202/22020405.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2202/22020405.pdf.

Viola:2001:UN

[634] Emanuele Viola. E-unifiability via narrowing. Lecture Notes in Computer Science, 2202:426–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2202/22020426.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2202/22020426.
pdf.

Anonymous:2001:AIo

[635] Anonymous. Author index. Lecture Notes in Computer Science, 2202:439– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2202/2202ind.pdf.

Davidsson:2001:CAS

[636] Paul Davidsson. Categories of artificial societies. Lecture Notes in Computer Science, 2203:1–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2203/22030001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2203/22030001.
pdf.

Fredriksson:2001:MPE

[637] Martin Fredriksson and Rune Gustavsson. A methodological perspective on engineering of agent societies. Lecture Notes in Computer Science, 2203:10– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2203/22030010.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2203/22030010.
pdf.

Vercouter:2001:DAD

[638] Laurent Vercouter. A distributed approach to design open multiagent systems. Lecture Notes in Computer Science, 2203:25–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2203/22030025.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2203/22030025.
pdf.

Cabri:2001:EIM

[639] Giacomo Cabri, Letizia Leonardi, Marco Mamei, and Franco Zambonelli. Engineering infrastructures for mobile organizations. Lecture Notes in Computer Science, 2203:39– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2203/22030039.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2203/22030039.
pdf.

DellAcqua:2001:PUA

[640] Pierangelo Dell’Acqua and Luís Moniz Pereira. Preferring and updating in abductive multi-agent systems. Lecture Notes in Computer Science, 2203:57– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2203/22030057.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2203/22030057.
pdf.

Aszalos:2001:RAF

[641] László Aszalós and Andreas Herzig. Reasoning about failure. Lecture Notes in Computer Science, 2203:74– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2203/22030074.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2203/22030074.
pdf.

Pallotta:2001:AOL

[642] Vincenzo Pallotta and Afzal Ballim. Agent-oriented language engineering for robust NLP. Lecture Notes in Computer Science, 2203:86– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2203/22030086.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2203/22030086.
pdf.

Torroni:2001:ELB

[643] Paolo Torroni and Francesca Toni. Extending a logic based one-to-one negotiation framework to one-to-many negotiation. Lecture Notes in Computer Science, 2203:105–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2203/22030105.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2203/22030105.
pdf.

Carlsson:2001:TCA

[644] Bengt Carlsson. The tragedy of the commons — arms race within peer-to-peer tools. Lecture Notes in Computer Science, 2203:119–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2203/22030119.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2203/22030119.
pdf.

Ambroszkiewicz:2001:AMS

[645] Stanislaw Ambroszkiewicz and Tomasz Nowak. Agentspace as a middleware for service integration. Lecture Notes in Computer Science, 2203:134– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2203/22030134.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2203/22030134.
pdf.

Balbo:2001:TMA

[646] Flavien Balbo and Suzanne Pinson. Toward a multi-agent modelling approach for urban public transportation systems. Lecture Notes in Computer Science, 2203:160–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2203/22030160.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2203/22030160.
pdf.

Perich:2001:IAA

[647] Filip Perich, Lalana Kagal, Harry Chen, Sovrin Tolia, Youyong Zou, Tim Finin, Anupam Joshi, Yun Peng, R. Scott Cost, and Charles Nicholas. ITTALKS: An application of agents in the semantic Web. Lecture Notes in Computer Science, 2203:175–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2203/22030175.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2203/22030175.
pdf.

Anonymous:2001:AIp

[648] Anonymous. Author index. Lecture Notes in Computer Science, 2203:195– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2203/22030195.pdf.

Bandelt:2001:IPM

[649] Hans-J. Bandelt. Invited presentation: Median hulls as Steiner hulls in rectilinear and molecular sequence spaces. Lecture Notes in Computer Science, 2204:1–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2204/22040001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2204/22040001.
pdf.

MeyeraufderHeide:2001:IPD

[650] Friedhelm Meyer auf der Heide. Invited presentation: Data management in networks. Lecture Notes in Computer Science, 2204:8–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2204/22040008.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2204/22040008.
pdf.

Bezrukov:2001:EIP

[651] Sergei L. Bezrukov and Robert Elsässer. Edge-isoperimetric problems for Cartesian powers of regular graphs. Lecture Notes in Computer Science, 2204:9–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2204/22040009.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2204/22040009.
pdf.

Caragiannis:2001:ACB

[652] Ioannis Caragiannis, Afonso Ferreira, Christos Kaklamanis, Stéphane Pérennes, Pino Persiano, and Hervé Rivano. Approximate constrained bipartite edge coloring. Lecture Notes in Computer Science, 2204:21–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2204/22040021.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2204/22040021.
pdf.

Chang:2001:MCT

[653] Maw-Shang Chang, Ton Kloks, and Chuan-Min Lee. Maximum clique transversals. Lecture Notes in Computer Science, 2204:32–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2204/22040032.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2204/22040032.
pdf.

Chang:2001:TDG

[654] Maw-Shang Chang and Haiko Müller. On the tree-degree of graphs. Lecture Notes in Computer Science, 2204:44– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2204/22040044.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2204/22040044.
pdf.

Chen:2001:CMV

[655] Jianer Chen and Iyad A. Kanj. On constrained minimum vertex covers of bipartite graphs: Improved algorithms. Lecture Notes in Computer Science, 2204:55–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2204/22040055.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2204/22040055.
pdf.

Cicerone:2001:KDH

[656] Serafino Cicerone, Gianluca D’Ermiliis, and Gabriele Di Stefano. (k, + )- distance-hereditary graphs. Lecture Notes in Computer Science, 2204:66– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2204/22040066.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2204/22040066.
                                        pdf.

Corneil:2001:RBC

[657] Derek G. Corneil and Udi Rotics. On the relationship between cliquewidth and treewidth. Lecture Notes in Computer Science, 2204:78–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2204/22040078.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2204/22040078.
                                        pdf.

Cornelsen:2001:PLC

[658] Sabine Cornelsen, Yefim Dinitz, and Dorothea Wagner. Planarity of the 2-level cactus model. Lecture Notes in Computer Science, 2204:91–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2204/22040091.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2204/22040091.
                                        pdf.

Dragan:2001:EAP

[659] Feodor F. Dragan. Estimating all pairs shortest paths in restricted graph families: a unified approach. Lecture Notes in Computer Science, 2204:103– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2204/22040103.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2204/22040103.
pdf.

Espelage:2001:HSN

[660] Wolfgang Espelage, Frank Gurski, and Egon Wanke. How to solve NPhard graph problems on clique-width bounded graphs in polynomial time. Lecture Notes in Computer Science, 2204:117–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2204/22040117.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2204/22040117.
pdf.

Feng:2001:FOS

[661] Haodi Feng. (g, f)-factorizations orthogonal to k subgraphs. Lecture Notes in Computer Science, 2204:129– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2204/22040129.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2204/22040129.
pdf.

Fertin:2001:SCG

[662] Guillaume Fertin, André Raspaud, and Bruce Reed. On star coloring of graphs. Lecture Notes in Computer Science, 2204:140–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2204/22040140.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2204/22040140.
                                        pdf.

Fiala:2001:GSC

[663] Jirí Fiala, Klaus Jansen, Van Bang Le, and Eike Seidel. Graph subcolorings: Complexity and algorithms. Lecture Notes in Computer Science, 2204:154– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2204/22040154.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2204/22040154.
                                        pdf.

Fomin:2001:APO

[664] Fedor V. Fomin and Hans L. Bodlaender. Approximation of pathwidth of outerplanar graphs. Lecture Notes in Computer Science, 2204:166– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2204/22040166.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2204/22040166.
                                        pdf.

Fomin:2001:MGG

[665] Fedor V. Fomin and Dimitrios M. Thilikos. On the monotonicity of games generated by symmetric submodular functions. Lecture Notes in Computer Science, 2204:177–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2204/22040177.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2204/22040177.
pdf.

Fuhrmann:2001:MHA

[666] Sven Fuhrmann, Sven Oliver Krumke, and Hans-Christoph Wirth. Multiple hotlink assignment. Lecture Notes in Computer Science, 2204:189– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2204/22040189.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2204/22040189.
pdf.

Gavoille:2001:SDS

[667] Cyril Gavoille, David Peleg, André Raspaud, and Eric Sopena. Small k-dominating sets in planar graphs with applications. Lecture Notes in Computer Science, 2204:201–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2204/22040201.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2204/22040201.
pdf.

Gropl:2001:LBA

[668] Clemens Gröpl, Stefan Hougardy, Till Nierhoff, and Hans Jürgen Prömel. Lower bounds for approximation algorithms for the Steiner tree problem. Lecture Notes in Computer Science, 2204:217–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2204/22040217.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2204/22040217.
                                        pdf.

Johansson:2001:AND

[669] Öjvind Johansson. log n-approximative NLC_{k}-decomposition in O(n^{2k+1}) time. Lecture Notes in Computer Science, 2204:229–240, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2204/22040229.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2204/22040229.pdf.

Kohler:2001:SFG

[670] Ekkehard Köhler, Derek G. Corneil, Stephan Olariu, and Lorna Stewart. On subfamilies of AT-free graphs. Lecture Notes in Computer Science, 2204: 241–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2204/22040241.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2204/22040241.
                                        pdf.

Kral:2001:CCG

[671] Daniel Král’, Jan Kratochvíl, Zsolt Tuza, and Gerhard J. Woeginger. Complexity of coloring graphs without forbidden induced subgraphs. Lecture Notes in Computer Science, 2204:254– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2204/22040254.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2204/22040254.
pdf.

Le:2001:SCL

[672] Van Bang Le and Bert Randerath. On stable cutsets in line graphs. Lecture Notes in Computer Science, 2204:263– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2204/22040263.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2204/22040263.
pdf.

Oh:2001:SMC

[673] Eunseuk Oh and Jianer Chen. On strong Menger-connectivity of star graphs. Lecture Notes in Computer Science, 2204:272–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2204/22040272.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2204/22040272.pdf.

Patrignani:2001:CMC

[674] Maurizio Patrignani and Maurizio Pizzonia. The complexity of the matching-cut problem. Lecture Notes in Computer Science, 2204:284–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2204/22040284.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2204/22040284.
pdf.

Pendavingh:2001:BGD

[675] Rudi Pendavingh, Petra Schuurman, and Gerhard J. Woeginger. De Bruijn graphs and DNA graphs. Lecture Notes in Computer Science, 2204:296– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2204/22040296.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2204/22040296.
pdf.

Puricella:2001:GGA

[676] Antonio Puricella and Iain A. Stewart. A generic greedy algorithm, partiallyordered graphs and NP-completeness. Lecture Notes in Computer Science, 2204:306–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2204/22040306.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2204/22040306.
pdf.

Wagler:2001:CAE

[677] Annegret Wagler. Critical and anticritical edges in perfect graphs. Lecture Notes in Computer Science, 2204: 317–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2204/22040317.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2204/22040317.
pdf.

Anonymous:2001:AIq

[678] Anonymous. Author index. Lecture Notes in Computer Science, 2204:329– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2204/2204auth.pdf.

Goodchild:2001:GLS

[679] Michael F. Goodchild. A geographer looks at spatial information theory. Lecture Notes in Computer Science, 2205:1–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2205/22050001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2205/22050001.
pdf.

Smith:2001:TG

[680] Barry Smith. True grid. Lecture Notes in Computer Science, 2205:14– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2205/22050014.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2205/22050014.
pdf.

Bittner:2001:TGP

[681] Thomas Bittner and Barry Smith. A taxonomy of granular partitions. Lecture Notes in Computer Science, 2205: 28–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2205/22050028.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2205/22050028.
pdf.

Kulik:2001:GTV

[682] Lars Kulik. A geometric theory of vague boundaries based on supervaluation. Lecture Notes in Computer Science, 2205:44–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2205/22050044.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2205/22050044.
pdf.

Ligozat:2001:WTT

[683] Gérard Ligozat. When tables tell it all: Qualitative spatial and temporal reasoning based on linear orderings. Lecture Notes in Computer Science, 2205: 60–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2205/22050060.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2205/22050060.
pdf.

Duckham:2001:CST

[684] Matt Duckham and Michael Worboys. Computational structure in three-valued nearness relations. Lecture Notes in Computer Science, 2205: 76–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2205/22050076.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2205/22050076.
pdf.

Hazarika:2001:QST

[685] Shyamanta M. Hazarika and Anthony G. Cohn. Qualitative spatiotemporal continuity. Lecture Notes in Computer Science, 2205:92–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2205/22050092.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2205/22050092.
pdf.

Bennett:2001:ASS

[686] Brandon Bennett. Application of supervaluation semantics to vaguely defined spatial concepts. Lecture Notes in Computer Science, 2205:108–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2205/22050108.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2205/22050108.
                                        pdf.

Frank:2001:SCS

[687] Andrew U. Frank, Steffen Bittner, and Martin Raubal. Spatial and cognitive simulation with multi-agent systems. Lecture Notes in Computer Science, 2205:124–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2205/22050124.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2205/22050124.
                                        pdf.

Edwards:2001:VTB

[688] Geoffrey Edwards. A virtual test bed in support of cognitively-aware geomatics technologies. Lecture Notes in Computer Science, 2205:140–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2205/22050140.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2205/22050140.
                                        pdf.

Fabrikant:2001:EUS

[689] Sara Irina Fabrikant. Evaluating the usability of the scale metaphor for querying semantic spaces. Lecture Notes in Computer Science, 2205:156– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2205/22050156.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2205/22050156.
pdf.

Hochmair:2001:SMB

[690] Hartwig Hochmair and Andrew U. Frank. A semantic map as basis for the decision process in the www navigation. Lecture Notes in Computer Science, 2205:173–188, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2205/22050173.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2205/22050173.
pdf.

Epstein:2001:PSL

[691] Susan L. Epstein, Bernard Moulin, Walid Chaker, Janice Glasgow, and Jeremi Gancet. Pragmatism and spatial layout design. Lecture Notes in Computer Science, 2205:189–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2205/22050189.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2205/22050189.
pdf.

Miller:2001:SFR

[692] Christy R. Miller and Gary L. Allen. Spatial frames of reference used in identifying direction of movement: An unexpected turn. Lecture Notes in Computer Science, 2205:206–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2205/22050206.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2205/22050206.
                                        pdf.

Sholl:2001:RSR

[693] M. Jeanne Sholl. The role of a selfreference system in spatial navigation. Lecture Notes in Computer Science, 2205:217–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2205/22050217.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2205/22050217.
                                        pdf.

Jefferies:2001:UGR

[694] M. E. Jefferies and W. K. Yeap. The utility of global representations in a cognitive map. Lecture Notes in Computer Science, 2205:233–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2205/22050233.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2205/22050233.
                                        pdf.

Talmy:2001:HSL

[695] Leonard Talmy. How spoken language and signed language structure space differently. Lecture Notes in Computer Science, 2205:247–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2205/22050247.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2205/22050247.
pdf.

Kray:2001:TPP

[696] Christian Kray, Jörg Baus, Hubert Zimmer, Harry Speiser, and Antonio Krüger. Two path prepositions: Along and past. Lecture Notes in Computer Science, 2205:263– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2205/22050263.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2205/22050263.
pdf.

Taylor:2001:AAS

[697] Holly A. Taylor, David H. Uttal, Joan Fisher, and Marshall Mazepa. Ambiguity in acquiring spatial representation from descriptions compared to depictions: The role of spatial orientation. Lecture Notes in Computer Science, 2205:278–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2205/22050278.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2205/22050278.
pdf.

Michon:2001:WWV

[698] Pierre-Emmanuel Michon and Michel Denis. When and why are visual landmarks used in giving directions? Lecture Notes in Computer Science, 2205: 292–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2205/22050292.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2205/22050292.
                                        pdf.

Steinhauer:2001:RAR

[699] Joe Heike Steinhauer, Tom Wiese, Christian Freksa, and Thomas Barkowsky. Recognition of abstract regions in cartographic maps. Lecture Notes in Computer Science, 2205:306–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2205/22050306.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2205/22050306.
                                        pdf.

Jones:2001:GIR

[700] Christopher B. Jones, Harith Alani, and Douglas Tudhope. Geographical information retrieval with ontologies of place. Lecture Notes in Computer Science, 2205:322–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2205/22050322.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2205/22050322.
                                        pdf.

Schlieder:2001:QSR

[701] C. Schlieder, T. Vögele, and U. Visser. Qualitative spatial representation for information retrieval by gazetteers. Lecture Notes in Computer Science, 2205:336–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2205/22050336.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2205/22050336.
pdf.

Behrmann:2001:SRU

[702] Marlene Behrmann and John Philbeck. Spatial representation and updating: Evidence from neuropsychological investigations. Lecture Notes in Computer Science, 2205:352–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2205/22050352.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2205/22050352.
pdf.

Barkowsky:2001:MPG

[703] Thomas Barkowsky. Mental processing of geographic knowledge. Lecture Notes in Computer Science, 2205:371– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2205/22050371.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2205/22050371.
pdf.

Fontaine:2001:SCP

[704] Sylvie Fontaine. Spatial cognition and the processing of verticality in underground environments. Lecture Notes in Computer Science, 2205:387– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2205/22050387.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2205/22050387.
                                        pdf.

Davies:2001:GPC

[705] Clare Davies and Eric Pederson. Grid patterns and cultural expectations in urban wayfinding. Lecture Notes in Computer Science, 2205:400–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2205/22050400.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2205/22050400.
                                        pdf.

Schmidtke:2001:HNR

[706] Hedda R. Schmidtke. The house is north of the river: Relative localization of extended objects. Lecture Notes in Computer Science, 2205:415– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2205/22050415.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2205/22050415.
                                        pdf.

Scivos:2001:DCD

[707] Alexander Scivos and Bernhard Nebel. Double-crossing: Decidability and computational complexity of a qualitative calculus for navigation. Lecture Notes in Computer Science, 2205:431– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2205/22050431.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2205/22050431.
pdf.

Knauff:2001:SRN

[708] Markus Knauff, Corinne Jola, and Gerhard Strube. Spatial reasoning: No need for visual information. Lecture Notes in Computer Science, 2205:447– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2205/22050447.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2205/22050447.
pdf.

Galton:2001:FTO

[709] Antony Galton. A formal theory of objects and fields. Lecture Notes in Computer Science, 2205:458–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2205/22050458.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2205/22050458.
pdf.

Camara:2001:WI

[710] Gilberto Câmara, Max J. Egenhofer, Frederico Fonseca, and Antônio Miguel Vieira Monteiro. What’s in an image? Lecture Notes in Computer Science, 2205:474–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2205/22050474.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2205/22050474.
                                        pdf.

Mark:2001:FOO

[711] David M. Mark, André Skupin, and Barry Smith. Features, objects, and other things: Ontological distinctions in the geographic domain. Lecture Notes in Computer Science, 2205:489– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2205/22050489.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2205/22050489.
                                        pdf.

Anonymous:2001:AIs

[712] Anonymous. Author index. Lecture Notes in Computer Science, 2205:503– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2205/2205auth.pdf.

Takagi:2001:IEC

[713] Hideyuki Takagi. Interactive evolutionary computation as humanized computational intelligence technology. Lecture Notes in Computer Science, 2206: 1–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2206/22060001.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2206/22060001.
                                        pdf.

Patzwahl:2001:MBF

[714] Steffen Patzwahl, Thomas Nacke, Dieter Frense, Dieter Beckmann, Klaus-Dietrich Kramer, Tobias Tautz, and Gerd-Rainer Vollmer. Microcontrollerbased fuzzy system to optimize the anaerobic digestion in biogas reactors. Lecture Notes in Computer Science, 2206:2–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2206/22060002.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2206/22060002.
pdf.

Dziouba:2001:DSC

[715] Tatiana A. Dziouba and Igor N. Rozenberg. The decision of service centres location problem in fuzzy conditions. Lecture Notes in Computer Science, 2206:11–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2206/22060011.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2206/22060011.
pdf.

Tamosiunaite:2001:ABI

[716] Minija Tamošiūnaitė and Erinija Prackevičienė. Application of the biologically inspired network for electroencephalogram analysis. Lecture Notes in Computer Science, 2206:18– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2206/22060018.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2206/22060018.
                                        pdf.

Martin:2001:DBT

[717] Jose L. Martin, Luciano Boquete, Carmen Diaz, Manuel Garcia de Leon, and Manuel Mazo. Discrimination between two mental states (rest and motor image) using wavelet transform and neural networks. Lecture Notes in Computer Science, 2206:28– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2206/22060028.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2206/22060028.
                                        pdf.

Kiseliova:2001:DMD

[718] Tatiana Kiseliova and Hajo Peters. Decision making in diagnosis of oral mucosa lesions. Lecture Notes in Computer Science, 2206:33– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2206/22060033.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2206/22060033.
                                        pdf.

Kacprzyk:2001:FLI

[719] Janusz Kacprzyk. Fuzzy logic and the Internet: Linguistic summarization of distributed sets of data. Lecture Notes in Computer Science, 2206:40– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2206/22060040.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2206/22060040.
pdf.

Leubner:2001:ACE

[720] Christian Leubner. Adaptive colorand edge-based image segmentation using fuzzy techniques. Lecture Notes in Computer Science, 2206:43–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2206/22060043.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2206/22060043.
pdf.

Hata:2001:CAD

[721] Yutaka Hata, Syoji Kobashi, Yasuhiro Tokimoto, Makoto Ishikawa, and Hitoshi Ishikawa. Computer aided diagnosis system of meniscal tears with T1 and T2 weighted MR images based on fuzzy inference. Lecture Notes in Computer Science, 2206:55– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2206/22060055.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2206/22060055.
pdf.

Soria-Frisch:2001:NPF

[722] Aureli Soria-Frisch. A new paradigm for fuzzy aggregation in multisensorial image processing. Lecture Notes in Computer Science, 2206:59– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2206/22060059.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2206/22060059.
                                        pdf.

Bodyanskiy:2001:ALA

[723] Yevgeniy Bodyanskiy, Vitaliy Kolodyazhniy, and Andreas Stephan. An adaptive learning algorithm for a neuro-fuzzy network. Lecture Notes in Computer Science, 2206:68–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2206/22060068.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2206/22060068.
                                        pdf.

Delgosha:2001:FPN

[724] Farshid Delgosha and Mohammad B. Menhaj. Fuzzy probabilistic neural networks: a practical approach to the implementation of Baysian [sic] classifier. Lecture Notes in Computer Science, 2206:76–85, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2206/22060076.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2206/22060076.
                                        pdf.

Bedzak:2001:FBP

[725] Miroslaw Bedzak. Fuzzy-η for back propagation networks. Lecture Notes in Computer Science, 2206:86–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2206/22060086.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2206/22060086.
pdf.

Rummler:2001:EJF

[726] Andreas Rummler and Gerd Scarbata. eaLib — A Java framework for implementation of evolutionary algorithms. Lecture Notes in Computer Science, 2206:92–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2206/22060092.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2206/22060092.
pdf.

Wijesoma:2001:EEF

[727] W. S. Wijesoma and K. R. S. Kodagoda. Experimental evaluation of fuzzy controllers for an outdoor AGV. Lecture Notes in Computer Science, 2206:103–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2206/22060103.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2206/22060103.
pdf.

Hugueney:2001:TSS

[728] Bernard Hugueney and Bernadette Bouchon-Meunier. Time-series segmentation and symbolic representation, from process-monitoring to datamining. Lecture Notes in Computer Science, 2206:118–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2206/22060118.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2206/22060118.pdf.

Kobashi:2001:BSR

[729] Syoji Kobashi, Yutaka Hata, Yuri T. Kitamura, Toshiaki Hayakata, and Toshio Yanagida. Brain state recognition using fuzzy C-means (FCM) clustering with near infrared spectroscopy (NIRS). Lecture Notes in Computer Science, 2206:124–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2206/22060124.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2206/22060124.
pdf.

Yardimci:2001:MBF

[730] A. Yardimci, A. Ferikoglu, and N. Hadimioglu. Microcontroller based fuzzy logic sevofluorane anesthesia control system. Lecture Notes in Computer Science, 2206:137–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2206/22060137.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2206/22060137.
pdf.

Szmidt:2001:IFS

[731] Eulalia Szmidt and Janusz Kacprzyk. Intuitionistic fuzzy sets in some medical applications. Lecture Notes in Computer Science, 2206:148–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2206/22060148.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2206/22060148.
pdf.

Palade:2001:ITN

[732] Vasile Palade, Daniel-Ciprian Neagu, and Ron J. Patton. Interpretation of trained neural networks by rule extraction. Lecture Notes in Computer Science, 2206:152–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2206/22060152.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2206/22060152.
pdf.

Allende:2001:NMI

[733] Héctor Allende, Claudio Moraga, and Rodrigo Salas. Neural model identification using local robustness analysis. Lecture Notes in Computer Science, 2206:162–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2206/22060162.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2206/22060162.
pdf.

Reuter:2001:CAP

[734] Matthias Reuter. Computing with activity patterns. Lecture Notes in Computer Science, 2206:174–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2206/22060174.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2206/22060174.
                                        pdf.

Seifipour:2001:GBA

[735] Navid Seifipour and Mohammad B. Menhaj. A GA-based algorithm with a very fast rate of convergence. Lecture Notes in Computer Science, 2206: 185–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2206/22060185.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2206/22060185.
                                        pdf.

Kern-Isberner:2001:RLC

[736] Gabriele Kern-Isberner. Representing and learning conditional information in possibility theory. Lecture Notes in Computer Science, 2206:194– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2206/22060194.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2206/22060194.
                                        pdf.

Botzheim:2001:ETM

[737] J. Botzheim, B. Hámori, and L. T. Kóczy. Extracting trapezoidal membership functions of a fuzzy rule system by bacterial algorithm. Lecture Notes in Computer Science, 2206:218– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2206/22060218.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2206/22060218.
pdf.

Dubois:2001:PTP

[738] Didier Dubois. Possibility theory, probability theory and multiple-valued logics: a clarification. Lecture Notes in Computer Science, 2206:228–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2206/22060228.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2206/22060228.
pdf.

Gacovski:2001:PPN

[739] Zoran M. Gacovski and Georgi M. Dimirovski. Parameterized Petri net reformulation of the fuzzy production system. Lecture Notes in Computer Science, 2206:229–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2206/22060229.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2206/22060229.
pdf.

Cock:2001:FTI

[740] Martine De Cock and E. E. Kerre. Fuzzy topologies induced by fuzzy relation based modifiers. Lecture Notes in Computer Science, 2206:239–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2206/22060239.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2206/22060239.
                                        pdf.

Kiendl:2001:AIF

[741] H. Kiendl and P. Krause. Advanced inference filter defuzzification. Lecture Notes in Computer Science, 2206:249– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2206/22060249.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2206/22060249.
                                        pdf.

Dyczkowski:2001:CSF

[742] Krzysztof Dyczkowski and Maciej Wygralak. On cardinality and singular fuzzy sets. Lecture Notes in Computer Science, 2206:261–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2206/22060261.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2206/22060261.
                                        pdf.

El-Bakry:2001:FID

[743] Hazem El-Bakry. Fast iris detection for personal verification using modular neural nets. Lecture Notes in Computer Science, 2206:269–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2206/22060269.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2206/22060269.
pdf.

El-Bakry:2001:RIA

[744] Hazem El-Bakry. A rotation invariant algorithm for recognition. Lecture Notes in Computer Science, 2206:284– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2206/22060284.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2206/22060284.
pdf.

Aizenberg:2001:ANN

[745] Igor Aizenberg, Ekaterina Myasnikova, Maria Samsonova, and John Reinitz. Application of the neural networks based on multi-valued neurons to classification of the images of gene expression patterns. Lecture Notes in Computer Science, 2206:291–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2206/22060291.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2206/22060291.
pdf.

Heidemann:2001:MPV

[746] Gunther Heidemann. A multi-purpose visual classification system. Lecture Notes in Computer Science, 2206:305– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2206/22060305.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2206/22060305.
                                        pdf.

Weicker:2001:PDR

[747] Karsten Weicker. Problem difficulty in real-valued dynamic problems. Lecture Notes in Computer Science, 2206:313– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2206/22060313.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2206/22060313.
                                        pdf.

Kokai:2001:GSC

[748] Gabriella Kókai. GeLog — A system combining genetic algorithm with inductive logic programming. Lecture Notes in Computer Science, 2206:326– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2206/22060326.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2206/22060326.
                                        pdf.

Takahashi:2001:DSE

[749] Arita Takahashi and Arkady Borisov. Decision strategies in evolutionary optimization. Lecture Notes in Computer Science, 2206:345–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2206/22060345.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2206/22060345.
pdf.

Borgulya:2001:CBE

[750] István Borgulya. A cluster-based evolutionary algorithm for multiobjective optimization. Lecture Notes in Computer Science, 2206:357–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2206/22060357.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2206/22060357.
pdf.

Nagamune:2001:IWE

[751] Kouki Nagamune, Yutaka Hata, and Kazuhiko Taniguchi. Inherent wave estimation on ultrasonic non-destructive testing using fuzzy inference. Lecture Notes in Computer Science, 2206:369– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2206/22060369.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2206/22060369.
pdf.

Hullermeier:2001:FAR

[752] Eyke Hüllermeier. Fuzzy association rules: Semantic issues and quality measures. Lecture Notes in Computer Science, 2206:380–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2206/22060380.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2206/22060380.
                                        pdf.

Huang:2001:CVS

[753] Chongfu Huang, Claudio Moraga, and Xingguang Yuan. Calculation vs. subjective assessment with respect to fuzzy probability. Lecture Notes in Computer Science, 2206:392– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2206/22060392.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2206/22060392.
                                        pdf.

Renners:2001:MOF

[754] Ingo Renners, Adolf Grauel, and Ernesto Saavedra. Methodology for optimizing fuzzy classifiers based on computational intelligence. Lecture Notes in Computer Science, 2206:412– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2206/22060412.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2206/22060412.
                                        pdf.

Reuter:2001:ASP

[755] Matthias Reuter. Analysing the structure of poly-crystalline materials by 2-dimensional DLS-spectra and neural nets. Lecture Notes in Computer Science, 2206:420–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2206/22060420.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2206/22060420.
pdf.

Varshavsky:2001:NMT

[756] Victor Varshavsky and Vyacheslav Marakhovsky. A neuron-MOS threshold element with switching capacitors. Lecture Notes in Computer Science, 2206:430–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2206/22060430.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2206/22060430.
pdf.

Neagoe:2001:NAC

[757] Victor Neagoe. A neural approach to compression of hyperspectral remote sensing imagery. Lecture Notes in Computer Science, 2206:436–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2206/22060436.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2206/22060436.
pdf.

Aizenberg:2001:MVU

[758] Igor Aizenberg. Multi-valued and universal binary neurons: New applications in intelligent image processing. Lecture Notes in Computer Science, 2206:450–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2206/22060450.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2206/22060450.
                                        pdf.

Dorado:2001:HTP

[759] Julian Dorado, Antonino Santos, Juan R. Rabuñal, Nieves Pedreira, and Alejandro Pazos. Hybrid twopopulation genetic algorithm. Lecture Notes in Computer Science, 2206:464– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2206/22060464.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2206/22060464.
                                        pdf.

Monsieurs:2001:RBG

[760] Patrick Monsieurs and Eddy Flerackers. Reducing bloat in genetic programming. Lecture Notes in Computer Science, 2206:471–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2206/22060471.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2206/22060471.
                                        pdf.

Schmiedle:2001:TMK

[761] Frank Schmiedle, Daniel Große, Rolf Drechsler, and Bernd Becker. Too much knowledge hurts: Acceleration of genetic programs for learning heuristics. Lecture Notes in Computer Science, 2206:479–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2206/22060479.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2206/22060479.
pdf.

Saavedra:2001:GBH

[762] Ernesto Saavedra, Ingo Renners, Adolf Grauel, H. J. Convey, and A. Razak. Gaussians-based hybrid system for prediction and classification. Lecture Notes in Computer Science, 2206:492– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2206/22060492.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2206/22060492.
pdf.

Bershtein:2001:CPF

[763] Leonid Bershtein and Alexander Bozhenuk. A color problem for fuzzy graph. Lecture Notes in Computer Science, 2206:500–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2206/22060500.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2206/22060500.pdf.

Thiele:2001:UIF

[764] Helmut Thiele. On the uniqueness of interpretations for fuzzy IF– THEN rule bases. Lecture Notes in Computer Science, 2206:506–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2206/22060506.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2206/22060506.
                                        pdf.

Pykacz:2001:QCF

[765] Jarosław Pykacz, Bart D’Hooghe, and Roman R. Zapatrin. Quantum computers as fuzzy computers. Lecture Notes in Computer Science, 2206:526– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2206/22060526.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2206/22060526.
                                        pdf.

Wong:2001:ICS

[766] K. W. Wong, L. T. Kóczy, T. D. Gedeon, A. Chong, and D. Tikk. Improvement of the cluster searching algorithm in Sugeno and Yasukawa’s qualitative modeling approach. Lecture Notes in Computer Science, 2206:536– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2206/22060536.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2206/22060536.
                                        pdf.

Falkowski:2001:CPI

[767] B.-J. Falkowski and J. Nietzschmann. Can perceptrons identify users of body lotion? Lecture Notes in Computer Science, 2206:550–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2206/22060550.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2206/22060550.
pdf.

Martyna:2001:RLC

[768] Jerzy Martyna. Reinforcement learning for control of traffic and access points in intelligent wireless ATM networks. Lecture Notes in Computer Science, 2206: 554–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2206/22060554.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2206/22060554.
pdf.

Bostelmann:2001:CTD

[769] Sabine Bostelmann. Categorization of time delay by means of neural networks. Lecture Notes in Computer Science, 2206:565–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2206/22060565.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2206/22060565.
pdf.

Dadhe:2001:NNT

[770] Kai Dadhe, Volker Roßmann, Kazim Durmus, and Sebastian Engell. Neural networks as a tool for gray box modelling in reactive distillation. Lecture Notes in Computer Science, 2206:576– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2206/22060576.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2206/22060576.
pdf.

Nolle:2001:SWA

[771] Lars Nolle, Alec Goodyear, Adrian A. Hopgood, Phil D. Picton, and Nicholas StJ. Braithwaite. On step width adaptation in simulated annealing for continuous parameter optimisation. Lecture Notes in Computer Science, 2206:589–598, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2206/22060589.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2206/22060589.
pdf.

Gemeinder:2001:GBS

[772] Marcus Gemeinder and Michael Gerke. GA-based search for paths with minimum energy consumption for mobile robot systems. Lecture Notes in Computer Science, 2206:599–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2206/22060599.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2206/22060599.
pdf.

Shayesteh:2001:NMG

[773] Mahrokh G. Shayesteh, Mohammad B. Menhaj, and Babak G. Nobary. A new modified genetic algorithm for multiuser detection in DS/CDMA systems. Lecture Notes in Computer Science, 2206:608–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2206/22060608.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2206/22060608.
pdf.

Villmann:2001:CCD

[774] Thomas Villmann and Conny Albani. Clustering of categoric data in medicine — application of evolutionary algorithms. Lecture Notes in Computer Science, 2206:619–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2206/22060619.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2206/22060619.
pdf.

Zadeh:2001:NDA

[775] Lotfi A. Zadeh. A new direction in AI — toward a computational theory of perceptions. Lecture Notes in Computer Science, 2206:628–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2206/22060628.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2206/22060628.
                                        pdf.

Presser:2001:PNU

[776] Gero Presser. Personalization of newsletters using multistage fuzzy inference. Lecture Notes in Computer Science, 2206:629–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2206/22060629.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2206/22060629.pdf.

Damiani:2001:FTX

[777] Ernesto Damiani, Barbara Oliboni, and Letizia Tanca. Fuzzy techniques for XML data smushing. Lecture Notes in Computer Science, 2206:637– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2206/22060637.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2206/22060637.
                                        pdf.

Bernard:2001:NFB

[778] Thomas Bernard, Markoto Sajidman, and Helge-Björn Kuntze. A new fuzzybased multi-objective optimization concept for process control systems. Lecture Notes in Computer Science, 2206:653–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/

0558/bibs/2206/22060653.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2206/22060653.
pdf.

Fernandez:2001:ISP

[779] Felipe Fernández and Julio Gutiérrez. An improved shape-preserving affine Takagi-Sugeno controller using splines. Lecture Notes in Computer Science, 2206:671–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2206/22060671.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2206/22060671.
pdf.

Botia:2001:MML

[780] Juan A. Botía, Antonio F. Gómez-Skarmeta, Mercedes Valdés, and Antonio Padilla. METALA: A Metalearning Architecture. Lecture Notes in Computer Science, 2206:688–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2206/22060688.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2206/22060688.
pdf.

Wijesoma:2001:LSV

[781] W. Sardha Wijesoma, Mingming Ma, and K. W. Yue. On-line signature verification using a computational intelligence approach. Lecture Notes in Computer Science, 2206:699–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2206/22060699.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2206/22060699.
                                        pdf.

Krodel:2001:ADT

[782] Michael Krödel and Klaus-Dieter Kuhnert. Autonomous driving through intelligent image processing and machine learning. Lecture Notes in Computer Science, 2206:712–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2206/22060712.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2206/22060712.
                                        pdf.

Kovacs:2001:SRC

[783] Szilveszter Kovács. SVD reduction in continuos environment reinforcement learning. Lecture Notes in Computer Science, 2206:719–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2206/22060719.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2206/22060719.
                                        pdf.

Cotta:2001:ADA

[784] Carlos Cotta and José M. Troya. Analyzing directed acyclic graph recombination. Lecture Notes in Computer Science, 2206:739–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/

series/0558/bibs/2206/22060739.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2206/22060739.pdf.

Uchida:2001:RSB

[785] Hayato Uchida, Yutaka Hata, Shinro Matsuura, Tadahiro Tsuchikawa, Yoshio Morotomi, and Hideyasu Aoyama. Rough set based knowledge discovery of interface for the Internet usage among Japanese elderly women. Lecture Notes in Computer Science, 2206:749–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2206/22060749.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2206/22060749.
pdf.

Kalina:2001:FLF

[786] Martin Kalina. Fuzzy limits and fuzzy nearness relation. Lecture Notes in Computer Science, 2206:755–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2206/22060755.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2206/22060755.
pdf.

vanderZwaag:2001:HDR

[787] Berend-Jan van der Zwaag. Handwritten digit recognition: a neural network demo. Lecture Notes in Computer Science, 2206:762–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2206/22060762.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2206/22060762.
                                        pdf.

Kromer:2001:NNN

[788] Thomas Kromer. New neural nets. Lecture Notes in Computer Science, 2206: 772–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2206/22060772.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2206/22060772.
                                        pdf.

Butkiewicz:2001:IFM

[789] Bohdan S. Butkiewicz. Inference in fuzzy models of physical processes. Lecture Notes in Computer Science, 2206: 782–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2206/22060782.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2206/22060782.
                                        pdf.

Surmann:2001:ACF

[790] Hartmut Surmann. About the combination of functional approaches and fuzzy reasoning. Lecture Notes in Computer Science, 2206:791–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2206/22060791.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2206/22060791.
pdf.

Slobodova:2001:EUN

[791] Anna Slobodová. Expected utility networks in transferable belief model. Lecture Notes in Computer Science, 2206: 800–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2206/22060800.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2206/22060800.
pdf.

Corcoran:2001:DCR

[792] Jonathan Corcoran, Ian D. Wilson, Owen M. Lewis, and J. Andrew Ware. Data clustering and rule abduction to facilitate crime hot spot prediction. Lecture Notes in Computer Science, 2206:807–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2206/22060807.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2206/22060807.
pdf.

Gemeinder:2001:CWM

[793] Marcus Gemeinder. Computing with words: Multi-objective GAS for approximate reasoning. Lecture Notes in Computer Science, 2206:822–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2206/22060822.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2206/22060822.
                                        pdf.

Khadjiev:2001:DAD

[794] Djavvat Khadjiev and Funda Karaçal. The description of all ∨-distributive triangular norms of lengths 2 and 3. Lecture Notes in Computer Science, 2206: 829–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2206/22060829.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2206/22060829.
                                        pdf.

Nakov:2001:LSA

[795] Preslav Nakov. Latent semantic analysis for German literature investigation. Lecture Notes in Computer Science, 2206:834–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2206/22060834.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2206/22060834.
                                        pdf.

Tipans:2001:SPI

[796] Eriks Tipans and Arkady Borisov. Some possibilities of improving the CORA classification algorithm. Lecture Notes in Computer Science, 2206: 842–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2206/22060842.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2206/22060842.
pdf.

Yilmaz:2001:CAS

[797] Oğuzhan Yılmaz, Günseli Görür, and Türkay Dereli. Computer aided selection of cutting parameters by using fuzzy logic. Lecture Notes in Computer Science, 2206:854–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2206/22060854.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2206/22060854.
pdf.

Neumann:2001:PMC

[798] Joachim Neumann, Görge Deerberg, Stefan Schlüter, and Hans Fahlenkamp. Process monitoring in chemical plants using neural networks. Lecture Notes in Computer Science, 2206:871–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2206/22060871.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2206/22060871.
pdf.

Dumitrescu:2001:FPP

[799] Mariana Dumitrescu and Toader Munteanu. Fuzzy probability and power system safety. Lecture Notes in Computer Science, 2206:885–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2206/22060885.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2206/22060885.
                                        pdf.

Dumitrescu:2001:FLS

[800] Mariana Dumitrescu and Toader Munteanu. A fuzzy logic system used in safety analysis. Lecture Notes in Computer Science, 2206:890–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2206/22060890.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2206/22060890.
                                        pdf.

Sosnowski:2001:CFR

[801] Zenon A. Sosnowski. Chaining of fuzzy rules in RETE network. Lecture Notes in Computer Science, 2206:895– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2206/22060895.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2206/22060895.
                                        pdf.

Calin:2001:FRC

[802] Marius Cãlin and Dan Gâlea. A fuzzy relation for comparing intervals. Lecture Notes in Computer Science, 2206: 904–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2206/22060904.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2206/22060904.
                                        pdf.

Kromer:2001:TFN

[803] Thomas Kromer. Tomography in fractal neural nets. Lecture Notes in Computer Science, 2206:917–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2206/22060917.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2206/22060917.
pdf.

Chen:2001:AAN

[804] Eric Y. Chen. AEGIS: An activenetwork-powered defense mechanism against DDoS attacks. Lecture Notes in Computer Science, 2207:1– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2207/22070001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2207/22070001.
pdf.

Lefevre:2001:ANS

[805] Laurent Lefèvre, Cong duc Pham, Pascale Primet, Bernard Tourancheau, Benjamin Gaidioz, Jean-Patrick Gelas, and Moufida Maimour. Active networking support for the Grid. Lecture Notes in Computer Science, 2207:16– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2207/22070016.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2207/22070016.
pdf.

Egawa:2001:FSP

[806] Takashi Egawa, Koji Hino, and Yohei Hasegawa. Fast and secure packet processing environment for per-packet QoS customization. Lecture Notes in Computer Science, 2207:34–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2207/22070034.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2207/22070034.
pdf.

Hasegawa:2001:PRT

[807] Toru Hasegawa, Shigehiro Ano, Koji Nakao, and Fumito Kubota. Programmable remote traffic monitoring method using active network approach. Lecture Notes in Computer Science, 2207:49–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2207/22070049.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2207/22070049.
pdf.

Carle:2001:DAV

[808] Georg Carle, Henning Sanneck, Sebastian Zander, and Long Le. Deploying an active voice application on a three-level active network node architecture. Lecture Notes in Computer Science, 2207:65–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2207/22070065.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2207/22070065.
pdf.

Simpson:2001:CSH

[809] Steven Simpson, Mark Banfield, Paul Smith, and David Hutchison. Component selection for heterogeneous active networking. Lecture Notes in Computer Science, 2207:84–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2207/22070084.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2207/22070084.
pdf.

Anagnostakis:2001:PNA

[810] Kostas G. Anagnostakis, Sotiris Ioannidis, Stefan Miltchev, and Jonathan M. Smith. Practical network applications on a lightweight active management environment. Lecture Notes in Computer Science, 2207:101–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2207/22070101.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2207/22070101.
pdf.

Tani:2001:ASM

[811] Seiichiro Tani, Toshiaki Miyazaki, and Noriyuki Takahashi. Adaptive stream multicast based on IP unicast and dynamic commercial attachment mechanism: An active network implementation. Lecture Notes in Computer Science, 2207:116–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2207/22070116.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2207/22070116.
                                        pdf.

Hicks:2001:CPS

[812] Michael Hicks, Jonathan T. Moore, and Scott Nettles. Compiling PLAN to SNAP. Lecture Notes in Computer Science, 2207:134–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2207/22070134.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2207/22070134.
                                        pdf.

Macian:2001:ONM

[813] Carlos Macián. The Octopus Network Model: Opening up the Internet to active and programmable network implementations. Lecture Notes in Computer Science, 2207:152– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2207/22070152.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2207/22070152.
                                        pdf.

Anonymous:2001:AIt

[814] Anonymous. Author index. Lecture Notes in Computer Science, 2207:165– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2207/22070165.pdf.

Mori:2001:MTC

[815] Kensaku Mori, Daisuke Deguchi, Jun ichi Hasegawa, Yasuhito Suenaga, Jun ichiro Toriwaki, Hirotsugu Takabatake, and Hiroshi Natori. A method for tracking the camera motion of real endoscope by epipolar geometry analysis and virtual endoscopy system. Lecture Notes in Computer Science, 2208: 1–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22080001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22080001.
pdf.

Nishikawa:2001:RTV

[816] Atsushi Nishikawa, Toshinori Hosoi, Kengo Koara, Daiji Negoro, Ayae Hikita, Shuichi Asano, Fumio Miyazaki, Mitsugu Sekimoto, Yasuhiro Miyake, Masayoshi Yasui, and Morito Monden. Real-time visual tracking of the surgeon’s face for laparoscopic surgery. Lecture Notes in Computer Science, 2208:9–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22080009.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22080009.
pdf.

Ron:2001:CBP

[817] Ofer Ron, Leo Joskowicz, Ariel Simkin, and Charles Milgrom. Computerbased periaxial rotation measurement for aligning fractured femur fragments: Method and preliminary results. Lecture Notes in Computer Science, 2208: 17–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22080017.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22080017.
pdf.

Gerig:2001:SVS

[818] Guido Gerig, Martin Styner, Martha E. Shenton, and Jeffrey A. Lieberman. Shape versus size: Improved understanding of the morphology of brain structures. Lecture Notes in Computer Science, 2208:24–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22080024.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22080024.
pdf.

Bouix:2001:HSA

[819] Sylvain Bouix, Jens C. Pruessner, Donald L. Collins, and Kaleem Siddiqi. Hippocampal shape analysis using medial surfaces. Lecture Notes in Computer Science, 2208:33– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22080033.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2208/22080033.
pdf.

Studholme:2001:DSC

[820] C. Studholme, V. Cardenas, N. Schuff, H. Rosen, B. Miller, and M. Weiner. Detecting spatially consistent structural differences in Alzheimer’s and fronto temporal dementia using deformation morphometry. Lecture Notes in Computer Science, 2208:41– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22080041.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22080041.
pdf.

Holden:2001:QSC

[821] Mark Holden, Julia A. Schnabel, and Derek L. G. Hill. Quantifying small changes in brain ventricular volume using non-rigid registration. Lecture Notes in Computer Science, 2208:49– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22080049.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22080049.
pdf.

Davies:2001:EMC

[822] Rhodri H. Davies, Tim F. Cootes, John C. Waterton, and Chris J. Taylor. An efficient method for constructing optimal statistical shape models. Lecture Notes in Computer Science, 2208: 57–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2208/22080057.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2208/22080057.
                                        pdf.

Hamarneh:2001:DOA

[823] Ghassan Hamarneh, Tim McInerney, and Demetri Terzopoulos. Deformable organisms for automatic medical image analysis. Lecture Notes in Computer Science, 2208:66–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2208/22080066.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2208/22080066.
                                        pdf.

Rueckert:2001:ACS

[824] D. Rueckert, A. F. Frangi, and J. A. Schnabel. Automatic construction of 3D statistical deformation models using non-rigid registration. Lecture Notes in Computer Science, 2208:77– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2208/22080077.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2208/22080077.
                                        pdf.

Gokturk:2001:LMA

[825] S. B. Göktürk, C. Tomasi, B. Acar, D. Paik, C. Beaulieu, and S. Napel. A learning method for automated polyp detection. Lecture Notes in Computer Science, 2208:85–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22080085.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22080085.
pdf.

Kita:2001:CSL

[826] Yasuyo Kita, Eriko Tohno, Ralph P. Highnam, and Michael Brady. A CAD system for 3D locating of lesions in mammogram. Lecture Notes in Computer Science, 2208:94–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22080094.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22080094.
pdf.

Kawata:2001:APN

[827] Y. Kawata, N. Niki, H. Ohmatsu, M. Kusumoto, R. Kakinuma, K. Mori, H. Nishiyama, K. Eguchi, M. Kaneko, and N. Moriyama. Analysis of pulmonary nodule evolutions using a sequence of three-dimensional thoracic CT images. Lecture Notes in Computer Science, 2208:103–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22080103.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22080103.
pdf.

Rohlfing:2001:IBN

[828] Torsten Rohlfing and Calvin R. Maurer, Jr. Intensity-based non-rigid registration using adaptive multilevel free-form deformation with an incompressibility constraint. Lecture Notes in Computer Science, 2208:111– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2208/22080111.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2208/22080111.
                                        pdf.

Haker:2001:MPM

[829] Steven Haker, Allen Tannenbaum, and Ron Kikinis. Mass preserving mappings and image registration. Lecture Notes in Computer Science, 2208:120– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2208/22080120.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2208/22080120.
                                        pdf.

Kauer:2001:IFE

[830] M. Kauer, V. Vuskovic, J. Dual, G. Szekely, and M. Bajka. Inverse finite element characterization of soft tissues. Lecture Notes in Computer Science, 2208:128–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2208/22080128.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2208/22080128.
pdf.

Brown:2001:MSS

[831] Joel Brown, Kevin Montgomery, Jean-Claude Latombe, and Michael Stephanides. A microsurgery simulation system. Lecture Notes in Computer Science, 2208:137–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22080137.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22080137.
pdf.

Nienhuys:2001:SSS

[832] Han-Wen Nienhuys and A. Frank van der Stappen. A surgery simulation supporting cuts and finite element deformation. Lecture Notes in Computer Science, 2208:145–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22080145.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22080145.
pdf.

Cebral:2001:PSS

[833] Juan R. Cebral, Rainald Löhner, Orlando Soto, Peter L. Choyke, and Peter J. Yim. Patient-specific simulation of carotid artery stenting using computational fluid dynamics. Lecture Notes in Computer Science, 2208:153– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2208/22080153.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2208/22080153.
                                        pdf.

Bullitt:2001:VRS

[834] Elizabeth Bullitt and Stephen Aylward. Volume rendering of segmented tubular objects. Lecture Notes in Computer Science, 2208:161–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2208/22080161.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2208/22080161.
                                        pdf.

Truyen:2001:CEA

[835] Roel Truyen, Thomas Deschamps, and Laurent D. Cohen. Clinical evaluation of an automatic path tracker for virtual colonoscopy. Lecture Notes in Computer Science, 2208:169–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2208/22080169.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2208/22080169.
                                        pdf.

Meijering:2001:EDT

[836] Erik Meijering, Wiro Niessen, Joachim Weickert, and Max Viergever. Evaluation of diffusion techniques for improved vessel visualization and quantification in three-dimensional rotational angiography. Lecture Notes in Computer Science, 2208:177–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22080177.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22080177.
pdf.

Mangin:2001:ECD

[837] J.-F. Mangin, C. Poupon, C. Clark, D. Le Bihan, and I. Bloch. Eddycurrent distortion correction and robust tensor estimation for MR diffusion imaging. Lecture Notes in Computer Science, 2208:186–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22080186.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22080186.
pdf.

Hahn:2001:EPR

[838] Klaus Hahn, Sergei Prigarin, and Benno Pütz. Edge preserving regularization and tracking for diffusion tensor imaging. Lecture Notes in Computer Science, 2208:195–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22080195.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22080195.
pdf.

VanLeemput:2001:SFP

[839] Koen Van Leemput, Frederik Maes, Dirk Vandermeulen, and Paul Suetens.

A statistical framework for partial volume segmentation. Lecture Notes in Computer Science, 2208:204–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22080204.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22080204.
pdf.

Elangovan:2001:SSD

[840] Vidya Elangovan and Ross T. Whitaker. From sinograms to surfaces: a direct approach to the segmentation of tomographic data. Lecture Notes in Computer Science, 2208:213–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22080213.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22080213.
pdf.

Sermesant:2001:EMM

[841] M. Sermesant, Y. Coudière, H. Delingette, N. Ayache, and J. A. Désidéri. An electro-mechanical model of the heart for cardiac image analysis. Lecture Notes in Computer Science, 2208:224– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22080224.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22080224.
pdf.

Surry:2001:MTD

[842] Kathleen J. M. Surry, Wendy L. Smith, Gregory R. Mills, Donal B. Downey, and Aaron Fenster. A mechanical, three-dimensional, ultrasound-guided breast biopsy apparatus. Lecture Notes in Computer Science, 2208:232– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22080232.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22080232.
pdf.

Rosenthal:2001:ARG

[843] Michael Rosenthal, Andrei State, Joohi Lee, Gentaro Hirota, Jeremy Ackerman, Kurtis Keller, Etta D. Pisano, Michael Jiroutek, Keith Muller, and Henry Fuchs. Augmented reality guidance for needle biopsies: a randomized, controlled trial in phantoms. Lecture Notes in Computer Science, 2208:240– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22080240.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22080240.
pdf.

Patriciu:2001:RKS

[844] Alexandru Patriciu, Stephen Solomon, Louis Kavoussi, and Dan Stoianovici. Robotic kidney and spine percutaneous procedures using a new laserbased CT registration method. Lecture Notes in Computer Science, 2208: 249–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22080249.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22080249.
pdf.

Hellier:2001:REI

[845] P. Hellier, C. Barillot, I. Corouge, B. Gibaud, G. Le Goualher, L. Collins, A. Evans, G. Malandain, and N. Ayache. Retrospective evaluation of intersubject brain registration. Lecture Notes in Computer Science, 2208:258– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22080258.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22080258.
pdf.

Warfield:2001:BEM

[846] Simon K. Warfield, Jan Rexilius, Petra S. Huppi, Terrie E. Inder, Erik G. Miller, William M. Wells III, Gary P. Zientara, Ferenc A. Jolesz, and Ron Kikinis. A binary entropy measure to assess nonrigid registration algorithms. Lecture Notes in Computer Science, 2208:266–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22080266.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22080266.
pdf.

Grova:2001:MVM

[847] Christophe Grova, Arnaud Biraben, Jean-Marie Scarabin, Pierre Jannin, Irène Buvat, Habib Benali, and Bernard Gibaud. A methodology to validate MRI/SPECT registration methods using realistic simulated SPECT data. Lecture Notes in Computer Science, 2208:275–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22080275.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22080275.
pdf.

Treece:2001:CPP

[848] Graham Treece, Richard Prager, Andrew Gee, and Laurence Berman. Correction of probe pressure artifacts in freehand 3D ultrasound. Lecture Notes in Computer Science, 2208:283– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22080283.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22080283.
pdf.

Rhode:2001:VVN

[849] Kawal Rhode, Gareth Ennew, Tryphon Lambrou, Alexander Seifalian, and David Hawkes. In-vitro validation of a novel model-based approach to the measurement of arterial blood flow waveforms from dynamic digital X-ray images. Lecture Notes in Computer Science, 2208:291–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2208/22080291.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2208/22080291.
                                        pdf.

Behiels:2001:RCH

[850] G. Behiels, F. Maes, D. Vandermeulen, and P. Suetens. Retrospective correction of the heel effect in hand radiographs. Lecture Notes in Computer Science, 2208:301– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2208/22080301.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2208/22080301.
                                        pdf.

Duchemin:2001:SSM

[851] Gilles Duchemin, Etienne Dombre, François Pierrot, Philippe Poignet, and Eric Dégoulange. SCALPP: a safe methodology to robotize skin harvesting. Lecture Notes in Computer Science, 2208:309–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2208/22080309.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2208/22080309.
                                        pdf.

Thakral:2001:SMA

[852] Anshul Thakral, Jeffrey Wallace, Damian Tomlin, Nikesh Seth, and Nitish V. Thakor. Surgical Motion Adaptive Robotic Technology (S.M.A.R.T): Taking the motion out of physiological motion. Lecture Notes in Computer Science, 2208:317–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22080317.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22080317.
pdf.

Gonzales:2001:TSR

[853] Adriana Vilchis Gonzales, Philippe Cinquin, Jocelyne Troccaz, Agnès Guerraz, Bernard Hennion, Franck Pellissier, Pierre Thorel, Fabien Courreges, Alain Gourdon, Gérard Poisson, Pierre Vieyres, Pierre Caron, Olivier Mérigeaux, Loïc Urbain, Cédric Daimo, Stéphane Lavallée, Philippe Arbeille, Marc Althuser, Jean-Marc Ayoubi, Bertrand Tondu, and Serge Ippolito. TER: a system for robotic tele-echography. Lecture Notes in Computer Science, 2208:326–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22080326.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22080326.
pdf.

Ma:2001:DAB

[854] Xin Ma, Zhong Fan, Chee-Kong Chui, Yiyu Cai, James H. Anderson, and Wieslaw L. Nowinski. Digital angioplasty balloon inflation device for interventional cardiovascular procedures. Lecture Notes in Computer Science, 2208:335–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2208/22080335.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2208/22080335.
                                        pdf.

Megali:2001:CAR

[855] Giuseppe Megali, Oliver Tonet, Cesare Stefanini, Mauro Boccadoro, Vassilios Papaspyropoulos, Licinio Angelini, and Paolo Dario. A computer-assisted robotic ultrasound-guided biopsy system for video-assisted surgery. Lecture Notes in Computer Science, 2208:343– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2208/22080343.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2208/22080343.
                                        pdf.

Cheng:2001:DBP

[856] Gang Cheng, Haisong Liu, Lydia Liao, and Yan Yu. Dynamic brachytherapy of the prostate under active image guidance. Lecture Notes in Computer Science, 2208:351–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2208/22080351.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2208/22080351.
                                        pdf.

Gomes:2001:CAS

[857] M. P. S. F. Gomes, A. R. W. Barrett, and B. L. Davies. Computerassisted soft-tissue surgery training and monitoring. Lecture Notes in Computer Science, 2208:360–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22080360.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22080360.
pdf.

Lehmann:2001:TDP

[858] G. Lehmann, A. Chiu, D. Gobbi, Y. Starreveld, D. Boyd, M. Drangova, and T. Peters. Towards dynamic planning and guidance of minimally invasive robotic cardiac bypass surgical procedures. Lecture Notes in Computer Science, 2208:368– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22080368.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22080368.
pdf.

Tomlin:2001:NRT

[859] Damian Tomlin, Jeffrey Wallace, Ralph Etienne-Cummings, and Nitish Thakor. Novel real-time tremor transduction technique for microsurgery. Lecture Notes in Computer Science, 2208:376–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2208/22080376.htm;   http:

                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2208/22080376.
                                        pdf.

Dutreuil:2001:CAD

[860] Julien Dutreuil, François Goulette, Claude Laurgeau, Jaime Clavero Zoreda, and Stefan Lundgren. Computer assisted dental implantology: a new method and a clinical validation. Lecture Notes in Computer Science, 2208:384–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2208/22080384.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2208/22080384.
                                        pdf.

Liao:2001:IOR

[861] Hongen Liao, Susumu Nakajima, Makoto Iwahara, Etsuko Kobayashi, Ichiro Sakuma, Naoki Yahagi, and Takeyoshi Dohi. Intra-operative realtime 3-D information display system based on integral videography. Lecture Notes in Computer Science, 2208:392– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2208/22080392.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2208/22080392.
                                        pdf.

Vogel:2001:TMR

[862] M. W. Vogel, Suprijanto, F. M. Vos, H. A. Vrooman, A. M. Vossepoel, and P. M. T. Pattynama. Towards motion-robust magnetic resonance thermometry. Lecture Notes in Computer Science, 2208:401–408, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22080401.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22080401.
pdf.

Kunz:2001:NCB

[863] Manuela Kunz, Matthias Strauss, Frank Langlotz, Georg Deuretzbacher, Wolfgang Rüther, and Lutz-Peter Nolte. A non-CT based total knee arthroplasty system featuring complete soft-tissue balancing. Lecture Notes in Computer Science, 2208:409– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22080409.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22080409.
pdf.

Shen:2001:SAP

[864] Dinggang Shen, Zhiqiang Lao, Jianchao Zeng, Edward H. Herskovits, Gabor Fichtinger, and Christos Davatzikos. A statistical atlas of prostate cancer for optimal biopsy. Lecture Notes in Computer Science, 2208:416– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22080416.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22080416.
pdf.

Serby:2001:NAC

[865] D. Serby, M. Harders, and G. Székely. A new approach to cutting into finite element models. Lecture Notes in Computer Science, 2208:425–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2208/22080425.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2208/22080425.
                                        pdf.

Xia:2001:MPS

[866] James Xia, Jaime Gateno, John Teichgraeber, and Andrew Rosen. Methodology of precise skull model creation. Lecture Notes in Computer Science, 2208: 434–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2208/22080434.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2208/22080434.
                                        pdf.

Dean:2001:MAS

[867] David Dean, Pengpeng Zhang, Andrew K. Metzger, Claudio Sibata, and Robert J. Maciunas. Medial axis seeding of a guided evolutionary simulated annealing (GESA) algorithm for automated gamma knife radiosurgery treatment planning. Lecture Notes in Computer Science, 2208:441– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2208/22080441.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2208/22080441.
pdf.

Wilson:2001:SFC

[868] Nathan Wilson, Kenneth Wang, Robert W. Dutton, and Charles Taylor. A software framework for creating patient specific geometric models from medical imaging data for simulation based medical planning of vascular surgery. Lecture Notes in Computer Science, 2208:449–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22080449.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22080449.
pdf.

Li:2001:ICS

[869] Zirui Li, Chee-Kong Chui, Yiyu Cai, James H. Anderson, and Wieslaw L. Nowinski. Interactive catheter shape modeling in interventional radiology simulation. Lecture Notes in Computer Science, 2208:457–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22080457.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22080457.
pdf.

Cai:2001:PEM

[870] Yiyu Cai, Chee-Kong Chui, Yaoping Wang, Zhenlan Wang, and James H. Anderson. Parametric eyeball model for interactive simulation of ophthalmologic surgery. Lecture Notes in Computer Science, 2208:465–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2208/22080465.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2208/22080465.
                                        pdf.

Zachow:2001:IOP

[871] Stefan Zachow, Evgeny Gladilin, Hans-Florian Zeilhofer, and Robert Sader. Improved 3D osteotomy planning in cranio-maxillofacial surgery. Lecture Notes in Computer Science, 2208:473– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2208/22080473.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2208/22080473.
                                        pdf.

Jolly:2001:CER

[872] Marie-Pierre Jolly. Combining edge, region, and shape information to segment the left ventricle in cardiac MR images. Lecture Notes in Computer Science, 2208:482–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2208/22080482.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2208/22080482.
                                        pdf.

Young:2001:VSV

[873] S. Young, V. Pekar, and J. Weese. Vessel segmentation for visualization of MRA with blood pool contrast agent. Lecture Notes in Computer Science, 2208:491–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22080491.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22080491.
pdf.

Bernard:2001:SAS

[874] Rok Bernard, Bostjan Likar, and Franjo Pernus. Segmenting articulated structures by hierarchical statistical modeling of shape, appearance, and topology. Lecture Notes in Computer Science, 2208:499–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22080499.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22080499.
pdf.

Srivastava:2001:VNS

[875] Siddharth Srivastava, Koen Van Leemput, Frederik Maes, Dirk Vandermeulen, and Paul Suetens. Validation of nonlinear spatial filtering to improve tissue segmentation of MR brain images. Lecture Notes in Computer Science, 2208:507–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22080507.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2208/22080507.
                                        pdf.

Gerig:2001:VNV

[876] Guido Gerig, Matthieu Jomier, and Miranda Chakos. Valmet: a new validation tool for assessing and improving 3D object segmentation. Lecture Notes in Computer Science, 2208:516– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2208/22080516.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2208/22080516.
                                        pdf.

Wei:2001:DDP

[877] Guo-Qing Wei, JianZhong Qian, and Helmuth Schramm. A dual dynamic programming approach to the detection of spine boundaries. Lecture Notes in Computer Science, 2208:524– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2208/22080524.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2208/22080524.
                                        pdf.

Sarrut:2001:GTA

[878] David Sarrut and Sébastien Clippe. Geometrical transformation approximation for 2D/3D intensity-based registration of portal images and CT scan. Lecture Notes in Computer Science, 2208:532–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/

0558/bibs/2208/22080532.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2208/22080532.
pdf.

Lotjonen:2001:EMU

[879] J. Lötjönen and T. Mäkelä. Elastic matching using a deformation sphere. Lecture Notes in Computer Science, 2208:541–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2208/22080541.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2208/22080541.
pdf.

Butz:2001:ARF

[880] Torsten Butz and Jean-Philippe Thiran. Affine registration with feature space mutual information. Lecture Notes in Computer Science, 2208:549– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22080549.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22080549.
pdf.

Makela:2001:NMR

[881] Timo Mäkelä, Patrick Clarysse, Jyrki Lötjönen, Outi Sipilä, Kirsi Lauerma, Helena Hänninen, Esa-Pekka Pyökkimies, Jukka Nenonen, Juhani Knuuti, Toivo Katila, and Isabelle E. Magnin. A new method for the registration of cardiac PET and MR images using deformable model based segmentation of the main thorax structures. Lecture Notes in Computer Science, 2208:

557–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22080557.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22080557.
pdf.

Jannin:2001:MSP

[882] P. Jannin, M. Raimbault, X. Morandi, and B. Gibaud. Modeling surgical procedures for multimodal imageguided neurosurgery. Lecture Notes in Computer Science, 2208:565–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22080565.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22080565.
pdf.

Schnabel:2001:GFN

[883] Julia A. Schnabel, Daniel Rueckert, Marcel Quist, Jane M. Blackall, Andy D. Castellano-Smith, Thomas Hartkens, Graeme P. Penney, Walter A. Hall, Haiying Liu, Charles L. Truwit, Frans A. Gerritsen, Derek L. G. Hill, and David J. Hawkes. A generic framework for non-rigid registration based on non-uniform multilevel free-form deformations. Lecture Notes in Computer Science, 2208:573– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22080573.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2208/22080573.
pdf.

Alker:2001:IREb

[884] Manfred Alker, Sönke Frantz, Karl Rohr, and H. Siegfried Stiehl. Improving the robustness in extracting 3D point landmarks from 3D medical images using parametric deformable models. Lecture Notes in Computer Science, 2208:582–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22080582.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22080582.
pdf.

Hein:2001:ICE

[885] Andreas Hein, Martin Klein, Tim C. Lueth, Jochen Queck, Malte Stien, Olaf Schermeier, and Juergen Bier. Integration and clinical evaluation of an interactive controllable robotic system for anaplastology. Lecture Notes in Computer Science, 2208:591–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22080591.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22080591.
pdf.

Brzeczko:2001:SAT

[886] Albert W. Brzeczko, Randal P. Goldberg, Russell H. Taylor, and Peter Evans. Smart alignment tool for knee MosaicPlasty surgery. Lecture Notes in Computer Science, 2208:599– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22080599.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22080599.
pdf.

Nakamura:2001:MDF

[887] Ryoichi Nakamura, Takeshi Oura, Etsuko Kobayashi, Ichiro Sakuma, Takeyoshi Dohi, Naoki Yahagi, Takayuki Tsuji, Mitsuo Shimada, and Makoto Hashizume. Multi-DOF forceps manipulator system for laparoscopic surgery — mechanism miniaturized & evaluation of new interface. Lecture Notes in Computer Science, 2208:606– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22080606.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22080606.
pdf.

Zivanovic:2001:DHR

[888] Aleksandar Zivanovic and Brian Davies. The development of a haptic robot to take blood samples from the forearm. Lecture Notes in Computer Science, 2208:614–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22080614.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22080614.
pdf.

Kiss:2001:CAD

[889] Gabriel Kiss, Johan Van Cleynenbreugel, Maarten Thomeer, Paul Suetens, and Guy Marchal. Computer aided diagnosis for virtual colonography. Lecture Notes in Computer Science, 2208:621–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2208/22080621.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2208/22080621.pdf.

Linguraru:2001:FHI

[890] Marius George Linguraru, Michael Brady, and Margaret Yam. Filtering h int images for the detection of microcalcifications. Lecture Notes in Computer Science, 2208:629– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22080629.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22080629.
pdf.

Acar:2001:UOF

[891] Burak Acar, Sandy Napel, David Paik, Burak Göktürk, Carlo Tomasi, and Christopher F. Beaulieu. Using optical flow fields for polyp detection in virtual colonoscopy. Lecture Notes in Computer Science, 2208:637– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22080637.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2208/22080637.
                                        pdf.

Vos:2001:NVM

[892] F. M. Vos, I. W. O. Serlie, R. E. van Gelder, F. H. Post, R. Truyen, F. A. Gerritsen, J. Stoker, and A. M. Vossepoel. A new visualization method for virtual colonoscopy. Lecture Notes in Computer Science, 2208:645–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2208/22080645.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2208/22080645.
                                        pdf.

Liu:2001:CDP

[893] Y. Liu, F. Dellaert, W. E. Rothfus, A. Moore, J. Schneider, and T. Kanade. Classification-driven pathological neuroimage retrieval using statistical asymmetry measures. Lecture Notes in Computer Science, 2208: 655–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2208/22080655.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2208/22080655.
                                        pdf.

Stoeckel:2001:CSI

[894] Jonathan Stoeckel, Grégoire Malandain, Octave Migneco, Pierre Malick Koulibaly, Philippe Robert, Nicholas Ayache, and Jacques Darcourt. Classification of SPECT images of normal subjects versus images of Alzheimer’s disease patients. Lecture Notes in Computer Science, 2208:666–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22080666.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22080666.
pdf.

Argotti:2001:TAR

[895] Yann Argotti, Valerie Outters, Larry Davis, Ami Sun, and Jannick P. Rolland. Technologies for augmented reality: Calibration for real-time superimposition on rigid and simpledeformable real objects. Lecture Notes in Computer Science, 2208:675–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22080675.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22080675.
pdf.

Stetten:2001:MRT

[896] George Stetten and Vikram Chib. Magnified real-time tomographic reflection. Lecture Notes in Computer Science, 2208:683–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2208/22080683.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2208/22080683.
pdf.

Buck:2001:SSL

[897] Stijn De Buck, Johan Van Cleynenbreugel, Indra Geys, Thomas Koninckx, Philippe R. Koninck, and Paul Suetens. A system to support laparoscopic surgery by augmented reality visualization. Lecture Notes in Computer Science, 2208:691–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22080691.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22080691.
pdf.

vanBemmel:2001:BPA

[898] C. M. van Bemmel, W. J. Niessen, O. Wink, B. Verdonck, and M. A. Viergever. Blood pool agent CE-MRA: Improved arterial visualization of the aortoiliac vasculature in the steadystate using first-pass data. Lecture Notes in Computer Science, 2208:699– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22080699.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22080699.
pdf.

Sauer:2001:HMD

[899] F. Sauer, A. Khamene, B. Bascle, and G. J. Rubino. A head-mounted display system for augmented reality image guidance: Towards clinical evaluation for iMRI-guided neurosurgery. Lecture Notes in Computer Science, 2208:707– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22080707.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2208/22080707.
pdf.

Lee:2001:NAR

[900] Sangyoon Lee, Gabor Fichtinger, and Gregory S. Chirikjian. Novel algorithms for robust registration of fiducials in CT and MRI. Lecture Notes in Computer Science, 2208:717– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22080717.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22080717.
pdf.

Betke:2001:ARL

[901] Margrit Betke, Harrison Hong, and Jane P. Ko. Automatic 3D registration of lung surfaces in computed tomography scans. Lecture Notes in Computer Science, 2208:725– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22080725.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22080725.
pdf.

Cachier:2001:MNR

[902] Pascal Cachier, Jean-François Mangin, Xavier Pennec, Denis Rivière, Dimitri Papadopoulos-Orfanos, Jean Régis, and Nicholas Ayache. Multisubject non-rigid registration of brain MRI using intensity and geometric features. Lecture Notes in Computer Science, 2208:734–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22080734.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22080734.
pdf.

Ourselin:2001:FHS

[903] S. Ourselin, E. Bardinet, D. Dormont, G. Malandain, A. Roche, N. Ayache, D. Tandé, K. Parain, and J. Yelnik. Fusion of histological sections and MR images: Towards the construction of an atlas of the human basal ganglia. Lecture Notes in Computer Science, 2208: 743–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22080743.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22080743.
pdf.

Granger:2001:RPS

[904] Sébastien Granger, Xavier Pennec, and Alexis Roche. Rigid pointsurface registration using an EM variant of ICP for computer guided oral implantology. Lecture Notes in Computer Science, 2208:752–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22080752.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22080752.
pdf.

Penney:2001:SIC

[905] G. P. Penney, P. J. Edwards, A. P. King, J. M. Blackall, P. G. Batchelor, and D. J. Hawkes. A stochastic iterative closest point algorithm (stochastICP). Lecture Notes in Computer Science, 2208:762–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22080762.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22080762.
pdf.

Wilkinson:2001:SPF

[906] Michael H. F. Wilkinson and Michel A. Westenberg. Shape preserving filament enhancement filtering. Lecture Notes in Computer Science, 2208:770– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22080770.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22080770.
pdf.

Ye:2001:FEA

[907] Xujiong Ye, J. Alison Noble, and Jerome Declerck. 3D freehand echocardiography for automatic left ventricle reconstruction and analysis based on multiple acoustic windows. Lecture Notes in Computer Science, 2208:778– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22080778.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2208/22080778.
                                        pdf.

Kerwin:2001:QVA

[908] William Kerwin, Chao Han, Baocheng Chu, Dongxiang Xu, Ying Luo, Jenq-Neng Hwang, Thomas Hatsukami, and Chun Yuan. A quantitative vascular analysis system for evaluation of atherosclerotic lesions by MRI. Lecture Notes in Computer Science, 2208: 786–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2208/22080786.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2208/22080786.
                                        pdf.

Hattery:2001:IML

[909] David Hattery, Victor Chernomordik, Amir Gandjbakhche, and Murray Loew. Imaging metabolism with light: Quantifying local fluorescence lifetime perturbation in tissuelike turbid media. Lecture Notes in Computer Science, 2208:795–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2208/22080795.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2208/22080795.
                                        pdf.

Sato:2001:LAT

[910] Yoshinobu Sato, Katsuyuki Nakanishi, Hisashi Tanaka, Takashi Nishii, Nobuhiko Sugano, Hironobu Nakamura, Takahiro Ochi, and Shinichi Tamura. Limits to the accuracy of 3D thickness measurement in magnetic resonance images. Lecture Notes in Computer Science, 2208:803–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22080803.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22080803.
pdf.

Prima:2001:MLE

[911] Sylvain Prima, Nicholas Ayache, Tom Barrick, and Neil Roberts. Maximum likelihood estimation of the bias field in MR brain images: Investigating different modelings of the imaging process. Lecture Notes in Computer Science, 2208:811–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22080811.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22080811.
pdf.

Bhalerao:2001:IVS

[912] Abhir Bhalerao, Elke Thönnes, Wilfrid Kendall, and Roland Wilson. Inferring vascular structure from 2D and 3D imagery. Lecture Notes in Computer Science, 2208:820–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22080820.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22080820.
pdf.

Stein:2001:SDM

[913] Benjamin Stein, Dimitri Lisin, Joseph Horowitz, Edward Riseman, and Gary Whitten. Statistical and deformable model approaches to the segmentation of MR imagery and volume estimation of stroke lesions. Lecture Notes in Computer Science, 2208:829– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2208/22080829.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2208/22080829.
                                        pdf.

Larsen:2001:QMS

[914] Rasmus Larsen, Hrafnkell Eiriksson, and Mikkel B. Stegmann. Q–MAF shape decomposition. Lecture Notes in Computer Science, 2208:837–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2208/22080837.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2208/22080837.
                                        pdf.

Wink:2001:VAD

[915] O. Wink, W. J. Niessen, B. Verdonck, and M. A. Viergever. Vessel axis determination using wave front propagation analysis. Lecture Notes in Computer Science, 2208:845– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2208/22080845.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2208/22080845.
pdf.

Desbleds-Mansard:2001:SIA

[916] Catherine Desbleds-Mansard, Alfred Anwander, Linda Chaabane, Maciej Orkisz, Bruno Neyran, Philippe C. Douek, and Isabelle E. Magnin. Size independent active contour model for blood vessel lumen quantification in high-resolution magnetic resonance images. Lecture Notes in Computer Science, 2208:854–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22080854.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22080854.
pdf.

Pizer:2001:SSF

[917] Stephen M. Pizer, Sarang Joshi, P. Thomas Fletcher, Martin Styner, Gregg Tracton, and James Z. Chen. Segmentation of single-figure objects by deformable M-reps. Lecture Notes in Computer Science, 2208:862–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22080862.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22080862.
pdf.

Degenhard:2001:HFA

[918] Andreas Degenhard, Christine Tanner, Carmel Hayes, David J. Hawkes, and Martin O. Leach. Half Fourier acquisition applied to time series analysis of contrast agent uptake. Lecture Notes in Computer Science, 2208:872– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22080872.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22080872.
pdf.

Thoraval:2001:ABF

[919] Laurent Thoraval, Jean-Paul Armspach, and Izzie Namer. Analysis of brain functional MRI time series based on continuous wavelet transform and stimulation-response coupling distance. Lecture Notes in Computer Science, 2208:881–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22080881.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22080881.
pdf.

Ledesma-Carbayo:2001:CMA

[920] María J. Ledesma-Carbayo, Jan Kybic, Manuel Desco, Andrés Santos, and Michael Unser. Cardiac motion analysis from ultrasound sequences using non-rigid registration. Lecture Notes in Computer Science, 2208:889– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22080889.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22080889.
pdf.

Cachia:2001:MCB

[921] A. Cachia, J.-F. Mangin, D. Rivière, N. Boddaert, A. Andrade, F. Kherif, P. Sonigo, D. Papadopoulos-Orfanos, M. Zilbovicius, J.-B. Poline, I. Bloch, F. Brunelle, and J. Régis. A mean curvature based primal sketch to study the cortical folding process from antenatal to adult brain. Lecture Notes in Computer Science, 2208:897–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22080897.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22080897.
pdf.

Fisher:2001:AER

[922] John W. Fisher III, Eric R. Cosman, Jr., Cindy Wible, and William M. Wells III. Adaptive entropy rates for fMRI time-series analysis. Lecture Notes in Computer Science, 2208:905– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22080905.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22080905.
pdf.

Arbel:2001:ANL

[923] Tal Arbel, Xavier Morandi, Roch M. Comeau, and D. Louis Collins. Automatic non-linear MRI-ultrasound registration for the correction of intraoperative brain deformations. Lecture Notes in Computer Science, 2208:913– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2208/22080913.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2208/22080913.
                                        pdf.

Rexilius:2001:NNR

[924] J. Rexilius, S. K. Warfield, C. R. G. Guttmann, X. Wei, R. Benson, L. Wolfson, M. Shenton, H. Handels, and R. Kikinis. A novel nonrigid registration algorithm and applications. Lecture Notes in Computer Science, 2208: 923–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2208/22080923.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2208/22080923.
                                        pdf.

Aylward:2001:APS

[925] Stephen R. Aylward, Sue Weeks, and Elizabeth Bullitt. Analysis of the parameter space of a metric for registering 3D vascular images. Lecture Notes in Computer Science, 2208:932– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2208/22080932.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2208/22080932.
                                        pdf.

Amin:2001:CMD

[926] Devin V. Amin, Takeo Kanade, Branislav Jaramaz, Anthony M. DiGioia III, Constantinos Nikou, Richard S. LaBarca, and James E. Moody, Jr. Calibration method for determining the physical location of the ultrasound image plane. Lecture Notes in Computer Science, 2208:940– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22080940.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22080940.
pdf.

Corouge:2001:NLL

[927] Isabelle Corouge, Christian Barillot, Pierre Hellier, Pierre Toulouse, and Bernard Gibaud. Non-linear local registration of functional data. Lecture Notes in Computer Science, 2208:948– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22080948.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22080948.
pdf.

Bardinet:2001:RRP

[928] E. Bardinet, A. C. F. Colchester, A. Roche, Y. Zhu, Y. He, S. Ourselin, B. Nailon, S. A. Hojjat, J. Ironside, S. Al-Sarraj, N. Ayache, and J. Wardlaw. Registration of reconstructed post mortem optical data with MR scans of the same patient. Lecture Notes in Computer Science, 2208:957– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22080957.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2208/22080957.
                                        pdf.

Kataoka:2001:MRB

[929] Hiroyuki Kataoka, Toshikatsu Washio, Michel Audette, and Kazuyuki Mizuhara. A model for relations between needle deflection, force, and thickness on needle penetration. Lecture Notes in Computer Science, 2208:966–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2208/22080966.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2208/22080966.
                                        pdf.

Ottensmeyer:2001:VDA

[930] Mark P. Ottensmeyer and J. Kenneth Salisbury, Jr. In vivo data acquisition instrument for solid organ mechanical property measurement. Lecture Notes in Computer Science, 2208: 975–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2208/22080975.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2208/22080975.
                                        pdf.

Mocanu:2001:PSS

[931] Daniel Mocanu, Joachim Kettenbach, Michael O. Sweeney, Bruce H. KenKnight, Ron Kikinis, and Solomon R. Eisenberg. Patient-specific simulation of internal defibrillation. Lecture Notes in Computer Science, 2208:983– 990, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22080983.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22080983.
pdf.

DeGroeve:2001:RPS

[932] Pieter De Groeve, Filip Schutyser, Johan Van Cleynenbreugel, and Paul Suetens. Registration of 3D photographs with spiral CT images for soft tissue simulation in maxillofacial surgery. Lecture Notes in Computer Science, 2208:991–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22080991.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22080991.
pdf.

Kukuk:2001:TP

[933] M. Kukuk, B. Geiger, and H. Müller. TBNA-protocols. Lecture Notes in Computer Science, 2208:997–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22080997.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22080997.
pdf.

Ahmed:2001:RHJ

[934] Moumen T. Ahmed, Ahmed H. Eid, and Aly A. Farag. 3D reconstruction of the human jaw: a new approach and improvements. Lecture Notes in Computer Science, 2208:1007– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22081007.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22081007.
pdf.

Suzuki:2001:RTS

[935] Naoki Suzuki, Asaki Hattori, Shigeyuki Suzuki, Max P. Baur, Andreas Hirner, Susumu Kobayashi, Yoji Yamazaki, and Yoshitaka Adachi. Real-time surgical simulation with haptic sensation as collaborated works between Japan and Germany. Lecture Notes in Computer Science, 2208:1015–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22081015.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22081015.
pdf.

ODonnell:2001:PBU

[936] Lauren O’Donnell, Carl-Fredrik Westin, W. Eric L. Grimson, Juan Ruiz-Alzola, Martha E. Shenton, and Ron Kikinis. Phase-based user-steered image segmentation. Lecture Notes in Computer Science, 2208:1022–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22081022.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22081022.
pdf.

Rodriguez-Florido:2001:CTR

[937] Miguel A. Rodriguez-Florido, Karl Krissian, Juan Ruiz-Alzola, and Carl-Fredrik Westin. Comparison of two restoration techniques in the context of 3D medical imaging. Lecture Notes in Computer Science, 2208:1031–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22081031.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22081031.
pdf.

Dawant:2001:RSM

[938] Benoit M. Dawant, Shiyan Pan, and Rui Li. Robust segmentation of medical images using geometric deformable models and a dynamic speed function. Lecture Notes in Computer Science, 2208:1040–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22081040.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22081040.
pdf.

Imielinska:2001:HSA

[939] Celina Imielinska, Dimitris Metaxas, Jayaram Udupa, Yinpeng Jin, and Ting Chen. Hybrid segmentation of anatomical data. Lecture Notes in Computer Science, 2208:1048–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22081048.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2208/22081048.
                                        pdf.

Boykov:2001:SDD

[940] Yuri Boykov, Vivian S. Lee, Henry Rusinek, and Ravi Bansal. Segmentation of dynamic N–D data sets via graph cuts using Markov models. Lecture Notes in Computer Science, 2208: 1058–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2208/22081058.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2208/22081058.
                                        pdf.

Pachai:2001:UAS

[941] Chahin Pachai, Yue Min Zhu, Charles R. G. Guttmann, Ron Kikinis, Ferenc A. Jolesz, Gérard Gimenez Jean-Claude Froment, Christian Confavreux, and Simon K. Warfield. Unsupervised and adaptive segmentation of multispectral 3D magnetic resonance images of human brain: a generic approach. Lecture Notes in Computer Science, 2208:1067–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2208/22081067.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2208/22081067.
                                        pdf.

Jaffray:2001:IGR

[942] David Jaffray, Marcel van Herk, Joos Lebesque, and Alvaro Martinez. Image guided radiotherapy of the prostate. Lecture Notes in Computer Science, 2208:1075–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22081075.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22081075.
pdf.

Lunn:2001:QCE

[943] Karen E. Lunn, Alex Hartov, Eric W. Hansen, Hai Sun, David W. Roberts, and Keith D. Paulsen. A quantitative comparison of edges in 3D intraoperative ultrasound and preoperative MR images of the brain. Lecture Notes in Computer Science, 2208:1081– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22081081.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22081081.
pdf.

Castellano-Smith:2001:CPS

[944] A. D. Castellano-Smith, T. Hartkens, J. Schnabel, D. R. Hose, H. Liu, W. A. Hall, C. L. Truwit, D. J. Hawkes, and D. L. G. Hill. Constructing patient specific models for correcting intraoperative brain deformation. Lecture Notes in Computer Science, 2208:1091– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22081091.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2208/22081091.
                                        pdf.

Paggetti:2001:IDE

[945] Cristiano Paggetti, Sandra Martelli, Laura Nofrini, and Paolo Vendruscolo. Interface design and evaluation for CAS systems. Lecture Notes in Computer Science, 2208:1099–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2208/22081099.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2208/22081099.
                                        pdf.

Hufner:2001:IRV

[946] T. Hüfner, M. Citak, S. Tarte, J. Geerling, T. Pohlemann, H. Rosenthal, L. P. Nolte, and C. Krettek. Independent registration and virtual controlled reduction of pelvic ring fractures. Lecture Notes in Computer Science, 2208:1107– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2208/22081107.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2208/22081107.
                                        pdf.

Sofer:2001:OPC

[947] Ariela Sofer, Jianchao Zeng, and Seong K. Mun. Optimization in prostate cancer detection. Lecture Notes in Computer Science, 2208:1114– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2208/22081114.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22081114.
pdf.

Taivalkoski:2001:AIB

[948] S. Taivalkoski, L. Jyrkinen, and O. Silvén. Automated identification and B-spline approximation of a profiling coil centerline from magnetic resonance images. Lecture Notes in Computer Science, 2208:1122–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22081122.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22081122.
pdf.

Tate:2001:PRA

[949] Peter M. Tate, Vladimir Lachine, Liqun Fu, Haniel Croitoru, and Marwan Sati. Performance and robustness of automatic fluoroscopic image calibration in a new computer assisted surgery system. Lecture Notes in Computer Science, 2208:1130–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22081130.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22081130.
pdf.

Liu:2001:ASN

[950] Alan Liu, Christoph Kaufmann, and Daigo Tanaka. An architecture for simulating needle-based surgical procedures. Lecture Notes in Computer Science, 2208:1137–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22081137.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22081137.
pdf.

Voirin:2001:CAH

[951] David Voirin, Yohan Payan, Miriam Amavizca, Antoine Leroy, Christian Létoublon, and Jocelyne Troccaz. Computer-aided hepatic tumour ablation. Lecture Notes in Computer Science, 2208:1145–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22081145.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22081145.
pdf.

VanCleynenbreugel:2001:IOT

[952] Johan Van Cleynenbreugel, Filip Schutyser, Chantal Malevez, Ellen Dhoore, Charbel BouSerhal, Reinhilde Jacobs, Paul Suetens, and Daniel van Steenberghe. Intra-operative transfer of planned zygomatic fixtures by personalized templates: a cadaver validation study. Lecture Notes in Computer Science, 2208:1147– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22081147.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2208/22081147.
pdf.

Surry:2001:PCB

[953] Kathleen J. M. Surry and Terry M. Peters. A PVA–C brain phantom derived from a high quality 3D MR data set. Lecture Notes in Computer Science, 2208:1149–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22081149.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22081149.
pdf.

Sato:2001:UIA

[954] Yoshinobu Sato, Masaki Miyamoto, Masahiko Nakamoto, Yoshikazu Nakajima, Mitsuo Shimada, Makoto Hashizume, and Shinichi Tamura. 3D ultrasound image acquisition using a magneto-optic hybrid sensor for laparoscopic surgery. Lecture Notes in Computer Science, 2208:1151–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22081151.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22081151.
pdf.

Ellis:2001:MIE

[955] R. E. Ellis, D. Kerr, J. F. Rudan, and L. Davidson. Minimally invasive excision of deep bone tumors. Lecture Notes in Computer Science, 2208:1154– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2208/22081154.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2208/22081154.
                                        pdf.

Kansy:2001:MNS

[956] K. Kansy, A. Schmitgen, M. Bublat, G. Grunst, M. Jungmann, P. Wisskirchen, M. Moche, G. Strauss, C. Trantakis, and T. Kahn. A multimodal navigation system for interventional MRI. Lecture Notes in Computer Science, 2208:1157–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2208/22081157.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2208/22081157.
                                        pdf.

Bestel:2001:BMM

[957] J. Bestel, F. Clément, and M. Sorine. A biomechanical model of muscle contraction. Lecture Notes in Computer Science, 2208:1159–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2208/22081159.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2208/22081159.
                                        pdf.

Lin:2001:FPS

[958] Wei te Lin and Richard A. Robb. A framework for patient-specific physicsbased simulation and interactive visualization of cardiac dynamics. Lecture Notes in Computer Science, 2208:1162– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22081162.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22081162.
pdf.

Nadkarni:2001:PCV

[959] Seemantini K. Nadkarni, Greg Mills, Derek R. Boughner, and Aaron Fenster. A pulsating coronary vessel phantom for two- and three-dimensional intravascular ultrasound studies. Lecture Notes in Computer Science, 2208:1164– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22081164.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22081164.
pdf.

Shute:2001:RAA

[960] Cameron A. Shute and Antony J. Hodgson. Repeatability and accuracy of ankle centre location estimates using a biaxial joint model. Lecture Notes in Computer Science, 2208:1166– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22081166.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22081166.
pdf.

Qatarneh:2001:WBA

[961] Sharif M. Qatarneh, Simo Hyödynmaa, Marilyn E. Noz, Gerald Q. Maguire Jr., Elissa L. Kramer, and Joakim Crafoord. A whole body atlas based segmentation for delineation of organs in radiation therapy planning. Lecture Notes in Computer Science, 2208:1168– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22081168.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22081168.
pdf.

Chen:2001:AMA

[962] Xuesong Chen, Chee-Kong Chui, Swee-Hin Teoh, Sim-Heng Ong, and Wieslaw L. Nowinski. Automatic modeling of anatomical structures for biomechanical analysis and visualization in a virtual spine workstation. Lecture Notes in Computer Science, 2208:1170– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22081170.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22081170.
pdf.

vanEngeland:2001:MBL

[963] Saskia van Engeland and Nico Karssemeijer. Matching breast lesions in multiple mammographic views. Lecture Notes in Computer Science, 2208:1172– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22081172.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2208/22081172.
pdf.

Rodriguez-Carranza:2001:ADL

[964] C. E. Rodríguez-Carranza and M. H. Loew. Automatic detection of large misregistrations of multimodality medical images. Lecture Notes in Computer Science, 2208:1174–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22081174.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22081174.
pdf.

Trobaugh:2001:RSU

[965] Jason W. Trobaugh and R. Martin Arthur. Registration of the spine using a physically-based image model for ultrasound. Lecture Notes in Computer Science, 2208:1176– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22081176.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22081176.
pdf.

Avants:2001:CMF

[966] Brian Avants, Marcelo Siqueira, and James Gee. Computing match functions for curves in IR^{2} and IR^{3} by refining polyline approximations. Lecture Notes in Computer Science, 2208:1178– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2208/22081178.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2208/22081178.
                                        pdf.

Harders:2001:MMA

[967] Matthias Harders and Gábor Székely. A multi-modal approach to segmentation of tubular structures. Lecture Notes in Computer Science, 2208:1180– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2208/22081180.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2208/22081180.
                                        pdf.

Barra:2001:SSN

[968] Vincent Barra, Jean-Jacques Lemaire, Franck Durif, and Jean-Yves Boire. Segmentation of the subthalamic nucleus in MR images using information fusion — A preliminary study for a computed-aided surgery of Parkinson’s disease. Lecture Notes in Computer Science, 2208:1183–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2208/22081183.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2208/22081183.
                                        pdf.

Bovenkamp:2001:CMA

[969] E. G. P. Bovenkamp, J. Dijkstra, J. G. Bosch, and J. H. C. Reiber. Collaborative multi-agent IVUS image segmentation. Lecture Notes in Computer Science, 2208:1185–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22081185.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22081185.
pdf.

Simonetti:2001:SCS

[970] A. W. Simonetti, R. Wehrens, and L. M. C. Buydens. Segmentation of chemical shift images with mixture modeling. Lecture Notes in Computer Science, 2208:1187–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22081187.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22081187.
pdf.

Duchesne:2001:ADF

[971] Simon Duchesne and D. Louis Collins. Analysis of 3D deformation fields for appearance-based segmentation. Lecture Notes in Computer Science, 2208:1189–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22081189.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22081189.
pdf.

Lievin:2001:MRF

[972] Marc Liévin, Nils Hanssen, Peter Zerfass, and Erwin Keeve. 3D Markov random fields and region growing for interactive segmentation of MR data. Lecture Notes in Computer Science, 2208:1191–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2208/22081191.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2208/22081191.
                                        pdf.

Yaniv:2001:LBP

[973] Z. Yaniv and L. Joskowicz. Long bone panoramas from fluoroscopic X-ray images. Lecture Notes in Computer Science, 2208:1193–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2208/22081193.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2208/22081193.
                                        pdf.

Worn:2001:ARO

[974] Heinz Wörn and Harald Hoppe. Augmented reality in the operating theatre of the future. Lecture Notes in Computer Science, 2208:1195–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2208/22081195.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2208/22081195.
                                        pdf.

Mizuta:2001:VVU

[975] Shinobu Mizuta, Ken ichi Kanda, and Tetsuya Matsuda. Volume visualization using gradient-based distance among voxels. Lecture Notes in Computer Science, 2208:1197–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22081197.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22081197.
pdf.

Tanaka:2001:ECB

[976] Daigo Tanaka, Alan Liu, and Christoph Kaufmann. The evaluation of the color blending function for the texture generation from photographs. Lecture Notes in Computer Science, 2208:1199– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22081199.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22081199.
pdf.

Kim:2001:AVR

[977] Tae-Young Kim, Byeong-Seok Shin, and Yeong Gil Shin. Anisotropic volume rendering using intensity interpolation. Lecture Notes in Computer Science, 2208:1201–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22081201.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22081201.
pdf.

Greenspan:2001:MIS

[978] Hayit Greenspan, Sharon Peled, Gal Oz, and Nahum Kiryati. MRI inter-slice reconstruction using superresolution. Lecture Notes in Computer Science, 2208:1204–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2208/22081204.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2208/22081204.pdf.

vanderGlas:2001:ACR

[979] Marjolein van der Glas, Frans Vos, and Albert Vossepoel. Assessment of center of rotation of the glenohumeral joint. Lecture Notes in Computer Science, 2208:1207–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2208/22081207.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2208/22081207.
                                        pdf.

Nobler:2001:PSE

[980] M. Nobler, S. Yu, B. Mensh, S. Lisanby, L. Alkalay, R. Van Heertum, E. Heyer, and H. Sackeim. PET studies of the effects of ECT on cerebral physiology. Lecture Notes in Computer Science, 2208:1210–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2208/22081210.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2208/22081210.
pdf.

Wallace:2001:FCM

[981] Jeffrey Wallace, Homayoun Mozaffari N., Li Pan, and Nitish V. Thakor. Fuzzy C-means clustering analysis to monitor tissue perfusion with near infrared imaging. Lecture Notes in Computer Science, 2208:1213–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22081213.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22081213.
pdf.

Cleynenbreugel:2001:MSB

[982] Johan Van Cleynenbreugel, Lars Bergmans, Martine Wevers, and Paul Lambrechts. A method for μCT based assessment of root canal instrumentation in endodontics research. Lecture Notes in Computer Science, 2208:1215– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22081215.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22081215.
pdf.

Kober:2001:RSB

[983] H. Kober, O. Ganslandt, Ch. Nimsky, M. Buchfelder, and R. Fahlbusch. Reconstruction of subcortical brain activity by spatially filtered MEG during epileptic seizures. Lecture Notes in Computer Science, 2208:1218–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2208/22081218.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2208/22081218.
                                        pdf.

Solanas:2001:EVC

[984] Eduard Solanas and Jean-Philippe Thiran. Exploiting voxel correlation for automated MRI bias field correction by conditional entropy minimization. Lecture Notes in Computer Science, 2208:1220–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2208/22081220.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2208/22081220.
                                        pdf.

Lyra:2001:SPE

[985] M. Lyra, K. Skouroliakou, C. Georgosopoulos, C. Stefanides, and J. Jordanou. Single photon emission computed tomography and 3 dimensional quantitative evaluation in renal scintigraphy. Lecture Notes in Computer Science, 2208:1222–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2208/22081222.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2208/22081222.
                                        pdf.

Gerard:2001:AAL

[986] O. Gerard, M. Fradkin, A. Collet Billon, M. Jacob, J.-M. Rouet, and S. Makram-Ebeid. Automatic analysis of the left ventricle in the time sequences of 3D echocardiographic images. Lecture Notes in Computer Science, 2208:1224–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22081224.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22081224.
pdf.

Dippel:2001:DFE

[987] Sabine Dippel. Displacement field estimation from CSPAMM images without segmentation of the myocardium. Lecture Notes in Computer Science, 2208:1226–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2208/22081226.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2208/22081226.
pdf.

Spreeuwers:2001:ADM

[988] Luuk Spreeuwers and Marcel Breeuwer. Automatic detection of myocardial boundaries in MR cardio perfusion images. Lecture Notes in Computer Science, 2208:1228–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22081228.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22081228.
pdf.

Rey:2001:USD

[989] David Rey, Jonathan Stoeckel, Grégoire Malandain, and Nicholas Ayache. Using SPM to detect evolving MS lesions. Lecture Notes in Computer Science, 2208:1232–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2208/22081232.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2208/22081232.
                                        pdf.

Welch:2001:SNE

[990] Edward Brian Welch, Armando Manduca, Roger C. Grimm, Heidi A. Ward, and Clifford R. Jack Jr. Spherical navigator echoes for full 3-D rigid body motion measurement in MRI. Lecture Notes in Computer Science, 2208:1235– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2208/22081235.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2208/22081235.
                                        pdf.

Suprijanto:2001:PUM

[991] Suprijanto, F. M. Vos, M. W. Vogel, A. M. Vossepoel, and H. M. Vrooman. Phase unwrapping for magnetic resonance thermometry. Lecture Notes in Computer Science, 2208:1237– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2208/22081237.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2208/22081237.
pdf.

Hartov:2001:SFV

[992] A. Hartov, H. Sun, D. W. Roberts, and K. D. Paulsen. A simplified field-of-view calibration technique for an operating microscope. Lecture Notes in Computer Science, 2208:1240– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22081240.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22081240.
pdf.

Holmes:2001:TUU

[993] David R. Holmes III, Brian J. Davis, and Richard A. Robb. Transurethral ultrasound: a new tool for diagnostic and therapeutic imaging in prostate cancer. Lecture Notes in Computer Science, 2208:1243–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22081243.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22081243.
pdf.

Geerling:2001:RRC

[994] J. Geerling, T. Hüfner, R. Raab, M. Citak, T. Pohlemann, H. Rosenthal, and C. Krettek. Resection of recurrent carcinomas in the pelvis — A new field of computer aided surgery. Lecture Notes in Computer Science, 2208:1245– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2208/22081245.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2208/22081245.
                                        pdf.

Hastreiter:2001:RAB

[995] P. Hastreiter, K. Engel, G. Soza, M. Bauer, M. Wolf, O. Ganslandt, R. Fahlbusch, G. Greiner, T. Ertl, and Ch. Nimsky. Remote analysis for brain shift compensation. Lecture Notes in Computer Science, 2208:1248– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2208/22081248.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2208/22081248.
                                        pdf.

Zeng:2001:DNI

[996] Jianchao Zeng, Ariela Sofer, and Seong K. Mun. Development of a new image-guided prostate biopsy system. Lecture Notes in Computer Science, 2208:1250–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2208/22081250.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2208/22081250.
                                        pdf.

Rosenthal:2001:ITA

[997] Michael Rosenthal, Susan Weeks, Stephen Aylward, Elizabeth Bullitt, and Henry Fuchs. Intraoperative tracking of anatomical structures using fluoroscopy and a vascular balloon catheter. Lecture Notes in Computer Science, 2208:1253–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22081253.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22081253.
pdf.

Kessler:2001:CUM

[998] Marc Kessler, Janelle Solock, Paul Archer, and Charles Meyer. Clinical use of a mutual informationbased automated image registration system for conformal radiotherapy treatment planning. Lecture Notes in Computer Science, 2208:1255–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22081255.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22081255.
pdf.

Schutyser:2001:VMS

[999] Filip Schutyser, Johan Van Cleynenbreugel, and Paul Suetens. Validation methods for soft tissue prediction in maxillofacial planning environments. Lecture Notes in Computer Science, 2208:1258–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22081258.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2208/22081258.
                                        pdf.

Lobregt:2001:DIP

[1000] Steven Lobregt, Ted Vuurberg, and Joost J. Schillings. Dental implant planning in EasyVision. Lecture Notes in Computer Science, 2208:1261–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2208/22081261.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2208/22081261.
                                        pdf.

Meier:2001:RTS

[1001] Ullrich Meier, Carlos Monserrat, Nils-Christian Parr, Francisco Javier García, and José Antonio Gil. Realtime simulation of minimally-invasive surgery with cutting based on boundary element methods. Lecture Notes in Computer Science, 2208:1263–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2208/22081263.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2208/22081263.
                                        pdf.

Chirani:2001:DRF

[1002] Reza Arbab Chirani, Jean-José Jacq, Christian Roux, and Philippe Meriot. 3-D reconstruction and functional analysis of the temporomandibular joint. Lecture Notes in Computer Science, 2208:1265–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2208/22081265.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2208/22081265.
pdf.

Azar:2001:MMP

[1003] Fred S. Azar, Dimitris N. Metaxas, and Mitchell D. Schnall. Methods for modeling and predicting mechanical deformations of the breast under external perturbations. Lecture Notes in Computer Science, 2208:1267– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22081267.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22081267.
pdf.

Schwartz:2001:NLS

[1004] Jean-Marc Schwartz, Ève Langelier, Christian Moisan, and Denis Laurendeau. Non-linear soft tissue deformations for the simulation of percutaneous surgeries. Lecture Notes in Computer Science, 2208:1271–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22081271.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22081271.
pdf.

Wachowiak:2001:CTS

[1005] Mark P. Wachowiak and Adel S. Elmaghraby. The continuous tabu search as an optimizer for 2D-to-3D biomedical image registration. Lecture Notes in Computer Science, 2208:1273–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2208/22081273.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2208/22081273.
                                        pdf.

Masutani:2001:MCF

[1006] Yoshitaka Masutani and Fumihiko Kimura. Modally controlled free form deformation for non-rigid registration in image-guided liver surgery. Lecture Notes in Computer Science, 2208:1275– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2208/22081275.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2208/22081275.
                                        pdf.

Anbeek:2001:TSR

[1007] P. Anbeek, K. L. Vincken, M. J. P. van Osch, J. P. W. Pluim, J. van der Grond, and M. A. Viergever. Twostep registration of subacute to hyperacute stroke MRIs. Lecture Notes in Computer Science, 2208:1279–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2208/22081279.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2208/22081279.
                                        pdf.

vanHerk:2001:EOM

[1008] M. van Herk, J. C. de Munck, M. T.

J. G. Groenewegen, A. R. Peters, and A. Touw. The effect of organ motion and image artifacts on monomodal volume registration. Lecture Notes in Computer Science, 2208:1281–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22081281.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22081281.
pdf.

West:2001:MLA

[1009] Jay B. West, Rasool Khadem, Shao Chin, and Ramin Shahidi. A maximum likelihood approach for 2-D to 3-D registration. Lecture Notes in Computer Science, 2208:1283–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22081283.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22081283.
pdf.

Cachier:2001:HTB

[1010] Pascal Cachier. How to trade off between regularization and image similarity in non-rigid registration ? Lecture Notes in Computer Science, 2208:1285– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22081285.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22081285.
pdf.

Behloul:2001:VER

[1011] F. Behloul, B. P. F. Lelieveldt, R. J. van der Geest, and J. H. C. Reiber. A virtual exploring robot for adaptive left ventricle contour detection in cardiac MR images. Lecture Notes in Computer Science, 2208:1287–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2208/22081287.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2208/22081287.
                                        pdf.

Jacq:2001:MBS

[1012] Jean-José Jacq and Christian Roux. A mesh-based shape analysis framework. Lecture Notes in Computer Science, 2208:1289–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2208/22081289.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2208/22081289.
                                        pdf.

Hahn:2001:CMB

[1013] Horst K. Hahn, Markus G. Lentschig, Burckhard Terwey, and Heinz-Otto Peitgen. Clinical MRI based volumetry: The cerebral ventricles. Lecture Notes in Computer Science, 2208:1291– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2208/22081291.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2208/22081291.
                                        pdf.

Pokric:2001:IPV

[1014] Maja Pokric, Neil Thacker, Marietta L. J. Scott, and Alan Jackson. The importance of partial voluming in multi-dimensional medical image segmentation. Lecture Notes in Computer Science, 2208:1293–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22081293.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22081293.
pdf.

Karangelis:2001:VDT

[1015] Grigoris Karangelis, Stelios Zimeras, Evelyn Firle, Min Wang, and Georgios Sakas. Volume definition tools for medical imaging applications. Lecture Notes in Computer Science, 2208:1295– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22081295.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22081295.
pdf.

Scahill:2001:AHS

[1016] Rachael I. Scahill, William R. Crum, and Nick C. Fox. Automated hippocampal segmentation by regional fluid registration of serial MRI: Validation and application in Alzheimer’s disease. Lecture Notes in Computer Science, 2208:1298–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2208/22081298.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2208/22081298.
                                        pdf.

Luo:2001:SSM

[1017] Hui Luo and Thomas O’Donnell. A 3D statistical shape model for the left ventricle of the heart. Lecture Notes in Computer Science, 2208:1300– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2208/22081300.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2208/22081300.
                                        pdf.

Boesecke:2001:RAD

[1018] R. Boesecke, J. Brief, J. Raczkowsky, O. Schorr, S. Daueber, R. Krempien, M. Treiber, T. Wetter, and S. Hassfeld. Robot assistant for dental implantology. Lecture Notes in Computer Science, 2208:1302–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2208/22081302.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2208/22081302.
                                        pdf.

Ruurda:2001:FLS

[1019] J. P. Ruurda, I. A. M. J. Broeders, R. K. J. Simmermacher, I. H. M. Borel Rinkes, and Th. J. M. V. van Vroonhoven. Feasibility of laparoscopic surgery assisted by a robotic telemanipulation system. Lecture Notes in Computer Science, 2208:1304–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22081304.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22081304.
pdf.

Krupa:2001:DSA

[1020] Alexandre Krupa, Michel de Mathelin, Christophe Doignon, Jacques Gangloff, Guillaume Morel, Luc Soler, and Jacques Marescaux. Development of semi-autonomous control modes in laparoscopic surgery using automatic visual servoing. Lecture Notes in Computer Science, 2208:1306–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22081306.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22081306.
pdf.

Engel:2001:SSR

[1021] Dirk Engel, Joerg Raczkowsky, and Heinz Woern. A supervising system for robot application in surgery. Lecture Notes in Computer Science, 2208:1308– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22081308.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22081308.
pdf.

Shi:2001:NDM

[1022] Xuemin Shi, Yi Guo, Shogo Fukushima, Kenshi Suzuki, Hisashi Uchiyama, and Ichiro Fukumoto. A new diagnostic method for Alzheimer’s disease using pupillary light reflex. Lecture Notes in Computer Science, 2208:1310– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22081310.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22081310.
pdf.

Yoshinaka:2001:RIR

[1023] Kiyoshi Yoshinaka, Ryo Sakamoto, and Ken Ikeuchi. Reduction of insertion resistance by means of vibration for catheters and endoscopes. Lecture Notes in Computer Science, 2208:1312– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22081312.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22081312.
pdf.

Sim:2001:HIP

[1024] Charlie Sim, Ming-Yeong Teo, Wan-Sing Ng, Carl Yap, Yong-Chong Loh, Tseng-Tsai Yeo, Shao-Ping Bai, and Charles Lo. Hexapod intervention planning for a robotic skullbase surgery system. Lecture Notes in Computer Science, 2208:1314–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22081314.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2208/22081314.
pdf.

Starkie:2001:AAC

[1025] Stephen J. Starkie and Brian L.

Davies. Advances in active constraints and their application to minimally invasive surgery. Lecture Notes in Computer Science, 2208:1316–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22081316.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22081316.
pdf.

Welch:2001:MCM

[1026] Edward Brian Welch and Armando Manduca. Motion correction of MRI from orthogonal k-space phase difference. Lecture Notes in Computer Science, 2208:1318–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2208/22081318.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2208/22081318.pdf.

Mourgues:2001:MCA

[1027] Fabien Mourgues, Frédéric Devernay, Grégoire Malandain, and Ève Coste-Manière. 3D+t modeling of coronary artery tree from standard non simultaneous angiograms. Lecture Notes in Computer Science, 2208:1320–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2208/22081320.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2208/22081320.
                                        pdf.

Mattes:2001:NTV

[1028] J. Mattes, J. Fieres, J. Beaudouin, D. Gerlich, J. Ellenberg, and R. Eils. New tools for visualization and quantification in dynamic processes: Application to the nuclear envelope dynamics during mitosis. Lecture Notes in Computer Science, 2208:1323–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2208/22081323.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2208/22081323.
                                        pdf.

Su:2001:FRB

[1029] Yi Su, Armando Manduca, Clifford R. Jack, E. Brian Welch, and Richard L. Ehman. Full 3D rigid body automatic motion correction of MRI images. Lecture Notes in Computer Science, 2208:1326–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2208/22081326.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2208/22081326.
                                        pdf.

Kakadiaris:2001:EML

[1030] I. A. Kakadiaris, A. Pednekar, G. Zouridakis, and K. Grigoriadis. Estimating the motion of the LAD: a simulation-based study. Lecture Notes in Computer Science, 2208:1328– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22081328.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22081328.
pdf.

Haber:2001:PDF

[1031] Idith Haber, Ron Kikinis, and Carl-Fredrik Westin. Phase-driven finite element model for spatio-temporal tracking in cardiac tagged MRI. Lecture Notes in Computer Science, 2208:1332– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22081332.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22081332.
pdf.

Keeve:2001:JES

[1032] Erwin Keeve, Thomas Jansen, Zdzislaw Krol, Lutz Ritter, Bartosz von Rymon-Lipinski, Robert Sader, Hans-Florian Zeilhofer, and Peter Zerfass. JULIUS — an extendable software framework for surgical planning and image-guided navigation. Lecture Notes in Computer Science, 2208:1336– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22081336.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22081336.
pdf.

Blackall:2001:SMR

[1033] J. M. Blackall, A. P. King, G. P. Penney, A. Adam, and D. J. Hawkes. A statistical model of respiratory motion and deformation of the liver. Lecture Notes in Computer Science, 2208:1338– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2208/22081338.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2208/22081338.
                                        pdf.

Teschner:2001:EST

[1034] Matthias Teschner, Sabine Girod, and Kevin Montgomery. Evaluation of soft-tissue model parameters. Lecture Notes in Computer Science, 2208:1341– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2208/22081341.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2208/22081341.
                                        pdf.

Stahl:2001:BPM

[1035] David J. Stahl, Jr. and Norberto Ezquerra. Bag-of-particles model for simulating tissue, organs, and flow. Lecture Notes in Computer Science, 2208:1343– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2208/22081343.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2208/22081343.
                                        pdf.

Bruyns:2001:SAD

[1036] Cynthia Bruyns, Simon Wildermuth, and Kevin Montgomery. Simulated animal dissection. Lecture Notes in Computer Science, 2208:1345–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22081345.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22081345.
pdf.

Wildermuth:2001:VP

[1037] Simon Wildermuth, Cynthia Bruyns, Kevin Montgomery, and Borut Marincek. Virtual polypectomy. Lecture Notes in Computer Science, 2208:1347– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22081347.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22081347.
pdf.

Bruyns:2001:RTI

[1038] Cynthia Bruyns, Steven Senger, Simon Wildermuth, Kevin Montgomery, and Richard Boyle. Real-time interactions using virtual tools. Lecture Notes in Computer Science, 2208:1349– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22081349.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22081349.
pdf.

Ferrant:2001:SBA

[1039] Matthieu Ferrant, Olivier Cuisenaire, Benoît Macq, Jean-Philippe Thiran, Martha E. Shenton, Ron Kikinis, and Simon K. Warfield. Surface based atlas matching of the brain using deformable surfaces and volumetric finite elements. Lecture Notes in Computer Science, 2208:1352–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22081352.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22081352.
pdf.

Dekker:2001:ECF

[1040] Niels Dekker, Lennert S. Ploeger, and Marcel van Herk. Evaluation of cost functions for gray value matching of 2D images in radiotherapy. Lecture Notes in Computer Science, 2208:1354– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22081354.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22081354.
pdf.

Behrenbruch:2001:MS

[1041] C. P. Behrenbruch, T. Kadir, and M. Brady. Mutual scale. Lecture Notes in Computer Science, 2208:1356– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22081356.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2208/22081356.
pdf.

Timoner:2001:FLE

[1042] Samson J. Timoner, W. Eric L. Grimson, Ron Kikinis, and William M. Wells III. Fast linear elastic matching without landmarks. Lecture Notes in Computer Science, 2208:1358–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22081358.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22081358.
pdf.

Khadem:2001:ECA

[1043] Rasool Khadem, Michael R. Bax, Jeremy A. Johnson, Eric P. Wilkinson, and Ramin Shahidi. Endoscope calibration and accuracy testing for 3D/2D image registration. Lecture Notes in Computer Science, 2208:1361– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22081361.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22081361.
pdf.

Vemuri:2001:FNR

[1044] B. C. Vemuri and J. Liu. Fast nonrigid multimodal image registration using local frequency maps. Lecture Notes in Computer Science, 2208:1363– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2208/22081363.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2208/22081363.
                                        pdf.

Nakajima:2001:AVC

[1045] Yoshikazu Nakajima, Toshihiko Sasama, Yoshinobu Sato, Takashi Nishii, Nobuhiko Sugano, Takashi Ishikawa, Kazuo Yonenobu, Takahiro Ochi, and Shinichi Tamura. Accuracy validation of cone-beam CT based registration. Lecture Notes in Computer Science, 2208:1366–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2208/22081366.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2208/22081366.
                                        pdf.

Ibanez:2001:MBD

[1046] Luis Ibáñez, Chafiaâ Hamitouche, Martial Boniou, and Christian Roux. Morphogenesis-based deformable models application to 3D medical image segmentation and analysis. Lecture Notes in Computer Science, 2208:1369– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2208/22081369.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2208/22081369.
                                        pdf.

Hult:2001:GLM

[1047] Roger Hult and Ewert Bengtsson. Grey-level morphology based segmentation of T1–MRI of the human cortex. Lecture Notes in Computer Science, 2208:1371–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22081371.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22081371.
pdf.

Gebhard:2001:ACM

[1048] Matthias Gebhard, Julian Mattes, and Roland Eils. An active contour model for segmentation based on cubic B-splines and gradient vector flow. Lecture Notes in Computer Science, 2208:1373–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22081373.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22081373.
pdf.

Fichtinger:2001:AVR

[1049] Gabor Fichtinger, Sheng Xu, Attila Tanacs, Kieran Murphy, Lee Myers, and Jeffery Williams. Approximate volumetric reconstruction from projected images. Lecture Notes in Computer Science, 2208:1376–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22081376.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22081376.
pdf.

Yoo:2001:ISA

[1050] T. S. Yoo and K. R. Subramanian. Implicit snakes: Active constrained implicit models. Lecture Notes in Computer Science, 2208:1379– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2208/22081379.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2208/22081379.
                                        pdf.

Nielsen:2001:TRP

[1051] Casper F. Nielsen and Peter J. Passmore. Towards a robust path growing algorithm for semi-automatic MRI segmentation. Lecture Notes in Computer Science, 2208:1382–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2208/22081382.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2208/22081382.
                                        pdf.

Reinhardt:2001:ISO

[1052] J. M. Reinhardt, J. Guo, L. Zhang, D. Bilgen, S. Hu, R. Uppaluri, R. M. Long, O. I. Saba, G. McLennan, M. Sonka, and E. A. Hoffman. Integrated system for objective assessment of global and regional lung structure. Lecture Notes in Computer Science, 2208:1384–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2208/22081384.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2208/22081384.
pdf.

Fukumoto:2001:CAD

[1053] Ichiro Fukumoto. Computer aided diagnosis for the alzheimer type dementia. Lecture Notes in Computer Science, 2208:1386–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2208/22081386.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2208/22081386.pdf.

Kubo:2001:CSA

[1054] M. Kubo, T. Yamamoto, Y. Kawata, N. Niki, K. Eguchi, H. Ohmatsu, R. Kakinuma, M. Kaneko, M. Kusumoto, N. Moriyama, K. Mori, and H. Nishiyama. CAD system for the assistance of comparative reading for lung cancer using serial helical CT images. Lecture Notes in Computer Science, 2208:1388– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22081388.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22081388.
pdf.

Varela:2001:CBT

[1055] Celia Varela, Nico Karssemeijer, and Pablo G. Tahoces. Classification of breast tumors on digital mammograms using Laws’ texture features. Lecture Notes in Computer Science, 2208:1391– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22081391.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22081391.
pdf.

Kawata:2001:CAD

[1056] Y. Kawata, N. Niki, H. Ohmatsu, M. Kusumoto, R. Kakinuma, K. Mori, H. Nishiyama, K. Eguchi, M. Kaneko, and N. Moriyama. Computeraided diagnosis of pulmonary nodules using three-dimensional thoracic CT images. Lecture Notes in Computer Science, 2208:1393–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22081393.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22081393.
pdf.

Bogoni:2001:OSB

[1057] Luca Bogoni, Jane C. Asmuth, David Hirvonen, Bojidar D. Madjarov, and Jeffrey W. Berger. Ophthalmic slitlamp-based computeraided diagnosis: Image processing foundations. Lecture Notes in Computer Science, 2208:1395–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22081395.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22081395.
pdf.

Jaume:2001:MSP

[1058] Sylvain Jaume, Matthieu Ferrant, Andreas Schreyer, Lennox Hoyte, Benoît Macq, Julia Fielding, Ron Kikinis, and Simon K. Warfield. Multiresolution signal processing on meshes for automatic pathological shape characterization. Lecture Notes in Computer Science, 2208:1398–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22081398.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22081398.
pdf.

Bhalerao:2001:IVM

[1059] Abhir Bhalerao and Paul Summers.

Interactive visualisation of MRI vector and tensor fields. Lecture Notes in Computer Science, 2208:1401–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22081401.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22081401.
pdf.

Pandya:2001:TMM

[1060] Abhilash Pandya, Mohamad Siadat, Lucia Zamorano, Jainxing Gong, Qinghang Li, James Maida, and Ioannis Kakadiaris. Tracking methods for medical augmented reality. Lecture Notes in Computer Science, 2208:1404– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2208/22081404.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2208/22081404.
                                        pdf.

Zhang:2001:TAV

[1061] S. Zhang, Ç. Demiralp, M. DaSilva, D. Keefe, D. Laidlaw, B. D. Greenberg, P. J. Basser, C. Pierpaoli, E. A. Chiocca, and T. S. Deisboeck. Toward application of virtual reality to visualization of DT–MRI volumes. Lecture Notes in Computer Science, 2208:1406– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2208/22081406.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2208/22081406.
                                        pdf.

Cuisenaire:2001:RVT

[1062] O. Cuisenaire, M. Ferrant, Y. Vandermeeren, E. Olivier, and B. Macq. Registration and visualization of transcranial magnetic stimulation on MR images. Lecture Notes in Computer Science, 2208:1409–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2208/22081409.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2208/22081409.
                                        pdf.

Koppel:2001:AIR

[1063] Dan Koppel, Yuan-Fang Wang, and Hua Lee. Automated image rectification in video-endoscopy. Lecture Notes in Computer Science, 2208:1412– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22081412.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22081412.
pdf.

Cocosco:2001:JIV

[1064] Chris A. Cocosco and Alan C. Evans.

Java Internet viewer: a WWW tool for remote 3D medical image data visualization and comparison. Lecture Notes in Computer Science, 2208:1415– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22081415.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22081415.
pdf.

Kumar:2001:ATL

[1065] Rajesh Kumar, Aaron C. Barnes, Gregory D. Hager, Patrick S. Jensen, and Russell H. Taylor. Applications of tasklevel augmentation for cooperative fine manipulation tasks in surgery. Lecture Notes in Computer Science, 2208:1417– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22081417.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22081417.
pdf.

Park:2001:VFR

[1066] Shinsuk Park, Robert D. Howe, and David F. Torchiana. Virtual fixtures for robotic cardiac surgery. Lecture Notes in Computer Science, 2208:1419– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22081419.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22081419.
pdf.

Benabid:2001:CUM

[1067] Alim-Louis Benabid, Dominque Hoffmann, Luc Court, Vincent Robert, Sébastien Burtin, Patrick Pittet, and Jörg Fischer. The clinical use of multi-modal resources (2D/3D/ statistics) for robot assisted functional neurosurgery. Lecture Notes in Computer Science, 2208:1421–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22081421.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22081421.
pdf.

Daeyoung:2001:ACP

[1068] Kim Daeyoung, Ryoichi Nakamura, Etsuko Kobayashi, Ichiro Sakuma, and Takeyoshi Dohi. Application of compact pneumatic actuators to laparoscopic manipulator. Lecture Notes in Computer Science, 2208:1424–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22081424.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2208/22081424.
pdf.

Berkelman:2001:PEC

[1069] Peter J. Berkelman, Daniel L. Rothbaum, Jaydeep Roy, Sam Lang, Louis L. Whitcomb, Greg Hager, Patrick S. Jensen, Eugene de Juan, Russell H. Taylor, and John K. Niparko. Performance evaluation of a cooperative manipulation microsurgical assistant robot applied to stapedotomy. Lecture Notes in Computer Science, 2208:1426–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22081426.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22081426.
pdf.

Goldberg:2001:MRS

[1070] Randal P. Goldberg, Mazilu Dumitru, Russell H. Taylor, and Dan Stoianovici. A modular robotic system for ultrasound image acquisition. Lecture Notes in Computer Science, 2208:1430– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2208/22081430.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2208/22081430.
pdf.

Kerschbaumer:2001:MSI

[1071] Fridun Kerschbaumer, Juergen Wahrburg, and Stephan Kuenzler. A mechatronic system for the implantation of the acetabular component in total hip alloarthroplasty. Lecture Notes in Computer Science, 2208:1433–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2208/22081433.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2208/22081433.
                                        pdf.

Selha:2001:OPP

[1072] Shaun Selha, Pierre Dupont, Robert Howe, and David Torchiana. Optimal port placement in robot-assisted coronary artery bypass grafting. Lecture Notes in Computer Science, 2208:1435– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2208/22081435.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2208/22081435.
                                        pdf.

Anonymous:2001:AIu

[1073] Anonymous. Author index. Lecture Notes in Computer Science, 2208:1437– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2208/2208indx.pdf.

Audestad:2001:TDE

[1074] Jan A. Audestad. Telecommunications, databases, and evolution. Lecture Notes in Computer Science, 2209: 1–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2209/22090001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2209/22090001.
pdf.

Adzic:2001:DWP

[1075] Jovanka Adzic, Valter Fiore, and Stefano Spelta. Data warehouse population platform. Lecture Notes in Computer Science, 2209:9–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2209/22090009.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2209/22090009.
pdf.

Wijnands:2001:ENC

[1076] J. E. P. Wijnands, S. J. Dijkstra, W. L. A. Derks, and W. Jonker. Experimenting NT cluster technology for massive CDR processing. Lecture Notes in Computer Science, 2209:19– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2209/22090019.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2209/22090019.
pdf.

Wu:2001:MSA

[1077] Pei-Hsin Wu, Wen-Chih Peng, and Ming-Syan Chen. Mining sequential alarm patterns in a telecommunication database. Lecture Notes in Computer Science, 2209:37–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2209/22090037.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2209/22090037.
                                        pdf.

Akinde:2001:GMJ

[1078] Michael O. Akinde and Michael H. Böhlen. Generalized MD-Joins: Evaluation and reduction to SQL. Lecture Notes in Computer Science, 2209:52– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2209/22090052.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2209/22090052.
                                        pdf.

Toman:2001:QPE

[1079] David Toman and Grant Weddell. Query processing in embedded control programs. Lecture Notes in Computer Science, 2209:68–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2209/22090068.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2209/22090068.
                                        pdf.

Lindstrom:2001:BRT

[1080] Jan Lindström and Tiina Niklander. Benchmark for real-time database systems for telecommunications. Lecture Notes in Computer Science, 2209:88– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2209/22090088.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2209/22090088.
pdf.

Bjornerstedt:2001:RBG

[1081] Anders Björnerstedt, Helena Ketoja, Johan Sintorn, and Martin Sköld. Replication between geographically separated clusters — an asynchronous scalable replication mechanism for very high availability. Lecture Notes in Computer Science, 2209:102–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2209/22090102.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2209/22090102.
pdf.

Zimmermann:2001:YDE

[1082] Roger Zimmermann, Kun Fu, Cyrus Shahabi, Didi Shu-Yuen Yao, and Hong Zhu. Yima: Design and evaluation of a streaming media system for residential broadband services. Lecture Notes in Computer Science, 2209:116– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2209/22090116.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2209/22090116.
pdf.

Georgalas:2001:QQB

[1083] Nektarios Georgalas. QuDAS: a QoS-based brokering architecture for data services. Lecture Notes in Computer Science, 2209:126–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2209/22090126.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2209/22090126.
                                        pdf.

Delot:2001:LDD

[1084] Thierry Delot, Pascal Déchamboux, Béatrice Finance, Yann Lepetit, and Gilles LeBrun. LDAP, databases and distributed objects: Towards a better integration. Lecture Notes in Computer Science, 2209:140–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2209/22090140.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2209/22090140.
                                        pdf.

Cochinwala:2001:NCU

[1085] Munir Cochinwala, Harald Hauser, and Naveen Suri. Network convergence using universal numbers: The UPT project. Lecture Notes in Computer Science, 2209:155–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2209/22090155.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2209/22090155.
                                        pdf.

Martin-Flatin:2001:TUI

[1086] Jean-Philippe Martin-Flatin. Toward universal information models in enterprise management. Lecture Notes in Computer Science, 2209:167–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2209/22090167.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2209/22090167.
pdf.

Anonymous:2001:AIv

[1087] Anonymous. Author index. Lecture Notes in Computer Science, 2209:179– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2209/2209indx.pdf.

Torresen:2001:TSI

[1088] J. Torresen. Two-step incremental evolution of a prosthetic hand controller based on digital logic gates. Lecture Notes in Computer Science, 2210:1– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2210/22100001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2210/22100001.
pdf.

Miller:2001:UEE

[1089] J. F. Miller and M. Hartmann. Untidy evolution: Evolving messy gates for fault tolerance. Lecture Notes in Computer Science, 2210:14–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2210/22100014.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2210/22100014.
                                        pdf.

Schnier:2001:EDC

[1090] T. Schnier and X. Yao. Evolutionary design calibration. Lecture Notes in Computer Science, 2210:26– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2210/22100026.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2210/22100026.
                                        pdf.

Iwata:2001:IGL

[1091] M. Iwata, I. Kajitani, Y. Liu, N. Kajihara, and T. Higuchi. Implementation of a gate-level evolvable hardware chip. Lecture Notes in Computer Science, 2210:38–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2210/22100038.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2210/22100038.
                                        pdf.

Schemmel:2001:VIA

[1092] J. Schemmel, K. Meier, and F. Schürmann. A VLSI implementation of an analog neural network suited for genetic algorithms. Lecture Notes in Computer Science, 2210:50–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2210/22100050.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2210/22100050.
pdf.

Langeheine:2001:ISN

[1093] J. Langeheine, J. Becker, S. Fölling, K. Meier, and J. Schemmel. Initial studies of a new VLSI field programmable transistor array. Lecture Notes in Computer Science, 2210:62– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2210/22100062.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2210/22100062.
pdf.

Restrepo:2001:EIS

[1094] H. F. Restrepo and D. Mange. An embryonics implementation of a selfreplicating universal Turing machine. Lecture Notes in Computer Science, 2210:74–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2210/22100074.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2210/22100074.
pdf.

Jackson:2001:AER

[1095] A. H. Jackson and A. M. Tyrrell. Asynchronous embryonics with reconfiguration. Lecture Notes in Computer Science, 2210:88–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2210/22100088.

                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2210/22100088.pdf.

Prodan:2001:EAC

[1096] L. Prodan, G. Tempesti, D. Mange, and A. Stauffer. Embryonics: Artificial cells driven by artificial DNA. Lecture Notes in Computer Science, 2210: 100–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2210/22100100.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2210/22100100.
                                        pdf.

Stauffer:2001:SRS

[1097] A. Stauffer, D. Mange, G. Tempesti, and C. Teuscher. A selfrepairing and self-healing electronic watch: The BioWatch. Lecture Notes in Computer Science, 2210:112–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2210/22100112.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2210/22100112.
                                        pdf.

Haddow:2001:SGS

[1098] P. C. Haddow, G. Tufte, and P. van Remortel. Shrinking the genotype: L-systems for EHW? Lecture Notes in Computer Science, 2210:128–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2210/22100128.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2210/22100128.
pdf.

Bradley:2001:MLD

[1099] D. W. Bradley and A. M. Tyrrell.

Multi-layered defence mechanisms: Architecture, implementation and demonstration of a hardware immune system. Lecture Notes in Computer Science, 2210:140–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2210/22100140.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2210/22100140.
pdf.

Tamplin:2001:ACW

[1100] M. R. Tamplin and A. Hamilton.

Ant circuit world: An ant algorithm MATLAB^{T M} toolbox for the design, visualisation and analysis of analogue circuits. Lecture Notes in Computer Science, 2210:151–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2210/22100151.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2210/22100151.
pdf.

Kim:2001:HLD

[1101] J. G. Kim, K g. Noh, and K. Park.

Human-like dynamic walking for a biped robot using genetic algorithm. Lecture Notes in Computer Science, 2210:159–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2210/22100159.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2210/22100159.
                                        pdf.

Islam:2001:EFE

[1102] Md. Monirul Islam, S. Terao, and K. Murase. Effect of fitness for the evolution of autonomous robots in an open-environment. Lecture Notes in Computer Science, 2210:171–181, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2210/22100171.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2210/22100171.
                                        pdf.

Islam:2001:IEA

[1103] Md. Monirul Islam, S. Terao, and K. Murase. Incremental evolution of autonomous robots for a complex task. Lecture Notes in Computer Science, 2210:182–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2210/22100182.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2210/22100182.
                                        pdf.

Yanai:2001:MAR

[1104] K. Yanai and H. Iba. Multi-agent robot learning by means of genetic programming: Solving an escape problem. Lecture Notes in Computer Science, 2210: 192–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2210/22100192.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2210/22100192.
pdf.

Fernandez:2001:PRC

[1105] F. Fernández, J. M. Sánchez, and M. Tomassini. Placing and routing circuits on FPGAs by means of parallel and distributed genetic programming. Lecture Notes in Computer Science, 2210:204–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2210/22100204.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2210/22100204.
pdf.

Eklund:2001:MPA

[1106] S. E. Eklund. A massively parallel architecture for linear machine code genetic programming. Lecture Notes in Computer Science, 2210:216– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2210/22100216.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2210/22100216.
pdf.

Korkin:2001:SOE

[1107] M. Korkin. Self-organized evolutionary process in sets of interdependent variables near the midpoint of phase transition in K-satisfiability. Lecture Notes in Computer Science, 2210:225– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2210/22100225.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2210/22100225.
                                        pdf.

Lohn:2001:EOY

[1108] J. D. Lohn, W. F. Kraus, D. S. Linden, and S. P. Colombano. Evolutionary optimization of Yagi-Uda antennas. Lecture Notes in Computer Science, 2210: 236–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2210/22100236.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2210/22100236.
                                        pdf.

Perez:2001:EDP

[1109] E. Islas Pérez, C. A. Coello Coello, and A. Hernández Aguirre. Extraction of design patterns from evolutionary algorithms using case-based reasoning. Lecture Notes in Computer Science, 2210: 244–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2210/22100244.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2210/22100244.
                                        pdf.

Leung:2001:BSC

[1110] K. S. Leung, K. H. Lee, and S. M. Cheang. Balancing samples’ contributions on GA learning. Lecture Notes in Computer Science, 2210:256– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2210/22100256.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2210/22100256.
pdf.

Sanchez:2001:SPO

[1111] Eduardo Sanchez, Andrés Pérez-Uribe, and Bertrand Mesot. Solving partially observable problems by evolution and learning of finite state machines. Lecture Notes in Computer Science, 2210: 267–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2210/22100267.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2210/22100267.
pdf.

Aguirre:2001:GBL

[1112] A. Hernández Aguirre, B. P. Buckles, and C. Coello Coello. GA-based learning of kDNF_{n}^{s} Boolean formulas. Lecture Notes in Computer Science, 2210: 279–290, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2210/22100279.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2210/22100279.
pdf.

Stoica:2001:PE

[1113] A. Stoica, R. Zebulum, and D. Keymeulen.

Polymorphic electronics. Lecture Notes in Computer Science, 2210:291–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2210/22100291.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2210/22100291.
                                        pdf.

Keymeulen:2001:IER

[1114] D. Keymeulen, R. Zebulum, A. Stoica, and M. Buehler. Initial experiments of reconfigurable sensor adapted by evolution. Lecture Notes in Computer Science, 2210:303–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2210/22100303.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2210/22100303.
                                        pdf.

Sakanashi:2001:LCM

[1115] H. Sakanashi, M. Iwata, and T. Higuchi. A lossless compression method for halftone images using evolvable hardware. Lecture Notes in Computer Science, 2210:314–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2210/22100314.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2210/22100314.pdf.

Nosato:2001:EOS

[1116] H. Nosato, Y. Kasai, T. Itatani, M. Murakawa, T. Furuya, and T. Higuchi. Evolvable optical systems and their applications. Lecture Notes in Computer Science, 2210:327– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2210/22100327.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2210/22100327.
pdf.

Anonymous:2001:AIw

[1117] Anonymous. Author index. Lecture Notes in Computer Science, 2210:341– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2210/2210auth.pdf.

Alexander:2001:HMS

[1118] Perry Alexander and Cindy Kong.

Heterogeneous modeling support for embedded systems design. Lecture Notes in Computer Science, 2211:1– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2211/22110001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2211/22110001.
pdf.

Alur:2001:HHM

[1119] R. Alur, T. Dang, J. Esposito, R. Fierro, Y. Hur, F. Ivančić, V. Kumar, I. Lee, P. Mishra, G. Pappas, and O. Sokolsky. Hierarchical hybrid modeling of embedded systems. Lecture Notes in Computer Science, 2211:14– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2211/22110014.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2211/22110014.
                                        pdf.

Benveniste:2001:SSI

[1120] Albert Benveniste. Some synchronization issues when designing embedded systems from components. Lecture Notes in Computer Science, 2211:32– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2211/22110032.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2211/22110032.
                                        pdf.

Berry:2001:SPT

[1121] Gérard Berry. Synchronous programming techniques for embedded systems: Present and future. Lecture Notes in Computer Science, 2211:50– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2211/22110050.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2211/22110050.
                                        pdf.

Broy:2001:RVE

[1122] Manfred Broy and Oscar Slotosch. From requirements to validated embedded systems. Lecture Notes in Computer Science, 2211:51–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2211/22110051.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2211/22110051.
pdf.

Butts:2001:USA

[1123] Ken Butts, Dave Bostic, Alongkrit Chutinan, Jeffrey Cook, Bill Milam, and Yanxin Wang. Usage scenarios for an automated model compiler. Lecture Notes in Computer Science, 2211:66– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2211/22110066.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2211/22110066.
pdf.

Caspi:2001:ECA

[1124] Paul Caspi. Embedded control: From asynchrony to synchrony and back. Lecture Notes in Computer Science, 2211:80–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2211/22110080.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2211/22110080.
pdf.

Cousot:2001:VES

[1125] Patrick Cousot and Radhia Cousot.

Verification of embedded software: Problems and perspectives. Lecture Notes in Computer Science, 2211:97– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2211/22110097.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2211/22110097.
pdf.

Culler:2001:NCA

[1126] David E. Culler, Jason Hill, Philip Buonadonna, Robert Szewczyk, and Alec Woo. A network-centric approach to embedded software for tiny devices. Lecture Notes in Computer Science, 2211:114–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2211/22110114.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2211/22110114.
pdf.

Donahue:2001:SAR

[1127] Steven M. Donahue, Matthew P. Hampton, Morgan Deters, Jonathan M. Nye, Ron K. Cytron, and Krishna M. Kavi. Storage allocation for realtime, embedded systems. Lecture Notes in Computer Science, 2211:131– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2211/22110131.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2211/22110131.
pdf.

deAlfaro:2001:ITC

[1128] Luca de Alfaro and Thomas A. Henzinger. Interface theories for component-based design. Lecture Notes in Computer Science, 2211:148–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2211/22110148.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2211/22110148.
pdf.

Henzinger:2001:GTT

[1129] Thomas A. Henzinger, Benjamin Horowitz, and Christoph Meyer Kirsch. Giotto: a time-triggered language for embedded programming. Lecture Notes in Computer Science, 2211:166– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2211/22110166.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2211/22110166.
pdf.

Taha:2001:DFP

[1130] Walid Taha, Paul Hudak, and Zhanyong Wan. Directions in functional programming for real(- time) applications. Lecture Notes in Computer Science, 2211:185–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2211/22110185.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2211/22110185.
pdf.

Jeffay:2001:RBR

[1131] Kevin Jeffay and Steve Goddard.

Rate-based resource allocation models for embedded systems. Lecture Notes in Computer Science, 2211:204– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2211/22110204.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2211/22110204.
                                        pdf.

Kopetz:2001:TSI

[1132] H. Kopetz. The temporal specification of interfaces in distributed real-time systems. Lecture Notes in Computer Science, 2211:223–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2211/22110223.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2211/22110223.
                                        pdf.

Lee:2001:SLT

[1133] Edward A. Lee and Yuhong Xiong. System-level types for componentbased design. Lecture Notes in Computer Science, 2211:237–??, 2001. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer-ny.com/link/
                                        service/series/0558/bibs/2211/22110237.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2211/22110237.pdf.

Malik:2001:ESI

[1134] Sharad Malik. Embedded software implementation tools for fully programmable application specific systems. Lecture Notes in Computer Science, 2211:254–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/

series/0558/bibs/2211/22110254.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2211/22110254.pdf.

Palem:2001:COA

[1135] Krishna V. Palem, Surendranath Talla, and Weng-Fai Wong. Compiler optimizations for adaptive EPIC processors. Lecture Notes in Computer Science, 2211:257–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2211/22110257.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2211/22110257.pdf.

Pree:2001:ESM

[1136] Wolfgang Pree and Alessandro Pasetti.

Embedded software market transformation through reusable frameworks. Lecture Notes in Computer Science, 2211:274–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2211/22110274.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2211/22110274.
pdf.

Rajkumar:2001:EEM

[1137] Ragunathan (Raj) Rajkumar. An end-to-end methodology for building embedded systems. Lecture Notes in Computer Science, 2211:287–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2211/22110287.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2211/22110287.
                                        pdf.

Beebee:2001:ISM

[1138] William S. Beebee, Jr. and Martin Rinard. An implementation of scoped memory for real-time Java. Lecture Notes in Computer Science, 2211:289– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2211/22110289.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2211/22110289.
                                        pdf.

Rushby:2001:BAS

[1139] John Rushby. Bus architectures for safety-critical embedded systems. Lecture Notes in Computer Science, 2211: 306–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2211/22110306.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2211/22110306.
                                        pdf.

Burch:2001:UML

[1140] Jerry R. Burch, Roberto Passerone, and Alberto L. Sangiovanni-Vincentelli. Using multiple levels of abstractions in embedded software design. Lecture Notes in Computer Science, 2211:324– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2211/22110324.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2211/22110324.
pdf.

Koo:2001:HAD

[1141] T. John Koo, Judy Liebman, Cedric Ma, and S. Shankar Sastry. Hierarchical approach for design of multi-vehicle multi-modal embedded software. Lecture Notes in Computer Science, 2211: 344–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2211/22110344.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2211/22110344.
pdf.

Schmidt:2001:ARM

[1142] Douglas C. Schmidt. Adaptive and reflective middleware for distributed realtime and embedded systems. Lecture Notes in Computer Science, 2211:361– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2211/22110361.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2211/22110361.
pdf.

Sifakis:2001:MRT

[1143] Joseph Sifakis. Modeling real-time systems-challenges and work directions. Lecture Notes in Computer Science, 2211:373–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2211/22110373.
htm;   http://link.springer-ny.

com/link/service/series/0558/papers/
2211/22110373.pdf.

Stankovic:2001:VTC

[1144] John A. Stankovic. VEST — A toolset for constructing and analyzing component based embedded systems. Lecture Notes in Computer Science, 2211:390– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2211/22110390.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2211/22110390.
pdf.

Sztipanovits:2001:ESC

[1145] Janos Sztipanovits and Gabor Karsai. Embedded software: Challenges and opportunities. Lecture Notes in Computer Science, 2211:403–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2211/22110403.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2211/22110403.
pdf.

Thiele:2001:ESN

[1146] Lothar Thiele, Samarjit Chakraborty, Matthias Gries, Alexander Maxiaguine, and Jonas Greutert. Embedded software in network processors — models and algorithms. Lecture Notes in Computer Science, 2211:416– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2211/22110416.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2211/22110416.
pdf.

Simsek:2001:DAD

[1147] Tunc Simsek and Pravin Varaiya. Design of autonomous, distributed systems. Lecture Notes in Computer Science, 2211:435–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2211/22110435.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2211/22110435.pdf.

Binns:2001:FSA

[1148] Pam Binns and Steve Vestal. Formalizing software architectures for embedded systems. Lecture Notes in Computer Science, 2211:451–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2211/22110451.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2211/22110451.
pdf.

Ferdinand:2001:RPW

[1149] Christian Ferdinand, Reinhold Heckmann, Marc Langenbach, Florian Martin, Michael Schmidt, Henrik Theiling, Stephan Thesing, and Reinhard Wilhelm. Reliable and precise WCET determination for a real-life processor. Lecture Notes in Computer Science, 2211:469–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2211/22110469.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2211/22110469.
                                        pdf.

Wirth:2001:ESR

[1150] Niklaus Wirth. Embedded systems and real-time programming. Lecture Notes in Computer Science, 2211:486– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2211/22110486.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2211/22110486.
                                        pdf.

Wolf:2001:ESV

[1151] Wayne Wolf. Embedded software for video. Lecture Notes in Computer Science, 2211:493–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2211/22110493.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2211/22110493.
                                        pdf.

Anonymous:2001:AIx

[1152] Anonymous. Author index. Lecture Notes in Computer Science, 2211:503– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2211/2211auth.pdf.

Pouzol:2001:DSM

[1153] Jean-Philippe Pouzol and Mireille Ducassé. From declarative signatures to misuse IDS. Lecture Notes in Computer Science, 2212:1– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2212/22120001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2212/22120001.
pdf.

Almgren:2001:AID

[1154] Magnus Almgren and Ulf Lindqvist.

Application-integrated data collection for security monitoring. Lecture Notes in Computer Science, 2212:22– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2212/22120022.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2212/22120022.
pdf.

Welz:2001:ITA

[1155] Marc Welz and Andrew Hutchison.

Interfacing trusted applications with intrusion detection systems. Lecture Notes in Computer Science, 2212:37– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2212/22120037.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2212/22120037.
pdf.

Valdes:2001:PAC

[1156] Alfonso Valdes and Keith Skinner. Probabilistic alert correlation. Lecture Notes in Computer Science, 2212:54– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2212/22120054.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2212/22120054.
                                        pdf.

Vigna:2001:DWH

[1157] Giovanni Vigna, Richard A. Kemmerer, and Per Blix. Designing a Web of highly-configurable intrusion detection sensors. Lecture Notes in Computer Science, 2212:69–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2212/22120069.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2212/22120069.
                                        pdf.

Debar:2001:ACI

[1158] Hervé Debar and Andreas Wespi. Aggregation and correlation of intrusiondetection alerts. Lecture Notes in Computer Science, 2212:85–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2212/22120085.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2212/22120085.
                                        pdf.

Cunningham:2001:ADS

[1159] Robert K. Cunningham and Craig S.

Stevenson. Accurately detecting source code of attacks that increase privilege. Lecture Notes in Computer Science, 2212:104–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2212/22120104.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2212/22120104.
pdf.

Williams:2001:CTC

[1160] Paul D. Williams, Kevin P. Anchor, John L. Bebo, Gregg H. Gunsch, and Gary D. Lamont. CDIS: Towards a computer immune system for detecting network intrusions. Lecture Notes in Computer Science, 2212:117– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2212/22120117.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2212/22120117.
pdf.

Sterne:2001:ARD

[1161] Dan Sterne, Kelly Djahandari, Brett Wilson, Bill Babson, Dan Schnackenberg, Harley Holliday, and Travis Reid. Autonomic response to distributed denial of service attacks. Lecture Notes in Computer Science, 2212: 134–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2212/22120134.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2212/22120134.
                                        pdf.

Johnston:2001:IPD

[1162] Steven R. Johnston. The impact of privacy and data protection legislation on the sharing of intrusion detection information. Lecture Notes in Computer Science, 2212:150–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2212/22120150.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2212/22120150.
                                        pdf.

Uppuluri:2001:ESB

[1163] Prem Uppuluri and R. Sekar. Experiences with specification-based intrusion detection. Lecture Notes in Computer Science, 2212:172–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2212/22120172.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2212/22120172.
                                        pdf.

Ko:2001:SHI

[1164] Calvin Ko, Paul Brutch, Jeff Rowe, Guy Tsafnat, and Karl Levitt. System health and intrusion monitoring using a hierarchy of constraints. Lecture Notes in Computer Science, 2212: 190–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2212/22120190.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2212/22120190.
pdf.

Anonymous:2001:AIy

[1165] Anonymous. Author index. Lecture Notes in Computer Science, 2212:205– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2212/2212auth.pdf.

Pierre:2001:GPS

[1166] Guillaume Pierre and Maarten van Steen. Globule: a platform for selfreplicating Web documents. Lecture Notes in Computer Science, 2213:1– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2213/22130001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2213/22130001.
pdf.

Tommasi:2001:AQQ

[1167] Franco Tommasi, Simone Molendini, Antonio Vilei, and Andrea Sponziello. Architecture of QOSMIC — A QoS manager for Internet connections. Lecture Notes in Computer Science, 2213: 12–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2213/22130012.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2213/22130012.
pdf.

Bagnulo:2001:QDI

[1168] Marcelo Bagnulo, Alberto Garcia-Martinez, David Larrabeiti, and Arturo Azcorra. A QoS-driven ISP selection mechanism for IPv6 multi-homed sites. Lecture Notes in Computer Science, 2213:23–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2213/22130023.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2213/22130023.
pdf.

Hesselman:2001:BMC

[1169] Cristian Hesselman and Henk Eertink. Broadcasting multimedia channels in future mobile systems. Lecture Notes in Computer Science, 2213:35–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2213/22130035.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2213/22130035.
pdf.

Kwon:2001:PVF

[1170] Jin B. Kwon and Heon Y. Yeom. Providing VCR functionality in staggered video broadcasting. Lecture Notes in Computer Science, 2213:44– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2213/22130044.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2213/22130044.
pdf.

Cores:2001:FSD

[1171] Fernando Cores, Ana Ripoll, and Emilio Luque. A fully scalable and distributed architecture for video-on-demand. Lecture Notes in Computer Science, 2213:59–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2213/22130059.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2213/22130059.
pdf.

Bouras:2001:MMM

[1172] Christos Bouras and A. Gkamas. A mechanism for multicast multimedia data with adaptive QoS characteristics. Lecture Notes in Computer Science, 2213:74–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2213/22130074.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2213/22130074.
pdf.

Lopes:2001:BMA

[1173] Rui J. Lopes, Laurent Mathy, and David Hutchison. Bandwidth measurements of ALM trees for content distribution. Lecture Notes in Computer Science, 2213:89–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2213/22130089.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2213/22130089.
pdf.

Diaz:2001:GNM

[1174] Michel Diaz, Roberto Canonico, Luis Costa, Serge Fdida, David Hutchison, Laurent Mathy, Andreas Meissner, Stephane Owezarski, Rolland Vida, and Lars Wolf. GCAP: a new multimedia multicast architecture for QoS. Lecture Notes in Computer Science, 2213: 103–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2213/22130103.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2213/22130103.
                                        pdf.

Konorski:2001:PFM

[1175] Jerzy Konorski. Protection of fairness for multimedia traffic streams in a noncooperative wireless LAN setting. Lecture Notes in Computer Science, 2213: 116–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2213/22130116.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2213/22130116.
                                        pdf.

Willekens:2001:AHR

[1176] Jeroen P. F. Willekens. Ad hoc routing in bluetooth. Lecture Notes in Computer Science, 2213:130–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2213/22130130.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2213/22130130.
                                        pdf.

Kormentzas:2001:OSP

[1177] George Kormentzas, D. Vergados, J. Soldatos, and G. Stassinopoulos. An open signaling protocol for wireless host mobility. Lecture Notes in Computer Science, 2213:145–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2213/22130145.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2213/22130145.
pdf.

Benz:2001:TIP

[1178] M. Benz, K. Feske, U. Hatnik, and P. Schwarz. TCP/IP protocol engine system simulation. Lecture Notes in Computer Science, 2213:155–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2213/22130155.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2213/22130155.
pdf.

Hui:2001:DIT

[1179] Albert K. T. Hui and Samuel T.

Chanson. Design and implementation of a transparent forward error correction coding daemon for unreliable transports. Lecture Notes in Computer Science, 2213:165–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2213/22130165.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2213/22130165.
                                        pdf.

Wu:2001:AST

[1180] Wei Wu, Yong Ren, and Xiuming Shan. An analytical study on TCP-friendly congestion control for multimedia streams. Lecture Notes in Computer Science, 2213:180–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2213/22130180.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2213/22130180.
                                        pdf.

Teeuw:2001:MBS

[1181] Wouter B. Teeuw and Dick A. C. Quartel. Model-based service creation in the friends project. Lecture Notes in Computer Science, 2213:192– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2213/22130192.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2213/22130192.
                                        pdf.

Wegdam:2001:VOS

[1182] Maarten Wegdam, Dirk-Jaap Plas, and Musa Unmehopa. Validation of the open service access API for UMTS application provisioning. Lecture Notes in Computer Science, 2213: 210–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2213/22130210.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2213/22130210.
pdf.

Widya:2001:EUQ

[1183] I. Widya, R. E. Stap, L. J. Teunissen, and B. F. Hopman. On the end-user QoS-awareness of a distributed service environment. Lecture Notes in Computer Science, 2213:222– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2213/22130222.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2213/22130222.
pdf.

Anonymous:2001:AIz

[1184] Anonymous. Author index. Lecture Notes in Computer Science, 2213:239– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2213/2213auth.pdf.

Amilhastre:2001:FMH

[1185] Jérôme Amilhastre, Philippe Janssen, and Marie-Catherine Vilarem. FA minimisation heuristics for a class of finite languages. Lecture Notes in Computer Science, 2214:1–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2214/22140001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2214/22140001.
pdf.

Andary:2001:SSE

[1186] Philippe Andary, Pascal Caron, Jean-Marc Champarnaud, Gérard Duchamp, Marianne Flouret, and Éric Laugerotte. SEA: a symbolic environment for automata theory. Lecture Notes in Computer Science, 2214:13– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2214/22140013.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2214/22140013.
pdf.

Bergeron:2001:ARS

[1187] Anne Bergeron and Riccardo Catalano. Analysis of reactive systems with n timers. Lecture Notes in Computer Science, 2214:27–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2214/22140027.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2214/22140027.
pdf.

Braune:2001:AGC

[1188] Beatrix Braune, Stephan Diehl, Andreas Kerren, and Reinhard Wilhelm. Animation of the generation and computation of finite automata for learning software. Lecture Notes in Computer Science, 2214:39– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2214/22140039.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2214/22140039.
pdf.

Calude:2001:MLA

[1189] Cristian S. Calude, Kai Salomaa, and Sheng Yu. Metric lexical analysis. Lecture Notes in Computer Science, 2214: 48–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2214/22140048.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2214/22140048.
pdf.

Campeanu:2001:SCB

[1190] C. Câmpeanu, K. Culik, Kai Salomaa, and Sheng Yu. State complexity of basic operations on finite languages. Lecture Notes in Computer Science, 2214: 60–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2214/22140060.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2214/22140060.
pdf.

Daciuk:2001:TUW

[1191] Jan Daciuk. Treatment of unknown words. Lecture Notes in Computer Science, 2214:71–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2214/22140071.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2214/22140071.pdf.

Eramian:2001:CEM

[1192] Mark G. Eramian. Computing entropy maps of finite-automaton-encoded binary images. Lecture Notes in Computer Science, 2214:81– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2214/22140081.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2214/22140081.
                                        pdf.

Giammarresi:2001:TDC

[1193] Dora Giammarresi, Jean-Luc Ponty, and Derick Wood. Thompson digraphs: a characterization. Lecture Notes in Computer Science, 2214:91– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2214/22140091.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2214/22140091.
                                        pdf.

Jurgensen:2001:FAE

[1194] Helmut Jürgensen and Ludwig Staiger. Finite automata encoding geometric figures. Lecture Notes in Computer Science, 2214:101–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2214/22140101.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2214/22140101.pdf.

Kiraz:2001:CSS

[1195] George Anton Kiraz. Compressed storage of sparse finite-state transducers. Lecture Notes in Computer Science, 2214:109–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2214/22140109.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2214/22140109.
pdf.

vanNoord:2001:ERE

[1196] Gertjan van Noord and Dale Gerdemann. An extendible regular expression compiler for finite-state approaches in natural language processing. Lecture Notes in Computer Science, 2214:122–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2214/22140122.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2214/22140122.pdf.

Paun:2001:MPM

[1197] Gheorghe Păun and Gabriel Thierrin.

Multiset processing by means of systems of finite state transducers. Lecture Notes in Computer Science, 2214: 140–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2214/22140140.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2214/22140140.
pdf.

Seuring:2001:SMO

[1198] M. Seuring and M. Gössel. A structural method for output compaction of sequential automata implemented as circuits. Lecture Notes in Computer Science, 2214:158–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2214/22140158.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2214/22140158.
pdf.

Trahtman:2001:AVL

[1199] A. N. Trahtman. An algorithm to verify local threshold testability of deterministic finite automata. Lecture Notes in Computer Science, 2214:164– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2214/22140164.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2214/22140164.
pdf.

Watson:2001:TAC

[1200] Bruce W. Watson. A taxonomy of algorithms for constructing minimal acyclic deterministic finite automata. Lecture Notes in Computer Science, 2214:174– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2214/22140174.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2214/22140174.
pdf.

Anonymous:2001:AIaa

[1201] Anonymous. Author index. Lecture Notes in Computer Science, 2214:183– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2214/2214auth.pdf.

Caires:2001:SLC

[1202] Luís Caires and Luca Cardelli. A spatial logic for concurrency. Lecture Notes in Computer Science, 2215:1– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2215/22150001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2215/22150001.
pdf.

Bugliesi:2001:BA

[1203] Michele Bugliesi, Giuseppe Castagna, and Silvia Crafa. Boxed ambients. Lecture Notes in Computer Science, 2215: 38–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2215/22150038.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2215/22150038.
pdf.

Hoshina:2001:TPC

[1204] Daisuke Hoshina, Eijiro Sumii, and Akinori Yonezawa. A typed process calculus for fine-grained resource access control in distributed computation. Lecture Notes in Computer Science, 2215:64–??, 2001. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2215/22150064.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2215/22150064.
pdf.

Abadi:2001:FEC

[1205] Martín Abadi and Jan Jürjens. Formal eavesdropping and its computational interpretation. Lecture Notes in Computer Science, 2215:82–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2215/22150082.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2215/22150082.
pdf.

Ueda:2001:RPC

[1206] Kazunori Ueda. Resource-passing concurrent programming. Lecture Notes in Computer Science, 2215:95– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2215/22150095.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2215/22150095.
pdf.

Laneve:2001:SD

[1207] Cosimo Laneve, Joachim Parrow, and Björn Victor. Solo diagrams. Lecture Notes in Computer Science, 2215:127– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

bibs/2215/22150127.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2215/22150127.
pdf.

Konig:2001:OES

[1208] Barbara König and Ugo Montanari.

Observational equivalence for synchronized graph rewriting with mobility. Lecture Notes in Computer Science, 2215:145–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2215/22150145.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2215/22150145.
pdf.

Nakano:2001:FPL

[1209] Hiroshi Nakano. Fixed-point logic with the approximation modality and its Kripke completeness. Lecture Notes in Computer Science, 2215:165– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2215/22150165.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2215/22150165.
pdf.

Leivant:2001:TPC

[1210] Daniel Leivant. Termination proofs and complexity certification. Lecture Notes in Computer Science, 2215:183– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2215/22150183.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2215/22150183.
                                        pdf.

Martin:2001:REA

[1211] Keye Martin. A renee equation for algorithmic complexity. Lecture Notes in Computer Science, 2215:201–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2215/22150201.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2215/22150201.
                                        pdf.

Pitts:2001:NLF

[1212] Andrew M. Pitts. Nominal logic: a first order theory of names and binding. Lecture Notes in Computer Science, 2215:219–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2215/22150219.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2215/22150219.
                                        pdf.

Hamana:2001:LPL

[1213] Makoto Hamana. A logic programming language based on binding algebras. Lecture Notes in Computer Science, 2215:243–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2215/22150243.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2215/22150243.
                                        pdf.

Galmiche:2001:PSC

[1214] Didier Galmiche and Daniel Méry.

Proof-search and countermodel generation in propositional BI logic. Lecture Notes in Computer Science, 2215: 263–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2215/22150263.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2215/22150263.
pdf.

Ogawa:2001:GLT

[1215] Mizuhito Ogawa. Generation of a linear time query processing algorithm based on well-quasi-orders. Lecture Notes in Computer Science, 2215:283– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2215/22150283.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2215/22150283.
pdf.

Paulin-Mohring:2001:MTA

[1216] Christine Paulin-Mohring. Modelisation of timed automata in coq. Lecture Notes in Computer Science, 2215: 298–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2215/22150298.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2215/22150298.
pdf.

Esparza:2001:MCL

[1217] Javier Esparza, Antonín Kučera, and Stefan Schwoon. Model-checking LTL with regular valuations for pushdown systems. Lecture Notes in Computer Science, 2215:316–339, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2215/22150316.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2215/22150316.
pdf.

Franzle:2001:WWE

[1218] Martin Fränzle. What will be eventually true of polynomial hybrid automata? Lecture Notes in Computer Science, 2215:340–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2215/22150340.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2215/22150340.
pdf.

Niehren:2001:NSS

[1219] Joachim Niehren and Tim Priesnitz. Non-structural subtype entailment in automata theory. Lecture Notes in Computer Science, 2215:360– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2215/22150360.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2215/22150360.
pdf.

Schnoebelen:2001:BOU

[1220] Ph. Schnoebelen. Bisimulation and other undecidable equivalences for lossy channel systems. Lecture Notes in Computer Science, 2215:385–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2215/22150385.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2215/22150385.
pdf.

Puhakka:2001:WCR

[1221] Antti Puhakka. Weakest congruence results concerning “any-lock”. Lecture Notes in Computer Science, 2215:400– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2215/22150400.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2215/22150400.
pdf.

Banerjee:2001:DCP

[1222] Anindya Banerjee, Nevin Heintze, and Jon G. Riecke. Design and correctness of program transformations based on control-flow analysis. Lecture Notes in Computer Science, 2215: 420–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2215/22150420.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2215/22150420.
pdf.

Bonsangue:2001:IIU

[1223] Marcello M. Bonsangue and Joost N. Kok. Infinite intersection and union types for the lazy lambda calculus. Lecture Notes in Computer Science, 2215: 448–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2215/22150448.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2215/22150448.
                                        pdf.

Yamagata:2001:SNS

[1224] Yoriyuki Yamagata. Strong normalization of second order symmetric lambda-mu calculus. Lecture Notes in Computer Science, 2215:459–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2215/22150459.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2215/22150459.
                                        pdf.

Wadler:2001:GRI

[1225] Philip Wadler. The Girard-Reynolds isomorphism. Lecture Notes in Computer Science, 2215:468–??, 2001. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer-ny.com/link/
                                        service/series/0558/bibs/2215/22150468.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2215/22150468.pdf.

Jackson:2001:LAO

[1226] Daniel Jackson and Alan Fekete. Lightweight analysis of object interactions. Lecture Notes in Computer Science, 2215:492–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2215/22150492.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2215/22150492.pdf.

Beringer:2001:TAP

[1227] Lennart Beringer. Typing assembly programs with explicit forwarding. Lecture Notes in Computer Science, 2215: 514–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2215/22150514.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2215/22150514.
pdf.

Serjantov:2001:UCR

[1228] Andrei Serjantov, Peter Sewell, and Keith Wansbrough. The UDP calculus: Rigorous semantics for real networking. Lecture Notes in Computer Science, 2215:535–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2215/22150535.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2215/22150535.
pdf.

Pierce:2001:UFS

[1229] Benjamin C. Pierce and Jérôme Vouillon. Unison: a file synchronizer and its specification. Lecture Notes in Computer Science, 2215:560–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2215/22150560.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2215/22150560.
                                        pdf.

Anonymous:2001:AIab

[1230] Anonymous. Author index. Lecture Notes in Computer Science, 2215:561– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2215/2215auth.pdf.

Seelam:2001:HBA

[1231] Nagasuresh Seelam, Pankaj Sethi, and Wu chi Feng. A hysteresis based approach for quality, frame rate, and buffer management for video streaming using TCP. Lecture Notes in Computer Science, 2216:1– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2216/22160001.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2216/22160001.
                                        pdf.

Son:2001:TSS

[1232] Sunghoon Son. A transmission scheme for streaming variable bit rate video over Internet. Lecture Notes in Computer Science, 2216:16– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2216/22160016.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2216/22160016.
pdf.

Wakamiya:2001:MVT

[1233] Naoki Wakamiya, Masaki Miyabayashi, Masayuki Murata, and Hideo Miyahara. MPEG-4 video transfer with TCP-friendly rate control. Lecture Notes in Computer Science, 2216:29– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2216/22160029.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2216/22160029.
pdf.

Vicente:2001:IRR

[1234] John Vicente and Andrew T. Campbell. IP radio resource control system. Lecture Notes in Computer Science, 2216:43–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2216/22160043.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2216/22160043.
pdf.

Iraqi:2001:IWI

[1235] Youssef Iraqi and Raouf Boutaba. Is it worth involving several cooperating agents for multimedia user’s admission in cellular networks? Lecture Notes in Computer Science, 2216:57– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2216/22160057.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2216/22160057.
                                        pdf.

Gao:2001:ARA

[1236] Jianbo Gao and Izhak Rubin. Analysis of random access protocol under bursty traffic. Lecture Notes in Computer Science, 2216:71–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2216/22160071.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2216/22160071.
                                        pdf.

Borsos:2001:PMV

[1237] Tamás Borsos. A practical model for VBR video traffic with applications. Lecture Notes in Computer Science, 2216:85–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2216/22160085.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2216/22160085.
                                        pdf.

Lecuire:2001:EQM

[1238] Vincent Lecuire, Francis Lepage, and Khalil Kammoun. Enhancing quality of MPEG video through partially reliable transport service in interactive application. Lecture Notes in Computer Science, 2216:96–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2216/22160096.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2216/22160096.
pdf.

Ahmed:2001:DMM

[1239] Toufik Ahmed, Guillaume Buridant, and Ahmed Mehaoua. Delivering of MPEG-4 multimedia content over next generation Internet. Lecture Notes in Computer Science, 2216:110– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2216/22160110.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2216/22160110.
pdf.

Lelescu:2001:VSS

[1240] Dan Lelescu and Dan Schonfeld. Video skimming and summarization based on principal component analysis. Lecture Notes in Computer Science, 2216:128– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2216/22160128.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2216/22160128.
pdf.

Molenkamp:2001:DRM

[1241] Gary Molenkamp, Michael Katchabaw, Hanan Lutfiyya, and Michael Bauer. Distributed resource management to support distributed application-specific quality of service. Lecture Notes in Computer Science, 2216:142–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2216/22160142.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2216/22160142.
                                        pdf.

Khalil:2001:IBB

[1242] Ibrahim Khalil and Torsten Braun. Implementation of a bandwidth broker for dynamic end-to-end capacity reservation over multiple Diffserv domains. Lecture Notes in Computer Science, 2216:160–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2216/22160160.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2216/22160160.
                                        pdf.

Reason:2001:ICQ

[1243] Johnathan M. Reason and David G. Messerschmitt. The impact of confidentiality on quality of service in heterogeneous voice over IP networks. Lecture Notes in Computer Science, 2216: 175–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2216/22160175.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2216/22160175.
                                        pdf.

Zhai:2001:QLB

[1244] Mingyu Zhai, Guanqun Gu, and Yuan Yuan. Queue length based fair queueing in core-stateless networks. Lecture Notes in Computer Science, 2216:193– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2216/22160193.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2216/22160193.
pdf.

Lee:2001:STU

[1245] Sung-Hyuck Lee, Seung-Joon Seok, Seung-Jin Lee, and Chul-Hee Kang. Study of TCP and UDP flows in a differentiated services network using two markers system. Lecture Notes in Computer Science, 2216:198– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2216/22160198.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2216/22160198.
pdf.

Elbiaze:2001:EBD

[1246] Halima Elbiaze, Tijani Chahed, Gérard Hébuterne, and Tulin Atmaca. Equation-based dynamic shaping algorithm. Lecture Notes in Computer Science, 2216:204–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2216/22160204.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2216/22160204.
pdf.

Zhao:2001:NCB

[1247] Li Zhao, Qi Wang, Yuwen He, Shiqiang Yang, and Yuzhuo Zhong.

A novel content-based video streaming algorithm for fine granular scalable coding. Lecture Notes in Computer Science, 2216:210–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2216/22160210.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2216/22160210.
pdf.

Rubin:2001:TSM

[1248] Izhak Rubin and Patrick Vincent. Topological synthesis of mobile backbone networks for managing ad hoc wireless networks. Lecture Notes in Computer Science, 2216:215–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2216/22160215.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2216/22160215.
pdf.

Ribeiro:2001:QMS

[1249] Marcelo Borges Ribeiro, Lisandro Zambenedetti Granville, Maria Janilce Bosquiroli Almeida, and Liane Margarida Rockenbach Tarouco. QoS monitoring system on IP networks. Lecture Notes in Computer Science, 2216:222–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2216/22160222.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2216/22160222.
pdf.

Li:2001:FSI

[1250] Hongjun Li and John S. Baras. A framework for supporting intelligent fault and performance management for communication networks. Lecture Notes in Computer Science, 2216:227– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2216/22160227.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2216/22160227.
pdf.

Zhang:2001:AGN

[1251] Zheng Zhang, Constantine Manikopoulos, and Jay Jorgenson. Architecture of generalized network service anomaly and fault thresholds. Lecture Notes in Computer Science, 2216:241– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2216/22160241.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2216/22160241.
pdf.

Sarac:2001:PSM

[1252] Kamil Saraç and Kevin C. Almeroth.

Providing scalable many-to-one feedback in multicast reachability monitoring systems. Lecture Notes in Computer Science, 2216:256–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2216/22160256.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2216/22160256.
                                        pdf.

Albaghdadi:2001:FEC

[1253] Mouayad Albaghdadi, Bruce Briley, Martha Evens, Rafid Sukkar, Mohammed Petiwala, and Mark Hamlen. A framework for event correlation in communication systems. Lecture Notes in Computer Science, 2216:271– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2216/22160271.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2216/22160271.
                                        pdf.

Harroud:2001:PBM

[1254] Hamid Harroud, Mouhsine Lakhdissi, Ahmed Karmouch, and Cliff Grossner. Policy-based management for multimedia collaborative services. Lecture Notes in Computer Science, 2216:285– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2216/22160285.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2216/22160285.
                                        pdf.

Chieng:2001:AED

[1255] David Chieng, Alan Marshall, Ivan Ho, and Gerard Parr. Agent-enhanced dynamic service level agreement in future network environments. Lecture Notes in Computer Science, 2216:299– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2216/22160299.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2216/22160299.
pdf.

Elarde:2001:WMC

[1256] Joseph V. Elarde and Gregory B.

Brewster. WEBARM: Mobile code based agent for Web Application Response Measurement — software implementations and analysis. Lecture Notes in Computer Science, 2216:313– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2216/22160313.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2216/22160313.
pdf.

Jagannathan:2001:DPR

[1257] Srinivasan Jagannathan and Kevin C.

Almeroth. The dynamics of price, revenue, and system utilization. Lecture Notes in Computer Science, 2216:329– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2216/22160329.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2216/22160329.
pdf.

Lim:2001:DPB

[1258] Koon-Seng Lim and Rolf Stadler.

Developing pattern-based management programs. Lecture Notes in Computer Science, 2216:345–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2216/22160345.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2216/22160345.
                                        pdf.

Elhajj:2001:SIB

[1259] Imad Elhajj, Ning Xi, Wai Keung Fung, Yun hui Liu, Wen J. Li, Tomoyuki Kaga, and Toshio Fukuda. Supermedia in Internet-based telerobotic operations. Lecture Notes in Computer Science, 2216:359–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2216/22160359.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2216/22160359.
                                        pdf.

Anonymous:2001:AIac

[1260] Anonymous. Author index. Lecture Notes in Computer Science, 2216:373– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2216/2216auth.pdf.

Holland:2001:ILM

[1261] Owen Holland. From the imitation of life to machine consciousness. Lecture Notes in Computer Science, 2217:1– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2217/22170001.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2217/22170001.
pdf.

Floreano:2001:ESN

[1262] Dario Floreano and Claudio Mattiussi.

Evolution of spiking neural controllers for autonomous vision-based robots. Lecture Notes in Computer Science, 2217:38–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2217/22170038.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2217/22170038.
pdf.

Pollack:2001:FTG

[1263] Jordan B. Pollack, Hod Lipson, Pablo Funes, and Gregory Hornby. First three generations of evolved robots. Lecture Notes in Computer Science, 2217:62– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2217/22170062.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2217/22170062.
pdf.

Brooks:2001:STL

[1264] Rodney A. Brooks. Steps towards living machines. Lecture Notes in Computer Science, 2217:72–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2217/22170072.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2217/22170072.
pdf.

Harvey:2001:AEC

[1265] Inman Harvey. Artificial evolution: a continuing SAGA. Lecture Notes in Computer Science, 2217:94–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2217/22170094.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2217/22170094.
                                        pdf.

Full:2001:UBI

[1266] Robert J. Full. Using biological inspiration to build artificial life that locomotes. Lecture Notes in Computer Science, 2217:110–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2217/22170110.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2217/22170110.
                                        pdf.

Mondada:2001:IBA

[1267] Francesco Mondada and Skye Legon. Interactions between art and mobile robotic system engineering. Lecture Notes in Computer Science, 2217:121– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2217/22170121.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2217/22170121.
                                        pdf.

Anonymous:2001:AIad

[1268] Anonymous. Author index. Lecture Notes in Computer Science, 2217:139– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2217/2217auth.pdf.

Kuo:2001:AAJ

[1269] Dean Kuo and Doug Palmer. Automated analysis of Java message service providers. Lecture Notes in Computer Science, 2218:1–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2218/22180001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2218/22180001.
pdf.

Eberhard:2001:EOC

[1270] John Eberhard and Anand Tripathi.

Efficient object caching for distributed Java RMI applications. Lecture Notes in Computer Science, 2218:15– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2218/22180015.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2218/22180015.
pdf.

Brebner:2001:EBB

[1271] Paul Brebner and Shuping Ran. Entity bean A, B, C’s: Enterprise Java beans commit options and caching. Lecture Notes in Computer Science, 2218:36– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2218/22180036.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2218/22180036.
pdf.

Reinstorf:2001:WBS

[1272] Timm Reinstorf, Rainer Ruggaber, Jochen Seitz, and Martina Zitterbart. A WAP-based session layer supporting distributed applications in nomadic environments. Lecture Notes in Computer Science, 2218:56– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2218/22180056.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2218/22180056.
pdf.

Rakotonirainy:2001:MRC

[1273] Andry Rakotonirainy, Jaga Indulska, Seng Wai Loke, and Arkady Zaslavsky. Middleware for reactive components: An integrated use of context, roles, and event based coordination. Lecture Notes in Computer Science, 2218:77– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2218/22180077.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2218/22180077.
pdf.

McKinley:2001:ECP

[1274] Philip K. McKinley, Udiyan I. Padmanabhan, and Nandagopal Ancha.

Experiments in composing proxy audio services for mobile users. Lecture Notes in Computer Science, 2218:99– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2218/22180099.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2218/22180099.
pdf.

Koster:2001:TTI

[1275] Rainer Koster, Andrew P. Black, Jie Huang, Jonathan Walpole, and Calton Pu. Thread transparency in information flow middleware. Lecture Notes in Computer Science, 2218:121– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2218/22180121.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2218/22180121.
pdf.

Sadou:2001:ASH

[1276] Salah Sadou, Gautier Koscielny, and Hafedh Mili. Abstracting services in a heterogeneous environment. Lecture Notes in Computer Science, 2218:141– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2218/22180141.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2218/22180141.
pdf.

Clarke:2001:ECM

[1277] Michael Clarke, Gordon S. Blair, Geoff Coulson, and Nikos Parlavantzas. An efficient component model for the construction of adaptive middleware. Lecture Notes in Computer Science, 2218: 160–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2218/22180160.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2218/22180160.
pdf.

Arregui:2001:RBT

[1278] Damián Arregui, François Pacull, and Jutta Willamowski. Rule-based transactional object migration over a reflective middleware. Lecture Notes in Computer Science, 2218:179–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2218/22180179.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2218/22180179.
pdf.

Houston:2001:CAS

[1279] Iain Houston, Mark C. Little, Ian Robinson, Santosh K. Shrivastava, and Stuart M. Wheater. The CORBA activity service framework for supporting extended transactions. Lecture Notes in Computer Science, 2218:197– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2218/22180197.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2218/22180197.
pdf.

Marsden:2001:FMA

[1280] Eric Marsden and Jean-Charles Fabre.

Failure mode analysis of CORBA service implementations. Lecture Notes in Computer Science, 2218:216–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2218/22180216.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2218/22180216.
pdf.

Kang:2001:RBV

[1281] Soon Ju Kang, Jun Ho Park, and Sung Ho Park. ROOM-BRIDGE: Vertically configurable network architecture and real-time middleware for interoperability between ubiquitous consumer devices in the home. Lecture Notes in Computer Science, 2218:232– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2218/22180232.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2218/22180232.
pdf.

Flinn:2001:REU

[1282] Jason Flinn, Eyal de Lara, Mahadev Satyanarayanan, Dan S. Wallach, and Willy Zwaenepoel. Reducing the energy usage of office applications. Lecture Notes in Computer Science, 2218: 252–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2218/22180252.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2218/22180252.
                                        pdf.

Nakajima:2001:SSA

[1283] Tatsuo Nakajima. System software for audio and visual networked home appliances on commodity operating systems. Lecture Notes in Computer Science, 2218:273–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2218/22180273.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2218/22180273.
                                        pdf.

Bacon:2001:ACT

[1284] Jean Bacon, Ken Moody, and Walt Yao. Access control and trust in the use of widely distributed services. Lecture Notes in Computer Science, 2218: 295–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2218/22180295.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2218/22180295.
                                        pdf.

Laumay:2001:PCS

[1285] Philippe Laumay, Eric Bruneton, Noël De Palma, and Sacha Krakowiak. Preserving causality in a scalable message-oriented middleware. Lecture Notes in Computer Science, 2218:311– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2218/22180311.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2218/22180311.
pdf.

Rowstron:2001:PSD

[1286] Antony Rowstron and Peter Druschel.

Pastry: Scalable, decentralized object location, and routing for largescale peer-to-peer systems. Lecture Notes in Computer Science, 2218:329– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2218/22180329.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2218/22180329.
pdf.

He:2001:PQC

[1287] Jun He, Matti A. Hiltunen, Mohan Rajagopalan, and Richard D. Schlichting. Providing QoS customization in distributed object systems. Lecture Notes in Computer Science, 2218:351– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2218/22180351.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2218/22180351.
pdf.

Wichadakul:2001:IAQ

[1288] Duangdao Wichadakul, Klara Nahrstedt, Xiaohui Gu, and Dongyan Xu. 2K^{Q+}: An integrated approach of QoS compilation and reconfigurable, component-based run-time middleware for the unified QoS management framework. Lecture Notes in Computer Science, 2218:373–394, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2218/22180373.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2218/22180373.
                                        pdf.

Anonymous:2001:AIae

[1289] Anonymous. Author index. Lecture Notes in Computer Science, 2218:395– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2218/2218auth.pdf.

Taft:2001:SG

[1290] S. Tucker Taft, Robert A. Duff, Randall L. Brukardt, and Erhard Ploedereder. Section 1: General. Lecture Notes in Computer Science, 2219:1– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2219/22190001.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2219/22190001.
                                        pdf.

Taft:2001:SLE

[1291] S. Tucker Taft, Robert A. Duff, Randall L. Brukardt, and Erhard Ploedereder. Section 2: Lexical elements. Lecture Notes in Computer Science, 2219:9–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2219/22190009.htm;   http:
                                        //link.springer-ny.com/link/service/

series/0558/papers/2219/22190009.
pdf.

Taft:2001:SDT

[1292] S. Tucker Taft, Robert A. Duff, Randall L. Brukardt, and Erhard Ploedereder. Section 3: Declarations and types. Lecture Notes in Computer Science, 2219:19–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2219/22190019.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2219/22190019.
pdf.

Taft:2001:SNE

[1293] S. Tucker Taft, Robert A. Duff, Randall L. Brukardt, and Erhard Ploedereder. Section 4: Names and expressions. Lecture Notes in Computer Science, 2219:77–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2219/22190077.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2219/22190077.
pdf.

Taft:2001:SSa

[1294] S. Tucker Taft, Robert A. Duff, Randall L. Brukardt, and Erhard Ploedereder. Section 5: Statements. Lecture Notes in Computer Science, 2219: 111–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2219/22190111.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2219/22190111.
                                        pdf.

Taft:2001:SSb

[1295] S. Tucker Taft, Robert A. Duff, Randall L. Brukardt, and Erhard Ploedereder. Section 6: Subprograms. Lecture Notes in Computer Science, 2219: 121–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2219/22190121.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2219/22190121.
                                        pdf.

Taft:2001:SP

[1296] S. Tucker Taft, Robert A. Duff, Randall L. Brukardt, and Erhard Ploedereder. Section 7: Packages. Lecture Notes in Computer Science, 2219:133– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2219/22190133.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2219/22190133.
                                        pdf.

Taft:2001:SVR

[1297] S. Tucker Taft, Robert A. Duff, Randall L. Brukardt, and Erhard Ploedereder. Section 8: Visibility rules. Lecture Notes in Computer Science, 2219: 147–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2219/22190147.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2219/22190147.
pdf.

Taft:2001:STS

[1298] S. Tucker Taft, Robert A. Duff, Randall L. Brukardt, and Erhard Ploedereder. Section 9: Tasks and synchronization. Lecture Notes in Computer Science, 2219:159–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2219/22190159.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2219/22190159.
pdf.

Taft:2001:SPS

[1299] S. Tucker Taft, Robert A. Duff, Randall L. Brukardt, and Erhard Ploedereder. Section 10: Program structure and compilation issues. Lecture Notes in Computer Science, 2219:187– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2219/22190187.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2219/22190187.
pdf.

Taft:2001:SE

[1300] S. Tucker Taft, Robert A. Duff, Randall L. Brukardt, and Erhard Ploedereder. Section 11: Exceptions. Lecture Notes in Computer Science, 2219: 199–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2219/22190199.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2219/22190199.
                                        pdf.

Taft:2001:SGU

[1301] S. Tucker Taft, Robert A. Duff, Randall L. Brukardt, and Erhard Ploedereder. Section 12: Generic units. Lecture Notes in Computer Science, 2219: 209–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2219/22190209.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2219/22190209.
                                        pdf.

Taft:2001:SRI

[1302] S. Tucker Taft, Robert A. Duff, Randall L. Brukardt, and Erhard Ploedereder. Section 13: Representation issues. Lecture Notes in Computer Science, 2219:225–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2219/22190225.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2219/22190225.
                                        pdf.

Taft:2001:ANP

[1303] S. Tucker Taft, Robert A. Duff, Randall L. Brukardt, and Erhard Ploedereder. Annex A (normative) predefined language environment. Lecture Notes in Computer Science, 2219:259– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2219/22190259.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2219/22190259.
pdf.

Taft:2001:ABN

[1304] S. Tucker Taft, Robert A. Duff, Randall L. Brukardt, and Erhard Ploedereder. Annex B (normative) interface to other languages. Lecture Notes in Computer Science, 2219:347– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2219/22190347.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2219/22190347.
pdf.

Taft:2001:ACN

[1305] S. Tucker Taft, Robert A. Duff, Randall L. Brukardt, and Erhard Ploedereder. Annex C (normative) systems programming. Lecture Notes in Computer Science, 2219:371–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2219/22190371.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2219/22190371.
pdf.

Taft:2001:ADN

[1306] S. Tucker Taft, Robert A. Duff, Randall L. Brukardt, and Erhard Ploedereder. Annex D (normative) real-time systems. Lecture Notes in Computer Science, 2219:387–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2219/22190387.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2219/22190387.
                                        pdf.

Taft:2001:AND

[1307] S. Tucker Taft, Robert A. Duff, Randall L. Brukardt, and Erhard Ploedereder. Annex E (normative) distributed systems. Lecture Notes in Computer Science, 2219:407–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2219/22190407.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2219/22190407.
                                        pdf.

Taft:2001:AFN

[1308] S. Tucker Taft, Robert A. Duff, Randall L. Brukardt, and Erhard Ploedereder. Annex F (normative) information systems. Lecture Notes in Computer Science, 2219:421–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2219/22190421.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2219/22190421.
                                        pdf.

Taft:2001:AGN

[1309] S. Tucker Taft, Robert A. Duff, Randall L. Brukardt, and Erhard Ploedereder. Annex G (normative) numerics. Lecture Notes in Computer Science, 2219:437–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2219/22190437.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2219/22190437.
pdf.

Taft:2001:AHN

[1310] S. Tucker Taft, Robert A. Duff, Randall L. Brukardt, and Erhard Ploedereder. Annex H (normative) safety and security. Lecture Notes in Computer Science, 2219:459–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2219/22190459.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2219/22190459.
pdf.

Taft:2001:AJN

[1311] S. Tucker Taft, Robert A. Duff, Randall L. Brukardt, and Erhard Ploedereder. Annex J (normative) obsolescent features. Lecture Notes in Computer Science, 2219:465–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2219/22190465.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2219/22190465.
pdf.

Taft:2001:AKI

[1312] S. Tucker Taft, Robert A. Duff, Randall L. Brukardt, and Erhard Ploedereder. Annex K (informative) language-defined attributes. Lecture Notes in Computer Science, 2219:471– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2219/22190471.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2219/22190471.
                                        pdf.

Taft:2001:AIL

[1313] S. Tucker Taft, Robert A. Duff, Randall L. Brukardt, and Erhard Ploedereder. Annex L (informative) language-defined pragmas. Lecture Notes in Computer Science, 2219:485– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2219/22190485.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2219/22190485.
                                        pdf.

Taft:2001:AMI

[1314] S. Tucker Taft, Robert A. Duff, Randall L. Brukardt, and Erhard Ploedereder. Annex M (informative) implementation-defined characteristics. Lecture Notes in Computer Science, 2219:487–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2219/22190487.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2219/22190487.pdf.

Taft:2001:AIG

[1315] S. Tucker Taft, Robert A. Duff, Randall L. Brukardt, and Erhard Ploedereder. Annex N (informative) glossary. Lecture Notes in Computer Science, 2219:493–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2219/22190493.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2219/22190493.
pdf.

Taft:2001:API

[1316] S. Tucker Taft, Robert A. Duff, Randall L. Brukardt, and Erhard Ploedereder. Annex P (informative) syntax summary. Lecture Notes in Computer Science, 2219:497–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2219/22190497.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2219/22190497.
pdf.

Anonymous:2001:I

[1317] Anonymous. Index. Lecture Notes in Computer Science, 2219:523–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2219/2219indx.pdf.

Luyten:2001:XBR

[1318] Kris Luyten and Karin Coninx.

An XML-based runtime user interface description language for mobile computing devices. Lecture Notes in Computer Science, 2220:1– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2220/22200001.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2220/22200001.
                                        pdf.

Mueller:2001:MBU

[1319] Andreas Mueller, Peter Forbrig, and Clemens Cap. Model-based user interface design using markup concepts. Lecture Notes in Computer Science, 2220:16–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2220/22200016.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2220/22200016.
                                        pdf.

Schneider:2001:AUI

[1320] Kevin A. Schneider and James R. Cordy. Abstract user interfaces: a model and notation to support plasticity in interactive systems. Lecture Notes in Computer Science, 2220:28– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2220/22200028.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2220/22200028.
                                        pdf.

Pribeanu:2001:TMC

[1321] Costin Pribeanu, Quentin Limbourg, and Jean Vanderdonckt. Task modelling for context-sensitive user interfaces. Lecture Notes in Computer Science, 2220:49–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2220/22200049.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2220/22200049.
pdf.

deFQVTurnell:2001:IUI

[1322] Maria de F. Q. V. Turnell, Alexandre Scaico, Marckson R. F. de Sousa, and Angelo Perkusich. Industrial user interface evaluation based on coloured Petri nets modelling and analysis. Lecture Notes in Computer Science, 2220:69– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2220/22200069.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2220/22200069.
pdf.

Navarre:2001:TSI

[1323] David Navarre, Philippe Palanque, Fabio Paternò, Carmen Santoro, and Rémi Bastide. A tool suite for integrating task and system models through scenarios. Lecture Notes in Computer Science, 2220:88–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2220/22200088.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2220/22200088.
pdf.

Du:2001:TPC

[1324] Min Du and David England. Temporal patterns for complex interaction design. Lecture Notes in Computer Science, 2220:114–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2220/22200114.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2220/22200114.pdf.

Garrido:2001:MDG

[1325] José Luis Garrido and Miguel Gea. Modelling dynamic group behaviours. Lecture Notes in Computer Science, 2220:128–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2220/22200128.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2220/22200128.
                                        pdf.

Doherty:2001:RAI

[1326] G. Doherty, M. Massink, and G. Faconti. Reasoning about interactive systems with stochastic models. Lecture Notes in Computer Science, 2220:144– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2220/22200144.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2220/22200144.
                                        pdf.

Limbourg:2001:TUT

[1327] Quentin Limbourg, Costin Pribeanu, and Jean Vanderdonckt. Towards uniformed task models in a modelbased approach. Lecture Notes in Computer Science, 2220:164–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2220/22200164.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2220/22200164.
pdf.

Sutcliffe:2001:HEW

[1328] Alistair Sutcliffe. Heuristic evaluation of Website attractiveness and usability. Lecture Notes in Computer Science, 2220:183–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2220/22200183.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2220/22200183.
pdf.

Thimbleby:2001:AS

[1329] Harold Thimbleby. Affordance and symmetry. Lecture Notes in Computer Science, 2220:199–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2220/22200199.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2220/22200199.
pdf.

Anonymous:2001:AIaf

[1330] Anonymous. Author index. Lecture Notes in Computer Science, 2220:219– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2220/2220auth.pdf.

Crovella:2001:PEH

[1331] Mark E. Crovella. Performance evaluation with heavy tailed distributions. Lecture Notes in Computer Science, 2221:1–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2221/22210001.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2221/22210001.
                                        pdf.

Harchol-Balter:2001:SSW

[1332] Mor Harchol-Balter, Nikhil Bansal, Bianca Schroeder, and Mukesh Agrawal. SRPT scheduling for Web servers. Lecture Notes in Computer Science, 2221: 11–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2221/22210011.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2221/22210011.
                                        pdf.

Yoo:2001:ESC

[1333] Andy B. Yoo and Morris A. Jette. An efficient and scalable coscheduling technique for large symmetric multiprocessor clusters. Lecture Notes in Computer Science, 2221:21–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2221/22210021.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2221/22210021.
pdf.

Gine:2001:CUM

[1334] Francesc Giné, Francesc Solsona, Porfidio Hernández, and Emilio Luque. Coscheduling under memory constraints in a NOW environment. Lecture Notes in Computer Science, 2221: 41–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2221/22210041.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2221/22210041.
pdf.

Bucur:2001:ICP

[1335] A. I. D. Bucur and D. H. J. Epema.

The influence of communication on the performance of co-allocation. Lecture Notes in Computer Science, 2221:66– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2221/22210066.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2221/22210066.
pdf.

Jackson:2001:CAM

[1336] David Jackson, Quinn Snell, and Mark Clement. Core algorithms of the maui scheduler. Lecture Notes in Computer Science, 2221:87–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2221/22210087.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2221/22210087.
                                        pdf.

Zhou:2001:DEC

[1337] B. B. Zhou and R. P. Brent. On the development of an efficient coscheduling system. Lecture Notes in Computer Science, 2221:103–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2221/22210103.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2221/22210103.
                                        pdf.

Suh:2001:EMP

[1338] G. Edward Suh, Larry Rudolph, and Srinivas Devadas. Effects of memory performance on parallel job scheduling. Lecture Notes in Computer Science, 2221:116–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2221/22210116.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2221/22210116.
                                        pdf.

Zhang:2001:IAP

[1339] Y. Zhang, H. Franke, J. E. Moreira, and A. Sivasubramaniam. An integrated approach to parallel scheduling using gang-scheduling, backfilling, and migration. Lecture Notes in Computer Science, 2221:133–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2221/22210133.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2221/22210133.
pdf.

Chiang:2001:CLS

[1340] Su-Hui Chiang and Mary K. Vernon. Characteristics of a large shared memory production workload. Lecture Notes in Computer Science, 2221:159– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2221/22210159.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2221/22210159.
pdf.

Feitelson:2001:MPJ

[1341] Dror G. Feitelson. Metrics for parallel job scheduling and their convergence. Lecture Notes in Computer Science, 2221:188–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2221/22210188.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2221/22210188.
pdf.

Anonymous:2001:AIag

[1342] Anonymous. Author index. Lecture Notes in Computer Science, 2221:207– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2221/2221auth.pdf.

Whitesides:2001:CRI

[1343] Sue Whitesides. Chain reconfiguration. the INs and outs, ups and downs of moving polygons and polygonal linkages. Lecture Notes in Computer Science, 2223:1–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2223/22230001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2223/22230001.
pdf.

Murota:2001:AMC

[1344] Kazuo Murota and Akihisa Tamura. Application of M-convex submodular flow problem to mathematical economics. Lecture Notes in Computer Science, 2223:14–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2223/22230014.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2223/22230014.
pdf.

Deng:2001:PTA

[1345] Xiaotie Deng, Haodi Feng, Pixing Zhang, and Hong Zhu. A polynomial time approximation scheme for minimizing total completion time of unbounded batch scheduling. Lecture Notes in Computer Science, 2223:26– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2223/22230026.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2223/22230026.
pdf.

Karuno:2001:PTA

[1346] Yoshiyuki Karuno and Hiroshi Nagamochi. A polynomial time approximation scheme for the multi-vehicle scheduling problem on a path with release and handling times. Lecture Notes in Computer Science, 2223:36– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2223/22230036.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2223/22230036.
pdf.

Chen:2001:SNS

[1347] Jianer Chen and Jingui Huang. Seminormal schedulings: Improvement on Goemans’ algorithm. Lecture Notes in Computer Science, 2223:48–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2223/22230048.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2223/22230048.
pdf.

Koga:2001:BST

[1348] Hisashi Koga. Balanced scheduling toward loss-free packet queuing and delay fairness. Lecture Notes in Computer Science, 2223:61–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2223/22230061.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2223/22230061.
                                        pdf.

Kim:2001:BUL

[1349] Jae-Hoon Kim and Kyung-Yong Chwa. Broadcasting with universal lists revisited: Using competitive analysis. Lecture Notes in Computer Science, 2223: 74–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2223/22230074.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2223/22230074.
                                        pdf.

Nomura:2001:AFD

[1350] Kumiko Nomura, Toshinori Yamada, and Shuichi Ueno. On adaptive fault diagnosis for multiprocessor systems. Lecture Notes in Computer Science, 2223:86–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2223/22230086.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2223/22230086.
                                        pdf.

Hashimoto:2001:LMA

[1351] Kenta Hashimoto, Toshinori Yamada, and Shuichi Ueno. On-line multicasting in all-optical networks. Lecture Notes in Computer Science, 2223:99– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2223/22230099.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2223/22230099.
pdf.

Nakano:2001:EF

[1352] Shin ichi Nakano. Enumerating floorplans with n rooms. Lecture Notes in Computer Science, 2223:107–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2223/22230107.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2223/22230107.
pdf.

Galbiati:2001:MMC

[1353] Giulia Galbiati. On min-max cycle bases. Lecture Notes in Computer Science, 2223:116–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2223/22230116.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2223/22230116.pdf.

Nagamochi:2001:MLV

[1354] Hiroshi Nagamochi and Toshimasa Ishii. On the minimum local-vertex-connectivity augmentation in graphs. Lecture Notes in Computer Science, 2223:124–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2223/22230124.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2223/22230124.
pdf.

Nagoya:2001:CGI

[1355] Takayuki Nagoya. Counting graph isomorphisms among chordal graphs with restricted clique number. Lecture Notes in Computer Science, 2223:136– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2223/22230136.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2223/22230136.
pdf.

Sadakane:2001:QAI

[1356] Kunihiko Sadakane, Norito Sugawara, and Takeshi Tokuyama. Quantum algorithms for intersection and proximity problems. Lecture Notes in Computer Science, 2223:148–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2223/22230148.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2223/22230148.
pdf.

Sun:2001:BAA

[1357] Zheng Sun and John Reif. BUSH-WHACK: An approximation algorithm for minimal paths through pseudo-Euclidean spaces. Lecture Notes in Computer Science, 2223:160–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2223/22230160.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2223/22230160.
pdf.

Chin:2001:AMT

[1358] Francis Y. L. Chin and Stanley P. Y. Fung. Approximation of minimum triangulation for polyhedron with bounded degrees. Lecture Notes in Computer Science, 2223:172–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2223/22230172.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2223/22230172.
pdf.

Schindelhauer:2001:TAW

[1359] Christian Schindelhauer and Birgitta Weber. Tree-approximations for the weighted cost-distance problem (extended abstract). Lecture Notes in Computer Science, 2223:185–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2223/22230185.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2223/22230185.
pdf.

Mizuki:2001:NSN

[1360] Takaaki Mizuki and Takao Nishizeki.

Necessary and sufficient numbers of cards for sharing secret keys on hierarchical groups. Lecture Notes in Computer Science, 2223:196–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2223/22230196.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2223/22230196.
                                        pdf.

Yamamura:2001:GAK

[1361] Akihiro Yamamura and Kaoru Kurosawa. Generic algorithms and key agreement protocols based on group actions. Lecture Notes in Computer Science, 2223:208–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2223/22230208.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2223/22230208.
                                        pdf.

Breutzmann:2001:BCN

[1362] Josef M. Breutzmann, David W. Juedes, and Jack H. Lutz. Baire category and nowhere differentiability for feasible real functions. Lecture Notes in Computer Science, 2223:219– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2223/22230219.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2223/22230219.
                                        pdf.

Horiyama:2001:TAC

[1363] Takashi Horiyama and Toshihide Ibaraki. Translation among CNFs, characteristic models and ordered binary decision diagrams. Lecture Notes in Computer Science, 2223:231–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2223/22230231.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2223/22230231.
pdf.

Ibarra:2001:RPS

[1364] Oscar H. Ibarra and Zhe Dang. On removing the pushdown stack in reachability constructions. Lecture Notes in Computer Science, 2223:244–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2223/22230244.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2223/22230244.
pdf.

Yamamoto:2001:NRA

[1365] Hiroaki Yamamoto. A new recognition algorithm for extended regular expressions. Lecture Notes in Computer Science, 2223:257–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2223/22230257.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2223/22230257.
pdf.

Koshiba:2001:PTA

[1366] Takeshi Koshiba. Polynomial-time algorithms for the equivalence for oneway quantum finite automata. Lecture Notes in Computer Science, 2223:268– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2223/22230268.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2223/22230268.
                                        pdf.

Ono:2001:IDS

[1367] Hirotaka Ono, Mutsunori Yagiura, and Toshihide Ibaraki. An index for the data size to extract decomposable structures in LAD. Lecture Notes in Computer Science, 2223:279– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2223/22230279.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2223/22230279.
                                        pdf.

Fellows:2001:PCM

[1368] Michael R. Fellows. Parameterized complexity: The main ideas and some research frontiers. Lecture Notes in Computer Science, 2223:291–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2223/22230291.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2223/22230291.
                                        pdf.

Biedl:2001:TBM

[1369] Therese Biedl, Erik D. Demaine, Christian A. Duncan, Rudolf Fleischer, and Stephen G. Kobourov. Tight bounds on maximal and maximum matchings. Lecture Notes in Computer Science, 2223:308–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2223/22230308.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2223/22230308.
pdf.

Nikolopoulos:2001:ROA

[1370] Stavros D. Nikolopoulos and Leonidas Palios. Recognition and orientation algorithms for P_{4}-comparability graphs. Lecture Notes in Computer Science, 2223:320–331, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2223/22230320.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2223/22230320.pdf.

Chen:2001:EAT

[1371] Danny Z. Chen and Xiaodong Wu. Efficient algorithms for k-terminal cuts on planar graphs. Lecture Notes in Computer Science, 2223:332–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2223/22230332.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2223/22230332.
pdf.

Galluccio:2001:PTA

[1372] Anna Galluccio and Guido Proietti. Polynomial time algorithms for edge-connectivity augmentation of Hamiltonian paths. Lecture Notes in Computer Science, 2223:345–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2223/22230345.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2223/22230345.
                                        pdf.

Albert:2001:API

[1373] Michael H. Albert, Robert E. L. Aldred, Mike D. Atkinson, and Derek A. Holton. Algorithms for pattern involvement in permutations. Lecture Notes in Computer Science, 2223:355– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2223/22230355.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2223/22230355.
                                        pdf.

Uno:2001:FAE

[1374] Takeaki Uno. A fast algorithm for enumerating bipartite perfect matchings. Lecture Notes in Computer Science, 2223:367–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2223/22230367.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2223/22230367.
                                        pdf.

Chen:2001:LSB

[1375] Bo Chen, Xiaotie Deng, and Wenan Zang. On-line scheduling a batch processing system to minimize total weighted job completion time. Lecture Notes in Computer Science, 2223:380– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2223/22230380.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2223/22230380.
pdf.

Erlebach:2001:CTA

[1376] Thomas Erlebach, Martin Gantenbein, Daniel Hürlimann, Gabriele Neyer, Aris Pagourtzis, Paolo Penna, Konrad Schlude, Kathleen Steinhöfel, David Scot Taylor, and Peter Widmayer. On the complexity of train assignment problems. Lecture Notes in Computer Science, 2223:390–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2223/22230390.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2223/22230390.
pdf.

Aspnes:2001:CTP

[1377] James Aspnes, Julia Hartling, Ming-Yang Kao, Junhyong Kim, and Gauri Shah. A combinatorial toolbox for protein sequence design and landscape analysis in the grand canonical model. Lecture Notes in Computer Science, 2223:403–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2223/22230403.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2223/22230403.
pdf.

Lyngso:2001:CCH

[1378] Rune B. Lyngsø and Christian N. S.

Pedersen. Complexity of comparing hidden Markov models. Lecture Notes in Computer Science, 2223:416– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2223/22230416.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2223/22230416.
                                        pdf.

Kao:2001:DSA

[1379] Ming-Yang Kao and Vijay Ramachandran. DNA self-assembly for constructing 3D boxes. Lecture Notes in Computer Science, 2223:429–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2223/22230429.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2223/22230429.
                                        pdf.

Gramm:2001:ESC

[1380] Jens Gramm, Rolf Niedermeier, and Peter Rossmanith. Exact solutions for CLOSEST STRING and related problems. Lecture Notes in Computer Science, 2223:441–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2223/22230441.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2223/22230441.
                                        pdf.

Chen:2001:TPI

[1381] Danny Z. Chen, Shuang Luan, and Jinhui Xu. Topological peeling and implementation. Lecture Notes in Computer Science, 2223:454–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2223/22230454.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2223/22230454.
pdf.

Chen:2001:ISM

[1382] Danny Z. Chen, Jie Wang, and Xiaodong Wu. Image segmentation with monotonicity and smoothness constraints. Lecture Notes in Computer Science, 2223:467–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2223/22230467.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2223/22230467.
pdf.

Lee:2001:OAS

[1383] Jae-Ha Lee, Sang-Min Park, and Kyung-Yong Chwa. Optimization algorithms for sweeping a polygonal region with mobile guards. Lecture Notes in Computer Science, 2223:480– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2223/22230480.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2223/22230480.
pdf.

Kovaleva:2001:AGS

[1384] Sofia Kovaleva and Frits C. R. Spieksma. Approximation of a geometric set covering problem. Lecture Notes in Computer Science, 2223:493– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2223/22230493.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2223/22230493.
                                        pdf.

Goldberg:2001:SPA

[1385] Andrew V. Goldberg. Shortest path algorithms: Engineering aspects. Lecture Notes in Computer Science, 2223: 502–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2223/22230502.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2223/22230502.
                                        pdf.

Zhou:2001:EAW

[1386] Xiao Zhou and Takao Nishizeki. Efficient algorithms for weighted colorings of series-parallel graphs. Lecture Notes in Computer Science, 2223:514– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2223/22230514.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2223/22230514.
                                        pdf.

Peinado:2001:GWA

[1387] Marcus Peinado. Go with the winners algorithms for cliques in random graphs. Lecture Notes in Computer Science, 2223:525–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2223/22230525.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2223/22230525.
pdf.

Fiala:2001:CPC

[1388] Jirí Fiala and Jan Kratochvíl. Complexity of partial covers of graphs. Lecture Notes in Computer Science, 2223: 537–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2223/22230537.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2223/22230537.
pdf.

Bodlaender:2001:GTM

[1389] Hans L. Bodlaender, Michael J.

Dinneen, and Bakhadyr Khoussainov. On game-theoretic models of networks. Lecture Notes in Computer Science, 2223:550–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2223/22230550.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2223/22230550.
pdf.

Jakoby:2001:CSB

[1390] Andreas Jakoby and Maciej Liskiewicz. The complexity of some basic problems for dynamic process graphs. Lecture Notes in Computer Science, 2223:562– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2223/22230562.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2223/22230562.
                                        pdf.

Lin:2001:DOQ

[1391] Xuemin Lin. Delay optimizations in quorum consensus. Lecture Notes in Computer Science, 2223:575–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2223/22230575.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2223/22230575.
                                        pdf.

Lee:2001:RSQ

[1392] Hyunyoung Lee and Jennifer L. Welch. Randomized shared queues applied to distributed optimization algorithms. Lecture Notes in Computer Science, 2223:587–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2223/22230587.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2223/22230587.
                                        pdf.

Brodnik:2001:MTQ

[1393] Andrej Brodnik and Johan Karlsson.

Multiprocess time queue. Lecture Notes in Computer Science, 2223:599– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2223/22230599.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2223/22230599.
pdf.

Poon:2001:LPW

[1394] Sheung-Hung Poon, Chan-Su Shin, Tycho Strijk, and Alexander Wolff. Labeling points with weights. Lecture Notes in Computer Science, 2223:610– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2223/22230610.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2223/22230610.
pdf.

Bremner:2001:SCQ

[1395] David Bremner, Ferran Hurtado, Suneeta Ramaswami, and Vera Sacristán. Small convex quadrangulations of point sets. Lecture Notes in Computer Science, 2223:623–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2223/22230623.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2223/22230623.
pdf.

Asano:2001:HCC

[1396] Tetsuo Asano and Takeshi Tokuyama. How to color a checkerboard with a given distribution — matrix rounding achieving low 2 × 2-discrepancy. Lecture Notes in Computer Science, 2223: 636–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2223/22230636.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2223/22230636.
pdf.

Garrido:2001:LSL

[1397] Mari Ángeles Garrido, Claudia Iturriaga, Alberto Márquez, José Ramón Portillo, Pedro Reyes, and Alexander Wolff. Labeling subway lines. Lecture Notes in Computer Science, 2223:649– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2223/22230649.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2223/22230649.
pdf.

Zhang:2001:CST

[1398] Louxin Zhang and Song Zhu. Complexity study on two clustering problems (extended abstract). Lecture Notes in Computer Science, 2223:660– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2223/22230660.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2223/22230660.
pdf.

Fujito:2001:MGA

[1399] Toshihiro Fujito and Tsuyoshi Okumura. A modified greedy algorithm for the set cover problem with weights 1 and 2. Lecture Notes in Computer Science, 2223:670–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2223/22230670.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2223/22230670.
pdf.

Zhao:2001:UFA

[1400] Liang Zhao, Hiroshi Nagamochi, and Toshihide Ibaraki. A unified framework for approximating multiway partition problems (extended abstract). Lecture Notes in Computer Science, 2223:682– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2223/22230682.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2223/22230682.
pdf.

Babel:2001:LAC

[1401] Luitpold Babel, Bo Chen, Hans Kellerer, and Vladimir Kotov. Online algorithms for cardinality constrained bin packing problems. Lecture Notes in Computer Science, 2223: 695–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2223/22230695.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2223/22230695.
                                        pdf.

Monostori:2001:SVS

[1402] Krisztián Monostori, Arkady Zaslavsky, and István Vajk. Suffix vector: a space-efficient suffix tree representation. Lecture Notes in Computer Science, 2223:707–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2223/22230707.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2223/22230707.
                                        pdf.

Hori:2001:FPM

[1403] Hideaki Hori, Shinichi Shimozono, Masayuki Takeda, and Ayumi Shinohara. Fragmentary pattern matching: Complexity, algorithms and applications for analyzing classic literary works. Lecture Notes in Computer Science, 2223:719–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2223/22230719.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2223/22230719.
                                        pdf.

Brodal:2001:CQD

[1404] Gerth Stølting Brodal, Rolf Fagerberg, and Christian N. S. Pedersen. Computing the quartet distance between evolutionary trees in time O(n log^{2} n). Lecture Notes in Computer Science, 2223: 731–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2223/22230731.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2223/22230731.
pdf.

Becker:2001:PTN

[1405] Ronald I. Becker, Yen-I Chiang, Isabella Lari, and Andrea Scozzari. The Cent-dian path problem on tree networks. Lecture Notes in Computer Science, 2223:743–755, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2223/22230743.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2223/22230743.
pdf.

Kranakis:2001:AHA

[1406] Evangelos Kranakis, Danny Krizanc, and Sunil Shende. Approximate hotlink assignment. Lecture Notes in Computer Science, 2223:756–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2223/22230756.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2223/22230756.
pdf.

Ku:2001:EAT

[1407] Shan-Chyun Ku, Chi-Jen Lu, Biing-Feng Wang, and Tzu-Chin Lin. Efficient algorithms for two generalized 2-median problems on trees. Lecture Notes in Computer Science, 2223:768– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2223/22230768.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2223/22230768.
                                        pdf.

Anonymous:2001:AIah

[1408] Anonymous. Author index. Lecture Notes in Computer Science, 2223:779– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2223/2223ind.pdf.

Chen:2001:CMC

[1409] Qiming Chen, Umeshwar Dayal, and Meichun Hsu. Conceptual modeling for collaborative E-business processes. Lecture Notes in Computer Science, 2224:1–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2224/22240001.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2224/22240001.
                                        pdf.

Enoki:2001:MNF

[1410] Kei ichi Enoki. “i-mode”, now & future. Lecture Notes in Computer Science, 2224:17–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2224/22240017.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2224/22240017.pdf.

Tsichritzis:2001:RRN

[1411] Dennis Tsichritzis. The role of research in the new economy. Lecture Notes in Computer Science, 2224:19– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2224/22240019.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2224/22240019.
pdf.

Bussler:2001:SBI

[1412] Christoph Bussler. Semantic B2B integration: Is modeling the superior approach over programming? Lecture Notes in Computer Science, 2224:26– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2224/22240026.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2224/22240026.
pdf.

Price:2001:MTC

[1413] Rosanne Price, Nectaria Tryfona, and Christian S. Jensen. Modeling topological constraints in spatial part-whole relationships. Lecture Notes in Computer Science, 2224:27– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2224/22240027.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2224/22240027.
pdf.

Ishikawa:2001:SDB

[1414] Yoshiharu Ishikawa and Hiroyuki Kitagawa. Source description-based approach for the modeling of spatial information integration. Lecture Notes in Computer Science, 2224:41– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2224/22240041.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2224/22240041.
pdf.

Behr:2001:TRC

[1415] Thomas Behr and Markus Schneider. Topological relationships of complex points and complex regions. Lecture Notes in Computer Science, 2224:56– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2224/22240056.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2224/22240056.
pdf.

Ram:2001:DDS

[1416] Sudha Ram, Richard T. Snodgrass, Vijay Khatri, and Yousub Hwang. DIS-TIL: a design support environment for conceptual modeling of spatiotemporal requirements. Lecture Notes in Computer Science, 2224:70–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2224/22240070.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2224/22240070.
pdf.

Griffiths:2001:TCM

[1417] Tony Griffiths, Alvaro A. A. Fernandes, Norman W. Paton, Keith T. Mason, Bo Huang, and Michael Worboys. Tripod: a comprehensive model for spatial and aspatial historical objects. Lecture Notes in Computer Science, 2224:84– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2224/22240084.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2224/22240084.
pdf.

Schneider:2001:DTP

[1418] Markus Schneider. A design of topological predicates for complex crisp and fuzzy regions. Lecture Notes in Computer Science, 2224:103–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2224/22240103.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2224/22240103.
pdf.

Rodriguez-Gianolli:2001:SAX

[1419] Patricia Rodríguez-Gianolli and John Mylopoulos. A semantic approach to XML-based data integration. Lecture Notes in Computer Science, 2224:117– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2224/22240117.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2224/22240117.
                                        pdf.

dosSantosMello:2001:RBC

[1420] Ronaldo dos Santos Mello and Carlos Alberto Heuser. A rule-based conversion of a DTD to a conceptual schema. Lecture Notes in Computer Science, 2224:133–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2224/22240133.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2224/22240133.
                                        pdf.

Mani:2001:SDM

[1421] Murali Mani, Dongwon Lee, and Richard R. Muntz. Semantic data modeling using XML schemas. Lecture Notes in Computer Science, 2224:149– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2224/22240149.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2224/22240149.
                                        pdf.

Yu:2001:MSA

[1422] Eric Yu, Lin Liu, and Ying Li. Modelling strategic actor relationships to support intellectual property management. Lecture Notes in Computer Science, 2224:164–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2224/22240164.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2224/22240164.
pdf.

Thalheim:2001:SCM

[1423] Bernhard Thalheim and Antje Düsterhöft.

SiteLang: Conceptual modeling of Internet sites. Lecture Notes in Computer Science, 2224:179–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2224/22240179.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2224/22240179.
pdf.

Karlapalem:2001:FWM

[1424] Kamalakar Karlapalem, Ajay R. Dani, and P. Radha Krishna. A frame work for modeling electronic contracts. Lecture Notes in Computer Science, 2224: 193–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2224/22240193.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2224/22240193.
pdf.

Detienne:2001:CTS

[1425] Virginie Detienne and Jean-Luc Hainaut. CASE tool support for temporal database design. Lecture Notes in Computer Science, 2224:208–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2224/22240208.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2224/22240208.
                                        pdf.

Tawil:2001:DSM

[1426] Abdel-Rahman H. Tawil, Nicholas J. Fiddian, and William A. Gray. Domain-specific metadata a key for building semantically-rich schema models. Lecture Notes in Computer Science, 2224:225–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2224/22240225.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2224/22240225.pdf.

Hartmann:2001:CIC

[1427] Sven Hartmann. Coping with inconsistent constraint specifications. Lecture Notes in Computer Science, 2224: 241–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2224/22240241.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2224/22240241.
                                        pdf.

Berti-Equille:2001:IBD

[1428] Laure Berti-Equille. Integration of biological data and quality-driven source negotiation. Lecture Notes in Computer Science, 2224:256–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2224/22240256.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2224/22240256.
pdf.

Cali:2001:ADI

[1429] Andrea Calì, Diego Calvanese, Giuseppe De Giacomo, and Maurizio Lenzerini. Accessing data integration systems through conceptual schemas. Lecture Notes in Computer Science, 2224:270– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2224/22240270.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2224/22240270.
pdf.

Peng:2001:RCH

[1430] Zhiyong Peng and Yahiko Kambayashi.

Resolving conflicts and handling replication during integration of multiple databases by object deputy model. Lecture Notes in Computer Science, 2224: 285–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2224/22240285.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2224/22240285.
pdf.

Hong:2001:EER

[1431] Seokjin Hong, Byoungho Song, and Sukho Lee. Efficient execution of range-aggregate queries in data warehouse environments. Lecture Notes in Computer Science, 2224:299–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2224/22240299.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2224/22240299.
                                        pdf.

Priebe:2001:PAC

[1432] Torsten Priebe and Günther Pernul. A pragmatic approach to conceptual modeling of OLAP security. Lecture Notes in Computer Science, 2224:311– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2224/22240311.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2224/22240311.
                                        pdf.

Theodoratos:2001:RAI

[1433] Dimitri Theodoratos, Theodore Dalamagas, Alkis Simitsis, and Manos Stavropoulos. A randomized approach for the incremental design of an evolving data warehouse. Lecture Notes in Computer Science, 2224:325– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2224/22240325.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2224/22240325.
                                        pdf.

Pinet:2001:SST

[1434] François Pinet and Ahmed Lbath. Semantics of stereotypes for type specification in UML: Theory and practice. Lecture Notes in Computer Science, 2224:339–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2224/22240339.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2224/22240339.
pdf.

Evermann:2001:TOB

[1435] Joerg Evermann and Yair Wand. Towards ontologically based semantics for UML constructs. Lecture Notes in Computer Science, 2224:354–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2224/22240354.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2224/22240354.
pdf.

Song:2001:DSD

[1436] Il-Yeol Song. Developing sequence diagrams in UML. Lecture Notes in Computer Science, 2224:368–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2224/22240368.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2224/22240368.
pdf.

Kraft:2001:THL

[1437] Peter Kraft and Jens Otto Sørensen.

Translation of a high-level temporal model into lower level models: Impact of modelling at different description levels. Lecture Notes in Computer Science, 2224:383–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2224/22240383.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2224/22240383.
                                        pdf.

Costal:2001:RTR

[1438] Dolors Costal, Antoni Olivé, and Ernest Teniente. Relationship type refinement in conceptual models with multiple classification. Lecture Notes in Computer Science, 2224:397–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2224/22240397.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2224/22240397.
                                        pdf.

Doerr:2001:MPW

[1439] Martin Doerr, Dimitris Plexousakis, and Chryssoyla Bekiari. A metamodel for part — whole relationships for reasoning on missing parts and reconstruction. Lecture Notes in Computer Science, 2224:412–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2224/22240412.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2224/22240412.
                                        pdf.

Embley:2001:DXD

[1440] David W. Embley and Wai Y. Mok. Developing XML documents with guaranteed “good” properties. Lecture Notes in Computer Science, 2224:426– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2224/22240426.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2224/22240426.
pdf.

Akoka:2001:DHD

[1441] Jacky Akoka, Isabelle Comyn-Wattiau, and Nicolas Prat. Dimension hierarchies design from UML generalizations and aggregations. Lecture Notes in Computer Science, 2224:442– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2224/22240442.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2224/22240442.
pdf.

Lepper:2001:MMN

[1442] Markus Lepper, Baltasar Trancón y Widemann, and Jacob Wieland. Minimize mark-up! Natural writing should guide the design of textual modeling frontends. Lecture Notes in Computer Science, 2224:456– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2224/22240456.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2224/22240456.
pdf.

Ralyte:2001:AMR

[1443] Jolita Ralyté and Colette Rolland. An approach for method reengineering. Lecture Notes in Computer Science, 2224:471–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2224/22240471.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2224/22240471.
                                        pdf.

Nakanishi:2001:MSM

[1444] Yoshihiro Nakanishi, Tatsuo Hirose, and Katsumi Tanaka. Modeling and structuring multiple perspective video for browsing. Lecture Notes in Computer Science, 2224:485–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2224/22240485.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2224/22240485.
                                        pdf.

Acharya:2001:MDO

[1445] Biswajit Acharya, Arun K. Majumdar, and Jayanta Mukherjee. Modeling dynamic objects in video databases: a logic based approach. Lecture Notes in Computer Science, 2224:499– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2224/22240499.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2224/22240499.
                                        pdf.

Sadiq:2001:PFW

[1446] Shazia Sadiq, Wasim Sadiq, and Maria Orlowska. Pockets of flexibility in workflow specification. Lecture Notes in Computer Science, 2224:513–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2224/22240513.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2224/22240513.
pdf.

Taveter:2001:AOE

[1447] Kuldar Taveter and Gerd Wagner.

Agent-oriented enterprise modeling based on business rules. Lecture Notes in Computer Science, 2224:527– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2224/22240527.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2224/22240527.
pdf.

Chiu:2001:TLM

[1448] Dickson K. W. Chiu. A three-layer model for workflow semantic recovery in an object-oriented environment. Lecture Notes in Computer Science, 2224: 541–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2224/22240541.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2224/22240541.
pdf.

Embley:2001:ROA

[1449] David W. Embley, Yiu-Kai Ng, and Li Xu. Recognizing ontology-applicable multiple-record Web documents. Lecture Notes in Computer Science, 2224:

555–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2224/22240555.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2224/22240555.
                                        pdf.

Schewe:2001:QWI

[1450] Klaus-Dieter Schewe. Querying Web information systems. Lecture Notes in Computer Science, 2224:571–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2224/22240571.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2224/22240571.
                                        pdf.

Anido:2001:CMF

[1451] Luis Anido, Martín Llamas, Manuel J. Fernández, Judith Rodríguez, Juan Santos, and Manuel Caeiro. A conceptual modelling framework for standards-driven Web-based distance learning. Lecture Notes in Computer Science, 2224:585–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2224/22240585.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2224/22240585.pdf.

Kunieda:2001:XSD

[1452] Takayuki Kunieda and Yuki Wakita. XML-schema dynamic mapped multimedia content description tool. Lecture Notes in Computer Science, 2224:

599–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2224/22240599.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2224/22240599.
pdf.

Jin:2001:CDM

[1453] Eunsook Jin, Jungmin Seo, Kwangjoon Lee, and Ju won Song. Constructing a data map of Korea Telecom. Lecture Notes in Computer Science, 2224: 602–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2224/22240602.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2224/22240602.
pdf.

Oba:2001:CSA

[1454] Tatsuo Oba. Competency of set analysis in CRM closed loop marketing. Lecture Notes in Computer Science, 2224: 604–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2224/22240604.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2224/22240604.
pdf.

Paul:2001:XBE

[1455] Ray Paul, Wei-Tek Tsai, Bing Li, and Xiaoying Bai. XML-based E2E test report management. Lecture Notes in Computer Science, 2224:607– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2224/22240607.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2224/22240607.
                                        pdf.

Horiuchi:2001:FRB

[1456] Hajime Horiuchi, Masaharu Obayashi, and Yoshihide Nagase. A framework for reusing business objects by multilevel object patterns. Lecture Notes in Computer Science, 2224:609– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2224/22240609.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2224/22240609.
                                        pdf.

Honishi:2001:MSB

[1457] Takashi Honishi, Gengo Suzuki, Nobuyuki Kobayashi, and Kazuya Konishi. A mediation system based on universal relation modeling. Lecture Notes in Computer Science, 2224: 611–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2224/22240611.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2224/22240611.
                                        pdf.

Anonymous:2001:AIai

[1458] Anonymous. Author index. Lecture Notes in Computer Science, 2224:613– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
papers/2224/2224indx.pdf.

Abe:2001:EI

[1459] Naoki Abe, Roni Khardon, and Thomas Zeugmann. Editors’ introduction. Lecture Notes in Computer Science, 2225:1–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2225/22250001.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2225/22250001.pdf.

Arikawa:2001:DSPa

[1460] Setsuo Arikawa. The discovery science project in japan. Lecture Notes in Computer Science, 2225:9– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2225/22250009.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2225/22250009.
pdf.

Angluin:2001:QRa

[1461] Dana Angluin. Queries revisited. Lecture Notes in Computer Science, 2225: 12–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2225/22250012.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2225/22250012.
pdf.

Cohen:2001:RBa

[1462] Paul R. Cohen, Tim Oates, Niall Adams, and Carole R. Beal. Robot baby 2001. Lecture Notes in Computer Science, 2225:32–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2225/22250032.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2225/22250032.pdf.

Darden:2001:DMCa

[1463] Lindley Darden. Discovering mechanisms: a computational philosophy of science perspective. Lecture Notes in Computer Science, 2225:57– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2225/22250057.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2225/22250057.
pdf.

Shneiderman:2001:IDTa

[1464] Ben Shneiderman. Inventing discovery tools: Combining information visualization with data mining. Lecture Notes in Computer Science, 2225:58– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2225/22250058.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2225/22250058.
pdf.

Yang:2001:LCB

[1465] Ke Yang. On learning correlated Boolean functions using statistical queries (extended abstract). Lecture Notes in Computer Science, 2225:59– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2225/22250059.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2225/22250059.
pdf.

Hatano:2001:SAM

[1466] Kohei Hatano. A simpler analysis of the multi-way branching decision tree boosting algorithm. Lecture Notes in Computer Science, 2225:77– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2225/22250077.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2225/22250077.
pdf.

Sima:2001:MQT

[1467] Jiří Šíma. Minimizing the quadratic training error of a sigmoid neuron is hard. Lecture Notes in Computer Science, 2225:92–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2225/22250092.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2225/22250092.pdf.

Sadohara:2001:LBF

[1468] Ken Sadohara. Learning of Boolean functions using support vector machines. Lecture Notes in Computer Science, 2225:106–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2225/22250106.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2225/22250106.pdf.

Balcazar:2001:RST

[1469] Jose Balcázar, Yang Dai, and Osamu Watanabe. A random sampling technique for training support vector machines. Lecture Notes in Computer Science, 2225:119–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2225/22250119.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2225/22250119.
                                        pdf.

Garg:2001:LCC

[1470] Ashutosh Garg and Dan Roth. Learning coherent concepts. Lecture Notes in Computer Science, 2225:135–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2225/22250135.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2225/22250135.
                                        pdf.

Kwek:2001:LIC

[1471] Stephen S. Kwek. Learning intermediate concepts. Lecture Notes in Computer Science, 2225:151–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2225/22250151.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2225/22250151.
pdf.

Dooly:2001:RVM

[1472] Daniel R. Dooly, Sally A. Goldman, and Stephen S. Kwek. Real-valued multiple-instance learning with queries. Lecture Notes in Computer Science, 2225:167–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2225/22250167.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2225/22250167.
pdf.

Kalnishkan:2001:LFC

[1473] Yuri Kalnishkan, Michael V. Vyugin, and Volodya Vovk. Loss functions, complexities, and the Legendre transformation. Lecture Notes in Computer Science, 2225:181–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2225/22250181.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2225/22250181.
pdf.

Vyugin:2001:NLI

[1474] Michael V. Vyugin and Vladimir V. V’yugin. Non-linear inequalities between predictive and Kolmogorov complexities. Lecture Notes in Computer Science, 2225:190–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2225/22250190.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2225/22250190.pdf.

Jain:2001:LST

[1475] Sanjay Jain and Frank Stephan. Learning by switching type of information. Lecture Notes in Computer Science, 2225:205–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2225/22250205.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2225/22250205.
                                        pdf.

Jain:2001:LHS

[1476] Sanjay Jain and Frank Stephan. Learning how to separate. Lecture Notes in Computer Science, 2225:219–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2225/22250219.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2225/22250219.
                                        pdf.

Jain:2001:LLU

[1477] Sanjay Jain, Yen Kaow Ng, and Tiong Seng Tay. Learning languages in a union. Lecture Notes in Computer Science, 2225:235–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2225/22250235.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2225/22250235.
pdf.

Zilles:2001:CII

[1478] Sandra Zilles. On the comparison of inductive inference criteria for uniform learning of finite classes. Lecture Notes in Computer Science, 2225:251– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2225/22250251.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2225/22250251.
pdf.

Mukouchi:2001:RLL

[1479] Yasuhito Mukouchi and Masako Sato.

Refutable language learning with a neighbor system. Lecture Notes in Computer Science, 2225:267–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2225/22250267.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2225/22250267.
pdf.

Jain:2001:LRF

[1480] Sanjay Jain, Efim Kinber, Rolf Wiehagen, and Thomas Zeugmann. Learning recursive functions refutably. Lecture Notes in Computer Science, 2225:283– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2225/22250283.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2225/22250283.
                                        pdf.

Merkle:2001:RLR

[1481] Wolfgang Merkle and Frank Stephan. Refuting learning revisited. Lecture Notes in Computer Science, 2225:299– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2225/22250299.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2225/22250299.
                                        pdf.

Arimura:2001:ELS

[1482] Hiroki Arimura, Hiroshi Sakamoto, and Setsuo Arikawa. Efficient learning of semi-structured data from queries. Lecture Notes in Computer Science, 2225: 315–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2225/22250315.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2225/22250315.
                                        pdf.

Lange:2001:EEF

[1483] Steffen Lange, Gunter Grieser, and Klaus P. Jantke. Extending elementary formal systems. Lecture Notes in Computer Science, 2225:332–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2225/22250332.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2225/22250332.
pdf.

Denis:2001:LRL

[1484] François Denis, Aurélien Lemay, and Alain Terlutte. Learning regular languages using RFSA. Lecture Notes in Computer Science, 2225:348–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2225/22250348.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2225/22250348.
pdf.

delaHiguera:2001:ILP

[1485] Colin de la Higuera and Jean-Christophe Janodet. Inference of ω-languages from prefixes. Lecture Notes in Computer Science, 2225:364– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2225/22250364.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2225/22250364.
pdf.

Anonymous:2001:AIaj

[1486] Anonymous. Author index. Lecture Notes in Computer Science, 2225:379– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        papers/2225/2225auth.pdf.

Arikawa:2001:DSPb

[1487] Setsuo Arikawa. The discovery science project in japan. Lecture Notes in Computer Science, 2226:1– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2226/22260001.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2226/22260001.
                                        pdf.

Darden:2001:DMCb

[1488] Lindley Darden. Discovering mechanisms: a computational philosophy of science perspective. Lecture Notes in Computer Science, 2226:3– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2226/22260003.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2226/22260003.
                                        pdf.

Angluin:2001:QRb

[1489] Dana Angluin. Queries revisited. Lecture Notes in Computer Science, 2226: 16–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2226/22260016.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2226/22260016.
                                        pdf.

Shneiderman:2001:IDTb

[1490] Ben Shneiderman. Inventing discovery tools: Combining information visualization with data mining. Lecture Notes in Computer Science, 2226:17– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2226/22260017.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2226/22260017.
pdf.

Cohen:2001:RBb

[1491] Paul R. Cohen, Tim Oates, Niall Adams, and Carole R. Beal. Robot baby 2001. Lecture Notes in Computer Science, 2226:29–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2226/22260029.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2226/22260029.pdf.

Bannai:2001:VVM

[1492] Hideo Bannai, Yoshinori Tamada, Osamu Maruyama, and Satoru Miyano. VML: a view modeling language for computational knowledge discovery. Lecture Notes in Computer Science, 2226:30–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2226/22260030.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2226/22260030.
pdf.

Dzeroski:2001:CDC

[1493] Sašo Džeroski and Pat Langley. Computational discovery of communicable knowledge: Symposium report. Lecture Notes in Computer Science, 2226: 45–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2226/22260045.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2226/22260045.
                                        pdf.

Fortes:2001:BNI

[1494] I. Fortes, J. L. Balcázar, and R. Morales. Bounding negative information in frequent sets algorithms. Lecture Notes in Computer Science, 2226:50–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2226/22260050.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2226/22260050.
                                        pdf.

Gama:2001:FT

[1495] João Gama. Functional trees. Lecture Notes in Computer Science, 2226:59– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2226/22260059.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2226/22260059.
                                        pdf.

Gorman:2001:SHS

[1496] Michael E. Gorman, Alexandra Kincannon, and Matthew M. Mehalik. Spherical horses and shared toothbrushes: Lessons learned from a workshop on scientific and technological thinking. Lecture Notes in Computer Science, 2226:74–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2226/22260074.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2226/22260074.
pdf.

Grundel:2001:CAN

[1497] Hans Gründel, Tino Naphtali, Christian Wiech, Jan-Marian Gluba, Maiken Rohdenburg, and Tobias Scheffer. Clipping and analyzing news using machine learning techniques. Lecture Notes in Computer Science, 2226:87– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2226/22260087.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2226/22260087.
pdf.

Horvath:2001:TDD

[1498] Tamás Horváth and Stefan Wrobel.

Towards discovery of deep and wide first-order structures: a case study in the domain of mutagenicity. Lecture Notes in Computer Science, 2226:100– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2226/22260100.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2226/22260100.
                                        pdf.

Ikeda:2001:EUP

[1499] Daisuke Ikeda, Yasuhiro Yamada, and Sachio Hirokawa. Eliminating useless parts in semi-structured documents using alternation counts. Lecture Notes in Computer Science, 2226: 113–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2226/22260113.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2226/22260113.
                                        pdf.

Iyer:2001:MBE

[1500] Naresh S. Iyer and John R. Josephson. Multicriterially best explanations. Lecture Notes in Computer Science, 2226: 128–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2226/22260128.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2226/22260128.
                                        pdf.

Kanda:2001:CAI

[1501] Kouta Kanda, Makoto Haraguchi, and Yoshiaki Okubo. Constructing approximate informative basis of association rules. Lecture Notes in Computer Science, 2226:141–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2226/22260141.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2226/22260141.
pdf.

Kise:2001:PBD

[1502] Koichi Kise, Markus Junker, Andreas Dengel, and Keinosuke Matsumoto. Passage-based document retrieval as a tool for text mining with user’s information needs. Lecture Notes in Computer Science, 2226:155– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2226/22260155.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2226/22260155.
pdf.

Kocabas:2001:AFR

[1503] Sakir Kocabas. Automated formulation of reactions and pathways in nuclear astrophysics: New results. Lecture Notes in Computer Science, 2226: 170–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2226/22260170.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2226/22260170.
pdf.

Kocabas:2001:IFE

[1504] Sakir Kocabas and Pat Langley. An integrated framework for extended discovery in particle physics. Lecture Notes in Computer Science, 2226:182– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2226/22260182.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2226/22260182.
                                        pdf.

Kostoff:2001:SD

[1505] Ronald N. Kostoff. Stimulating discovery. Lecture Notes in Computer Science, 2226:196–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2226/22260196.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2226/22260196.pdf.

Mahidadia:2001:AMD

[1506] Ashesh Mahidadia and Paul Compton. Assisting model-discovery in neuroendocrinology. Lecture Notes in Computer Science, 2226:214–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2226/22260214.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2226/22260214.
                                        pdf.

Martin:2001:GTD

[1507] Eric Martin, Arun Sharma, and Frank Stephan. A general theory of deduction, induction, and learning. Lecture Notes in Computer Science, 2226:228– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2226/22260228.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2226/22260228.
                                        pdf.

Maruyama:2001:LCR

[1508] Osamu Maruyama, Takayoshi Shoudai, Emiko Furuichi, Satoru Kuhara, and Satoru Miyano. Learning conformation rules. Lecture Notes in Computer Science, 2226:243–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2226/22260243.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2226/22260243.
pdf.

Matsumura:2001:KNV

[1509] Naohiro Matsumura, Yukio Ohsawa, and Mitsuru Ishizuka. Knowledge navigation on visualizing complementary documents. Lecture Notes in Computer Science, 2226:258–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2226/22260258.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2226/22260258.
pdf.

Matsuo:2001:KEK

[1510] Yutaka Matsuo, Yukio Ohsawa, and Mitsuru Ishizuka. KeyWorld: Extracting keywords from a document as a small world. Lecture Notes in Computer Science, 2226:271–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2226/22260271.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2226/22260271.
                                        pdf.

Murata:2001:MDP

[1511] Tsuyoshi Murata. A method for discovering purified Web communities. Lecture Notes in Computer Science, 2226: 282–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2226/22260282.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2226/22260282.
                                        pdf.

Ouyang:2001:DCM

[1512] Ming Ouyang, John Case, and Joan Burnside. Divide and conquer machine learning for a genomics analogy problem (progress report). Lecture Notes in Computer Science, 2226: 290–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2226/22260290.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2226/22260290.
                                        pdf.

Phillips:2001:TMS

[1513] Joseph Phillips. Towards a method of searching a diverse theory space for scientific discovery. Lecture Notes in Computer Science, 2226:304–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2226/22260304.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2226/22260304.
pdf.

Robardet:2001:ELS

[1514] Céline Robardet and Fabien Feschet. Efficient local search in conceptual clustering. Lecture Notes in Computer Science, 2226:323–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2226/22260323.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2226/22260323.
pdf.

Saito:2001:CRQ

[1515] Kazumi Saito, Pat Langley, Trond Grenager, Christopher Potter, Alicia Torregrosa, and Steven A. Klooster. Computational revision of quantitative scientific models. Lecture Notes in Computer Science, 2226:336–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2226/22260336.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2226/22260336.
pdf.

Sugimoto:2001:EDE

[1516] Noriko Sugimoto, Hiroki Ishizaka, and Takeshi Shinohara. An efficient derivation for elementary formal systems based on partial unification. Lecture Notes in Computer Science, 2226:350– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2226/22260350.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2226/22260350.
                                        pdf.

Suzuki:2001:WCA

[1517] Einoshin Suzuki. Worst-case analysis of rule discovery. Lecture Notes in Computer Science, 2226:365–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2226/22260365.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2226/22260365.
                                        pdf.

Taniguchi:2001:MSS

[1518] Katsuaki Taniguchi, Hiroshi Sakamoto, Hiroki Arimura, Shinichi Shimozono, and Setsuo Arikawa. Mining semistructured data by path expressions. Lecture Notes in Computer Science, 2226:378–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2226/22260378.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2226/22260378.
                                        pdf.

Todorovski:2001:TRE

[1519] Ljupčo Todorovski and Sašo Džeroski. Theory revision in equation discovery. Lecture Notes in Computer Science, 2226:389–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2226/22260389.htm;   http:
                                        //link.springer-ny.com/link/service/

series/0558/papers/2226/22260389.
pdf.

Ueda:2001:STA

[1520] Nobuhisa Ueda and Taisuke Sato. Simplified training algorithms for hierarchical hidden Markov models. Lecture Notes in Computer Science, 2226:401– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2226/22260401.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2226/22260401.
pdf.

Yamamoto:2001:DRE

[1521] Koichiro Yamamoto, Masayuki Takeda, Ayumi Shinohara, Tomoko Fukuda, and Ichirō Nanri. Discovering repetitive expressions and affinities from anthologies of classical Japanese poems. Lecture Notes in Computer Science, 2226:416–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2226/22260416.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2226/22260416.
pdf.

Hiraishi:2001:WSR

[1522] Hironori Hiraishi, Hisayoshi Kato, Naonori Ohtsuka, and Fumio Mizoguchi. Web site rating and improvement based on hyperlink structure. Lecture Notes in Computer Science, 2226:429–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2226/22260429.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2226/22260429.
                                        pdf.

Hirao:2001:PAF

[1523] Masahiro Hirao, Shunsuke Inenaga, Ayumi Shinohara, Masayuki Takeda, and Setsuo Arikawa. A practical algorithm to find the best episode patterns. Lecture Notes in Computer Science, 2226:435–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2226/22260435.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2226/22260435.
                                        pdf.

Hochheiser:2001:IET

[1524] Harry Hochheiser and Ben Shneiderman. Interactive exploration of time series data. Lecture Notes in Computer Science, 2226:441–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2226/22260441.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2226/22260441.
                                        pdf.

Lele:2001:CRU

[1525] Shreevardhan Lele, Bruce Golden, Kimberly Ozga, and Edward Wasil. Clustering rules using empirical similarity of support sets. Lecture Notes in Computer Science, 2226:447–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2226/22260447.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2226/22260447.
pdf.

Simina:2001:CLC

[1526] Marin Simina, Michael E. Gorman, and Janet L. Kolodner. Computational lessons from a cognitive study of invention. Lecture Notes in Computer Science, 2226:452–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2226/22260452.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2226/22260452.
pdf.

Tanaka:2001:CBF

[1527] Yuzuru Tanaka and Tsuyoshi Sugibuchi. Component-based framework for virtual information materialization. Lecture Notes in Computer Science, 2226:458–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2226/22260458.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2226/22260458.
pdf.

Tang:2001:DAS

[1528] Lida Tang and Ben Shneiderman.

Dynamic aggregation to support pattern discovery: a case study with Web logs. Lecture Notes in Computer Science, 2226:464–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2226/22260464.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2226/22260464.
                                        pdf.

Ueno:2001:SPM

[1529] Genta Ueno, Nagatomo Nakamura, and Tomoyuki Higuchi. Separation of photoelectrons via multivariate Maxwellian mixture model. Lecture Notes in Computer Science, 2226: 470–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2226/22260470.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2226/22260470.
                                        pdf.

vandenBosch:2001:LDD

[1530] Alexander P. M. van den Bosch. Logic of drug discovery: a descriptive model of a practice in neuropharmacology. Lecture Notes in Computer Science, 2226:476–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2226/22260476.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2226/22260476.
                                        pdf.

Yamada:2001:SRE

[1531] Yasuhiro Yamada, Daisuke Ikeda, and Sachio Hirokawa. SCOOP: a record extractor without knowledge on input. Lecture Notes in Computer Science, 2226:482–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2226/22260482.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2226/22260482.
pdf.

Zak:2001:MAM

[1532] Premysl Zak, Pavel Spacil, and Jaroslava Halova. Meta-analysis of mutagenes discovery. Lecture Notes in Computer Science, 2226:488–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2226/22260488.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2226/22260488.
pdf.

Anonymous:2001:AIak

[1533] Anonymous. Author index. Lecture Notes in Computer Science, 2226:493– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2226/2226auth.pdf.

Massey:2001:URM

[1534] J. L. Massey. The ubiquity of Reed–Muller codes. Lecture Notes in Computer Science, 2227:1–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2227/22270001.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2227/22270001.
                                        pdf.

Pless:2001:SDC

[1535] V. Pless. Self-dual codes-theme and variations. Lecture Notes in Computer Science, 2227:13–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2227/22270013.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2227/22270013.
                                        pdf.

Shokrollahi:2001:DDS

[1536] A. Shokrollahi. Design of differential space-time codes using group theory. Lecture Notes in Computer Science, 2227:22–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2227/22270022.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2227/22270022.
                                        pdf.

Sudan:2001:IEC

[1537] M. Sudan. Ideal error-correcting codes: Unifying algebraic and numbertheoretic algorithms. Lecture Notes in Computer Science, 2227:36–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2227/22270036.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2227/22270036.
                                        pdf.

Baliga:2001:SDC

[1538] A. Baliga and J. Chua. Selfdual codes using image restoration techniques. Lecture Notes in Computer Science, 2227:46–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2227/22270046.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2227/22270046.
pdf.

Cadic:2001:LCT

[1539] E. Cadic, J. C. Carlach, G. Olocco, A. Otmani, and J. P. Tillich. Low complexity tail-biting trellises of selfdual codes of length 24, 32 and 40 over GF (2) and Z_{4} of large minimum distance. Lecture Notes in Computer Science, 2227:57–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2227/22270057.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2227/22270057.
pdf.

Dey:2001:FQL

[1540] B. Kumar Dey and B. Sundar Rajan. F_{q}-linear cyclic codes over F_{q^{m}}: DFT characterization. Lecture Notes in Computer Science, 2227:67– 76, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2227/22270067.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2227/22270067.
                                        pdf.

Berger:2001:CPR

[1541] Thierry P. Berger and Louis de Maximy. Cyclic projective Reed–Muller codes. Lecture Notes in Computer Science, 2227:77–81, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2227/22270077.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2227/22270077.pdf.

Laihonen:2001:CIS

[1542] T. Laihonen and S. Ranto. Codes identifying sets of vertices. Lecture Notes in Computer Science, 2227:82– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2227/22270082.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2227/22270082.
                                        pdf.

Schaathun:2001:DGW

[1543] H. G. Schaathun. Duality and greedy weights of linear codes and projective multisets. Lecture Notes in Computer Science, 2227:92–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2227/22270092.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2227/22270092.
                                        pdf.

Betsumiya:2001:TIC

[1544] K. Betsumiya, M. Harada, and A. Munemasa. Type II codes over F_{2^{r}}. Lecture Notes in Computer Science, 2227:102–111, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2227/22270102.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2227/22270102.pdf.

Gupta:2001:SSC

[1545] M. K. Gupta, D. G. Glynn, and T. A.

Gulliver. On senary simplex codes. Lecture Notes in Computer Science, 2227: 112–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2227/22270112.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2227/22270112.
pdf.

Gulliver:2001:ODC

[1546] T. A. Gulliver and M. Harada. Optimal double circulant Z_{4}-codes. Lecture Notes in Computer Science, 2227: 122–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2227/22270122.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2227/22270122.
pdf.

Guruswami:2001:CCN

[1547] V. Guruswami. Constructions of codes from number fields. Lecture Notes in Computer Science, 2227:129– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2227/22270129.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2227/22270129.
pdf.

Horimoto:2001:GHW

[1548] H. Horimoto and K. Shiromoto. On generalized Hamming weights for codes over finite chain rings. Lecture Notes in Computer Science, 2227:141–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2227/22270141.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2227/22270141.
pdf.

Kelarev:2001:IRW

[1549] A. Kelarev and O. Sokratova. Information rates and weights of codes in structural matrix rings. Lecture Notes in Computer Science, 2227:151– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2227/22270151.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2227/22270151.
pdf.

Geil:2001:HC

[1550] O. Geil and T. Høholdt. On hyperbolic codes. Lecture Notes in Computer Science, 2227:159–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2227/22270159.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2227/22270159.
pdf.

Sakata:2001:FIM

[1551] S. Sakata. On fast interpolation method for Guruswami–Sudan list decoding of one-point algebraicgeometry codes. Lecture Notes in Computer Science, 2227:172–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2227/22270172.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2227/22270172.
pdf.

Rodriguez-Palanquex:2001:CGC

[1552] M. C. Rodríguez-Palánquex, L. J.

García-Villalba, and I. Luengo-Velasco. Computing the genus of a class of curves. Lecture Notes in Computer Science, 2227:182–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2227/22270182.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2227/22270182.
pdf.

Gutierrez:2001:IMP

[1553] J. Gutierrez and D. Gomez-Perez.

Iterations of multivariate polynomials and discrepancy of pseudorandom numbers. Lecture Notes in Computer Science, 2227:192–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2227/22270192.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2227/22270192.
                                        pdf.

Parker:2001:ELB

[1554] M. G. Parker. Even length binary sequence families with low negaperiodic autocorrelation. Lecture Notes in Computer Science, 2227:200–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2227/22270200.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2227/22270200.
                                        pdf.

Parraud:2001:NEA

[1555] P. Parraud. On the non-existence of (almost-)perfect quaternary sequences. Lecture Notes in Computer Science, 2227:210–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2227/22270210.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2227/22270210.
                                        pdf.

Peinado:2001:MP

[1556] A. Peinado, F. Montoya, J. Muñoz, and A. J. Yuste. Maximal periods of x + c in F_{q}. Lecture Notes in Computer Science, 2227:219–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2227/22270219.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2227/22270219.
pdf.

Udaya:2001:ACF

[1557] P. Udaya and S. Boztaş. On the aperiodic correlation function of Galois ring m-sequences. Lecture Notes in Computer Science, 2227:229–238, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2227/22270229.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2227/22270229.
pdf.

Wang:2001:EMM

[1558] L. Wang. Euclidean modules and multisequence synthesis. Lecture Notes in Computer Science, 2227:239–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2227/22270239.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2227/22270239.
pdf.

Charnes:2001:HBF

[1559] C. Charnes, M. Rötteler, and T. Beth.

On homogeneous bent functions. Lecture Notes in Computer Science, 2227: 249–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2227/22270249.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2227/22270249.
                                        pdf.

Encheva:2001:PIC

[1560] S. Encheva and G. Cohen. Partially identifying codes for copyright protection. Lecture Notes in Computer Science, 2227:260–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2227/22270260.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2227/22270260.
                                        pdf.

Shparlinski:2001:GHN

[1561] I. E. Shparlinski. On the generalised hidden number problem and bit security of XTR. Lecture Notes in Computer Science, 2227:268–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2227/22270268.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2227/22270268.
                                        pdf.

Ustimenko:2001:CGT

[1562] V. Ustimenko. CRYPTIM: Graphs as tools for symmetric encryption. Lecture Notes in Computer Science, 2227: 278–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2227/22270278.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2227/22270278.
pdf.

Alvarez:2001:ACC

[1563] V. Álvarez, J. A. Armario, M. D.

Frau, and P. Real. An algorithm for computing cocyclic matrices developed over some semidirect products. Lecture Notes in Computer Science, 2227: 287–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2227/22270287.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2227/22270287.
pdf.

Giesbrecht:2001:ALI

[1564] M. Giesbrecht, M. Jacobson, Jr., and A. Storjohann. Algorithms for large integer matrix problems. Lecture Notes in Computer Science, 2227:297– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2227/22270297.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2227/22270297.
pdf.

Honkala:2001:IVE

[1565] I. Honkala, M. G. Karpovsky, and S. Litsyn. On the identification of vertices and edges using cycles. Lecture Notes in Computer Science, 2227:308– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2227/22270308.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2227/22270308.
                                        pdf.

Balakirsky:2001:ASD

[1566] V. B. Balakirsky. On algebraic soft decision decoding of cyclic binary codes. Lecture Notes in Computer Science, 2227:315–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2227/22270315.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2227/22270315.
                                        pdf.

Byrne:2001:LDS

[1567] E. Byrne. Lifting decoding schemes over a Galois ring. Lecture Notes in Computer Science, 2227:323–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2227/22270323.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2227/22270323.
                                        pdf.

Kasami:2001:SCM

[1568] T. Kasami, H. Tokushige, and Y. Kaji. Sufficient conditions on most likely local sub-codewords in recursive maximum likelihood decoding algorithms. Lecture Notes in Computer Science, 2227:333–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2227/22270333.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2227/22270333.
pdf.

Kuijper:2001:UST

[1569] M. Kuijper, M. van Dijk, H. Hollmann, and J. Oostveen. A unifying systemtheoretic framework for errors-and-erasures Reed–Solomon decoding. Lecture Notes in Computer Science, 2227: 343–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2227/22270343.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2227/22270343.
pdf.

Wadayama:2001:ACR

[1570] T. Wadayama. An algorithm for computing rejection probability of MLD with threshold test over BSC. Lecture Notes in Computer Science, 2227:353– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2227/22270353.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2227/22270353.
pdf.

Boulier:2001:CCS

[1571] F. Boulier and S. Neut. Cartan’s characters and stairs of characteristic sets. Lecture Notes in Computer Science, 2227:363–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2227/22270363.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2227/22270363.
                                        pdf.

Gaudry:2001:IQJ

[1572] P. Gaudry and É. Schost. On the invariants of the quotients of the Jacobian of a curve of genus 2. Lecture Notes in Computer Science, 2227:373– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2227/22270373.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2227/22270373.
                                        pdf.

Guidi:2001:ACP

[1573] A. M. Guidi, A. J. Grant, and S. S. Pietrobon. Algebraic constructions for PSK space-time coded modulation. Lecture Notes in Computer Science, 2227:387–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2227/22270387.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2227/22270387.
                                        pdf.

Anonymous:2001:AIal

[1574] Anonymous. Author index. Lecture Notes in Computer Science, 2227:397– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2227/2227auth.pdf.

Koukopoulos:2001:SIH

[1575] D. K. Koukopoulos, S. E. Nikoletseas, and Paul G. Spirakis. Stability issues in heterogeneous and FIFO networks under the adversarial queueing model. Lecture Notes in Computer Science, 2228:3– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2228/22280003.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2228/22280003.
pdf.

Kosaraju:2001:MAM

[1576] S. Rao Kosaraju. Mesh algorithms for multiplication and division. Lecture Notes in Computer Science, 2228:17– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2228/22280017.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2228/22280017.
pdf.

Chandra:2001:CRD

[1577] Punit Chandra and Ajay D. Kshemkalyani.

Compact routing in directed networks with stretch factor of two. Lecture Notes in Computer Science, 2228:24– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2228/22280024.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2228/22280024.
pdf.

Subramani:2001:PSA

[1578] K. Subramani. Parametric scheduling — algorithms and complexity. Lecture Notes in Computer Science, 2228:36– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2228/22280036.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2228/22280036.
                                        pdf.

Panda:2001:EAC

[1579] B. S. Panda and Sajal K. Das. An efficient algorithm for computing lower bounds on time and processors for scheduling precedence graphs on multicomputer systems. Lecture Notes in Computer Science, 2228:47– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2228/22280047.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2228/22280047.
                                        pdf.

Streit:2001:JSH

[1580] Achim Streit. On job scheduling for HPC-clusters and the dynP scheduler. Lecture Notes in Computer Science, 2228:58–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2228/22280058.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2228/22280058.
                                        pdf.

Al-Omari:2001:ASF

[1581] R. Al-Omari, Arun K. Somani, and G. Manimaran. An adaptive scheme for fault-tolerant scheduling of soft real-time tasks in multiprocessor systems. Lecture Notes in Computer Science, 2228:68–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2228/22280068.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2228/22280068.
pdf.

Gropp:2001:LSM

[1582] William D. Gropp. Learning from the success of MPI. Lecture Notes in Computer Science, 2228:81–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2228/22280081.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2228/22280081.
pdf.

Lewandowski:2001:GSP

[1583] J. L. V. Lewandowski, W. W. Lee, and Z. Lin. Gyrokinetic simulations of plasma turbulence on massively parallel computers. Lecture Notes in Computer Science, 2228:95– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2228/22280095.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2228/22280095.
pdf.

Chronopoulos:2001:PKT

[1584] Anthony T. Chronopoulos and Andrey B. Kucherov. A parallel Krylov-type method for nonsymmetric linear systems. Lecture Notes in Computer Science, 2228:104–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2228/22280104.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2228/22280104.
pdf.

Ganguly:2001:ECA

[1585] Niloy Ganguly, Arijit Das, Pradipta Maji, and Biplab K. Sikdar. Evolving cellular automata based associative memory for pattern recognition. Lecture Notes in Computer Science, 2228: 115–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2228/22280115.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2228/22280115.
pdf.

Hariharan:2001:EPA

[1586] Bhanu Hariharan and Srinivas Aluru. Efficient parallel algorithms and software for compressed octrees with applications to hierarchical methods. Lecture Notes in Computer Science, 2228: 125–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2228/22280125.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2228/22280125.
pdf.

Salamon:2001:CSI

[1587] Victor Salamon, Paul Lu, Ben Watson, Dima Brodsky, and Dave Gomboc. A case study of improving memory locality in polygonal model simplification: Metrics and performance. Lecture Notes in Computer Science, 2228: 137–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2228/22280137.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2228/22280137.
pdf.

Sarkar:2001:HPS

[1588] Vivek Sarkar and Julian Dolby.

High-performance scalable Java virtual machines. Lecture Notes in Computer Science, 2228:151–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2228/22280151.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2228/22280151.
pdf.

Gibson:2001:SVM

[1589] Courtney R. Gibson and Angelos Bilas. Shared virtual memory clusters with next-generation interconnection networks and wide compute nodes. Lecture Notes in Computer Science, 2228: 167–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2228/22280167.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2228/22280167.
                                        pdf.

Shahabuddin:2001:SPR

[1590] Johara Shahabuddin, Abhay Chrungoo, Vishu Gupta, Sandeep Juneja, Sanjiv Kapoor, and Arun Kumar. Stream-packing: Resource allocation in Web server farms with a QoS guarantee. Lecture Notes in Computer Science, 2228:182–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2228/22280182.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2228/22280182.
                                        pdf.

Ozer:2001:WMT

[1591] Emre Özer, Thomas M. Conte, and Saurabh Sharma. Weld: a multithreading technique towards latencytolerant VLIW processors. Lecture Notes in Computer Science, 2228:192– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2228/22280192.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2228/22280192.
                                        pdf.

Aggarwal:2001:PDV

[1592] Aneesh Aggarwal and Manoj Franklin. Putting data value predictors to work in fine-grain parallel processors. Lecture Notes in Computer Science, 2228: 204–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2228/22280204.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2228/22280204.
pdf.

Aragon:2001:CEB

[1593] Juan L. Aragón, José González, José M. García, and Antonio González. Confidence estimation for branch prediction reversal. Lecture Notes in Computer Science, 2228:214–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2228/22280214.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2228/22280214.
pdf.

Ravindran:2001:RPP

[1594] Rajiv Ravindran and Rajat Moona.

Retargetable program profiling using high level processor models. Lecture Notes in Computer Science, 2228:224– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2228/22280224.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2228/22280224.
pdf.

Cociorva:2001:TAS

[1595] D. Cociorva, J. Wilkins, G. Baumgartner, P. Sadayappan, J. Ramanujam, M. Nooijen, D. Bernholdt, and R. Harrison. Towards automatic synthesis of high-performance codes for electronic structure calculations: Data locality optimization. Lecture Notes in Computer Science, 2228:237–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2228/22280237.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2228/22280237.
                                        pdf.

Sivathanu:2001:BAF

[1596] Muthian Sivathanu, Venkateshwaran Venkataramani, and Remzi H. Arpaci-Dusseau. Block asynchronous I/O: a flexible infrastructure for userlevel filesystems. Lecture Notes in Computer Science, 2228:249–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2228/22280249.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2228/22280249.
                                        pdf.

T:2001:TOS

[1597] Subramania Sharma T. and Matthew J. Thazhuthaveetil. TWLinuX: Operating system support for optimistic parallel discrete event simulation. Lecture Notes in Computer Science, 2228:262– 271, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2228/22280262.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2228/22280262.
                                        pdf.

Ahn:2001:LCG

[1598] JinHo Ahn, Sung-Gi Min, and Chong-Sun Hwang. Low-cost garbage collection for causal message logging. Lecture Notes in Computer Science, 2228: 272–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2228/22280272.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2228/22280272.
pdf.

Jaleel:2001:IPI

[1599] Aamer Jaleel and Bruce Jacob. Improving the precise interrupt mechanism of software-managed TLB miss handlers. Lecture Notes in Computer Science, 2228:282–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2228/22280282.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2228/22280282.
pdf.

Manjunath:2001:HCA

[1600] K. V. Manjunath and R. Govindarajan. Hidden costs in avoiding false sharing in software DSMs. Lecture Notes in Computer Science, 2228:294– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2228/22280294.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2228/22280294.
pdf.

Braun:2001:HCG

[1601] Tracy D. Braun, Howard Jay Siegel, and Anthony A. Maciejewski. Heterogeneous computing: Goals, methods, and open problems. Lecture Notes in Computer Science, 2228:307– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2228/22280307.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2228/22280307.
pdf.

De:2001:MAC

[1602] Swades De and Sajal K. Das. Maximum achievable capacity gain through traffic load balancing in cellular radio networks: a practical perspective. Lecture Notes in Computer Science, 2228: 321–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2228/22280321.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2228/22280321.
pdf.

Jha:2001:PEM

[1603] Rahul Jha and Sridhar Iyer. Performance evaluation of mobile agents for E-commerce applications. Lecture Notes in Computer Science, 2228:331– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2228/22280331.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2228/22280331.
pdf.

Fernandez:2001:PER

[1604] Juan Fernández, José M. García, and José Duato. Performance evaluation of real-time communication services on high-speed LANs under topology changes. Lecture Notes in Computer Science, 2228:341–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2228/22280341.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2228/22280341.
pdf.

Sivakumar:2001:WCP

[1605] Mahesh Sivakumar and Suresh Subramaniam. Wavelength conversion placement and wavelength assignment in WDM optical networks. Lecture Notes in Computer Science, 2228:351– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2228/22280351.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2228/22280351.
pdf.

Smitha:2001:ILT

[1606] Smitha, Inkoo Kim, and A. L. Narasimha Reddy. Identifying long-term highbandwidth flows at a router. Lecture Notes in Computer Science, 2228:361– 371, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2228/22280361.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2228/22280361.
                                        pdf.

Manjunath:2001:VLP

[1607] D. Manjunath and Biplab Sikdar. Variable length packet switches: Input queued fabrics with finite buffers, speedup, and parallelism. Lecture Notes in Computer Science, 2228:372– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2228/22280372.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2228/22280372.
                                        pdf.

Kitsuregawa:2001:WMP

[1608] Masaru Kitsuregawa, Iko Pramudiono, and Katsumi Takahashi. Web mining is parallel. Lecture Notes in Computer Science, 2228:385–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2228/22280385.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2228/22280385.
                                        pdf.

Bannister:2001:OBI

[1609] Joe Bannister, Joe Touch, Purushotham Kamath, and Aatash Patel. An optical booster for Internet routers. Lecture Notes in Computer Science, 2228:399–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2228/22280399.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2228/22280399.
pdf.

Kumar:2001:ITE

[1610] Anurag Kumar. Intelligent traffic engineering of internets: Towards a model-based approach. Lecture Notes in Computer Science, 2228:414–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2228/22280414.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2228/22280414.
pdf.

Marsan:2001:PAD

[1611] Marco Ajmone Marsan, Marco Gribaudo, Michela Meo, and Matteo Sereno. Performance analysis of data services over GPRS. Lecture Notes in Computer Science, 2228:425–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2228/22280425.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2228/22280425.
pdf.

Anonymous:2001:AIam

[1612] Anonymous. Author index. Lecture Notes in Computer Science, 2228:437– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2228/2228auth.pdf.

Schnorr:2001:SBD

[1613] Claus Peter Schnorr. Security of blind discrete log signatures against interactive attacks. Lecture Notes in Computer Science, 2229:1–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2229/22290001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2229/22290001.
pdf.

Liu:2001:IIM

[1614] Dongxi Liu, Xiaoyong Li, and Yingcai Bai. An intelligent intruder model for security protocol analysis. Lecture Notes in Computer Science, 2229:13– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2229/22290013.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2229/22290013.
pdf.

Gupta:2001:PPG

[1615] Kishan Chand Gupta and Subhamoy Maitra. Primitive polynomials over GF(2) — A cryptologic approach. Lecture Notes in Computer Science, 2229: 23–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2229/22290023.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2229/22290023.
pdf.

Yang:2001:USO

[1616] Bo Yang, Shixiong Zhu, and Yumin Wang. Unconditionally-secure oblivious transfer. Lecture Notes in Computer Science, 2229:35–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2229/22290035.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2229/22290035.
pdf.

Chang:2001:CIU

[1617] Chin-Chen Chang and Iuon-Chang Lin. Cryptanalysis of the improved user efficient blind signatures. Lecture Notes in Computer Science, 2229:42– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2229/22290042.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2229/22290042.
pdf.

Chang:2001:TFG

[1618] Chin-Chen Chang and Kuo-Feng Hwang. Towards the forgery of a group signature without knowing the group center’s secret. Lecture Notes in Computer Science, 2229:47– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2229/22290047.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2229/22290047.
pdf.

Bo:2001:EID

[1619] Xiaochen Bo, Lincheng Shen, and Wensen Chang. Evaluation of the image degradation for a typical watermarking algorithm in the block-DCT domain. Lecture Notes in Computer Science, 2229:52–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2229/22290052.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2229/22290052.
pdf.

Aoki:2001:CWA

[1620] Kazumaro Aoki, Fumitaka Hoshino, and Tetsutaro Kobayashi. A cyclic window algorithm for ECC defined over extension fields. Lecture Notes in Computer Science, 2229:62–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2229/22290062.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2229/22290062.
pdf.

Zhang:2001:FSM

[1621] Fangguo Zhang, Futai Zhang, and Yumin Wang. Fast scalar multiplication on the Jacobian of a family of hyperelliptic curves. Lecture Notes in Computer Science, 2229:74– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2229/22290074.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2229/22290074.
pdf.

Ye:2001:ATD

[1622] Dingfeng Ye, Junhui Yang, Zongduo Dai, and Haiwen Ou. Attacks on two digital signature schemes based on error correcting codes. Lecture Notes in Computer Science, 2229:84– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2229/22290084.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2229/22290084.
pdf.

Matsuura:2001:DDO

[1623] Kanta Matsuura. A derivative of digital objects and estimation of default risks in electronic commerce. Lecture Notes in Computer Science, 2229:90– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2229/22290090.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2229/22290090.
pdf.

Shim:2001:NAS

[1624] Young-Chul Shim. A new approach for secure multicast routing in a large scale network. Lecture Notes in Computer Science, 2229:95–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2229/22290095.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2229/22290095.
                                        pdf.

Kim:2001:TLS

[1625] Hee-Wan Kim, Hae-Kyung Rhee, Tai M. Chung, Young Ik Eom, and Ung-Mo Kim. A transaction lengthsensitive protocol based on altruistic locking for multilevel secure database systems. Lecture Notes in Computer Science, 2229:107–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2229/22290107.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2229/22290107.
                                        pdf.

Cho:2001:DUR

[1626] Sungbaek Cho and Zbigniew Ciechanowicz. Dealing with uncertainties in risk analysis using belief functions. Lecture Notes in Computer Science, 2229:119– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2229/22290119.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2229/22290119.
                                        pdf.

Hitchens:2001:RXD

[1627] Michael Hitchens and Vijay Varadharajan. RBAC for XML document stores. Lecture Notes in Computer Science, 2229:131–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2229/22290131.htm;   http:

//link.springer-ny.com/link/service/
series/0558/papers/2229/22290131.
pdf.

Zhang:2001:CIS

[1628] Xian-Mo Zhang and Josef Pieprzyk.

Cheating immune secret sharing. Lecture Notes in Computer Science, 2229: 144–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2229/22290144.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2229/22290144.
pdf.

Samid:2001:ESR

[1629] Gideon Samid. Encryption sticks (randomats). Lecture Notes in Computer Science, 2229:150–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2229/22290150.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2229/22290150.pdf.

Song:2001:ANL

[1630] Zhimin Song and Sihan Qing. Applying NCP logic to the analysis of SSL 3.0. Lecture Notes in Computer Science, 2229:155–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2229/22290155.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2229/22290155.
pdf.

Herwono:2001:PWI

[1631] Ian Herwono and Ingo Liebhardt. Performance of WTLS and its impact on an M-commerce transaction. Lecture Notes in Computer Science, 2229:167– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2229/22290167.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2229/22290167.
                                        pdf.

Ribeiro:2001:EOS

[1632] Carlos Ribeiro, André Zúquete, and Paulo Ferreira. Enforcing obligation with security monitors. Lecture Notes in Computer Science, 2229:172– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2229/22290172.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2229/22290172.
                                        pdf.

Ning:2001:ESI

[1633] Peng Ning and Yiqun Lisa Yin. Efficient software implementation for finite field multiplication in normal basis. Lecture Notes in Computer Science, 2229:177–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2229/22290177.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2229/22290177.
                                        pdf.

Zhou:2001:PLI

[1634] Jianying Zhou and Chunfu Tan. Playing lottery on the Internet. Lecture Notes in Computer Science, 2229:189– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2229/22290189.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2229/22290189.
pdf.

Bao:2001:PPT

[1635] Feng Bao and Robert Deng. Privacy protection for transactions of digital goods. Lecture Notes in Computer Science, 2229:202–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2229/22290202.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2229/22290202.
pdf.

Xu:2001:ECA

[1636] Jie lü Xu, Han liang Xu, Yan Wang, and Shu-Wang Lü. Equivalent characterizations and applications of multi-output correlation-immune Boolean functions. Lecture Notes in Computer Science, 2229:214–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2229/22290214.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2229/22290214.
pdf.

Wang:2001:TUR

[1637] Guilin Wang, Sihan Qing, Mingsheng Wang, and Zhanfei Zhou. Threshold undeniable RSA signature scheme. Lecture Notes in Computer Science, 2229: 221–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2229/22290221.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2229/22290221.
                                        pdf.

Hwang:2001:TSB

[1638] Min-Shiang Hwang, Cheng-Chi Lee, and Yuan-Liang Tang. Two simple batch verifying multiple digital signatures. Lecture Notes in Computer Science, 2229:233–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2229/22290233.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2229/22290233.
                                        pdf.

He:2001:SAR

[1639] Yeping He and Sihan Qing. Square attack on reduced Camellia cipher. Lecture Notes in Computer Science, 2229: 238–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2229/22290238.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2229/22290238.
                                        pdf.

You:2001:GEC

[1640] Lin You, Yi Xian Yang, and Chun Qi Zhang. Generalization of elliptic curve digital signature schemes. Lecture Notes in Computer Science, 2229:246– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2229/22290246.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2229/22290246.
pdf.

Crispo:2001:RAA

[1641] Bruno Crispo and Giancarlo Ruffo.

Reasoning about accountability within delegation. Lecture Notes in Computer Science, 2229:251–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2229/22290251.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2229/22290251.pdf.

Pan:2001:NDH

[1642] Gang Pan, Yijun Wu, and Zhaohui Wu. A novel data hiding method for two-color images. Lecture Notes in Computer Science, 2229:261–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2229/22290261.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2229/22290261.
pdf.

Lee:2001:ISP

[1643] C.-H. Lee, X. Deng, and H. Zhu. An identification scheme provably secure against reset attack. Lecture Notes in Computer Science, 2229:271– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2229/22290271.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2229/22290271.
pdf.

Kunnari:2001:ESI

[1644] Sanna Kunnari. Estimating the scalability of the Internet key exchange. Lecture Notes in Computer Science, 2229: 280–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2229/22290280.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2229/22290280.
pdf.

Kuninobu:2001:EIF

[1645] Shigeta Kuninobu, Yoshiaki Takata, Hiroyuki Seki, and Katsuro Inoue. An efficient information flow analysis of recursive programs based on a lattice model of security classes. Lecture Notes in Computer Science, 2229:292– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2229/22290292.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2229/22290292.
pdf.

Ye:2001:DDS

[1646] Baoqing Ye. Defeating denial-of-service attacks on the Internet. Lecture Notes in Computer Science, 2229: 304–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2229/22290304.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2229/22290304.
pdf.

Zhang:2001:RBA

[1647] Dianlong Zhang, Harald Lukhaub, and Werner Zorn. A role-based access control model and implementation for data-centric enterprise applications. Lecture Notes in Computer Science, 2229:316–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2229/22290316.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2229/22290316.
pdf.

Permpoontanalarp:2001:UMV

[1648] Yongyuth Permpoontanalarp and Chaiwat Rujimethabhas. A unified methodology for verification and synthesis of firewall configurations. Lecture Notes in Computer Science, 2229: 328–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2229/22290328.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2229/22290328.
pdf.

Chen:2001:QND

[1649] Yan Chen, Adam Bargteil, David Bindel, Randy H. Katz, and John Kubiatowicz. Quantifying network denial of service: a location service case study. Lecture Notes in Computer Science, 2229:340–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2229/22290340.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2229/22290340.
                                        pdf.

Nieto:2001:PKC

[1650] Juan Manuel González Nieto, Colin Boyd, and Ed Dawson. A public key cryptosystem based on the subgroup membership problem. Lecture Notes in Computer Science, 2229:352– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2229/22290352.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2229/22290352.
                                        pdf.

Hong:2001:NSM

[1651] Ki-Yoong Hong and Chul Kim. On a network security model for the secure information flow on multilevel secure network. Lecture Notes in Computer Science, 2229:364–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2229/22290364.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2229/22290364.
pdf.

Luo:2001:NRB

[1652] Wenjian Luo, Xianbin Cao, and Xufa Wang. NIDS research based on artificial immunology. Lecture Notes in Computer Science, 2229:371–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2229/22290371.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2229/22290371.
pdf.

Canovas:2001:APA

[1653] Oscar Cánovas and Antonio F.

Gómez. AMBAR protocol: Access management based on authorization reduction. Lecture Notes in Computer Science, 2229:376–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2229/22290376.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2229/22290376.
pdf.

Zou:2001:CRT

[1654] Xukai Zou, Byrav Ramamurthy, and Spyros S. Magliveras. Chinese remainder theorem based hierarchical access control for secure group communication. Lecture Notes in Computer Science, 2229:381–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2229/22290381.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2229/22290381.
                                        pdf.

Wang:2001:DMA

[1655] Yan Wang and Kian-Lee Tan. Dispatching mobile agents with secure routes in parallel. Lecture Notes in Computer Science, 2229:386–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2229/22290386.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2229/22290386.
                                        pdf.

Chen:2001:TSS

[1656] Yu Chen, Qian Fang, Zhihui Du, Zhenchun Huang, and Sanli Li. TH-SMS: Security management system in advanced computational infrastructure. Lecture Notes in Computer Science, 2229:398–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2229/22290398.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2229/22290398.pdf.

Lang:2001:CMS

[1657] Ulrich Lang, Dieter Gollmann, and Rudolf Schreiner. Cryptography and middleware security. Lecture Notes in Computer Science, 2229:408–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2229/22290408.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2229/22290408.
pdf.

Zeng:2001:CHR

[1658] Kencheng Zeng, Chung-Huang Yang, and T. R. N. Rao. Cryptanalysis of the Hwang-Rao secret errorcorrecting code schemes. Lecture Notes in Computer Science, 2229:419– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2229/22290419.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2229/22290419.
pdf.

Disson:2001:RBM

[1659] Eric Disson, Danielle Boulanger, and Gilles Dubois. A role-based model for access control in database federations. Lecture Notes in Computer Science, 2229:429–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2229/22290429.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2229/22290429.
pdf.

Li:2001:UID

[1660] Hongpei Li, Lianli Chang, and Xinmei Wang. A useful intrusion detection system prototype to monitor multiprocesses based on system calls. Lecture Notes in Computer Science, 2229: 441–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2229/22290441.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2229/22290441.
                                        pdf.

Park:2001:DNP

[1661] Hee-Un Park and Im-Yeong Lee. A digital nominative proxy signature scheme for mobile communication. Lecture Notes in Computer Science, 2229:451– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2229/22290451.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2229/22290451.
                                        pdf.

Yi:2001:HSM

[1662] Mi Ra Yi and Tae Ho Cho. Hierarchical simulation model with animation for large network security. Lecture Notes in Computer Science, 2229: 456–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2229/22290456.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2229/22290456.
                                        pdf.

Maitland:2001:FEC

[1663] Greg Maitland and Colin Boyd. Fair electronic cash based on a group signature scheme. Lecture Notes in Computer Science, 2229:461–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2229/22290461.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2229/22290461.
pdf.

Wu:2001:FED

[1664] Chuan-Kun Wu and Vijay Varadharajan. Fair exchange of digital signatures with offline trusted third party. Lecture Notes in Computer Science, 2229: 466–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2229/22290466.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2229/22290466.
pdf.

Park:2001:STC

[1665] Jong Sou Park, Jang-Se Lee, Hwan Kuk Kim, Jeong-Rye Jeong, Dong-Bok Yeom, and Sung-Do Chi. SE-CUSIM: a tool for the cyberattack simulation. Lecture Notes in Computer Science, 2229:471–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2229/22290471.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2229/22290471.
pdf.

Li:2001:NSA

[1666] Yifa Li. A new semantics of authentication logic. Lecture Notes in Computer Science, 2229:476–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2229/22290476.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2229/22290476.
                                        pdf.

Park:2001:RFW

[1667] Ji Hwan Park, Sook Ee Jeong, and Chang Soo Kim. Robust and fragile watermarking techniques for documents using bi-directional diagonal profiles. Lecture Notes in Computer Science, 2229:483–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2229/22290483.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2229/22290483.
                                        pdf.

Gram:2001:ROS

[1668] Seenil Gram. Redundancy, obscurity, self-containment & independence. Lecture Notes in Computer Science, 2229: 495–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2229/22290495.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2229/22290495.
                                        pdf.

Anonymous:2001:AIan

[1669] Anonymous. Author index. Lecture Notes in Computer Science, 2229:503– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2229/2229auth.pdf.

Arts:2001:MCM

[1670] T. Arts, P. Bovendeerd, A. van der Toorn, L. Geerts, R. Kerckhoffs, and F. Prinzen. Modules in cardiac modeling: Mechanics, circulation, and depolarization wave. Lecture Notes in Computer Science, 2230: 1–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2230/22300001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2230/22300001.
pdf.

Mari:2001:GMH

[1671] Jean-Luc Mari, Laurent Astart, and Jean Sequeira. Geometrical modeling of the heart and its main vessels. Lecture Notes in Computer Science, 2230: 10–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2230/22300010.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2230/22300010.
pdf.

Veistera:2001:RBE

[1672] Henri Veisterä and Jyrki Lötjönen.

Reconstructing 3D boundary element heart models from 2D biplane fluoroscopy. Lecture Notes in Computer Science, 2230:17–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2230/22300017.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2230/22300017.
                                        pdf.

Dydenko:2001:ISE

[1673] Igor Dydenko, Denis Friboulet, and Isabelle E. Magnin. Introducing spectral estimation for boundary detection in echographic radiofrequency images. Lecture Notes in Computer Science, 2230:24–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2230/22300024.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2230/22300024.
                                        pdf.

Mourad:2001:GMF

[1674] Ayman Mourad, Luc Biard, Denis Caillerie, Pierre-Simon Jouk, Annie Raoult, Nicolas Szafran, and Yves Usson. Geometrical modelling of the fibre organization in the human left ventricle. Lecture Notes in Computer Science, 2230: 32–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2230/22300032.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2230/22300032.
                                        pdf.

Horacek:2001:CMH

[1675] B. Milan Horáček, Kim Simelius, Rok Hren, and Jukka Nenonen. Challenges in modelling human heart’s total excitation. Lecture Notes in Computer Science, 2230:39–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2230/22300039.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2230/22300039.
pdf.

Dhooge:2001:TDU

[1676] Jan D’hooge, Fadi Jamal, Bart Bijnens, Jan Thoen, Frans Van de Werf, George R. Sutherland, and Paul Suetens. Two-dimensional ultrasonic strain rate measurement of the human heart in vivo. Lecture Notes in Computer Science, 2230:47– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2230/22300047.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2230/22300047.
pdf.

Moreau:2001:DFE

[1677] Valérie Moreau, Laurent D. Cohen, and Denis Pellerin. Deformation field estimation for the cardiac wall using Doppler tissue imaging. Lecture Notes in Computer Science, 2230:53– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2230/22300053.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2230/22300053.
pdf.

Allouche:2001:NKM

[1678] Cyril F. Allouche, S. Makam, Nicholas Ayache, and Hervé Delingette. A new kinetic modeling scheme for the human left ventricle wall motion with MR-tagging imaging. Lecture Notes in Computer Science, 2230:61– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2230/22300061.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2230/22300061.
pdf.

Clarysse:2001:IQA

[1679] Patrick Clarysse, Pierre Croisille, Luc Bracoud, and Isabelle E. Magnin. Integrated quantitative analysis of tagged magnetic resonance images. Lecture Notes in Computer Science, 2230:69– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2230/22300069.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2230/22300069.
pdf.

McVeigh:2001:MVW

[1680] Elliot McVeigh, Owen Faris, Dan Ennis, Patrick Helm, and Frank Evans. Measurement of ventricular wall motion, epicardial electrical mapping, and myocardial fiber angles in the same heart. Lecture Notes in Computer Science, 2230:76– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2230/22300076.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2230/22300076.
pdf.

Makela:2001:DMB

[1681] Timo Mäkelä, Quoc-Cuong Pham, Patrick Clarysse, Jyrki Lötjönen, Kirsi Lauerma, Helena Hänninen, Juhani Knuuti, Toivo Katila, and Isabelle E. Magnin. A 3-D modelbased approach for the PET-functional and MR-anatomical cardiac imaging data fusion. Lecture Notes in Computer Science, 2230:83–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2230/22300083.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2230/22300083.
pdf.

Frouin:2001:RSF

[1682] Frédérique Frouin, Paté Boubacar, Vincent Frouin, Alain De Cesare, Andrew Todd-Pokropek, Pascal Merlet, and Alain Herment. 3D regularisation and segmentation of factor volumes to process PET H_{2}O myocardial perfusion studies. Lecture Notes in Computer Science, 2230:91– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2230/22300091.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2230/22300091.
pdf.

Nahrendorf:2001:VAR

[1683] M. Nahrendorf, K.-H. Hiller, K. Hu, C. Waller, F. Wiesmann, J. Ruff, G. Ertl, A. Haase, and W. R. Bauer. In vivo assessment of rat hearts with and without myocardial infarction by cine NMR — comparison of the NMR method to invasive techniques and application to intervention studies. Lecture Notes in Computer Science, 2230: 97–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2230/22300097.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2230/22300097.
pdf.

Muller:2001:DSA

[1684] Chantal Muller, Michèle Rombaut, and Marc Janier. Dempster Shafer approach for high level data fusion applied to the assessment of myocardial viability. Lecture Notes in Computer Science, 2230:104–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2230/22300104.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2230/22300104.
pdf.

McCulloch:2001:ECM

[1685] Andrew D. McCulloch, Derrick Sung, Mary Ellen Thomas, and Anushka Michailova. Experimental and computational modeling of cardiac electromechanical coupling. Lecture Notes in Computer Science, 2230:113–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2230/22300113.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2230/22300113.
pdf.

Ayache:2001:TMB

[1686] N. Ayache, D. Chapelle, F. Clément, Y. Coudière, H. Delingette, J. A. Désidéri, M. Sermesant, M. Sorine, and José M. Urquiza. Towards modelbased estimation of the cardiac electromechanical activity from ECG signals and ultrasound images. Lecture Notes in Computer Science, 2230:120– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2230/22300120.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2230/22300120.
pdf.

Chapelle:2001:PBM

[1687] Dominique Chapelle, Frédérique Clément, Frank Génot, Patrick Le Tallec, Michel Sorine, and José M. Urquiza. A physiologically-based model for the active cardiac muscle contraction. Lecture Notes in Computer Science, 2230: 128–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2230/22300128.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2230/22300128.
pdf.

Claus:2001:PST

[1688] Piet Claus, Bart Bijnens, Frank Weidemann, Christoph Dommke, Virginie Bito, Frank Heinzel, Karin Sipido, Ivan De Scheerder, Frank E. Rademakers, and George R. Sutherland. Post-systolic thickening in ischaemic myocardium: a simple mathematical model for simulating regional deformation. Lecture Notes in Computer Science, 2230:134–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2230/22300134.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2230/22300134.
pdf.

Simelius:2001:SAP

[1689] Kim Simelius, Jukka Nenonen, and B. Milan Horáček. Simulation of anisotropic propagation in the myocardium with a hybrid bidomain model. Lecture Notes in Computer Science, 2230:140–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2230/22300140.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2230/22300140.pdf.

Tilg:2001:IEF

[1690] Bernhard Tilg, Robert Modre, Gerald Fischer, Friedrich Hanser, Bernd Messnarz, and Franz Xaver Roithinger. Imaging of electrical function within the human atrium and ventricle from paced ECG mapping data. Lecture Notes in Computer Science, 2230:148– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2230/22300148.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2230/22300148.
pdf.

Anonymous:2001:AIao

[1691] Anonymous. Author index. Lecture Notes in Computer Science, 2230:157– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2230/2230auth.pdf.

Pasetti:2001:IC

[1692] Alessandro Pasetti. 1 introduction and context. Lecture Notes in Computer Science, 2231:1–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2231/22310001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2231/22310001.
pdf.

Pasetti:2001:AOC

[1693] Alessandro Pasetti. 2 attitude and orbit control systems (AOCS). Lecture Notes in Computer Science, 2231:17– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2231/22310017.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2231/22310017.
pdf.

Pasetti:2001:SF

[1694] Alessandro Pasetti. 3 software frameworks. Lecture Notes in Computer Science, 2231:29–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2231/22310029.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2231/22310029.pdf.

Pasetti:2001:FIC

[1695] Alessandro Pasetti. 4 framelets and implementation cases. Lecture Notes in Computer Science, 2231:49–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2231/22310049.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2231/22310049.
pdf.

Pasetti:2001:FS

[1696] Alessandro Pasetti. 5 framework specification. Lecture Notes in Computer Science, 2231:69–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2231/22310069.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2231/22310069.pdf.

Pasetti:2001:FD

[1697] Alessandro Pasetti. 6 framework design. Lecture Notes in Computer Science, 2231:79–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2231/22310079.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2231/22310079.pdf.

Pasetti:2001:UP

[1698] Alessandro Pasetti. 7 the user’s perspective. Lecture Notes in Computer Science, 2231:99–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2231/22310099.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2231/22310099.pdf.

Pasetti:2001:GSA

[1699] Alessandro Pasetti. 8 general structure of the AOCS framework. Lecture Notes in Computer Science, 2231:111– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2231/22310111.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2231/22310111.
pdf.

Pasetti:2001:GDP

[1700] Alessandro Pasetti. 9 general design principles. Lecture Notes in Computer Science, 2231:131–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2231/22310131.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2231/22310131.
pdf.

Pasetti:2001:SMF

[1701] Alessandro Pasetti. 10 the system management framelet. Lecture Notes in Computer Science, 2231:141–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2231/22310141.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2231/22310141.
pdf.

Pasetti:2001:OMF

[1702] Alessandro Pasetti. 11 the object monitoring framelet. Lecture Notes in Computer Science, 2231:147–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2231/22310147.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2231/22310147.
pdf.

Pasetti:2001:OMM

[1703] Alessandro Pasetti. 12 the operational mode management framelet. Lecture Notes in Computer Science, 2231:159– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2231/22310159.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2231/22310159.
pdf.

Pasetti:2001:ICF

[1704] Alessandro Pasetti. 13 the intercomponent communication framelet. Lecture Notes in Computer Science, 2231: 167–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

bibs/2231/22310167.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2231/22310167.
pdf.

Pasetti:2001:SDP

[1705] Alessandro Pasetti. 14 the sequential data processing framelet. Lecture Notes in Computer Science, 2231:183– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2231/22310183.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2231/22310183.
pdf.

Pasetti:2001:AUF

[1706] Alessandro Pasetti. 15 the AOCS unit framelet. Lecture Notes in Computer Science, 2231:193–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2231/22310193.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2231/22310193.pdf.

Pasetti:2001:RMF

[1707] Alessandro Pasetti. 16 the reconfiguration management framelet. Lecture Notes in Computer Science, 2231:209– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2231/22310209.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2231/22310209.
pdf.

Pasetti:2001:MMF

[1708] Alessandro Pasetti. 17 the manoeuvre management framelet. Lecture Notes in Computer Science, 2231:217– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2231/22310217.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2231/22310217.
                                        pdf.

Pasetti:2001:FDM

[1709] Alessandro Pasetti. 18 the failure detection management framelet. Lecture Notes in Computer Science, 2231:223– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2231/22310223.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2231/22310223.
                                        pdf.

Pasetti:2001:FRM

[1710] Alessandro Pasetti. 19 the failure recovery management framelet. Lecture Notes in Computer Science, 2231:233– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2231/22310233.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2231/22310233.
                                        pdf.

Pasetti:2001:TMFa

[1711] Alessandro Pasetti. 20 the telecommand management framelet. Lecture Notes in Computer Science, 2231:243– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2231/22310243.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2231/22310243.
pdf.

Pasetti:2001:TMFb

[1712] Alessandro Pasetti. 21 the telemetry management framelet. Lecture Notes in Computer Science, 2231:253– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2231/22310253.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2231/22310253.
pdf.

Pasetti:2001:CMF

[1713] Alessandro Pasetti. 22 the controller management framelet. Lecture Notes in Computer Science, 2231:263– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2231/22310263.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2231/22310263.
pdf.

Pasetti:2001:FIP

[1714] Alessandro Pasetti. 23 the framework instantiation process. Lecture Notes in Computer Science, 2231:271– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2231/22310271.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2231/22310271.
                                        pdf.

Pasetti:2001:A

[1715] Alessandro Pasetti. Appendix. Lecture Notes in Computer Science, 2231: 279–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2231/22310279.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2231/22310279.
                                        pdf.

Pasetti:2001:R

[1716] Alessandro Pasetti. References. Lecture Notes in Computer Science, 2231: 285–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2231/22310285.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2231/22310285.
                                        pdf.

Anonymous:2001:Iq

[1717] Anonymous. Index. Lecture Notes in Computer Science, 2231:291–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2231/2231indx.pdf.

Lazcano:2001:PBS

[1718] Amaia Lazcano and Gustavo Alonso. Process based E-services. Lecture Notes in Computer Science, 2232:1– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2232/22320001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2232/22320001.
pdf.

Heuser:2001:DRM

[1719] Stephan Heuser. Digital rights management — dealmaker for E-business? Lecture Notes in Computer Science, 2232:11–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2232/22320011.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2232/22320011.
pdf.

Buchmann:2001:SNW

[1720] Alejandro Buchmann. E-services: The next wave of Internet-based applications. Lecture Notes in Computer Science, 2232:13–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2232/22320013.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2232/22320013.pdf.

Garcia-Reinoso:2001:NCC

[1721] Jaime García-Reinoso, Javier Vales-Alonso, Francisco J. González-Castaño, Luis Anido-Rifón, and Pedro S. Rodríguez-Hernández. A new Mcommerce concept: m-mall. Lecture Notes in Computer Science, 2232:14– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2232/22320014.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2232/22320014.
                                        pdf.

dosSantos:2001:BCS

[1722] Simone C. dos Santos, Sérgio Angelim, and Silvio R. L. Meira. Building comparison-shopping brokers on the Web. Lecture Notes in Computer Science, 2232:26–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2232/22320026.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2232/22320026.
                                        pdf.

Piccinelli:2001:TMS

[1723] Giacomo Piccinelli, Cesare Stefanelli, and David Trastour. Trusted mediation for E-service provision in electronic marketplaces. Lecture Notes in Computer Science, 2232:39–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2232/22320039.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2232/22320039.
                                        pdf.

Schummer:2001:GMS

[1724] Till Schümmer. GAMA-mall — shopping in communities. Lecture Notes in Computer Science, 2232:51– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2232/22320051.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2232/22320051.
pdf.

Eymann:2001:MMF

[1725] Torsten Eymann. Markets without makers — A framework for decentralized economic coordination in multiagent systems. Lecture Notes in Computer Science, 2232:63–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2232/22320063.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2232/22320063.
pdf.

Golle:2001:ISP

[1726] Philippe Golle, Kevin Leyton-Brown, Ilya Mironov, and Mark Lillibridge. Incentives for sharing in peerto-peer networks. Lecture Notes in Computer Science, 2232:75–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2232/22320075.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2232/22320075.
pdf.

Wrona:2001:MPS

[1727] Konrad Wrona, Marko Schuba, and Guido Zavagli. Mobile payments — state of the art and open problems. Lecture Notes in Computer Science, 2232: 88–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2232/22320088.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2232/22320088.
                                        pdf.

Vogt:2001:USC

[1728] Holger Vogt, Henning Pagnia, and Felix C. Gärtner. Using smart cards for fair exchange. Lecture Notes in Computer Science, 2232:101–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2232/22320101.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2232/22320101.
                                        pdf.

Buttyan:2001:REF

[1729] Levente Buttyán and Jean-Pierre Hubaux. Rational exchange — A formal model based on game theory. Lecture Notes in Computer Science, 2232: 114–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2232/22320114.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2232/22320114.
                                        pdf.

Kugler:2001:EPP

[1730] Dennis Kügler. Enabling privacy protection in E-commerce applications. Lecture Notes in Computer Science, 2232:127–??, 2001. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2232/22320127.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2232/22320127.
pdf.

Csirik:2001:FFS

[1731] János A. Csirik, Michael L. Littman, Satinder Singh, and Peter Stone. FAucS: An FCC spectrum auction simulator for autonomous bidding agents. Lecture Notes in Computer Science, 2232:139–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2232/22320139.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2232/22320139.
pdf.

Byde:2001:DPM

[1732] Andrew Byde. A dynamic programming model for algorithm design in simultaneous auctions. Lecture Notes in Computer Science, 2232:152–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2232/22320152.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2232/22320152.
pdf.

Aberg:2001:UML

[1733] Johan Aberg, Nahid Shahmehri, and Dennis Maciuszek. User modelling for live help systems. Lecture Notes in Computer Science, 2232:164–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2232/22320164.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2232/22320164.
                                        pdf.

Adomavicius:2001:MRS

[1734] Gediminas Adomavicius and Alexander Tuzhilin. Multidimensional recommender systems: a data warehousing approach. Lecture Notes in Computer Science, 2232:180–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2232/22320180.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2232/22320180.
                                        pdf.

Bartelt:2001:MCT

[1735] Andreas Bartelt and Winfried Lamersdorf. A multi-criteria taxonomy of business models in electronic commerce. Lecture Notes in Computer Science, 2232:193–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2232/22320193.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2232/22320193.
                                        pdf.

Thede:2001:IGD

[1736] Anke Thede, Albrecht Schmidt, and Christian Merz. Integration of goods delivery supervision into Ecommerce supply chain. Lecture Notes in Computer Science, 2232:206–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2232/22320206.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2232/22320206.
pdf.

Minsky:2001:SRI

[1737] Naftaly H. Minsky and Victoria Ungureanu. Scalable regulation of interenterprise electronic commerce. Lecture Notes in Computer Science, 2232: 219–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2232/22320219.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2232/22320219.
pdf.

Anonymous:2001:AIap

[1738] Anonymous. Author index. Lecture Notes in Computer Science, 2232:233– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2232/2232auth.pdf.

Henderson:2001:LUB

[1739] Tristan Henderson. Latency and user behaviour on a multiplayer game server. Lecture Notes in Computer Science, 2233:1–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2233/22330001.
htm;   http://link.springer-ny.

com/link/service/series/0558/papers/
2233/22330001.pdf.

Ratnasamy:2001:ALM

[1740] Sylvia Ratnasamy, Mark Handley, Richard Karp, and Scott Shenker. Application-level multicast using contentaddressable networks. Lecture Notes in Computer Science, 2233:14–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2233/22330014.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2233/22330014.
pdf.

Rowstron:2001:SDL

[1741] Antony Rowstron, Anne-Marie Kermarrec, Miguel Castro, and Peter Druschel. Scribe: The design of a large-scale event notification infrastructure. Lecture Notes in Computer Science, 2233:30–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2233/22330030.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2233/22330030.
pdf.

Ganesh:2001:SPP

[1742] Ayalvadi J. Ganesh, Anne-Marie Kermarrec, and Laurent Massoulié. Scamp: Peer-to-peer lightweight membership service for large-scale group communication. Lecture Notes in Computer Science, 2233:44–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2233/22330044.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2233/22330044.
pdf.

Widmer:2001:EFV

[1743] Jörg Widmer and Thomas Fuhrmann.

Extremum feedback for very large multicast groups. Lecture Notes in Computer Science, 2233:56–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2233/22330056.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2233/22330056.
pdf.

Mathy:2001:OTB

[1744] Laurent Mathy, Roberto Canonico, and David Hutchison. An overlay tree building control protocol. Lecture Notes in Computer Science, 2233:76– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2233/22330076.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2233/22330076.
pdf.

Azar:2001:MBA

[1745] Yossi Azar, Meir Feder, Eyal Lubetzky, Doron Rajwan, and Nadav Shulman. The multicast bandwidth advantage in serving a Web site. Lecture Notes in Computer Science, 2233:88– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2233/22330088.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2233/22330088.
                                        pdf.

Byers:2001:SPA

[1746] John Byers and Gu-In Kwon. STAIR: Practical AIMD multirate multicast congestion control. Lecture Notes in Computer Science, 2233:100–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2233/22330100.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2233/22330100.
                                        pdf.

Busson:2001:ITS

[1747] Anthony Busson, Jean Louis Rougier, and Daniel Kofman. Impact of tree structure on retransmission efficiency for TRACK. Lecture Notes in Computer Science, 2233:113–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2233/22330113.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2233/22330113.
                                        pdf.

Amir:2001:FAA

[1748] Yair Amir, Cristina Nita-Rotaru, and Jonathan R. Stanton. Framework for authentication and access control of client-server group communication systems. Lecture Notes in Computer Science, 2233:128–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2233/22330128.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2233/22330128.
pdf.

Wang:2001:SIM

[1749] Ning Wang and George Pavlou. Scalable IP multicast sender access control for bi-directional trees. Lecture Notes in Computer Science, 2233:141– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2233/22330141.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2233/22330141.
pdf.

Rafaeli:2001:EEP

[1750] Sandro Rafaeli, Laurent Mathy, and David Hutchison. EHBT: An efficient protocol for group key management. Lecture Notes in Computer Science, 2233:159–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2233/22330159.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2233/22330159.
pdf.

Fei:2001:AMI

[1751] Aiguo Fei, Junhong Cui, Mario Gerla, and Michalis Faloutsos. Aggregated multicast with inter-group tree sharing. Lecture Notes in Computer Science, 2233:172–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2233/22330172.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2233/22330172.
                                        pdf.

Adler:2001:TLI

[1752] Micah Adler, Tian Bu, Ramesh K. Sitaraman, and Don Towsley. Tree layout for internal network characterizations in multicast networks. Lecture Notes in Computer Science, 2233:189– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2233/22330189.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2233/22330189.
                                        pdf.

Anonymous:2001:AIaq

[1753] Anonymous. Author index. Lecture Notes in Computer Science, 2233:205– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2233/22330205.pdf.

Afsarmanesh:2001:PGV

[1754] Hamideh Afsarmanesh, Ersin Kaletas, and Louis O. Hertzberger. The potential of Grid, virtual laboratories and virtual organizations for bio-sciences. Lecture Notes in Computer Science, 2234:1–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2234/22340001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2234/22340001.
pdf.

Charron-Bost:2001:APF

[1755] Bernadette Charron-Bost. Agreement problems in fault-tolerant distributed systems. Lecture Notes in Computer Science, 2234:10–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2234/22340010.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2234/22340010.
pdf.

Grosky:2001:NSG

[1756] William I. Grosky and Rong Zhao. Negotiating the semantic gap: From feature maps to semantic landscapes. Lecture Notes in Computer Science, 2234: 33–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2234/22340033.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2234/22340033.
pdf.

Jenei:2001:IRB

[1757] Sándor Jenei. Inference in rulebased systems by interpolation and extrapolation revisited. Lecture Notes in Computer Science, 2234:53–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2234/22340053.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2234/22340053.
                                        pdf.

Kaklamanis:2001:RAW

[1758] Christos Kaklamanis. Recent advances in wavelength routing. Lecture Notes in Computer Science, 2234:58– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2234/22340058.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2234/22340058.
                                        pdf.

Laforenza:2001:MGC

[1759] Domenico Laforenza. From metacomputing to Grid computing: Evolution or revolution? Lecture Notes in Computer Science, 2234:73–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2234/22340073.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2234/22340073.
                                        pdf.

Lambert:2001:KBC

[1760] Simon Lambert. Knowledge-based control systems. Lecture Notes in Computer Science, 2234:75–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2234/22340075.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2234/22340075.
pdf.

vanLeeuwen:2001:BTL

[1761] Jan van Leeuwen and Jiří Wiedermann. Beyond the Turing limit: Evolving interactive systems. Lecture Notes in Computer Science, 2234:90– 109, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2234/22340090.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2234/22340090.
pdf.

Santoro:2001:DCA

[1762] Nicola Santoro. Distributed computations by autonomous mobile robots. Lecture Notes in Computer Science, 2234:110–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2234/22340110.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2234/22340110.
pdf.

Slobodova:2001:FVM

[1763] Anna Slobodová. Formal verification methods for industrial hardware design. Lecture Notes in Computer Science, 2234:116–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2234/22340116.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2234/22340116.pdf.

Watanabe:2001:HCC

[1764] Osamu Watanabe. How can computer science contribute to knowledge discovery? Lecture Notes in Computer Science, 2234:136–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2234/22340136.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2234/22340136.
                                        pdf.

Aho:2001:AIK

[1765] Isto Aho. On the approximability of interactive knapsack problems. Lecture Notes in Computer Science, 2234:152– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2234/22340152.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2234/22340152.
                                        pdf.

Argon:2001:MCC

[1766] Pablo Argón, Giorgio Delzanno, Supratik Mukhopadhyay, and Andreas Podelski. Model checking communication protocols. Lecture Notes in Computer Science, 2234:160–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2234/22340160.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2234/22340160.
                                        pdf.

Beran:2001:PDB

[1767] Martin Beran. Pipelined decomposable BSP computers. Lecture Notes in Computer Science, 2234:171–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2234/22340171.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2234/22340171.
pdf.

Bonner:2001:QVP

[1768] Richard Bonner, Rusins Freivalds, and Maksim Kravtsev. Quantum versus probabilistic one-way finite automata with counter. Lecture Notes in Computer Science, 2234:181–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2234/22340181.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2234/22340181.
pdf.

Brim:2001:HER

[1769] Lubos Brim, Ivana Cerná, Pavel Krcál, and Radek Pelánek. How to employ reverse search in distributed single source shortest paths. Lecture Notes in Computer Science, 2234:191– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2234/22340191.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2234/22340191.
pdf.

Brim:2001:MAS

[1770] Lubos Brim, David Gilbert, Jean-Marie Jacquet, and Mojmír Kretínský. Multi-agent systems as concurrent constraint processes. Lecture Notes in Computer Science, 2234:201–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2234/22340201.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2234/22340201.
                                        pdf.

Pasquale:2001:AOP

[1771] Adriano Di Pasquale and Enrico Nardelli. ADST: An order preserving scalable distributed data structure with constant access costs. Lecture Notes in Computer Science, 2234:211– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2234/22340211.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2234/22340211.
                                        pdf.

Fernau:2001:ALR

[1772] Henning Fernau. Approximative learning of regular languages. Lecture Notes in Computer Science, 2234:223– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2234/22340223.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2234/22340223.
                                        pdf.

Freivalds:2001:QFS

[1773] Rūsiņš Freivalds and Andreas Winter.

Quantum finite state transducers. Lecture Notes in Computer Science, 2234: 233–242, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2234/22340233.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2234/22340233.
pdf.

Galambos:2001:LDI

[1774] Leo Galambos. Lemmatizer for document information retrieval systems in JAVA. Lecture Notes in Computer Science, 2234:243–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2234/22340243.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2234/22340243.
pdf.

Gebala:2001:RPA

[1775] Maciej Gebala. The reconstruction of polyominoes from approximately orthogonal projections. Lecture Notes in Computer Science, 2234:253–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2234/22340253.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2234/22340253.
pdf.

Jayanti:2001:BLB

[1776] Prasad Jayanti, King Tan, Gregory Friedland, and Amir Katz. Bounding Lamport’s bakery algorithm. Lecture Notes in Computer Science, 2234:261– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2234/22340261.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2234/22340261.
                                        pdf.

Kocsor:2001:FIC

[1777] András Kocsor and János Csirik. Fast independent component analysis in kernel feature spaces. Lecture Notes in Computer Science, 2234:271–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2234/22340271.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2234/22340271.
                                        pdf.

Kralovic:2001:MVG

[1778] Rastislav Královič. On majority voting games in trees. Lecture Notes in Computer Science, 2234:282–291, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2234/22340282.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2234/22340282.
                                        pdf.

Kralovic:2001:TSC

[1779] Richard Královič. Time and space complexity of reversible pebbling. Lecture Notes in Computer Science, 2234:292– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2234/22340292.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2234/22340292.
pdf.

Plachetka:2001:HRS

[1780] Tomas Plachetka, Olaf Schmidt, and Frank Albracht. The HiQoS rendering system. Lecture Notes in Computer Science, 2234:304–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2234/22340304.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2234/22340304.
pdf.

Platek:2001:TWR

[1781] Martin Plátek. Two-way restarting automata and J-monotonicity. Lecture Notes in Computer Science, 2234:316– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2234/22340316.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2234/22340316.
pdf.

Sawa:2001:PHE

[1782] Zdenek Sawa and Petr Jancar. P - hardness of equivalence testing on finite-state processes. Lecture Notes in Computer Science, 2234:326–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2234/22340326.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2234/22340326.
pdf.

Pratt:2001:SGP

[1783] Vaughan R. Pratt. Software geography: Physical and economic aspects. Lecture Notes in Computer Science, 2234: 336–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2234/22340336.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2234/22340336.
pdf.

Anonymous:2001:AIar

[1784] Anonymous. Author index. Lecture Notes in Computer Science, 2234:347– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2234/2234auth.pdf.

Arroyo:2001:SBL

[1785] Fernando Arroyo, Angel V. Baranda, Juan Castellanos, Carmen Luengo, and Luis F. Mingo. Structures and biolanguage to simulate transition P systems on digital computers. Lecture Notes in Computer Science, 2235:1– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2235/22350001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2235/22350001.
pdf.

Banatre:2001:GCR

[1786] Jean-Pierre Banâtre, Pascal Fradet, and Daniel Le Métayer. Gamma and the chemical reaction model: Fifteen years after. Lecture Notes in Computer Science, 2235:17–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2235/22350017.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2235/22350017.
pdf.

Bottoni:2001:VMR

[1787] Paolo Bottoni, Bernd Meyer, and Francesco Parisi Presicce. Visual multiset rewriting: Applications to diagram parsing and reasoning. Lecture Notes in Computer Science, 2235:45– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2235/22350045.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2235/22350045.
pdf.

Csuhaj-Varju:2001:MA

[1788] Erzsébet Csuhaj-Varjú, Carlos Martín-Vide, and Victor Mitrana. Multiset automata. Lecture Notes in Computer Science, 2235:69–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2235/22350069.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2235/22350069.
                                        pdf.

Dassow:2001:PMI

[1789] Jürgen Dassow. Parikh mapping and iteration. Lecture Notes in Computer Science, 2235:85–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2235/22350085.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2235/22350085.
                                        pdf.

Dovier:2001:MCP

[1790] Agostino Dovier, Carla Piazza, and Gianfranco Rossi. Multiset constraints and P systems. Lecture Notes in Computer Science, 2235:103– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2235/22350103.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2235/22350103.
                                        pdf.

Kudlek:2001:TFM

[1791] Manfred Kudlek, Carlos Martín-Vide, and Gheorghe Păun. Toward a formal macroset theory. Lecture Notes in Computer Science, 2235:123–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2235/22350123.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2235/22350123.
pdf.

Kudlek:2001:NFG

[1792] Manfred Kudlek and Victor Mitrana.

Normal forms of grammars, finite automata, abstract families, and closure properties of multiset languages. Lecture Notes in Computer Science, 2235: 135–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2235/22350135.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2235/22350135.
pdf.

Lamperti:2001:MDS

[1793] Gianfranco Lamperti, Michele Melchiori, and Marina Zanella. On multisets in database systems. Lecture Notes in Computer Science, 2235:147– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2235/22350147.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2235/22350147.
pdf.

Marcus:2001:TM

[1794] Solomon Marcus. Tolerance multisets.

Lecture Notes in Computer Science, 2235:217–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2235/22350217.htm;   http:
//link.springer-ny.com/link/service/

                                        series/0558/papers/2235/22350217.
                                        pdf.

Miyamoto:2001:FMT

[1795] Sadaaki Miyamoto. Fuzzy multisets and their generalizations. Lecture Notes in Computer Science, 2235:225– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2235/22350225.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2235/22350225.
                                        pdf.

Mutyam:2001:URS

[1796] Madhu Mutyam and Kamala Krithivasan. Universality results for some variants of P systems. Lecture Notes in Computer Science, 2235:237–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2235/22350237.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2235/22350237.
                                        pdf.

Nishida:2001:MKS

[1797] Taishin Yasunobu Nishida. Multiset and K-subset transforming systems. Lecture Notes in Computer Science, 2235:255–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2235/22350255.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2235/22350255.
                                        pdf.

Obtulowicz:2001:PSA

[1798] Adam Obtułowicz. On P systems with active membranes solving the integer factorization problem in a polynomial time. Lecture Notes in Computer Science, 2235:267–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2235/22350267.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2235/22350267.
pdf.

Reisig:2001:LTM

[1799] Wolfgang Reisig. The linear theory of multiset based dynamic systems. Lecture Notes in Computer Science, 2235: 287–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2235/22350287.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2235/22350287.
pdf.

Suzuki:2001:ALA

[1800] Yasuhiro Suzuki, Yoshi Fujiwara, Junji Takabayashi, and Hiroshi Tanaka. Artificial life applications of a class of P systems: Abstract rewriting systems on multisets. Lecture Notes in Computer Science, 2235:299–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2235/22350299.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2235/22350299.
pdf.

Syropoulos:2001:MM

[1801] Apostolos Syropoulos. Mathematics of multisets. Lecture Notes in Computer Science, 2235:347–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2235/22350347.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2235/22350347.
                                        pdf.

Anonymous:2001:AIas

[1802] Anonymous. Author index. Lecture Notes in Computer Science, 2235:359– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2235/2235auth.pdf.

Drira:2001:CI

[1803] K. Drira, A. Martelli, and T. Villemur. Chapter 1. introduction. Lecture Notes in Computer Science, 2236: 3–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2236/22360003.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2236/22360003.
                                        pdf.

Baurens:2001:CRD

[1804] B. Baurens, P. Chilaev, V. Krivtsov, and V. Volochinov. Chapter 2. research and development projects. Lecture Notes in Computer Science, 2236: 7–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2236/22360007.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2236/22360007.
pdf.

Avino:2001:CRE

[1805] D. Avino, R. Becchini, P. Chilaev, H. Follet, V. Krivtsov, A. Martelli, and V. Volochinov. Chapter 3. relevant existing practices. Lecture Notes in Computer Science, 2236:15– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2236/22360015.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2236/22360015.
pdf.

Becchini:2001:CM

[1806] R. Becchini, P. Chilaev, V. Krivtsov, I. Viglietti, and V. Volochinov. Chapter 4. middleware. Lecture Notes in Computer Science, 2236:41–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2236/22360041.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2236/22360041.
pdf.

Drira:2001:CPD

[1807] K. Drira, M. Molina, O. Nabuco, L. M. Rodriguez-Peralta, and T. Villemur. Chapter 5. product data and workflow management. Lecture Notes in Computer Science, 2236:107–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2236/22360107.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2236/22360107.
                                        pdf.

Baurens:2001:CC

[1808] B. Baurens, L. Costa, A. Dracinschi, S. Fdida, V. Roca, and R. Vida. Chapter 6. communications. Lecture Notes in Computer Science, 2236:153– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2236/22360153.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2236/22360153.
                                        pdf.

Baurens:2001:CG

[1809] B. Baurens. Chapter 7. groupware. Lecture Notes in Computer Science, 2236: 177–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2236/22360177.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2236/22360177.
                                        pdf.

Anonymous:2001:AAP

[1810] Anonymous. Appendix A1. product classification summary. Lecture Notes in Computer Science, 2236:263– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2236/22360263.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2236/22360263.
pdf.

Anonymous:2001:AAA

[1811] Anonymous. Appendix A2. acronyms and abbreviations. Lecture Notes in Computer Science, 2236:271–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2236/22360271.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2236/22360271.
pdf.

Anonymous:2001:Re

[1812] Anonymous. References. Lecture Notes in Computer Science, 2236:275– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2236/22360275.pdf.

Anonymous:2001:AIat

[1813] Anonymous. Author index. Lecture Notes in Computer Science, 2236:281– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2236/2236auth.pdf.

Colmerauer:2001:SMC

[1814] A. Colmerauer. Solving the multiplication constraint in several approximation spaces. Lecture Notes in Computer Science, 2237:1–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2237/22370001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2237/22370001.
pdf.

Kowalski:2001:LRD

[1815] R. Kowalski. Is logic really dead or only just sleeping? Lecture Notes in Computer Science, 2237:2– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2237/22370002.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2237/22370002.
pdf.

Cousot:2001:DSP

[1816] P. Cousot. Design of syntactic program transformations by abstract interpretation of semantic transformations. Lecture Notes in Computer Science, 2237: 4–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2237/22370004.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2237/22370004.
pdf.

Gupta:2001:TSC

[1817] A. Gupta. X-tegration — some cross-enterprise thoughts. Lecture Notes in Computer Science, 2237:6– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2237/22370006.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2237/22370006.
pdf.

Wielemaker:2001:BRL

[1818] J. Wielemaker. Building real-life applications with prolog. Lecture Notes in Computer Science, 2237:7– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2237/22370007.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2237/22370007.
pdf.

VillemontedelaClergerie:2001:NLT

[1819] É. Villemonte de la Clergerie. Natural language tabular parsing. Lecture Notes in Computer Science, 2237:8– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2237/22370008.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2237/22370008.
pdf.

Ueda:2001:CLC

[1820] K. Ueda. A close look at constraintbased concurrency. Lecture Notes in Computer Science, 2237:9–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2237/22370009.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2237/22370009.
pdf.

Subrahmanian:2001:PDL

[1821] V. S. Subrahmanian. Probabilistic databases and logic programming. Lecture Notes in Computer Science, 2237: 10–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2237/22370010.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2237/22370010.
pdf.

Castro:2001:UMM

[1822] L. F. Castro and V. S. Costa. Understanding memory management in prolog systems. Lecture Notes in Computer Science, 2237:11–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2237/22370011.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2237/22370011.
pdf.

Villaverde:2001:PPI

[1823] K. Villaverde, E. Pontelli, H. Guo, and G. Gupta. PALS: An orparallel implementation of prolog on Beowulf architectures. Lecture Notes in Computer Science, 2237:27–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2237/22370027.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2237/22370027.
pdf.

Rocha:2001:TEC

[1824] R. Rocha, F. Silva, and V. S. Costa.

On a tabling engine that can exploit or-parallelism. Lecture Notes in Computer Science, 2237:43–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2237/22370043.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2237/22370043.
pdf.

Beldiceanu:2001:RCO

[1825] N. Beldiceanu and M. Carlsson.

Revisiting the cardinality operator and introducing the cardinality-path constraint family. Lecture Notes in Computer Science, 2237:59–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2237/22370059.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2237/22370059.
pdf.

Holzbaur:2001:OCC

[1826] C. Holzbaur, M. García de la Banda, D. Jeffery, and P. J. Stuckey. Optimizing compilation of constraint handling rules. Lecture Notes in Computer Science, 2237:74–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2237/22370074.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2237/22370074.
pdf.

GarciadelaBanda:2001:BCS

[1827] M. García de la Banda, D. Jeffery, K. Marriott, N. Nethercote, P. J. Stuckey, and C. Holzbaur. Building constraint solvers with HAL. Lecture Notes in Computer Science, 2237:90– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2237/22370090.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2237/22370090.
                                        pdf.

Mazur:2001:PAW

[1828] N. Mazur, P. Ross, G. Janssens, and M. Bruynooghe. Practical aspects for a working compile time garbage collection system for mercury. Lecture Notes in Computer Science, 2237:105– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2237/22370105.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2237/22370105.
                                        pdf.

Howe:2001:PBF

[1829] J. M. Howe and A. King. Positive Boolean functions as multiheaded clauses. Lecture Notes in Computer Science, 2237:120–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2237/22370120.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2237/22370120.
                                        pdf.

Codish:2001:HPG

[1830] M. Codish, S. Genaim, H. Søndergaard, and P. J. Stuckey. Higher-precision groundness analysis. Lecture Notes in Computer Science, 2237:135–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2237/22370135.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2237/22370135.
pdf.

Guo:2001:SBC

[1831] H.-F. Guo, C. R. Ramakrishnan, and I. V. Ramakrishnan. Speculative beats conservative justification. Lecture Notes in Computer Science, 2237: 150–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2237/22370150.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2237/22370150.
pdf.

Basu:2001:LSB

[1832] S. Basu, M. Mukund, C. R. Ramakrishnan, I. V. Ramakrishnan, and R. Verma. Local and symbolic bisimulation using tabled constraint logic programming. Lecture Notes in Computer Science, 2237:166– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2237/22370166.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2237/22370166.
pdf.

Guo:2001:SSI

[1833] H.-F. Guo and G. Gupta. A simple scheme for implementing tabled logic programming systems based on dynamic reordering of alternatives. Lecture Notes in Computer Science, 2237: 181–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2237/22370181.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2237/22370181.
                                        pdf.

Lonc:2001:FPC

[1834] Z. Lonc and M. Truszczyński. Fixedparameter complexity of semantics for logic programs. Lecture Notes in Computer Science, 2237:197–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2237/22370197.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2237/22370197.
                                        pdf.

Denecker:2001:UWF

[1835] M. Denecker, N. Pelov, and M. Bruynooghe. Ultimate well-founded and stable semantics for logic programs with aggregates. Lecture Notes in Computer Science, 2237:212–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2237/22370212.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2237/22370212.pdf.

Kumar:2001:AFP

[1836] K. Narayan Kumar, C. R. Ramakrishnan, and S. A. Smolka. Alternating fixed points in Boolean equation systems as preferred stable models. Lecture Notes in Computer Science, 2237: 227–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2237/22370227.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2237/22370227.
pdf.

Erdem:2001:FTP

[1837] E. Erdem and V. Lifschitz. Fages’ theorem for programs with nested expressions. Lecture Notes in Computer Science, 2237:242–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2237/22370242.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2237/22370242.pdf.

Orejas:2001:SNL

[1838] F. Orejas, E. Pasarella, and E. Pino.

Semantics of normal logic programs with embedded implications. Lecture Notes in Computer Science, 2237:255– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2237/22370255.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2237/22370255.
pdf.

Medina:2001:MAL

[1839] J. Medina, M. Ojeda-Aciego, and P. Vojtás. A multi-adjoint logic approach to abductive reasoning. Lecture Notes in Computer Science, 2237:269– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2237/22370269.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2237/22370269.
pdf.

Drabent:2001:PCC

[1840] W. Drabent and M. Miłkowska. Proving correctness and completeness of normal programs — A declarative approach. Lecture Notes in Computer Science, 2237:284–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2237/22370284.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2237/22370284.
pdf.

Kaneiwa:2001:OSR

[1841] K. Kaneiwa and S. Tojo. An order-sorted resolution with implicitly negative sorts. Lecture Notes in Computer Science, 2237:300–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2237/22370300.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2237/22370300.
pdf.

Banbara:2001:LPF

[1842] M. Banbara, K.-S. Kang, T. Hirai, and N. Tamura. Logic programming in a fragment of intuitionistic temporal linear logic. Lecture Notes in Computer Science, 2237:315– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2237/22370315.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2237/22370315.
pdf.

Almendros-Jimenez:2001:CMF

[1843] J. M. Almendros-Jiménez, A. Becerra-Terón, and J. Sánchez-Hernández. A computational model for functional logic deductive databases. Lecture Notes in Computer Science, 2237:331– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2237/22370331.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2237/22370331.
pdf.

Greco:2001:LPA

[1844] G. Greco, S. Greco, and E. Zumpano.

A logic programming approach to the integration, repairing and querying of inconsistent databases. Lecture Notes in Computer Science, 2237:348– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2237/22370348.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2237/22370348.
                                        pdf.

Anonymous:2001:AIau

[1845] Anonymous. Author index. Lecture Notes in Computer Science, 2237:365– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2237/2237auth.pdf.

Eremin:2001:HBD

[1846] Andrew Eremin and Mark Wallace. Hybrid Benders decomposition algorithms in constraint logic programming. Lecture Notes in Computer Science, 2239:1–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2239/22390001.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2239/22390001.
                                        pdf.

Thorsteinsson:2001:BCH

[1847] Erlendur S. Thorsteinsson. Branchand-check: a hybrid framework integrating mixed integer programming and constraint logic programming. Lecture Notes in Computer Science, 2239: 16–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2239/22390016.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2239/22390016.
                                        pdf.

Abdennadher:2001:TIC

[1848] Slim Abdennadher and Christophe Rigotti. Towards inductive constraint solving. Lecture Notes in Computer Science, 2239:31–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2239/22390031.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2239/22390031.
pdf.

Epstein:2001:CLC

[1849] Susan L. Epstein and Eugene Freuder.

Collaborative learning for constraint solving. Lecture Notes in Computer Science, 2239:46–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2239/22390046.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2239/22390046.pdf.

Benoist:2001:TSC

[1850] Thierry Benoist, Eric Bourreau, Yves Caseau, and Benoît Rottembourg. Towards stochastic constraint programming: a study of online multi-choice knapsack with deadlines. Lecture Notes in Computer Science, 2239:61– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2239/22390061.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2239/22390061.
pdf.

Focacci:2001:GCF

[1851] Filippo Focacci and Michaela Milano. Global cut framework for removing symmetries. Lecture Notes in Computer Science, 2239:77–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2239/22390077.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2239/22390077.
                                        pdf.

Fahle:2001:SB

[1852] Torsten Fahle, Stefan Schamberger, and Meinolf Sellmann. Symmetry breaking. Lecture Notes in Computer Science, 2239:93–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2239/22390093.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2239/22390093.
                                        pdf.

Dubois:2001:NEC

[1853] Olivier Dubois and Gilles Dequen. The non-existence of (3, 1, 2)-conjugate orthogonal idempotent Latin square of order 10. Lecture Notes in Computer Science, 2239:108–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2239/22390108.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2239/22390108.pdf.

Aguirre:2001:RSB

[1854] Alfonso San Miguel Aguirre and Moshe Y. Vardi. Random 3-SAT and BDDs: The plot thickens further. Lecture Notes in Computer Science, 2239: 121–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2239/22390121.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2239/22390121.
pdf.

Bejar:2001:CSS

[1855] Ramón Béjar, Alba Cabiscol, Cèsar Fernàndez, Felip Manyà, and Carla Gomes. Capturing structure with satisfiability. Lecture Notes in Computer Science, 2239:137–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2239/22390137.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2239/22390137.pdf.

Zhang:2001:PTB

[1856] Weixiong Zhang. Phase transitions and backbones of 3-SAT and maximum 3-SAT. Lecture Notes in Computer Science, 2239:153–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2239/22390153.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2239/22390153.
pdf.

Mamoulis:2001:SNB

[1857] Nikos Mamoulis and Kostas Stergiou. Solving non-binary CSPs using the hidden variable encoding. Lecture Notes in Computer Science, 2239:168– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2239/22390168.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2239/22390168.
                                        pdf.

Pesant:2001:FAS

[1858] Gilles Pesant. A filtering algorithm for the stretch constraint. Lecture Notes in Computer Science, 2239:183– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2239/22390183.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2239/22390183.
                                        pdf.

Bockmayr:2001:NFP

[1859] Alexander Bockmayr, Nicolai Pisaruk, and Abderrahmane Aggoun. Network flow problems in constraint programming. Lecture Notes in Computer Science, 2239:196–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2239/22390196.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2239/22390196.
                                        pdf.

Beldiceanu:2001:PMC

[1860] Nicolas Beldiceanu. Pruning for the minimum constraint family and for the number of distinct values constraint family. Lecture Notes in Computer Science, 2239:211–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2239/22390211.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2239/22390211.
pdf.

Gent:2001:CPA

[1861] Ian P. Gent, Robert W. Irving, David F. Manlove, Patrick Prosser, and Barbara M. Smith. A constraint programming approach to the stable marriage problem. Lecture Notes in Computer Science, 2239:225– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2239/22390225.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2239/22390225.
pdf.

Choi:2001:CSR

[1862] Chiu Wo Choi, Martin Henz, and Ka Boon Ng. Components for state restoration in tree search. Lecture Notes in Computer Science, 2239:240– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2239/22390240.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2239/22390240.
                                        pdf.

Wolf:2001:ACH

[1863] Armin Wolf. Adaptive constraint handling with CHR in Java. Lecture Notes in Computer Science, 2239:256– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2239/22390256.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2239/22390256.
                                        pdf.

Silaghi:2001:CMA

[1864] Marius-Călin Silaghi, Djamila Sam-Haroud, and Boi Faltings. Consistency maintenance for ABT. Lecture Notes in Computer Science, 2239:271– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2239/22390271.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2239/22390271.
                                        pdf.

Delzanno:2001:CBV

[1865] Giorgio Delzanno and Tevfik Bultan. Constraint-based verification of client-server protocols. Lecture Notes in Computer Science, 2239:286– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2239/22390286.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2239/22390286.
                                        pdf.

Palamidessi:2001:TCC

[1866] Catuscia Palamidessi and Frank D. Valencia. A temporal concurrent constraint programming calculus. Lecture Notes in Computer Science, 2239:302– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2239/22390302.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2239/22390302.
pdf.

Meseguer:2001:LBN

[1867] Pedro Meseguer, Javier Larrosa, and Martì Sánchez. Lower bounds for non-binary constraint optimization problems. Lecture Notes in Computer Science, 2239:317–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2239/22390317.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2239/22390317.
pdf.

Regin:2001:NLB

[1868] Jean-Charles Régin, Thierry Petit, Christian Bessière, and Jean-François Puget. New lower bounds of constraint violations for overconstrained problems. Lecture Notes in Computer Science, 2239:332–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2239/22390332.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2239/22390332.
                                        pdf.

Dechter:2001:GSM

[1869] Rina Dechter, Kalev Kask, and Javier Larrosa. A general scheme for multiple lower bound computation in constraint optimization. Lecture Notes in Computer Science, 2239:346–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2239/22390346.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2239/22390346.
                                        pdf.

Marriott:2001:SDC

[1870] Kim Marriott, Peter Moulder, Peter J. Stuckey, and Alan Borning. Solving disjunctive constraints for interactive graphical applications. Lecture Notes in Computer Science, 2239:361– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2239/22390361.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2239/22390361.
                                        pdf.

Beldiceanu:2001:SGP

[1871] Nicolas Beldiceanu and Mats Carlsson. Sweep as a generic pruning technique applied to the non-overlapping rectangles constraint. Lecture Notes in Computer Science, 2239:377–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2239/22390377.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2239/22390377.
pdf.

Beldiceanu:2001:NOC

[1872] Nicolas Beldiceanu, Qi Guo, and Sven Thiel. Non-overlapping constraints between convex polytopes. Lecture Notes in Computer Science, 2239:392– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2239/22390392.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2239/22390392.
pdf.

Chen:2001:FMH

[1873] Hubie Chen, Carla Gomes, and Bart Selman. Formal models of heavy-tailed behavior in combinatorial search. Lecture Notes in Computer Science, 2239: 408–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2239/22390408.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2239/22390408.
pdf.

Armando:2001:PTL

[1874] Alessandro Armando, Felice Peccia, and Silvio Ranise. The phase transition of the linear inequalities problem. Lecture Notes in Computer Science, 2239: 422–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2239/22390422.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2239/22390422.
                                        pdf.

Kolaitis:2001:SPT

[1875] Phokion G. Kolaitis and Thomas Raffill. In search of a phase transition in the AC-matching problem. Lecture Notes in Computer Science, 2239:433– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2239/22390433.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2239/22390433.
                                        pdf.

Petit:2001:SFA

[1876] Thierry Petit, Jean-Charles Régin, and Christian Bessière. Specific filtering algorithms for over-constrained problems. Lecture Notes in Computer Science, 2239:451–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2239/22390451.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2239/22390451.pdf.

Meseguer:2001:SRD

[1877] Pedro Meseguer and Martì Sánchez. Specializing Russian doll search. Lecture Notes in Computer Science, 2239: 464–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2239/22390464.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2239/22390464.
pdf.

Swain:2001:CAP

[1878] Martin T. Swain and Graham J. L.

Kemp. A CLP approach to the protein side-chain placement problem. Lecture Notes in Computer Science, 2239:479– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2239/22390479.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2239/22390479.
pdf.

Backofen:2001:FCB

[1879] Rolf Backofen and Sebastian Will.

Fast, constraint-based threading of HPsequences to hydrophobic cores. Lecture Notes in Computer Science, 2239: 494–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2239/22390494.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2239/22390494.
pdf.

Henz:2001:OFP

[1880] Martin Henz, Edgar Tan, and Roland Yap. One flip per clock cycle. Lecture Notes in Computer Science, 2239: 509–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2239/22390509.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2239/22390509.
pdf.

Michel:2001:SCF

[1881] Claude Michel, Michel Rueher, and Yahia Lebbah. Solving constraints over floating-point numbers. Lecture Notes in Computer Science, 2239:524– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2239/22390524.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2239/22390524.
pdf.

Janssen:2001:OPP

[1882] Micha Janssen, Pascal Van Hentenryck, and Yves Deville. Optimal pruning in parametric differential equations. Lecture Notes in Computer Science, 2239:539–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2239/22390539.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2239/22390539.
pdf.

Azevedo:2001:ICP

[1883] Francisco Azevedo and Pedro Barahona. Interaction of constraint programming and local search for optimisation problems. Lecture Notes in Computer Science, 2239:554–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2239/22390554.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2239/22390554.
pdf.

Bennaceur:2001:PKA

[1884] Hachemi Bennaceur and Mohamed-Salah Affane. Partition-k-AC: An efficient filtering technique combining domain partition and arc consistency. Lecture Notes in Computer Science, 2239:560–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2239/22390560.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2239/22390560.
pdf.

Bessiere:2001:NBV

[1885] Christian Bessière, Assef Chmeiss, and Lakhdar Saïs. Neighborhood-based variable ordering heuristics for the constraint satisfaction problem. Lecture Notes in Computer Science, 2239:565– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2239/22390565.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2239/22390565.
pdf.

Cadoli:2001:EPB

[1886] Marco Cadoli. The expressive power of binary linear programming. Lecture Notes in Computer Science, 2239:570– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2239/22390570.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2239/22390570.
pdf.

Colton:2001:CGA

[1887] Simon Colton and Ian Miguel. Constraint generation via automated theory formation. Lecture Notes in Computer Science, 2239:575–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2239/22390575.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2239/22390575.
                                        pdf.

Easton:2001:TTP

[1888] Kelly Easton, George Nemhauser, and Michael Trick. The traveling tournament problem description and benchmarks. Lecture Notes in Computer Science, 2239:580–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2239/22390580.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2239/22390580.
                                        pdf.

Freuder:2001:DEI

[1889] Eugene C. Freuder, Chavalit Likitvivatanavong, and Richard J. Wallace. Deriving explanations and implications for constraint satisfaction problems. Lecture Notes in Computer Science, 2239:585–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2239/22390585.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2239/22390585.
pdf.

Freuder:2001:GTI

[1890] Eugene C. Freuder and Barry O’Sullivan.

Generating tradeoffs for interactive constraint-based configuration. Lecture Notes in Computer Science, 2239:590– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2239/22390590.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2239/22390590.
pdf.

Gasca:2001:SCB

[1891] Rafael M. Gasca, Juan A. Ortega, and Miguel Toro. Structural constraintbased modeling and reasoning with basic configuration cells. Lecture Notes in Computer Science, 2239:595– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2239/22390595.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2239/22390595.
pdf.

Granvilliers:2001:COC

[1892] Laurent Granvilliers and Eric Monfroy. Composition operators for constraint propagation: An application to choco. Lecture Notes in Computer Science, 2239:600–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2239/22390600.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2239/22390600.
                                        pdf.

Hirsch:2001:SBS

[1893] Edward A. Hirsch and Arist Kojevnikov. Solving Boolean satisfiability using local search guided by unit clause elimination. Lecture Notes in Computer Science, 2239:605–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2239/22390605.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2239/22390605.
                                        pdf.

Katsirelos:2001:GCC

[1894] George Katsirelos and Fahiem Bacchus. GAC on conjunctions of constraints. Lecture Notes in Computer Science, 2239:610–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2239/22390610.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2239/22390610.
                                        pdf.

Smith:2001:DMP

[1895] Barbara M. Smith. Dual models of permutation problems. Lecture Notes in Computer Science, 2239:615– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2239/22390615.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2239/22390615.
pdf.

Solnon:2001:BLS

[1896] Christine Solnon. Boosting local search with artificial ants. Lecture Notes in Computer Science, 2239:620– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2239/22390620.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2239/22390620.
pdf.

vanBeek:2001:FOI

[1897] Peter van Beek and Kent Wilken.

Fast optimal instruction scheduling for single-issue processors with arbitrary latencies. Lecture Notes in Computer Science, 2239:625–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2239/22390625.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2239/22390625.
pdf.

Ekelin:2001:ESH

[1898] Cecilia Ekelin and Jan Jonsson. Evaluation of search heuristics for embedded system scheduling problems. Lecture Notes in Computer Science, 2239: 640–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2239/22390640.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2239/22390640.
pdf.

Fromherz:2001:ISS

[1899] Markus P. J. Fromherz and James V. Mahoney. Interpreting sloppy stick figures with constraint-based subgraph matching. Lecture Notes in Computer Science, 2239:655–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2239/22390655.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2239/22390655.
pdf.

Verfaillie:2001:SSO

[1900] Gérard Verfaillie and Michel Lemaître. Selecting and scheduling observations for agile satellites: Some lessons from the constraint reasoning community point of view. Lecture Notes in Computer Science, 2239:670–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2239/22390670.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2239/22390670.
pdf.

Modi:2001:DDC

[1901] Pragnesh Jay Modi, Hyuckchul Jung, Milind Tambe, Wei-Min Shen, and Shriniwas Kulkarni. A dynamic distributed constraint satisfaction approach to resource allocation. Lecture Notes in Computer Science, 2239:685– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2239/22390685.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2239/22390685.
pdf.

Mattioli:2001:COF

[1902] Juliette Mattioli, Nicolas Museux, J. Jourdan, Pierre Savéant, and Simon de Givry. A constraint optimization framework for mapping a digital signal processing application onto a parallel architecture. Lecture Notes in Computer Science, 2239:701– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2239/22390701.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2239/22390701.
pdf.

Voudouris:2001:IST

[1903] Christos Voudouris, Raphael Dorne, David Lesaint, and Anne Liret. iOpt: a software toolkit for heuristic search methods. Lecture Notes in Computer Science, 2239:716–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2239/22390716.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2239/22390716.
pdf.

Merchez:2001:APS

[1904] Sylvain Merchez, Christophe Lecoutre, and Frederic Boussemart. AbsCon: a prototype to solve CSPs with abstraction. Lecture Notes in Computer Science, 2239:730–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2239/22390730.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2239/22390730.
                                        pdf.

Vancza:2001:CEM

[1905] József Váncza and András Márkus. A constraint engine for manufacturing process planning. Lecture Notes in Computer Science, 2239:745–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2239/22390745.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2239/22390745.
                                        pdf.

Beckwith:2001:DDI

[1906] Amy M. Beckwith and Berthe Y. Choueiry. On the dynamic detection of interchangeability in finite constraint satisfaction problems. Lecture Notes in Computer Science, 2239:760– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2239/22390760.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2239/22390760.
                                        pdf.

Drake:2001:AGI

[1907] Lyndon Drake. Automatic generation of implied clauses for SAT. Lecture Notes in Computer Science, 2239:761– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2239/22390761.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2239/22390761.
pdf.

Fioravanti:2001:VIS

[1908] Fabio Fioravanti. Verification of infinite-state systems by specialization of CLP programs. Lecture Notes in Computer Science, 2239:762–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2239/22390762.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2239/22390762.
pdf.

Gavanelli:2001:POC

[1909] Marco Gavanelli. Partially ordered constraint optimization problems. Lecture Notes in Computer Science, 2239:763– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2239/22390763.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2239/22390763.
pdf.

Gennari:2001:TCS

[1910] Rosella Gennari. Translations for comparing soft frameworks. Lecture Notes in Computer Science, 2239:764– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2239/22390764.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2239/22390764.
                                        pdf.

Harris:2001:CSC

[1911] Mitchell A. Harris. Counting satisfiable k-CNF formulas. Lecture Notes in Computer Science, 2239:765– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2239/22390765.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2239/22390765.
                                        pdf.

Hnich:2001:HLM

[1912] Brahim Hnich. High-level modelling and reformulation of constraint satisfaction problems. Lecture Notes in Computer Science, 2239:766–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2239/22390766.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2239/22390766.
                                        pdf.

Jung:2001:DCS

[1913] Hyuckchul Jung. Distributed constraint satisfaction as a computational model of negotiation via argumentation. Lecture Notes in Computer Science, 2239:767–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/

series/0558/bibs/2239/22390767.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2239/22390767.pdf.

Kilborn:2001:AAU

[1914] Erik Kilborn. Aircraft assignment using constraint programming. Lecture Notes in Computer Science, 2239:768– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2239/22390768.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2239/22390768.
pdf.

Kiziltan:2001:LHC

[1915] Zeynep Kiziltan. Labelling heuristics for CSP application domains. Lecture Notes in Computer Science, 2239:769– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2239/22390769.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2239/22390769.
pdf.

Lynce:2001:ISA

[1916] Inês Lynce and João Marques-Silva.

Improving SAT algorithms by using search pruning techniques. Lecture Notes in Computer Science, 2239:770– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2239/22390770.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2239/22390770.
pdf.

McDonald:2001:OSB

[1917] Iain McDonald. Optimum symmetry breaking in CSPs using group theory. Lecture Notes in Computer Science, 2239:771–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2239/22390771.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2239/22390771.
pdf.

Bessiere:2001:DDB

[1918] Christian Bessière, Arnold Maestre, and Pedro Meseguer. Distributed dynamic backtracking. Lecture Notes in Computer Science, 2239:772–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2239/22390772.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2239/22390772.
pdf.

Modi:2001:CPD

[1919] Pragnesh Jay Modi. Constraint programming for distributed resource allocation. Lecture Notes in Computer Science, 2239:773–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2239/22390773.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2239/22390773.
pdf.

Neagu:2001:ECS

[1920] Nicoleta Neagu. Exploiting the CSP structure by interchangeability. Lecture Notes in Computer Science, 2239: 774–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2239/22390774.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2239/22390774.
pdf.

Panati:2001:CPT

[1921] Andrea Panati. Constraint processing techniques for model-based reasoning about dynamic systems. Lecture Notes in Computer Science, 2239:776– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2239/22390776.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2239/22390776.
pdf.

Ringwelski:2001:DCS

[1922] Georg Ringwelski. Distributed constraint satisfaction with cooperating asynchronous solvers. Lecture Notes in Computer Science, 2239:777–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2239/22390777.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2239/22390777.
pdf.

Rousseau:2001:BNR

[1923] Louis-Martin Rousseau, Gilles Pesant, and Michel Gendreau. Building negative reduced cost paths using constraint programming. Lecture Notes in Computer Science, 2239:778– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2239/22390778.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2239/22390778.
pdf.

Salido:2001:INB

[1924] Miguel A. Salido and Federico Barber. An incremental and non-binary CSP solver: The hyperpolyhedron search algorithm. Lecture Notes in Computer Science, 2239:779–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2239/22390779.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2239/22390779.
pdf.

Schwarz:2001:PSG

[1925] Sibylle Schwarz. Partial stable generated models of generalized logic programs with constraints. Lecture Notes in Computer Science, 2239:781– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2239/22390781.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2239/22390781.
pdf.

Seelisch:2001:HCP

[1926] Frank Seelisch. Heterogeneous constraint problems. Lecture Notes in Computer Science, 2239:783–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2239/22390783.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2239/22390783.
pdf.

Sheridan:2001:CSE

[1927] Daniel Sheridan. Comparing SAT encodings for model checking. Lecture Notes in Computer Science, 2239:784– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2239/22390784.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2239/22390784.
pdf.

Silaghi:2001:ASN

[1928] Marius-Călin Silaghi, Stefan Sabău, Djamila Sam-Haroud, and Boi Faltings. Asynchronous search for numeric DisCSPs. Lecture Notes in Computer Science, 2239:785–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2239/22390785.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2239/22390785.
pdf.

Valencia:2001:TCC

[1929] Frank D. Valencia. Temporal concurrent constraint programming. Lecture Notes in Computer Science, 2239:786– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2239/22390786.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2239/22390786.
                                        pdf.

Anonymous:2001:AIav

[1930] Anonymous. Author index. Lecture Notes in Computer Science, 2239:787– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2239/2239auth.pdf.

Roth:2001:RSC

[1931] Volker Roth. On the robustness of some cryptographic protocols for mobile agent protection. Lecture Notes in Computer Science, 2240:1– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2240/22400001.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2240/22400001.
                                        pdf.

Tan:2001:TRM

[1932] Hock Kim Tan and Luc Moreau. Trust relationships in a mobile agent system. Lecture Notes in Computer Science, 2240:15–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2240/22400015.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2240/22400015.
pdf.

Fischmeister:2001:EST

[1933] Sebastian Fischmeister, Giovanni Vigna, and Richard A. Kemmerer. Evaluating the security of three Javabased mobile agent systems. Lecture Notes in Computer Science, 2240:31– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2240/22400031.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2240/22400031.
pdf.

Hannotin:2001:FSV

[1934] Xavier Hannotin, Paolo Maggi, and Riccardo Sisto. Formal specification and verification of mobile agent data integrity properties: a case study. Lecture Notes in Computer Science, 2240: 42–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2240/22400042.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2240/22400042.
pdf.

Carbunar:2001:LR

[1935] Bogdan Carbunar, Marco Tulio Valente, and Jan Vitek. Lime revisited. Lecture Notes in Computer Science, 2240:54–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2240/22400054.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2240/22400054.
                                        pdf.

Brandt:2001:DAM

[1936] Raimund Brandt and Helmut Reiser. Dynamic adaptation of mobile agents in heterogeneous environments. Lecture Notes in Computer Science, 2240: 70–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2240/22400070.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2240/22400070.
                                        pdf.

Gendelman:2001:FFA

[1937] Eugene Gendelman, Lubomir F. Bic, and Michael B. Dillencourt. Fast file access for fast agents. Lecture Notes in Computer Science, 2240:88– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2240/22400088.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2240/22400088.
                                        pdf.

Satoh:2001:FER

[1938] Ichiro Satoh. Flying emulator: Rapid building and testing of networked applications for mobile computers. Lecture Notes in Computer Science, 2240: 103–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2240/22400103.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2240/22400103.
pdf.

Thati:2001:CAH

[1939] Prasannaa Thati, Po-Hao Chang, and Gul Agha. Crawlets: Agents for high performance Web search engines. Lecture Notes in Computer Science, 2240: 119–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2240/22400119.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2240/22400119.
pdf.

Feng:2001:EMB

[1940] Xinyu Feng, Jiannong Cao, Jian Lü, and Henry Chan. An efficient mailboxbased algorithm for message delivery in mobile agent systems. Lecture Notes in Computer Science, 2240:135– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2240/22400135.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2240/22400135.
pdf.

Haller:2001:UPS

[1941] Klaus Haller and Heiko Schuldt. Using predicates for specifying targets of migration and messages in a peer-to-peer mobile agent environment. Lecture Notes in Computer Science, 2240: 152–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2240/22400152.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2240/22400152.
                                        pdf.

Roth:2001:SSG

[1942] Volker Roth and Jan Peters. A scalable and secure global tracking service for mobile agents. Lecture Notes in Computer Science, 2240:169– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2240/22400169.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2240/22400169.
                                        pdf.

Bettini:2001:TSM

[1943] Lorenzo Bettini and Rocco De Nicola. Translating strong mobility into weak mobility. Lecture Notes in Computer Science, 2240:182–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2240/22400182.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2240/22400182.
                                        pdf.

Illmann:2001:TMM

[1944] Torsten Illmann, Tilman Krueger, Frank Kargl, and Michael Weber. Transparent migration of mobile agents using the Java platform debugger architecture. Lecture Notes in Computer Science, 2240:198–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2240/22400198.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2240/22400198.
pdf.

Villazon:2001:PRR

[1945] Alex Villazón and Walter Binder.

Portable resource reification in Javabased mobile agent systems. Lecture Notes in Computer Science, 2240:213– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2240/22400213.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2240/22400213.
pdf.

Gray:2001:MAV

[1946] Robert S. Gray, David Kotz, Ronald A.

Peterson, Joyce Barton, Daria Chacón, Peter Gerken, Martin Hofmann, Jeffrey Bradshaw, Maggie Breedy, Renia Jeffers, and Niranjan Suri. Mobile-agent versus client/server performance: Scalability in an information-retrieval task. Lecture Notes in Computer Science, 2240:229–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2240/22400229.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2240/22400229.
pdf.

Dikaiakos:2001:PEM

[1947] Marios Dikaiakos, Melinos Kyriakou, and George Samaras. Performance evaluation of mobile-agent middleware: a hierarchical approach. Lecture Notes in Computer Science, 2240:244– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2240/22400244.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2240/22400244.
                                        pdf.

Xie:2001:SMT

[1948] Rong Xie, Daniela Rus, and Cliff Stein. Scheduling multi-task agents. Lecture Notes in Computer Science, 2240:260– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2240/22400260.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2240/22400260.
                                        pdf.

Anonymous:2001:AIaw

[1949] Anonymous. Author index. Lecture Notes in Computer Science, 2240:277– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2240/2240auth.pdf.

Martin:2001:GMI

[1950] Alexander Martin. General mixed integer programming: Computational issues for branch-and-cut algorithms. Lecture Notes in Computer Science, 2241:1–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/

0558/bibs/2241/22410001.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2241/22410001.
pdf.

Balas:2001:PLC

[1951] Egon Balas. Projection and lifting in combinatorial optimization. Lecture Notes in Computer Science, 2241:26– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2241/22410026.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2241/22410026.
pdf.

Pochet:2001:MPM

[1952] Yves Pochet. Mathematical programming models and formulations for deterministic production planning problems. Lecture Notes in Computer Science, 2241:57–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2241/22410057.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2241/22410057.
pdf.

Lemarechal:2001:LR

[1953] Claude Lemaréchal. Lagrangian relaxation. Lecture Notes in Computer Science, 2241:112–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2241/22410112.
htm;   http://link.springer-ny.

com/link/service/series/0558/papers/
2241/22410112.pdf.

Elf:2001:BCA

[1954] Matthias Elf, Carsten Gutwenger, Michael Jünger, and Giovanni Rinaldi. Branch-and-cut algorithms for combinatorial optimization and their implementation in ABACUS. Lecture Notes in Computer Science, 2241:157– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2241/22410157.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2241/22410157.
pdf.

Ladanyi:2001:BCP

[1955] Laszlo Ladányi, Ted K. Ralphs, and Leslie E. Trotter, Jr. Branch, cut, and price: Sequential and parallel. Lecture Notes in Computer Science, 2241:223– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2241/22410223.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2241/22410223.
pdf.

Applegate:2001:TCW

[1956] David Applegate, Robert Bixby, Vašek Chvátal, and William Cook. TSP cuts which do not conform to the template paradigm. Lecture Notes in Computer Science, 2241:261–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

bibs/2241/22410261.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2241/22410261.
pdf.

Anonymous:2001:AIax

[1957] Anonymous. Author index. Lecture Notes in Computer Science, 2241:305– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2241/2241auth.pdf.

Gannon:2001:GAD

[1958] Dennis Gannon. Grid application design using software components and Web services. Lecture Notes in Computer Science, 2242:1– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2242/22420001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2242/22420001.
pdf.

Verma:2001:DIC

[1959] Snigdha Verma, Manish Parashar, Jarek Gawor, and Gregor von Laszewski. Design and implementation of a CORBA commodity grid kit. Lecture Notes in Computer Science, 2242: 2–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2242/22420002.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2242/22420002.
pdf.

Denis:2001:THP

[1960] Alexandre Denis, Christian Pérez, and Thierry Priol. Towards high performance CORBA and MPI middlewares for grid computing. Lecture Notes in Computer Science, 2242:14– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2242/22420014.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2242/22420014.
pdf.

Furmento:2001:IGE

[1961] Nathalie Furmento, Anthony Mayer, Stephen McGough, Steven Newhouse, Tony Field, and John Darlington. An integrated grid environment for component applications. Lecture Notes in Computer Science, 2242:26–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2242/22420026.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2242/22420026.
pdf.

Cannataro:2001:KGH

[1962] Mario Cannataro, Domenico Talia, and Paolo Trunfio. KNOWLEDGE GRID: High performance knowledge discovery services on the grid. Lecture Notes in Computer Science, 2242:38– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2242/22420038.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2242/22420038.
pdf.

Iamnitchi:2001:FDR

[1963] Adriana Iamnitchi and Ian Foster.

On fully decentralized resource discovery in grid environments. Lecture Notes in Computer Science, 2242:51– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2242/22420051.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2242/22420051.
pdf.

Lee:2001:ASG

[1964] Byoung-Dai Lee and Jon B. Weissman.

An adaptive service grid architecture using dynamic replica management. Lecture Notes in Computer Science, 2242:63–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2242/22420063.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2242/22420063.
pdf.

Ranganathan:2001:IDR

[1965] Kavitha Ranganathan and Ian Foster.

Identifying dynamic replication strategies for a high-performance data grid. Lecture Notes in Computer Science, 2242:75–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2242/22420075.htm;   http:
//link.springer-ny.com/link/service/

series/0558/papers/2242/22420075.
pdf.

Weissman:2001:ESR

[1966] Jon B. Weissman and Pramod Srinivasan. Ensemble scheduling: Resource co-allocation on the computational grid. Lecture Notes in Computer Science, 2242:87–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2242/22420087.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2242/22420087.
pdf.

Katramatos:2001:JCG

[1967] Dimitrios Katramatos, Marty Humphrey, Andrew Grimshaw, and Steve Chapin. JobQueue: a computational Gridwide queuing system. Lecture Notes in Computer Science, 2242:99–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2242/22420099.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2242/22420099.
pdf.

Sarangi:2001:SMG

[1968] Anuraag Sarangi, Alok Shriram, and Avinash Shankar. A scheduling model for grid computing systems. Lecture Notes in Computer Science, 2242:111– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2242/22420111.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2242/22420111.
pdf.

Beck:2001:EVE

[1969] Micah Beck, Terry Moore, and James S. Plank. Exposed versus encapsulated approaches to grid service architecture. Lecture Notes in Computer Science, 2242:124–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2242/22420124.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2242/22420124.
pdf.

Hacker:2001:MAM

[1970] Thomas J. Hacker and Brian D.

Athey. A methodology for account management in grid computing environments. Lecture Notes in Computer Science, 2242:133–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2242/22420133.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2242/22420133.
pdf.

Sundaram:2001:PEF

[1971] Babu Sundaram and Barbara M. Chapman. Policy engine: a framework for authorization, accounting policy specification and evaluation in grids. Lecture Notes in Computer Science, 2242: 145–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2242/22420145.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2242/22420145.
                                        pdf.

Vraalsen:2001:PCP

[1972] Fredrik Vraalsen, Ruth A. Aydt, Celso L. Mendes, and Daniel A. Reed. Performance contracts: Predicting and monitoring grid application behavior. Lecture Notes in Computer Science, 2242:154–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2242/22420154.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2242/22420154.
                                        pdf.

Yarrow:2001:PLD

[1973] Maurice Yarrow, Karen M. McCann, Edward Tejni, and Adrian DeVivo. Production-level distributed parametric study capabilities for the grid. Lecture Notes in Computer Science, 2242: 166–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2242/22420166.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2242/22420166.
                                        pdf.

Lueking:2001:DED

[1974] Lee Lueking, Lauri Loebel-Carpenter, Wyatt Merritt, Carmenita Moore, Ruth Pordes, Igor Terekhov, Sinisa Veseli, Matt Vranicar, Steve White, and Vicky White. The D0 experiment data grid — SAM. Lecture Notes in Computer Science, 2242:177– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2242/22420177.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2242/22420177.
pdf.

Anonymous:2001:AIay

[1975] Anonymous. Author index. Lecture Notes in Computer Science, 2242:185– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2242/2242auth.pdf.

Dominguez:2001:AAD

[1976] Eladio Domínguez and Angel R.

Francés. An axiomatic approach to digital topology. Lecture Notes in Computer Science, 2243:3–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2243/22430003.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2243/22430003.
pdf.

Kothe:2001:GPT

[1977] Ullrich Köthe. Generic programming techniques that make planar cell complexes easy to use. Lecture Notes in Computer Science, 2243:17– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2243/22430017.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2243/22430017.
                                        pdf.

Kovalevsky:2001:ADS

[1978] Vladimir Kovalevsky. Algorithms and data structures for computer topology. Lecture Notes in Computer Science, 2243:38–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2243/22430038.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2243/22430038.
                                        pdf.

Matveev:2001:CPM

[1979] Sergei Matveev. Computer presentation of 3-manifolds. Lecture Notes in Computer Science, 2243:59–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2243/22430059.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2243/22430059.
                                        pdf.

Tsaur:2001:CMD

[1980] Rueiher Tsaur and Michael B. Smyth. “continuous” multifunctions in discrete spaces with applications to fixed point theory. Lecture Notes in Computer Science, 2243:75–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2243/22430075.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2243/22430075.
pdf.

Andres:2001:SDS

[1981] Eric Andres, Rodolphe Breton, and Pascal Lienhardt. SpaMod: Design of a spatial modeling tool. Lecture Notes in Computer Science, 2243:91– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2243/22430091.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2243/22430091.
pdf.

Brun:2001:ICP

[1982] Luc Brun and Walter Kropatsch. Introduction to combinatorial pyramids. Lecture Notes in Computer Science, 2243:108–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2243/22430108.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2243/22430108.
pdf.

Danovaro:2001:RVB

[1983] Emanuele Danovaro, Leila De Floriani, Paola Magillo, and Enrico Puppo. Representing vertex-based simplicial multi-complexes. Lecture Notes in Computer Science, 2243:129– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2243/22430129.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2243/22430129.
                                        pdf.

Kenmochi:2001:DPL

[1984] Yukiko Kenmochi and Atsushi Imiya. Discrete polyhedrization of a lattice point set. Lecture Notes in Computer Science, 2243:150–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2243/22430150.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2243/22430150.
                                        pdf.

Zunic:2001:DPE

[1985] Joviša Zunic. Digital partitions encoding. Lecture Notes in Computer Science, 2243:163–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2243/22430163.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2243/22430163.pdf.

Alpers:2001:SID

[1986] Andreas Alpers, Peter Gritzmann, and Lionel Thorens. Stability and instability in discrete tomography. Lecture Notes in Computer Science, 2243:175– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2243/22430175.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2243/22430175.
                                        pdf.

Bhattacharya:2001:PLM

[1987] Prabir Bhattacharya, Azriel Rosenfeld, and Isaac Weiss. Point-to-line mappings and Hough transforms. Lecture Notes in Computer Science, 2243:187– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2243/22430187.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2243/22430187.
pdf.

Eckhardt:2001:DLD

[1988] Ulrich Eckhardt. Digital lines and digital convexity. Lecture Notes in Computer Science, 2243:209–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2243/22430209.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2243/22430209.
pdf.

Imiya:2001:CFD

[1989] Atsushi Imiya. Curvature flow in discrete space. Lecture Notes in Computer Science, 2243:229–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2243/22430229.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2243/22430229.
pdf.

Ronse:2001:HSC

[1990] Christian Ronse and Mohamed Tajine.

Hausdorff sampling of closed sets into a boundedly compact space. Lecture Notes in Computer Science, 2243:250– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2243/22430250.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2243/22430250.
pdf.

Webster:2001:CCD

[1991] Julian Webster. Cell complexes and digital convexity. Lecture Notes in Computer Science, 2243:272–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2243/22430272.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2243/22430272.
pdf.

Bulow:2001:ASP

[1992] Thomas Bülow and Reinhard Klette. Approximation of 3D shortest polygons in simple cube curves. Lecture Notes in Computer Science, 2243:285– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2243/22430285.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2243/22430285.
pdf.

Coeurjolly:2001:SLE

[1993] David Coeurjolly, Isabelle Debled-Rennesson, and Olivier Teytaud. Segmentation and length estimation of 3D discrete curves. Lecture Notes in Computer Science, 2243:299–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2243/22430299.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2243/22430299.
pdf.

Klette:2001:MCG

[1994] Reinhard Klette. Multigrid convergence of geometric features. Lecture Notes in Computer Science, 2243:318– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2243/22430318.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2243/22430318.
pdf.

Noakes:2001:LEC

[1995] Lyle Noakes, Ryszard Kozera, and Reinhard Klette. Length estimation for curves with different samplings. Lecture Notes in Computer Science, 2243: 339–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2243/22430339.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2243/22430339.
pdf.

Noakes:2001:DLF

[1996] Lyle Noakes and Ryszard Kozera.

The 2-D leap-frog: Integrability, noise, and digitization. Lecture Notes in Computer Science, 2243:352–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2243/22430352.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2243/22430352.
                                        pdf.

Sloboda:2001:AJS

[1997] Fridrich Sloboda and Bedrich Zatko. On approximation of Jordan surfaces in 3D. Lecture Notes in Computer Science, 2243:365–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2243/22430365.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2243/22430365.
                                        pdf.

Roerdink:2001:SMC

[1998] Jos B. T. M. Roerdink and Henk Bekker. Similarity measure computation of convex polyhedra revisited. Lecture Notes in Computer Science, 2243: 389–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2243/22430389.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2243/22430389.
                                        pdf.

Svensson:2001:RSS

[1999] Stina Svensson. Reversible surface skeletons of 3D objects by iterative thinning of distance transforms. Lecture Notes in Computer Science, 2243: 400–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2243/22430400.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2243/22430400.
pdf.

Toriwaki:2001:DTS

[2000] Jun ichiro Toriwaki and Kensaku Mori.

Distance transformation and skeletonization of 3D pictures and their applications to medical images. Lecture Notes in Computer Science, 2243:412– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2243/22430412.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2243/22430412.
pdf.

VanHorebeek:2001:ALB

[2001] Johan Van Horebeek and Ernesto Tapia-Rodriguez. About the limiting behaviour of iterated robust morphological operators. Lecture Notes in Computer Science, 2243:429–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2243/22430429.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2243/22430429.
pdf.

Veelaert:2001:CWC

[2002] Peter Veelaert. Collinearity and weak collinearity in the digital plane. Lecture Notes in Computer Science, 2243: 439–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2243/22430439.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2243/22430439.
                                        pdf.

Anonymous:2001:AIaz

[2003] Anonymous. Author index. Lecture Notes in Computer Science, 2243:455– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2243/2243ind.pdf.

Pottosin:2001:PEP

[2004] Igor V. Pottosin. A. P. Ershov — A pioneer and a leader of national programming. Lecture Notes in Computer Science, 2244:1–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2244/22440001.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2244/22440001.
                                        pdf.

Podlovchenko:2001:LPE

[2005] Rimma I. Podlovchenko. A. A. Lyapunov and A. P. Ershov in the theory of program schemes and the development of its logic concepts. Lecture Notes in Computer Science, 2244: 8–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2244/22440008.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2244/22440008.
pdf.

Gurevich:2001:ASM

[2006] Yuri Gurevich. The abstract state machine paradigm: What is in and what is out. Lecture Notes in Computer Science, 2244:24–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2244/22440024.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2244/22440024.
pdf.

Lavrov:2001:AU

[2007] Svyatoslav S. Lavrov. On algorithmic unsolvability. Lecture Notes in Computer Science, 2244:25–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2244/22440025.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2244/22440025.
pdf.

Groote:2001:RBD

[2008] Jan Friso Groote and Hans Zantema. Resolution and binary decision diagrams cannot simulate each other polynomially. Lecture Notes in Computer Science, 2244:33–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2244/22440033.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2244/22440033.
pdf.

Shilov:2001:EMC

[2009] Nikolai V. Shilov and Kwang Yi. On expressive and model checking power of propositional program logics. Lecture Notes in Computer Science, 2244: 39–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2244/22440039.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2244/22440039.
pdf.

Baar:2001:EDL

[2010] Thomas Baar, Bernhard Beckert, and Peter H. Schmitt. An extension of dynamic logic for modelling OCL’s @pre operator. Lecture Notes in Computer Science, 2244:47– 54, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2244/22440047.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2244/22440047.
pdf.

Vyatkin:2001:OAE

[2011] Valeriy Vyatkin. Optimal algorithms of event-driven re-evaluation of Boolean functions. Lecture Notes in Computer Science, 2244:55–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2244/22440055.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2244/22440055.
pdf.

Ioustinova:2001:TSS

[2012] Natalia Ioustinova and Natalia Sidorova.

A transformation of SDL specifications – A step towards the verification. Lecture Notes in Computer Science, 2244: 64–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2244/22440064.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2244/22440064.
pdf.

Mukhopadhyay:2001:AWB

[2013] Supratik Mukhopadhyay and Andreas Podelski. Accurate widenings and boundedness properties of timed systems. Lecture Notes in Computer Science, 2244:79–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2244/22440079.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2244/22440079.
pdf.

Riazanov:2001:ASB

[2014] Alexandre Riazanov and Andrei Voronkov. Adaptive saturationbased reasoning. Lecture Notes in Computer Science, 2244:95–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2244/22440095.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2244/22440095.
                                        pdf.

Eschbach:2001:VAD

[2015] Robert Eschbach. A verification approach for distributed abstract state machines. Lecture Notes in Computer Science, 2244:109–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2244/22440109.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2244/22440109.
                                        pdf.

Ehm:2001:TCC

[2016] Thorsten Ehm. Transformational construction of correct pointer algorithms. Lecture Notes in Computer Science, 2244:116–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2244/22440116.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2244/22440116.
                                        pdf.

Akama:2001:TFP

[2017] Kiyoshi Akama, Hidekatsu Koike, and Hiroshi Mabuchi. A theoretical foundation of program synthesis by equivalent transformation. Lecture Notes in Computer Science, 2244:131–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2244/22440131.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2244/22440131.
pdf.

Akama:2001:ETS

[2018] Kiyoshi Akama, Hidekatsu Koike, and Hiroshi Mabuchi. Equivalent transformation by safe extension of data structures. Lecture Notes in Computer Science, 2244:140–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2244/22440140.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2244/22440140.
pdf.

Sabelfeld:2001:STF

[2019] Viktor Sabelfeld, Christian Blumenröhr, and Kai Kapp. Semantics and transformations in formal synthesis at system level. Lecture Notes in Computer Science, 2244:149– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2244/22440149.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2244/22440149.
pdf.

Harf:2001:APS

[2020] Mait Harf, Kristiina Kindel, Vahur Kotkas, Peep Küngas, and Enn Tyugu. Automated program synthesis for Java programming language. Lecture Notes in Computer Science, 2244:157–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2244/22440157.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2244/22440157.
                                        pdf.

Mosses:2001:VPL

[2021] Peter D. Mosses. The varieties of programming language semantics. Lecture Notes in Computer Science, 2244:165– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2244/22440165.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2244/22440165.
                                        pdf.

Heldal:2001:BTA

[2022] Rogardt Heldal and John Hughes. Binding-time analysis for polymorphic types. Lecture Notes in Computer Science, 2244:191–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2244/22440191.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2244/22440191.pdf.

Mogensen:2001:ICE

[2023] Torben Æ. Mogensen. An investigation of compact and efficient number representations in the pure lambda calculus. Lecture Notes in Computer Science, 2244:205–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2244/22440205.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2244/22440205.
pdf.

Virbitskaite:2001:OST

[2024] Irina B. Virbitskaite. Observational semantics for timed event structures. Lecture Notes in Computer Science, 2244:214–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2244/22440214.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2244/22440214.
pdf.

Sabelfeld:2001:ISS

[2025] Andrei Sabelfeld. The impact of synchronisation on secure information flow in concurrent programs. Lecture Notes in Computer Science, 2244:225– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2244/22440225.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2244/22440225.
pdf.

Sokolov:2001:DPT

[2026] Valery A. Sokolov and Eugeny A.

Timofeev. Dynamical priorities without time measurement and modification of the TCP. Lecture Notes in Computer Science, 2244:240–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2244/22440240.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2244/22440240.
pdf.

Astesiano:2001:AUL

[2027] Egidio Astesiano, Maura Cerioli, and Gianna Reggio. From ADT to UML-like modelling. Lecture Notes in Computer Science, 2244:246–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2244/22440246.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2244/22440246.
pdf.

Roubtsova:2001:TUS

[2028] Ella E. Roubtsova, Jan van Katwijk, Ruud C. M. de Rooij, and Hans Toetenel. Transformation of UML specification to XTG. Lecture Notes in Computer Science, 2244:247–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2244/22440247.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2244/22440247.
pdf.

Farwer:2001:SAT

[2029] Berndt Farwer and Irina Lomazova. A systematic approach towards objectbased Petri net formalisms. Lecture Notes in Computer Science, 2244:255– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2244/22440255.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2244/22440255.
pdf.

Kozura:2001:UCP

[2030] Vitaly E. Kozura. Unfoldings of coloured Petri nets. Lecture Notes in Computer Science, 2244:268–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2244/22440268.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2244/22440268.
pdf.

Wang:2001:NBM

[2031] Shengyuan Wang, Jian Yu, and Chongyi Yuan. A net-based multitier behavior inheritance modelling method. Lecture Notes in Computer Science, 2244:279–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2244/22440279.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2244/22440279.pdf.

Petrenko:2001:SBT

[2032] Alexander K. Petrenko. Specification based testing: Towards practice. Lecture Notes in Computer Science, 2244: 287–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2244/22440287.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2244/22440287.
pdf.

Bourdonov:2001:JSE

[2033] Igor B. Bourdonov, Alexey V. Demakov, Andrew A. Jarov, Alexander S. Kossatchev, Victor V. Kuliamin, Alexander K. Petrenko, and Sergey V. Zelenov. Java specification extension for automated test development. Lecture Notes in Computer Science, 2244: 301–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2244/22440301.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2244/22440301.
pdf.

Jurjens:2001:SBT

[2034] Jan Jürjens and Guido Wimmel. Specification-based testing of firewalls. Lecture Notes in Computer Science, 2244:308–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2244/22440308.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2244/22440308.
pdf.

Terekhov:2001:AVI

[2035] Andrey N. Terekhov and Len Erlikh. Academic vs. industrial software engineering: Closing the gap. Lecture Notes in Computer Science, 2244:317– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2244/22440317.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2244/22440317.
pdf.

Koznov:2001:MRM

[2036] Dmitry Koznov, Konstantin Romanovsky, and Alexei Nikitin. A method for recovery and maintenance of software architecture. Lecture Notes in Computer Science, 2244:324–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2244/22440324.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2244/22440324.
pdf.

Boulytchev:2001:ESR

[2037] Dmitry Boulytchev and Dmitry Lomov. An empirical study of retargetable compilers. Lecture Notes in Computer Science, 2244:328–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2244/22440328.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2244/22440328.
pdf.

Lellahi:2001:CDM

[2038] Kazem Lellahi. Conceptual data modeling: An algebraic viewpoint. Lecture Notes in Computer Science, 2244:336– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2244/22440336.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2244/22440336.
pdf.

Greco:2001:IMC

[2039] Sergio Greco, Luigi Pontieri, and Ester Zumpano. Integrating and managing conflicting data. Lecture Notes in Computer Science, 2244:349–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2244/22440349.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2244/22440349.
pdf.

Zarri:2001:KEA

[2040] Gian Piero Zarri. A knowledge engineering approach to deal with ‘narrative’ multimedia documents. Lecture Notes in Computer Science, 2244:363– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2244/22440363.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2244/22440363.
pdf.

Sazonov:2001:UAC

[2041] Vladimir Sazonov. Using agents for concurrent querying of Weblike databases via a hyper-set-theoretic approach. Lecture Notes in Computer Science, 2244:378–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2244/22440378.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2244/22440378.
pdf.

Cortesi:2001:RBA

[2042] Agostino Cortesi, Baudouin Le Charlier, and Sabina Rossi. Reexecutionbased analysis of logic programs with delay declarations. Lecture Notes in Computer Science, 2244:395–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2244/22440395.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2244/22440395.
pdf.

Bruynooghe:2001:EMA

[2043] Maurice Bruynooghe, Wim Vanhoof, and Michael Codish. Pos(T ): Analyzing dependencies in typed logic programs. Lecture Notes in Computer Science, 2244:406–420, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2244/22440406.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2244/22440406.
pdf.

Jhi:2001:PTT

[2044] Yoon-Chan Jhi, Ki-Chang Kim, Kemal Ebcioglu, and Yong Surk Lee. A prolog tailoring technique on an epilog tailored procedure. Lecture Notes in Computer Science, 2244:421– 433, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2244/22440421.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2244/22440421.
                                        pdf.

Ushakov:2001:HCS

[2045] Dmitry Ushakov. Hierarchical constraint satisfaction based on subdefinite models. Lecture Notes in Computer Science, 2244:434–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2244/22440434.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2244/22440434.
                                        pdf.

Telerman:2001:UCS

[2046] Vitaly Telerman. Using constraint solvers in CAD/CAM systems. Lecture Notes in Computer Science, 2244: 446–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2244/22440446.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2244/22440446.
                                        pdf.

Granvilliers:2001:GIS

[2047] Laurent Granvilliers and Eric Monfroy. A graphical interface for solver cooperations. Lecture Notes in Computer Science, 2244:457–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2244/22440457.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2244/22440457.
                                        pdf.

Nikitchenko:2001:ACN

[2048] Nikolaj S. Nikitchenko. Abstract computability of non-deterministic programs over various data structures. Lecture Notes in Computer Science, 2244:468–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2244/22440468.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2244/22440468.
pdf.

Bonfante:2001:LTO

[2049] Guillaume Bonfante, Jean-Yves Marion, and Jean-Yves Moyen. On lexicographic termination ordering with space bound certifications. Lecture Notes in Computer Science, 2244:482– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2244/22440482.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2244/22440482.
pdf.

Korovina:2001:GCA

[2050] Margarita V. Korovina and Oleg V.

Kudinov. Generalised computability and applications to hybrid systems. Lecture Notes in Computer Science, 2244:494–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2244/22440494.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2244/22440494.
pdf.

Weiss:2001:ETT

[2051] Roland Weiss and Volker Simonis. Exploring template template parameters. Lecture Notes in Computer Science, 2244:500–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2244/22440500.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2244/22440500.
                                        pdf.

Mikheev:2001:CCM

[2052] Vitaly V. Mikheev and Stanislav A. Fedoseev. Compiler-cooperative memory management in Java. Lecture Notes in Computer Science, 2244:511– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2244/22440511.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2244/22440511.
                                        pdf.

Birngruber:2001:SCL

[2053] Dietrich Birngruber. A software composition language and its implementation. Lecture Notes in Computer Science, 2244:519–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2244/22440519.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2244/22440519.pdf.

Kalnins:2001:EDL

[2054] Audris Kalnins, Karlis Podnieks, Andris Zarins, Edgars Celms, and Janis Barzdins. Editor definition language and its implementation. Lecture Notes in Computer Science, 2244:530– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2244/22440530.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2244/22440530.
pdf.

Rodionov:2001:OSM

[2055] Alexey S. Rodionov and Dmitry V.

Leskov. Oberon-2 as successor of Modula-2 in simulation. Lecture Notes in Computer Science, 2244:538–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2244/22440538.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2244/22440538.
pdf.

Anonymous:2001:AIba

[2056] Anonymous. Author index. Lecture Notes in Computer Science, 2244:547– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2244/2244auth.pdf.

Allender:2001:WWC

[2057] Eric Allender. When worlds collide: Derandomization, lower bounds, and Kolmogorov complexity. Lecture Notes in Computer Science, 2245:1– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2245/22450001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2245/22450001.
pdf.

Arora:2001:ASG

[2058] Sanjeev Arora. Approximation schemes for geometric NP-hard problems: a survey. Lecture Notes in Computer Science, 2245:16–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2245/22450016.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2245/22450016.
pdf.

Harel:2001:CUR

[2059] David Harel and Yehuda Koren. On clustering using random walks. Lecture Notes in Computer Science, 2245: 18–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2245/22450018.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2245/22450018.
pdf.

Stirling:2001:IDD

[2060] Colin Stirling. An introduction to decidability of DPDA equivalence. Lecture Notes in Computer Science, 2245: 42–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2245/22450042.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2245/22450042.
pdf.

Zwick:2001:SPB

[2061] Uri Zwick. Semidefinite programming based approximation algorithms. Lecture Notes in Computer Science, 2245: 57–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2245/22450057.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2245/22450057.
pdf.

Agrawal:2001:HSP

[2062] Manindra Agrawal. Hard sets and pseudo-random generators for constant depth circuits. Lecture Notes in Computer Science, 2245:58–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2245/22450058.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2245/22450058.
pdf.

Agrawal:2001:FOI

[2063] Manindra Agrawal. The first-order isomorphism theorem. Lecture Notes in Computer Science, 2245:70–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2245/22450070.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2245/22450070.
                                        pdf.

Anderson:2001:TOB

[2064] Richard Anderson, Sampath Kannan, Howard Karloff, and Richard E. Ladner. Thresholds and optimal binary comparison search trees. Lecture Notes in Computer Science, 2245:83– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2245/22450083.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2245/22450083.
                                        pdf.

Brim:2001:DLM

[2065] Luboš Brim, Ivana Černá, Pavel Krčál, and Radek Pelánek. Distributed LTL model checking based on negative cycle detection. Lecture Notes in Computer Science, 2245:96–107, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2245/22450096.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2245/22450096.
                                        pdf.

Calcagno:2001:CCR

[2066] Cristiano Calcagno, Hongseok Yang, and Peter W. O’Hearn. Computability and complexity results for a spatial assertion language for data structures. Lecture Notes in Computer Science, 2245:108–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2245/22450108.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2245/22450108.
pdf.

Chen:2001:UND

[2067] Jianer Chen, Donald K. Friesen, Weijia Jia, and Iyad A. Kanj. Using nondeterminism to design deterministic algorithms. Lecture Notes in Computer Science, 2245:120–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2245/22450120.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2245/22450120.
pdf.

Dang:2001:LVR

[2068] Zhe Dang, Oscar H. Ibarra, and Pierluigi San Pietro. Liveness verification of reversal-bounded multicounter machines with a free counter. Lecture Notes in Computer Science, 2245:132– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2245/22450132.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2245/22450132.
pdf.

Dold:2001:MVC

[2069] Axel Dold and Vincent Vialard. A mechanically verified compiling specification for a Lisp compiler. Lecture Notes in Computer Science, 2245:144– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2245/22450144.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2245/22450144.
                                        pdf.

Fisman:2001:BRM

[2070] Dana Fisman and Amir Pnueli. Beyond regular model checking. Lecture Notes in Computer Science, 2245:156– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2245/22450156.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2245/22450156.
                                        pdf.

Forster:2001:RBC

[2071] Jürgen Forster, Matthias Krause, Satyanarayana V. Lokam, Rustam Mubarakzjanov, Niels Schmitt, and Hans Ulrich Simon. Relations between communication complexity, linear arrangements, and computational complexity. Lecture Notes in Computer Science, 2245:171–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2245/22450171.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2245/22450171.
                                        pdf.

Gupta:2001:OOS

[2072] Neelima Gupta, Sumit Chopra, and Sandeep Sen. Optimal, outputsensitive algorithms for constructing upper envelope of line segments in parallel. Lecture Notes in Computer Science, 2245:183–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2245/22450183.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2245/22450183.
pdf.

Guruswami:2001:LDE

[2073] Venkatesan Guruswami. List decoding from erasures: Bounds and code constructions. Lecture Notes in Computer Science, 2245:195–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2245/22450195.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2245/22450195.
pdf.

Jaggi:2001:VLE

[2074] Neeraj Jaggi and K. Gopinath. Verification of a leader election algorithm in timed asynchronous systems. Lecture Notes in Computer Science, 2245: 207–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2245/22450207.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2245/22450207.
pdf.

Jakoby:2001:EAF

[2075] Andreas Jakoby and Christian Schindelhauer. Efficient addition on field programmable gate arrays. Lecture Notes in Computer Science, 2245:219– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2245/22450219.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2245/22450219.
                                        pdf.

Krishnan:2001:DMD

[2076] Radha Krishnan and Balaji Raghavachari. The directed minimum-degree spanning tree problem. Lecture Notes in Computer Science, 2245:232–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2245/22450232.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2245/22450232.
                                        pdf.

Lukovszki:2001:EBR

[2077] Tamás Lukovszki, Anil Maheshwari, and Norbert Zeh. I/O-efficient batched range counting and its applications to proximity problems. Lecture Notes in Computer Science, 2245:244–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2245/22450244.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2245/22450244.
                                        pdf.

Madhusudan:2001:BMS

[2078] P. Madhusudan and B. Meenakshi. Beyond message sequence graphs. Lecture Notes in Computer Science, 2245: 256–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2245/22450256.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2245/22450256.
pdf.

Mastrolilli:2001:GTO

[2079] Monaldo Mastrolilli. Grouping techniques for one machine scheduling subject to precedence constraints. Lecture Notes in Computer Science, 2245:268– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2245/22450268.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2245/22450268.
pdf.

Nielsen:2001:PDT

[2080] Mogens Nielsen, Vladimiro Sassone, and Jirí Srba. Properties of distributed timed-arc Petri nets. Lecture Notes in Computer Science, 2245:280– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2245/22450280.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2245/22450280.
pdf.

Peled:2001:FV

[2081] Doron Peled, Amir Pnueli, and Lenore Zuck. From falsification to verification. Lecture Notes in Computer Science, 2245:292–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2245/22450292.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2245/22450292.
                                        pdf.

Plaku:2001:PRB

[2082] Erion Plaku and Igor E. Shparlinski. On polynomial representations of Boolean functions related to some number theoretic problems. Lecture Notes in Computer Science, 2245:305– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2245/22450305.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2245/22450305.
                                        pdf.

Pnueli:2001:RAE

[2083] Amir Pnueli, Yoav Rodeh, and Ofer Shtrichman. Range allocation for equivalence logic. Lecture Notes in Computer Science, 2245:317–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2245/22450317.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2245/22450317.
                                        pdf.

Tiwari:2001:RCG

[2084] Ashish Tiwari. Rewrite closure for ground and cancellative AC theories. Lecture Notes in Computer Science, 2245:334–??, 2001. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2245/22450334.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2245/22450334.
pdf.

Anonymous:2001:AIbb

[2085] Anonymous. Author index. Lecture Notes in Computer Science, 2245:347– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2245/2245auth.pdf.

Falcone:2001:IBT

[2086] Rino Falcone, Munindar Singh, and Yao-Hua Tan. Introduction: Bringing together humans and artificial agents in cyber-societies: a new field of trust research. Lecture Notes in Computer Science, 2246:1– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2246/22460001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2246/22460001.
pdf.

Kramer:2001:TRT

[2087] Roderick M. Kramer. Trust rules for trust dilemmas: How decision makers think and act in the shadow of doubt. Lecture Notes in Computer Science, 2246:9–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2246/22460009.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2246/22460009.
                                        pdf.

McKnight:2001:TDD

[2088] D. Harrison McKnight and Norman L. Chervany. Trust and distrust definitions: One bite at a time. Lecture Notes in Computer Science, 2246:27– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2246/22460027.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2246/22460027.
                                        pdf.

Falcone:2001:SCD

[2089] Rino Falcone and Cristiano Castelfranchi. The socio-cognitive dynamics of trust: Does trust create trust? Lecture Notes in Computer Science, 2246: 55–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2246/22460055.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2246/22460055.
                                        pdf.

Barber:2001:BRP

[2090] K. Suzanne Barber and Joonoo Kim. Belief revision process based on trust: Agents evaluating reputation of information sources. Lecture Notes in Computer Science, 2246:73–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2246/22460073.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2246/22460073.
pdf.

Nooteboom:2001:ATC

[2091] Bart Nooteboom, Tomas Klos, and Renè Jorna. Adaptive trust and co-operation: An agent-based simulation approach. Lecture Notes in Computer Science, 2246:83–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2246/22460083.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2246/22460083.
pdf.

Witkowski:2001:EBE

[2092] Mark Witkowski, Alexander Artikis, and Jeremy Pitt. Experiments in building experiential trust in a society of objective-trust based agents. Lecture Notes in Computer Science, 2246:111– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2246/22460111.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2246/22460111.
pdf.

Birk:2001:LT

[2093] Andreas Birk. Learning to trust. Lecture Notes in Computer Science, 2246: 133–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2246/22460133.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2246/22460133.
                                        pdf.

Mukherjee:2001:LMT

[2094] Rajatish Mukherjee, Bikramjit Banerjee, and Sandip Sen. Learning mutual trust. Lecture Notes in Computer Science, 2246:145–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2246/22460145.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2246/22460145.
                                        pdf.

Mass:2001:DTO

[2095] Yosi Mass and Onn Shehory. Distributed trust in open multi-agent systems. Lecture Notes in Computer Science, 2246:159–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2246/22460159.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2246/22460159.pdf.

Yu:2001:MTS

[2096] Eric Yu and Lin Liu. Modelling trust for system design using the i * strategic actors framework. Lecture Notes in Computer Science, 2246:175– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2246/22460175.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2246/22460175.
pdf.

Anonymous:2001:AIbc

[2097] Anonymous. Author index. Lecture Notes in Computer Science, 2246:195– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2246/2246auth.pdf.

Canteaut:2001:CFD

[2098] A. Canteaut. Cryptographic functions and design criteria for block ciphers. Lecture Notes in Computer Science, 2247:1–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2247/22470001.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2247/22470001.
pdf.

Domingo-Ferrer:2001:MAR

[2099] J. Domingo-Ferrer. Mobile agent route protection through hash-based mechanisms. Lecture Notes in Computer Science, 2247:17–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2247/22470017.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2247/22470017.pdf.

Kuribayashi:2001:NAF

[2100] M. Kuribayashi and H. Tanaka. A new anonymous fingerprinting scheme with high enciphering rate. Lecture Notes in Computer Science, 2247:30– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2247/22470030.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2247/22470030.
pdf.

Sarkar:2001:PAE

[2101] P. Sarkar and P. J. Schellenberg. A parallel algorithm for extending cryptographic hash functions. Lecture Notes in Computer Science, 2247:40– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2247/22470040.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2247/22470040.
pdf.

Goi:2001:IHF

[2102] B.-M. Goi, M. U. Siddiqi, and H.- T. Chuah. Incremental hash function based on pair chaining & modular arithmetic combining. Lecture Notes in Computer Science, 2247:50– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2247/22470050.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2247/22470050.
pdf.

Gupta:2001:MPP

[2103] K. C. Gupta and S. Maitra. Multiples of primitive polynomials over GF(2). Lecture Notes in Computer Science, 2247:62–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2247/22470062.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2247/22470062.
pdf.

Kim:2001:FGC

[2104] J. M. Kim, I. Yie, S. I. Oh, H. D. Kim, and J. Ryu. Fast generation of cubic irreducible polynomials for XTR. Lecture Notes in Computer Science, 2247: 73–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2247/22470073.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2247/22470073.
pdf.

Pieprzyk:2001:CPS

[2105] J. Pieprzyk and X.-M. Zhang. Cheating prevention in secret sharing over GF ( p t ). Lecture Notes in Computer Science, 2247:79–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2247/22470079.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2247/22470079.
pdf.

Miri:2001:ASM

[2106] S. A. Miri and V. K. Murty.

An application of sieve methods to elliptic curves. Lecture Notes in Computer Science, 2247:91–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2247/22470091.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2247/22470091.
                                        pdf.

Baier:2001:ECP

[2107] H. Baier. Elliptic curves of prime order over optimal extension fields for use in cryptography. Lecture Notes in Computer Science, 2247:99– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2247/22470099.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2247/22470099.
                                        pdf.

Ciet:2001:SFC

[2108] M. Ciet, J.-J. Quisquater, and F. Sica. A secure family of composite finite fields suitable for fast implementation of elliptic curve cryptography. Lecture Notes in Computer Science, 2247:108– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2247/22470108.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2247/22470108.
                                        pdf.

Sarkar:2001:FIC

[2109] P. Sarkar and D. R. Stinson. Frameproof and IPP codes. Lecture Notes in Computer Science, 2247:117–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2247/22470117.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2247/22470117.
pdf.

Safavi-Naini:2001:LAC

[2110] R. Safavi-Naini, H. Wang, and C. Xing.

Linear authentication codes: Bounds and constructions. Lecture Notes in Computer Science, 2247:127–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2247/22470127.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2247/22470127.
pdf.

Kremer:2001:SRC

[2111] S. Kremer and O. Markowitch. Selective receipt in certified E-mail. Lecture Notes in Computer Science, 2247: 136–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2247/22470136.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2247/22470136.
pdf.

Maitra:2001:SDD

[2112] S. Maitra and D. P. Mukherjee. Spatial domain digital watermarking with buyer authentication. Lecture Notes in Computer Science, 2247:149–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2247/22470149.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2247/22470149.
                                        pdf.

Lee:2001:EPA

[2113] B. Lee, K. Kim, and J. Ma. Efficient public auction with one-time registration and public verifiability. Lecture Notes in Computer Science, 2247:162– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2247/22470162.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2247/22470162.
                                        pdf.

Viswanathan:2001:AIS

[2114] K. Viswanathan, C. Boyd, and E. Dawson. An analysis of integrity services in protocols. Lecture Notes in Computer Science, 2247:175–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2247/22470175.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2247/22470175.
                                        pdf.

Bedi:2001:CNF

[2115] S. S. Bedi and N. R. Pillai. Cryptanalysis of the nonlinear FeedForward generator. Lecture Notes in Computer Science, 2247:188–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2247/22470188.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2247/22470188.
pdf.

Maurer:2001:AGW

[2116] M. Maurer, A. Menezes, and E. Teske.

Analysis of the GHS Weil descent attack on the ECDLP over characteristic two finite fields of composite degree. Lecture Notes in Computer Science, 2247:195–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2247/22470195.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2247/22470195.
pdf.

Youssef:2001:CIM

[2117] A. Youssef and G. Gong. Cryptanalysis of Imai and Matsumoto scheme B asymmetric cryptosystem. Lecture Notes in Computer Science, 2247:214– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2247/22470214.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2247/22470214.
pdf.

Mu:2001:RSB

[2118] Y. Mu and V. Varadharajan. Robust and secure broadcasting. Lecture Notes in Computer Science, 2247:223– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2247/22470223.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2247/22470223.
                                        pdf.

Srinathan:2001:TOP

[2119] K. Srinathan, C. P. Rangan, and V. Kamakoti. Toward optimal player weights in secure distributed protocols. Lecture Notes in Computer Science, 2247:232–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2247/22470232.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2247/22470232.
                                        pdf.

Maitra:2001:APC

[2120] S. Maitra. Autocorrelation properties of correlation immune Boolean functions. Lecture Notes in Computer Science, 2247:242–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2247/22470242.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2247/22470242.pdf.

Fedorova:2001:CHN

[2121] M. Fedorova and Y. Tarannikov. On the constructing of highly nonlinear resilient Boolean functions by means of special matrices. Lecture Notes in Computer Science, 2247:254–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2247/22470254.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2247/22470254.
pdf.

Ramesh:2001:TAE

[2122] D. Ramesh. A twin algorithm for efficient generation of digital signatures. Lecture Notes in Computer Science, 2247:267–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2247/22470267.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2247/22470267.
pdf.

Okamoto:2001:ESF

[2123] T. Okamoto, M. Tada, and A. Miyaji.

Efficient “on the fly” signature schemes based on integer factoring. Lecture Notes in Computer Science, 2247:275– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2247/22470275.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2247/22470275.
pdf.

Kholosha:2001:CCS

[2124] A. Kholosha. Clock-controlled shift registers and generalized geffe keystream generator. Lecture Notes in Computer Science, 2247:287–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2247/22470287.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2247/22470287.
pdf.

Chowdhury:2001:ESI

[2125] S. Chowdhury and S. Maitra. Efficient software implementation of linear feedback shift registers. Lecture Notes in Computer Science, 2247:297– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2247/22470297.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2247/22470297.
pdf.

Lim:2001:CSS

[2126] S. Lim, S. Kim, I. Yie, and J. Kim. Comments on a signature scheme based on the third order LFSR proposed at ACISP2001. Lecture Notes in Computer Science, 2247:308–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2247/22470308.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2247/22470308.
pdf.

Shujun:2001:PRB

[2127] L. Shujun, M. Xuanqin, and C. Yuanlong. Pseudo-random bit generator based on couple chaotic systems and its applications in streamcipher cryptography. Lecture Notes in Computer Science, 2247:316–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

bibs/2247/22470316.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2247/22470316.
pdf.

Samid:2001:RDC

[2128] G. Samid. Re-dividing complexity between algorithms and keys. Lecture Notes in Computer Science, 2247:330– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2247/22470330.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2247/22470330.
pdf.

Kiltz:2001:TBC

[2129] E. Kiltz. A tool box of cryptographic functions related to the Diffie– Hellman function. Lecture Notes in Computer Science, 2247:339–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2247/22470339.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2247/22470339.
pdf.

Anonymous:2001:AIbd

[2130] Anonymous. Author index. Lecture Notes in Computer Science, 2247:351– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2247/2247auth.pdf.

Gentry:2001:CNS

[2131] Craig Gentry, Jakob Jonsson, Jacques Stern, and Michael Szydlo. Cryptanalysis of the NTRU Signature Scheme (NSS) from Eurocrypt 2001. Lecture Notes in Computer Science, 2248:1– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2248/22480001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2248/22480001.
pdf.

Nguyen:2001:ISA

[2132] Phong Q. Nguyen and Igor E. Shparlinski. On the insecurity of a server-aided RSA protocol. Lecture Notes in Computer Science, 2248:21– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2248/22480021.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2248/22480021.
pdf.

Boneh:2001:MIH

[2133] Dan Boneh, Shai Halevi, and Nick Howgrave-Graham. The modular inversion hidden number problem. Lecture Notes in Computer Science, 2248: 36–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2248/22480036.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2248/22480036.
pdf.

Hopper:2001:SHI

[2134] Nicholas J. Hopper and Manuel Blum. Secure human identification protocols. Lecture Notes in Computer Science, 2248:52–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2248/22480052.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2248/22480052.
pdf.

Lenstra:2001:USM

[2135] Arjen K. Lenstra. Unbelievable security. matching AES security using public key systems. Lecture Notes in Computer Science, 2248:67– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2248/22480067.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2248/22480067.
pdf.

Muller:2001:PPT

[2136] Siguna Müller. A probable prime test with very high confidence for n ≡ 1 mod 4. Lecture Notes in Computer Science, 2248:87–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2248/22480087.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2248/22480087.
pdf.

Thome:2001:CDL

[2137] Emmanuel Thomé. Computation of discrete logarithms in F_{2^{607}}. Lecture Notes in Computer Science, 2248:107– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2248/22480107.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2248/22480107.
                                        pdf.

Stam:2001:SX

[2138] Martijn Stam and Arjen K. Lenstra. Speeding up XTR. Lecture Notes in Computer Science, 2248:125–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2248/22480125.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2248/22480125.
                                        pdf.

Cha:2001:EIB

[2139] Jae Choon Cha, Ki Hyoung Ko, Sang Jin Lee, Jae Woo Han, and Jung Hee Cheon. An efficient implementation of braid groups. Lecture Notes in Computer Science, 2248:144– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2248/22480144.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2248/22480144.
                                        pdf.

Courtois:2001:HAM

[2140] Nicolas T. Courtois, Matthieu Finiasz, and Nicolas Sendrier. How to achieve a McEliece-based digital signature scheme. Lecture Notes in Computer Science, 2248:157–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2248/22480157.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2248/22480157.
pdf.

Silverberg:2001:ETT

[2141] Alice Silverberg, Jessica Staddon, and Judy L. Walker. Efficient traitor tracing algorithms using list decoding. Lecture Notes in Computer Science, 2248: 175–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2248/22480175.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2248/22480175.
pdf.

Sugita:2001:SRV

[2142] Makoto Sugita, Kazukuni Kobara, and Hideki Imai. Security of reduced version of the block cipher Camellia against truncated and impossible differential cryptanalysis. Lecture Notes in Computer Science, 2248:193– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2248/22480193.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2248/22480193.
                                        pdf.

Hong:2001:KIA

[2143] Deukjo Hong, Jaechul Sung, Seokhie Hong, Wonil Lee, Sangjin Lee, Jongin Lim, and Okyeon Yi. Known-IV attacks on triple modes of operation of block ciphers. Lecture Notes in Computer Science, 2248:208– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2248/22480208.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2248/22480208.
                                        pdf.

Patarin:2001:GAF

[2144] Jacques Patarin. Generic attacks on Feistel schemes. Lecture Notes in Computer Science, 2248:222–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2248/22480222.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2248/22480222.
                                        pdf.

Satoh:2001:CRH

[2145] Akashi Satoh, Sumio Morioka, Kohji Takano, and Seiji Munetoh. A compact Rijndael hardware architecture with S-box optimization. Lecture Notes in Computer Science, 2248:239– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2248/22480239.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2248/22480239.
pdf.

Kang:2001:PSK

[2146] Ju-Sung Kang, Sang-Uk Shin, Dowon Hong, and Okyeon Yi. Provable security of KASUMI and 3GPP encryption mode f8. Lecture Notes in Computer Science, 2248:255–271, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2248/22480255.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2248/22480255.
pdf.

Wong:2001:EMA

[2147] Duncan S. Wong and Agnes H.

Chan. Efficient and mutually authenticated key exchange for low power computing devices. Lecture Notes in Computer Science, 2248:272–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2248/22480272.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2248/22480272.
pdf.

Bresson:2001:PAG

[2148] Emmanuel Bresson, Olivier Chevassut, and David Pointcheval. Provably authenticated group Diffie–Hellman key exchange — the dynamic case. Lecture Notes in Computer Science, 2248: 290–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2248/22480290.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2248/22480290.
                                        pdf.

Fouque:2001:FDT

[2149] Pierre-Alain Fouque and Jacques Stern. Fully distributed threshold RSA under standard assumptions. Lecture Notes in Computer Science, 2248:310– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2248/22480310.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2248/22480310.
                                        pdf.

Lysyanskaya:2001:AST

[2150] Anna Lysyanskaya and Chris Peikert. Adaptive security in the threshold setting: From cryptosystems to signature schemes. Lecture Notes in Computer Science, 2248:331–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2248/22480331.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2248/22480331.
                                        pdf.

Fouque:2001:TCS

[2151] Pierre-Alain Fouque and David Pointcheval. Threshold cryptosystems secure against chosen-ciphertext attacks. Lecture Notes in Computer Science, 2248:351– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2248/22480351.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2248/22480351.
pdf.

Chang:2001:OPE

[2152] Yan-Cheng Chang and Chi-Jen Lu.

Oblivious polynomial evaluation and oblivious neural learning. Lecture Notes in Computer Science, 2248:369– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2248/22480369.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2248/22480369.
pdf.

Liskov:2001:MIC

[2153] Moses Liskov, Anna Lysyanskaya, Silvio Micali, Leonid Reyzin, and Adam Smith. Mutually independent commitments. Lecture Notes in Computer Science, 2248:385–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2248/22480385.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2248/22480385.
pdf.

Courtois:2001:EZK

[2154] Nicolas T. Courtois. Efficient zeroknowledge authentication based on a linear algebra problem MinRank. Lecture Notes in Computer Science, 2248: 402–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2248/22480402.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2248/22480402.
                                        pdf.

Cohen:2001:RRC

[2155] Tzafrir Cohen, Joe Kilian, and Erez Petrank. Responsive round complexity and concurrent zero-knowledge. Lecture Notes in Computer Science, 2248: 422–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2248/22480422.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2248/22480422.
                                        pdf.

Haastad:2001:PCA

[2156] Johan Håstad and Mats Näslund. Practical construction and analysis of pseudo-randomness primitives. Lecture Notes in Computer Science, 2248:442– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2248/22480442.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2248/22480442.
                                        pdf.

Tarannikov:2001:ACC

[2157] Yuriy Tarannikov, Peter Korolev, and Anton Botev. Autocorrelation coefficients and correlation immunity of Boolean functions. Lecture Notes in Computer Science, 2248:460–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2248/22480460.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2248/22480460.
pdf.

Gaudry:2001:EKP

[2158] Pierrick Gaudry and Nicolas Gürel. An extension of Kedlaya’s point-counting algorithm to superelliptic curves. Lecture Notes in Computer Science, 2248: 480–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2248/22480480.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2248/22480480.
pdf.

Galbraith:2001:SCC

[2159] Steven D. Galbraith. Supersingular curves in cryptography. Lecture Notes in Computer Science, 2248:495– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2248/22480495.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2248/22480495.
pdf.

Boneh:2001:SSW

[2160] Dan Boneh, Ben Lynn, and Hovav Shacham. Short signatures from the Weil pairing. Lecture Notes in Computer Science, 2248:514–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2248/22480514.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2248/22480514.
                                        pdf.

Verheul:2001:SBC

[2161] Eric R. Verheul. Self-blindable credential certificates from the Weil pairing. Lecture Notes in Computer Science, 2248:533–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2248/22480533.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2248/22480533.
                                        pdf.

Rivest:2001:HLS

[2162] Ronald L. Rivest, Adi Shamir, and Yael Tauman. How to leak a secret. Lecture Notes in Computer Science, 2248: 552–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2248/22480552.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2248/22480552.
                                        pdf.

Bellare:2001:KPP

[2163] Mihir Bellare, Alexandra Boldyreva, Anand Desai, and David Pointcheval. Key-privacy in public-key encryption. Lecture Notes in Computer Science, 2248:566–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2248/22480566.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2248/22480566.
                                        pdf.

Abe:2001:PSF

[2164] Masayuki Abe and Miyako Ohkubo.

Provably secure fair blind signatures with tight revocation. Lecture Notes in Computer Science, 2248:583–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2248/22480583.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2248/22480583.
pdf.

Anonymous:2001:AIbe

[2165] Anonymous. Author index. Lecture Notes in Computer Science, 2248:603– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2248/2248auth.pdf.

Nagi:2001:I

[2166] Khaled Nagi. 1 introduction. Lecture Notes in Computer Science, 2249:1– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2249/22490001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2249/22490001.
pdf.

Nagi:2001:AS

[2167] Khaled Nagi. 2 application scenario.

Lecture Notes in Computer Science, 2249:5–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/

0558/bibs/2249/22490005.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2249/22490005.
pdf.

Nagi:2001:OAT

[2168] Khaled Nagi. 3 overview of agent technology. Lecture Notes in Computer Science, 2249:21–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2249/22490021.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2249/22490021.
pdf.

Nagi:2001:OTP

[2169] Khaled Nagi. 4 overview of transaction processing. Lecture Notes in Computer Science, 2249:41–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2249/22490041.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2249/22490041.
pdf.

Nagi:2001:PAT

[2170] Khaled Nagi. 5 proposed approach: Transactional agents. Lecture Notes in Computer Science, 2249:63–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2249/22490063.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2249/22490063.
pdf.

Nagi:2001:ATM

[2171] Khaled Nagi. 6 the agent transaction model. Lecture Notes in Computer Science, 2249:79–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2249/22490079.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2249/22490079.pdf.

Nagi:2001:RGM

[2172] Khaled Nagi. 7 robustness guaranteeing mechanisms. Lecture Notes in Computer Science, 2249:117–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2249/22490117.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2249/22490117.
pdf.

Nagi:2001:IEA

[2173] Khaled Nagi. 8 interacting with the execution agent. Lecture Notes in Computer Science, 2249:135–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2249/22490135.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2249/22490135.
pdf.

Nagi:2001:SS

[2174] Khaled Nagi. 9 simulation study. Lecture Notes in Computer Science, 2249: 151–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2249/22490151.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2249/22490151.
pdf.

Nagi:2001:SR

[2175] Khaled Nagi. 10 simulation results. Lecture Notes in Computer Science, 2249:161–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2249/22490161.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2249/22490161.
pdf.

Nagi:2001:SFW

[2176] Khaled Nagi. 11 summary and future work. Lecture Notes in Computer Science, 2249:183–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2249/22490183.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2249/22490183.pdf.

Nagi:2001:AAI

[2177] Khaled Nagi. Appendix I: Agent implementation prototype. Lecture Notes in Computer Science, 2249:187– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2249/22490187.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2249/22490187.
pdf.

Anonymous:2001:SId

[2178] Anonymous. Subject index. Lecture Notes in Computer Science, 2249:193– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2249/2249indx.pdf.

Anonymous:2001:Rf

[2179] Anonymous. References. Lecture Notes in Computer Science, 2249:203– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2249/2249ref.pdf.

Hodkinson:2001:MFF

[2180] Ian Hodkinson, Frank Wolter, and Michael Zakharyaschev. Monodic fragments of first-order temporal logics: 2000–2001 A.D. Lecture Notes in Computer Science, 2250:1– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2250/22500001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2250/22500001.
pdf.

Kupferman:2001:BS

[2181] Orna Kupferman and Moshe Y. Vardi.

On bounded specifications. Lecture Notes in Computer Science, 2250:24– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2250/22500024.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2250/22500024.
pdf.

Schneider:2001:IAG

[2182] Klaus Schneider. Improving automata generation for linear temporal logic by considering the automaton hierarchy. Lecture Notes in Computer Science, 2250:39–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2250/22500039.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2250/22500039.
pdf.

Diekert:2001:LTL

[2183] Volker Diekert and Paul Gastin. Local temporal logic is expressively complete for cograph dependence alphabets. Lecture Notes in Computer Science, 2250: 55–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2250/22500055.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2250/22500055.
pdf.

Berwanger:2001:GMC

[2184] Dietmar Berwanger and Erich Grädel. Games and model checking for guarded logics. Lecture Notes in Computer Science, 2250:70–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2250/22500070.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2250/22500070.
pdf.

Georgieva:2001:CSE

[2185] Lilia Georgieva, Ullrich Hustadt, and Renate A. Schmidt. Computational space efficiency and minimal model generation for guarded formulae. Lecture Notes in Computer Science, 2250: 85–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2250/22500085.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2250/22500085.
pdf.

Alechina:2001:LOC

[2186] Natasha Alechina and Brian Logan. Logical omniscience and the cost of deliberation. Lecture Notes in Computer Science, 2250:100–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2250/22500100.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2250/22500100.
pdf.

Sardina:2001:LCH

[2187] Sebastian Sardiña. Local conditional high-level robot programs. Lecture Notes in Computer Science, 2250:110– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2250/22500110.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2250/22500110.
                                        pdf.

Engelhardt:2001:RTS

[2188] Kai Engelhardt, Ron van der Meyden, and Yoram Moses. A refinement theory that supports reasoning about knowledge and time. Lecture Notes in Computer Science, 2250:125– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2250/22500125.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2250/22500125.
                                        pdf.

Letz:2001:PMG

[2189] Reinhold Letz and Gernot Stenz. Proof and model generation with disconnection tableaux. Lecture Notes in Computer Science, 2250:142–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2250/22500142.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2250/22500142.
                                        pdf.

Horton:2001:CNE

[2190] Joseph D. Horton. Counting the number of equivalent binary resolution proofs. Lecture Notes in Computer Science, 2250:157–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2250/22500157.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2250/22500157.
pdf.

deNivelle:2001:STN

[2191] Hans de Nivelle. Splitting through new proposition symbols. Lecture Notes in Computer Science, 2250:172– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2250/22500172.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2250/22500172.
pdf.

Lynch:2001:CLS

[2192] Christopher Lynch and Barbara Morawska. Complexity of linear standard theories. Lecture Notes in Computer Science, 2250:186–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2250/22500186.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2250/22500186.
pdf.

Baaz:2001:HTP

[2193] Matthias Baaz, Agata Ciabattoni, and Christian G. Fermüller. Herbrand’s theorem for Prenex Gödel logic and its consequences for theorem proving. Lecture Notes in Computer Science, 2250: 201–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2250/22500201.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2250/22500201.
                                        pdf.

Baader:2001:UDL

[2194] Franz Baader and Ralf Küsters. Unification in a description logic with transitive closure of roles. Lecture Notes in Computer Science, 2250:217– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2250/22500217.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2250/22500217.
                                        pdf.

Perrier:2001:IMP

[2195] Guy Perrier. Intuitionistic multiplicative proof nets as models of directed acyclic graph descriptions. Lecture Notes in Computer Science, 2250:233– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2250/22500233.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2250/22500233.
                                        pdf.

Luo:2001:CTC

[2196] Yong Luo and Zhaohui Luo. Coherence and transitivity in coercive subtyping. Lecture Notes in Computer Science, 2250:249–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2250/22500249.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2250/22500249.
pdf.

Schurmann:2001:TTA

[2197] Carsten Schürmann. A type-theoretic approach to induction with higherorder encodings. Lecture Notes in Computer Science, 2250:266–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2250/22500266.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2250/22500266.
pdf.

Smaus:2001:APT

[2198] Jan-Georg Smaus. Analysis of polymorphically typed logic programs using ACI-unification. Lecture Notes in Computer Science, 2250:282–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2250/22500282.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2250/22500282.
pdf.

Koshimura:2001:MGB

[2199] Miyuki Koshimura, Hiroshi Fujita, and Ryuzo Hasegawa. Model generation with Boolean constraints. Lecture Notes in Computer Science, 2250:299– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2250/22500299.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2250/22500299.
                                        pdf.

Afshordel:2001:FOA

[2200] Bijan Afshordel, Thomas Hillenbrand, and Christoph Weidenbach. First-order atom definitions extended. Lecture Notes in Computer Science, 2250:309– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2250/22500309.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2250/22500309.
                                        pdf.

Rasmussen:2001:APS

[2201] Thomas Marthedal Rasmussen. Automated proof support for interval logics. Lecture Notes in Computer Science, 2250:320–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2250/22500320.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2250/22500320.
                                        pdf.

Leivant:2001:FPF

[2202] Daniel Leivant. The functions provable by first order abstraction. Lecture Notes in Computer Science, 2250: 330–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2250/22500330.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2250/22500330.
                                        pdf.

Brunnler:2001:LSC

[2203] Kai Brünnler and Alwen Fernanto Tiu.

A local system for classical logic. Lecture Notes in Computer Science, 2250: 347–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2250/22500347.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2250/22500347.
pdf.

Rintanen:2001:PIU

[2204] Jussi Rintanen. Partial implicit unfolding in the Davis–Putnam procedure for quantified Boolean formulae. Lecture Notes in Computer Science, 2250:362– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2250/22500362.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2250/22500362.
pdf.

Walsh:2001:PPC

[2205] Toby Walsh. Permutation problems and channelling constraints. Lecture Notes in Computer Science, 2250:377– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2250/22500377.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2250/22500377.
pdf.

delVal:2001:SBP

[2206] Alvaro del Val. Simplifying binary propositional theories into connected components twice as fast. Lecture Notes in Computer Science, 2250:392– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2250/22500392.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2250/22500392.
pdf.

Eiter:2001:RAE

[2207] Thomas Eiter, Michael Fink, Giuliana Sabbatini, and Hans Tompits. Reasoning about evolving nonmonotonic knowledge bases. Lecture Notes in Computer Science, 2250:407–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2250/22500407.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2250/22500407.
pdf.

Behrend:2001:ECW

[2208] Andreas Behrend. Efficient computation of the well-founded model using update propagation. Lecture Notes in Computer Science, 2250:422– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2250/22500422.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2250/22500422.
pdf.

Amato:2001:ICB

[2209] Gianluca Amato and James Lipton. Indexed categories and bottom-up semantics of logic programs. Lecture Notes in Computer Science, 2250:438– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2250/22500438.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2250/22500438.
pdf.

Lopez-Fraguas:2001:FLP

[2210] F. J. López-Fraguas and J. Sánchez-Hernández. Functional logic programming with failure: a set-oriented view. Lecture Notes in Computer Science, 2250:455–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2250/22500455.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2250/22500455.
pdf.

Kreutzer:2001:OSF

[2211] Stephan Kreutzer. Operational semantics for fixed-point logics on constraint databases. Lecture Notes in Computer Science, 2250:470–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2250/22500470.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2250/22500470.
pdf.

Munoz:2001:ENU

[2212] Susana Muñoz, Juan José Moreno, and Manuel Hermenegildo. Efficient negation using abstract interpretation. Lecture Notes in Computer Science, 2250: 485–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2250/22500485.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2250/22500485.
                                        pdf.

Boulme:2001:CSF

[2213] Sylvain Boulmé and Grégoire Hamon. Certifying synchrony for free. Lecture Notes in Computer Science, 2250:495– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2250/22500495.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2250/22500495.
                                        pdf.

McMath:2001:CEW

[2214] David McMath, Marianna Rozenfeld, and Richard Sommer. A computer environment for writing ordinary mathematical proofs. Lecture Notes in Computer Science, 2250:507– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2250/22500507.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2250/22500507.
                                        pdf.

Serebrenik:2001:TMP

[2215] Alexander Serebrenik and Danny De Schreye. On termination of metaprograms. Lecture Notes in Computer Science, 2250:517–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2250/22500517.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2250/22500517.pdf.

Borralleras:2001:MHO

[2216] Cristina Borralleras and Albert Rubio. A monotonic higher-order semantic path ordering. Lecture Notes in Computer Science, 2250:531–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2250/22500531.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2250/22500531.
pdf.

Baumgartner:2001:EWE

[2217] Robert Baumgartner, Sergio Flesca, and Georg Gottlob. The elog Web extraction language. Lecture Notes in Computer Science, 2250:548–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2250/22500548.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2250/22500548.
pdf.

Franconi:2001:CDR

[2218] Enrico Franconi, Antonio Laureti Palma, Nicola Leone, Simona Perri, and Francesco Scarcello. Census data repair: a challenging application of disjunctive logic programming. Lecture Notes in Computer Science, 2250:561– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2250/22500561.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2250/22500561.
                                        pdf.

Bagnara:2001:BFF

[2219] Roberto Bagnara, Enea Zaffanella, Roberta Gori, and Patricia M. Hill. Boolean functions for finitetree dependencies. Lecture Notes in Computer Science, 2250:579–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2250/22500579.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2250/22500579.
                                        pdf.

Comini:2001:HTA

[2220] Marco Comini, Roberta Gori, and Giorgio Levi. How to transform an analyzer into a verifier. Lecture Notes in Computer Science, 2250:595– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2250/22500595.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2250/22500595.
                                        pdf.

Yang:2001:AMR

[2221] Rong Yang and Steve Gregory. Andorra model revised: Introducing nested domain variables and a targeted search. Lecture Notes in Computer Science, 2250:610–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2250/22500610.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2250/22500610.
pdf.

Arieli:2001:CCD

[2222] Ofer Arieli, Bert Van Nuffelen, Marc Denecker, and Maurice Bruynooghe. Coherent composition of distributed knowledge-bases through abduction. Lecture Notes in Computer Science, 2250:624–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2250/22500624.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2250/22500624.
pdf.

Fermuller:2001:TRA

[2223] Christian G. Fermüller, Georg Moser, and Richard Zach. Tableaux for reasoning about atomic updates. Lecture Notes in Computer Science, 2250:639– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2250/22500639.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2250/22500639.
pdf.

Serebrenik:2001:ITC

[2224] Alexander Serebrenik and Danny De Schreye. Inference of termination conditions for numerical loops in prolog.

Lecture Notes in Computer Science, 2250:654–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2250/22500654.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2250/22500654.
                                        pdf.

Lucas:2001:TRS

[2225] Salvador Lucas. Termination of rewriting with strategy annotations. Lecture Notes in Computer Science, 2250:669– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2250/22500669.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2250/22500669.
                                        pdf.

Genaim:2001:ITC

[2226] Samir Genaim and Michael Codish. Inferring termination conditions for logic programs using backwards analysis. Lecture Notes in Computer Science, 2250:685–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2250/22500685.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2250/22500685.
                                        pdf.

Genet:2001:RAT

[2227] Thomas Genet and Valérie Viet Triem Tong. Reachability analysis of term rewriting systems with timbuk. Lecture Notes in Computer Science, 2250: 695–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2250/22500695.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2250/22500695.
pdf.

Vanhoof:2001:BTA

[2228] Wim Vanhoof and Maurice Bruynooghe.

Binding-time annotations without binding-time analysis. Lecture Notes in Computer Science, 2250:707–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2250/22500707.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2250/22500707.
pdf.

Monroy:2001:CFP

[2229] Raúl Monroy. Concept formation via proof planning failure. Lecture Notes in Computer Science, 2250:723– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2250/22500723.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2250/22500723.
pdf.

Anonymous:2001:AIbf

[2230] Anonymous. Author index. Lecture Notes in Computer Science, 2250:737– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2250/2250auth.pdf.

Daugman:2001:PIR

[2231] John Daugman, OBE. Personal identification in real-time by wavelet analysis of iris patterns (abstract only). Lecture Notes in Computer Science, 2251: 1, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2251/22510001.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2251/22510001.
                                        pdf.

Torresani:2001:HRA

[2232] Bruno Torresani. Hybrid representations of audiophonic signals. Lecture Notes in Computer Science, 2251:2– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2251/22510002.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2251/22510002.
                                        pdf.

Wickerhauser:2001:SDA

[2233] M. Victor Wickerhauser. Singularity detection from autocovariance via wavelet packets. Lecture Notes in Computer Science, 2251:3–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2251/22510003.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2251/22510003.
                                        pdf.

Schremmer:2001:EEB

[2234] Claudia Schremmer. Empirical evaluation of boundary policies for waveletbased image coding. Lecture Notes in Computer Science, 2251:4–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2251/22510004.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2251/22510004.
pdf.

Guoxiang:2001:IFB

[2235] Song Guoxiang and Wang Weiwei.

Image-feature based second generation watermarking in wavelet domain. Lecture Notes in Computer Science, 2251: 16–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2251/22510016.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2251/22510016.
pdf.

Kim:2001:SPM

[2236] Wonkoo Kim and Ching-Chung Li. A study on preconditioning multiwavelet systems for image compression. Lecture Notes in Computer Science, 2251: 22–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2251/22510022.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2251/22510022.
pdf.

Ling:2001:RBA

[2237] Wing kuen Ling and Peter Kwong shun Tam. Reduction of blocking artifacts in both spatial domain and transformed domain. Lecture Notes in Computer Science, 2251:37– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2251/22510037.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2251/22510037.
pdf.

Ling:2001:SFS

[2238] Wing kuen Ling and Peter Kwong shun Tam. Simple and fast subband de-blocking technique by discarding the high band signals. Lecture Notes in Computer Science, 2251:44– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2251/22510044.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2251/22510044.
pdf.

Lutai:2001:MSD

[2239] Guan Lütai and Lu Feng. A method with scattered data spline and wavelets for image compression. Lecture Notes in Computer Science, 2251:49–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2251/22510049.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2251/22510049.
pdf.

Liu:2001:WBP

[2240] Li-Chang Liu, Jong-Chih Chien, Henry Y. Chuang, and Ching-Chung Li. A wavelet-based preprocessing for moving object segmentation in video sequences. Lecture Notes in Computer Science, 2251:54–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2251/22510054.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2251/22510054.
pdf.

Jerome:2001:EZW

[2241] Mbainaibeye Jérôme and Noureddine Ellouze. Embedded zerotree wavelet coding of image sequence. Lecture Notes in Computer Science, 2251:65– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2251/22510065.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2251/22510065.
pdf.

Marpe:2001:WBV

[2242] Detlev Marpe, Thomas Wiegand, and Hans L. Cycon. Waveletbased video compression using longterm memory motion-compensated prediction and context-based adaptive arithmetic coding. Lecture Notes in Computer Science, 2251:76–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2251/22510076.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2251/22510076.
                                        pdf.

Ueno:2001:WFI

[2243] Yoshito Ueno. Wavelets and fractal image compression based on their self-similarity of the space-frequency plane of images. Lecture Notes in Computer Science, 2251:87–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2251/22510087.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2251/22510087.
                                        pdf.

Heinrich:2001:IMH

[2244] Stefan Heinrich, Fred J. Hickernell, and Rong-Xian Yue. Integration of multivariate Haar wavelet series. Lecture Notes in Computer Science, 2251:99– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2251/22510099.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2251/22510099.
                                        pdf.

Han-zhang:2001:ACW

[2245] Qu Han-zhang, Xu Chen, and Zhao Ruizhen. An application of continuous wavelet transform in differential equations. Lecture Notes in Computer Science, 2251:107–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2251/22510107.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2251/22510107.
pdf.

Curran:2001:SBW

[2246] Paul F. Curran and Gary Mc-Darby. Stability of biorthogonal wavelet bases in L_{2}(R). Lecture Notes in Computer Science, 2251:117–128, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2251/22510117.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2251/22510117.
pdf.

Yang:2001:CDE

[2247] Lihua Yang, Xinge You, Robert M.

Haralick, Ihsin T. Phillips, and Yuan Y. Tang. Characterization of Dirac edge with new wavelet transform. Lecture Notes in Computer Science, 2251:129– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2251/22510129.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2251/22510129.
pdf.

Shen:2001:WAN

[2248] Youjian Shen and Wei Lin. Wavelet algorithm for the numerical solution of plane elasticity problem. Lecture Notes in Computer Science, 2251:139– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2251/22510139.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2251/22510139.
                                        pdf.

Guoxiang:2001:TNM

[2249] Song Guoxiang and Zhao Ruizhen. Three novel models of threshold estimator for wavelet coefficients. Lecture Notes in Computer Science, 2251:145– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2251/22510145.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2251/22510145.
                                        pdf.

Li:2001:PWP

[2250] Mingqi Li, Qicong Peng, and Shouming Zhong. The PSD of the waveletpacket modulation. Lecture Notes in Computer Science, 2251:151–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2251/22510151.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2251/22510151.
                                        pdf.

Yang:2001:OMD

[2251] Shouzhi Yang, Zhengxing Cheng, and Hongyong Wang. Orthogonal multiwavelets with dilation factor a. Lecture Notes in Computer Science, 2251: 157–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2251/22510157.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2251/22510157.
pdf.

Kubo:2001:WBI

[2252] Masaaki Kubo, Zaher Aghbari, Kun Seok Oh, and Akifumi Makinouchi. A wavelet-based image indexing, clustering, and retrieval technique based on edge feature. Lecture Notes in Computer Science, 2251:164–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2251/22510164.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2251/22510164.
pdf.

Tan:2001:WAS

[2253] Chew Lim Tan, Ruini Cao, and Peiyi Shen. Wavelet applications in segmentation of handwriting in archival documents. Lecture Notes in Computer Science, 2251:176–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2251/22510176.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2251/22510176.
pdf.

Kouzani:2001:WPL

[2254] Abbas Z. Kouzani and S. H. Ong.

Wavelet packets for lighting-effects determination. Lecture Notes in Computer Science, 2251:188–??, 2001. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.com/link/

                                        service/series/0558/bibs/2251/22510188.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2251/22510188.pdf.

Ma:2001:TIF

[2255] Kun Ma and Xiaoou Tang. Translationinvariant face feature estimation using discrete wavelet transform. Lecture Notes in Computer Science, 2251: 200–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2251/22510200.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2251/22510200.
                                        pdf.

Guan:2001:TEB

[2256] Yujing Guan and Lixin Zhang. Text extraction based on nonlinear frame. Lecture Notes in Computer Science, 2251: 211–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2251/22510211.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2251/22510211.
                                        pdf.

Qiang:2001:WME

[2257] Wang Qiang, Hu Weiping, Hu Jianping, and Hu Kai. A wavelet multiresolution edge analysis method for recovery of depth from defocused images. Lecture Notes in Computer Science, 2251:217–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2251/22510217.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2251/22510217.
pdf.

Chen:2001:CFN

[2258] Hanlin Chen and Silong Peng. Construction of finite non-separable orthogonal filter banks with linear phase and its application in image segmentation. Lecture Notes in Computer Science, 2251:223–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2251/22510223.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2251/22510223.
pdf.

Tang:2001:MSD

[2259] Yuan Y. Tang, Yuhua Hou, Jinping Song, and Xiaoyi Yang. Mixture-state document segmentation using waveletdomain hidden Markov tree models. Lecture Notes in Computer Science, 2251:230–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2251/22510230.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2251/22510230.
pdf.

Law:2001:SER

[2260] Tak Ming Law. Some experiment results on feature analyses of stroke sequence free matching algorithms for online Chinese character recognition. Lecture Notes in Computer Science, 2251: 237–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2251/22510237.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2251/22510237.
                                        pdf.

Law:2001:ADA

[2261] Tak Ming Law. Automatic detection algorithm of connected segments for online Chinese character recognition. Lecture Notes in Computer Science, 2251: 242–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2251/22510242.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2251/22510242.
                                        pdf.

Weiping:2001:SSD

[2262] Hu Weiping and Robert Linggard. Speech signal deconvolution using wavelet filter banks. Lecture Notes in Computer Science, 2251:248–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2251/22510248.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2251/22510248.
                                        pdf.

Borgosz:2001:PJA

[2263] Jan Borgosz and Boguslaw Cyganek. A proposal of jitter analysis based on a wavelet transform. Lecture Notes in Computer Science, 2251:257– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2251/22510257.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2251/22510257.
pdf.

Yu:2001:SGW

[2264] Weichuan Yu, Gerald Sommer, and Kostas Daniilidis. Skewness of Gabor wavelets and source signal separation. Lecture Notes in Computer Science, 2251:269–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2251/22510269.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2251/22510269.
pdf.

MacCallum:2001:AWT

[2265] M. MacCallum and A. E. A. Almaini.

The application of the wavelet transform to polysomnographic signals. Lecture Notes in Computer Science, 2251: 284–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2251/22510284.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2251/22510284.
pdf.

Su:2001:WTM

[2266] Wei Su, Hong Ma, Yuan Yan Tang, and Michio Umeda. Wavelet transform method of waveform estimation for Hilbert transform of fractional stochastic signals with noise. Lecture Notes in Computer Science, 2251:296– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2251/22510296.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2251/22510296.
                                        pdf.

Zhao:2001:MKF

[2267] Juan Zhao, Hong Ma, Zhi sheng You, and Michio Umeda. Multiscale Kalman filtering of fractal signals using wavelet transform. Lecture Notes in Computer Science, 2251:305– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2251/22510305.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2251/22510305.
                                        pdf.

Li:2001:GAC

[2268] Jian Ping Li and Yuan Yan Tang. General analytic construction for wavelet low-passed filters. Lecture Notes in Computer Science, 2251:314–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2251/22510314.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2251/22510314.
                                        pdf.

Liu:2001:DAS

[2269] Yishu Liu, Jinyu Cen, Qian Sun, and Lihua Yang. A design of automatic speech playing system based on wavelet transform. Lecture Notes in Computer Science, 2251:321–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2251/22510321.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2251/22510321.
pdf.

Yang:2001:GDW

[2270] Lihua Yang, Qiuhui Chen, and Yuan Y. Tang. General design of wavelet high-pass filters from reconstructional symbol. Lecture Notes in Computer Science, 2251:326–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2251/22510326.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2251/22510326.
pdf.

Ling:2001:RPR

[2271] Wing kuen Ling and Peter Kwong-Shun Tam. Realization of perfect reconstruction non-uniform filter banks via a tree structure. Lecture Notes in Computer Science, 2251:331–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2251/22510331.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2251/22510331.
pdf.

Ling:2001:SDT

[2272] Wing kuen Ling and Peter Kwong-Shun Tam. Set of decimators for tree structure filter banks. Lecture Notes in Computer Science, 2251:336– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2251/22510336.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2251/22510336.
                                        pdf.

Ling:2001:SPR

[2273] Wing kuen Ling and Peter Kwong-Shun Tam. Set of perfect reconstruction non-uniform filter banks via a tree structure. Lecture Notes in Computer Science, 2251:341–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2251/22510341.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2251/22510341.
                                        pdf.

Raihan:2001:JTF

[2274] Sharif Md. Raihan, Yi Wen, and Bing Zeng. Joint time-frequency distributions for business cycle analysis. Lecture Notes in Computer Science, 2251: 347–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2251/22510347.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2251/22510347.
                                        pdf.

Razak:2001:DDW

[2275] Zaidi Razak and Mashkuri Yaacob. The design of discrete wavelet transformation chip. Lecture Notes in Computer Science, 2251:359–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2251/22510359.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2251/22510359.
pdf.

Ahmadi:2001:PIW

[2276] H. Ahmadi, R. Tafreshi, F. Sassani, and G. Dumont. On the performance of informative wavelets for classification and diagnosis of machine faults. Lecture Notes in Computer Science, 2251: 369–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2251/22510369.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2251/22510369.
pdf.

Ong:2001:WBA

[2277] S. H. Ong and A. Z. Kouzani. A wavelet-based ammunition Doppler radar system. Lecture Notes in Computer Science, 2251:382–??, 2001. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.com/link/
service/series/0558/bibs/2251/22510382.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2251/22510382.pdf.

Li:2001:AWA

[2278] Jian Ping Li, Shang An Yan, and Yuan Yan Tang. The application of wavelet analysis method to civil infrastructure health monitoring. Lecture Notes in Computer Science, 2251: 393–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2251/22510393.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2251/22510393.
                                        pdf.

Ling:2001:PPW

[2279] Wing kuen Ling and Peter Kwung-Shun Tam. Piecewise periodized wavelet transform and its realization, properties and applications. Lecture Notes in Computer Science, 2251:398– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2251/22510398.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2251/22510398.
                                        pdf.

Zunze:2001:WTA

[2280] Hou Zunze. Wavelet transform and its application to decomposition of gravity anomalies. Lecture Notes in Computer Science, 2251:404–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2251/22510404.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2251/22510404.
                                        pdf.

Boying:2001:CIP

[2281] Wu Boying, Liu Shaohui, and Deng Zhongxing. Computations of inverse problem by using wavelet in multi-layer soil. Lecture Notes in Computer Science, 2251:411–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2251/22510411.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2251/22510411.
pdf.

Lu:2001:WAC

[2282] Feng Lu, Zhaoxia Yang, and Yuesheng Li. Wavelets approach in choosing adaptive regularization parameter. Lecture Notes in Computer Science, 2251:418–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2251/22510418.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2251/22510418.
pdf.

Zhao:2001:DSC

[2283] Jing Zhao, Xiu Wen Yang, Jian Ping Li, and Yuan Yan Tang. DNA sequences classification based on wavelet packet analysis. Lecture Notes in Computer Science, 2251:424–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2251/22510424.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2251/22510424.
pdf.

Yang:2001:AWT

[2284] Xiu Wen Yang, Jing Zhao, Jian Ping Li, Jing Liu, and Shun Peng Zeng. The application of the wavelet transform to the prediction of gas zones. Lecture Notes in Computer Science, 2251:430– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2251/22510430.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2251/22510430.
                                        pdf.

Zhang:2001:PMB

[2285] Zeyin Zhang and Daren Huang. Parameterizations of M-band biorthogonal wavelets. Lecture Notes in Computer Science, 2251:435–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2251/22510435.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2251/22510435.pdf.

Anonymous:2001:AIbg

[2286] Anonymous. Author index. Lecture Notes in Computer Science, 2251:449– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2251/2251ind.pdf.

Fukuda:2001:IAI

[2287] Toshio Fukuda. Intelligent adaptive interface. Lecture Notes in Computer Science, 2252:1–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2252/22520001.

htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2252/22520001.pdf.

Ishida:2001:DCB

[2288] Toru Ishida. Digital city: Bridging technologies and humans. Lecture Notes in Computer Science, 2252:2– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2252/22520002.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2252/22520002.
pdf.

Kanade:2001:DRE

[2289] Takeo Kanade. Detection, recognition, and expression analysis of human faces. Lecture Notes in Computer Science, 2252:3–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2252/22520003.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2252/22520003.
pdf.

Khatib:2001:DIH

[2290] Oussama Khatib. Digital interactivity: Human-centered robotics, haptics, and simulation. Lecture Notes in Computer Science, 2252:4– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2252/22520004.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2252/22520004.
pdf.

Yu:2001:SME

[2291] Hong Heather Yu, Peng Yin, Samuel Cheng, Xiangyang Kong, Alexander Gelman, and Robert S. Fish. Smart media: Empower media with active data hiding. Lecture Notes in Computer Science, 2252:5–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2252/22520005.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2252/22520005.
pdf.

Takahashi:2001:CPN

[2292] Hiroto Takahashi and Yasuo Yonezawa. Congestion prospect of network traffics by fluctuation of Internet transmission time. Lecture Notes in Computer Science, 2252:17–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2252/22520017.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2252/22520017.
pdf.

Horiguchi:2001:NHR

[2293] Yukio Horiguchi and Tetsuo Sawaragi. Naturalistic human-robot collaboration mediated by shared communicational modality in teleoperation system. Lecture Notes in Computer Science, 2252:24–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2252/22520024.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2252/22520024.
pdf.

Sae-Tang:2001:WPT

[2294] Suwimol Sae-Tang and Vatcharaporn Esichaikul. Web personalization techniques for E-commerce. Lecture Notes in Computer Science, 2252:36– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2252/22520036.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2252/22520036.
pdf.

Huang:2001:SNA

[2295] Runhe Huang, Jianhua Ma, and Qun Jin. A shopping negotiation agent that adapts to user preferences. Lecture Notes in Computer Science, 2252:45– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2252/22520045.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2252/22520045.
pdf.

Goderis:2001:CML

[2296] Sofie Goderis, Gustavo Rossi, Andres Fortier, Juan Cappi, and Daniel Schwabe. Combining metalevel and logic-based constructs in Web personalization. Lecture Notes in Computer Science, 2252:57–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2252/22520057.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2252/22520057.
                                        pdf.

Murugesan:2001:WPO

[2297] San Murugesan and Annamalai Ramanathan. Web personalisation — an overview. Lecture Notes in Computer Science, 2252:65–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2252/22520065.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2252/22520065.
                                        pdf.

OdoricodeMoraesFilho:2001:BOW

[2298] Carlos Alberto Odorico de Moraes Filho and Francilene Procpio Garcia. Business-oriented Web personalization: a decision-making approach. Lecture Notes in Computer Science, 2252:77– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2252/22520077.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2252/22520077.
                                        pdf.

Li:2001:PIA

[2299] Weihua Li and Xiaodong Fu. A personalized interface agent with feedback evaluation. Lecture Notes in Computer Science, 2252:83–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2252/22520083.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2252/22520083.
pdf.

Marriott:2001:VDT

[2300] Andrew Marriott, Simon Beard, John Stallo, and Quoc Huynh. VHML — directing a talking head. Lecture Notes in Computer Science, 2252:90– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2252/22520090.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2252/22520090.
pdf.

Beard:2001:BIT

[2301] Simon Beard, Donald Reid, and Russell Shepherdson. Believable and interactive talking heads for websites: MetaFace and MPEG-4. Lecture Notes in Computer Science, 2252:101– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2252/22520101.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2252/22520101.
pdf.

Arafa:2001:BMM

[2302] Yasmine Arafa and Abe Mamdani. Building multi-modal personal sales agents as interfaces to E-commerce applications. Lecture Notes in Computer Science, 2252:113– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2252/22520113.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2252/22520113.
                                        pdf.

Chau:2001:FVM

[2303] Chun Keung Chau, Chak Shun Lai, and Bertram Emil Shi. Feature vs. model based vocal tract length normalization for a speech recognitionbased interactive toy. Lecture Notes in Computer Science, 2252:134–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2252/22520134.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2252/22520134.
                                        pdf.

Lam:2001:GAB

[2304] Stephen Wang-Cheung Lam and Kelvin Yuen-Hin Ho. A genetic algorithm based approach to the synthesis of three dimensional morphing sequences. Lecture Notes in Computer Science, 2252:144–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2252/22520144.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2252/22520144.
                                        pdf.

Sim:2001:CSA

[2305] Kwang Mong Sim and Weng Hong Sun. A comparative study of ANT-based optimization for dynamic routing. Lecture Notes in Computer Science, 2252:

153–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2252/22520153.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2252/22520153.
pdf.

Munoz-Melendez:2001:WKC

[2306] Angélica Muñoz-Meléndez and Alexis Drogoul. What kind of cooperation is required by situated agents? Lecture Notes in Computer Science, 2252:165– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2252/22520165.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2252/22520165.
pdf.

Kang:2001:FTS

[2307] YunHee Kang, HoSang Ham, and ChongSun Hwang. A fault-tolerant scheme of multi-agent system for worker agents. Lecture Notes in Computer Science, 2252:171–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2252/22520171.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2252/22520171.
pdf.

Katai:2001:DCM

[2308] Osamu Katai, Kentaro Toda, and Hiroshi Kawakami. Decentralized control of multi-agent systems based on modal logics and extended higher order Petri nets. Lecture Notes in Computer Science, 2252:182–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2252/22520182.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2252/22520182.
                                        pdf.

Erdan:2001:IFE

[2309] Gu Erdan, Xu Duanqing, and Chen Chun. An improved foreground extraction approach. Lecture Notes in Computer Science, 2252:191–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2252/22520191.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2252/22520191.
                                        pdf.

Li:2001:ODS

[2310] Xingshan Li and Chunwei Yuan. Object detection simulating visual attention. Lecture Notes in Computer Science, 2252:199–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2252/22520199.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2252/22520199.pdf.

El-Bakry:2001:FFD

[2311] Hazem M. El-Bakry. Fast face detection using neural networks and image decomposition. Lecture Notes in Computer Science, 2252:205–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2252/22520205.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2252/22520205.
pdf.

Huang:2001:MFE

[2312] Ding Huang, Wei Lin, and Hong Yan. Modeling of facial expressions using NURBS curves. Lecture Notes in Computer Science, 2252:216–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2252/22520216.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2252/22520216.
pdf.

Choi:2001:ENA

[2313] Samuel P. M. Choi, Jiming Liu, and Sheung-Ping Chan. Evolutionary negotiation in agent-mediated commerce. Lecture Notes in Computer Science, 2252:224–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2252/22520224.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2252/22520224.
pdf.

Chun:2001:OAB

[2314] Andy Hon Wai Chun and Rebecca Y. M. Wong. Optimizing agent-based negotiations with branch-and-bound. Lecture Notes in Computer Science, 2252:235–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2252/22520235.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2252/22520235.
                                        pdf.

Chau:2001:EFC

[2315] Chi Kin Chau and Kwang Mong Sim. Engineering fuzzy constraint satisfaction agents for multi-user timetable scheduling. Lecture Notes in Computer Science, 2252:244–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2252/22520244.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2252/22520244.
                                        pdf.

Mashita:2001:XBD

[2316] Naoaki Mashita, Hiroshi Shigeno, and Yutaka Matsushita. An XMLbased distance learning system capable of conveying information on “LECTURE EVENT”. Lecture Notes in Computer Science, 2252:255–267, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2252/22520255.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2252/22520255.
                                        pdf.

Macias:2001:ATB

[2317] José A. Macías and Pablo Castells. An authoring tool for building adaptive learning guidance systems on the Web. Lecture Notes in Computer Science, 2252:268–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2252/22520268.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2252/22520268.
pdf.

Notsu:2001:FCI

[2318] Akira Notsu, Osamu Katai, and Hiroshi Kawakami. A framework of caring interaction by a network model of client’s concepts based on a nursing theory and naive psychological approaches. Lecture Notes in Computer Science, 2252:279–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2252/22520279.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2252/22520279.
pdf.

Zhang:2001:PDB

[2319] Shaobai Zhang, Xiefeng Cheng, and Zhiquan Feng. Parallel design based on neural network. Lecture Notes in Computer Science, 2252:291–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2252/22520291.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2252/22520291.
pdf.

Law:2001:FPP

[2320] Tak-Ming Law. Fair play protocol. Lecture Notes in Computer Science, 2252: 297–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2252/22520297.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2252/22520297.
                                        pdf.

Mild:2001:CFM

[2321] Andreas Mild and Thomas Reutterer. Collaborative filtering methods for binary market basket data analysis. Lecture Notes in Computer Science, 2252: 302–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2252/22520302.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2252/22520302.
                                        pdf.

Kim:2001:IDC

[2322] Young-Jun Kim. An implementation and design of COMOR system for OOM reuse. Lecture Notes in Computer Science, 2252:314–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2252/22520314.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2252/22520314.
                                        pdf.

Yi:2001:AAR

[2323] Wei Yi. Automatic aircraft recognition using maximum likelihood ratio test.

Lecture Notes in Computer Science, 2252:321–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2252/22520321.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2252/22520321.
pdf.

Ye:2001:MLA

[2324] Nong Ye and Xiangyang Li. A machine learning algorithm based on supervised clustering and classification. Lecture Notes in Computer Science, 2252:327– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2252/22520327.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2252/22520327.
pdf.

Li:2001:VPG

[2325] Xiaodong Li. Visualization of a parallel genetic algorithm in real time. Lecture Notes in Computer Science, 2252: 335–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2252/22520335.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2252/22520335.
pdf.

Carlsson:2001:RFN

[2326] Bengt Carlsson and Rune Gustavsson. The rise and fall of napster — an evolutionary approach. Lecture Notes in Computer Science, 2252:347– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2252/22520347.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2252/22520347.
                                        pdf.

Simhon:2001:EHD

[2327] Saul Simhon and Gregory Dudek. On the elaboration of hand-drawn sketches. Lecture Notes in Computer Science, 2252:355–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2252/22520355.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2252/22520355.pdf.

Zhao:2001:ITM

[2328] Guoying Zhao and Xinyuan Huang. The introduction of three methods generating stereoscopic image. Lecture Notes in Computer Science, 2252:365– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2252/22520365.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2252/22520365.
                                        pdf.

Wang:2001:KOI

[2329] Shangfei Wang, Enhong Chen, Jing Hu, and Xufa Wang. Kanseioriented image retrieval. Lecture Notes in Computer Science, 2252:377– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2252/22520377.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2252/22520377.
pdf.

Tavares:2001:IIC

[2330] T. A. Tavares, A. S. Araújo, and G. L. Souza Filho. ICSPACE — an Internet Cultural Space. Lecture Notes in Computer Science, 2252:389– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2252/22520389.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2252/22520389.
pdf.

Anonymous:2001:AIbh

[2331] Anonymous. Author index. Lecture Notes in Computer Science, 2252:401– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2252/2252auth.pdf.

Nishida:2001:SID

[2332] Toyoaki Nishida. 1. social intelligence design — an overview. Lecture Notes in Computer Science, 2253:3– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2253/22530003.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2253/22530003.
pdf.

Ohguro:2001:FTS

[2333] Takeshi Ohguro, Kazuhiro Kuwabara, Tatsuo Owada, and Yoshinari Shirai. 2. FaintPop: In touch with the social relationships. Lecture Notes in Computer Science, 2253:11– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2253/22530011.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2253/22530011.
                                        pdf.

Nijholt:2001:VEV

[2334] A. Nijholt. 3. from virtual environment to virtual community. Lecture Notes in Computer Science, 2253:19– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2253/22530019.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2253/22530019.
                                        pdf.

Thomas:2001:CIT

[2335] John C. Thomas. 4. collaborative innovation tools. Lecture Notes in Computer Science, 2253:27–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2253/22530027.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2253/22530027.
                                        pdf.

Fruchter:2001:BBI

[2336] R. Fruchter. 5. bricks & bits & interaction. Lecture Notes in Computer Science, 2253:35–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2253/22530035.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2253/22530035.pdf.

Cardon:2001:DMA

[2337] Alain Cardon. 6. A distributed multi-agent system for the selfevaluation of dialogs. Lecture Notes in Computer Science, 2253:43–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2253/22530043.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2253/22530043.
pdf.

Fukuhara:2001:POC

[2338] Tomohiro Fukuhara, Toyoaki Nishida, and Shunsuke Uemura. 7. public opinion channel: a system for augmenting social intelligence of a community. Lecture Notes in Computer Science, 2253: 51–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2253/22530051.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2253/22530051.
pdf.

Nakata:2001:EPD

[2339] Keiichi Nakata. 8. enabling public discourse. Lecture Notes in Computer Science, 2253:59–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2253/22530059.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2253/22530059.pdf.

Luehrs:2001:IDD

[2340] R. Luehrs, T. Malsch, and K. Voss. 9. Internet, discourses, and democracy. Lecture Notes in Computer Science, 2253:67–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2253/22530067.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2253/22530067.
                                        pdf.

Fujihara:2001:HES

[2341] Nobuhiko Fujihara. 10. how to evaluate social intelligence design. Lecture Notes in Computer Science, 2253:75– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2253/22530075.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2253/22530075.
                                        pdf.

Namatame:2001:O

[2342] Akira Namatame. 11. overview. Lecture Notes in Computer Science, 2253: 85–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2253/22530085.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2253/22530085.
pdf.

Kurahashi:2001:ANE

[2343] Setsuya Kurahashi and Takao Terano. 12. analyzing norm emergence in communal sharing via agentbased simulation. Lecture Notes in Computer Science, 2253:88–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2253/22530088.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2253/22530088.
pdf.

Takadama:2001:TCP

[2344] Keiki Takadama and Katsunori Shimohara. 13. toward cumulative progress in agent-based simulation. Lecture Notes in Computer Science, 2253:99– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2253/22530099.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2253/22530099.
pdf.

Izumi:2001:CAC

[2345] Kiyoshi Izumi. 14. complexity of agents and complexity of markets. Lecture Notes in Computer Science, 2253:110– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2253/22530110.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2253/22530110.
                                        pdf.

Sato:2001:UMP

[2346] Hiroshi Sato, Hiroyuki Matsui, Isao Ono, Hajime Kita, Takao Terano, Hiroshi Deguchi, and Yoshinori Shiozawa.

15. U-mart project: Learning economic principles from the bottom by both human and software agents. Lecture Notes in Computer Science, 2253:121– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2253/22530121.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2253/22530121.
                                        pdf.

Fukumoto:2001:MOG

[2347] Rikiya Fukumoto and Hajime Kita. 16. A multi-objective genetic algorithm approach to construction of trading agents for artificial market study. Lecture Notes in Computer Science, 2253: 132–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2253/22530132.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2253/22530132.
                                        pdf.

Mizuta:2001:ABS

[2348] Hideyuki Mizuta and Yoshiki Yamagata. 17. agent-based simulation for economic and environmental studies. Lecture Notes in Computer Science, 2253:142–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2253/22530142.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2253/22530142.
pdf.

Aruka:2001:AGE

[2349] Yuji Aruka. 18. Avatamsaka game experiment as a nonlinear Polya urn process. Lecture Notes in Computer Science, 2253:153–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2253/22530153.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2253/22530153.pdf.

Suzuki:2001:EPA

[2350] Keji Suzuki. 19. effects of punishment into actions in social agents. Lecture Notes in Computer Science, 2253:162– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2253/22530162.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2253/22530162.
pdf.

Yamashita:2001:ANG

[2351] Tomohisa Yamashita, Hidenori Kawamura, Masahito Yamamoto, and Azuma Ohuchi. 20. analysis of norms game with mutual choice. Lecture Notes in Computer Science, 2253:174– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2253/22530174.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2253/22530174.
                                        pdf.

Cho:2001:CCE

[2352] Sung-Bae Cho. 21. cooperative coevolution of multi-agents. Lecture Notes in Computer Science, 2253:185– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2253/22530185.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2253/22530185.
                                        pdf.

Sato:2001:SIK

[2353] Kazuyo Sato and Akira Namatame. 22. social interaction as knowledge trading games. Lecture Notes in Computer Science, 2253:195–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2253/22530195.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2253/22530195.
                                        pdf.

Kurumatani:2001:WTL

[2354] Koichi Kurumatani and Azuma Ohuchi. 23. world trade league as a standard problem for multi-agent economics — concept and background. Lecture Notes in Computer Science, 2253:208–??, 2001. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2253/22530208.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2253/22530208.
pdf.

Deguchi:2001:VES

[2355] Hiroshi Deguchi, Takao Terano, Koichi Kurumatani, Taro Yuzawa, Shigeji Hashimoto, Hiroyuki Matsui, Akio Sashima, and Toshiyuki Kaneda. 24. virtual economy simulation and gaming –an agent based approach–. Lecture Notes in Computer Science, 2253: 218–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2253/22530218.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2253/22530218.
pdf.

Iba:2001:BEF

[2356] Takashi Iba, Yohei Takabe, Yoshihide Chubachi, Junichiro Tanaka, Kenichi Kamihashi, Ryunosuke Tsuya, Satomi Kitano, Masaharu Hirokane, and Yoshiaki Matsuzawa. 25. boxed economy foundation model: Model framework for agent-based economic simulations. Lecture Notes in Computer Science, 2253:227–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2253/22530227.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2253/22530227.
pdf.

Tsumoto:2001:WRS

[2357] Shusaku Tsumoto, Shoji Hirano, and Masahiro Inuiguchi. 26. workshop on rough set theory and granular computing — summary. Lecture Notes in Computer Science, 2253:239–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2253/22530239.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2253/22530239.
                                        pdf.

Pawlak:2001:BTR

[2358] Zdzisław Pawlak. 27. Bayes’ theorem revised — the rough set view. Lecture Notes in Computer Science, 2253: 240–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2253/22530240.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2253/22530240.
                                        pdf.

Skowron:2001:TIS

[2359] Andrzej Skowron. 28. toward intelligent systems: Calculi of information granules. Lecture Notes in Computer Science, 2253:251–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2253/22530251.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2253/22530251.
                                        pdf.

Pal:2001:SCP

[2360] Sankar K. Pal. 29. soft computing pattern recognition: Principles, integrations, and data mining. Lecture Notes in Computer Science, 2253:261– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2253/22530261.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2253/22530261.
pdf.

Guo:2001:IUL

[2361] P. Guo and Hideo Tanaka. 30. identifying upper and lower possibility distributions with rough set concept. Lecture Notes in Computer Science, 2253: 272–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2253/22530272.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2253/22530272.
pdf.

Polkowski:2001:FIS

[2362] Lech Polkowski. 31. on fractals in information systems: The first step. Lecture Notes in Computer Science, 2253: 278–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2253/22530278.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2253/22530278.
pdf.

Miyamoto:2001:GFM

[2363] Sadaaki Miyamoto. 32. generalizations of fuzzy multisets for including infiniteness. Lecture Notes in Computer Science, 2253:283–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2253/22530283.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2253/22530283.
pdf.

Koga:2001:FMM

[2364] Takatsugu Koga, Sadaaki Miyamoto, and Osamu Takata. 33. fuzzy c-means and mixture distribution model for clustering based on L_{1}-space. Lecture Notes in Computer Science, 2253:289– 294, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2253/22530289.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2253/22530289.
pdf.

Inuiguchi:2001:RSU

[2365] Masahiro Inuiguchi and Tetsuzo Tanino. 34. on rough sets under generalized equivalence relations. Lecture Notes in Computer Science, 2253:295– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2253/22530295.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2253/22530295.
pdf.

Sakai:2001:TPD

[2366] Hiroshi Sakai. 35. two procedures for dependencies among attributes in a table with non-deterministic information: a summary. Lecture Notes in Computer Science, 2253:301–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2253/22530301.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2253/22530301.
pdf.

Song:2001:AES

[2367] Chi-Hwa Song and Won Don Lee.

36. an application of extended simulated annealing algorithm to generate the learning data set for speech recognition system. Lecture Notes in Computer Science, 2253:306–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2253/22530306.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2253/22530306.
pdf.

Intan:2001:GRS

[2368] Rolly Intan, Masao Mukaidono, and Y. Y. Yao. 37. generalization of rough sets with α-coverings of the universe induced by conditional probability relations. Lecture Notes in Computer Science, 2253:311–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2253/22530311.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2253/22530311.
                                        pdf.

Yao:2001:MOR

[2369] Y. Y. Yao and Ying Sai. 38. on mining ordering rules. Lecture Notes in Computer Science, 2253:316–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2253/22530316.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2253/22530316.
                                        pdf.

Tanaka:2001:NAM

[2370] Hideo Tanaka, Kazutomi Sugihara, and Yutaka Maeda. 39. nonadditive measures by interval probability functions. Lecture Notes in Computer Science, 2253:322–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2253/22530322.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2253/22530322.
                                        pdf.

Nguyen:2001:SCC

[2371] Ngoc Thanh Nguyen. 40. susceptibility to consensus of conflict profiles. Lecture Notes in Computer Science, 2253: 327–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2253/22530327.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2253/22530327.
pdf.

Nguyen:2001:AIS

[2372] Hung Son Nguyen, Andrzej Skowron, and Marcin S. Szczuka. 41. analysis of image sequences for the unmanned aerial vehicle. Lecture Notes in Computer Science, 2253:333–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2253/22530333.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2253/22530333.
pdf.

Maheswari:2001:VPRb

[2373] V. Uma Maheswari, Arul Siromoney, and K. M. Mehata. 42. the variable precision rough set inductive logic programming model and Web usage graphs. Lecture Notes in Computer Science, 2253:339–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2253/22530339.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2253/22530339.
pdf.

Entani:2001:OPW

[2374] Tomoe Entani, Hidetomo Ichihashi, and Hideo Tanaka. 43. optimistic priority weights with an interval comparison matrix. Lecture Notes in Computer Science, 2253:344–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2253/22530344.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2253/22530344.
                                        pdf.

Deja:2001:RST

[2375] Rafa Deja and Dominik Ślȩzak. 44. rough set theory in conflict analysis. Lecture Notes in Computer Science, 2253:349–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2253/22530349.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2253/22530349.
                                        pdf.

Liu:2001:DID

[2376] Chunnian Liu and Ning Zhong. 45. dealing with imperfect data by RS– ILP. Lecture Notes in Computer Science, 2253:354–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2253/22530354.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2253/22530354.pdf.

Skowron:2001:EPU

[2377] Andrzej Skowron, Jarosław Stepaniuk, and James F. Peters. 46. extracting patterns using information granules: a brief introduction. Lecture Notes in Computer Science, 2253:359– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2253/22530359.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2253/22530359.
pdf.

Slezak:2001:CMB

[2378] Dominik Ślȩzak. 47. classification models based on approximate Bayesian networks. Lecture Notes in Computer Science, 2253:364–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2253/22530364.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2253/22530364.
pdf.

Shinkawa:2001:IAC

[2379] Yoshiyuki Shinkawa and Masao J.

Matsumoto. 48. identifying adaptable components — A rough sets style approach. Lecture Notes in Computer Science, 2253:370–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2253/22530370.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2253/22530370.
pdf.

Pawlak:2001:RMI

[2380] Zdzisław Pawlak, James F. Peters, Andrzej Skowron, Z. Suraj, S. Ramanna, and M. Borkowski. 49. rough measures and integrals: a brief introduction. Lecture Notes in Computer Science, 2253:375–379, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2253/22530375.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2253/22530375.
                                        pdf.

Lin:2001:ARS

[2381] T. Y. Lin and Eric Louie. 50. association rules in semantically rich relations: Granular computing approach. Lecture Notes in Computer Science, 2253:380– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2253/22530380.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2253/22530380.
                                        pdf.

Murai:2001:NFG

[2382] Tetsuya Murai, Michinori Nakata, and Yoshiharu Sato. 51. A note on filtration and granular reasoning. Lecture Notes in Computer Science, 2253:385– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2253/22530385.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2253/22530385.
                                        pdf.

Murai:2001:NCL

[2383] Tetsuya Murai, Michinori Nakata, and Yoshiharu Sato. 52. A note on conditional logic and association rules. Lecture Notes in Computer Science, 2253: 390–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2253/22530390.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2253/22530390.
pdf.

Freeman:2001:ASI

[2384] Rachel L. Freeman, Jerzy W. Grzymala-Busse, Laura A. Riffel, and Stephen R. Schroeder. 53. analysis of selfinjurious behavior by the LERS data mining system. Lecture Notes in Computer Science, 2253:395–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2253/22530395.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2253/22530395.
pdf.

Hirano:2001:CMN

[2385] Shoji Hirano, Shusaku Tsumoto, Tomohiro Okuzaki, and Yutaka Hata.

54. A clustering method for nominal and numerical data based on rough set theory. Lecture Notes in Computer Science, 2253:400–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2253/22530400.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2253/22530400.
pdf.

Kanasugi:2001:DAR

[2386] Akinori Kanasugi. 55. A design of architecture for rough set processor. Lecture Notes in Computer Science, 2253: 406–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2253/22530406.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2253/22530406.
                                        pdf.

Ohsawa:2001:SCD

[2387] Yukio Ohsawa. 56. the scope of chance discovery. Lecture Notes in Computer Science, 2253:413–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2253/22530413.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2253/22530413.
                                        pdf.

McBurney:2001:CDU

[2388] Peter McBurney and Simon Parsons. 57. chance discovery using dialectical argumentation. Lecture Notes in Computer Science, 2253:414–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2253/22530414.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2253/22530414.
                                        pdf.

Prendinger:2001:MCC

[2389] Helmut Prendinger and Mitsuru Ishizuka. 58. methodological considerations on chance discovery. Lecture Notes in Computer Science, 2253:425– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2253/22530425.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2253/22530425.
pdf.

Matsumura:2001:FDC

[2390] Naohiro Matsumura, Yukio Ohsawa, and Mitsuru Ishizuka. 59. future directions of communities on the Web. Lecture Notes in Computer Science, 2253: 435–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2253/22530435.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2253/22530435.
pdf.

Matsuo:2001:DSW

[2391] Yutaka Matsuo, Yukio Ohsawa, and Mitsuru Ishizuka. 60. A document as a small world. Lecture Notes in Computer Science, 2253:444–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2253/22530444.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2253/22530444.
pdf.

Sunayama:2001:SSC

[2392] Wataru Sunayama and Masahiko Yachida. 61. support system for creative activity by information acquirement through Internet. Lecture Notes in Computer Science, 2253:449– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2253/22530449.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2253/22530449.
                                        pdf.

Shibata:2001:ASL

[2393] Hirohito Shibata and Koichi Hori. 62. an approach to support long-term creative thinking and its feasibility. Lecture Notes in Computer Science, 2253: 455–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2253/22530455.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2253/22530455.
                                        pdf.

Shoji:2001:CDC

[2394] Hiroko Shoji and Koichi Hori. 63. chance discovery by creative communicators observed in real shopping behavior. Lecture Notes in Computer Science, 2253:462–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2253/22530462.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2253/22530462.
                                        pdf.

Horiguchi:2001:RCD

[2395] Tomoya Horiguchi and Tsukasa Hirashima. 64. the role of counterexamples in discovery learning environment: Awareness of the chance for learning. Lecture Notes in Computer Science, 2253:468–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2253/22530468.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2253/22530468.
pdf.

Kato:2001:IDM

[2396] Yoshikiyo Kato, Takehisa Yairi, and Koichi Hori. 65. integrating data mining techniques and design information management for failure prevention. Lecture Notes in Computer Science, 2253:475–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2253/22530475.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2253/22530475.
pdf.

Ohsawa:2001:APD

[2397] Yukio Ohsawa and Yumiko Nara. 66.

action proposal as discovery of context. Lecture Notes in Computer Science, 2253:481–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2253/22530481.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2253/22530481.
pdf.

Nishimura:2001:RST

[2398] Takuichi Nishimura and Ryuichi Oka.

67. retrieval of similar time-series patterns for chance discovery. Lecture Notes in Computer Science, 2253:486– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2253/22530486.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2253/22530486.
                                        pdf.

Torra:2001:FKB

[2399] Vicenç Torra. 68. fuzzy knowledge based systems and chance discovery. Lecture Notes in Computer Science, 2253:491–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2253/22530491.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2253/22530491.
                                        pdf.

Washio:2001:JKC

[2400] Takashi Washio. 69. JSAI KDD challenge 2001: JKDD01. Lecture Notes in Computer Science, 2253:499– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2253/22530499.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2253/22530499.
                                        pdf.

Hatazawa:2001:KDS

[2401] Hiromitsu Hatazawa, Hidenao Abe, Mao Komori, Yoshiaki Tachibana, and Takahira Yamaguchi. 70. knowledge discovery support from a meningoencephalitis dataset using an automatic composition tool for inductive applications. Lecture Notes in Computer Science, 2253:500–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2253/22530500.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2253/22530500.
pdf.

Ho:2001:EMK

[2402] T. B. Ho, S. Kawasaki, and D. D.

Nguyen. 71. extracting meningitis knowledge by integration of rule induction and association mining. Lecture Notes in Computer Science, 2253:508– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2253/22530508.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2253/22530508.
pdf.

Ikeda:2001:BAM

[2403] Takayuki Ikeda, Takashi Washio, and Hiroshi Motoda. 72. basket analysis on meningitis data. Lecture Notes in Computer Science, 2253:516– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2253/22530516.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2253/22530516.
pdf.

Niimi:2001:EGP

[2404] Ayahiko Niimi and Eiichiro Tazaki. 73.

extended genetic programming using apriori algorithm for rule discovery. Lecture Notes in Computer Science, 2253:525–??, 2001. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2253/22530525.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2253/22530525.
                                        pdf.

Okada:2001:MKD

[2405] Takashi Okada. 74. medical knowledge discovery on the meningoencephalitis diagnosis studied by the cascade model. Lecture Notes in Computer Science, 2253:533–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2253/22530533.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2253/22530533.
                                        pdf.

Zhong:2001:MDM

[2406] Ning Zhong, Ju-Zhen Dong, and Setsuo Ohsuga. 75. meningitis data mining by cooperatively using GDT– RS and RSBR. Lecture Notes in Computer Science, 2253:541–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2253/22530541.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2253/22530541.
                                        pdf.

Anonymous:2001:AIbi

[2407] Anonymous. Author index. Lecture Notes in Computer Science, 2253:549– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2253/2253auth.pdf.

Anonymous:2001:SIe

[2408] Anonymous. Subject index. Lecture Notes in Computer Science, 2253:551– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2253/2253subject.pdf.

Garlan:2001:ADF

[2409] David Garlan. Aura: Distractionfree ubiquitous computing. Lecture Notes in Computer Science, 2254:1– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2254/22540001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2254/22540001.
pdf.

Greenberg:2001:SCI

[2410] Saul Greenberg. Supporting casual interaction between intimate collaborators. Lecture Notes in Computer Science, 2254:3–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2254/22540003.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2254/22540003.pdf.

Raskin:2001:TAI

[2411] Jef Raskin. Turning the art of interface design into engineering. Lecture Notes in Computer Science, 2254:5– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2254/22540005.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2254/22540005.
pdf.

Paterno:2001:TUI

[2412] Fabio Paternò. Towards a UML for interactive systems. Lecture Notes in Computer Science, 2254:7– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2254/22540007.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2254/22540007.
pdf.

Metzker:2001:IAS

[2413] Eduard Metzker and Michael Offergeld. An interdisciplinary approach for successfully integrating human-centered design methods into development processes practiced by industrial software development organizations. Lecture Notes in Computer Science, 2254:19– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2254/22540019.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2254/22540019.
pdf.

Kaindl:2001:USW

[2414] Hermann Kaindl and Rudolf Jezek. From usage scenarios to widget classes.

Lecture Notes in Computer Science, 2254:35–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2254/22540035.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2254/22540035.
pdf.

Bass:2001:ESA

[2415] Len Bass and Bonnie E. John. Evaluating software architectures for usability. Lecture Notes in Computer Science, 2254:37–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2254/22540037.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2254/22540037.
pdf.

Jambon:2001:ISS

[2416] Francis Jambon, Patrick Girard, and Yamine Aït-ameur. Interactive system safety and usability enforced with the development process. Lecture Notes in Computer Science, 2254:39– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2254/22540039.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2254/22540039.
pdf.

Curzon:2001:DMC

[2417] Paul Curzon and Ann Blandford. Detecting multiple classes of user errors. Lecture Notes in Computer Science, 2254:57–??, 2001. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2254/22540057.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2254/22540057.
                                        pdf.

Roussel:2001:ENU

[2418] Nicolas Roussel. Exploring new uses of video with VideoSpace. Lecture Notes in Computer Science, 2254:73– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2254/22540073.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2254/22540073.
                                        pdf.

Willans:2001:PPI

[2419] James S. Willans and Michael D. Harrison. Prototyping pre-implementation designs of virtual environment behaviour. Lecture Notes in Computer Science, 2254:91–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2254/22540091.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2254/22540091.pdf.

Grolaux:2001:QMD

[2420] Donatien Grolaux, Peter Van Roy, and Jean Vanderdonckt. QTk — A mixed declarative/procedural approach for designing executable user interfaces. Lecture Notes in Computer Science, 2254:109–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2254/22540109.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2254/22540109.
pdf.

Dubois:2001:CAR

[2421] Emmanuel Dubois, Laurence Nigay, and Jocelyne Troccaz. Consistency in augmented reality systems. Lecture Notes in Computer Science, 2254:111– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2254/22540111.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2254/22540111.
pdf.

Baker:2001:HEG

[2422] Kevin Baker, Saul Greenberg, and Carl Gutwin. Heuristic evaluation of groupware based on the mechanics of collaboration. Lecture Notes in Computer Science, 2254:123–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2254/22540123.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2254/22540123.
pdf.

Henninger:2001:OLM

[2423] Scott Henninger. An organizational learning method for applying usability guidelines and patterns. Lecture Notes in Computer Science, 2254:141– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2254/22540141.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2254/22540141.
                                        pdf.

Bergman:2001:PAD

[2424] Lawrence D. Bergman, Tatiana Kichkaylo, Guruduth Banavar, and Jeremy Sussman. Pervasive application development and the WYSIWYG pitfall. Lecture Notes in Computer Science, 2254:157–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2254/22540157.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2254/22540157.pdf.

Calvary:2001:URF

[2425] Gaëlle Calvary, Joëlle Coutaz, and David Thevenin. A unifying reference framework for the development of plastic user interfaces. Lecture Notes in Computer Science, 2254:173– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2254/22540173.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2254/22540173.
                                        pdf.

Zimmerman:2001:BUC

[2426] Guy Zimmerman, Julie Barnes, and Laura Leventhal. Building usercontrolled 3D models and animations for inherently-3D construction tasks: Which tool, which representation? Lecture Notes in Computer Science, 2254:193–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2254/22540193.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2254/22540193.
pdf.

Salzman:2001:UVC

[2427] Tim Salzman, Szymon Stachniak, and Wolfgang Stürzlinger. Unconstrained vs. constrained 3D scene manipulation. Lecture Notes in Computer Science, 2254:207–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2254/22540207.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2254/22540207.
pdf.

Kettebekov:2001:TNG

[2428] Sanshzar Kettebekov and Rajeev Sharma. Toward natural gesture/ speech control of a large display. Lecture Notes in Computer Science, 2254: 221–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2254/22540221.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2254/22540221.
pdf.

MacKenzie:2001:ETI

[2429] I. Scott MacKenzie and Shaidah Jusoh. An evaluation of two input devices for remote pointing. Lecture Notes in Computer Science, 2254:235– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2254/22540235.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2254/22540235.
                                        pdf.

MacGregor:2001:DMM

[2430] Carolyn MacGregor and Alice Thomas. Does multi-modal feedback help in everyday computing tasks? Lecture Notes in Computer Science, 2254:251– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2254/22540251.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2254/22540251.
                                        pdf.

Roth:2001:ISH

[2431] Jörg Roth. Information sharing with handheld appliances. Lecture Notes in Computer Science, 2254:263–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2254/22540263.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2254/22540263.
                                        pdf.

Gray:2001:DLM

[2432] Philip Gray and Meurig Sage. Dynamic links for mobile connected context-sensitive systems. Lecture Notes in Computer Science, 2254:281– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2254/22540281.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2254/22540281.
pdf.

Renevier:2001:MCA

[2433] Philippe Renevier and Laurence Nigay. Mobile collaborative augmented reality: The augmented stroll. Lecture Notes in Computer Science, 2254:299– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2254/22540299.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2254/22540299.
pdf.

Gray:2001:MUS

[2434] Philip Gray and Daniel Salber. Modelling and using sensed context information in the design of interactive applications. Lecture Notes in Computer Science, 2254:317–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2254/22540317.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2254/22540317.
pdf.

Choi:2001:DAW

[2435] Andrew Choi and Hanan Lutfiyya. Delivering adaptive Web content based on client computing resources. Lecture Notes in Computer Science, 2254:337– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2254/22540337.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2254/22540337.
                                        pdf.

Makarainen:2001:HCN

[2436] Minna Mäkäräinen, Johanna Tiitola, and Katja Konkka. How cultural needs affect user interface design? Lecture Notes in Computer Science, 2254:357– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2254/22540357.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2254/22540357.
                                        pdf.

Anonymous:2001:AIbj

[2437] Anonymous. Author index. Lecture Notes in Computer Science, 2254:359– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2254/2254auth.pdf.

Abts:2001:CBS

[2438] Chris Abts. COTS-based systems (CBS) functional density – A heuristic for better CBS design. Lecture Notes in Computer Science, 2255:1– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2255/22550001.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2255/22550001.
pdf.

Albert:2001:MCC

[2439] Cecilia Albert and Lisa Brownsword.

Meeting the challenges of commercialoff-the-shelf (COTS) products: The information technology solutions evolution process (ITSEP). Lecture Notes in Computer Science, 2255:10– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2255/22550010.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2255/22550010.
pdf.

Baker:2001:LLI

[2440] Thomas G. Baker. Lessons learned integrating COTS into systems. Lecture Notes in Computer Science, 2255:21– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2255/22550021.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2255/22550021.
pdf.

Ballurio:2001:RRC

[2441] Keith Ballurio, Betsy Scalzo, and Lou Rose. Risk reduction in COTS software selection with BASIS. Lecture Notes in Computer Science, 2255:31– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2255/22550031.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2255/22550031.
                                        pdf.

Benguria:2001:ECU

[2442] Gorka Benguria, Ana Belén García, David Sellier, and Sandy Tay. European COTS user working group: Analysis of the common problems and current practices of the European COTS users. Lecture Notes in Computer Science, 2255:44–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2255/22550044.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2255/22550044.
                                        pdf.

Burgues:2001:CSC

[2443] X. Burgués, C. Estay, X. Franch, J. A. Pastor, and C. Quer. Combined selection of COTS components. Lecture Notes in Computer Science, 2255:54– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2255/22550054.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2255/22550054.
                                        pdf.

Davis:2001:IEI

[2444] L. Davis and Rose Gamble. Identifying evolvability for integration. Lecture Notes in Computer Science, 2255:65– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2255/22550065.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2255/22550065.
pdf.

Dean:2001:IDS

[2445] John C. Dean and Li Li. Issues in developing security wrapper technology for COTS software products. Lecture Notes in Computer Science, 2255:76– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2255/22550076.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2255/22550076.
pdf.

Comella-Dorda:2001:PCS

[2446] Santiago Comella-Dorda, John C.

Dean, Edwin Morris, and Patricia Oberndorf. A process for COTS software product evaluation. Lecture Notes in Computer Science, 2255:86– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2255/22550086.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2255/22550086.
pdf.

Earl:2001:FHS

[2447] Anthony Earl. Five hurdles to the successful adoption of component-based COTS in a corporate setting. Lecture Notes in Computer Science, 2255:97– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2255/22550097.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2255/22550097.
                                        pdf.

Gao:2001:BTS

[2448] Jerry Gao, Kamal Gupta, Shalini Gupta, and Simon Shim. On building testable software components. Lecture Notes in Computer Science, 2255: 108–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2255/22550108.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2255/22550108.
                                        pdf.

Gorton:2001:SAP

[2449] Ian Gorton and Anna Liu. Streamlining the acquisition process for largescale COTS middleware components. Lecture Notes in Computer Science, 2255:122–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2255/22550122.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2255/22550122.
                                        pdf.

Gregor:2001:SPA

[2450] Sallie Gregor, Joseph Hutson, and Colleen Oresky. Storyboard process to assist in requirements verification and adaptation to capabilities inherent in COTS. Lecture Notes in Computer Science, 2255:132– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2255/22550132.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2255/22550132.
pdf.

ShallerHornstein:2001:RPC

[2451] Rhoda Shaller Hornstein and John K.

Willoughby. Realizing the potential for COTS utilization: a work in progress. Lecture Notes in Computer Science, 2255:142–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2255/22550142.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2255/22550142.
pdf.

Maiden:2001:RPG

[2452] N. A. M. Maiden, H. Kim, and Cornelius Ncube. Rethinking process guidance for selecting software components. Lecture Notes in Computer Science, 2255:151–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2255/22550151.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2255/22550151.
pdf.

Morisio:2001:DCC

[2453] Maurizio Morisio and Marco Torchiano. Definition and classification of COTS: a proposal. Lecture Notes in Computer Science, 2255:165–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2255/22550165.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2255/22550165.
                                        pdf.

Ncube:2001:LCD

[2454] Cornelius Ncube and John C. Dean. The limitations of current decisionmaking techniques in the procurement of COTS software components. Lecture Notes in Computer Science, 2255: 176–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2255/22550176.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2255/22550176.
                                        pdf.

Quang:2001:CBS

[2455] Nguyen Thanh Quang and Isabelle Comyn-Wattiau. COTS-based system engineering: The linguistics approach. Lecture Notes in Computer Science, 2255:188–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2255/22550188.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2255/22550188.
                                        pdf.

Payton:2001:MIS

[2456] Jamie Payton, Gerður Jónsdóttir, Daniel Flagg, and Rose Gamble. Merging integration solutions for architecture and security mismatch. Lecture Notes in Computer Science, 2255:199– 208, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2255/22550199.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2255/22550199.
pdf.

Pfarr:2001:ICG

[2457] Thomas Pfarr and James E. Reis.

The integration of COTS/GOTS within NASA’s HST command and control system. Lecture Notes in Computer Science, 2255:209–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2255/22550209.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2255/22550209.
pdf.

Seacord:2001:RCS

[2458] Robert C. Seacord. Replaceable components and the service provider interface. Lecture Notes in Computer Science, 2255:222–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2255/22550222.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2255/22550222.
pdf.

Semancik:2001:SAF

[2459] Susan K. Semancik and Annette M.

Conger. The standard autonomous file server, a customized, off-the-shelf success story. Lecture Notes in Computer Science, 2255:234–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2255/22550234.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2255/22550234.
                                        pdf.

Smith:2001:ILS

[2460] James D. Smith and Duane Hybertson. Implementing large-scale COTS reengineering within the United States Department of Defense. Lecture Notes in Computer Science, 2255:245–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2255/22550245.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2255/22550245.
                                        pdf.

Anonymous:2001:AIbk

[2461] Anonymous. Author index. Lecture Notes in Computer Science, 2255:257– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2255/2255auth.pdf.

Abbass:2001:MPE

[2462] H. A. Abbass. A memetic Pareto evolutionary approach to artificial neural networks. Lecture Notes in Computer Science, 2256:1–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2256/22560001.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2256/22560001.
pdf.

Antoniou:2001:RDD

[2463] Grigoris Antoniou and David Billington. Relating defeasible and default logic. Lecture Notes in Computer Science, 2256:13–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2256/22560013.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2256/22560013.pdf.

Rueda:2001:RMP

[2464] Luis Rueda and B. John Oommen. Resolving Minsky’s Paradox: The d-dimensional normal distribution case. Lecture Notes in Computer Science, 2256:25–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2256/22560025.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2256/22560025.
pdf.

Beaumont:2001:SOT

[2465] Matthew Beaumont, Abdul Sattar, Michael Maher, and John Thornton. Solving overconstrained temporal reasoning problems. Lecture Notes in Computer Science, 2256:37– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2256/22560037.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2256/22560037.
                                        pdf.

Beckwith:2001:HLI

[2466] Amy M. Beckwith, Berthe Y. Choueiry, and Hui Zou. How the level of interchangeability embedded in a finite constraint satisfaction problem affects the performance of search. Lecture Notes in Computer Science, 2256:50– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2256/22560050.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2256/22560050.
                                        pdf.

Boxer:2001:TLN

[2467] Paul A. Boxer. Towards learning naive physics by visual observation: Qualitative spatial representations. Lecture Notes in Computer Science, 2256:62– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2256/22560062.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2256/22560062.
                                        pdf.

Bracken:2001:PAS

[2468] N. J. Bracken and R. I. (Bob) McKay. Phased array sub-beam optimisation. Lecture Notes in Computer Science, 2256:71–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2256/22560071.htm;   http:
                                        //link.springer-ny.com/link/service/

series/0558/papers/2256/22560071.
pdf.

Bui:2001:GFE

[2469] The Duy Bui, Dirk Heylen, Mannes Poel, and Anton Nijholt. Generation of facial expressions from emotion using a fuzzy rule based system. Lecture Notes in Computer Science, 2256:83– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2256/22560083.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2256/22560083.
pdf.

Byun:2001:EFC

[2470] Yungcheol Byun, Sungsoo Yoon, Yeongwoo Choi, Gyeonghwan Kim, and Yillbyung Lee. An efficient form classification method using partial matching. Lecture Notes in Computer Science, 2256:95–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2256/22560095.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2256/22560095.
pdf.

Cameron-Jones:2001:GDS

[2471] Mike Cameron-Jones. Gradient descent style leveraging of decision trees and stumps for misclassification cost performance. Lecture Notes in Computer Science, 2256:107–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2256/22560107.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2256/22560107.
                                        pdf.

Chopra:2001:PPD

[2472] Samir Chopra and Dongmo Zhang. Postdiction problems in dynamic logic. Lecture Notes in Computer Science, 2256:119–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2256/22560119.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2256/22560119.
                                        pdf.

Corbett:2001:MRO

[2473] Dan Corbett. A method for reasoning with ontologies represented as conceptual graphs. Lecture Notes in Computer Science, 2256:130–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2256/22560130.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2256/22560130.
                                        pdf.

Debenham:2001:STD

[2474] John Debenham. Selection of tasks and delegation of responsibility in a multiagent system for emergent process management. Lecture Notes in Computer Science, 2256:142–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2256/22560142.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2256/22560142.
pdf.

Foo:2001:SMR

[2475] Norman Foo, Abhaya Nayak, Maurice Pagnucco, and Dongmo Zhang. State minimization re-visited. Lecture Notes in Computer Science, 2256:153– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2256/22560153.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2256/22560153.
pdf.

French:2001:DQP

[2476] Tim French. Decidability of quantified propositional branching time logics. Lecture Notes in Computer Science, 2256:165–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2256/22560165.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2256/22560165.
pdf.

Kee:2001:ITI

[2477] Gyundo Kee, Yungcheol Byun, Kwanyong Lee, and Yillbyung Lee. Improved techniques for an iris recognition system with high performance. Lecture Notes in Computer Science, 2256:177– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2256/22560177.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2256/22560177.
                                        pdf.

Kendall:2001:IAP

[2478] Graham Kendall and Mark Willdig. An investigation of an adaptive poker player. Lecture Notes in Computer Science, 2256:189–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2256/22560189.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2256/22560189.pdf.

Khan:2001:AAK

[2479] Abdus Salam Khan and Achim Hoffmann. Acquiring adaptation knowledge for CBR with MIKAS. Lecture Notes in Computer Science, 2256:201– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2256/22560201.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2256/22560201.
                                        pdf.

Kim:2001:RTR

[2480] Jong Bae Kim, Hye Sun Park, Min Ho Park, and Hang Joon Kim. A realtime region-based motion segmentation using adaptive thresholding and K-means clustering. Lecture Notes in Computer Science, 2256:213–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2256/22560213.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2256/22560213.
pdf.

Kim:2001:ACC

[2481] Kyungsu Kim, Jeonghee Jeon, WanKyoo Choi, Pankoo Kim, and Yo-Sung Ho. Automatic cell classification in human’s peripheral blood images based on morphological image processing. Lecture Notes in Computer Science, 2256:225–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2256/22560225.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2256/22560225.
pdf.

Kim:2001:FCA

[2482] Mihye Kim and Paul Compton. Formal concept analysis for domain-specific document retrieval systems. Lecture Notes in Computer Science, 2256:237– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2256/22560237.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2256/22560237.
pdf.

Kowalczyk:2001:LSA

[2483] Adam Kowalczyk and Bhavani Raskutti.

Learner’s self-assessment: a case study of SVM for information retrieval. Lecture Notes in Computer Science, 2256: 249–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2256/22560249.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2256/22560249.
                                        pdf.

Kryssanov:2001:SMC

[2484] Victor V. Kryssanov, Masayuki Okabe, Koh Kakusho, and Michihiko Minoh. A semiotic model of communication and its implications for the digital city development. Lecture Notes in Computer Science, 2256:261– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2256/22560261.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2256/22560261.
                                        pdf.

Hoffmann:2001:SCK

[2485] Achim Hoffmann, Rex Kwok, and Paul Compton. Simulations for comparing knowledge acquisition and machine learning. Lecture Notes in Computer Science, 2256:273–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2256/22560273.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2256/22560273.
                                        pdf.

Hulme:2001:AGA

[2486] Daniel Hulme and Shuxiang Xu. Application of genetic algorithms to the optimisation of neural network configuration for stock market forecasting. Lecture Notes in Computer Science, 2256:

285–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2256/22560285.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2256/22560285.
pdf.

Lang:2001:UES

[2487] Jérôme Lang, Pierre Marquis, and Mary-Anne Williams. Updating epistemic states. Lecture Notes in Computer Science, 2256:297–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2256/22560297.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2256/22560297.
pdf.

Lee:2001:FTC

[2488] Michael D. Lee. Fast text classification using sequential sampling processes. Lecture Notes in Computer Science, 2256:309–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2256/22560309.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2256/22560309.
pdf.

Lister:2001:AMC

[2489] Kendall Lister and Leon Sterling.

Agents in a multi-cultural world: Towards ontological reconciliation. Lecture Notes in Computer Science, 2256: 321–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2256/22560321.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2256/22560321.
                                        pdf.

Liu:2001:LFR

[2490] Chuchang Liu. Logical foundations for reasoning about trust in secure digital communication. Lecture Notes in Computer Science, 2256:333–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2256/22560333.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2256/22560333.
                                        pdf.

Monekosso:2001:PQP

[2491] Ndedi Monekosso and Paolo Remagnino. Phe-Q: a pheromone based Q-learning. Lecture Notes in Computer Science, 2256:345–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2256/22560345.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2256/22560345.
                                        pdf.

Murakami:2001:SKP

[2492] Tomoko Murakami, Ryohei Orihara, and Naomichi Sueda. Specification of Kansei patterns in an adaptive perceptual space. Lecture Notes in Computer Science, 2256:356–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2256/22560356.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2256/22560356.
pdf.

Norling:2001:EJA

[2493] Emma Norling and Frank E. Ritter.

Embodying the JACK agent architecture. Lecture Notes in Computer Science, 2256:368–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2256/22560368.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2256/22560368.pdf.

Nowak:2001:TOP

[2494] Chris Nowak and Richard Raban. Towards an ontology of part-of. Lecture Notes in Computer Science, 2256:378– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2256/22560378.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2256/22560378.
pdf.

Padmanabhan:2001:AME

[2495] Vineet Padmanabhan, Guido Governatori, and Abdul Sattar. Actions made explicit in BDI. Lecture Notes in Computer Science, 2256:390– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2256/22560390.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2256/22560390.
                                        pdf.

Pfahringer:2001:WBA

[2496] Bernhard Pfahringer, Geoffrey Holmes, and Gabi Schmidberger. Wrapping boosters against noise. Lecture Notes in Computer Science, 2256:402–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2256/22560402.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2256/22560402.
                                        pdf.

Reynolds:2001:CTM

[2497] Mark Reynolds. Continuous temporal models. Lecture Notes in Computer Science, 2256:414–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2256/22560414.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2256/22560414.pdf.

Saba:2001:LCK

[2498] Walid S. Saba. Language and commonsense knowledge. Lecture Notes in Computer Science, 2256:426–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2256/22560426.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2256/22560426.
                                        pdf.

Seo:2001:SNS

[2499] Hee Suk Seo and Tae Ho Cho. Simulation of network security with collaboration among IDS models. Lecture Notes in Computer Science, 2256:438– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2256/22560438.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2256/22560438.
pdf.

Starkie:2001:PSD

[2500] Bradford Starkie. Programming spoken dialogs using grammatical inference. Lecture Notes in Computer Science, 2256:449–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2256/22560449.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2256/22560449.
pdf.

Song:2001:TGP

[2501] Andy Song, Thomas Loveard, and Vic Ciesielski. Towards genetic programming for texture classification. Lecture Notes in Computer Science, 2256:461– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2256/22560461.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2256/22560461.
pdf.

Stumptner:2001:UDI

[2502] Markus Stumptner. Using design information to identify structural software faults. Lecture Notes in Computer Science, 2256:473–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2256/22560473.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2256/22560473.
pdf.

Szarowicz:2001:AAA

[2503] Adam Szarowicz, Juan Amiguet-Vercher, Peter Forte, Jonathan Briggs, Petros Gelepithis, and Paolo Remagnino. The application of AI to automatically generated animation. Lecture Notes in Computer Science, 2256: 487–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2256/22560487.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2256/22560487.
pdf.

Thielscher:2001:PNA

[2504] Michael Thielscher. Planning with noisy actions. Lecture Notes in Computer Science, 2256:495–??, 2001. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.com/link/
service/series/0558/bibs/2256/22560495.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2256/22560495.pdf.

Versteeg:2001:GJD

[2505] Steven Versteeg and Alan Blair.

Getting the job done in a hostile environment. Lecture Notes in Computer Science, 2256:507–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2256/22560507.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2256/22560507.
pdf.

Vo:2001:SQP

[2506] Quoc Bao Vo and Norman Y. Foo.

Solving the qualification problem. Lecture Notes in Computer Science, 2256: 519–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2256/22560519.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2256/22560519.
pdf.

Wallis:2001:DMC

[2507] Peter Wallis, Helen Mitchard, Jyotsna Das, and Damian O’Dea. Dialogue modelling for a conversational agent. Lecture Notes in Computer Science, 2256:532–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2256/22560532.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2256/22560532.
pdf.

Webb:2001:CEC

[2508] Geoffrey I. Webb. Candidate elimination criteria for lazy Bayesian rules. Lecture Notes in Computer Science, 2256:545–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2256/22560545.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2256/22560545.
                                        pdf.

Winikoff:2001:SDI

[2509] Michael Winikoff, Lin Padgham, and James Harland. Simplifying the development of intelligent agents. Lecture Notes in Computer Science, 2256:557– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2256/22560557.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2256/22560557.
                                        pdf.

Wobcke:2001:OSP

[2510] Wayne Wobcke. An operational semantics for a PRS-like agent architecture. Lecture Notes in Computer Science, 2256:569–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2256/22560569.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2256/22560569.
                                        pdf.

Yoon:2001:MUD

[2511] Sungsoo Yoon, Yungcheol Byun, Gyeonghwan Kim, Yeongwoo Choi, and Yillbyung Lee. A model of unconstrained digit recognition based on hypothesis testing and data reconstruction. Lecture Notes in Computer Science, 2256:581–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2256/22560581.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2256/22560581.
pdf.

Zhang:2001:CQD

[2512] Chengqi Zhang and Shichao Zhang.

Collecting quality data for database mining. Lecture Notes in Computer Science, 2256:593–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2256/22560593.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2256/22560593.pdf.

Zhang:2001:FPE

[2513] Songmao Zhang and Geoffrey I. Webb.

Further pruning for efficient association rule discovery. Lecture Notes in Computer Science, 2256:605–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2256/22560605.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2256/22560605.
pdf.

Zhang:2001:PDP

[2514] Shichao Zhang and Chengqi Zhang.

Pattern discovery in probabilistic databases. Lecture Notes in Computer Science, 2256:619–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2256/22560619.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2256/22560619.pdf.

Zhang:2001:CLP

[2515] Yan Zhang. The complexity of logic program updates. Lecture Notes in Computer Science, 2256:631–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2256/22560631.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2256/22560631.
                                        pdf.

Zhang:2001:RMI

[2516] Zili Zhang and Chengqi Zhang. Reasoning with multimedia information using symbolic projection. Lecture Notes in Computer Science, 2256:643– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2256/22560643.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2256/22560643.
                                        pdf.

Dubois:2001:CRB

[2517] Didier Dubois. Confidence relations as a basis for uncertainty modeling, plausible reasoning, and belief revision. Lecture Notes in Computer Science, 2256:

655–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2256/22560655.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2256/22560655.
pdf.

Anonymous:2001:AIbl

[2518] Anonymous. Author index. Lecture Notes in Computer Science, 2256:665– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2256/2256auth.pdf.

Meadows:2001:UDL

[2519] Catherine Meadows. Using a declarative language to build an experimental analysis tool. Lecture Notes in Computer Science, 2257:1–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2257/22570001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2257/22570001.
pdf.

Dahl:2001:HTY

[2520] Veronica Dahl. How to talk to your computer so that it will listen. Lecture Notes in Computer Science, 2257: 3–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2257/22570003.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2257/22570003.
                                        pdf.

Boyer:2001:STO

[2521] Robert S. Boyer and J. Strother Moore. Single-threaded objects in ACL2. Lecture Notes in Computer Science, 2257: 9–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2257/22570009.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2257/22570009.
                                        pdf.

Jayaraman:2001:MES

[2522] Bharat Jayaraman and Pallavi Tambay. Modeling engineering structures with constrained objects. Lecture Notes in Computer Science, 2257:28– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2257/22570028.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2257/22570028.
                                        pdf.

Liang:2001:CCH

[2523] Chuck C. Liang. Compiler construction in higher order logic programming. Lecture Notes in Computer Science, 2257: 47–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2257/22570047.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2257/22570047.
                                        pdf.

Torgersson:2001:DPC

[2524] Olof Torgersson. Declarative programming and clinical medicine. Lecture Notes in Computer Science, 2257:64– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2257/22570064.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2257/22570064.
pdf.

Gupta:2001:SBF

[2525] Gopal Gupta, Hai-Feng Guo, Arthur I.

Karshmer, Enrico Pontelli, Juan Raymundo Iglesias, Desh Ranjan, Brook Milligan, Nayana Datta, Omar El Khatib, Mohammed Noamany, and Xinhong Zhou. Semantics-based filtering: Logic programming’s killer app? Lecture Notes in Computer Science, 2257:82–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2257/22570082.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2257/22570082.
pdf.

Johansson:2001:LSR

[2526] Erik Johansson and Konstantinos Sagonas. Linear scan register allocation in a high-performance Erlang compiler. Lecture Notes in Computer Science, 2257:101–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2257/22570101.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2257/22570101.
                                        pdf.

Rhiger:2001:CEP

[2527] Morten Rhiger. Compiling embedded programs to byte code. Lecture Notes in Computer Science, 2257:120– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2257/22570120.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2257/22570120.
                                        pdf.

Lammel:2001:TCG

[2528] Ralf Lämmel and Joost Visser. Typed combinators for generic traversal. Lecture Notes in Computer Science, 2257: 137–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2257/22570137.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2257/22570137.
                                        pdf.

Wan:2001:EDF

[2529] Zhanyong Wan, Walid Taha, and Paul Hudak. Event-driven FRP. Lecture Notes in Computer Science, 2257:155– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2257/22570155.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2257/22570155.
                                        pdf.

Erwig:2001:AAO

[2530] Martin Erwig and Margaret Burnett.

Adding apples and oranges. Lecture Notes in Computer Science, 2257:173– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2257/22570173.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2257/22570173.
pdf.

Thiemann:2001:WCS

[2531] Peter Thiemann. WASH/CGI: Serverside Web scripting with sessions and typed, compositional forms. Lecture Notes in Computer Science, 2257:192– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2257/22570192.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2257/22570192.
pdf.

Kiselyov:2001:BXP

[2532] Oleg Kiselyov. A better XML parser through functional programming. Lecture Notes in Computer Science, 2257: 209–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2257/22570209.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2257/22570209.
pdf.

Karczmarczuk:2001:FAT

[2533] Jerzy Karczmarczuk. Functional approach to texture generation. Lecture Notes in Computer Science, 2257:225– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2257/22570225.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2257/22570225.
pdf.

Gallagher:2001:AIN

[2534] John P. Gallagher and Germán Puebla. Abstract interpretation over non-deterministic finite tree automata for set-based analysis of logic programs. Lecture Notes in Computer Science, 2257:243–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2257/22570243.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2257/22570243.
pdf.

Shen:2001:HLG

[2535] Kish Shen, Joachim Schimpf, Stefano Novello, and Josh Singer. A high-level generic interface to external programming languages for ECLiPSe. Lecture Notes in Computer Science, 2257:262– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2257/22570262.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2257/22570262.
pdf.

Bunus:2001:DSD

[2536] Peter Bunus and Peter Fritzson. A debugging scheme for declarative equation based modeling languages. Lecture Notes in Computer Science, 2257: 280–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2257/22570280.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2257/22570280.
pdf.

Vandeginste:2001:SOP

[2537] Ruben Vandeginste, Konstantinos Sagonas, and Bart Demoen. Segment order preserving and generational garbage collection for prolog. Lecture Notes in Computer Science, 2257:299–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2257/22570299.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2257/22570299.
pdf.

Yang:2001:EEC

[2538] Rong Yang and Steve Gregory. Exploiting efficient control and data structures in logic programs. Lecture Notes in Computer Science, 2257:318– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2257/22570318.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2257/22570318.
pdf.

Castro:2001:SRC

[2539] Luís F. Castro, Terrance Swift, and David S. Warren. Suspending and resuming computations in engines for SLG evaluation. Lecture Notes in Computer Science, 2257:332–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2257/22570332.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2257/22570332.
                                        pdf.

Anonymous:2001:AIbm

[2540] Anonymous. Author index. Lecture Notes in Computer Science, 2257:351– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2257/2257ind.pdf.

Bueno:2001:APC

[2541] Francisco Bueno and the CLIP Group. Agent programming in ciao prolog. Lecture Notes in Computer Science, 2258:1–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2258/22580001.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2258/22580001.
                                        pdf.

Flach:2001:MRD

[2542] Peter A. Flach. Multi-relational data mining: a perspective. Lecture Notes in Computer Science, 2258:3– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2258/22580003.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2258/22580003.
pdf.

Dhar:2001:CGO

[2543] Vasant Dhar. A comparison of GLOWER and other machine learning methods for investment decision making. Lecture Notes in Computer Science, 2258:5–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2258/22580005.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2258/22580005.
pdf.

Amado:2001:PID

[2544] Nuno Amado, João Gama, and Fernando Silva. Parallel implementation of decision tree learning algorithms. Lecture Notes in Computer Science, 2258: 6–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2258/22580006.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2258/22580006.
pdf.

Brazdil:2001:RRC

[2545] Pavel Brazdil, Carlos Soares, and Rui Pereira. Reducing rankings of classifiers by eliminating redundant classifiers. Lecture Notes in Computer Science, 2258:14–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2258/22580014.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2258/22580014.
                                        pdf.

Ferrer-Troyano:2001:NPN

[2546] Francisco J. Ferrer-Troyano, Jesús S. Aguilar-Ruiz, and José C. Riquelme. Non-parametric nearest neighbor with local adaptation. Lecture Notes in Computer Science, 2258:22–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2258/22580022.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2258/22580022.
                                        pdf.

Gamallo:2001:SRA

[2547] Pablo Gamallo, Alexandre Agustini, and Gabriel P. Lopes. Selection restrictions acquisition from corpora. Lecture Notes in Computer Science, 2258:30– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2258/22580030.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2258/22580030.
                                        pdf.

Jovanoski:2001:CRL

[2548] Viktor Jovanoski and Nada Lavrač. Classification rule learning with APRIORI-C. Lecture Notes in Computer Science, 2258:44–??, 2001. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer-ny.com/link/

service/series/0558/bibs/2258/22580044.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2258/22580044.pdf.

Nascimento:2001:PMF

[2549] S. Nascimento, B. Mirkin, and F. Moura-Pires. Proportional membership in fuzzy clustering as a model of ideal types. Lecture Notes in Computer Science, 2258:52–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2258/22580052.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2258/22580052.
pdf.

Mayer:2001:ECS

[2550] Helmut A. Mayer and Roland Schwaiger.

Evolution of cubic spline activation functions for artificial neural networks. Lecture Notes in Computer Science, 2258:63–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2258/22580063.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2258/22580063.
pdf.

Silva:2001:MDC

[2551] Joaquim Silva, João Mexia, Carlos A.

Coelho, and Gabriel Lopes. Multilingual document clustering, topic extraction and data transformations. Lecture Notes in Computer Science, 2258:74– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2258/22580074.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2258/22580074.
                                        pdf.

Soares:2001:SBR

[2552] Carlos Soares, Johann Petrak, and Pavel Brazdil. Sampling-based relative landmarks: Systematically test-driving algorithms before choosing. Lecture Notes in Computer Science, 2258:88– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2258/22580088.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2258/22580088.
                                        pdf.

Tapia:2001:RAE

[2553] Elizabeth Tapia, José Carlos González, Javier García-Villalba, and Julio Villena. Recursive adaptive ECOC models. Lecture Notes in Computer Science, 2258:96–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2258/22580096.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2258/22580096.pdf.

Torgo:2001:SEC

[2554] Luis Torgo. A study on endcut preference in least squares regression trees. Lecture Notes in Computer Science, 2258:104–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2258/22580104.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2258/22580104.
pdf.

Almeida:2001:UDK

[2555] Pedro de Almeida and Luís Torgo.

The use of domain knowledge in feature construction for financial time series prediction. Lecture Notes in Computer Science, 2258:116–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2258/22580116.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2258/22580116.
pdf.

Hellstrom:2001:OSR

[2556] Thomas Hellström. Optimizing the Sharpe ratio for a rank based trading system. Lecture Notes in Computer Science, 2258:130–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2258/22580130.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2258/22580130.
pdf.

Antunes:2001:CKA

[2557] Luis Antunes, João Faria, and Helder Coelho. Choice: The key for autonomy. Lecture Notes in Computer Science, 2258:142–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2258/22580142.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2258/22580142.
                                        pdf.

Bourne:2001:DEC

[2558] Rachel A. Bourne, Karen Shoop, and Nicholas R. Jennings. Dynamic evaluation of coordination mechanisms for autonomous agents. Lecture Notes in Computer Science, 2258:155–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2258/22580155.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2258/22580155.
                                        pdf.

DellAcqua:2001:EMA

[2559] Pierangelo Dell’Acqua, João Alexandre Leite, and Luís Moniz Pereira. Evolving multi-agent viewpoints — an architecture. Lecture Notes in Computer Science, 2258:169–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2258/22580169.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2258/22580169.
                                        pdf.

DellAcqua:2001:EAU

[2560] Pierangelo Dell’Acqua and Luís Moniz Pereira. Enabling agents to update their knowledge and to prefer. Lecture Notes in Computer Science, 2258:183– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2258/22580183.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2258/22580183.
pdf.

Dignum:2001:MAS

[2561] Virginia Dignum and Frank Dignum.

Modelling agent societies: Coordination frameworks and institutions. Lecture Notes in Computer Science, 2258:191–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2258/22580191.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2258/22580191.pdf.

Malheiro:2001:ADB

[2562] Benedita Malheiro and Eugénio Oliveira. Argumentation as distributed belief revision: Conflict resolution in decentralised co-operative multi-agent systems. Lecture Notes in Computer Science, 2258:205–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2258/22580205.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2258/22580205.
pdf.

Reis:2001:SRS

[2563] Joaquim Reis and Nuno Mamede.

Scheduling, re-scheduling and communication in the multi-agent extended enterprise environment. Lecture Notes in Computer Science, 2258:

219–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2258/22580219.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2258/22580219.
                                        pdf.

Rocha:2001:EIF

[2564] Ana Paula Rocha and Eugénio Oliveira. Electronic institutions as a framework for agents’ negotiation and mutual commitment. Lecture Notes in Computer Science, 2258:232– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2258/22580232.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2258/22580232.
                                        pdf.

Trajkovski:2001:IBA

[2565] Goran Trajkovski. An imitationbased approach to modeling homogeneous agents societies. Lecture Notes in Computer Science, 2258:246– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2258/22580246.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2258/22580246.
                                        pdf.

Blackburn:2001:SCH

[2566] Patrick Blackburn, Jaap Kamps, and Maarten Marx. Situation calculus as hybrid logic: First steps. Lecture Notes in Computer Science, 2258:253– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2258/22580253.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2258/22580253.
pdf.

Leite:2001:MSL

[2567] João Alexandre Leite. A modified semantics for LUPS. Lecture Notes in Computer Science, 2258:261–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2258/22580261.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2258/22580261.
pdf.

Leite:2001:UMD

[2568] João Alexandre Leite, José Júlio Alferes, and Luís Moniz Pereira. On the use of multi-dimensional dynamic logic programming to represent societal agents’ viewpoints. Lecture Notes in Computer Science, 2258:276– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2258/22580276.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2258/22580276.
pdf.

Medina:2001:PSM

[2569] Jesús Medina, Manuel Ojeda-Aciego, and Peter Vojtás. A procedural semantics for multi-adjoint logic programming. Lecture Notes in Computer Science, 2258:290–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2258/22580290.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2258/22580290.
                                        pdf.

Pacheco:2001:RRT

[2570] Julio Pacheco, M. a Teresa Escrig, and Francisco Toledo. Representing and reasoning on three-dimensional qualitative orientation point objects. Lecture Notes in Computer Science, 2258: 298–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2258/22580298.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2258/22580298.
                                        pdf.

Pearce:2001:EEL

[2571] David Pearce, Hans Tompits, and Stefan Woltran. Encodings for equilibrium logic and logic programs with nested expressions. Lecture Notes in Computer Science, 2258:306–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2258/22580306.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2258/22580306.
                                        pdf.

Broda:2001:CFG

[2572] Sabine Broda and Luís Damas. A context-free grammar representation for normal inhabitants of types in TA_{λ}. Lecture Notes in Computer Science, 2258:321–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2258/22580321.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2258/22580321.
pdf.

Cravo:2001:PBR

[2573] Maria R. Cravo, João P. Cachopo, Ana C. Cachopo, and João P. Martins. Permissive belief revision. Lecture Notes in Computer Science, 2258: 335–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2258/22580335.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2258/22580335.
pdf.

Cruz:2001:GHC

[2574] Jorge Cruz and Pedro Barahona.

Global hull consistency with local search for continuous constraint solving. Lecture Notes in Computer Science, 2258:349–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2258/22580349.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2258/22580349.pdf.

Lynce:2001:TPC

[2575] Inês Lynce, Luís Baptista, and João Marques-Silva. Towards provably complete stochastic search algorithms for satisfiability. Lecture Notes in Computer Science, 2258:363–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2258/22580363.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2258/22580363.
                                        pdf.

Oliveira:2001:CCB

[2576] Elias Oliveira and Barbara M. Smith. A combined constraint-based search method for single-track railway scheduling problem. Lecture Notes in Computer Science, 2258:371– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2258/22580371.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2258/22580371.
                                        pdf.

Garrido:2001:TPS

[2577] Antonio Garrido, Eva Onaindía, and Federico Barber. A temporal planning system for time-optimal planning. Lecture Notes in Computer Science, 2258: 379–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2258/22580379.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2258/22580379.
                                        pdf.

Onaindia:2001:SEM

[2578] Eva Onaindia, Oscar Sapena, Laura Sebastia, and Eliseo Marzal. SimPlanner: An execution-monitoring system for replanning in dynamic worlds. Lecture Notes in Computer Science, 2258: 393–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2258/22580393.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2258/22580393.
pdf.

Prasad:2001:HHK

[2579] Bhanu Prasad. Hybrid hierarchical knowledge organization for planning. Lecture Notes in Computer Science, 2258:401–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2258/22580401.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2258/22580401.
pdf.

Sebastia:2001:SOS

[2580] L. Sebastia, E. Onaindia, and E. Marzal. STeLLa : An optimal sequential and parallel planner. Lecture Notes in Computer Science, 2258:409– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2258/22580409.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2258/22580409.
pdf.

Anonymous:2001:AIbn

[2581] Anonymous. Author index. Lecture Notes in Computer Science, 2258:417– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2258/2258auth.pdf.

Fluhrer:2001:WKS

[2582] Scott Fluhrer, Itsik Mantin, and Adi Shamir. Weaknesses in the key scheduling algorithm of RC4. Lecture Notes in Computer Science, 2259:1– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2259/22590001.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2259/22590001.
                                        pdf.

Hawkes:2001:PCS

[2583] Philip Hawkes, Frank Quick, and Gregory G. Rose. A practical cryptanalysis of SSC2. Lecture Notes in Computer Science, 2259:25–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2259/22590025.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2259/22590025.
                                        pdf.

Fluhrer:2001:AES

[2584] Scott Fluhrer and Stefan Lucks. Analysis of the E_{0} encryption system. Lecture Notes in Computer Science, 2259: 38–48, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2259/22590038.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2259/22590038.
pdf.

Youssef:2001:BFL

[2585] Amr Youssef and Guang Gong.

Boolean functions with large distance to all bijective monomials: N odd case. Lecture Notes in Computer Science, 2259:49–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2259/22590049.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2259/22590049.
pdf.

Pasalic:2001:LCC

[2586] Enes Pasalic and Subhamoy Maitra.

Linear codes in constructing resilient functions with high nonlinearity. Lecture Notes in Computer Science, 2259: 60–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2259/22590060.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2259/22590060.
pdf.

Iwata:2001:NCR

[2587] Tetsu Iwata, Takayuki Yoshiwara, and Kaoru Kurosawa. New covering radius of Reed–Muller codes for t-resilient functions. Lecture Notes in Computer Science, 2259:75–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2259/22590075.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2259/22590075.
                                        pdf.

DArco:2001:GZZ

[2588] Paolo D’Arco and Douglas Stinson. Generalized zig-zag functions and oblivious transfer reductions. Lecture Notes in Computer Science, 2259:87– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2259/22590087.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2259/22590087.
                                        pdf.

Ferguson:2001:SAR

[2589] Niels Ferguson, Richard Schroeppel, and Doug Whiting. A simple algebraic representation of Rijndael. Lecture Notes in Computer Science, 2259: 103–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2259/22590103.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2259/22590103.
                                        pdf.

Keliher:2001:IUB

[2590] Liam Keliher, Henk Meijer, and Stafford Tavares. Improving the upper bound on the maximum average linear hull probability for Rijndael. Lecture Notes in Computer Science, 2259: 112–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2259/22590112.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2259/22590112.
pdf.

Kiayias:2001:PRB

[2591] Aggelos Kiayias and Moti Yung. Polynomial reconstruction based cryptography. Lecture Notes in Computer Science, 2259:129–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2259/22590129.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2259/22590129.pdf.

King:2001:IIE

[2592] Brian King. An improved implementation of elliptic curves over GF (2) when using projective point arithmetic. Lecture Notes in Computer Science, 2259: 134–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2259/22590134.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2259/22590134.
pdf.

Coron:2001:FGP

[2593] Jean-Sébastien Coron, David M’Raïhi, and Christophe Tymen. Fast generation of pairs (k, [k]P) for Koblitz elliptic curves. Lecture Notes in Computer Science, 2259:151–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2259/22590151.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2259/22590151.
                                        pdf.

Moller:2001:AME

[2594] Bodo Möller. Algorithms for multi-exponentiation. Lecture Notes in Computer Science, 2259:165–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2259/22590165.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2259/22590165.
                                        pdf.

Hess:2001:TTH

[2595] Florian Hess, Gadiel Seroussi, and Nigel P. Smart. Two topics in hyperelliptic cryptography. Lecture Notes in Computer Science, 2259:181–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2259/22590181.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2259/22590181.
                                        pdf.

Raddum:2001:DAR

[2596] Håvard Raddum and Lars R. Knudsen. A differential attack on reduced-round SC2000. Lecture Notes in Computer Science, 2259:190–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2259/22590190.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2259/22590190.
                                        pdf.

Junod:2001:CMA

[2597] Pascal Junod. On the complexity of Matsui’s attack. Lecture Notes in Computer Science, 2259:199–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2259/22590199.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2259/22590199.
pdf.

Kuhn:2001:RWR

[2598] Fabian Kuhn and René Struik. Random walks revisited: Extensions of Pollard’s rho algorithm for computing multiple discrete logarithms. Lecture Notes in Computer Science, 2259:212– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2259/22590212.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2259/22590212.
pdf.

Reyhani-Masoleh:2001:FNB

[2599] Arash Reyhani-Masoleh and M. Anwar Hasan. Fast normal basis multiplication using general purpose processors. Lecture Notes in Computer Science, 2259:230–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2259/22590230.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2259/22590230.
pdf.

Khachatrian:2001:FMI

[2600] Gurgen H. Khachatrian, Melsik K. Kuregian, Karen R. Ispiryan, and James L. Massey. Fast multiplication of integers for publickey applications. Lecture Notes in Computer Science, 2259:245–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2259/22590245.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2259/22590245.
                                        pdf.

Akishita:2001:FSS

[2601] Toru Akishita. Fast simultaneous scalar multiplication on elliptic curve with Montgomery form. Lecture Notes in Computer Science, 2259:255–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2259/22590255.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2259/22590255.
                                        pdf.

Sakai:2001:PMS

[2602] Yasuyuki Sakai and Kouichi Sakurai. On the power of multidoubling in speeding up elliptic scalar multiplication. Lecture Notes in Computer Science, 2259:268–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2259/22590268.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2259/22590268.
pdf.

Gong:2001:GPK

[2603] Guang Gong, Lein Harn, and Huapeng Wu. The GH publickey cryptosystem. Lecture Notes in Computer Science, 2259:284–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2259/22590284.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2259/22590284.
pdf.

Lim:2001:XEG

[2604] Seongan Lim, Seungjoo Kim, Ikkwon Yie, Jaemoon Kim, and Hongsub Lee. XTR extended to GF(p^{6m}). Lecture Notes in Computer Science, 2259:301– 312, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2259/22590301.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2259/22590301.
pdf.

Nguyen:2001:TFL

[2605] Phong Q. Nguyen. The two faces of lattices in cryptology. Lecture Notes in Computer Science, 2259:313– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2259/22590313.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2259/22590313.
pdf.

denBoer:2001:NTT

[2606] Bert den Boer, Bart Van Rompay, Bart Preneel, and Joos Vandewalle. New (Two-Track-)MAC based on the two trails of RIPEMD. Lecture Notes in Computer Science, 2259:314–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2259/22590314.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2259/22590314.
                                        pdf.

Blomer:2001:KRI

[2607] Johannes Blömer and Alexander May. Key revocation with interval cover families. Lecture Notes in Computer Science, 2259:325–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2259/22590325.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2259/22590325.
                                        pdf.

Mao:2001:TRC

[2608] Wenbo Mao. Timed-release cryptography. Lecture Notes in Computer Science, 2259:342–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2259/22590342.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2259/22590342.pdf.

Anonymous:2001:AIbo

[2609] Anonymous. Author index. Lecture Notes in Computer Science, 2259:359– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2259/2259auth.pdf.

AlJabri:2001:SDA

[2610] A. Al Jabri. A statistical decoding algorithm for general linear block codes. Lecture Notes in Computer Science, 2260:1–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2260/22600001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2260/22600001.
pdf.

Lange:2001:UEP

[2611] Christoph Lange and Andreas Ahrens.

On the undetected error probability for shortened Hamming codes on channels with memory. Lecture Notes in Computer Science, 2260:9– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2260/22600009.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2260/22600009.
pdf.

Siap:2001:CWE

[2612] Irfan Siap. The complete weight enumerator for codes over M_{\×∫} (F_{⨿}). Lecture Notes in Computer Science, 2260:

20–26, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2260/22600020.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2260/22600020.
pdf.

Yoshida:2001:FIK

[2613] Maki Yoshida and Toru Fujiwara. Further improvement of Kumar-Rajagopalan-Sahai coding constructions for blacklisting problem. Lecture Notes in Computer Science, 2260:27– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2260/22600027.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2260/22600027.
pdf.

Hirst:2001:SSI

[2614] Simon Hirst and Bahram Honary. A simple soft-input/soft-output decoder for Hamming codes. Lecture Notes in Computer Science, 2260:38– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2260/22600038.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2260/22600038.
pdf.

vonWillich:2001:TIT

[2615] Manfred von Willich. A technique with an information-theoretic basis for protecting secret data from differential power attacks. Lecture Notes in Computer Science, 2260:44– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2260/22600044.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2260/22600044.
pdf.

Brincat:2001:KRA

[2616] Karl Brincat and Chris J. Mitchell. Key recovery attacks on MACs based on properties of cryptographic APIs. Lecture Notes in Computer Science, 2260: 63–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2260/22600063.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2260/22600063.
pdf.

Smart:2001:ESE

[2617] N. P. Smart. The exact security of ECIES in the generic group model. Lecture Notes in Computer Science, 2260: 73–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2260/22600073.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2260/22600073.
pdf.

Filiol:2001:NUS

[2618] Eric Filiol and Caroline Fontaine.

A new ultrafast stream cipher design: COS ciphers. Lecture Notes in Computer Science, 2260:85–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2260/22600085.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2260/22600085.
                                        pdf.

Joye:2001:RTS

[2619] Marc Joye and Jean-Jacques Quisquater. On Rabin-type signatures. Lecture Notes in Computer Science, 2260:99– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2260/22600099.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2260/22600099.
                                        pdf.

Kim:2001:SAC

[2620] Seungjoo Kim, Jung Hee Cheon, Marc Joye, Seongan Lim, Masahiro Mambo, Dongho Won, and Yuliang Zheng. Strong adaptive chosen-ciphertext attacks with memory dump (or: The importance of the order of decryption and validation). Lecture Notes in Computer Science, 2260:114–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2260/22600114.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2260/22600114.
                                        pdf.

Galan-Simon:2001:MLD

[2621] F. Javier Galán-Simón, Edgar Martínez-Moro, and Juan G. Tena-Ayuso. Majority-logic-decodable cyclic arithmeticmodular AN-codes in 1, 2, and L steps. Lecture Notes in Computer Science, 2260:128–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2260/22600128.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2260/22600128.
pdf.

MacKay:2001:ACR

[2622] David J. C. MacKay. Almost-certainly runlength-limiting codes. Lecture Notes in Computer Science, 2260:138–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2260/22600138.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2260/22600138.
pdf.

vanMourik:2001:WVM

[2623] Jort van Mourik, David Saad, and Yoshiyuki Kabashima. Weight vs. magnetization enumerator for gallager codes. Lecture Notes in Computer Science, 2260:148–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2260/22600148.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2260/22600148.
pdf.

Paire:2001:GCD

[2624] J. T. Paire, P. Coulton, and P. G.

Farrell. Graph configurations and decoding performance. Lecture Notes in Computer Science, 2260:158–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2260/22600158.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2260/22600158.
                                        pdf.

Benachour:2001:LCC

[2625] P. Benachour, P. G. Farrell, and Bahram Honary. A line code construction for the adder channel with rates higher than time-sharing. Lecture Notes in Computer Science, 2260: 166–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2260/22600166.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2260/22600166.
                                        pdf.

Al-Dabbagh:2001:STS

[2626] Ahmed Al-Dabbagh and Michael Darnell. The synthesis of TD-sequences and their application to multifunctional communication systems. Lecture Notes in Computer Science, 2260:176–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2260/22600176.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2260/22600176.
                                        pdf.

Nikova:2001:IDB

[2627] Svetla Nikova and Ventzislav Nikov. Improvement of the Delsarte bound for τ-designs in finite polynomial metric spaces. Lecture Notes in Computer Science, 2260:191–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2260/22600191.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2260/22600191.pdf.

Li:2001:SPD

[2628] Shujun Li, Qi Li, Wenmin Li, Xuanqin Mou, and Yuanlong Cai. Statistical properties of digital piecewise linear chaotic maps and their roles in cryptography and pseudo-random coding. Lecture Notes in Computer Science, 2260: 205–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2260/22600205.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2260/22600205.
pdf.

Daemen:2001:WTD

[2629] Joan Daemen and Vincent Rijmen.

The wide trail design strategy. Lecture Notes in Computer Science, 2260: 222–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2260/22600222.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2260/22600222.
pdf.

Borselius:2001:UTS

[2630] Niklas Borselius, Chris J. Mitchell, and Aaron Wilson. Undetachable threshold signatures. Lecture Notes in Computer Science, 2260:239–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2260/22600239.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2260/22600239.
pdf.

Schindler:2001:IDC

[2631] Werner Schindler, François Koeune, and Jean-Jacques Quisquater. Improving divide and conquer attacks against cryptosystems by better error detection/correction strategies. Lecture Notes in Computer Science, 2260:245– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2260/22600245.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2260/22600245.
pdf.

Rantos:2001:KRS

[2632] Konstantinos Rantos and Chris J. Mitchell. Key recovery scheme interoperability — A protocol for mechanism negotiation. Lecture Notes in Computer Science, 2260:268–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2260/22600268.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2260/22600268.
pdf.

Prissette:2001:USK

[2633] Cyril Prissette. Unconditionally secure key agreement protocol. Lecture Notes in Computer Science, 2260:277– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2260/22600277.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2260/22600277.
pdf.

Komninos:2001:ESC

[2634] N. Komninos, Bahram Honary, and Michael Darnell. An efficient stream cipher Alpha1 for mobile and wireless devices. Lecture Notes in Computer Science, 2260:294–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2260/22600294.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2260/22600294.
pdf.

Masol:2001:ILC

[2635] Viktoriya Masol. Investigation of linear codes possessing some extra properties. Lecture Notes in Computer Science, 2260:301–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2260/22600301.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2260/22600301.
pdf.

Saad:2001:SPL

[2636] David Saad, Yoshiyuki Kabashima, Tatsuto Murayama, and Renato Vicente. Statistical physics of low density parity check error correcting codes. Lecture Notes in Computer Science, 2260:307–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2260/22600307.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2260/22600307.
pdf.

Giuliani:2001:GLI

[2637] Kenneth J. Giuliani and Guang Gong. Generating large instances of the Gong-Harn cryptosystem. Lecture Notes in Computer Science, 2260:317–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2260/22600317.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2260/22600317.
pdf.

Crouch:2001:LAR

[2638] P. A. Crouch and J. H. Davenport. Lattice attacks on RSAencrypted IP and TCP. Lecture Notes in Computer Science, 2260:329– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2260/22600329.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2260/22600329.
pdf.

Parker:2001:SBS

[2639] Matthew G. Parker. Spectrally bounded sequences, codes, and states: Graph constructions and entanglement. Lecture Notes in Computer Science, 2260:339–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2260/22600339.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2260/22600339.pdf.

Geiselmann:2001:AAP

[2640] Willi Geiselmann, Rainer Steinwandt, and Thomas Beth. Attacking the affine parts of SFLASH. Lecture Notes in Computer Science, 2260:355– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2260/22600355.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2260/22600355.
pdf.

Cocks:2001:IBE

[2641] Clifford Cocks. An identity based encryption scheme based on quadratic residues. Lecture Notes in Computer Science, 2260:360–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2260/22600360.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2260/22600360.pdf.

Batina:2001:AWD

[2642] Lejla Batina and Geeke Muurling. Another way of doing RSA cryptography in hardware. Lecture Notes in Computer Science, 2260:364–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2260/22600364.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2260/22600364.
pdf.

Hernandez:2001:DTR

[2643] Julio César Hernández, José María Sierra, Arturo Ribagorda, Benjamín Ramos, and J. C. Mex-Perera. Distinguishing TEA from a random permutation: Reduced round versions of TEA do not have the SAC or do not generate random numbers. Lecture Notes in Computer Science, 2260:374– 377, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2260/22600374.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2260/22600374.
pdf.

Tabatabaian:2001:NSP

[2644] Seyed J. Tabatabaian, Sam Ikeshiro, Murat Gumussoy, and Mungal S. Dhanda. A new search pattern in multiple residue method (MRM) and its importance in the cryptanalysis of the RSA. Lecture Notes in Computer Science, 2260:378–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2260/22600378.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2260/22600378.
pdf.

Jongkook:2001:NUS

[2645] Lee Jongkook, Ryu Shiryong, Kim Jeungseop, and Yoo Keeyoung. A new undeniable signature scheme using smart cards. Lecture Notes in Computer Science, 2260:387–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2260/22600387.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2260/22600387.
pdf.

Lev:2001:NBB

[2646] Alexandr Y. Lev, Yuliy A. Lev, and Vyacheslav N. Okhrymenko. Non-binary block inseparable errors control codes. Lecture Notes in Computer Science, 2260:395–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2260/22600395.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2260/22600395.
pdf.

Leveiller:2001:CNF

[2647] Sabine Leveiller, Joseph Boutros, Philippe Guillot, and Gilles Zémor. Cryptanalysis of nonlinear filter generators with {0, 1}-metric Viterbi decoding. Lecture Notes in Computer Science, 2260:402–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2260/22600402.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2260/22600402.pdf.

Anonymous:2001:AIbp

[2648] Anonymous. Author index. Lecture Notes in Computer Science, 2260:415– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2260/2260auth.pdf.

Muller:2001:I

[2649] Peter Müller. 1. introduction. Lecture Notes in Computer Science, 2262: 1–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2262/22620001.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2262/22620001.
                                        pdf.

Muller:2001:MUT

[2650] Peter Müller. 2. mojave and the universe type system. Lecture Notes in Computer Science, 2262:39–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2262/22620039.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2262/22620039.
                                        pdf.

Muller:2001:SM

[2651] Peter Müller. 3. the semantics of mojave. Lecture Notes in Computer Science, 2262:77–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2262/22620077.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2262/22620077.
pdf.

Muller:2001:MSVa

[2652] Peter Müller. 4. modular specification and verification of functional behavior. Lecture Notes in Computer Science, 2262:123–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2262/22620123.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2262/22620123.
pdf.

Muller:2001:MSVb

[2653] Peter Müller. 5. modular specification and verification of frame properties. Lecture Notes in Computer Science, 2262:143–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2262/22620143.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2262/22620143.
pdf.

Muller:2001:MSVc

[2654] Peter Müller. 6. modular specification and verification of type invariants. Lecture Notes in Computer Science, 2262: 195–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2262/22620195.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2262/22620195.
pdf.

Muller:2001:C

[2655] Peter Müller. 7. conclusion. Lecture Notes in Computer Science, 2262:213– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2262/22620213.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2262/22620213.
pdf.

Muller:2001:FBN

[2656] Peter Müller. A. formal background and notations. Lecture Notes in Computer Science, 2262:223–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2262/22620223.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2262/22620223.
pdf.

Muller:2001:BPT

[2657] Peter Müller. B. predefined type declarations. Lecture Notes in Computer Science, 2262:227–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2262/22620227.

htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2262/22620227.pdf.

Muller:2001:CE

[2658] Peter Müller. C. examples. Lecture Notes in Computer Science, 2262:229– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2262/22620229.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2262/22620229.
pdf.

Muller:2001:DAL

[2659] Peter Müller. D. auxiliary lemmas, proofs, and models. Lecture Notes in Computer Science, 2262:237–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2262/22620237.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2262/22620237.
pdf.

Anonymous:2001:B

[2660] Anonymous. Bibliography. Lecture Notes in Computer Science, 2262:271– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2262/22620271.pdf.

Anonymous:2001:LF

[2661] Anonymous. List of figures. Lecture Notes in Computer Science, 2262:285– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2262/22620285.pdf.

Anonymous:2001:Ir

[2662] Anonymous. Index. Lecture Notes in Computer Science, 2262:287–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2262/22620287.pdf.

Selic:2001:I

[2663] Bran Selic. Introduction. Lecture Notes in Computer Science, 2263:1– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2263/22630001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2263/22630001.
pdf.

Clark:2001:MMF

[2664] Tony Clark, Andy Evans, and Stuart Kent. A meta-model facility for a family of UML constraint languages. Lecture Notes in Computer Science, 2263: 4–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2263/22630004.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2263/22630004.
pdf.

Schurr:2001:NTC

[2665] Andy Schürr. A new type checking approach for OCL version 2.0 ? Lecture Notes in Computer Science, 2263:21– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2263/22630021.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2263/22630021.
pdf.

Richters:2001:OSS

[2666] Mark Richters and Martin Gogolla.

OCL: Syntax, semantics, and tools. Lecture Notes in Computer Science, 2263:42–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2263/22630042.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2263/22630042.
pdf.

Hennicker:2001:PMO

[2667] Rolf Hennicker, Heinrich Hussmann, and Michel Bidoit. On the precise meaning of OCL constraints. Lecture Notes in Computer Science, 2263:69– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2263/22630069.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2263/22630069.
pdf.

Gogolla:2001:EUC

[2668] Martin Gogolla and Mark Richters.

Expressing UML class diagrams properties with OCL. Lecture Notes in Computer Science, 2263:85–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2263/22630085.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2263/22630085.
                                        pdf.

Cook:2001:AMO

[2669] Steve Cook, Anneke Kleppe, Richard Mitchell, Bernhard Rumpe, Jos Warmer, and Alan Wills. The Amsterdam Manifesto on OCL. Lecture Notes in Computer Science, 2263:115– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2263/22630115.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2263/22630115.
                                        pdf.

Flake:2001:OER

[2670] Stephan Flake and Wolfgang Mueller. An OCL extension for real-time constraints. Lecture Notes in Computer Science, 2263:150–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2263/22630150.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2263/22630150.pdf.

Knapman:2001:SCV

[2671] John Knapman. Statistical constraints and verification. Lecture Notes in Computer Science, 2263:172– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2263/22630172.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2263/22630172.
pdf.

Rumpe:2001:BNP

[2672] Bernhard Rumpe. ≪Java≫OCL based on new presentation of the OCLsyntax. Lecture Notes in Computer Science, 2263:189–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2263/22630189.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2263/22630189.pdf.

Kleppe:2001:SOA

[2673] Anneke Kleppe and Jos Warmer. The semantics of the OCL action clause. Lecture Notes in Computer Science, 2263:213–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2263/22630213.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2263/22630213.
pdf.

Kent:2001:CT

[2674] Stuart Kent and John Howse. Constraint trees. Lecture Notes in Computer Science, 2263:228–??, 2001. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.com/link/
service/series/0558/bibs/2263/22630228.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2263/22630228.pdf.

Sendall:2001:UOU

[2675] Shane Sendall and Alfred Strohmeier. Using OCL and UML to specify system behavior. Lecture Notes in Computer Science, 2263:250–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2263/22630250.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2263/22630250.
                                        pdf.

Anonymous:2001:AIbq

[2676] Anonymous. Author index. Lecture Notes in Computer Science, 2263:281– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2263/2263auth.pdf.

Hromkovic:2001:RCP

[2677] Juraj Hromkovic. Randomized communication protocols (A survey). Lecture Notes in Computer Science, 2264: 1–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2264/22640001.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2264/22640001.
                                        pdf.

Mahnig:2001:OMR

[2678] Thilo Mahnig and Heinz Mühlenbein. Optimal mutation rate using Bayesian priors for estimation of distribution algorithms. Lecture Notes in Computer Science, 2264:33–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2264/22640033.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2264/22640033.
pdf.

Fitzpatrick:2001:EAS

[2679] Stephen Fitzpatrick and Lambert Meertens. An experimental assessment of a stochastic, anytime, decentralized, soft colourer for sparse graphs. Lecture Notes in Computer Science, 2264: 49–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2264/22640049.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2264/22640049.
pdf.

Sauerhoff:2001:RBP

[2680] Martin Sauerhoff. Randomized branching programs. Lecture Notes in Computer Science, 2264:65–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2264/22640065.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2264/22640065.
pdf.

Codognet:2001:YAL

[2681] Philippe Codognet and Daniel Diaz.

Yet another local search method for constraint solving. Lecture Notes in Computer Science, 2264:73–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2264/22640073.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2264/22640073.
                                        pdf.

Meloni:2001:EAS

[2682] Carlo Meloni. An evolutionary algorithm for the sequence coordination in furniture production. Lecture Notes in Computer Science, 2264:91– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2264/22640091.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2264/22640091.
                                        pdf.

Poland:2001:ESS

[2683] Jan Poland, Kosmas Knödler, Alexander Mitterer, Thomas Fleischhauer, Frank Zuber-Goos, and Andreas Zell. Evolutionary search for smooth maps in motor control unit calibration. Lecture Notes in Computer Science, 2264: 107–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2264/22640107.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2264/22640107.
                                        pdf.

Sorkin:2001:SNR

[2684] Gregory B. Sorkin. Some notes on random satisfiability. Lecture Notes in Computer Science, 2264:117–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2264/22640117.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2264/22640117.
pdf.

Naumann:2001:PSA

[2685] Uwe Naumann and Peter Gottschling.

Prospects for simulated annealing algorithms in automatic differentiation. Lecture Notes in Computer Science, 2264:131–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2264/22640131.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2264/22640131.
pdf.

Behnke:2001:OSS

[2686] Henning Behnke, Michael Kolonko, Ulrich Mertins, and Stefan Schnitter. Optimization and simulation: Sequential packing of flexible objects using evolutionary algorithms. Lecture Notes in Computer Science, 2264:145– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2264/22640145.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2264/22640145.
pdf.

Zeugmann:2001:SFL

[2687] Thomas Zeugmann. Stochastic finite learning. Lecture Notes in Computer Science, 2264:155–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2264/22640155.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2264/22640155.
                                        pdf.

Gavalda:2001:SSA

[2688] Ricard Gavaldà and Osamu Watanabe. Sequential sampling algorithms: Unified analysis and lower bounds. Lecture Notes in Computer Science, 2264: 173–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2264/22640173.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2264/22640173.
                                        pdf.

Damaschke:2001:ALR

[2689] Peter Damaschke. Approximate location of relevant variables under the crossover distribution. Lecture Notes in Computer Science, 2264:189–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2264/22640189.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2264/22640189.
                                        pdf.

Anonymous:2001:AIbr

[2690] Anonymous. Author index. Lecture Notes in Computer Science, 2264:203– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
papers/2264/2264auth.pdf.

Bardohl:2001:IRB

[2691] R. Bardohl, M. Große-Rhode, and M. Simeoni. Interactive rule-based specification with an application to visual language definition. Lecture Notes in Computer Science, 2267:1– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2267/22670001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2267/22670001.
pdf.

Bidoit:2001:AA

[2692] Michel Bidoit and Alexandre Boisseau. Algebraic abstractions. Lecture Notes in Computer Science, 2267:21– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2267/22670021.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2267/22670021.
pdf.

Caleiro:2001:CLP

[2693] C. Caleiro, P. Mateus, J. Ramos, and A. Sernadas. Combining logics: Parchments revisited. Lecture Notes in Computer Science, 2267:48– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2267/22670048.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2267/22670048.
pdf.

Costa:2001:CIB

[2694] J. Félix Costa and H. Lourenço. Canonical institutions of behaviour. Lecture Notes in Computer Science, 2267:71–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2267/22670071.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2267/22670071.
pdf.

Dosch:2001:LTS

[2695] Walter Dosch and Sönke Magnussen. The Lübeck transformation system: a transformation system for equational higher order algebraic specifications. Lecture Notes in Computer Science, 2267:85–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2267/22670085.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2267/22670085.
pdf.

Gajewsky:2001:CMM

[2696] Maike Gajewsky and Francesco Parisi-Presicce. On the compatibility of model and model-class transformations. Lecture Notes in Computer Science, 2267: 109–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2267/22670109.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2267/22670109.
pdf.

Harman:2001:VSP

[2697] N. A. Harman. Verifying a simple pipelined microprocessor using maude. Lecture Notes in Computer Science, 2267:128–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2267/22670128.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2267/22670128.
pdf.

Hoffman:2001:VAS

[2698] Piotr Hoffman. Verifying architectural specifications. Lecture Notes in Computer Science, 2267:152–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2267/22670152.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2267/22670152.
pdf.

Ledoux:2001:GMC

[2699] Franck Ledoux, Agnès Arnould, Pascale Le Gall, and Yves Bertrand. Geometric modelling with CASL. Lecture Notes in Computer Science, 2267:176– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2267/22670176.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2267/22670176.
pdf.

Lopes:2001:CAC

[2700] Antónia Lopes, Michel Wermelinger, and José Luiz Fiadeiro. A compositional approach to connector construction. Lecture Notes in Computer Science, 2267:201–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2267/22670201.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2267/22670201.pdf.

Mossakowski:2001:IIS

[2701] Till Mossakowski and Bartek Klin. Institution independent static analysis for C ASL. Lecture Notes in Computer Science, 2267:221–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2267/22670221.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2267/22670221.
                                        pdf.

Orejas:2001:TLS

[2702] Fernando Orejas, Hartmut Ehrig, and Elvira Pino. Tight and loose semantics for transformation systems. Lecture Notes in Computer Science, 2267: 238–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2267/22670238.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2267/22670238.
                                        pdf.

Pawlowski:2001:PAC

[2703] Wiesław Pawłowski. Presentations for abstract context institutions. Lecture Notes in Computer Science, 2267:256– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2267/22670256.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2267/22670256.
pdf.

Poernomo:2001:PPP

[2704] Iman Poernomo, John N. Crossley, and Martin Wirsing. Programs, proofs and parametrized specifications. Lecture Notes in Computer Science, 2267:280– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2267/22670280.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2267/22670280.
pdf.

Roggenbach:2001:TTS

[2705] Markus Roggenbach and Lutz Schröder.

Towards trustworthy specifications I: Consistency checks. Lecture Notes in Computer Science, 2267:305–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2267/22670305.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2267/22670305.
pdf.

Sannella:2001:CFI

[2706] Donald Sannella. The common framework initiative for algebraic specification and development of software: Recent progress. Lecture Notes in Computer Science, 2267:328–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2267/22670328.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2267/22670328.
                                        pdf.

Anonymous:2001:AIbs

[2707] Anonymous. Author index. Lecture Notes in Computer Science, 2267:345– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2267/2267auth.pdf.

Kaliski:2001:HFF

[2708] Burton S. Kaliski Jr. On hash function firewalls in signature schemes. Lecture Notes in Computer Science, 2271: 1–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2271/22710001.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2271/22710001.
                                        pdf.

Joye:2001:OAD

[2709] Marc Joye, Jean-Jacques Quisquater, Sung-Ming Yen, and Moti Yung. Observability analysis — detecting when improved cryptosystems fail. Lecture Notes in Computer Science, 2271:17– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2271/22710017.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2271/22710017.
pdf.

Walter:2001:PBM

[2710] Colin D. Walter. Precise bounds for montgomery modular multiplication and some potentially insecure RSA moduli. Lecture Notes in Computer Science, 2271:30–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2271/22710030.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2271/22710030.
pdf.

Batina:2001:MPH

[2711] Lejla Batina and Geeke Muurling.

Montgomery in practice: How to do it more efficiently in hardware. Lecture Notes in Computer Science, 2271: 40–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2271/22710040.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2271/22710040.
pdf.

Walter:2001:MER

[2712] Colin D. Walter. M IST : An efficient, randomized exponentiation algorithm for resisting power analysis. Lecture Notes in Computer Science, 2271:53– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2271/22710053.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2271/22710053.
                                        pdf.

Wolkerstorfer:2001:AIA

[2713] Johannes Wolkerstorfer, Elisabeth Oswald, and Mario Lamberger. An ASIC implementation of the AES SBoxes. Lecture Notes in Computer Science, 2271:67–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2271/22710067.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2271/22710067.
                                        pdf.

Fischlin:2001:ICN

[2714] Marc Fischlin. On the impossibility of constructing non-interactive statistically-secret protocols from any trapdoor one-way function. Lecture Notes in Computer Science, 2271:79– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2271/22710079.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2271/22710079.
                                        pdf.

Fischlin:2001:RPB

[2715] Marc Fischlin and Roger Fischlin. The representation problem based on factoring. Lecture Notes in Computer Science, 2271:96–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2271/22710096.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2271/22710096.pdf.

Black:2001:CAF

[2716] John Black and Phillip Rogaway. Ciphers with arbitrary finite domains. Lecture Notes in Computer Science, 2271:114–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2271/22710114.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2271/22710114.
pdf.

Miyaji:2001:KPC

[2717] Atsuko Miyaji, Masao Nonaka, and Yoshinori Takii. Known plaintext correlation attack against RC5. Lecture Notes in Computer Science, 2271:131– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2271/22710131.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2271/22710131.
pdf.

Micali:2001:MR

[2718] Silvio Micali and Ronald L. Rivest.

Micropayments revisited. Lecture Notes in Computer Science, 2271:149– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2271/22710149.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2271/22710149.
                                        pdf.

Jakobsson:2001:PC

[2719] Markus Jakobsson, Ari Juels, and Phong Q. Nguyen. Proprietary certificates. Lecture Notes in Computer Science, 2271:164–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2271/22710164.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2271/22710164.pdf.

Ateniese:2001:SRC

[2720] Giuseppe Ateniese and Cristina Nita-Rotaru. Stateless-recipient certified E-mail system based on verifiable encryption. Lecture Notes in Computer Science, 2271:182–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2271/22710182.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2271/22710182.
                                        pdf.

Galbraith:2001:RBU

[2721] Steven D. Galbraith, Wenbo Mao, and Kenneth G. Paterson. RSA-based undeniable signatures for general moduli. Lecture Notes in Computer Science, 2271:200–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/

0558/bibs/2271/22710200.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2271/22710200.
pdf.

Maitland:2001:COF

[2722] Greg Maitland and Colin Boyd. Cooperatively formed group signatures. Lecture Notes in Computer Science, 2271:218–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2271/22710218.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2271/22710218.
pdf.

Micali:2001:TSS

[2723] Silvio Micali and Ronald L. Rivest.

Transitive signature schemes. Lecture Notes in Computer Science, 2271:236– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2271/22710236.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2271/22710236.
pdf.

Johnson:2001:HSS

[2724] Robert Johnson, David Molnar, Dawn Song, and David Wagner. Homomorphic signature schemes. Lecture Notes in Computer Science, 2271:244– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2271/22710244.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2271/22710244.
pdf.

Coron:2001:GGE

[2725] Jean-Sébastien Coron, Helena Handschuh, Marc Joye, Pascal Paillier, David Pointcheval, and Christophe Tymen. GEM: a Generic chosenciphertext secure Encryption Method. Lecture Notes in Computer Science, 2271:263–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2271/22710263.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2271/22710263.
pdf.

Abe:2001:SSP

[2726] Masayuki Abe. Securing “Encryption + Proof of Knowledge” in the random oracle model. Lecture Notes in Computer Science, 2271:277–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2271/22710277.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2271/22710277.
pdf.

Cheng:2001:NPT

[2727] Qi Cheng and Shigenori Uchiyama. Nonuniform polynomial time algorithm to solve decisional Diffie– Hellman problem in finite fields under conjecture. Lecture Notes in Computer Science, 2271:290–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2271/22710290.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2271/22710290.
pdf.

Lu:2001:SKE

[2728] Cheng-Fen Lu and ShiuhPyng Winston Shieh. Secure key-evolving protocols for discrete logarithm schemes. Lecture Notes in Computer Science, 2271: 300–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2271/22710300.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2271/22710300.
pdf.

Anonymous:2001:AIbt

[2729] Anonymous. Author index. Lecture Notes in Computer Science, 2271:311– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2271/2271auth.pdf.

Allan:2001:PIR

[2730] James Allan. Perspectives on information retrieval and speech. Lecture Notes in Computer Science, 2273:1– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2273/22730001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2273/22730001.
pdf.

Brown:2001:CRT

[2731] Eric W. Brown and Anni R. Coden. Capitalization recovery for text. Lecture Notes in Computer Science, 2273: 11–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2273/22730011.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2273/22730011.
pdf.

Ibrahimov:2001:CIT

[2732] Oktay Ibrahimov, Ishwar Sethi, and Nevenka Dimitrova. Clustering of imperfect transcripts using a novel similarity measure. Lecture Notes in Computer Science, 2273:23–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2273/22730023.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2273/22730023.
pdf.

Desilets:2001:EKS

[2733] Alain Désilets, Berry de Bruijn, and Joel Martin. Extracting keyphrases from spoken audio documents. Lecture Notes in Computer Science, 2273:36– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2273/22730036.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2273/22730036.
pdf.

Ries:2001:SCT

[2734] Klaus Ries. Segmenting conversations by topic, initiative, and style. Lecture Notes in Computer Science, 2273:51– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2273/22730051.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2273/22730051.
pdf.

Huang:2001:ECI

[2735] Jing Huang, Geoffrey Zweig, and Mukund Padmanabhan. Extracting caller information from voicemail. Lecture Notes in Computer Science, 2273: 67–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2273/22730067.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2273/22730067.
pdf.

Sanderson:2001:SHT

[2736] Mark Sanderson and Xiao Mang Shou. Speech and hand transcribed retrieval. Lecture Notes in Computer Science, 2273:78–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2273/22730078.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2273/22730078.
pdf.

Kim:2001:USR

[2737] Jinmook Kim and Douglas W. Oard. The use of speech retrieval systems: a study design. Lecture Notes in Computer Science, 2273:86– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2273/22730086.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2273/22730086.
pdf.

Fujii:2001:SDT

[2738] Atsushi Fujii, Katunobu Itou, and Tetsuya Ishikawa. Speech-driven text retrieval: Using target IR collections for statistical language model adaptation in speech recognition. Lecture Notes in Computer Science, 2273:94– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2273/22730094.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2273/22730094.
pdf.

Brown:2001:WFR

[2739] Eric W. Brown and Anni R. Coden. WASABI: Framework for real-time speech analysis applications (demo). Lecture Notes in Computer Science, 2273:105–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2273/22730105.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2273/22730105.
pdf.

Anonymous:2001:AIbu

[2740] Anonymous. Author index. Lecture Notes in Computer Science, 2273:109– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2273/2273auth.pdf.

Sakurai:2001:NSS

[2741] Kouichi Sakurai and Tsuyoshi Takagi. New semantically secure public-key cryptosystems from the RSA-primitive. Lecture Notes in Computer Science, 2274:1–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2274/22740001.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2274/22740001.
pdf.

Coron:2001:OCC

[2742] Jean-Sébastien Coron, Helena Handschuh, Marc Joye, Pascal Paillier, David Pointcheval, and Christophe Tymen. Optimal chosen-ciphertext secure encryption of arbitrary-length messages. Lecture Notes in Computer Science, 2274:17–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2274/22740017.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2274/22740017.pdf.

Koshiba:2001:SRS

[2743] Takeshi Koshiba. On sufficient randomness for secure public-key cryptosystems. Lecture Notes in Computer Science, 2274:34–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2274/22740034.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2274/22740034.pdf.

Kurosawa:2001:MRP

[2744] Kaoru Kurosawa. Multi-recipient public-key encryption with shortened ciphertext. Lecture Notes in Computer Science, 2274:48–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2274/22740048.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2274/22740048.
pdf.

Hanaoka:2001:EUS

[2745] Goichiro Hanaoka, Junji Shikata, Yuliang Zheng, and Hideki Imai. Efficient and unconditionally secure digital signatures and a security analysis of a multireceiver authentication code. Lecture Notes in Computer Science, 2274: 64–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2274/22740064.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2274/22740064.
pdf.

Baek:2001:FPS

[2746] Joonsang Baek, Ron Steinfeld, and Yuliang Zheng. Formal proofs for the security of signcryption. Lecture Notes in Computer Science, 2274:80– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2274/22740080.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2274/22740080.
pdf.

Maitland:2001:PSR

[2747] Greg Maitland and Colin Boyd. A provably secure restrictive partially blind signature scheme. Lecture Notes in Computer Science, 2274:99– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2274/22740099.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2274/22740099.
pdf.

Abe:2001:SPA

[2748] Masayuki Abe and Koutarou Suzuki.

M +1-st price auction using homomorphic encryption. Lecture Notes in Computer Science, 2274:115–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2274/22740115.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2274/22740115.
pdf.

Damgaard:2001:CST

[2749] Ivan Damgård and Mads Jurik. Client/ server tradeoffs for online elections. Lecture Notes in Computer Science, 2274:125–??, 2001. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2274/22740125.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2274/22740125.
                                        pdf.

Kiayias:2001:STE

[2750] Aggelos Kiayias and Moti Yung. Selftallying elections and perfect ballot secrecy. Lecture Notes in Computer Science, 2274:141–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2274/22740141.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2274/22740141.pdf.

Tzeng:2001:EOT

[2751] Wen-Guey Tzeng. Efficient 1-out-n oblivious transfer schemes. Lecture Notes in Computer Science, 2274:159– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2274/22740159.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2274/22740159.
                                        pdf.

Kurosawa:2001:LCI

[2752] Kaoru Kurosawa and Takuya Yoshida. Linear code implies public-key traitor tracing. Lecture Notes in Computer Science, 2274:172–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2274/22740172.

htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2274/22740172.pdf.

Lee:2001:DSA

[2753] Chan H. Lee, Xiaotie Deng, and Huafei Zhu. Design and security analysis of anonymous group identification protocols. Lecture Notes in Computer Science, 2274:188–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2274/22740188.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2274/22740188.
pdf.

Quisquater:2001:STS

[2754] Michaël Quisquater, Bart Preneel, and Joos Vandewalle. On the security of the threshold scheme based on the Chinese Remainder Theorem. Lecture Notes in Computer Science, 2274:199– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2274/22740199.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2274/22740199.
pdf.

Courtois:2001:SUS

[2755] Nicolas Courtois, Louis Goubin, Willi Meier, and Jean-Daniel Tacier. Solving underdefined systems of multivariate quadratic equations. Lecture Notes in Computer Science, 2274:211– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2274/22740211.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2274/22740211.
                                        pdf.

Lenstra:2001:SFR

[2756] Arjen K. Lenstra and Igor E. Shparlinski. Selective forgery of RSA signatures with fixed-pattern padding. Lecture Notes in Computer Science, 2274: 228–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2274/22740228.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2274/22740228.
                                        pdf.

Kobara:2001:NCP

[2757] Kazukuni Kobara and Hideki Imai. New chosen-plaintext attacks on the one-wayness of the modified McEliece PKC proposed at Asiacrypt 2000. Lecture Notes in Computer Science, 2274: 237–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2274/22740237.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2274/22740237.
                                        pdf.

Novak:2001:SBA

[2758] Roman Novak. SPA-based adaptive chosen-ciphertext attack on RSA implementation. Lecture Notes in Computer Science, 2274:252–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2274/22740252.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2274/22740252.
pdf.

Schindler:2001:CTP

[2759] Werner Schindler. A combined timing and power attack. Lecture Notes in Computer Science, 2274:263–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2274/22740263.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2274/22740263.
pdf.

Izu:2001:FPE

[2760] Tetsuya Izu and Tsuyoshi Takagi. A fast parallel elliptic curve multiplication resistant against side channel attacks. Lecture Notes in Computer Science, 2274:280–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2274/22740280.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2274/22740280.
pdf.

Preneel:2001:NES

[2761] Bart Preneel. New European Schemes for Signature, Integrity and Encryption (NESSIE): a status report. Lecture Notes in Computer Science, 2274: 297–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2274/22740297.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2274/22740297.
                                        pdf.

Park:2001:IMM

[2762] Young-Ho Park, Sangho Oh, Sangjin Lee, Jongin Lim, and Maenghee Sung. An improved method of multiplication on certain elliptic curves. Lecture Notes in Computer Science, 2274:310– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2274/22740310.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2274/22740310.
                                        pdf.

Park:2001:ADI

[2763] Young-Ho Park, Sangtae Jeong, Chang Han Kim, and Jongin Lim. An alternate decomposition of an integer for faster point multiplication on certain elliptic curves. Lecture Notes in Computer Science, 2274:323– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2274/22740323.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2274/22740323.
                                        pdf.

Brier:2001:WEC

[2764] Éric Brier and Marc Joye. Weierstraß elliptic curves and side-channel attacks. Lecture Notes in Computer Science, 2274:335–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2274/22740335.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2274/22740335.
pdf.

Joye:2001:OWC

[2765] Marc Joye and Sung-Ming Yen. Oneway cross-trees and their applications. Lecture Notes in Computer Science, 2274:346–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2274/22740346.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2274/22740346.
pdf.

Juels:2001:RKG

[2766] Ari Juels and Jorge Guajardo.

RSA key generation with verifiable randomness. Lecture Notes in Computer Science, 2274:357–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2274/22740357.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2274/22740357.
pdf.

Joye:2001:NMM

[2767] Marc Joye and Sung-Ming Yen. New minimal modified radix- r representation with applications to smart cards. Lecture Notes in Computer Science, 2274:375–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2274/22740375.htm;   http:

                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2274/22740375.
                                        pdf.

Anonymous:2001:AIbv

[2768] Anonymous. Author index. Lecture Notes in Computer Science, 2274:385– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2274/2274auth.pdf.

Dubois:2001:NPR

[2769] D. Dubois, H. Prade, and L. Ughetto. A new perspective on reasoning with fuzzy rules. Lecture Notes in Computer Science, 2275:1–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2275/22750001.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2275/22750001.
                                        pdf.

Furuhashi:2001:IFM

[2770] T. Furuhashi. On interpretability of fuzzy models. Lecture Notes in Computer Science, 2275:12–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2275/22750012.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2275/22750012.
                                        pdf.

Intan:2001:DSF

[2771] R. Intan and M. Mukaidono. Degree of similarity in fuzzy partition. Lecture Notes in Computer Science, 2275: 20–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2275/22750020.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2275/22750020.
pdf.

Raha:2001:FSD

[2772] S. Raha and S. Hossain. Fuzzy set in default reasoning. Lecture Notes in Computer Science, 2275:27– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2275/22750027.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2275/22750027.
pdf.

Koczy:2001:IHR

[2773] L. T. Kóczy and L. Muresan. Interpolation in hierarchical rule-bases with normal conclusions. Lecture Notes in Computer Science, 2275:34– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2275/22750034.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2275/22750034.
pdf.

Kim:2001:DSA

[2774] Y.-C. Kim, S.-B. Cho, and S.-R.

Oh. The Dempster-Shafer approach to map-building for an autonomous mobile robot with fuzzy controller. Lecture Notes in Computer Science, 2275:

40–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2275/22750040.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2275/22750040.
                                        pdf.

Ionita:2001:FMA

[2775] S. Ionita and E. Sofron. The fuzzy model for aircraft landing control. Lecture Notes in Computer Science, 2275: 47–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2275/22750047.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2275/22750047.
                                        pdf.

Himavathi:2001:INF

[2776] S. Himavathi and B. Umamaheswari. Implementation of nonlinear fuzzy models using microcontrollers. Lecture Notes in Computer Science, 2275:55– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2275/22750055.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2275/22750055.
                                        pdf.

Mudi:2001:GAF

[2777] R. K. Mudi, K. Majumdar, and C. Dey. A gain adaptive fuzzy logic controller. Lecture Notes in Computer Science, 2275:62–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2275/22750062.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2275/22750062.
pdf.

Choi:2001:TLC

[2778] W. K. Choi, H. S. Yoon, K. Kim, I. Y. Chung, and S. J. Lee. A traffic light controlling FLC considering the traffic congestion. Lecture Notes in Computer Science, 2275:69– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2275/22750069.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2275/22750069.
pdf.

Hirota:2001:VRS

[2779] K. Hirota, F. Dong, K. Chen, and Y. Takama. Vehicle routing, scheduling and dispatching system based on HIMS model. Lecture Notes in Computer Science, 2275:76–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2275/22750076.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2275/22750076.
pdf.

Myung:2001:SBF

[2780] H.-C. Myung and Z. Z. Bien. Similarity between fuzzy multi-objective control and eligibility. Lecture Notes in Computer Science, 2275:85–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2275/22750085.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2275/22750085.
                                        pdf.

Moitra:2001:FGP

[2781] B. N. Moitra and B. B. Pal. A fuzzy goal programming approach for solving bilevel programming problems. Lecture Notes in Computer Science, 2275:91– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2275/22750091.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2275/22750091.
                                        pdf.

Deb:2001:MHE

[2782] S. K. Deb, B. Bhattacharyya, and S. K. Sorkhel. Material handling equipment selection by fuzzy multi-criteria decision making methods. Lecture Notes in Computer Science, 2275:99– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2275/22750099.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2275/22750099.
                                        pdf.

Pawlak:2001:RSV

[2783] Z. Pawlak. The rough set view on Bayes’ theorem. Lecture Notes in Computer Science, 2275:106–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2275/22750106.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2275/22750106.
pdf.

Sugeno:2001:VCW

[2784] M. Sugeno. Verbalizing computers — A way to everyday language computing. Lecture Notes in Computer Science, 2275:117–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2275/22750117.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2275/22750117.
pdf.

Bien:2001:SCB

[2785] Z. Z. Bien, J.-B. Kim, D.-J. Kim, J.- S. Han, and J.-H. Do. Soft computing based emotion/intention reading for service robot. Lecture Notes in Computer Science, 2275:121–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2275/22750121.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2275/22750121.
pdf.

Intan:2001:GRM

[2786] R. Intan and M. Mukaidono. Generalization of rough membership function based on α-coverings of the universe. Lecture Notes in Computer Science, 2275:129–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2275/22750129.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2275/22750129.
                                        pdf.

Taneja:2001:CQQ

[2787] H. C. Taneja. A class of quantitativequalitative measures of directeddivergence. Lecture Notes in Computer Science, 2275:136–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2275/22750136.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2275/22750136.pdf.

Ghosh:2001:RDT

[2788] S. Ghosh and S. S. Alam. (α, β) reduction of decision table: a rough approach. Lecture Notes in Computer Science, 2275:141–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2275/22750141.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2275/22750141.
                                        pdf.

Tiwari:2001:WTB

[2789] A. K. Tiwari and K. K. Shukla. Wavelet transform based fuzzy inference system for power quality classification. Lecture Notes in Computer Science, 2275:148–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2275/22750148.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2275/22750148.
pdf.

De:2001:FNT

[2790] A. De and N. Chatterjee. A fuzzyneural technique for flashover diagnosis of winding insulation in transformers. Lecture Notes in Computer Science, 2275:156–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2275/22750156.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2275/22750156.
pdf.

Mukherjee:2001:FIF

[2791] A. Mukherjee, P. P. Adhikari, and P. K. Nandi. Feature identification for fuzzy logic based adaptive Kalman filtering. Lecture Notes in Computer Science, 2275:163–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2275/22750163.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2275/22750163.
pdf.

Franke:2001:SBS

[2792] K. Franke and J. Ruiz del Solar. Softbiometrics: Soft-computing technologies for biometric-applications. Lecture Notes in Computer Science, 2275:171– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2275/22750171.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2275/22750171.
                                        pdf.

Navarrete:2001:CSB

[2793] P. Navarrete and J. Ruiz del Solar. Comparative study between different eigenspace-based approaches for face recognition. Lecture Notes in Computer Science, 2275:178–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2275/22750178.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2275/22750178.
                                        pdf.

Franke:2001:SSV

[2794] K. Franke, Y.-N. Zhang, and M. Köppen. Static signature verification employing a kosko-neuro-fuzzy approach. Lecture Notes in Computer Science, 2275:185– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2275/22750185.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2275/22750185.
                                        pdf.

Roy:2001:CAN

[2795] P. K. Roy, J. P. Miller, and D. Dutta Majumder. A control analysis of neuronal information processing: a study of electrophysiological experimentation and non-equilibrium information theory. Lecture Notes in Computer Science, 2275:191–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2275/22750191.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2275/22750191.
pdf.

Yamakawa:2001:MNS

[2796] T. Yamakawa and K. Horio. Modeling of nonlinear systems by employing self-organization and evaluation — SOR network. Lecture Notes in Computer Science, 2275:204–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2275/22750204.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2275/22750204.
pdf.

Ganguly:2001:CNL

[2797] N. Ganguly, P. Maji, A. Das, B. K. Sikdar, and P. Pal Chaudhuri. Characterization of non-linear cellular automata model for pattern recognition. Lecture Notes in Computer Science, 2275:214– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2275/22750214.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2275/22750214.
pdf.

Parui:2001:CNN

[2798] S. K. Parui and A. Datta. On convergence of a neural network model computing MSC. Lecture Notes in Computer Science, 2275:221–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2275/22750221.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2275/22750221.
                                        pdf.

Bhattacharya:2001:RHB

[2799] U. Bhattacharya, T. K. Das, A. Datta, S. K. Parui, and B. B. Chaudhuri. Recognition of handprinted bangla numerals using neural network models. Lecture Notes in Computer Science, 2275:228–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2275/22750228.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2275/22750228.
                                        pdf.

Sung:2001:OSM

[2800] S.-K. Sung, J.-W. Jung, J.-T. Lee, W.-J. Choi, and S.-J. Ji. Optimal synthesis method for binary neural network using NETLA. Lecture Notes in Computer Science, 2275:236– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2275/22750236.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2275/22750236.
                                        pdf.

Chakraborty:2001:NNB

[2801] B. Chakraborty. A neural network based seafloor classification using acoustic backscatter. Lecture Notes in Computer Science, 2275:245– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2275/22750245.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2275/22750245.
pdf.

Chakraborty:2001:DRB

[2802] D. Chakraborty and N. R. Pal.

Designing rule-based classifiers with on-line feature selection: a neurofuzzy approach. Lecture Notes in Computer Science, 2275:251–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2275/22750251.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2275/22750251.
pdf.

Golob:2001:DNF

[2803] M. Golob and B. Tovornik. Decomposed neuro-fuzzy ARX model. Lecture Notes in Computer Science, 2275: 260–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2275/22750260.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2275/22750260.
pdf.

Tham:2001:WFS

[2804] C.-W. Tham, S.-H. Tian, and L. Ding.

Weather forecasting system based on satellite imageries using neurofuzzy techniques. Lecture Notes in Computer Science, 2275:267–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2275/22750267.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2275/22750267.
                                        pdf.

Velayutham:2001:ESP

[2805] C. S. Velayutham, S. Paul, and S. Kumar. Evolutionary subsethood product fuzzy neural network. Lecture Notes in Computer Science, 2275:274– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2275/22750274.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2275/22750274.
                                        pdf.

Yildiran:2001:VLB

[2806] U. Yildiran and O. Kaynak. VSS learning based intelligent control of a bioreactor system. Lecture Notes in Computer Science, 2275:281–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2275/22750281.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2275/22750281.
                                        pdf.

Bezdek:2001:SNA

[2807] J. C. Bezdek and R. J. Hathaway. Some notes on alternating optimization. Lecture Notes in Computer Science, 2275:

288–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2275/22750288.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2275/22750288.
pdf.

Lin:2001:NSS

[2808] C.-T. Lin, D.-J. Liu, R.-C. Wu, and G.- D. Wu. Noisy speech segmentation/ enhancement with multiband analysis and neural fuzzy networks. Lecture Notes in Computer Science, 2275:301– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2275/22750301.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2275/22750301.
pdf.

Ryu:2001:TOF

[2809] J. Ryu and S.-B. Cho. Towards optimal feature and classifier for gene expression classification of cancer. Lecture Notes in Computer Science, 2275: 310–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2275/22750310.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2275/22750310.
pdf.

Tran:2001:FCM

[2810] D. Tran and M. Wagner. Fuzzy Cmeans clustering-based speaker verification. Lecture Notes in Computer Science, 2275:318–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2275/22750318.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2275/22750318.pdf.

Tran:2001:NCB

[2811] D. Tran and M. Wagner. Noise clustering-based speaker verification. Lecture Notes in Computer Science, 2275:325–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2275/22750325.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2275/22750325.
                                        pdf.

Dimitriadou:2001:CSF

[2812] E. Dimitriadou, A. Weingessel, and K. Hornik. A combination scheme for fuzzy clustering. Lecture Notes in Computer Science, 2275:332–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2275/22750332.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2275/22750332.
                                        pdf.

Mali:2001:CSD

[2813] K. Mali and S. Mitra. Clustering of symbolic data and its validation. Lecture Notes in Computer Science, 2275: 339–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2275/22750339.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2275/22750339.
pdf.

Tran:2001:GFH

[2814] D. Tran and M. Wagner. Generalised fuzzy hidden Markov models for speech recognition. Lecture Notes in Computer Science, 2275:345– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2275/22750345.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2275/22750345.
pdf.

Chowdhury:2001:GKF

[2815] P. Roy Chowdhury and K. K.

Shukla. On generalization and Kfold cross validation performance of MLP trained with EBPDT. Lecture Notes in Computer Science, 2275:352– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2275/22750352.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2275/22750352.
pdf.

DiGesu:2001:PIS

[2816] V. Di Gesú and S. Roy. Pictorial indexes and soft image distances. Lecture Notes in Computer Science, 2275: 360–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2275/22750360.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2275/22750360.
                                        pdf.

Kumar:2001:SCU

[2817] S. S. Kumar and B. N. Chatterji. Stereo correspondence using a fuzzy approach. Lecture Notes in Computer Science, 2275:367–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2275/22750367.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2275/22750367.
                                        pdf.

Prasad:2001:IBI

[2818] M. V. N. K. Prasad, K. K. Shukla, and R. N. Mukherjee. Implementation of BTTC image compression algorithm using fuzzy technique. Lecture Notes in Computer Science, 2275:375– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2275/22750375.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2275/22750375.
                                        pdf.

Soria-Frisch:2001:AIP

[2819] A. Soria-Frisch. Applications of the ILF paradigm in image processing. Lecture Notes in Computer Science, 2275:382– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2275/22750382.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2275/22750382.
pdf.

Garain:2001:ODD

[2820] U. Garain and B. B. Chaudhuri. On OCR of degraded documents using fuzzy multifactorial analysis. Lecture Notes in Computer Science, 2275:388– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2275/22750388.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2275/22750388.
pdf.

Kumar:2001:BML

[2821] R. Kumar and P. Rockett. A bootstrapped modular learning approach for scaling and generalisation of greylevel corner detection. Lecture Notes in Computer Science, 2275:395–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2275/22750395.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2275/22750395.
pdf.

Jawahar:2001:TFC

[2822] C. V. Jawahar and P. J. Narayanan.

Towards fuzzy calibration. Lecture Notes in Computer Science, 2275:401– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2275/22750401.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2275/22750401.
pdf.

Tolt:2001:FSB

[2823] G. Tolt and I. Kalaykov. Fuzzysimilarity-based image noise cancellation. Lecture Notes in Computer Science, 2275:408–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2275/22750408.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2275/22750408.pdf.

Mohan:2001:GFA

[2824] A. Mohan and K. Deb. Geneticfuzzy approach in robot motion planning revisited: Rigorous testing and towards an implementation. Lecture Notes in Computer Science, 2275:414– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2275/22750414.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2275/22750414.
pdf.

Pal:2001:EAR

[2825] T. Pal. Evolutionary approaches to rule extraction for fuzzy logic controllers. Lecture Notes in Computer Science, 2275:421–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2275/22750421.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2275/22750421.
pdf.

Pawar:2001:RGI

[2826] P. Pawar and G. Nagaraja. Regular grammatical inference: a genetic algorithm approach. Lecture Notes in Computer Science, 2275:429– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2275/22750429.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2275/22750429.
pdf.

Nanda:2001:PCS

[2827] P. K. Nanda, D. P. Muni, and P. Kanungo. Parallelized crowding scheme using a new interconnection model. Lecture Notes in Computer Science, 2275:436–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2275/22750436.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2275/22750436.
pdf.

Ghosh:2001:GFA

[2828] S. Ghosh and A. Ghosh. A GA– FUZZY approach to evolve Hopfield type optimum networks for object extraction. Lecture Notes in Computer Science, 2275:444–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2275/22750444.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2275/22750444.
pdf.

Krishnapuram:2001:SCC

[2829] R. Krishnapuram, M. Kumar, J. Basak, and V. Jain. Soft computing in E-commerce. Lecture Notes in Computer Science, 2275:450–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2275/22750450.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2275/22750450.
                                        pdf.

Loia:2001:IMB

[2830] V. Loia, P. Luongo, S. Senatore, and M. I. Sessa. Info-miner: Bridging agent technology with approximate information retrieval. Lecture Notes in Computer Science, 2275:459– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2275/22750459.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2275/22750459.
                                        pdf.

Barone:2001:FPF

[2831] J. M. Barone. Fuzzy points and fuzzy prototypes. Lecture Notes in Computer Science, 2275:466–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2275/22750466.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2275/22750466.
                                        pdf.

Srivastava:2001:SIC

[2832] A. K. Srivastava. Some injective cogenerators in fuzzy topology. Lecture Notes in Computer Science, 2275:471– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2275/22750471.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2275/22750471.
pdf.

Congxin:2001:DDC

[2833] W. Congxin, W. Guixiang, and W. Cong. Derivative and differential of convex fuzzy valued functions and application. Lecture Notes in Computer Science, 2275:478–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2275/22750478.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2275/22750478.
pdf.

Srivastava:2001:TFA

[2834] A. K. Srivastava and S. P. Tiwari. A topology for fuzzy automata. Lecture Notes in Computer Science, 2275:485– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2275/22750485.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2275/22750485.
pdf.

Maleki:2001:FNL

[2835] H. R. Maleki, H. Mishmast N., and M. Mashinchi. Fuzzy number linear programming: a probabilistic approach. Lecture Notes in Computer Science, 2275:491–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2275/22750491.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2275/22750491.
pdf.

Mi:2001:CGF

[2836] J. Mi and W. Zhang. Composition of general fuzzy approximation spaces. Lecture Notes in Computer Science, 2275:497–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2275/22750497.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2275/22750497.
pdf.

Chen:2001:LUA

[2837] D. Chen and W. Zhang. The lower and upper approximations of fuzzy sets in a fuzzy group. Lecture Notes in Computer Science, 2275:502–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2275/22750502.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2275/22750502.
pdf.

Ameri:2001:FHS

[2838] R. Ameri and M. M. Zahedi. Tfuzzy hyperalgebraic systems. Lecture Notes in Computer Science, 2275:509– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2275/22750509.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2275/22750509.
pdf.

Bhaumik:2001:SWF

[2839] R. N. Bhaumik. On some weaker forms of fuzzy compactness. Lecture Notes in Computer Science, 2275:515– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2275/22750515.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2275/22750515.
pdf.

Bhakat:2001:SRF

[2840] S. K. Bhakat. Some results on fuzzy commutative and fuzzy cyclic subgroups. Lecture Notes in Computer Science, 2275:520–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2275/22750520.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2275/22750520.
pdf.

Taheri:2001:FHT

[2841] S. M. Taheri and J. Behboodian.

Fuzzy hypotheses testing with fuzzy data: a Bayesian approach. Lecture Notes in Computer Science, 2275:527– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2275/22750527.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2275/22750527.
pdf.

Anonymous:2001:AIbw

[2842] Anonymous. Author index. Lecture Notes in Computer Science, 2275:535– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2275/2275auth.pdf.

Sag:2001:MEP

[2843] Ivan A. Sag, Timothy Baldwin, Francis Bond, Ann Copestake, and Dan Flickinger. Multiword expressions: a pain in the neck for NLP. Lecture Notes in Computer Science, 2276:1– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2276/22760001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2276/22760001.
pdf.

Hausser:2001:HOS

[2844] Roland Hausser. A hypothesis on the origin of the sign types. Lecture Notes in Computer Science, 2276:16– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

bibs/2276/22760016.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2276/22760016.
pdf.

Loukanova:2001:QIS

[2845] Roussanka Loukanova. Quantification and intensionality in situation semantics. Lecture Notes in Computer Science, 2276:32–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2276/22760032.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2276/22760032.
pdf.

Loukanova:2001:GQS

[2846] Roussanka Loukanova. Generalized quantification in situation semantics. Lecture Notes in Computer Science, 2276:46–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2276/22760046.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2276/22760046.
pdf.

Safar:2001:SLT

[2847] Éva Sáfár and Ian Marshall. Sign language translation via DRT and HPSG. Lecture Notes in Computer Science, 2276:58–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2276/22760058.htm;   http:
//link.springer-ny.com/link/service/

series/0558/papers/2276/22760058.
pdf.

Helbig:2001:MES

[2848] H. Helbig and C. Gnörlich. Multilayered extended semantic networks as a language for meaning representation in NLP systems. Lecture Notes in Computer Science, 2276:69– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2276/22760069.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2276/22760069.
pdf.

Horiguchi:2001:CSJ

[2849] Atsushi Horiguchi, Seiji Tsuchiya, Kazuhide Kojima, Hirokazu Watabe, and Tsukasa Kawaoka. Constructing a sensuous judgment system based on conceptual processing. Lecture Notes in Computer Science, 2276:86– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2276/22760086.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2276/22760086.
pdf.

Knapp:2001:TNL

[2850] Melanie Knapp and Jens Woch. Towards a natural language driven automated help desk. Lecture Notes in Computer Science, 2276:96–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

bibs/2276/22760096.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2276/22760096.
pdf.

Wilks:2001:LT

[2851] Yorick Wilks and Roberta Catizone. Lexical tuning. Lecture Notes in Computer Science, 2276:106–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2276/22760106.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2276/22760106.
pdf.

Pedersen:2001:BMW

[2852] Ted Pedersen. A baseline methodology for word sense disambiguation. Lecture Notes in Computer Science, 2276: 126–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2276/22760126.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2276/22760126.
pdf.

Banerjee:2001:ALA

[2853] Satanjeev Banerjee and Ted Pedersen. An adapted lesk algorithm for word sense disambiguation using WordNet. Lecture Notes in Computer Science, 2276:136–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2276/22760136.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2276/22760136.
                                        pdf.

Suarez:2001:FSA

[2854] Armando Suárez and Manuel Palomar. Feature selection analysis for maximum entropy-based WSD. Lecture Notes in Computer Science, 2276: 146–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2276/22760146.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2276/22760146.
                                        pdf.

Montoyo:2001:CSU

[2855] Andrés Montoyo, Armando Suárez, and Manuel Palomar. Combining supervised-unsupervised methods for word sense disambiguation. Lecture Notes in Computer Science, 2276:156– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2276/22760156.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2276/22760156.
                                        pdf.

Soler:2001:PWU

[2856] Susana Soler and Andrés Montoyo. A proposal for WSD using semantic similarity. Lecture Notes in Computer Science, 2276:165–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2276/22760165.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2276/22760165.
pdf.

Mitkov:2001:NFA

[2857] Ruslan Mitkov, Richard Evans, and Constantin Orasan. A new, fully automatic version of Mitkov’s knowledgepoor pronoun resolution method. Lecture Notes in Computer Science, 2276: 168–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2276/22760168.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2276/22760168.
pdf.

Peral:2001:PAG

[2858] Jesús Peral and Antonio Ferrández.

Pronominal anaphora generation in an English–Spanish MT approach. Lecture Notes in Computer Science, 2276: 187–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2276/22760187.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2276/22760187.
pdf.

Klebanov:2001:ULP

[2859] Beata Klebanov and Peter Wiemer-Hastings. Using LSA for pronominal anaphora resolution. Lecture Notes in Computer Science, 2276:197– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2276/22760197.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2276/22760197.
                                        pdf.

Meza:2001:SAV

[2860] Ivan Meza and Luis Pineda. The Spanish auxiliary verb system in HPSG. Lecture Notes in Computer Science, 2276:200–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2276/22760200.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2276/22760200.
                                        pdf.

Bolshakov:2001:SSR

[2861] Igor A. Bolshakov. Surface syntactic relations in Spanish. Lecture Notes in Computer Science, 2276:210– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2276/22760210.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2276/22760210.
                                        pdf.

Blache:2001:PIF

[2862] Philippe Blache and David-Olivier Azulay. Parsing ill-formed inputs with constraint graphs. Lecture Notes in Computer Science, 2276:220–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2276/22760220.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2276/22760220.
                                        pdf.

Araujo:2001:PST

[2863] Lourdes Araujo. Part-of-speech tagging with evolutionary algorithms. Lecture Notes in Computer Science, 2276:230– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2276/22760230.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2276/22760230.
pdf.

Grana:2001:FMT

[2864] Jorge Graña, Fco. Mario Barcala, and Jesús Vilares. Formal methods of tokenization for partof-speech tagging. Lecture Notes in Computer Science, 2276:240–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2276/22760240.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2276/22760240.
pdf.

Jimenez:2001:SPT

[2865] Héctor Jiménez and Guillermo Morales.

Sepe: a POS tagger for Spanish. Lecture Notes in Computer Science, 2276: 250–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2276/22760250.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2276/22760250.
pdf.

Kogut:2001:FST

[2866] Dariusz J. Kogut. Fuzzy set tagging. Lecture Notes in Computer Science, 2276:260–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2276/22760260.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2276/22760260.
pdf.

Calzolari:2001:TSM

[2867] Nicoletta Calzolari, Antonio Zampolli, and Alessandro Lenci. Towards a standard for a multilingual lexical entry: The EAGLES/ISLE initiative. Lecture Notes in Computer Science, 2276:264– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2276/22760264.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2276/22760264.
pdf.

Bolshakov:2001:QCH

[2868] Igor A. Bolshakov, Sofia N. Galicia-Haro, and Alexander Gelbukh. Quantitative comparison of homonymy in Spanish EuroWordNet and traditional dictionaries. Lecture Notes in Computer Science, 2276:280–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2276/22760280.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2276/22760280.
pdf.

Gelbukh:2001:CSR

[2869] Alexander Gelbukh, Grigori Sidorov, and Liliana Chanona-Hernández. Compilation of a Spanish representative corpus. Lecture Notes in Computer Science, 2276:285–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2276/22760285.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2276/22760285.pdf.

Casillas:2001:AMT

[2870] Arantza Casillas and Raquel Martínez.

Aligning multiword terms using a hybrid approach. Lecture Notes in Computer Science, 2276:289–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2276/22760289.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2276/22760289.
pdf.

Gelbukh:2001:ASD

[2871] Alexander Gelbukh and Grigori Sidorov. Automatic selection of defining vocabulary in an explanatory dictionary. Lecture Notes in Computer Science, 2276:300–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2276/22760300.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2276/22760300.
pdf.

Harbusch:2001:INL

[2872] Karin Harbusch and Jens Woch. Integrated natural language generation with schema-tree adjoining grammars. Lecture Notes in Computer Science, 2276:304–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2276/22760304.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2276/22760304.
pdf.

Casillas:2001:EBD

[2873] Arantza Casillas and Raquel Martínez. Experiments with a bilingual document generation environment. Lecture Notes in Computer Science, 2276:314– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2276/22760314.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2276/22760314.
pdf.

Shirado:2001:CMC

[2874] Tamotsu Shirado and Hitoshi Isahara. A computational model of change in politeness with the addition of word endings. Lecture Notes in Computer Science, 2276:318–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2276/22760318.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2276/22760318.
pdf.

Suleymanov:2001:TMI

[2875] Dj.Sh. Suleymanov. Tartar morphology implementation. Lecture Notes in Computer Science, 2276:327–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2276/22760327.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2276/22760327.
pdf.

Uraga:2001:AGP

[2876] Esmeralda Uraga and Luis Pineda.

Automatic generation of pronunciation lexicons for Spanish. Lecture Notes in Computer Science, 2276:330– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2276/22760330.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2276/22760330.
pdf.

Mihalcea:2001:DRL

[2877] Rada F. Mihalcea. Diacritics restoration: Learning from letters versus learning from words. Lecture Notes in Computer Science, 2276:339–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2276/22760339.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2276/22760339.
pdf.

Feldman:2001:CSI

[2878] Ronen Feldman, Yonatan Aumann, Michal Finkelstein-Landau, Eyal Hurvitz, Yizhar Regev, and Ariel Yaroshevich. A comparative study of information extraction strategies. Lecture Notes in Computer Science, 2276:349– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2276/22760349.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2276/22760349.
pdf.

Rinaldi:2001:AET

[2879] Fabio Rinaldi, Michael Hess, Diego Mollá, Rolf Schwitter, James Dowdall, Gerold Schneider, and Rachel Fournier. Answer extraction in technical domains. Lecture Notes in Computer Science, 2276:360–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2276/22760360.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2276/22760360.
pdf.

Rodriguez:2001:AEN

[2880] Carlos Rodríguez. Automatic extraction of non-standard lexical data for a metalinguistic information database. Lecture Notes in Computer Science, 2276:370–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2276/22760370.htm;   http:
//link.springer-ny.com/link/service/

series/0558/papers/2276/22760370.
pdf.

Llopis:2001:TSE

[2881] Fernando Llopis, Antonio Ferrández, and José Luis Vicedo. Text segmentation for efficient information retrieval. Lecture Notes in Computer Science, 2276:373–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2276/22760373.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2276/22760373.
pdf.

Vilares:2001:USD

[2882] Jesús Vilares, Fco. Mario Barcala, and Miguel A. Alonso. Using syntactic dependency-pairs conflation to improve retrieval performance in Spanish. Lecture Notes in Computer Science, 2276: 381–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2276/22760381.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2276/22760381.
pdf.

Kuo:2001:MDS

[2883] June-Jei Kuo, Hung-Chia Wung, Chuan-Jie Lin, and Hsin-Hsi Chen. Multi-document summarization using informative words and its evaluation with a QA system. Lecture Notes in Computer Science, 2276:391– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2276/22760391.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2276/22760391.
                                        pdf.

Zizka:2001:ASI

[2884] Jan Žižka and Aleš Bourek. Automated selection of interesting medical text documents by the TEA text analyzer. Lecture Notes in Computer Science, 2276:402–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2276/22760402.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2276/22760402.
                                        pdf.

Zhou:2001:CDC

[2885] Shuigeng Zhou and Jihong Guan. Chinese documents classification based on N-grams. Lecture Notes in Computer Science, 2276:405–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2276/22760405.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2276/22760405.pdf.

Steinberger:2001:CLD

[2886] Ralf Steinberger, Bruno Pouliquen, and Johan Hagman. Cross-lingual document similarity calculation using the multilingual thesaurus EUROVOC. Lecture Notes in Computer Science, 2276:415–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2276/22760415.htm;   http:

//link.springer-ny.com/link/service/
series/0558/papers/2276/22760415.
pdf.

Makagonov:2001:EFT

[2887] Pavel Makagonov and Mikhail Alexandrov. Empirical formula for testing word similarity and its application for constructing a word frequency list. Lecture Notes in Computer Science, 2276: 425–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2276/22760425.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2276/22760425.
pdf.

Akhtar:2001:ATA

[2888] Shazia Akhtar, Ronan G. Reilly, and John Dunnion. AutoMarkup: a tool for automatically marking up text documents. Lecture Notes in Computer Science, 2276:433–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2276/22760433.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2276/22760433.
pdf.

Sierra:2001:IRP

[2889] Gerardo Sierra and Rodrigo Alarcón.

Identification of recurrent patterns to extract definitory contexts. Lecture Notes in Computer Science, 2276:436– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2276/22760436.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2276/22760436.
                                        pdf.

Vazquez:2001:SMM

[2890] Sonia Vázquez, Maria Carmen Calle, Susana Soler, and Andrés Montoyo. Specification marks method: Design and implementation. Lecture Notes in Computer Science, 2276:439–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2276/22760439.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2276/22760439.
                                        pdf.

Anonymous:2001:AIbx

[2891] Anonymous. Author index. Lecture Notes in Computer Science, 2276:443– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2276/2276auth.pdf.

Arieli:2001:MPR

[2892] Ofer Arieli and Marc Denecker. Modeling paraconsistent reasoning by classical logic. Lecture Notes in Computer Science, 2284:1–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2284/22840001.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2284/22840001.
                                        pdf.

Banerjee:2001:ITA

[2893] R. N. Banerjee and A. Bujosa.

Implementing term algebra syntactic unification in free modules over certain rings. Lecture Notes in Computer Science, 2284:15–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2284/22840015.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2284/22840015.
pdf.

Bertossi:2001:ATD

[2894] Leopoldo Bertossi and Camilla Schwind.

Analytic tableaux and database repairs: Foundations. Lecture Notes in Computer Science, 2284:32–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2284/22840032.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2284/22840032.
pdf.

Biskup:2001:CQE

[2895] Joachim Biskup and Piero Bonatti. Controlled query evaluation for known policies by combining lying and refusal. Lecture Notes in Computer Science, 2284:49–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2284/22840049.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2284/22840049.
pdf.

Amo:2001:LFI

[2896] Sandra de Amo, Walter A. Carnielli, and Jo ao Marcos. A logical framework for integrating inconsistent information in multiple databases. Lecture Notes in Computer Science, 2284:67– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2284/22840067.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2284/22840067.
                                        pdf.

Demetrovics:2001:FDP

[2897] János Demetrovics, Gyula O. H. Katona, and Dezső Miklós. Functional dependencies in presence of errors. Lecture Notes in Computer Science, 2284: 85–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2284/22840085.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2284/22840085.
                                        pdf.

Huynh:2001:CMC

[2898] V. N. Huynh, Y. Nakamori, T. B. Ho, and G. Resconi. A context model for constructing membership functions of fuzzy concepts based on modal logic. Lecture Notes in Computer Science, 2284:93–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2284/22840093.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2284/22840093.
pdf.

Kern-Isberner:2001:PCP

[2899] Gabriele Kern-Isberner. The principle of conditional preservation in belief revision. Lecture Notes in Computer Science, 2284:105–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2284/22840105.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2284/22840105.
pdf.

Koch:2001:QRS

[2900] Christoph Koch. Query rewriting with symmetric constraints. Lecture Notes in Computer Science, 2284:130– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2284/22840130.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2284/22840130.
pdf.

Li:2001:PEM

[2901] Yingjiu Li, Sencun Zhu, Lingyu Wang, and Sushil Jajodia. A privacyenhanced microaggregation method. Lecture Notes in Computer Science, 2284:148–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2284/22840148.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2284/22840148.
pdf.

Link:2001:TTT

[2902] Sebastian Link. Towards a tailored theory of consistency enforcement in databases. Lecture Notes in Computer Science, 2284:160–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2284/22840160.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2284/22840160.
                                        pdf.

Maimon:2001:ISL

[2903] Oded Maimon and Lior Rokach. Improving supervised learning by feature decomposition. Lecture Notes in Computer Science, 2284:178–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2284/22840178.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2284/22840178.
                                        pdf.

Tichler:2001:ETD

[2904] Krisztián Tichler. Extremal theorems for databases. Lecture Notes in Computer Science, 2284:197–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2284/22840197.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2284/22840197.
                                        pdf.

Torres:2001:RDH

[2905] José María Turull Torres. Relational databases and homogeneity in logics with counting. Lecture Notes in Computer Science, 2284:212–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2284/22840212.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2284/22840212.
pdf.

Hegner:2001:UUS

[2906] Stephen J. Hegner. Uniqueness of update strategies for database views. Lecture Notes in Computer Science, 2284: 230–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2284/22840230.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2284/22840230.
pdf.

Cirulis:2001:KRE

[2907] Jānis Cīrulis. Knowledge representation in extended Pawlak’s information systems: Algebraic aspects. Lecture Notes in Computer Science, 2284:250– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2284/22840250.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2284/22840250.
pdf.

Engel:2001:MSD

[2908] Konrad Engel and Sven Hartmann. Minimal sample databases for global cardinality constraints. Lecture Notes in Computer Science, 2284:268–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2284/22840268.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2284/22840268.
                                        pdf.

Anonymous:2001:AIby

[2909] Anonymous. Author index. Lecture Notes in Computer Science, 2284:289– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2284/2284auth.pdf.

Berstel:2001:BGT

[2910] Jean Berstel and Luc Boasson. Balanced grammars and their languages. Lecture Notes in Computer Science, 2300:3–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2300/23000003.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2300/23000003.
                                        pdf.

Diekert:2001:SLP

[2911] Volker Diekert and Paul Gastin. Safety and liveness properties for real traces and a direct translation from LTL to monoids. Lecture Notes in Computer Science, 2300:26– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2300/23000026.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2300/23000026.
pdf.

Engelfriet:2001:DOS

[2912] Joost Engelfriet. The delta operation:

From strings to trees to strings. Lecture Notes in Computer Science, 2300: 39–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2300/23000039.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2300/23000039.
pdf.

Halava:2001:ISM

[2913] Vesa Halava and Tero Harju. Infinite solutions of marked post correspondence problem. Lecture Notes in Computer Science, 2300:57–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2300/23000057.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2300/23000057.
pdf.

Karhumaki:2001:BPA

[2914] Juhani Karhumäki and Ion Petre.

The branching point approach to Conway’s problem. Lecture Notes in Computer Science, 2300:69–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2300/23000069.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2300/23000069.
                                        pdf.

Luca:2001:SSQ

[2915] Aldo de Luca. A survey of some quantitative approaches to the notion of information. Lecture Notes in Computer Science, 2300:77–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2300/23000077.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2300/23000077.
                                        pdf.

Mateescu:2001:NT

[2916] Alexandru Mateescu and Arto Salomaa. Nondeterministic trajectories. Lecture Notes in Computer Science, 2300:96–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2300/23000096.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2300/23000096.
                                        pdf.

Restivo:2001:BPI

[2917] Antonio Restivo and Sergio Salemi. Binary patterns in infinite binary words. Lecture Notes in Computer Science, 2300:107–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2300/23000107.htm;   http:
                                        //link.springer-ny.com/link/service/

series/0558/papers/2300/23000107.
pdf.

Kreowski:2001:SST

[2918] Hans-Jörg Kreowski. A sight-seeing tour of the computational landscape of graph transformation. Lecture Notes in Computer Science, 2300:119– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2300/23000119.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2300/23000119.
pdf.

Janssens:2001:LAS

[2919] Dirk Janssens. Local action systems and DPO graph transformation. Lecture Notes in Computer Science, 2300: 138–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2300/23000138.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2300/23000138.
pdf.

Baldan:2001:BEG

[2920] Paolo Baldan, Andrea Corradini, and Ugo Montanari. Bisimulation equivalences for graph grammars. Lecture Notes in Computer Science, 2300:158– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2300/23000158.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2300/23000158.
pdf.

Ehrig:2001:HLN

[2921] Hartmut Ehrig, Kathrin Hoffmann, Julia Padberg, Paolo Baldan, and Reiko Heckel. High-level net processes. Lecture Notes in Computer Science, 2300: 191–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2300/23000191.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2300/23000191.
                                        pdf.

Beek:2001:PNC

[2922] Maurice ter Beek and Jetty Kleijn. Petri net control for grammar systems. Lecture Notes in Computer Science, 2300:220–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2300/23000220.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2300/23000220.
                                        pdf.

Thiagarajan:2001:RES

[2923] P. S. Thiagarajan. Regular event structures and finite Petri nets: a conjecture. Lecture Notes in Computer Science, 2300:244–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2300/23000244.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2300/23000244.pdf.

Engels:2001:TTA

[2924] Gregor Engels and Luuk Groenewegen. Towards team-automata-driven objectoriented collaborative work. Lecture Notes in Computer Science, 2300:257– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2300/23000257.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2300/23000257.
pdf.

Esparza:2001:GP

[2925] Javier Esparza. Grammars as processes. Lecture Notes in Computer Science, 2300:277–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2300/23000277.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2300/23000277.pdf.

Nielsen:2001:TCC

[2926] Mogens Nielsen and Frank D. Valencia.

Temporal concurrent constraint programming: Applications and behavior. Lecture Notes in Computer Science, 2300:298–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2300/23000298.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2300/23000298.
pdf.

Bottoni:2001:RPS

[2927] Paolo Bottoni, Anna Labella, Carlos Martín-Vide, and Gheorghe Paun. Rewriting P systems with conditional communication. Lecture Notes in Computer Science, 2300:325–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2300/23000325.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2300/23000325.
                                        pdf.

Head:2001:AAF

[2928] Tom Head. An aqueous algorithm for finding the bijections contained in a binary relation. Lecture Notes in Computer Science, 2300:354–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2300/23000354.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2300/23000354.
                                        pdf.

Hoogeboom:2001:UBR

[2929] Hendrik Jan Hoogeboom and Nikè van Vugt. Upper bounds for restricted splicing. Lecture Notes in Computer Science, 2300:361–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2300/23000361.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2300/23000361.
                                        pdf.

Kari:2001:CID

[2930] Lila Kari, Rob Kitto, and Gabriel Thierrin. Codes, involutions, and DNA encodings. Lecture Notes in Computer Science, 2300:376–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2300/23000376.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2300/23000376.
pdf.

Prescott:2001:DMC

[2931] David M. Prescott. DNA manipulations in ciliates. Lecture Notes in Computer Science, 2300:394–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2300/23000394.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2300/23000394.
pdf.

Yokomori:2001:MPS

[2932] Takashi Yokomori, Yasubumi Sakakibara, and Satoshi Kobayashi. A magic pot: Self-assembly computation revisited. Lecture Notes in Computer Science, 2300:418–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2300/23000418.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2300/23000418.
pdf.

Anonymous:2001:AIbz

[2933] Anonymous. Author index. Lecture Notes in Computer Science, 2300:431– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        papers/2300/2300indx.pdf.